A Complete Bibliography of Publications in Fisheries

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 August 2024 Version 1.20

Title word cross-reference

#25DaysofFishmas [3054]. #6 [1446]. #Welcome [2324].

15 [869]. ^{nd} [1190]. ^{st} [1112]. ^{th} [1190, 2446]. δ [869].

-i [101].

//fisheries.org/meetings [1619].

100-Watt [2821]. 101 [1930, 1649]. 12th [2117]. 13th [2629, 2646]. 140th [907]. 142nd [1268]. 143rd [1449]. 144th [1576, 1658]. 145th [1873, 1603, 1619, 1629, 1656]. 146th [2089, 1832, 1845, 1860, 1870, 1921, 1996]. 147th [2037, 2051, 2072, 2086, 2109, 2280]. 148 [2505]. 148th [2239, 2255, 2376, 2440]. 149th [2456]. 150 [2575, 2602, 2588, 2670]. 150th [2673, 2777, 2347, 2721]. 151st [2754, 2770, 2882, 2913]. 151th [2870]. 153rd [3239, 3165]. 17th [2119]. 1870 [2435]. 1870s [635]. 1872 [928]. 1880s [635]. 19 [2693, 2809, 2909]. 19-=23 [1268]. 1926 [182]. 1936 [2680]. 1980’s [43]. 1985 [195]. 1990s [298].

2003 [1665]. 2008 [1660]. 2010 [905, 962, 1176]. 2011 [963, 961, 1009, 1064, 1119, 1037, 1053, 996, 1022, 1106, 1092, 1079, 1102]. 2012 [1072, 1101, 1180, 1128, 1182, 1235, 1148, 1291, 1153, 1260, 1136, 1208, 1223, 1167, 1194, 1192, 1209, 1232, 1246, 1289, 1277, 1262, 1248, 1179, 1221, 1274, 1268, 1303, 1450]. 2012/Journal [1192]. 2013 [1242, 1256, 1269, 1351, 1317, 1338, 1394, 1458, 1305, 1326, 1355, 1378, 1408, 1441, 1315, 1363, 1423, 1588, 1571, 1472, 1776]. 2014 [1558, 1506, 1524, 1579, 1592, 1622, 1661, 1487, 1544, 1607, 1650, 1670, 1686, 1600, 1744]. 2015 [1662, 1815, 1704, 1739, 1769, 1831, 1861, 1878, 1875, 1723, 1755, 1784, 1846, 1911, 1912, 1895, 1823, 1873, 1781]. 2015.fisheries.org [1629, 1656]. 2016 [1959, 1972, 2043, 1958, 2027, 2059, 2094, 1943, 1997, 2073, 2111, 2089, 2024]. 2017 [2037, 2062, 2091, 2169, 2200, 2126, 2109, 2172, 2238, 2142, 2157, 2202, 2223, 2268, 2286, 2125, 2187, 2254, 2304, 2265]. 2018 [2239, 2255, 2290, 2307, 2321, 2335, 2346, 2358, 2371, 2383, 2393, 2407, 2420, 2433, 2354, 2404, 2332, 2343, 2380, 2390, 2318, 2368]. 2019 [2461, 2500, 2564, 2580, 2593, 2607]. 2020 [2636, 2682, 2731, 2746, 2760, 2775, 2781, 2785, 2798, 2661, 2799, 2786, 2755, 2739]. 2021 [2754, 2770, 2870, 2803, 2816, 2840, 2888, 2904, 2918, 2929, 2952, 2858, 2963, 2881, 2935, 2911, 2936, 2831, 3199]. 2022 [3091, 2957, 2967, 2984, 2999, 3013, 3024, 3037, 3049, 3060, 3070, 3083, 2980, 3020, 3117, 3102, 3113, 2986, 3088, 3129, 3007]. 2023 [3210, 3095, 3107, 3121, 3135, 3146, 3159, 3169, 3180, 3194, 3207, 3218, 3233, 3129, 3227, 3268]. 2024 [3246, 3259, 3274, 3253]. 20th [1204]. 21st [1281, 2041, 1471, 584, 380, 3261]. 21st-Century [1281]. 22nd [1205].

3 [2446]. 33431 [168].

4.0 [2264]. 42nd [2423]. 48105 [141].

5th [3115, 1824].

63rd [1778]. 6th [1126].

712 [175]. 75th [1202].

8 [612, 1683].

98195 [216].

= [2162]. =23 [1268].

A. [1142, 1577, 2911]. Abbreviating [1908]. Abdicating [884]. Abnormal [689]. Abound [2309]. Absolutely [1852]. Abstracts [1534, 1576, 1996, 2119]. Abundance [2165, 1157, 2481, 1097, 947, 3101, 2452]. Academic [2877, 2879, 2018, 1635, 1049, 1213]. Academic-Government [1213]. Academy [61]. Accelerated [1565]. Accept [3017, 2794]. Accepting [1299]. Access [172, 1601, 492]. Accomodations [994]. Accomplished [2384]. Accomplishments [3163]. Account [1496]. Accountability [334, 759]. Accounting [768]. Accreditation [389]. accuracy [649]. Achieving [1098]. Acid [124, 191]. Acidification [296, 121]. Acidified [120, 233, 123, 126]. ACL [1095]. ACL/AM [1095]. Acoustic [1613, 2923, 901, 2922]. Acronyms [3271]. Across [2339, 2710, 1214, 2395, 2121]. Acrylic [2955]. Act [16, 1658, 1307, 2856, 2864, 602, 659, 756, 821, 1431, 2867, 755, 2427, 1646, 1766, 2203, 3029, 738, 2409, 2697, 3002, 67]. Act-listed [821]. Acting [1427, 584]. Action [2891, 1368, 626, 2565, 2890, 421, 3224, 2184, 2209, 2852, 2210, 2930, 1843, 1966]. Actions [2048, 2550, 1904, 875, 3201]. Activation [2910]. Active [860, 3248]. Activities [846, 3213]. Activity [2534]. Actually [2107]. Adams [1540]. Adaptation [2895, 2413, 1779]. Adapting [1909, 1983]. Adaptions [2166]. Adaptive [1909, 1296, 3264, 525, 1970, 902, 2299]. Adapts [2672]. Add [2104]. Added [1790]. Adding [2146]. Addition [1528, 123]. Address [2136, 2228]. Addresses [1439]. Addressing [879, 1110, 1070, 999, 3187, 1970, 1239]. Adequate [35]. Adirondack [125]. Administration [1953, 1949]. Administrators [2508, 809]. Adopt [643]. Adopt-A-Trout [643]. Adopting [1172]. Adoption [2610]. Adopts [1300]. Adriatic [1404]. Adult [851, 574, 1227]. Advance [1614]. Advanced [609, 2531, 2518, 1429, 1240, 2260]. Advances [416, 2906]. Advancing [1076, 1024, 2337]. Adventure [3073, 1799]. Adventures [1797, 1822]. Advice [2016, 1810]. Advisor [1160]. Advisories [2744, 290]. Advisory [364, 2300]. Advocacy [2338, 2190, 2197, 2194, 2195, 2323]. Advocate [1529, 2191, 2776]. Advocating [2196]. Aerial [724, 3078, 2962, 761]. Affect [1761, 2892]. Affecting [2313]. Affiliates [1039]. Africa [1555]. African [441]. AFS [905, 963, 1201, 1921, 2062, 2126, 2882, 2913, 1932, 1720, 1777, 3129, 1744, 2376, 632, 2411, 3210, 908, 1112, 1060, 1017, 979, 995, 1033, 1071, 1114, 1100, 980, 1180, 1220, 1127, 1173, 1148, 1151, 1178, 1233, 1164, 1311, 1324, 1336, 1351, 1327, 1354, 1365, 1377, 1424, 1440, 1466, 1481, 1498, 1521, 1668, 1662, 1489, 1508, 1525, 1545, 1581, 1792, 1690, 1691, 1812, 1903, 2091, 2575, 2602, 2478, 2588, 2949, 3033, 2951, 2945, 3142, 3117, 3154, 3228, 3214, 3254, 1435, 1511, 1791, 1918, 2435, 2777, 2628, 3272, 1749, 1728, 1745]. AFS [2101, 2130, 2115, 1871, 1901, 1933, 2696, 1264, 989, 1271, 2468, 2308, 2324, 787, 2975, 2038, 2687, 964, 2469, 2148, 2976, 3283, 1866, 1948, 1930, 1977, 1081, 1094, 1225, 1122, 1941, 2824, 2599, 1020, 2532, 2986, 2464, 1039, 1444, 1889, 2644, 3196, 1349, 2871, 1149, 1335, 3288, 2669, 1245, 2114, 2205, 3282, 2004, 2322, 2372, 2291, 2359, 2272, 2258, 2226, 3008, 1216, 1393, 2427, 2672, 2153, 2314, 2630, 2642, 2749, 2713, 3130, 1967, 1141, 279, 1611, 1596, 1708, 2825, 1074, 2194, 2505, 1102, 1161, 2315, 3125, 3198, 1746, 1711, 1710, 977, 2725, 2426, 2638, 3131, 2645, 1399, 2610]. AFS [2831, 1239, 1131, 1132, 1457, 1362, 1372, 1640, 3257, 2670, 1902, 1934, 2154, 2977, 1788, 2260, 2409, 2747, 2899, 3153, 1143, 1314, 2726, 3281]. AFS/Sea [1314]. After [1928, 913, 1939, 1089, 2764, 818]. Again [3085, 2743]. Against [2891, 2890, 2701, 2897]. Age [707, 1390, 3279, 815, 1254, 1736, 3043, 2597, 2034, 649, 749]. Agencie [400]. Agencies [682, 894, 1430, 937, 310, 139, 816, 410, 2359, 929, 252, 448]. Agency [192, 881, 2387, 403, 593, 2807, 809, 1552, 3212, 426]. Agent [33]. Ages [1611]. Aggregations [613]. Aging [816, 3238]. Agreements [250]. Agricultural [2508, 175, 104]. agriculture [605]. Ahead [1311, 1324, 1336, 1392, 1237]. Ahhh [1864]. Aid [2386, 2688]. Aiming [1715]. Aims [1738]. Air [1827]. Airborne [2250]. Aircraft [2424]. Airliner [1275]. al [1500]. Alabama [2022, 1244, 2036, 1731]. Alan [1015]. Alaska [962, 444, 1538, 637, 2569, 513, 2455, 3175, 2149, 2192, 2352, 152, 2299, 2551, 2553, 2991]. Alaska-Now [152]. Alaskan [873]. Albatross [2658]. Alberta [1780]. albus [2650, 107]. Alert [1323]. Alewife [3069, 1266]. Alexander [2749]. Algae [2080]. Algorithm [758]. Alien [894, 2082, 3150, 1186]. Alike [554, 1820]. Alleles [2107]. Allen [1542, 1104, 1321]. Allied [1244]. Allocations [449]. Allowing [2531]. Alluvial [2056]. Ally [2823]. Already [3052]. also [2084]. Alter [1886, 1820]. Altered [832]. Altering [1771]. Alternate [1982, 3271, 340]. Alternative [802, 1432, 69]. Alternatives [416]. Alters [1819]. Alumni [2868]. Always [2196, 3026]. AM [1095]. Amazing [2721]. Amazon [3225, 694, 2164, 2833]. Ambassador [2899]. Amber [2080]. Amendment [1017, 1095]. America [3141, 2864, 3002, 1631, 969, 712, 946, 701, 2483, 419, 337, 1212, 660, 565, 732, 581, 2809, 601, 1971, 230, 276, 448, 242]. American [822, 899, 961, 1059, 952, 1003, 1021, 1032, 1090, 1116, 970, 997, 1048, 1103, 1128, 1150, 1165, 1192, 1209, 1232, 1246, 1289, 1134, 1179, 1221, 1274, 1268, 1303, 1326, 1355, 1378, 1408, 1441, 1315, 1363, 1423, 1603, 1619, 1629, 1656, 1472, 1506, 1524, 1579, 1592, 1622, 1487, 1544, 1607, 1650, 1670, 1832, 1845, 1860, 1870, 1686, 1704, 1739, 1769, 1831, 1861, 1878, 1723, 1755, 1784, 1846, 2072, 2086, 1926, 1912, 1958, 2027, 2059, 2094, 1895, 1943, 1997, 2073, 2109, 2142, 2157, 2202, 2223, 2268]. American [2286, 2125, 2187, 2254, 2332, 2343, 2380, 2390, 2318, 2368, 2456, 2952, 3091, 3116, 3128, 2574, 3129, 2504, 3019, 907, 976, 1129, 1283, 1284, 1190, 1380, 1454, 1453, 1664, 1663, 1813, 1874, 1876, 1825, 1802, 2037, 2051, 2070, 2090, 2088, 2169, 2200, 2239, 2282, 2283, 2255, 2442, 2446, 2443, 2511, 2613, 2615, 2616, 2479, 2673, 2754, 2783, 2770, 3059, 3239, 3242, 516, 2721, 1565, 1390, 366, 1313, 3152, 2363, 1826, 576, 1158, 1157, 895, 2313, 1923, 854, 203, 779, 2006, 2481, 866, 573, 1981, 2122, 749, 2209, 2547, 243, 546]. American [589, 1979, 2895, 2210, 1614, 2195, 1828, 3221, 1300, 2724, 1980, 218, 3153]. americana [146]. Amezcua [957]. Amia [2802, 2231]. Among [3263, 1762, 2807, 73, 235, 200, 278, 2660, 2284, 42, 2300, 640, 2991]. Ampat [2232]. Amphibians [2183]. Amplified [3063, 2959, 2987, 2944, 3086, 3097, 3124, 3137, 3172, 3197, 3249]. Amputee [1907]. Amur [33]. Anadromous [287, 730, 423, 2234, 154, 1990, 1820, 445, 1920, 2506]. Analog [800]. Analogy [270]. Analysis [1125, 2922, 777, 1752, 2120, 291, 383, 1872, 826, 1683, 359, 29, 2312, 654, 3201, 2053, 315, 1214]. Analytical [140]. Ancestry [2413]. Anchovy [2752]. Ancient [2659]. Anderson [2496]. Anemia [1631, 1632]. Anesthetic [989, 765]. Anesthetic/Sedative [989]. Anesthetics [159, 1820]. Angle [1700, 956, 1532]. Angler [631, 3004, 281, 623, 2425, 361, 2701, 3074, 2744, 2532, 114, 561, 562, 436, 2924, 1254, 1797, 2584, 3110, 1780, 188, 700, 533, 1566]. Angler-assisted [562]. Angler-based [561]. Anglers [554, 1389, 2425, 235, 2657, 752, 1554, 593, 2737, 2465, 2300, 2454, 3262, 667]. Angling [330, 1171, 802, 2671, 2748, 1809, 1767, 58, 3126, 2934, 2534]. Animal [951, 1077, 1260, 1150, 1192, 1338, 1394, 1458, 1579, 1815, 1926, 1972, 2043, 2111, 2172, 2238, 2304, 2354, 2404, 157]. Ann [141]. Anne [1587]. Anniversary [1202, 2347, 2721, 1204]. Announced [2661]. Announcements [911]. Annual [3210, 907, 1112, 1060, 1120, 962, 1180, 1220, 1190, 1268, 1603, 1619, 1629, 1656, 1588, 1576, 1813, 1877, 1832, 1845, 1860, 1870, 1921, 2037, 2051, 2092, 2072, 2086, 1996, 2169, 2239, 2287, 2109, 2255, 2446, 2511, 2456, 2673, 2754, 2770, 2680, 2882, 2913, 2870, 2949, 3033, 2993, 3090, 3089, 3142, 3239, 3165, 3254, 1476, 1918, 1523, 1658, 1873, 2089, 2280, 3098, 2308, 1617, 1778, 2687, 1217, 2532, 3123, 2423, 2871, 1205, 3008, 1967, 2024, 1161, 1449, 2376, 2831, 2504, 1132, 2440, 1230, 3235, 2200, 2155, 2341, 2523, 2673, 2781, 1591]. Annual [1795]. Answer [1994]. Answered [1448]. Answers [1647]. Anthropogenic [1361, 679, 3149]. Anti [1171]. Anti-Angling [1171]. Anticipated [1993, 267, 266]. Anticipating [245]. Anticipation [1191]. Antimycin [634]. Anyone [913]. Apex [1746]. Apostle [2749]. App [2085]. Appalachian [320, 1389, 740]. Applicants [2126]. Application [1129, 1380, 224, 50, 3226, 2264, 761, 114, 454, 815, 163, 1780, 27, 2424, 608]. Applications [1662, 2055, 2250, 491, 2924, 219, 2767, 644, 1938]. Applied [124]. Applying [987, 3127]. Appointment [1129, 1380]. Appreciate [1563]. Appreciation [1635, 2118]. Approach [802, 522, 2130, 2218, 3064, 1681, 1110, 2585, 756, 596, 1070, 336, 227, 386, 355, 3127, 538, 762, 1722, 1920, 1788, 2844, 1702]. Approaches [939, 2243, 2005, 1469, 1470, 2351, 1418, 2907, 935, 929, 3042, 1995, 262, 706]. Appropriate [568, 1500, 1501]. Approved [1589]. Approves [1095]. April [3113, 2786, 3199, 2755, 3007, 1009, 1182, 1209, 1232, 1355, 1378, 1579, 1769, 2027, 2202, 2335, 2380, 2489, 2667, 2830, 2984, 3135, 2469]. Aqua [1589]. Aquacultural [2702]. Aquaculture [1643, 2624, 1627, 1645, 1369, 1681, 1649, 1638, 1070, 86, 2011, 3151, 255, 31, 1642, 1646, 137, 1633, 1634, 1711, 1775, 1760, 1820, 167, 2632, 1239, 1362, 1628, 2735, 2747, 1649, 2871, 1192, 952, 1003, 1032, 1090, 1150, 1232, 1289, 1326, 1378, 1441, 1506, 1579, 1704, 1769, 1861, 1926, 2027, 2094, 2142, 2202, 2268, 2332, 2380]. Aquaculture-Themed [1628]. Aquacultures [1244]. Aquaponics [1636, 1635]. Aquaria [2337]. Aquarium [973]. Aquatic [801, 718, 951, 1077, 1260, 1150, 1192, 1338, 1394, 1458, 1467, 1579, 1815, 1926, 1972, 2043, 2111, 2172, 2238, 2304, 2354, 2404, 2891, 2980, 3031, 1678, 2723, 602, 492, 1512, 1978, 1931, 306, 2890, 805, 540, 169, 2134, 170, 33, 2328, 2853, 1986, 724, 2116, 1635, 1070, 120, 2166, 826, 1892, 1762, 2455, 3224, 2400, 479, 741, 3100, 436, 104, 2184, 378, 1904, 893, 355, 2399, 2118, 42, 1949, 2099, 1995, 1428, 2398, 917, 167, 2397, 166, 1083, 2892, 1843, 2768, 2276, 1132, 1325, 1297, 165, 1844, 2292, 761]. Arapaima [1822]. Arbor [141]. Archives [1034, 182, 490]. Arctic [3179, 1886, 3188, 160, 2181, 3175, 2149, 1804]. Arden [2524]. Area [2960]. Areal [723]. Areas [688, 1027, 2531, 863, 421, 993, 2209, 2210, 700, 1026]. Arena [2337]. Aren’t [3251]. Arising [2300]. Arizona [205, 2230, 2229]. Armor [2698]. Armored [2698]. Arowana [3134]. Array [907]. arrogance [502]. Arse [3287]. Art [2131, 2641, 2361, 2163, 1714, 2158, 2284]. Article [1, 2, 4, 6, 9, 10, 11, 12, 13, 15, 19, 21, 22, 24, 26, 30, 32, 34, 36, 41, 48, 52, 53, 54, 56, 59, 62, 64, 70, 72, 74, 76, 81, 83, 84, 87, 91, 93, 95, 96, 99, 102, 106, 108, 110, 112, 113, 116, 119, 127, 130, 134, 138, 142, 145, 148, 151, 153, 156, 158, 161, 164, 171, 174, 178, 181, 184, 186, 190, 193, 197, 201, 204, 207, 210, 214, 217, 221, 225, 228, 231, 234, 237, 241]. Article [244, 248, 251, 254, 258, 263, 271, 272, 275, 280, 283, 289, 292, 297, 302, 309, 312, 316, 319, 322, 325, 328, 332, 335, 338, 341, 344, 346, 348, 350, 352, 354, 357, 360, 362, 365, 368, 371, 374, 376, 379, 382, 385, 388, 392, 395, 397, 399, 402, 405, 407, 408, 412, 415, 418, 420, 422, 425, 428, 431, 434, 437, 440, 443, 446, 450, 453, 456, 458, 459, 462, 465, 468, 469, 471, 472, 474, 477, 480, 483, 485, 487, 489, 493]. Article [496, 499, 501, 503, 506, 509, 512, 515, 518, 521, 524, 526, 530, 534, 537, 539, 542, 545, 548, 552, 555, 560, 563, 566, 569, 572, 575, 578, 580, 582, 585, 588, 591, 594, 597, 600, 603, 604, 607, 610, 614, 617, 620, 622, 625, 628, 630, 633, 636, 639, 642, 645, 648, 650, 653, 656, 658, 661, 663, 666, 669, 672, 675, 677, 680, 683, 685, 687, 690, 693, 696, 699, 702, 705, 708, 710, 713, 716, 719, 721, 726, 729, 733, 736]. Article [739, 742, 745, 748, 751, 754, 757, 760, 763, 766, 770, 773, 775, 778, 781, 784, 788, 790, 793, 796, 799, 803, 807, 810, 814, 817, 820, 823, 825, 828, 831, 834, 837, 840, 844, 847, 850, 853, 855, 858, 861, 865, 868, 871, 874, 877, 882, 885, 889, 892, 896, 942, 945]. Artificial [2537, 486, 461, 460, 1761, 619, 835, 236, 1639, 632]. Artificially [869]. Artisanal [2415, 2039]. Artist [2213, 1955]. Artists [1177]. Asia [1349]. Asian [2374, 1415]. Aspects [277, 2123]. Assays [1837]. Assemblage [1332]. Assemblages [2922, 1187, 1981, 404, 615]. Assembling [3174]. Assess [1794, 925, 1775, 1747, 168, 1855]. Assessing [718, 523, 930, 417, 701, 2701, 1872, 2085, 1267, 149, 433, 1990, 538, 1014, 704, 55, 1702, 3149, 761]. Assessment [2750, 1504, 1515, 879, 878, 416, 1842, 649, 50, 2834, 2585, 634, 358, 208, 895, 1957, 2011, 779, 862, 89, 1570, 1187, 236, 2312, 589, 695, 279, 97, 725, 762, 845, 290, 2100, 380, 2833, 1598, 90, 69, 1640, 553, 505, 1069]. Assessments [1124, 1145, 1856, 936]. Assignment [3279]. Assist [75, 1214]. assisted [383, 562]. Associated [930, 2671, 2401, 596, 723, 1922, 538, 651, 809, 504]. Association [1823, 1616]. Associations [281]. Assumptions [590, 735]. Assurance [439]. At-Sea [1798]. Atlantic [2239, 2255, 843, 2308, 750, 98, 686, 2341, 914, 930, 1041, 177, 771, 3115, 2137, 2570, 1095, 179, 1894, 2631, 236, 635, 1344, 294, 229, 1332, 2440, 123, 2309]. Atlas [2601, 60]. Atop [1741]. Attempts [121]. Attend [2140, 3123, 1349, 2831, 2670]. Attendance [2313]. Attendee [1727]. Attention [1325]. Attitudes [73, 631, 114, 139, 2584, 700]. Attraction [486]. Attractions [2882]. Attractive [2180]. Auburn [1244]. Audience [2337, 1950]. August [1180, 1190, 1268, 1831, 2037, 2051, 1959, 2239, 2109, 2255, 2673, 2754, 2770, 3037, 3020, 1270, 2265, 2911, 3268, 2936, 2755, 1064, 1235, 1246, 1408, 2383, 2542, 2720, 2876, 3180]. Austen [3272]. Australia [1432, 1820, 798, 891, 1241]. Australian [238, 1384]. Australis [2294]. Authentication [1922]. Authors [904, 960, 1130, 1301]. Authorship [2879]. Autocidal [838]. AUV [1344]. Available [1683]. Avoid [1414]. Avoiding [1793]. Award [1283, 1284, 1453, 1663, 1876, 2091, 2283, 2443, 2616, 2784, 1744, 3125, 2126, 2132, 3045]. Awardees [2133]. Awards [961, 976, 1128, 1284, 1454, 1453, 1664, 1663, 1874, 1876, 2090, 2282, 2283, 2442, 2443, 2615, 2616, 2783, 3242, 2709, 2952, 3091, 1314]. Awareness [1462, 2275, 2036, 767, 2300, 818, 2102].

B [1189, 2330, 2935]. Baby [2247]. BACI [3201]. Back [1665, 100, 1630, 1914, 1928, 1962, 1975, 1999, 2029, 2045, 2061, 2075, 2096]. Backpack [2821]. Bag [1775]. Bahamian [2596]. Bain [1189]. Baird [592]. Bait [2425, 894, 1532, 3, 436, 414, 3240]. Baited [115]. Baitfish [351]. Balancing [3066, 863]. Baltimore [2754, 2770, 2882, 2913, 2870, 2831]. Ban [58]. Banking [926]. Bans [3277]. Bar [1788]. Baremore [1841]. Barotrauma [1499, 3055, 2714]. Barrens [1713]. Barrier [2600, 2108]. Barriers [3008]. Barton [1522]. Base [364]. Based [2891, 1533, 1175, 1759, 1871, 2890, 2218, 2078, 2134, 777, 2264, 1419, 808, 2373, 2532, 2570, 813, 895, 2521, 676, 2244, 3126, 774, 2176, 1734, 2194, 2807, 1775, 1868, 1598, 2571, 704, 2409, 396, 2509, 561, 797, 873, 759, 495, 2833]. Baselines [2377]. Bashing [2735]. Basic [124]. Basics [2627]. Basin [2834, 1097, 2586, 299, 2230, 2375, 1783, 2167, 3264, 659, 1027, 1485, 2509, 2058, 925, 2340, 154, 829, 3238, 2439, 2229, 1098, 1722, 691, 1227, 2844, 706]. Basins [1018, 2121]. Basis [1980]. Bass [623, 2671, 103, 681, 20, 3101, 2387, 2749, 2948, 3265, 3082, 1348, 728, 1793, 2764, 3174, 398, 2806, 2454, 2453, 1716, 753, 467, 504, 670, 173]. Basses [135, 797]. Bathwater [2247]. Baton [3182]. Battle [543, 2697]. Bay [2455, 822, 286, 177, 2535, 804]. Bayesian [2329]. Bayous [1783]. Be [963, 16, 1773, 2823, 2657, 155, 1838, 1794, 2762, 1853, 1727, 917, 2896, 1775, 1852, 1674, 605, 2102, 883, 2292, 2422, 1148, 3222]. Beach [3076]. Bears [2674, 752, 2127]. Beasts [369]. Beautification [2118]. Beauty [2305]. Beaver [1198, 1741]. Beeman [1540]. Beginning [3054, 1555]. Beginnings [3051]. Behalf [1382]. Behavior [238, 1819, 851, 1761, 1254, 1361, 1780]. Behavioral [2166, 2738, 2584, 2948, 2040]. Behind [1560, 2213]. Behnke [1502]. Being [1148, 3062, 1790, 917, 498]. Belize [613]. Belonging [3003]. Below [2305]. Beltway [1606]. Bench [1804]. Benchmarking [2121]. Beneath [3225]. Benefit [2360]. Benefits [1172, 626, 1484, 2401, 1613, 1644, 1245, 562, 1393, 2084, 809, 2806]. Benign [1158]. Benthic [863]. Bergman [1940]. Bering [327]. Berkeley [1087, 1257, 1282, 1451, 1452, 1661, 1875, 2055, 2252, 2428, 2587, 2753, 2912, 3056, 3211, 1660, 1662, 1659]. Best [2132, 2777, 1936, 1675, 2108, 2894, 2822, 1314]. Bet [1954]. Bethesda [1446]. Better [1541, 2674, 1397, 1411, 159, 1837, 2897]. Between [387, 611, 1074, 747, 881, 2629, 2646, 2822, 849, 2454, 1362, 1370, 2228]. Beyond [2654, 1312, 2108, 2099, 1807]. Bias [2518, 3027]. Biased [2762]. Biases [1517]. Big [3004, 1866, 2104, 1994]. Bigford [2479]. Billfishes [2647]. Billings [3044]. Billion [1538]. Billion-Dollar [1538]. Billy [1620]. BioBlitz [2036]. Biodata [2182]. Biodiverse [2388]. Biodiversity [1851, 306, 303, 2232, 573, 2184, 355, 2738, 2337, 3054, 2398, 308, 1213, 441, 307, 305]. Bioelectrical [1872]. Bioenergetic [2767]. Bioenergetics [2264, 318]. Biofiltration [2180]. Biographical [3229]. Biological [1504, 2010, 1553, 1783, 783, 33, 2715, 304, 277, 356, 55]. Biological-Objectives [783]. Biologist [3059, 631, 1133, 118, 2536, 400, 2007, 203]. Biologists [794, 2010, 954, 253, 866, 442, 2780, 2162]. Biology [969, 2087, 2545, 2723, 2763, 956, 946, 851, 1312, 463, 2994, 2868, 1007, 1522]. Biomass [2698]. Biota [718, 436]. Biotelemetry [2302]. Biotic [89, 219, 466]. Biphenyl [8]. Birds [128]. Bireley [2643]. Birth [2193]. Bit [2683]. Bite [2866]. Bite-Shortened [2866]. Black [3206, 2671, 3101, 2749, 135, 3082, 797, 1348, 2806, 2453, 1716, 504, 670, 173, 3213, 2454]. Blacksburg [1205]. Blends [1298]. Blitz [2591]. Bloaters [2391]. Blocked [1914]. Blogging [1654, 1853]. Blooms [887]. Blue [2049, 1512, 1836, 1642, 2778, 2789, 2344]. Blue-Ribbon [2049, 1836]. Bluefin [1404]. Bluefish [1657]. Bluegill [3012, 1775]. Blues [1043]. Board [966, 982, 1710]. Boat [1752, 3200, 1598]. Boat-Based [1598]. Boating [551]. Boats [917]. Bob [3186, 1420, 1271]. Boca [168]. Body [100, 1469, 1470, 1771]. Bomb [2597]. Bonar [2479, 1901]. Bonefish [2052, 2181]. Bonefishes [2711]. Boogers [1006]. Book [1604, 1522, 3284, 1540, 1015, 2846, 141, 1539, 3267, 978, 2185, 1568, 2662, 2586, 1117, 2192]. Books [3284]. Booth [1351, 1903]. Boots [887]. Border [2710]. Borders [1149]. born [1127]. Borrowing [3279]. Both [1145, 3173, 1868]. Bottom [50, 2712, 2533, 2521]. Bottom-Trawl [50]. Bottomfish [3079]. Boulder [1006]. Boundaries [2531, 2075, 3262]. Bow [2049]. Bowfin [2288, 2797, 2802, 2231]. Bowfins [313]. Bowfishing [2974, 2932]. Bowker [2479, 2468]. Box [1045]. Boyle [2221]. Bradley [2935]. Brazil [1265, 2164, 1332]. Brazilian [3225]. Breach [2886, 3153]. Break [3015, 37]. Breaking [373, 1100]. Brett [3044]. Brian [2153]. Bridge [2038, 2899]. Bridge-Building [2038]. Bridging [3188, 2454]. Brief [2699, 2038]. Briefings [1749]. Briggs [876]. Bring [2008, 2323]. Bringing [1618, 1547, 1273, 2247, 1843]. Brink [1099]. Bristol [2455, 2535]. British [887, 3284, 689, 1343, 476, 2467, 445, 2883, 665]. Broad [2018, 881, 3175, 2819, 240]. Broad-Scale [2819]. Broadening [1677]. Broader [2328, 2734, 3262]. Brook [1854, 740, 1712, 1295, 2277]. Browns [1652]. Bruce [1522, 3227, 3268, 1955]. Bruner [2989]. Brushes [1714]. Buck [1073]. Buckets [436]. Budding [2162]. Budget [2348, 349]. Buffalo [2940]. Bugas [2662]. Build [794, 2463, 3139, 3042]. Builder [2725]. Building [3040, 2038, 1818, 1066, 1229, 3280, 3129, 759]. Bull [833, 2405, 2339, 859, 2815, 533]. Bumpy [2193]. Burbot [2665]. Burkett [1690]. Burrowing [238]. Business [1778, 2856]. BVET [706]. By-catch [295, 294]. Bycatch [494, 2105, 2104, 1414]. Bypass [605]. Bytes [1514].

C [2071, 3113, 2662]. C. [1782]. ca [950]. Cabecera [811]. Calcified [3043]. Calculating [2042]. Calculation [100]. Calendar [905, 906, 1063, 1705, 2389, 2560, 2576, 2589, 2603, 2617, 2622, 2847, 2953, 2964, 2981, 3046, 3057, 3080, 3092, 3104, 3118, 3132, 3143, 3156, 3166, 3177, 3191, 3204, 3215, 3230, 3243, 3255, 3269, 3285]. Calendars [963]. California [2583, 1388, 508, 423, 993, 3078, 3076, 355, 497, 2301, 273, 510, 944, 692, 1775, 1793, 605, 2573, 553]. Call [1113, 1086, 1101, 1284, 1285, 1256, 1269, 1453, 1406, 1455, 1421, 1438, 1603, 1619, 1629, 1656, 1663, 1662, 1534, 1832, 1845, 1860, 1870, 1876, 2051, 2070, 2072, 2086, 2091, 2055, 2126, 2283, 2269, 2255, 2443, 2575, 2602, 2588, 2616, 2784, 3203, 626, 2522, 2134, 988, 826, 1397, 3224, 2660, 3138, 1966]. Calling [1710]. Calls [2747]. calva [2802, 2231]. Camera [1767]. Cameras [1843]. Campaigns [1968]. Campbell [1374]. Campus [3248]. Can [939, 16, 1563, 2653, 1215, 2183, 1312, 366, 2582, 337, 1838, 1794, 310, 888, 2907, 2242, 3240, 284, 2762, 3072, 2809, 1989, 1968, 917, 2930, 2958, 2482, 1775, 1760, 1731, 883, 1920, 2533, 583, 370, 605]. Canada [257, 2699, 1681, 1110, 771, 2710, 723, 1070, 1431, 755, 1360, 1908, 1334, 351, 535, 1780, 797, 929, 621, 1014, 1766, 427, 818, 2856, 345, 2302]. Canadian [1553, 1369, 2598, 2867, 1633, 1984, 1213, 665]. Canal [33]. Candidate [1139, 1140, 1690, 1691, 1901, 2468, 2469, 3283, 2824, 2644, 3282, 2153, 2314, 3130, 2825, 2315, 3131, 2645, 1902, 2154, 2976, 2977]. Candidates [1933, 3129, 1934, 2975]. Can’t [1935]. Cape [1555]. capitalization [1337]. Capitol [1749, 1399, 2864]. Captains [195]. Captive [1404, 3029, 3111].

Capture [922]. Capture-Recapture [922]. Carbon [1512, 2855, 2276]. Carcass [800]. Carcasses [540, 869]. Care [3176, 1359, 2671, 2412]. Career [3171, 2018, 175, 253, 1810]. Careers [2004]. Cares [366]. Caribbean [3141, 3116, 3128, 3019, 2832, 792]. Carlos [1556]. Carolina [1752, 295, 82]. Carp [196, 2487, 3232, 2374, 317, 764, 5, 1415]. carpio [2487]. Case [819, 2002, 1565, 623, 851, 664, 3074, 79, 1677, 1070, 508, 318, 1308, 1599, 2822, 1780, 2880, 2844, 2206]. Casimir [3113]. Caspian [3264]. Cast [1096, 1693]. Casts [3262]. Catalog [2924]. Catalysts [1386]. Catastrophe [2869]. Catch [3004, 1752, 2701, 1001, 2657, 2412, 1761, 1186, 2085, 3087, 1267, 2168, 2519, 947, 3110, 590, 1807, 294, 2465, 2294, 295]. Catch-and-Release [1761, 3087, 1267, 2168, 2519]. Catchability [20]. Catchable [2878, 208]. Catches [3200]. Catching [1807]. Category [2596]. Catfish [631, 2698, 1445, 1417, 535, 1216, 2778, 1868, 2833, 2789]. Cating [749]. Caught [2465, 670]. Causation [1891]. Cause [2068]. Cause-and-Effect [2068]. Caused [2891, 2890]. Causes [2152, 564]. Causing [1533]. Cautionary [3223]. Caviar [2797]. Ceiling [373]. Celebrate [2347, 2609]. Celebrating [2280, 3280, 2960, 2609]. Celebration [2777]. Census [1240, 2424]. Center [911, 965, 981, 1959, 791, 1486, 2020]. Centered [1829, 935, 1253]. Central [519, 414, 279, 2573, 553, 488, 261]. Century [2639, 1281, 2437, 2041, 243, 298, 1471, 584, 3043, 380, 3261]. cerebralis [470]. Certainly [1719]. Certification [1802, 1930, 1020, 2752, 3261, 2006, 1245]. Certified [1031, 1377]. Chains [2068]. Challenge [1678, 1511, 198, 842, 136]. Challenges [1311, 1324, 1336, 1392, 923, 1910, 626, 1237, 2167, 2687, 3062, 1469, 1470, 1051, 2702, 2521, 2600, 3173, 2122, 2197, 2041, 2711, 2099, 2611, 3162, 2865]. Chamberlain [203]. Champion [1391]. Champlain [1565]. Change [2891, 1678, 1917, 2683, 2890, 1992, 1993, 1987, 599, 1809, 1977, 2507, 265, 270, 3263, 269, 3222, 1982, 1439, 267, 3175, 1985, 2893, 1981, 2069, 1218, 1601, 2336, 1904, 2387, 298, 2818, 1991, 1979, 2895, 1322, 1984, 1348, 1856, 266, 1989, 1968, 2150, 2676, 1868, 2251, 1399, 264, 2892, 3041, 1980, 1779, 2123, 1988, 2422, 2491, 2899, 2897, 928, 950]. Changed [1800]. Changes [296, 1993, 1794, 1337, 1956, 1887, 2451]. Changing [1100, 647, 1986, 1752, 3188, 1108, 1155, 18, 1695, 1055, 3187, 857, 1990, 2889, 1983, 1893, 3017]. Channel [1868]. Chapter [3210, 905, 963, 962, 2687, 2532, 2464, 2871, 3008, 1967, 977, 1230, 3228, 3214, 1217, 1205, 1204]. Chapters [1493]. Char [3179, 859, 160, 2690]. Characteristics [1143]. Characterizing [1808, 1859]. Charles [3102]. Charr [117]. Charter [1752, 727, 195, 321, 455, 587, 1750, 1280, 1751, 1598]. Charting [2479]. Chatwin [3284]. Chemical [2053]. Chemicals [2657, 1361]. Chemistry [122]. Chemosensation [1361]. Chesapeake [286]. Child [3176]. Children [1294]. Chimeras [2883]. China [1735]. Chinese [2178]. Chinook [3258, 574, 3188, 541, 1775, 827, 2795, 553]. Chloramine [1589]. Chloramine-T [1589]. Choice [1676, 541]. Choices [3004, 787]. CHONe [1213]. Chronology [651]. Chub [1867, 1796, 2203, 1857]. Cichlid [2185].

CIMTAN [1369]. Citations [2465]. CITES [792]. Cities [1072, 1101, 1190, 1113, 1086, 1164, 1268]. Citizen [1936, 3099, 1777, 3076, 3072, 3079]. Citizenry [2275]. City [1921, 1903, 1959, 2239, 2255, 1406, 1455, 1421, 1438, 1505, 1935, 2341, 1658, 2089, 2308, 2440, 2309]. Civil [2668]. Clam [2659]. Clapp [1501]. Claramunt [1501]. Clark [435]. clarkii [3223, 3126, 3126]. Class [2763, 2661, 172, 1696, 2465, 1868, 3174]. Classes [3045, 2808]. Classic [60, 2477]. Classics [1852, 1793]. Classification [57, 2520, 226, 3140]. Classifying [336]. Classroom [3040, 1636, 2136, 1635, 941]. Claytor [2333]. Clean [602, 2697]. Clear [801, 2747]. Clear-water [801]. Clearinghouse [1699]. Climate [2891, 1678, 1917, 2683, 2890, 1992, 1993, 1987, 1809, 1986, 1977, 2507, 3188, 265, 270, 3263, 269, 3222, 1982, 1439, 1281, 267, 3187, 3175, 1985, 2893, 1981, 2069, 2898, 2336, 211, 3238, 1990, 2889, 2894, 1991, 1983, 1979, 2895, 1322, 1984, 3017, 266, 1989, 1868, 1399, 264, 2892, 1980, 1779, 2123, 1988, 2491, 2899, 2897, 2855]. Climate-Friendly [2855]. Climbing [1665]. Clint [2536]. Clinton [2658]. Close [2927]. Cluster [1884]. Co [3214, 2879, 1259, 2302, 2812]. Co-Authorship [2879]. Co-hosted [3214]. Co-Management [2302]. co-production [2812]. Co-winners [1259]. Coalition [949]. Coarse [776]. Coast [1719, 3073, 513, 1801, 1985, 848, 611, 1158, 179, 3225]. Coastal [806, 2712, 267, 2168, 476, 3126, 3078, 1344, 692, 478, 804]. Cod [686]. Code [2478, 1172]. Coeur [1217]. Cofradías [2494]. Coho [822, 2921, 1485, 779, 869, 478]. Cohosted [3228]. Colapso [811]. Coldwater [571, 3187]. Collaborating [993, 2337]. Collaboration [2135, 2295, 2710, 1613, 1957, 949, 2660, 2822]. Collaborative [3064, 1081, 2494, 3078, 1074, 1698, 1111]. Collapse [621, 811, 1415, 750]. Colleagues [366]. Collect [2182]. Collected [1871]. Collection [3063, 2959, 2987, 2906, 2944, 3086, 2909, 3097, 3124, 3137, 3172, 3197, 3249]. Collections [2688]. College [1844]. Colleges [183]. Collision [955]. Color [3100]. Colorado [2230, 1027, 2340, 2229, 2375, 250, 746, 2916, 856, 577]. Colorful [2319]. Columbia [2586, 887, 482, 2167, 3264, 689, 574, 1485, 925, 476, 154, 829, 1343, 1098, 2066, 2736, 2467, 1722, 445, 2883, 691, 1227, 2844, 520, 665, 3284]. Columbus [2673, 2673, 2670]. Combining [3140]. Come [2777, 3221, 2504]. Commandments [808]. Comment [2230, 2246, 101, 2247, 1131, 1457, 1433]. Commentary [1632, 3221]. Comments [2248]. Commercial [1125, 2401, 860, 86, 3278, 694, 2437, 1042, 2352, 2948, 2765, 3127, 529, 140, 2932, 2833, 94, 2572, 1679]. Commission [2729, 98, 168, 2327]. Committee [1129, 1380, 1109, 2574, 1626, 2479, 3129, 2427, 2454, 632]. Common [2487, 2510, 1697, 1564, 1337, 1337, 3237]. Commons [1533, 172]. Communicate [2744, 2408]. Communicating [1447, 806, 1964, 1011, 2351, 1302, 2490, 2894, 3054, 824, 2633]. Communication [2010, 1390, 2324, 647, 2687, 2990, 1762, 857, 1393, 2818, 1536, 1465, 2493, 2934, 1905, 3252, 2034, 2899]. Communication-Online [1465]. Communication/Cooperation [1762]. Communications [1611, 1788, 2260]. Communicators [1757]. Communities [1125, 296, 1681, 3148, 647, 863, 612, 89, 2596, 1026, 2360]. Community [806, 1586, 880, 3280, 2920, 2266, 695, 759, 776, 789, 136]. Community-based [759]. Comparative [812, 1752, 2120]. Compare [1871]. Compared [179]. Comparing [703]. Comparison [400, 611, 723, 1199]. Comparisons [461, 771, 20]. Compatibility [143]. compatible [605]. Compensation [723]. Competition [200, 172]. Competitive [2671, 660, 276, 277]. Competitiveness [895]. Competitors [568]. Complement [2842]. Completed [144]. Completes [1640]. Complex [558, 2136, 3263]. Compliance [250, 767]. Components [678]. Composition [1752]. Compounds [8]. Comprehensive [144, 239, 1722, 2844]. Computer [17, 383, 318, 774, 163]. Computer-assisted [383]. Computers [1495]. Computing [2654]. Concentrate [2452]. Concept [367, 830, 902, 2182, 608]. Concepts [2399]. Conceptual [718, 103, 2373, 3139]. Concern [1665, 421, 242]. Concerns [426, 846, 246, 304, 299, 1647]. Conch [792]. Condit [1939]. Condition [1872, 179, 1747, 1069, 824]. Conditions [1783, 925, 3127, 762, 640]. Conduct [2478]. Conducting [115, 384]. Conference [1206, 1467, 1572, 1602, 1618, 1786, 2162, 1504, 1515, 1534, 2613, 2423]. Conferences [2253, 2285, 2124, 2141, 2156, 2171, 2186, 2201, 2222, 2236, 2267, 2303, 2317, 2331, 2342, 2353, 2367, 2379, 2403, 2416, 2429, 2444, 2457, 2470, 2485, 2497, 2512, 2525, 2538, 2648, 2663, 2678, 2691, 2703, 2716, 2727, 2740, 2757, 2787, 2771, 2813, 2800, 2826, 2836, 2859, 2872, 2884, 2900, 2914, 2925, 2938, 2996, 3010, 3021, 3034, 3067]. Conferencing [3030]. Conflict [881, 715, 1068]. Conflicts [623, 747]. confluentus [2405, 2815]. Confronting [2677, 1175]. Congress [3253, 2661, 2858, 2881, 1126, 3141, 3116, 3128, 2023, 3276, 3019, 3002]. Connect [2943]. Connecting [1401, 2676, 356]. Connection [1108]. Connections [1818, 1081, 1169, 2098]. Connectivity [1789]. Connects [1342, 2008, 2734]. Consequence [720, 769]. Consequences [2152, 2067, 1993, 701, 120, 1868, 564]. Conservation [2217, 939, 2414, 2639, 969, 1087, 1257, 1282, 1451, 1452, 1661, 1875, 2055, 2252, 2428, 2587, 2753, 2912, 3056, 3211, 1578, 1676, 2386, 2910, 2167, 2134, 2232, 1027, 859, 673, 1818, 2585, 756, 1075, 2137, 863, 2311, 611, 2400, 1973, 854, 933, 1360, 1585, 2340, 1084, 955, 1417, 3173, 2234, 2122, 2700, 1970, 2898, 162, 1035, 2427, 2490, 893, 1646, 2337, 3265, 2084, 1098, 1971, 1322, 2711, 1614, 475, 2100, 2398, 2235, 2520, 2933, 98, 613, 1144, 1746, 1793, 926, 643, 2263, 1566, 1938, 2994, 1598, 2797, 3138, 1213, 1659, 427, 818, 2454, 2453]. Conservation [1857, 1716, 305, 579, 3111, 1214, 345, 824, 1026, 1660, 2566, 67, 2812, 2179, 1573, 2245]. Conservationist [2213]. Conserve [2183, 1775]. Conserving [3055, 943, 664, 308, 1698]. Consider [1372]. Considerations [2083, 1987, 3264, 2748, 1501, 673, 863, 101, 629, 115, 2427, 295, 140]. Consistency [364]. Constantine [1347]. Constitutional [1017].

Constraints [97, 798]. Constructed [720]. Construction [131]. Constructivist [731]. Consultant [2021]. Consultants [255]. Consulting [118]. Consumed [2510]. Consumption [364, 2744, 291, 58, 206, 290, 2300]. Containing [596, 160]. Containment [646]. Contaminant [416]. Contaminated [416]. Contemporary [2796, 999, 2688, 2768]. Content [315]. Contentious [2633]. Contest [1299, 1259, 1821, 1216]. Context [1537, 880, 902, 2398]. Contextual [1215]. Continental [2898]. Continental-Scale [2898]. Continue [1734, 2348]. Continued [2697]. Continues [1820]. Continuing [1019, 1190, 1352, 1543, 1753, 1959, 2169, 2355, 2511, 3006, 3155, 3115, 403, 2364]. Continuum [1627]. Contrasting [1058, 478]. Contribute [1389, 2733, 2930]. Contribution [1637, 141, 168, 216]. Contributions [540, 3074, 612]. Control [33, 439, 317, 1428, 1252, 3201, 838, 577, 165, 605]. Controls [1198]. Controversy [709, 623, 329, 430]. Convene [3019]. Convention [1486]. Conversation [2608, 2244, 2192, 2505]. Conversion [805]. Convincing [1447]. Cool [1853]. Coolwater [285]. Coop [2898]. Cooperation [430, 278, 1762, 411]. Cooperative [783, 860, 133, 194]. Coordinated [2639, 567]. Coping [1760]. Coregonines [3193]. Coregonus [3175]. Cormorants [2233]. Coronavirus [2683]. Correction [1385, 1490, 1685, 1703, 1754, 1862, 1869, 1942, 1960, 2237, 2540, 2634]. Corrective [16]. Correlation [1891]. COSEWIC [1014]. Cost [2878, 2120, 187, 895, 29, 1708, 704, 2365, 919]. Cost-Effectiveness [895]. Could [1729, 2451]. Council [659]. Counterintuitive [875]. Counterpoint [1773]. Countries [31, 308]. Counts [1712]. Couple [2188]. Course [2479, 3155, 618, 955, 857, 2476]. Courses [2012]. Court [697, 2349]. Court-Ordered [697]. Courtenay [1542]. Cover [1955]. Coverage [1859]. COVID [2693, 3161, 2893, 2809, 2958, 2909]. COVID-19 [2693, 2809, 2909]. Cowan [2936]. Cows [1583]. Crab [543, 14, 2352, 1852]. Crabs [1925]. Cracks [3052]. Craft [1362]. Craig [1854]. Cramer [3185]. Crappie [2718, 3206, 3201]. Crappies [3110]. Crash [977]. Crater [2339]. Crawfish [3]. Crayfish [894, 2082, 3150, 972, 285]. Crayfishes [1084, 573, 427, 818]. Create [2920, 1837]. Created [1655]. Creates [2349, 832]. Creating [2328, 2295, 2599, 880, 3043]. Creation [570]. Creational [2450]. Creative [1836, 2641]. Creativity [1905]. Credibility [881, 1020, 2247]. Credible [2493]. Crediting [926]. Creek [2795, 3069]. Creel [599, 2924, 797]. CreelCat [2924]. Crested [2233]. Crisis [2809, 1760]. Criswell [2185]. Criteria [63, 703, 546, 82, 3018, 856]. Critical [2466, 1110, 2266, 3005, 2426, 347]. Critically [3225]. Crop [3]. Cross [1467, 2295, 1868]. Cross-Class [1868]. Cross-Institutional [2295]. Cross-Sectoral [1467]. Crossing [1484, 3002]. Crossings [720]. Crossroad [1043]. Crossroads [273]. Crowd [2400, 3072]. Crowd-Sourced [2400]. Crown [1959]. Croxton [2469]. Crushed [122]. Ctenopharyngodon [5]. Culprit [1676]. Cultural [638, 1601, 356, 564]. Culture [1522, 3040, 1793, 1346, 69, 1132, 3111]. Culverts [1676, 2084]. Cumulative [813, 3001]. Curator [1763]. Current [1643, 1533, 23, 2386, 2038, 550, 304, 3151, 816, 1074, 2583]. Currents [958]. Curricula [129]. Curry [1139]. Curtis [2658]. Curves [2767]. Cutthroat [2414, 2639, 2983, 3273, 3223, 3263, 424, 2501, 752, 3126, 2971, 226]. Cycle [176]. Cycles [342]. Cyprinus [2487]. Czech [3151].

D [1350, 3267, 3268, 2662]. Daggernose [2483, 3225]. Dahl [1733]. Dakota [2582, 2610, 2534]. Dale [1690]. d’Alene [1217]. Dall [323]. Dam [2621, 1557, 2066, 1450, 1820, 2638, 1939, 2886]. Damages [704]. Dams [730, 3153]. Damsel [1888]. Daniel [1684]. Danish [1275, 2632]. Dark [2145, 2021, 1833, 1928]. Darter [3023, 2240, 2758, 2036, 2955]. Data [1125, 2750, 224, 1871, 3099, 50, 1041, 2425, 364, 3279, 2906, 1767, 460, 79, 2275, 1124, 2293, 1794, 2085, 1308, 2351, 359, 2907, 2842, 3075, 3072, 1519, 1348, 1938, 2833, 1855, 1880, 2909, 2042, 768, 1994]. Data-Limited [2833]. Database [2057, 827, 1069]. Databases [2400, 1601, 1994]. Datasets [2832]. Daughter [1945]. David [1073, 1732]. Day [1085, 2620, 2734, 2697, 1483, 3280, 1701]. Dead [1677]. Deadhead [862]. Deadliest [2294]. Deadline [2119]. Deadly [1612]. Debate [1883, 1041]. Debby [3288]. Debris [1729, 862]. Decade [3280]. Decades [2573, 3111]. Deceits [884]. December [2995, 2979, 2739, 1119, 1291, 1303, 1472, 1686, 1912, 2111, 2433, 2304, 2607, 2775, 2929, 3083, 3233]. Decent [2462]. Decide [2845]. Deciding [568]. Decision [1435, 426, 987, 239, 2600, 2069, 2108, 654, 1450, 1322, 452, 75]. Decision-making [426, 239]. Decisions [833, 939, 2199, 2015, 2363, 2640]. Deckhand [1623]. Decks [527]. Declaration [1665, 1966]. Declared [848]. Decline [1353, 576, 752, 488, 728, 613, 2764, 3174]. Declines [2971, 1665]. Decontamination [1252]. Deep [2158, 532, 2173]. Deep-sea [532]. Deeper [2531]. Deepwater [1089, 1042, 1702]. Deer [3069]. Defensible [1970]. Deficit [550]. Defilippi [2479]. Defilippi-Simpson [2479]. Define [546]. Defining [916, 1890]. Definition [612, 2655]. Degradation [920]. Degree [1429]. DEIJ [3109]. DEIJA [3184]. Déjà [2743]. Del [811]. Delaware [914, 2168]. Delphi [870]. Delta [3036, 2301, 1793]. Demand [419]. Demographic [298]. Demographics [2549]. Demonstrate [40]. Demonstrates [172, 3184]. Demonstration [845]. Denise [1605]. Department [1244]. Dependent [103, 128, 612, 872, 2596]. Depensation [2854]. Depleted [386]. Depredation [2714]. Dequine [1162]. derived [476]. Descender [2714]. Descending [3055]. Deschutes [2418]. Describing [2039]. Descriptions [3155, 2482]. Design [3216, 311, 2083, 1892, 359, 703, 82, 2084]. Designated [1212]. Designing [3077]. Designs [1484, 294]. Detailed [3202]. Detect [3042]. Detected [1692]. Detecting [1370]. Detection [2107, 2120, 922, 1892]. Determination [815, 749]. Determine [63, 498]. Determining [1501, 2053]. Dettmers [1500]. Develop [1213, 2751]. Developed [498, 1214]. Developing [1279, 851, 551, 2050, 671, 1970, 31, 1536, 252, 3174, 49, 308, 2797]. Development [2537, 2890, 1499, 609, 2082, 1298, 1075, 2295, 2509, 1124, 222, 175, 253, 1149, 2924, 236, 800, 2685, 229, 827, 1735, 856, 1679, 1720]. Developments [177]. Devices [3202, 717, 3055, 2714, 294]. Devils [870, 180]. Diadromous [854, 2350, 581, 789]. Diagnosing [715]. Diagnosis [386]. Dialog [1904]. Diane [2211]. Dichotomy [1999]. Dick [1668]. Dictionary [3229, 1124]. Did [1706, 2077, 2205]. Didymo [887, 1252]. Diem [2718]. Diet [1692]. Diets [1820]. Differ [787]. Difference [2594]. Differences [330, 946, 1826, 2300, 640]. Different [939, 2923, 121]. Differential [869]. Difficult [1461, 832, 1775]. Difficulties [574]. Diffuse [1970]. Digest [1482]. Digital [1390, 1416, 2356, 2672, 3043]. Digitizing [100]. Dilemma [486, 1890, 259]. Dimension [2040]. Dimensions [1808, 448]. Dingell [92]. Dioxide [2276]. Direct [988, 3017, 2794]. Directional [897]. Directions [1537, 3277, 816, 2907]. Directive [1537]. Director [1720, 989, 974, 950]. Directors [2508]. Disaster [848, 549, 2360]. Disasters [1191]. Discarded [1675, 947]. Discarding [2712]. Discharge [606]. Discipline [2689, 640]. Disciplines [989, 3188]. Discovered [3223, 424, 2173]. Discoveries [435]. Discrepancy [2505]. Discussions [3063, 2959, 2987, 2944, 3086, 3097, 3124, 3137, 3172, 3197, 3249]. Disease [2677, 1258, 3263, 525, 1747, 3174]. Diseases [1403, 1589]. Diseases-What [1589]. Disinfection [1760]. Disparate [2584]. Dispelling [517]. Dispersal [1157, 670]. Disrupted [1275]. Disrupters [1793]. Disruptive [2302]. Distance [941]. Distortion [1552]. Distress [1888]. Distribution [470, 707, 2120, 730, 579]. Distributions [758, 2057]. Distributive [324]. Districts [39]. Disturbance [2583, 2399]. Diverse [907, 3040, 3016, 2018, 2547, 2733]. Diversion [1775, 27]. Diversions [141]. Diversity [2879, 2545, 2548, 2543, 519, 2970, 2557, 3277, 2550, 2549, 2908, 304, 2558, 2555, 2114, 1468, 2209, 2291, 2556, 1694, 1231, 2554, 497, 2852, 2948, 3265, 2546, 1971, 2210, 1828, 356, 2453, 1716, 2544, 579]. Division [905, 963, 2070, 2062, 632, 3153, 3210, 3228, 3214, 519, 1778, 279, 3198]. DNA [2120, 1349, 1375, 1699, 1922, 827, 370]. Do [1609, 2273, 1202, 1250, 2199, 554, 1004, 2701, 460, 1411, 1528, 3271, 1652, 2191, 2732, 1496, 891, 2008, 370, 2228, 1173]. Docks [940]. Documentaries [1287]. Does [2945, 1621, 2931, 2714, 270, 2557, 2908, 2558, 2555, 629, 2556, 2554, 1161, 2452, 1828, 1905, 2892, 913, 883, 867]. Doesn’t [1918]. Doing [888, 2291, 1853]. Dollar [1538]. Dollars [1932]. Dolly [859]. Domestication [1793]. Dominica [2582]. Donald [2785]. Done [917]. Donna [992]. Donors [3117]. Don’t [963, 3222, 2247]. Donut [327]. Dots [1401]. Double [2233, 101]. Double-Crested [2233]. Doubt [1041, 2032]. Doug [3272]. Doughnuts [1873]. Douglas [2185]. Down [1805, 1882, 692]. Downfall [2165]. Downstream [2906]. Dox [1350]. Doxtater [1350]. Dr. [1147, 1374, 1542, 1321]. Draft [1558]. Dragonfishes [2173]. Drainage [1018, 42, 958]. Drainages [488]. Dredging [500]. Dressed [2499]. Drift [2742]. Driven [2387]. Drivers [1572, 1601]. Driving [1809, 3200]. Drones [1570]. Drop [891, 2192]. Drought [1978, 1775, 2991]. Drug [1448]. Drum [3168, 624]. Dual [1393]. dubio [2032]. Dues [2226]. During [2693, 2696, 2671, 2853, 3196, 243, 2809, 2991, 1838, 2893]. Duties [139]. Dwarf [160]. Dwelling [954, 200]. Dynamic [473]. Dynamics [2978, 1553, 2026, 2058, 2266]. Dysfunction [249].

Earliest [815]. Early [2107, 1892, 815, 175, 253, 772, 1810]. Earning [1528]. Earthquake [2103]. East [2103, 1255]. Eastern [337, 3065]. Easy [2088, 994, 1884, 1148]. Eat [1436, 1004, 1240, 1807]. Eating [1376]. Eats [1220]. Eclipse [2280]. Ecological [794, 1018, 2067, 2686, 1310, 638, 1013, 1532, 1066, 1184, 1229, 1211, 1484, 863, 596, 2702, 1375, 464, 629, 299, 1361, 2084, 538, 651, 849, 3163, 2806, 1297, 2040, 1859, 3112]. Ecologist [1945]. Ecologists [1149, 2162]. Ecology [1100, 2249, 1765, 2136, 1829, 859, 278, 2659, 999, 1011, 1024, 1039, 1055, 949, 1585, 351, 1417, 2114, 2738, 3265, 984, 2737, 2150, 457]. Economic [180, 1515, 2067, 1553, 1484, 291, 1794, 1644, 955, 447, 105, 2675, 398, 277, 2806, 2534, 2365]. Economics [1266, 457]. Economy [1475]. Ecosystem [801, 375, 2674, 2978, 1175, 1383, 1494, 2653, 2218, 1369, 2078, 2466, 189, 777, 1266, 2992, 1419, 808, 2509, 2570, 508, 2026, 1391, 676, 1418, 2244, 2176, 2198, 2583, 584, 1734, 902, 2399, 873, 2807, 451, 2794, 2475, 1702, 2484, 372, 550]. Ecosystem-Based [1175, 2218, 2078, 777, 1419, 808, 2570, 676, 2244, 2176, 1734, 2807, 2509, 873]. Ecosystems [1467, 1678, 2146, 2067, 540, 3066, 701, 120, 2455, 1531, 544, 3005, 547, 1083, 2892, 932, 761]. ed [1792, 1791]. Edge [943, 1099]. edited [1522, 1540]. Edition [1337, 3237, 2396, 2294, 2586]. Editor [1662, 1398, 2033]. Editorial [1710, 966, 982]. Editors [1568, 1604, 2568]. Education [1019, 1190, 1206, 1352, 1543, 1753, 1959, 2169, 2355, 2511, 3006, 3155, 2002, 1748, 1992, 44, 2656, 46, 1615, 491, 1108, 18, 2005, 2311, 731, 358, 393, 133, 45, 2641, 43, 2013, 2552, 1035, 403, 1333, 240, 47, 132, 643, 49, 2553, 505]. Educational [381, 199, 2007, 2311, 2808, 2011, 473, 2465]. Educators [2008]. Edward [1219]. Eel [516, 1565, 1665, 576, 1828, 2112]. Eels [1665, 2481]. Effect [2877, 3266, 2068, 301, 58, 695, 327]. Effective [568, 988, 79, 1838, 1461, 2521, 1995, 3252, 1679, 2812]. Effectively [1775]. Effectiveness [1910, 1909, 3064, 697, 812, 895, 63, 252, 1252, 2465]. Effects [375, 1678, 1597, 1809, 3188, 3028, 1484, 270, 122, 500, 813, 1470, 1281, 267, 1985, 1981, 268, 1361, 1599, 3110, 1979, 2895, 679, 266, 2452, 121, 2930, 3001, 2958, 124, 1855, 55, 1300, 1988, 510]. Efficiency [2750, 1085]. Efficient [1747]. Effort [1836, 1709, 2120, 1767, 1186, 1751, 194]. Efforts [2414, 1939, 1638, 1322, 168, 2454, 1300, 1698, 3111]. Egg [2527]. Eggs [2966, 3145]. Ego [1886]. Eight [1323, 2573]. Eighth [3237]. Eiler [1540]. Eiserman [1911]. Elasmobranchii [557]. Elbows [3287]. Electrofishing [3232, 2120, 3200, 2821, 510, 2685]. Electroimmobilization [2611]. Electronic [3202, 2362]. Electronics [2701]. Element [2176]. Elements [2400]. Elephant [785]. elevation [655]. Eleven [1068]. Elusive [2483, 830]. Elwha [372]. Embracing [2338, 2302, 3030]. Embrey [2886]. Embryonic [1721]. Emended [101]. Emergence [238, 3065]. Emergency [2066]. Emergent [2892]. Emerging [1567, 1900, 2910, 1681, 2748, 2689, 1416, 2193, 1793, 547, 2453]. Emmeline [2726]. Emotional [3062]. Emphases [531]. Emphasize [426]. Empirical [3064, 2854]. Employers [1334, 2008]. Employment [1206, 1748, 1108, 18]. Empower [2326]. Emulating [1198]. Enable [2747]. Encourages [1300]. End [2079]. Endangered [2902, 205, 756, 595, 369, 1630, 525, 447, 433, 162, 2036, 510, 1144, 1793, 926, 2916, 3018, 577, 218, 242, 738, 665, 3225, 659, 821, 2203, 3029]. Endeavor [897]. Endemic [2102, 1555, 738]. Endings [101]. Endocrine [1793]. Energy [805, 540, 2385]. Engage [2638]. Engaged [2337]. Engagement [1125, 2197, 1904, 2517, 2751]. Engaging [2535, 2737, 2147]. Engineering [2551]. England [73, 860, 2219]. Enhance [1296, 1308, 2209, 2210]. Enhanced [1518]. Enhancement [502, 1742, 363]. Enhancing [2207, 1635, 3220, 1859]. Enigma [101]. Enjoy [3026]. Enough [1492, 463, 641]. Ensuring [1644, 249]. Entrepreneurship [2641]. Entry [73]. Envelopment [1125]. Environment [3161, 1695, 2166, 473, 305]. Environmental [16, 118, 1076, 2120, 2712, 2311, 999, 1051, 1375, 245, 60, 1699, 507, 929, 1634, 2251, 7, 1191, 936]. Environmentally [1518]. Environments [805]. envisioning [1957]. Episodic [2412, 233]. Equality [3235]. Equator [1991]. Equity [3277, 2852]. Era [1369, 1035, 2676]. Eradication [1838]. Eric [2214]. Erie [321, 455, 587, 1280]. Erratum [2810]. Error [1371]. Escapement [586, 624, 873]. Escapements [444]. essay [2934]. Essential [1358, 540, 2400, 2176, 498, 3162]. Essig [1140, 2479]. Establish [586, 1839]. Established [1681]. Establishing [2412]. Establishment [1157, 764]. Estimate [2510, 3126, 947, 2362]. Estimates [1837]. Estimating [870, 574, 396, 813, 2151, 3087, 678]. Estimation [550, 706, 649]. Estuaries [85, 1838]. Estuarine [567, 144, 267, 872, 1985, 581]. Estuarine-Dependent [872]. Estuario [811]. Estuary [728, 914, 2164, 811]. Etheostoma [2758]. Ethic [1273]. Ethical [2574]. Ethics [2532, 1528, 2190, 457]. Ethnography [2535]. Eurasian [2857, 394]. Europe [771, 573]. European [1537, 2989]. Eustis [3106]. Evaluate [1125]. Evaluating [2218, 783, 2340, 2387, 3110, 762, 3201, 795, 28]. Evaluation [1533, 288, 586, 1371, 697, 894, 3150, 1124, 500, 925, 317, 829, 82, 1432, 1241, 349, 744]. Evaluations [3064]. Even [37, 2762]. Event [2734, 1100]. Events [906, 963, 1063, 959, 975, 990, 1008, 1016, 1036, 1052, 1078, 1091, 1105, 1118, 1135, 1152, 1166, 1181, 1193, 1207, 1222, 1234, 1247, 1261, 1276, 1290, 1304, 1316, 1328, 1339, 1356, 1364, 1379, 1395, 1409, 1425, 1442, 1459, 1473, 1488, 1507, 1526, 1546, 1559, 1580, 1593, 1608, 1624, 1651, 1671, 1687, 1724, 1740, 1756, 1770, 1785, 1814, 1830, 1847, 1863, 1879, 1896, 1913, 1927, 1944, 1961, 1974, 1998, 2028, 2044, 2060, 2074, 2095, 2253, 2285, 2110, 2303, 2317, 2331, 2342, 2353, 2367, 2379, 2403, 2416, 2429, 2444, 2457]. Events [2470, 2485, 2497, 2512, 2525, 2538, 2648, 2663, 2678, 2691, 2703, 2716, 2727, 2740, 2757, 2787, 2771, 2882, 2813, 2800, 2826, 2836, 2859, 2872, 2884, 2900, 2914, 2925, 2938, 2996, 3010, 3021, 3034, 3067, 2671, 2412]. Ever [3209]. Everyday [1809]. Everyone [1689, 1539]. Everything [1838, 1240, 1852]. Evidence [2891, 2890, 3064, 2134, 550, 730, 1694, 2854, 1775, 869, 605, 2495]. Evidence-Based [2134]. Evidenced [2165, 3213]. evides [2955]. Evolution [2536, 859, 1001, 2035, 3213, 2994]. Evolve [2003, 1820]. Evolving [1419, 2521]. Exactly [1359]. Exam [2950]. Examination [1826]. Examines [3064]. Examining [2082, 2006]. Example [1909, 1809, 331, 2494, 1097, 2880, 795]. Excavation [82]. Excellence [1954, 1744]. Except [1760]. Exciting [2906]. Exclusive [2675, 581]. Executive [1068]. Exemplified [872, 3053]. Exercise [172]. Exercises [2808, 1333]. Exhibit [2361]. Exhibitors [3033]. Exhibits [2913]. Existing [460]. Exotic [516, 698, 168, 743]. Expand [1297]. Expanded [1600]. Expanding [2736, 2260, 3111]. Expansion [146, 2778]. Expect [1989, 2376]. Expectations [1401, 1477, 3250, 198, 1635]. Expecting [1174]. Expeditions [435]. Expenditures [3029]. Expensive [1161]. Experience [2545, 717, 1798, 310, 2819, 1333, 1014, 2610, 1719]. Experienced [2162]. Experiences [2582, 2274, 3100, 2197]. Experiencing [3266, 2412]. Experiential [2136]. Experiment [2508, 175, 484, 123]. Experimental [851, 2923, 2572]. Experiments [115, 635, 1868]. Expert [909]. Expertise [1995, 1111, 2228]. Experts [791]. Explain [2584]. Explaining [330]. Explicit [1322, 300]. Exploitation [285, 94]. Explore [2068]. Exploring [1505, 2658, 881, 2351, 2733, 1804, 883]. Export [2106]. Exposed [2657]. Exposure [1771, 2300]. Expression [1771]. Extend [2395]. Extension [133, 78, 49]. Extent [723, 528]. External [384]. Extinction [1533, 638, 546, 581, 556, 1099, 1014]. Extinctions [243]. Extirpation [1485]. Extraction [1767, 1889, 1457]. Extractive [2180]. Extraordinary [979, 964]. Eye [2472, 893].

F [2025, 2132, 1162, 2786, 3199]. Face [2099, 1775, 2251]. Facebook [1496]. Faceoff [2431]. Facilitated [2509]. Facilitates [1149]. Facilitating [1074]. Facilities [339]. Fact [2530]. Factor [813]. Factors [2313, 866, 1751, 2573]. Facts [536]. Failure [2373]. Fajen [2025]. Fall [1375, 2591]. Falls [1617]. Fame [980]. Families [647]. Family [1216]. FAMS [1683]. FAO [1407, 1642]. Far [2149]. Farmed [2675]. Farms [1820]. Fate [1907]. Father [2642]. Fauci [2805]. Fauna [353]. Favor [2032]. Favorite [1837]. FDA [1448, 1589]. Feasibility [123]. Feature [761, 762]. Features [336]. February [2995, 2911, 3007, 1153, 1317, 1524, 1739, 1958, 2157, 2307, 2343, 2461, 2636, 2803, 2957, 3107, 3259]. Federal [1005, 1062, 124, 426, 637, 2339, 169, 2007, 155, 45, 925, 1428, 194]. Federally [848]. Federally-Declared [848]. Federation [1523]. Feedbacks [2040]. Feeds [340, 1775]. Felipe [957]. Fellow [2126]. Fellows [1825]. Fellowship [1087, 1257, 1282, 1451, 1452, 1661, 1875, 2055, 2252, 2428, 2587, 2753, 2912, 3056, 3211, 1937, 1776, 1176, 1450, 1660]. Fence [2195]. Fenske [1937, 1776, 1176, 1450, 2054]. Fester [1623]. Festival [2036]. Fetterolf [1556]. Few [2594, 1696, 2260]. Fewer [2012]. Fickle [3222]. FiCli [3222]. Fiction [369]. Field [288, 2723, 224, 1215, 3284, 2853, 396, 862, 2923, 830, 115, 1046, 1696, 2022, 2662]. Fielder [1732]. Fieldwork [2909]. Fight [2897]. Figures [2599]. Film [2328, 1287]. Filming [2655]. Filth [1216]. Final [1656, 1870, 2086]. Financing [1836]. Find [1160]. Finding [1810, 3262]. Finfish [1681, 369, 532, 294, 765]. Finfisheries [179]. Finfishermen [73]. Fingerlings [3206]. Finish [3002]. Fire [1273]. First [2639, 1406, 1603, 1832, 2575, 2602, 3203, 1435, 1323, 3126, 570, 2173, 2934, 215, 298, 2278, 2480]. Fiscal [2084]. Fischer [2185]. Fish [2217, 794, 909, 1028, 1243, 1483, 1577, 1854, 2729, 288, 1547, 1621, 199, 3004, 224, 568, 1873, 311, 707, 1359, 1551, 1819, 1917, 1900, 1931, 2101, 2259, 2177, 2227, 958, 2922, 626, 1871, 2010, 1382, 1448, 300, 954, 1499, 1538, 1371, 1637, 339, 609, 296, 1388, 2569, 1987, 50, 747, 3202, 2466, 717, 851, 2463, 2366, 25, 2107, 364, 1027, 3115, 1030, 128, 1058, 396, 1916, 2007, 2120, 1840, 1258, 761, 1586, 100, 2510, 2689, 2712, 182, 2612, 720, 1937, 698, 3200, 2906, 2744, 291, 460, 3222, 122, 246, 247, 836, 596, 2412, 723, 331, 2866, 1761]. Fish [1838, 1733, 1872, 813, 2820, 2166, 233, 2151, 2349, 2908, 63, 1186, 1444, 3, 139, 1431, 1439, 1654, 1973, 3280, 318, 888, 3087, 2423, 301, 245, 89, 2871, 1894, 2631, 1714, 2019, 780, 3065, 1516, 866, 2350, 955, 1187, 447, 1387, 1302, 1981, 2700, 1970, 816, 141, 154, 857, 206, 1554, 1226, 829, 1097, 1174, 703, 29, 1361, 1954, 2036, 920, 435, 1990, 497, 746, 1922, 2688, 1309, 581, 556, 3075, 514, 510, 2266, 2733, 484, 3267, 97, 2809, 2895, 538, 2066, 2780, 929, 2972, 1322, 678, 875, 42, 2032, 593, 762, 404, 3017, 290, 475, 536, 3005]. Fish [2100, 2452, 782, 2482, 2022, 498, 340, 776, 789, 2932, 1837, 1820, 1763, 2261, 2173, 191, 615, 51, 168, 2093, 2833, 1735, 3029, 3018, 2397, 90, 1855, 1833, 3252, 2300, 363, 704, 2436, 27, 7, 1332, 55, 838, 3041, 2102, 1132, 1300, 1325, 1640, 1589, 2434, 2365, 2735, 2572, 2620, 3044, 2973, 2734, 577, 1920, 2506, 2231, 1069, 691, 2121, 839, 867, 1026, 2360, 932, 3149, 3225, 786, 744, 743, 1827, 2729, 2264, 2933, 2192, 1482]. fish-gill [743]. Fish-related [857]. Fish-Seed [2735]. fish-stocking [744]. FishBase [1519]. Fisher [2842]. Fisheries [2179, 220, 381, 939, 2654, 2796, 1125, 2750, 907, 906, 938, 899, 898, 961, 976, 1120, 1063, 959, 975, 990, 1008, 1016, 1036, 1052, 1078, 1091, 1105, 1118, 1059, 1021, 1116, 1100, 970, 997, 1048, 1103, 1128, 1129, 1283, 1284, 1190, 1244, 1135, 1152, 1166, 1181, 1193, 1207, 1222, 1234, 1247, 1261, 1276, 1290, 1165, 1209, 1246, 1126, 1163, 1134, 1179, 1221, 1274, 1268, 1380, 1454, 1453, 1304, 1316, 1328, 1339, 1356, 1364, 1379, 1395, 1409, 1425, 1442, 1459, 1303, 1355, 1408, 1407, 1315, 1363, 1423, 1664]. Fisheries [1603, 1619, 1629, 1656, 1663, 1473, 1488, 1507, 1526, 1546, 1559, 1580, 1593, 1608, 1624, 1651, 1671, 1572, 1602, 1504, 1515, 1534, 1472, 1524, 1592, 1622, 1487, 1544, 1607, 1650, 1670, 1813, 1874, 1832, 1845, 1860, 1870, 1876, 1687, 1724, 1740, 1756, 1770, 1785, 1814, 1830, 1847, 1863, 1879, 1825, 1686, 1739, 1831, 1878, 1802, 1723, 1755, 1784, 1846, 2037, 2051, 2070, 2090, 2072, 2086, 2088, 1896, 1913, 1927, 1944, 1961, 1974, 1998, 2028, 2044, 2060, 2074, 2095, 2087, 1912, 1958, 2059, 1895, 1943, 1997, 2073]. Fisheries [2169, 2200, 2239, 2282, 2126, 2109, 2283, 2110, 2157, 2223, 2286, 2255, 2125, 2187, 2254, 2442, 2446, 2443, 2343, 2390, 2318, 2368, 2511, 2613, 2615, 2456, 2616, 2479, 2754, 2783, 2770, 2624, 2952, 3091, 3239, 3242, 3141, 3116, 3128, 3253, 2877, 2879, 1172, 1171, 1533, 367, 1606, 1133, 1523, 426, 2545, 2723, 2763, 2537, 202, 1658, 2023, 1175, 1286, 23, 707, 878, 416, 1749, 2002, 2146, 2415, 958, 2721, 2658, 1319, 682, 989, 884, 192, 400, 954, 1553, 637, 3077, 80, 1748, 1992, 2207, 2508, 2686, 2961, 2699, 956, 44, 3040]. Fisheries [1540, 1310, 623, 697, 1500, 17, 2656, 1390, 287, 2078, 387, 2015, 2466, 2038, 783, 1110, 2134, 2463, 2823, 3063, 2959, 2987, 494, 177, 3016, 2328, 2250, 1798, 894, 2329, 128, 46, 988, 1532, 366, 2018, 2582, 1430, 2779, 1419, 1818, 1835, 1964, 2001, 1899, 2507, 881, 361, 2855, 634, 491, 511, 657, 1108, 1066, 1184, 1229, 1211, 1196, 334, 430, 265, 808, 1265, 2295, 722, 937, 314, 2689, 1313, 2712, 439, 3074, 270, 1638, 2518, 910, 2005, 2311, 2808, 79, 1684, 269, 2944, 3003, 2275, 1416, 731, 358, 2661, 2858, 2363, 1826, 1695, 2016, 393, 2570, 3277]. Fisheries [1953, 2990, 208, 1070, 2598, 3062, 133, 2494, 45, 895, 922, 1176, 310, 1891, 2550, 2549, 278, 2313, 999, 1011, 1024, 1039, 1055, 949, 1528, 1510, 3224, 2867, 987, 454, 2752, 232, 815, 1882, 2012, 2641, 3220, 3261, 3276, 411, 2017, 1349, 3087, 1287, 2846, 1429, 2006, 222, 209, 175, 267, 253, 38, 3100, 2871, 2835, 2437, 2574, 1570, 1478, 1495, 2019, 3129, 780, 1334, 2660, 43, 1644, 528, 676, 772, 866, 2535, 143, 2013, 2923, 3175, 711, 2819, 2893, 2039, 629, 1418, 2244, 324, 2519, 3151, 1050, 1302, 2122, 488, 816, 830, 432, 159, 2176, 1468, 410]. Fisheries [2907, 1554, 378, 1042, 1601, 2881, 2004, 2209, 2336, 2547, 2242, 442, 3008, 2041, 1035, 935, 2103, 163, 1642, 1531, 1904, 3086, 532, 1742, 1599, 284, 2946, 298, 274, 403, 2021, 2842, 2765, 546, 3075, 2894, 1680, 129, 406, 1240, 1141, 150, 240, 47, 3072, 2326, 1471, 2811, 1465, 1098, 1074, 1734, 644, 1983, 2895, 2210, 654, 873, 3164, 3184, 2099, 621, 2299, 984, 2737, 1893, 252, 1697, 1005, 1062, 2467, 2962, 2195, 2685, 1766, 3109, 1968, 463, 531, 590, 2919, 1956, 2520, 2807, 1568, 2150, 2854, 389, 2932, 3053, 1852, 1793, 377, 451, 1199, 1146, 249]. Fisheries [2079, 662, 2833, 2325, 1552, 347, 2182, 2632, 1399, 176, 549, 1833, 2595, 2865, 564, 2950, 1843, 2008, 1966, 2014, 429, 2856, 216, 2768, 2504, 856, 1384, 2909, 2276, 2751, 1300, 1589, 2462, 2868, 3019, 2424, 2832, 343, 1297, 136, 768, 1370, 2040, 583, 1844, 3212, 2724, 1321, 448, 668, 218, 457, 640, 2123, 2020, 1988, 1007, 2409, 2385, 2609, 2451, 3153, 1143, 2178, 1111, 505, 3097, 3124, 3137, 3172, 3197, 3249, 2869, 492, 2812, 452, 605, 1618, 1431, 1766, 61, 1015, 904, 966, 982, 960, 1130, 1292, 1301, 1456, 1669, 1877, 2092, 2287, 2290, 2307, 2321]. Fisheries [2335, 2346, 2358, 2371, 2383, 2393, 2407, 2420, 2433, 2449, 2461, 2474, 2489, 2500, 2503, 2529, 2542, 2564, 2580, 2593, 2607, 2626, 2636, 2652, 2667, 2682, 2695, 2707, 2720, 2731, 2746, 2760, 2775, 2791, 2803, 2816, 2830, 2840, 2851, 2863, 2876, 2888, 2904, 2918, 2929, 2942, 2957, 2967, 2984, 2999, 3013, 3024, 3037, 3049, 3060, 3070, 3083, 3095, 3107, 3121, 3135, 3146, 3159, 3169, 3180, 3194, 3207, 3218, 3233, 3246, 3259, 3274, 3183, 985, 2567, 2669, 1887, 2396, 3236]. Fisheries-Achieving [1098]. fisheries.org [1619]. Fisherman [155, 2595]. Fishermen [887, 291, 860, 206, 947, 1111]. Fisheroes [2673]. Fishers [1808, 1982, 3278, 891]. Fisherwomen [2811]. Fishery [180, 3059, 1678, 1353, 223, 2978, 323, 1494, 286, 554, 449, 777, 1752, 1808, 394, 571, 40, 198, 606, 2509, 852, 611, 2866, 421, 1957, 2026, 1431, 203, 694, 2085, 2521, 282, 160, 65, 528, 2519, 359, 141, 105, 260, 2427, 1646, 1557, 2584, 2948, 295, 544, 589, 695, 3072, 767, 1432, 183, 57, 593, 2299, 132, 613, 2960, 313, 1760, 2764, 3174, 398, 809, 2263, 49, 547, 1598, 684, 2797, 92, 1415, 2596, 8, 1241, 327, 467, 1679, 67, 880, 1683, 2312, 67]. Fishery-Dependent [2596]. Fishery-Induced [1415]. Fishes [2601, 2135, 3284, 2910, 1993, 2183, 109, 558, 876, 1013, 1586, 3028, 337, 2057, 2510, 3152, 1719, 730, 3112, 1403, 815, 86, 854, 1600, 872, 1908, 149, 2340, 2104, 2122, 488, 774, 317, 268, 162, 243, 2056, 2284, 2642, 2713, 546, 1240, 1979, 3043, 1057, 1984, 3042, 679, 1014, 1614, 2185, 266, 2611, 2235, 121, 2362, 2933, 1144, 1938, 168, 811, 1682, 1725, 1848, 1864, 1701, 1771, 2662, 441, 2102, 2655, 1564, 2106, 353, 579, 1555, 1980, 218, 242, 520, 738, 2123, 2726, 1337, 3237, 2586]. FishGen.net [2377]. Fishician [2950]. Fishing [1792, 2693, 2415, 1806, 1805, 3148, 647, 806, 2748, 1389, 627, 1532, 2779, 551, 2710, 618, 2401, 846, 1616, 863, 527, 2494, 1562, 1089, 612, 3278, 2481, 660, 1437, 1464, 727, 321, 455, 587, 1750, 993, 1280, 2352, 841, 1751, 3127, 1780, 2809, 2591, 1856, 598, 1893, 495, 692, 2465, 1820, 276, 277, 809, 1566, 798, 891, 1131, 1300, 2450, 2709, 2112, 28, 3262]. FishSmart [935]. FISHstory [3077]. Fishy [1757]. Fit [1905, 583]. Five [1373, 2103, 355, 2630]. Five-tiered [355]. Flannelmouth [2375]. FLAs [3271]. Flathead [189]. Flexible [1773, 2520]. Flight [3258]. Flint [1731]. Flood [1484, 605]. Floodplain [926]. Floodplains [2484]. Florida [2037, 2051, 168, 2601, 1125, 3074, 3073, 2631, 65, 2387, 498, 51, 168, 1598, 66]. Floristic [1555]. Flow [288, 1978, 1965, 2843, 250, 239, 2767, 2972, 845, 230, 1552, 7, 936, 608]. Flows [1216, 497]. Flux [2096]. Fly [2693]. Flying [2569]. FMS [1954]. Focus [1618, 1143]. Focused [2975, 3281]. Focuses [1325]. Following [2596]. Food [1467, 1806, 1805, 2779, 179, 3, 1672, 2149, 2733, 692, 776, 869, 932]. Footsteps [2054]. Forage [1584, 2032, 3005, 1890]. Foray [1434]. Forecasting [758, 270]. Forest [192, 331, 175, 2780, 475, 2398, 2397, 2396, 1332, 2394, 2405, 2849, 334, 2400, 175, 253, 2395]. Forests [734, 2401]. Forgetting [1296]. Forging [1050]. Forgiveness [1738]. Forgotten [1885, 2612, 1677]. Form [1351, 2748, 1903]. Former [2165]. Forming [1376, 1414, 522]. Formosa [595]. Fortunately [1747]. Forum [2946]. Forward [1375, 903, 2396]. Fossil [1889, 1725]. Fostering [1075]. Foundation [2373, 240, 35]. Foundational [1493]. Foundations [1568]. Founding [2630]. Fountain [1737]. Four [2088, 1302, 3153]. Fourth [1113, 1285, 1455, 2261]. Fractured [2066]. Fracturing [1133, 1295]. Fragmentation [1057]. Fragmented [1923]. Framework [2414, 718, 2386, 2207, 3226, 417, 1537, 1973, 1766, 1069, 2747]. Frances [1542, 1321]. Francisco [822, 811, 811]. Frank [1620]. Franks [1801]. Fraser [1343, 689, 851, 2766, 2302]. Fred [1911]. Frederic [203]. Free [1595, 1680]. Fresh [2601, 1937, 168]. Freshwater [3168, 523, 85, 1871, 2135, 2386, 2508, 3064, 17, 109, 2245, 1013, 2250, 3028, 2057, 2510, 550, 1695, 1439, 854, 1360, 3175, 351, 573, 1187, 488, 816, 285, 2337, 3054, 1099, 1984, 1614, 452, 3017, 266, 531, 2235, 2482, 2854, 2662, 441, 2909, 2424, 1564, 1370, 353, 579, 345, 761, 2569]. Freshwater-Inflow-to-Estuaries [85]. Freshwaters [660, 893]. Friedrich [2799]. Friendly [2767, 2855]. Frogfishes [2738]. Front [2480]. Frontier [2569, 2910]. Frontiers [1100, 999, 1011, 1024, 1039, 1055, 949, 984]. Frontloading [1191]. Fry [2742]. Fs [1302]. Fuca [3284]. Fuel [1889, 1085]. Full [1, 2, 4, 6, 9, 10, 11, 12, 13, 15, 19, 21, 22, 24, 26, 30, 32, 34, 36, 41, 48, 52, 53, 54, 56, 59, 62, 64, 70, 72, 74, 76, 81, 83, 84, 87, 91, 93, 95, 96, 99, 102, 106, 108, 110, 112, 113, 116, 119, 127, 130, 134, 138, 142, 145, 148, 151, 153, 156, 158, 161, 164, 171, 174, 178, 181, 184, 186, 190, 193, 197, 201, 204, 207, 210, 214, 217, 221, 225, 228, 231, 234, 237, 241]. Full [244, 248, 251, 254, 258, 263, 271, 272, 275, 280, 283, 289, 292, 297, 302, 309, 312, 316, 319, 322, 325, 328, 332, 335, 338, 341, 344, 346, 348, 350, 352, 354, 357, 360, 362, 365, 368, 371, 374, 376, 379, 382, 385, 388, 392, 395, 397, 399, 402, 405, 407, 408, 412, 415, 418, 420, 422, 425, 428, 431, 434, 437, 440, 443, 446, 450, 453, 456, 458, 459, 462, 465, 468, 469, 471, 472, 474, 477, 480, 483, 485, 487, 489, 493]. Full [496, 499, 501, 503, 506, 509, 512, 515, 518, 521, 524, 526, 530, 534, 537, 539, 542, 545, 548, 552, 555, 560, 563, 566, 569, 572, 575, 578, 580, 582, 585, 588, 591, 594, 597, 600, 603, 604, 607, 610, 614, 617, 620, 622, 625, 628, 630, 633, 636, 639, 642, 645, 648, 650, 653, 656, 658, 661, 663, 666, 669, 672, 675, 677, 680, 683, 685, 687, 690, 693, 696, 699, 702, 705, 708, 710, 713, 716, 719, 721, 726, 729, 733, 736]. Full [739, 742, 745, 748, 751, 754, 757, 760, 763, 766, 770, 773, 775, 778, 781, 784, 788, 790, 793, 796, 799, 803, 807, 810, 814, 817, 820, 823, 825, 828, 831, 834, 837, 840, 844, 847, 850, 853, 855, 858, 861, 865, 868, 871, 874, 877, 882, 885, 889, 892, 896, 942, 945, 912, 915, 918, 921, 924, 927, 931, 934, 948, 967, 983, 998, 1010, 1023, 1038, 1054, 1065, 1080, 1093, 1107, 1121, 1137, 1154, 1168, 1183, 1195, 1210, 1224, 1236, 1249, 1263]. Full [1278, 1293, 1306, 1318, 1329, 1340, 1357, 1366, 1381, 1396, 1410, 1426, 1443, 1460, 1474, 1491, 1509, 1527, 1549, 1561, 1582, 1594, 1610, 1625, 1653, 1673, 1688, 1707, 1726, 1743, 1758, 1772, 1787, 1817, 1834, 1849, 1865, 1881, 1898, 1915, 1929, 1947, 1963, 1976, 2000, 2030, 2046, 2063, 2076, 2097, 2113, 2128, 2144, 2159, 2174, 2189, 2204, 2225, 2241, 2257, 2271]. Function [2068]. Functional [238, 1736, 2173]. Functioning [1225]. Fund [2245]. Funded [92, 175]. Funding [220, 722, 2974, 1428, 531, 2932]. Fundy [177]. furcatus [2789]. Further [2637]. Future [1422, 2722, 1115, 23, 2415, 956, 1310, 2015, 2038, 1076, 2232, 2582, 2001, 419, 334, 314, 270, 2050, 910, 1684, 3277, 293, 1470, 1281, 2752, 86, 3278, 2006, 2835, 816, 2258, 1035, 1531, 1590, 152, 1074, 1949, 2099, 2960, 2251, 2632, 2595, 1952, 216, 2751, 1132, 3281]. Future-Focused [3281].

G [1571, 2798]. Gage [1517]. Gains [723]. Game [1058, 3112, 3075, 168, 2476]. Gammon [3007]. Ganga [2834]. Gap [2698, 2629, 2646, 2454, 2899]. Gaps [2748]. Gar [2224]. Garbage [1550]. Gardens [2659]. Gars [313]. Gary [3267, 1350, 2316, 2154]. Gas [131, 1457]. Gasoline [1835]. Gastropods [1360]. Gather [3098, 2127]. Gatherings [907]. Geechee [1616]. Gender [1826, 3027, 1203, 2505]. Gene [1771]. General [2820]. Generated [2534]. Generating [2192]. Generation [3278, 2017, 2013, 2004, 1146, 2014, 1143, 1349]. Generations [2017]. Generic [2857, 2989]. Generous [1575]. Genes [1633]. Genetic [417, 758, 673, 1404, 933, 2377, 2312, 1938, 343, 262, 2123]. Genetics [358, 1349, 1375, 222, 162, 1868]. Genomics [851]. Genus [2104, 3009]. Geochemical [2795]. Geographic [1430, 491, 454, 1680]. Geological [866]. Geomorphology [42]. Georgia [3284, 2921, 862, 1204]. Georgian [804]. Gerald [2265]. Get [1595, 3272, 68, 3288, 2077, 1950, 3026, 3257, 2533, 1164]. Getting [1330, 1400, 2226]. Giant [2639]. Gibson [2138]. Gift [1880]. Gifts [1288]. Gila [3158, 2230, 2229]. Gilchrist [2344]. gill [743]. Gilt [2955]. GIS [1535]. Give [2199, 37]. Given [641]. Gives [1482]. Giving [1880]. Glance [1574, 1781]. Glass [373]. Glimmer [232]. Global [1572, 1602, 1172, 1842, 1499, 1058, 269, 479, 2702, 2835, 2105, 1418, 2700, 268, 2711, 2676, 264, 3111, 1618, 1504, 1515, 1534]. Globally [1658, 1307, 1427, 584, 2408]. Globe [2463]. Go [2845, 2273, 1445, 2637, 2308, 1388, 2031, 2919, 1848, 2950, 2008]. Goal [75]. Goals [2796, 23, 1501, 821, 860, 3196, 97, 1098, 873, 1500]. Goes [1820]. Going [212]. Golden [2834, 944, 2133, 3045]. Goliath [2631]. Good [2015, 2233, 2820, 2521, 2675, 1160, 2034]. Goodbye [2933]. gorbuscha [78]. Gordon [3199]. Gorge [508]. Governance [1602, 1110, 2299, 2812]. Government [1738, 1213]. Governmental [836]. Grace [1667]. Grad [2845]. Graduate [1810, 1160, 1858, 1199]. Graduates [2016]. Grand [2122, 2680, 3239, 3103, 3123, 3235]. Grande [2844]. Grant [1314, 1633]. Graphical [2068]. Grass [196, 764, 5]. Grassroots [2323]. Gravel [507]. Gray [259]. Grayling [2149]. Grazing [375, 1583, 536]. Great [2459, 2089, 1477, 146, 1500, 1798, 3115, 1894, 1738, 78, 141, 268, 1967, 1696, 2467, 1018, 2569, 1266, 2715, 2509, 1776, 421, 2908, 1892, 1212, 3278, 1308, 727, 1750, 2069, 2103, 2108, 2118, 1057, 77]. Greater [1867, 334]. Greatest [2585]. Greatness [1714]. Green [141, 1148, 1273, 864]. greet [1127]. Greeters [3076]. Grossman [3267]. Groundfish [326, 848, 611, 2219]. Groundwater [211, 2697]. Group [1098, 1325]. Grouper [558, 2631, 613, 2597]. Grouse [1867]. Growing [2101]. Grown [2878]. Growth [300, 1884, 20, 955, 433, 31, 1642, 484, 2747]. Grunion [3076].

Guest [989]. Guidance [1640]. Guide [2601, 904, 960, 1130, 1301, 1435, 3284, 1540, 1004, 2148, 2019, 1718, 2399, 2662, 2586, 2599]. Guidelines [904, 2574, 285, 1213, 218, 1600]. Guiding [1679]. Guild [367, 1057]. Guilds [2494]. Gulf [1678, 1801, 1089, 1694, 1125, 686, 2447, 750, 956, 688, 1041, 1613, 1794, 2521, 1042, 692, 1702]. Gullah [1616]. Gullah/Geechee [1616]. Gut [2689]. Gutting [1431]. Guy [2213]. Guyana [1822]. Gyotaku [2973, 2284].

H [991, 1540, 2314, 2936, 2315, 2221]. Habit [1672]. Habitat [939, 2870, 923, 602, 522, 157, 1401, 1342, 1359, 1931, 2101, 3251, 3223, 2078, 761, 1537, 720, 527, 1124, 723, 336, 200, 2908, 1431, 1973, 919, 925, 528, 342, 261, 2176, 920, 2767, 279, 97, 57, 725, 762, 266, 1722, 2100, 498, 776, 2150, 1837, 2093, 549, 2247, 704, 429, 632, 466, 1069, 1295, 123, 2844]. Habitat-Based [704]. Habitat-forming [522]. Habitats [523, 1655, 331, 149, 1187, 940, 1775, 1938]. Hagerty [2793]. Hagfishery [256]. Hairy [2188]. Halamid(R) [1589]. Half [1742]. Half-Shell [1742]. Halibut [2327, 2865]. HALK [3279]. Hall [1954, 980]. HaMAR [1325]. Hand [2919]. Handle [2412]. Handling [2866]. Hands [2562, 1798]. Hands-On [1798]. Hansen [1345]. Happens [1774]. Hard [1798, 1141, 1614]. Hard-Part [1614]. Harder [2531, 2047]. Harmful [2657, 1158]. Harness [3072]. Harnessing [2275]. Harsh [2166]. Hartman [3199]. Haruspication [1873]. Harvest [1621, 3064, 2779, 2701, 411, 2437, 1639]. Harvests [3066]. Harvey [2213]. Hash [1652]. Hatched [2966]. Hatcheries [1641, 199, 712, 709, 293, 1644, 29, 1098, 902, 1639, 1132, 1325, 2365, 517, 1325]. Hatchery [1243, 1541, 1577, 2921, 652, 3223, 2878, 1721, 737, 570, 849, 2736, 3213, 363, 2573, 1325, 1640, 2735, 1854, 2262, 1840, 1098]. Hatchery-Bashing [2735]. Hatchery-Origin [1325, 1640]. Hatchery-Reared [1721]. Hatching [2369]. Haul [2234]. Haven [1986]. Having [2608]. Hawaii [3079]. Hawk [2729]. Haystack [1344]. He/Him [3257]. Head [2173, 2950]. Headliners [1142, 1156, 1170, 1185, 1197, 1238, 1251, 1212, 1123]. Headlines [1434]. Headwater [2466, 701, 1951]. Health [951, 1637, 1987, 497, 475, 1199, 1192, 1077, 1260, 1150, 1338, 1394, 1458, 1579, 1815, 1926, 1972, 2043, 2111, 2172, 2238, 2304, 2354, 2404]. Healthy [423, 1083, 1213]. Hearing [1109]. Heat [1682, 2991]. Heaven [1951]. Heighten [1728]. Helicopter [977]. Help [3004, 2347, 1215, 2183, 366, 1020, 331, 3079, 2533, 2899, 3002]. Helping [794]. Hemdal [1763]. Hemisphere [1657]. Henshall [2749]. her [3288]. Here [2273, 2077, 2031, 2008]. Hermaphoditic [1856]. Hero [1345, 2410]. Heroine [2054]. Herring [3065]. Hickok [3227]. Hidden [1808, 1516, 3240]. Hide [1554]. Hiding [2688]. Hierarchical [336, 1069]. Hierarchies [1552]. Higgins [2472]. High [2723, 689, 2272, 2751, 1239, 2655, 655]. High-elevation [655]. High-Engagement [2751]. High-Priority [2272].

Highland [2960]. Highlighting [1744]. Highlights [899, 1059, 2417, 2430, 2445, 2458, 2471, 2486, 2498, 2513, 2526, 2539, 2561, 2577, 2590, 2604, 2618, 2623, 2649, 2664, 2679, 2692, 2704, 2717, 2728, 2741, 2756, 2788, 2772, 2814, 2801, 2827, 2837, 2848, 2860, 2873, 2885, 2901, 2915, 2926, 2939, 2954, 2965, 2982, 2997, 3011, 3022, 3035, 3047, 3058, 3068, 3081, 3093, 3105, 3119, 3133, 3144, 3157, 3167, 3178, 3192, 3205, 3231, 3244, 3190, 3256, 3270, 3286, 2280, 1102]. Highly [2058]. Hiking [1741]. Hilborn [1539, 1539]. Hill [1749, 1399, 2864]. Hilling [2662]. Hills [3213]. Him [3257]. Hinder [2711]. Hindering [1677]. Hires [937]. Hiring [1413]. Historic [2696, 2339, 550]. Historical [2165, 2057, 730, 65, 1694, 2688, 797, 2397, 2724]. History [2624, 819, 2696, 2658, 887, 1700, 2699, 2038, 1076, 270, 179, 772, 2535, 1497, 3151, 532, 406, 1519, 2194, 1893, 2452, 229, 3163, 2586, 2533]. Hitch [1741]. Hitch-Hiking [1741]. Hobikiami [88]. Hogan [2620]. Hogarth [3114]. Hogfish [2459]. Holarctic [1829]. Hold [1161]. Holds [1217, 1109, 1205, 1204]. Hole [870, 2552, 327]. Holiday [1288]. Holidays [2609]. Home [2323]. Homer [1073]. Homogenization [615]. Homogenizing [404]. Hong [2361]. Honorable [1886, 1216]. Honorary [1271]. Honoring [2054]. Hook [2425, 2866, 968, 949, 984, 2865]. Hope [2683, 232, 2905, 2864]. Hopes [1655]. Hopkins [2662]. Horizon [1089, 1042, 1702]. Horseshoe [543]. hosted [3214]. Hosting [2934]. Hosts [1749]. Hot [2232]. Hotdogs [1376]. Hotel [994]. House [1109]. HSRG [1641]. http [1619]. Hughes [1548]. Human [2891, 2890, 783, 1808, 955, 2733, 549, 448]. Human-as [783]. Human-Caused [2891, 2890]. Humanity [2361, 136]. Humankind [1312]. Humans [2674]. Humpback [1741]. Hunt [2220, 1420]. Huron [1751]. Hurricane [1218, 832, 2596]. Hutton [2022, 2087, 2545, 2723, 2763, 2937, 3098, 3266, 2868, 1007]. Hydraulic [1133, 2247, 1295]. Hydro [1499]. Hydroelectric [609]. Hydroelectricity [2880]. Hydrokinetic [805]. Hydrology [832]. HydroNet [1083]. Hydropower [339, 730, 1518, 252, 1083]. Hydrosystem [333]. Hyperbole [2778]. Hypotheses [1806]. Hypoxia [1531].

Ice [954, 2701, 2710, 2352, 2008]. Iceland [229]. IchthyMaps [2057]. Ichthyologists [3229]. Ichthyology [618]. Ichthyopedia [3229]. Icon [2971]. Ictalurus [2789]. Idaho [2687, 2532, 2464, 2871, 3008, 1967, 3210, 1617, 1217, 2971, 273, 42, 977]. Idea [2049, 1696, 2934]. Ideas [618, 2934]. idella [5]. Identification [2601, 1027, 1404, 1309, 2586]. Identify [1836, 533, 2795]. Identifying [2425, 2585, 1982, 774, 1793]. Identities [3100]. If [2653, 2016, 1837, 1820, 3042]. Ignore [2675]. Ignoring [2067]. II [1062]. IKONOS [804]. Illegal [888]. Illinois [1226, 1415]. Illustrated [3284, 3267, 1893]. Illustration [2802]. Illustrations [182]. Illustrator [2684]. Imagery [804]. Images [1767, 893]. Imagine [1729]. Imaging [3110].

Immediate [989, 1096]. Immediate-Release [989, 1096]. Impact [939, 180, 1728, 177, 1794, 1982, 1267, 1042, 211, 398, 2934, 94, 2909, 3149]. Impact-a-thon [2934]. Impacted [2766]. Impacts [718, 1806, 805, 2218, 1013, 2328, 846, 1794, 3112, 245, 752, 141, 1557, 1984, 1856, 940, 832, 1760, 1747, 216, 1295, 1980, 1702, 2395, 2385, 3225]. Impairment [3187]. Impedance [1872]. Impede [1734]. Impending [792]. Imperative [1368]. Imperiled [854, 2426, 353]. Implant [3202, 717]. Implement [2736]. Implementation [1027, 826, 933, 191, 2571, 349]. Implementing [1766, 2934]. Implications [196, 23, 1371, 1110, 494, 1517, 2507, 2712, 269, 576, 2058, 476, 1361, 1035, 298, 451, 1598, 2797]. Importance [367, 157, 1700, 461, 638, 3171, 1517, 40, 182, 291, 1267, 206, 3184]. Important [963, 1765, 1884, 1697, 1760, 1837, 1370]. Impounded [3101]. Impoundment [3110]. Impoundments [71, 135, 744]. Imprinting [2910, 1721]. Improve [609, 2701, 1068, 2906, 2275, 2349, 2641, 391, 484, 1938, 1920]. Improved [2671, 3028, 826, 2520, 2219, 883, 1859]. Improves [2750]. Improving [922, 1762, 1145, 1519, 252, 2102]. Impspired [3267]. IMTA [1681]. In-Person [2905]. Inability [533]. Inadvertent [3223]. Inappropriately [720]. Inaugural [1720]. Inc. [2179]. Incivility [2970]. include [1337]. includes [2479]. Including [1795, 1681]. Inclusion [2879, 2543, 2931, 3277, 2209, 2852, 2210]. Inclusive [3040, 2547, 2920]. Inconnu [1906]. Incorporate [2136]. Incorporates [1995]. Incorporating [2078, 1633]. Incorporation [869]. Increase [624, 3264, 2714, 460, 955, 1904, 2452]. Increased [2067, 1157, 818]. Increasing [1675, 1252, 2228]. Increasingly [447, 1647]. Incredible [3279]. Incremental [288, 239]. Independent [791]. Index [1120, 1877, 2092, 2287, 219, 274, 1793, 824, 880]. India [2834, 1228]. Indicator [1736]. Indicators [602, 117, 584]. Individual [813, 2930]. Individual-Based [813]. Indonesia [2249]. Induced [58, 1415]. Inductees [1720]. Industrial [1597]. Industries [3278]. Industry [894, 3150, 611, 727, 2974, 321, 455, 587, 1750, 1280, 14, 1633, 3111]. Inexpensive [2655]. Infections [3126]. Infectious [1631, 1403, 1632]. Infer [2832]. Inference [2329]. Inferences [922]. InFish [2700]. Inflow [85]. Influence [3200, 895, 2493]. Influences [1554, 466, 2123]. Influencing [866, 1751, 2573]. Influx [476]. Inform [2217, 833, 988, 821, 2895]. Information [2289, 2306, 2320, 2334, 2345, 2357, 2370, 2382, 2392, 2406, 2419, 2432, 2448, 2460, 2473, 2488, 2502, 2515, 2528, 2541, 2563, 2579, 2592, 2606, 2625, 2635, 2651, 2666, 2681, 2694, 2706, 2719, 2730, 2745, 2759, 2761, 2774, 2790, 2804, 2817, 2829, 2839, 2850, 2862, 2875, 2887, 2903, 2917, 2928, 2941, 2956, 2968, 2985, 3000, 3014, 3025, 3038, 3050, 3061, 3071, 3084, 3096, 3108, 3122, 3136, 3147, 3160, 3170, 3181, 3195, 3208, 3219, 3234, 3247, 3260, 3275, 492, 396, 1430, 1964, 491, 2275, 852, 1011, 454, 2019, 255, 1254, 857]. Information [1309, 1680, 2100, 1552, 1859, 616]. Information-based [396].

Informed [2570, 1531]. Informing [2570, 2427]. Infrastructure [2259, 1484]. Infusing [731]. Ingestion [1855]. Inhibit [1761]. Initial [1884, 2313]. Initiative [3278, 546]. Initiatives [1362, 2864]. Injuries [704]. Inland [2796, 2388, 1572, 1602, 1618, 1504, 1515, 1534, 626, 2961, 1993, 881, 734, 1265, 722, 1982, 2819, 1981, 2122, 2700, 2924, 2041, 2765, 129, 1983, 1979, 2150, 2932, 809, 2833, 2595, 2950, 1966, 1980, 2123, 1988]. Inner [3005]. Innovations [647, 2671, 137, 2736]. Innovative [2545, 1638, 2005, 935, 1995]. Input [1591, 2198]. Inquiry [1297]. Ins [1776]. Insight [2313]. Insightful [3016]. Insights [2855, 1429, 2013, 3175, 3198, 1746]. Inspirational [2048]. Inspiring [2723]. Institute [2245]. Institutional [2295, 640]. Institutions [1075]. Instream [288, 1965, 2843, 250, 239, 507, 2767, 845, 230, 7, 608]. Instructor [1145]. Instrumenting [2923]. Insurance [1199]. Intake [1597]. Integrate [2069]. Integrated [1369, 1681, 189, 2351, 1344, 1788]. Integrating [2537, 2992, 673, 1211, 421, 454, 2641, 2907, 662]. Integrative [2207]. Integrity [1386, 304, 89, 219, 2193, 615, 466]. Intensively [1910, 3250, 1909]. Inter [299]. Inter-basin [299]. Interacting [806]. Interactions [387, 3263, 528, 651, 849, 2228]. Interactive [1215]. Interbasin [436]. Interdependent [528]. Interdisciplinary [2002, 851, 40, 3174]. Interest [1899]. Interface [1312]. Interim [659]. Interior [1485, 925]. Interjurisdictional [2495]. Intern [2065, 2595]. Internal [3212]. International [1618, 2126, 2278, 2135, 2038, 3115, 1829, 1169, 2009, 3276, 1417, 2098, 1824, 2629, 2646, 2117, 2480, 1820, 2261, 165, 1988, 1286, 2327, 1590, 98]. Internet [492, 1254]. Interplay [1554]. Interpretation [724, 761]. Intersection [1043]. Intersectionality [3100]. Interspecific [200]. Intertidal [1928]. Intervention [568]. Interview [1126, 2310, 2496, 2793, 2536, 1684, 1764, 1801, 2684, 1796, 1391, 1954, 1732, 1763, 1914, 2620, 2868]. Introduced [2010, 212, 1857, 1844, 839]. Introducing [1825, 2723, 2324, 2767]. Introduction [426, 2543, 2328, 3150, 2329, 390, 1070, 1158, 2190, 1497, 2209, 1680, 1979, 1628]. Introductions [169, 170, 876, 894, 2082, 514, 51, 167, 166, 753, 165, 218]. Invader [2220, 2715]. Invaders [3112, 3240]. Invasion [1829, 973, 401]. Invasions [2559, 3238, 1989, 615]. Invasive [2698, 2218, 3263, 1838, 1158, 1157, 1892, 741, 3240, 886, 1428, 917, 838, 2276, 1415]. Invasivore [2778]. Inventory [2569]. Inverse [513]. Invertebrates [1681, 913]. Invertivorous [1647]. Investigating [590, 2385]. Investing [2870]. Investments [2508]. Invisible [621]. Involvement [2640]. Iowa [2240]. Iron [1510]. Ironhorse [2375]. Isaac [2297]. Island [887, 2659]. Islands [3228, 3214, 378]. Isotope [586]. Issue [1, 2, 4, 6, 9, 10, 11, 12, 13, 15, 19, 21, 22, 24, 26, 30, 32, 34, 36, 41, 48, 52, 53, 54, 56, 59, 62, 64, 70, 72, 74, 76, 81, 83, 84, 87, 91, 93, 95, 96, 99, 102, 106, 108, 110, 112, 113, 116, 119, 127, 130, 134, 138, 142, 145, 148, 151, 153, 156, 158, 161, 164, 171, 174, 178, 181, 184, 186, 190, 193, 197, 201, 204, 207, 210, 214, 217, 221, 225, 228, 231, 234, 237, 241]. Issue [244, 248, 251, 254, 258, 263, 271, 272, 275, 280, 283, 289, 292, 297, 302, 309, 312, 316, 319, 322, 325, 328, 332, 335, 338, 341, 344, 346, 348, 350, 352, 354, 357, 360, 362, 365, 368, 371, 374, 376, 379, 382, 385, 388, 392, 395, 397, 399, 402, 405, 407, 408, 412, 415, 418, 420, 422, 425, 428, 431, 434, 437, 440, 443, 446, 450, 453, 456, 458, 459, 462, 465, 468, 469, 471, 472, 474, 477, 480, 483, 485, 487, 489, 493]. Issue [496, 499, 501, 503, 506, 509, 512, 515, 518, 521, 524, 526, 530, 534, 537, 539, 542, 545, 548, 552, 555, 560, 563, 566, 569, 572, 575, 578, 580, 582, 585, 588, 591, 594, 597, 600, 603, 604, 607, 610, 614, 617, 620, 622, 625, 628, 630, 633, 636, 639, 642, 645, 648, 650, 653, 656, 658, 661, 663, 666, 669, 672, 675, 677, 680, 683, 685, 687, 690, 693, 696, 699, 702, 705, 708, 710, 713, 716, 719, 721, 726, 729, 733, 736]. Issue [739, 742, 745, 748, 751, 754, 757, 760, 763, 766, 770, 773, 775, 778, 781, 784, 788, 790, 793, 796, 799, 803, 807, 810, 814, 817, 820, 823, 825, 828, 831, 834, 837, 840, 844, 847, 850, 853, 855, 858, 861, 865, 868, 871, 874, 877, 882, 885, 889, 892, 896, 942, 945, 912, 915, 918, 921, 924, 927, 931, 934, 948, 967, 983, 998, 1010, 1023, 1038, 1054, 1065, 1080, 1093, 1107, 997, 1121, 1137, 1154, 1168, 1183, 1195, 1210, 1224, 1236, 1249]. Issue [1263, 1278, 1293, 1306, 1318, 1329, 1340, 1357, 1366, 1381, 1396, 1410, 1426, 1443, 1460, 1474, 1491, 1509, 1527, 1549, 1561, 1582, 1594, 1610, 1625, 1653, 1673, 1688, 1707, 1726, 1743, 1758, 1772, 1787, 1817, 1834, 1849, 1865, 1881, 1898, 1915, 1929, 1947, 1963, 1976, 2000, 2030, 2046, 2063, 2076, 2097, 2113, 2128, 2144, 2159, 2174, 2189, 2204, 2225, 2241, 2257, 2271, 2268, 2125, 2289, 2306, 2320, 2334, 2345, 2357, 2370, 2382, 2392, 2406, 2419, 2432, 2304, 2318, 2448, 2460, 2473, 2488, 2502, 2515, 2528]. Issue [2541, 2563, 2579, 2592, 2606, 2625, 2635, 2651, 2666, 2681, 2694, 2706, 2719, 2730, 2745, 2759, 2761, 2774, 2790, 2804, 2817, 2829, 2839, 2850, 2862, 2875, 2887, 2903, 2917, 2928, 2941, 2956, 2968, 2985, 3000, 3014, 3025, 3038, 3050, 3061, 3071, 3084, 3096, 3108, 3122, 3136, 3147, 3160, 3170, 3181, 3195, 3208, 3219, 3234, 3247, 3260, 3275, 85, 2543, 2190, 1979, 1793, 2426, 1628]. Issues [1641, 1584, 1745, 303, 1110, 1765, 1070, 999, 222, 43, 561, 410, 2272, 507, 285, 475, 1775, 809, 1325, 1640, 520, 2348]. Ivy [1841].

J [1140, 1542, 1502, 2769, 3284, 2524, 2662, 1321]. JAAH [1662]. Jack [1350, 1162, 2612]. Jackson [2479]. Jady [1542]. James [2785, 1270, 3088, 2749, 2936, 3007]. Janice [2054]. January [1192, 3227, 1136, 1179, 1305, 1326, 1315, 1506, 1487, 1704, 1723, 1943, 2142, 2125, 2290, 2332, 2318, 2449, 2626, 2791, 2942, 3095, 3246]. Japan [2103, 256, 88]. JASM [3031]. Jason [1391]. Jay [2185, 1763]. Jefferts [1666]. Jellyfish [2067]. Jerry [2265]. Jersey [2239, 2255]. Jesse [1691, 1941]. Jim [2936, 2479, 2468, 1801]. Jo [1542]. Job [911, 965, 981, 2780]. Jobs [1009, 1064, 1119, 1037, 1053, 996, 1022, 1106, 1092, 1079, 1182, 1235, 1291, 1153, 1136, 1208, 1223, 1167, 1194, 1277, 1262, 1248, 1317, 1305]. John [1540, 2138, 1350, 2479, 1162, 2732, 1924]. Johnny [2735]. Johns [2662]. Johnson [92]. Join [1122, 2363, 2986, 1557]. Joint [3210, 2613, 2173, 2980, 3031, 2504]. jointly [175]. Jon [2496]. Jones [3186, 2868]. Jordyn [2868]. Joseph [1542, 1088, 1969]. Journal [899, 951, 1059, 952, 1003, 1021, 1032, 1090, 1116, 1150, 1165, 1192, 1209, 1232, 1246, 1289, 1303, 1326, 1355, 1378, 1408, 1441, 1472, 1506, 1524, 1579, 1592, 1622, 1686, 1704, 1739, 1769, 1831, 1861, 1878, 1926, 1912, 1958, 2027, 2059, 2094, 2142, 2157, 2202, 2223, 2268, 2286, 2417, 2430, 2445, 2296, 2332, 2343, 2380, 2390, 2458, 2471, 2486, 2498, 2513, 2526, 2539, 2561, 2577, 2590, 2604, 2618, 2623, 2649, 2664, 2679, 2692, 2704, 2717, 2728, 2741, 2756, 2788, 2772, 2814, 2801, 2827, 2837, 2848, 2860, 2873]. Journal [2885, 2901, 2915, 2926, 2939, 2954, 2965, 2982, 2997, 3011, 3022, 3035, 3047, 3058, 3068, 3081, 3093, 3105, 3119, 3133, 3144, 3157, 3167, 3178, 3192, 3205, 3231, 3244, 3216, 3256, 3270, 3286, 175, 1192, 1077, 1260, 1150, 1338, 1394, 1458, 1579, 1815, 1926, 1972, 2043, 2111, 2172, 2238, 2304, 2354, 2404]. Journalism [1434]. Journals [2411, 1530, 3154, 895, 2669, 857, 2258, 1746, 1711]. Journey [3196, 3267]. Jr [2185, 3227, 2936]. Jr. [1556, 1620, 2366, 1782]. JSFS [1548]. Juan [3284]. Judgment [659, 845]. Julie [2479]. July [3024, 1270, 2786, 2739, 1053, 1223, 1607, 2073, 2268, 2254, 2371, 2529, 2707, 2863, 3169]. Juncture [347]. June [3013, 3088, 2935, 3227, 3268, 1037, 1260, 1208, 1394, 1579, 1592, 1815, 2043, 2059, 2238, 2223, 2358, 2404, 2390, 2503, 2695, 2851, 3159]. Junior [2087, 2545, 2723, 2763, 2868, 1007]. Just [1551, 2733, 1727, 2260]. Justice [449, 324]. Justifying [497]. Juvenile [3264, 68, 901, 869, 2053].

Kaiser [1126]. Kambouris [2709]. Kamchatka [2674]. Kansas [1921, 1903, 1959, 2089, 1935, 3110, 173]. Kayak [2709]. KC [2024]. Keep [2653, 1738, 1858, 2365]. Keeping [3073, 3079]. Keeps [1880]. Keith [1666]. Kells [2662, 2684]. Kenai [2299, 481]. Kenneth [1744]. Key [3003, 2494, 2818, 2100, 2348]. Keys [1111]. Killed [1503]. Kills [233]. King [2352, 2192]. Kings [3188]. Klamath [730, 659]. Klein [1667]. Klein-MacPhee [1667]. Klickitat [287]. Knew [2016, 1906]. Know [1541, 3272, 954, 1719, 2016, 3288, 1706, 1950, 3257, 913, 1539]. Knowing [3220]. Knowledge [1851, 2812, 988, 1528, 2535, 1111]. Known [846, 3043, 1868]. Known-Age [3043]. Knowns [264]. Koike [1955]. Kolz [2911]. Kong [2361]. Kootenai [525]. Korea [619, 835]. Korean [1523]. Kristen [1733]. Kurile [2674]. Kuskokwim [3188].

L [992, 1139, 1420, 2316, 2979, 2963]. Lab [1804]. Labels [3148]. Laboratories [140]. Laboratory [224, 2923, 141, 1333, 168, 2020, 1801]. LACFC [3190]. Lacustrine [150]. Lahontan [2639]. Lake [2742, 3245, 2339, 189, 1776, 424, 38, 160, 39, 752, 2819, 2069, 890, 2439, 3127, 404, 484, 776, 180, 2391, 3106, 2674, 1565, 714, 3263, 2026, 424, 1630, 321, 455, 587, 1280, 105, 1751, 2333, 832, 398, 49, 94]. Lakes [1500, 1030, 2779, 2531, 2509, 120, 304, 1973, 125, 433, 141, 940, 2520, 2534, 632, 401, 1018, 146, 1266, 2715, 1776, 421, 1892, 3278, 1308, 727, 78, 1750, 2069, 2108, 77, 141].

Lamprey [2705, 2167, 2522, 638, 2715, 1698, 1227]. Lampreys [969, 946]. Land [2339, 1783, 1312, 1273, 466]. Land-Water-Humankind [1312]. Lander [1775]. Landing [612]. Landlocked [595, 1803]. Lands [637, 925, 2398]. Landscape [1573, 1198, 1517, 857, 1564]. Landscape-Scale [1564]. Landsman [1914, 2868]. Language [1809]. Lapsed [891]. Large [1573, 177, 1094, 1586, 278, 862, 3065, 829, 135, 2736, 2930, 547, 795, 2121, 1026, 641]. Large-Scale [177, 829, 1026, 278, 795]. Largemouth [103, 681, 20, 467]. Largest [3138]. Largetooth [2164]. Larimore [1768]. Larry [1768, 2911]. Larson [2316]. Larvae [1404, 1760]. Larval [2423]. Last [1713]. Lasting [1362]. Lately [3209]. Later [1665, 2103]. Latin [3141, 3116, 3128, 3019]. Launching [2811]. Laurence [2799]. Laurent [212]. Lavett [1782]. Law [25, 1937, 2193, 928]. Lawrence [2911]. Lawsuits [3029]. Lead [846, 1131, 1300]. Leaders [1567, 1020, 2050, 3162]. Leadership [1563, 3182, 2853, 1075, 949, 3196, 1218, 984, 3281]. Leading [1100, 1463, 1055]. Leads [2639, 1777]. Leaky [3052]. Leanne [2315]. Learn [2627, 1695, 310, 1968]. Learned [848, 812, 3016, 829, 2811, 2150, 2477, 864]. Learning [2136, 618, 390, 731, 1776, 1695, 1635, 1145, 403, 1821, 2326, 941]. Least [1868, 1857]. Lee [2054, 3102]. Legacy [1900, 1279, 1002, 2725, 1659, 2112]. Legal [2507]. Legends [2598]. Legislation [900, 953, 971, 986, 1012, 1025, 1040, 1082, 1056, 1067, 230, 1000]. Lembeh [2249]. Length [797, 467]. Length-based [797]. Lengths [1371, 100, 2510]. Lens [2879, 3276, 2724]. Lepidogalaxias [238]. Leslie [1219]. Less [1809, 1049]. Lessons [2805, 1867, 2415, 848, 812, 3016, 1001, 2990, 1084, 573, 829, 532, 2811, 2150, 1760, 2251, 2477, 441, 577, 864]. Let [2841, 2943, 2374]. Letter [1331, 2033]. Letters [3027, 1398]. Level [2386, 1267, 886, 1807, 1712, 2008]. Levels [586, 3226, 550, 3127]. Leverage [2084]. Leveraging [1994]. Lewis [435]. lewisi [3223]. Lexicon [1596]. LGBTQ [2920]. Liaison [2115]. Liberating [3262]. Lice [3126]. Life [2522, 1076, 2170, 943, 1800, 104, 2284, 1240, 3267, 1519, 651, 2452, 176]. Life-span [651]. Lifelong [403]. Lifestage [870]. Lifestyle [787]. Light [2011, 3235]. Like [2931, 2792, 3054]. Lime [123]. Limestone [122]. Liming [120, 125]. Limit [822, 779, 1300]. Limited [73, 2833]. Limiting [2198]. Limits [185, 599, 2412, 135, 797, 1775, 467]. Limnologist [2316]. Linda [2643]. Lindsey [3113]. Line [1503, 989, 974, 950, 3002, 2521]. Lines [1434, 2356]. Lingering [2711, 2958]. Link [1391]. Linkages [1370]. Linking [2674, 851, 2509, 1322, 873, 932]. Links [886]. Lion [2766]. Lionfish [2218, 2220, 1747]. listed [821]. Listen [2633]. Listening [782, 1848]. Literally [1774]. Literature [3222, 315, 259]. Literatures [378]. Little [1311, 1324, 1336, 1351, 1242, 1256, 1269, 1648, 1449]. Littoral [3127, 940]. Liturgy [1177]. Live [2671, 1532, 3240, 3110, 1085, 2147]. Live-Imaging [3110]. Live-Release [2671]. Live-Streaming [2147]. lived [558]. Livelihoods [1467]. Lives [1659]. Livestock [375, 1583]. Living [1528, 1099, 1725]. Loan [1738]. Lobster [68, 1729, 65, 947, 1760, 66, 69, 67]. Local [963, 881, 1712, 2832]. Localized [795]. Locally [1658, 1427, 584, 2408]. Location [2908, 2452]. Locations [1517, 1721]. Lochie [1542]. Logger [155]. Logo [2945]. Logperch [3048]. Logs [862]. Long [2339, 558, 2992, 2026, 1226, 1696, 3201, 2121, 2476, 592]. Long-lived [558]. Long-Term [2339, 2992, 2026, 1226, 1696, 3201, 2121]. Longer [2531]. Longhurst [1015]. Longitudinal [2780]. Longlining [1803]. Look [2722, 2763, 2931, 1258, 1323, 3054]. Looking [903]. Loss [720, 1386, 1694]. Losses [723]. Lost [916, 2867]. Lot [3054]. Lota [2665]. Lots [3004]. Louisiana [1638]. Low [2855, 919, 2104]. Low-Carbon [2855]. Low-cost [919]. Lower [2230, 146, 1783, 78, 2039, 2056, 2439, 1343, 2229, 3153]. Lure [1855, 668]. Lure-size [668].

M [1556, 1911, 3088, 2662]. Maceina [2769]. Machines [2701]. MacIntire [175]. Macintosh [383]. Macon [1204]. MacPhee [1667]. Macrophytes [484]. maculatus [2381]. Made [223, 1884]. Magazine [2567, 2669, 3236]. Magnuson [2427, 1646, 2409]. Mahseer [2834]. Mail [2567, 2669, 3236]. Maine [686, 213, 1613, 86, 789, 1760, 2880, 753]. Mainstem [746, 2066]. Maintaining [895, 2228]. Makah [2865]. Make [1709, 2637, 2199, 2015, 2328, 68, 1160, 2323]. Making [1435, 2492, 2594, 2696, 1818, 390, 987, 2600, 2337, 3043, 1450, 452, 2930, 75, 2206, 426, 239]. Male [3179]. Man [101]. Manage [1266, 2082, 1695, 1308, 2519, 284, 1428, 2932, 168]. Manageable [1514]. Managed [2026, 260, 1775, 3029, 3018]. Management [833, 939, 196, 969, 1100, 667, 2674, 1678, 631, 223, 367, 2537, 202, 1658, 1175, 23, 1522, 1494, 1749, 2146, 2243, 2386, 1909, 843, 192, 185, 637, 1371, 281, 3077, 303, 296, 2207, 2686, 2218, 1645, 623, 697, 714, 17, 747, 946, 2078, 3264, 647, 1110, 2134, 2463, 1765, 189, 494, 3226, 2220, 777, 68, 205, 1266, 1501, 128, 988, 2018, 1430, 1419, 394, 571, 1818, 2507, 881, 361, 2855, 1001, 1258, 657, 1068, 430, 314, 2373, 1537, 439, 3074, 18, 846, 1937, 910, 502, 79, 2509, 1677, 320, 1776, 852, 2570, 576, 421, 2598, 2494, 114, 1176]. Management [1891, 2349, 999, 1011, 1024, 1039, 1055, 949, 1461, 2058, 1391, 987, 3220, 411, 2521, 925, 835, 160, 525, 676, 772, 2974, 2819, 1084, 629, 2244, 561, 1417, 324, 433, 3151, 1050, 2069, 2122, 1970, 2971, 432, 154, 2176, 410, 1554, 162, 2198, 535, 2041, 219, 935, 2427, 2117, 285, 1646, 298, 274, 295, 557, 3075, 150, 3072, 584, 2809, 559, 409, 1734, 1983, 2399, 654, 1450, 873, 2711, 875, 3164, 944, 1995, 2299, 1253, 1614, 984, 452, 1893, 290, 2962, 2685, 832, 2398, 531, 1890, 2517, 2520, 2807, 613, 2150, 229, 313, 3053, 3213, 377, 451, 2263]. Management [1938, 249, 3221, 662, 2325, 684, 3201, 1552, 2797, 2595, 2865, 2453, 1857, 216, 1384, 2276, 2751, 2495, 1325, 795, 2832, 1370, 2506, 262, 583, 315, 481, 1859, 448, 824, 640, 2484, 2409, 616, 2228, 2302, 3140, 67, 2869, 2812, 1831, 1059, 1021, 1116, 1165, 1209, 1246, 1303, 1355, 1408, 1472, 1524, 1592, 1622, 1686, 1739, 1878, 1912, 1958, 2059, 2157, 2223, 2286, 2343, 2390]. Managerial [139, 373]. Managers [1500, 1517, 198, 611, 175, 2019, 1699, 2822, 3072]. Managing [568, 416, 3250, 3150, 511, 842, 740, 1695, 1635, 2702, 3065, 2198, 2842, 589, 484, 3017, 2040]. Mandate [2161]. Mandates [583]. Mangroves [2093]. Mania [1514]. Manipulation [122]. Many [1492, 3271]. Map [2116, 2244, 1344, 804]. Mapping [2068, 919, 857, 378]. Marcellus [1295]. March [1192, 3113, 3088, 3114, 2936, 996, 1167, 1338, 1544, 1755, 1972, 2172, 2119, 2187, 2321, 2354, 2368, 2474, 2652, 2816, 2967, 3121, 3274]. Marcos [1577]. Margaret [991, 2479, 2314]. Mariculture [213]. Marietta [60]. Marine [1087, 1257, 1282, 1451, 1452, 1648, 1661, 1875, 2055, 2252, 2428, 2587, 2753, 2912, 3056, 3211, 2537, 2677, 2213, 1749, 2243, 1681, 688, 624, 806, 876, 2232, 1419, 265, 502, 863, 1729, 1015, 369, 282, 3100, 528, 476, 993, 1990, 544, 546, 581, 556, 1240, 1014, 3005, 1634, 700, 1775, 547, 684, 166, 347, 1213, 1659, 3252, 429, 305, 2106, 583, 1660, 640]. Marine-derived [476]. marinus [2705]. Maritime [2361]. Mark [963, 1132, 1189]. Markers [343]. Market [756]. Markets [3066]. Marking [635, 593]. Martin [60]. Maryland [2754, 2770, 2882, 2913]. Mason [1542]. Massachusetts [2552]. Masses [2147]. Match [390]. Materials [505]. Matt [2216]. Matter [2908, 629, 1828]. Matters [2107, 769, 2490]. Maximize [2148]. May [624, 2657, 2963, 2265, 2326, 2467, 3026, 1837, 1674, 1022, 1194, 1221, 1363, 1558, 1784, 1997, 2346, 2500, 2682, 2840, 2999, 3146]. Maybe [1551, 2308]. Maynard [1714]. Mazatlan [1560]. McCloud [592]. McIntyre [1350]. McKenzie [1097]. Me [2089, 1535, 2969, 2475]. Mead [105]. Meal [340]. Mean [2467, 1807]. Means [956, 1760, 1589]. Measurement [1371, 100]. Measures [16, 359, 165]. Measuring [224]. Meat [1482]. Mechanical [1586, 1838]. Mechanics [482]. Mechanisms [2548]. Mechanistic [851]. Media [1569, 1390, 2327, 1335, 1393, 893, 3054, 2326, 1465, 1479, 2325, 2454, 2034]. Meditation [1672]. Mediterranean [2494]. Meehan [2725]. Meet [1373, 1812, 2599, 3030, 1127]. Meeting [3210, 1112, 979, 1180, 1190, 1164, 1268, 1603, 1619, 1629, 1656, 1520, 1505, 1832, 1845, 1860, 1870, 1921, 2037, 2051, 2072, 2086, 2169, 2200, 2239, 2109, 2255, 2446, 2511, 2456, 2673, 2754, 2770, 2680, 2781, 2882, 2913, 2870, 2993, 3090, 3089, 2980, 3031, 3142, 3239, 3165, 3254, 1658, 1873, 2089, 2280, 1341, 2308, 1617, 1778, 964, 2148, 1217, 612, 3123, 1349, 2041, 1727, 1204, 1639, 2504, 907, 962, 1220, 1548, 1576, 1486, 2070, 2062, 1996, 2155, 2341, 2523, 2673, 2949, 3033, 1932, 1523, 1823, 1795, 2687, 2532, 2464, 2871, 1205]. Meeting [3008, 1967, 2024, 1449, 2376, 1816, 2831, 1132, 2440, 1230, 3235]. Meetings [2478, 1476, 1773, 1358, 2148, 1138, 2313, 1011, 2905, 905, 963, 1918, 1161, 1619]. Meets [2302]. Mega [3152]. Mega-Risk [3152]. Megadroughts [3152]. Mekong [2039, 3173, 2833]. Member [1405, 2363, 1790, 2133, 1271]. Members [963, 1033, 1071, 1114, 1044, 1061, 980, 1151, 1178, 1373, 1349, 1245, 2004, 977, 1201, 1233, 1327, 1354, 1365, 1377, 1424, 1440, 1489, 1508, 1525, 1545, 1581]. Members/Newly [1377]. Membership [1811, 3045, 866, 1393, 1904, 1708, 1143, 3085]. Memoriam [3186, 1073, 1088, 1147, 1189, 1620, 1542, 1571, 1502, 2995, 1162, 2785, 2769, 2979, 3102, 2798, 3185, 3113, 2799, 3088, 2524, 2963, 2935, 3114, 3227, 2911, 3268, 2936, 1104, 2786, 3199, 2755, 2739, 3007]. Men [787]. Mention [1886, 1216]. Mentor [3209]. Mentoring [1279, 3171, 2853, 897, 253, 2780, 1952, 2768]. Mercury [58, 2300]. Mercury-Induced [58]. Merging [2907]. Merits [2364]. MesoHABSIM [608]. Message [2323]. Messaging [2778]. Metaphor [2778]. Method [870, 988, 1908, 1097, 29, 2821, 704, 28, 749]. Methodological [3042]. Methodology [288, 239, 1990, 725, 762, 926]. Methods [523, 1871, 2135, 2083, 3202, 25, 644, 678, 3109, 498, 90, 2610, 583]. Metrics [1533, 2042]. Metropolises [1471]. Mexican [2855, 1697]. Mexico [1125, 2230, 1678, 2165, 750, 688, 1041, 1089, 2521, 1908, 488, 1042, 1694, 2229, 1702, 1643]. MFCMA [202]. Michael [1542, 2995, 2769]. Michel [1126]. Michigan [141, 2391, 3239, 714, 772]. Micro [2748]. Micro-Fishing [2748]. Microbial [1681]. Microbiome [2689]. Microbiota [1516]. Microbrew [1837]. Microchemistry [1614]. Microcomputer [224, 100]. Microthemes [391]. Mid [3171, 2751, 1095]. Mid- [2751]. Mid-Atlantic [1095]. Mid-Career [3171]. Middle [179]. Midnight [2166]. Midshipman [3149]. Midway [2479]. Midwestern [3150, 492, 80]. Might [1296]. Migrating [2773]. Migration [1483, 689, 1838, 3280, 872, 2495, 2620, 2734]. Migrations [851, 1894]. Migrators [2143]. Migratory [2910, 2058]. Milestones [1132]. Millennial [1143]. Millennials [1035]. Mind [1858]. Minds [1738]. Mini [1004, 988]. Mini-Reviews [988]. Mini-travel [1004]. Minimal [1760]. Minimize [2866]. Mining [1397, 1889, 507, 1141, 1457, 928]. Minneapolis [1180, 1127, 1173, 1164, 1268]. Minnesota [599, 1973, 767, 404, 940]. Minnow [926]. Minnows [215]. Minorities [1904]. Minority [1075]. Minority-Serving [1075]. Miracles [198]. Misconception [849]. Mislabeling [1922, 3225]. Mismanagement [1015]. Mismatch [2600]. Mismeasure [615]. Miss [963, 1795, 1046, 2024]. Missing [2532]. Mission [993]. Mississippi [1783, 1801, 2437, 2056, 1018, 606, 2058, 232, 1212, 3238, 2263]. Missouri [2439, 2089, 894, 187, 227, 333, 114, 2640, 841]. Misuse [311]. Mitigating [1470]. Mitigation [2714, 227, 63, 252, 121, 124, 191, 2360]. Mixed [3109, 343, 768]. Mixture [2312]. MN [1268]. MO [1921]. Mobile [1780]. Model [103, 813, 1212, 3139, 2299, 1852, 2263, 3221, 2312]. Modeler [2366]. Modeling [1494, 2264, 2329, 318, 1683, 2767, 452, 1775]. Models [300, 1076, 2992, 2151, 2247, 768, 608]. Modern [2865]. Modernize [2906]. Mola [2998]. Monetary [704]. Money [1476, 2833]. Monitor [993, 2932]. Monitored [1910, 3250, 1909]. Monitoring [2750, 2067, 413, 2843, 574, 2992, 1550, 2819, 1226, 829, 1370, 2121, 173, 2385, 1679, 786]. Monocultures [1681]. Montana [189, 1313, 250, 533, 643]. Month [1955]. Moore [2726]. Mora [1840]. Moratorium [589]. Morone [146]. Morphoedaphic [1793]. Morphological [1404]. Morphology [238]. Morphometric [2834]. Morrison [1587]. Mortalities [958, 914]. Mortality [2083, 689, 2412, 2151, 2481, 1267, 678, 504, 2991]. Morton [203]. Mosquitofish [3120]. Most [2388, 1709, 1795, 787, 68, 1630, 2337, 1160]. Mothers [2713]. Motion [1850]. Motivates [2425]. Motivation [235]. Motivational [2464]. Motivations [361, 2948, 1566]. Motives [114]. Mount [1216]. Mountain [2143, 1281]. Mountains [320, 261]. Move [3240, 1682]. Moved [917]. Movement [2166, 1253, 2734]. Movement-centered [1253]. Movements [2217]. Moving [917, 1807]. MSU [1937]. MSUAFS [1313]. MSY [35]. Much [1827, 2233, 2047, 1708]. Mucus [1516]. Mudbass [2477]. Mudminnow [1585, 2499]. Muhlfeld [2536]. Multi [1369, 1681, 813, 897, 484, 492]. Multi-Directional [897]. Multi-Factor [813]. Multi-lake [484]. Multi-state [492]. Multi-Trophic [1369, 1681]. Multidisciplinary [1716]. Multilevel [768]. Multiple [543, 1806, 2582, 143, 1098, 886, 2398]. Multiple-Use [2398, 543]. Multiples [1174]. Multispecies [284]. Multistate [2151]. Multivariate [644]. Murphy [991, 2479, 2153, 2314]. Museum [2361]. Muskellunge [1765, 933]. Mussel [2386, 213, 401]. Mussels [337, 2675, 1099, 345]. Must [1359, 1202, 1411, 2896, 3236, 2422]. Must-Do [1202]. Mutualism [2131]. My [1945, 1886, 3196, 1799, 3267]. mykiss [2773, 3223]. Myself [2205]. Mystery [2533]. Myths [1734, 517]. Myxobolus [470].

N [869]. Nagid [2214]. NAJFM [1820]. Name [2857, 101, 2989, 3041]. Names [109, 1908, 1337, 3237, 1697, 226, 2533]. Napping [1741]. Narrative [2559, 2897]. Narratives [1919]. Narrow [2018]. NAS [1323]. Nassau [613]. nasus [3175]. Natal [2053]. Nation [1949, 2099, 2395]. National [1854, 2405, 2849, 1823, 1553, 2339, 2463, 364, 551, 1840, 1070, 1109, 826, 2011, 2056, 2413, 2100, 1083, 92, 429, 307, 1573, 16, 659, 734, 2401, 612, 1187, 2398, 1297, 1069]. Nationalities [1621]. Nationwide [3277]. Native [822, 3223, 2183, 747, 1027, 734, 3152, 3112, 1186, 423, 424, 2340, 2971, 268, 746, 1971, 2972, 475, 2933, 643, 615, 2806, 51, 2551, 5, 579, 1026, 2778, 168]. natura [2032]. Natural [1206, 2921, 303, 461, 3226, 1615, 18, 2412, 1109, 1461, 2455, 2640, 1231, 2842, 2399, 2972, 2807, 869, 2595, 69, 2586, 2228]. Naturalists [1177]. Nature [1759, 1919, 2841, 2032]. Nature-Based [1759]. Navigable [919]. Navigate [1748]. Navigating [2687]. Nays [486]. NCIFD [3212]. Neal [2973]. Near [2339]. Nearly [2639]. Nearshore [3099, 842, 3078, 260]. Need [2891, 2065, 2259, 2208, 2890, 989, 843, 3202, 2585, 2373, 2311, 208, 278, 63, 3271, 2972, 3164, 2194, 307, 2495, 3153]. Needed [2010, 159, 2697]. Needlefish [1272]. Needs [381, 1115, 2167, 2671, 1765, 937, 2137, 358, 139, 2011, 209, 261, 2198, 274, 403, 1539, 2611, 3005, 613, 1855, 549, 1295, 824, 2566, 505]. Neglect [249]. Neglected [369]. Nelson [1088]. Nematode [516]. NEPA [16]. NerkaSim [473]. Net [2605, 286, 1634, 1747, 3262]. Net-Pen [1634]. Netflix [2990]. Neth [2071]. Nets [1693]. Netting [2742]. Network [1369, 1110, 777, 1225, 1070, 2700, 1083, 1213, 2879, 1297]. Networking [3171, 2710]. Networks [1310, 1081, 1066, 1184, 1229, 1211, 1196, 544, 2326, 1127]. Neutralization [126]. Nevada [2446, 2613, 205]. Never [37]. Newly [1377]. News [908, 898, 1466, 1481, 1498, 1521, 1286, 2480, 1633, 2034]. Next [2975, 3278, 2017, 2013, 2004, 1949, 1146, 2008, 2014, 1349]. NHDPlus [1859]. Nicaragua [94]. Night [1275]. Nile [2838]. niloticus [2838]. Nine [2209, 2210]. Nitrogen [586]. NMFS’ [2244, 1095]. No [1621, 1431, 1888, 141, 544, 3262, 216]. No-take [544]. NOAA [1798, 1095, 1799, 2020]. NOAA/NMFS [1095]. Noah [1540]. Noakes [2785]. Noise [1819, 2170, 3149]. Nome [2352]. Nomenclatural [101]. Nominating [2975, 3129]. Nominations [1284, 1453, 1663, 1876, 2091, 2283, 2443, 2616, 2784, 3281]. Non [2910, 1058, 3112, 2806, 51, 2778, 168]. Non-Game [1058, 3112]. Non-Native [2806, 51, 2778, 168]. Non-Salmonid [2910]. Nonfederal [339]. Nonindigenous [758, 573, 514]. Nonlethal [1692]. Nonmember [2363]. Nonmigratory [111]. Nonnative [747, 2107, 1013, 701, 1586, 2349, 752, 577]. Nonpermanent [3198]. Nontraditional [3075, 3072]. Norman [1571]. Norris [623]. North [1631, 701, 337, 279, 1971, 822, 969, 202, 712, 843, 519, 946, 1752, 366, 3188, 419, 1001, 1923, 854, 779, 660, 573, 1981, 2122, 565, 732, 2149, 414, 243, 295, 581, 82, 2809, 601, 1979, 2895, 1614, 230, 276, 3221, 1214, 1980, 448, 242, 1059, 952, 1003, 1021, 1032, 1090, 1116, 1150, 1165, 1192, 1209, 1232, 1246, 1289, 1303, 1326, 1355, 1378, 1408, 1441, 1472, 1506, 1524, 1579, 1592, 1622, 1686, 1704, 1739, 1769, 1831, 1861, 1878, 1926, 1912, 1958, 2027, 2059, 2094, 2142, 2157, 2202, 2223, 2268, 2286, 2332, 2343]. North [2380, 2390]. Northeast [1718, 401, 1792, 3055, 550]. Northeastern [1778, 326, 2675, 2807]. Northern [1765, 438, 1432, 601, 1241, 73, 2483, 2712, 1089, 1657, 2583, 2960, 2164]. Northwest [1186, 356, 522, 1910, 3251, 652, 750, 177, 155, 550, 1124, 331, 1157, 293, 423, 1630, 972, 386, 2198, 549, 1757, 973, 2020]. Norway [233, 121]. Note [2568, 2103]. Notes [1786]. Notropis [2381]. Nova [123]. Novel [3099, 919, 1868, 1747]. November [1911, 2882, 2913, 2870, 2785, 3114, 1106, 1277, 1670, 1895, 2420, 2593, 2760, 2918, 3070, 3218]. NSERC [1083]. Nuisance [317]. Number [966, 982, 1059, 1003, 1021, 1032, 1090, 1116, 1048, 1103, 1260, 1150, 1165, 1192, 1209, 1232, 1246, 1289, 1134, 1179, 1221, 1274, 1338, 1394, 1458, 1303, 1326, 1355, 1378, 1408, 1441, 1315, 1363, 1423, 1472, 1506, 1524, 1579, 1592, 1622, 1487, 1544, 1607, 1650, 1670, 1815, 1686, 1704, 1739, 1769, 1831, 1861, 1878, 1723, 1755, 1784, 1846, 1926, 1972, 2043, 1912, 1958, 2027, 2059, 2094, 1895, 1943, 1997, 2073, 2111, 2172, 2238, 2142, 2157, 2202, 2223, 2286, 2187, 2254, 2290, 2307, 2321, 2335, 2346]. Number [2358, 2371, 2383, 2393, 2407, 2420, 2433, 2354, 2404, 2332, 2343, 2380, 2390, 2368, 2449, 2461, 2474, 2489, 2500, 2503, 2529, 2542, 2564, 2580, 2593, 2607, 2626, 2636, 2652, 2667, 2682, 2695, 2707, 2720, 2731, 2746, 2760, 2775, 2791, 2803, 2816, 2830, 2840, 2851, 2863, 2876, 2888, 2904, 2918, 2929, 2942, 2957, 2967, 2984, 2999, 3013, 3024, 3037, 3049, 3060, 3070, 3083, 3095, 3107, 3121, 3135, 3146, 3159, 3169, 3180, 3194, 3207, 3218, 3233, 3246, 3259, 3274, 1773, 168]. Numbers [1412, 2104]. Nursery [144, 433]. Nurture [2841]. Nutrient [550, 476, 800]. Nutrients [540].

Oases [104]. Obama [1331]. Objective [2820]. Objectives [2537, 783, 925, 1370]. Obligations [1639]. Observation [845]. Observations [1089, 135, 2991]. Observatory [1344, 1297]. Observed [122, 527]. Obstacles [2958]. Occupancy [3101]. Ocean [2677, 567, 1110, 2170, 1109, 2835, 2019, 342, 1344, 2994, 3162, 843, 750, 1110]. Oceans [2415, 2531, 1213, 1682]. October [1911, 2446, 2456, 2785, 2979, 2963, 1092, 1289, 1262, 1441, 1622, 1861, 1878, 2094, 2286, 2407, 2580, 2746, 2904, 3060, 3207]. Ode [2668, 1701]. Odyssey [2686]. off [326]. Offer [416, 1746]. Offering [2012]. Offers [1647]. Office [2022, 1372]. Officer [1948]. Offs [832]. Offshore [1646, 2385, 2747]. Ohio [2673, 195, 321, 455, 587, 1280, 3163]. Oil [956, 1638, 1794, 1089, 1042, 131, 1457, 1702]. Ok [3026, 2308]. Okaloosa [3023, 2758]. okaloosae [2758]. Okeechobee [832]. Oken [2989]. Oklahoma [398, 3053]. Olaf [2799]. Old [1694, 2571, 2865, 583]. Olfactory [2910]. Oligotrophication [564]. Olympic [2499]. Omernik [3088]. Omnibus [1095]. Onboard [1975]. Oncorhynchus [2773, 3223, 78, 3126]. One [2082, 1916, 1089, 1387, 2889, 3267, 1693, 1868, 2251, 2734]. One-Day [2734]. Oneida [2026]. Ones [3026]. Ongoing [1362]. Online [2377, 1465, 2108]. Ontario [1565, 741, 2819, 49]. Op [1792, 1791]. Op-ed [1792, 1791]. Open [2311, 2808, 2858, 2293, 172, 1535, 1601]. Open-Access [172, 1601]. Opening [3008]. Operating [3127, 902]. Operation [3200, 131, 1640]. Operational [323, 286]. Opinion [833, 3139, 2629, 2646]. Opinions [631, 717]. Opportunistic [1613, 2832]. Opportunities [426, 2065, 2115, 2135, 1798, 2714, 358, 411, 2924, 589, 2611, 3030, 124, 1779, 2360, 2309]. Opportunity [189, 2349, 2675, 136]. Opt [3236]. Opt-In [3236]. Optic [2922]. Optimism [1945]. Optimum [432, 35]. Oral [2535]. Order [42]. Ordered [697]. Oregon [2418, 1867, 2339, 647, 806, 1796, 342, 1097, 273, 2203, 1230]. Oreochromis [2838]. ORERP [3163]. Organic [1633]. Organism [2116]. Organismal [1361]. Organisms [169, 170, 167, 166, 165]. Organization [39, 2019]. Organizational [531]. Organizations [2035, 3164, 1566, 2228]. Organizing [2477]. Origin [1325, 1640]. Original [1714]. Origins [2053, 2795]. Orr [3284]. Orth [2662]. Osprey [2874]. Other [1681, 2170, 2559, 3220, 2864, 179, 317, 533]. Otolith [1824, 2053]. Otoliths [1880]. Otter [2712]. Otto [2025]. Our [1373, 1575, 1709, 1728, 2161, 2415, 2347, 1341, 2199, 1081, 1155, 1169, 2531, 1011, 1177, 3271, 2920, 284, 2889, 1949, 2260]. Outbreak [2677]. Outcome [2897]. Outcomes [2476]. Outdoor [1212]. Outlaw [841]. Outreach [551, 1615, 2295, 2005, 495, 2454]. Outs [1776]. Overcome [2600]. Overdue [928]. Overfished [916]. Overfishing [819, 529, 1539]. Overlooked [3112, 2040]. Overlooking [1564]. Overview [682, 571, 1265, 2574, 2819, 561, 2809, 601, 2298, 90, 92, 441, 353]. Owens [508]. Own [2439]. Owner [223]. Ownership [282]. Oxbow [1030]. Oyster [1591]. Ozone [1760].

P [1690, 2185, 2662]. Pacific [550, 1719, 2327, 1214, 3228, 3214, 444, 202, 522, 1910, 2152, 586, 3251, 652, 2083, 567, 540, 2167, 638, 1076, 3055, 1266, 1001, 155, 821, 550, 1124, 513, 1158, 1157, 1186, 423, 1630, 972, 386, 2198, 581, 273, 559, 2298, 1971, 1697, 473, 2251, 549, 1757, 973, 1698, 262, 1227, 2991, 1718]. Package [2293, 2042, 3212]. Pad [100]. Paddlefish [3059, 819, 80, 2058, 2702, 1253, 3053, 2495]. Page [966, 982, 1914, 1928, 1962, 1975, 1999, 2029, 2045, 2061, 2075, 2096]. pages [3284, 2185, 2662]. Painted [2270]. Painting [2973]. Paired [1946]. Pairs [665]. Paleochannels [2056]. Paleoclimate [1657]. Pallid [2369, 2650, 1677, 107]. Panama [33]. Pandemic [2877, 2893, 2809]. Paper [2119, 2132, 1632, 1889, 175, 1433]. Paperback [3284, 3267]. Papers [1113, 1086, 1101, 1285, 1256, 1269, 1406, 1455, 1421, 1438, 1603, 1619, 1629, 1656, 1832, 1845, 1860, 1870, 2070, 2072, 2086, 2269, 2255, 3203, 2051]. Paradigm [1721, 1677, 2933, 547, 2453]. Paradise [2052]. Paradox [801]. Parasite [516, 2688]. Parasitic [946]. Parentage [2571]. Parentage-Based [2571]. Parenthesis [2079]. Park [2262, 2339, 1187, 2413]. Parkinson [1258]. Parks [114, 495]. Parrish [992]. Part [1614, 1062]. Partial [872]. Participant [1727]. Participate [2425]. Participation [235, 2359, 1074, 798]. Participatory [2535]. Partner [1633]. Partners [281]. Partnership [2339, 2552, 1414, 1213, 2553, 1362]. Partnerships [1818, 2400, 2736, 1376]. Pascagoula [1212]. Paso [2185]. Pass [3002]. Passage [909, 339, 609, 574, 2116, 2906, 2820, 2066, 1735]. Passing [1499, 3182]. Passive [2351]. Past [1668, 3266, 2038, 1281, 1531, 1590, 1821, 2960, 2632, 592, 243]. Pasture [2246, 2081, 2248]. Patenting [247]. Path [1810, 2517]. Pathway [973, 2812]. Pathways [3150, 2018, 1982, 3008]. Patronyms [101]. Patterns [2083, 235, 1186, 31, 3101, 514]. Patterson [2215]. Paul [1180, 1268, 2139, 1127, 1173, 1164, 2071, 2212, 1796]. Pauly [1684]. PCB [8]. PDF [912, 915, 918, 921, 924, 927, 931, 934, 948, 967, 983, 998, 1010, 1023, 1038, 1054, 1065, 1080, 1093, 1107, 1121, 1137, 1154, 1168, 1183, 1195, 1210, 1224, 1236, 1249, 1263, 1278, 1293, 1306, 1318, 1329, 1340, 1357, 1366, 1381, 1396, 1410, 1426, 1443, 1460, 1474, 1491, 1509, 1527, 1549, 1561, 1582, 1594, 1610, 1625, 1653, 1673, 1688, 1707, 1726, 1743, 1758, 1772, 1787, 1817, 1834, 1849, 1865, 1881, 1898, 1915, 1929, 1947, 1963, 1976, 2000, 2030, 2046, 2063, 2076, 2097, 2113, 2128, 2144]. PDF [2159, 2174, 2189, 2204, 2225, 2241, 2257, 2271]. Peacock [3265]. Pearl [1591, 147]. Pearlymussels [2472]. Pearsons [2310]. Peces [811]. Peebles [2211]. Peer [1826, 1632, 883]. Peer-Review [1826]. Pelagic [863, 3078, 811]. Pelagicos [811]. Pen [1634, 1747]. Pennsylvania [296, 1084, 2185]. People [1028, 2870, 1492, 891, 2734, 2127, 1127]. Per- [2657]. Perceived [2806]. Perception [1637]. Perceptions [1401, 860, 1904, 2842, 2493, 2807, 2596]. Perch [2861, 146, 714, 872]. Percid [3115, 2261]. Percidae [2857, 2989]. Percina [2955]. Performance [2878, 703, 1634]. Peripheral [943]. Periscope [2147]. Permanent [1747]. Persistent [2907]. Person [2905]. Personal [2658, 2927, 2323]. Personnel [223]. Perspective [2877, 1133, 202, 3266, 1383, 306, 2961, 169, 170, 46, 1829, 1419, 147, 45, 464, 65, 2974, 1418, 2103, 1991, 47, 409, 2118, 1984, 2364, 1803, 167, 684, 166, 2768, 2178]. Perspectives [311, 2508, 2569, 2463, 1937, 2363, 611, 1953, 1176, 1050, 2387, 944, 2962, 3109, 377, 2397, 2424, 1988]. Peruvian [694]. Pet [3150, 2108]. Peter [1940, 3257]. Petroleum [2922]. Petromyzon [2705]. PFAS [2657]. Pflieger [2979]. pH [1760]. Ph.D. [1502]. PHABSIM [2246, 2081, 2248, 2364, 2247]. Pharmaceuticals [1771]. Phenology [1868]. Phillip [1104]. Philosophical [170]. Philosophies [1893]. Philosophy [1171]. Photo [774, 2934, 1914, 1928, 1962, 1975, 1999, 2029, 2045, 2061, 2075, 2096]. Photo-Based [774]. Photo-essay [2934]. Photograph [724, 761]. Photographer [2212, 3189]. Photographic [2165]. Photography [2310, 2496, 2279]. Photos [1694]. Physical [703]. Physiological [1980]. Physiology [851, 1991, 1868, 1346]. Picke [2861]. Picke-Perch [2861]. Pickled [2045]. Piece [1791]. Pieces [3174]. Pietsch [3284]. Pike [2828, 1765, 438, 601]. Pikeminnow [2916]. Pikeperch [2857]. Pikeperches [2989]. Pinder [2662]. Pink [1678, 78, 77]. Pioneer [203, 1321]. Pipeline [3052]. Pipelines [131]. Piper [1271]. Piscicides [913]. Piscivores [1674]. Piscivorous [2510, 1186]. PIT [2472, 2293, 922]. Pitfalls [31, 1696]. PITR [2293]. Place [979, 964, 2373, 2350, 2533]. Place-Based [2373]. Placement [835]. Places [2149]. Plain [2688]. Plainfin [3149]. Plains [2908, 268, 2118, 1057]. Plan [2479, 567, 124, 3018, 429]. Plankton [383]. Planners [478]. Planning [1939, 323, 682, 673, 40, 821, 421, 2100, 384, 2709, 608]. Plans [1970, 2184]. Plant [3026, 1775, 27]. Plant-Based [1775]. Planted [869]. Planting [655]. Plants [55]. Plastic [2567, 1855]. Platforms [2922, 1613]. Plea [2067, 1475, 57]. Pleas [536]. Plenary [1576, 1996, 2200, 2378, 2913, 3032, 3176, 2661]. Plenty [1482]. Plight [2612]. Plot [916, 359]. Plus [950]. Podcast [2811]. Poets [1177]. Point [1773]. Poised [1297]. Poles [1991]. Policies [712, 697, 3065, 71, 1536, 2809, 1199, 665]. Policy [1096, 1602, 16, 1627, 1745, 2130, 2161, 306, 989, 884, 192, 2748, 988, 1043, 269, 246, 247, 1794, 900, 1109, 953, 971, 1012, 1025, 1040, 1082, 1056, 1067, 1889, 222, 245, 2244, 2272, 886, 2919, 2426, 2638, 307, 1131, 1300, 1457, 2323, 2309, 2476, 1000]. Policymakers [2638]. Political [1968]. Politicization [2228]. Politics [2621, 2117, 2301, 1968, 2228]. Pollock [1538, 327]. Polluted [147]. Pollution [117]. Polybags [2669]. Polychlorinated [8]. Polyculture [3]. Polyfluoroalkyl [2657]. Polyodon [80, 3053]. Pond [3217, 223, 301, 315]. Pool [2533]. Poor [2612]. Popular [1780]. Population [2366, 576, 2058, 1267, 1630, 525, 955, 1226, 678, 1143]. Population-Level [1267]. Populations [2414, 707, 2152, 1842, 2656, 930, 2107, 771, 943, 2834, 664, 813, 1923, 386, 1981, 2439, 544, 510, 875, 42, 3079, 1855, 1857, 94, 750]. Port [1648]. Portfolio [2414, 2399]. Portions [947]. Portland [1730, 1932, 1873, 1781, 1904]. Ports [1892]. Pose [3152]. Position [1889]. Positioning [479]. Possible [16]. Post [2878, 3161, 835, 2530]. Post-COVID [3161]. Post-Fact [2530]. Post-Placement [835]. Post-Release [2878]. Postcard [2506]. Poster [2119, 1314]. Postmodern [1719]. Postponed [2777]. Postrestoration [3042]. Postsmolt [1884]. Poststocking [1296]. Potable [143]. Potential [199, 805, 1993, 177, 894, 758, 634, 846, 612, 1308, 2340, 268, 1042, 285, 2767, 940, 728, 2880, 1295, 141]. Potentials [31]. Potomac [958, 1793]. Pound [286]. Pound-Net [286]. Power [213, 177, 1850, 2275, 3072, 27, 55, 1370]. Practical [596, 406, 538]. Practice [1172, 1633, 2611]. Practices [712, 192, 2463, 1761, 933, 2894, 2822, 230, 1598, 111]. Practitioner [2364]. Prawn [1432, 1241]. Precautionary [583]. Precedent [2106]. Precipitation [124, 191]. Predation [2923, 2766]. Predator [1296, 2270]. Predators [568, 1746]. Predatory [746, 3201]. Predict [1097]. Predictable [2747]. Predicted [3029]. Predicting [257, 1310, 1712]. Preferences [2300, 667]. Preliminary [586]. Prelude [1281]. Premature [2991]. Prepare [2006]. Preparing [1311, 1324, 1336, 1392, 1237, 2015, 2001, 2009, 2013, 2004, 2014]. Presence [1712]. Present [1684, 2960, 2632]. Presentation [2313, 1314]. Preservation [1308]. Preserving [1714]. President [1126, 1668, 1901, 2975, 2976, 1902, 2154, 2977, 1139, 1140, 1331, 1320, 1690, 1691, 1933, 2468, 2469, 3283, 1941, 2824, 1480, 968, 949, 2644, 3129, 2205, 3282, 2153, 2314, 3130, 2825, 984, 2315, 3131, 2645, 1934, 3281]. Prespawn [2083]. Press [3267, 2185, 2662]. Pressure [58]. Pressures [1386, 1216]. Prevalence [1844]. Prevent [1428, 1890]. Prevention [1844]. Preview [2155, 2341, 2523, 2673, 2993]. Prey [103, 337, 2510, 1186]. Prey-Dependent [103]. Priceless [2516]. Pricing [2516]. Primacy [2295]. Primer [1171, 954, 2689]. Primitive [2231]. Principles [1679]. Print [2955]. Printing [2973]. Prior [730]. Priorities [222]. Prioritize [3164]. Prioritizing [2414, 2108]. Priority [410, 2272]. Private [2339, 1334, 1245, 255, 47, 744]. Privatized [519]. pro [2032]. Proactive [416, 843]. Probiotics [1775]. Problem [894, 2641, 409, 886, 2571]. Problem-solving [409]. Problematic [2374]. Problems [930, 2136, 2907, 809]. Procedural [449]. Procedure [55]. Procedures [2637, 286, 816, 3164]. Process [1435, 202, 819, 2975, 2082, 1826, 1145, 507, 883, 1314]. Processes [522, 954, 2484]. Processing [3278, 163, 1922]. Produces [60]. Production [2921, 157, 486, 550, 606, 460, 513, 301, 476, 433, 1837, 2365, 932, 2812]. Productive [2251]. Productivity [282, 342]. Products [1922]. Profession [2545, 1319, 2550, 2549, 3271, 3008]. Professional [833, 1802, 1720, 3161, 1930, 1298, 1066, 1229, 1155, 2363, 1762, 3261, 2006, 3139, 2893, 1790, 2700, 1005, 1062, 132, 2751, 259]. Professionalism [139]. Professionals [381, 1031, 1377, 1987, 2015, 554, 3171, 2001, 2295, 1416, 3224, 2017, 2006, 2660, 2013, 2004, 1715, 1035, 403, 2894, 1074, 984, 1968, 1146, 2325, 2014, 3019, 1143]. Professions [3100, 1231]. Proficiencies [937]. Program [1019, 1190, 1352, 1543, 1753, 1802, 1959, 2656, 2425, 334, 1777, 860, 671, 933, 2006, 2552, 1226, 2736, 384, 495, 2932, 3163, 2553, 2020, 643, 1825, 2087, 2545, 2723, 2763, 3266, 1212, 1738, 3075, 98, 191, 92, 2551, 2868, 173, 1007, 2899]. Programmatic [531]. Programming [3277, 2079, 75]. Programs [1533, 602, 257, 2386, 1721, 571, 2295, 722, 596, 1953, 209, 561, 562, 829, 279, 538, 651, 183, 188, 1428, 389, 1199, 363, 2768, 2751, 1844, 1679]. Progress [1910, 3264, 1630, 1734, 2844]. Prohibited [2168]. Project [2108, 2452, 2856, 175, 3077, 287, 3173]. Projects [339, 723, 2820, 711, 829, 442, 82]. Promise [1775]. Promote [2545, 2295, 2700, 1083, 867]. Promotes [2710]. Proof [2182]. Propagation [933, 3029, 1639]. Properties [260, 2892]. Property [529]. Proportion [511]. Proportional [802]. Proposal [599, 2150]. Proposed [2209, 2210, 2451]. Proposition [1511]. Prospects [86, 1415, 738]. Prospectus [873]. Protect [856, 2697, 665]. Protected [688, 863, 993, 700]. Protecting [819, 1253]. Protection [909, 707, 1883, 1431, 355, 556, 1766, 2856]. Protection/Restoration [1883]. Protections [2867]. Protein [340]. Protocol [2821]. Protocols [1844]. Prove [2299]. Provide [3099, 3126, 2907, 2864]. Provides [1020, 2100, 2360]. Providing [1986]. Provincial [816, 593, 448]. Provisions [1766]. Psyche [2350]. Public [712, 637, 1637, 1043, 246, 247, 3277, 1583, 245, 1334, 2350, 1302, 2629, 2646, 2036, 764, 893, 570, 2337, 2102, 1647, 2365]. Publication [2574]. Publications [2574]. Publish [1386]. Published [2132, 3284, 3139, 3267]. Publishing [2846, 2934]. Puget [3284]. Pulsed [413]. Pupfish [3106, 870]. Purpose [1700, 39]. Purposes [2365]. Pursuant [738]. Pursuing [2115, 1500]. Pursuit [1298, 830]. Push [2722]. Pushing [2531]. Put [2049, 2246, 2081, 2248, 590]. Put-and-Take [590]. Putman [3185]. Puts [1803]. Putting [2871, 1630].

Q&A [1126, 1604, 1567, 2213, 1933, 1941, 1684, 1733, 1841, 1764, 2212, 1796, 1568, 1732, 1763, 1955, 1969, 1716, 1589, 1924, 1934]. Quadra [2659]. Qualifying [1314]. Qualitative [2842]. Quality [1645, 1783, 1993, 1537, 439, 2509, 2660, 342, 1497, 466, 28, 2697]. Quantifying [2779, 724, 761, 2518]. Quantitative [1842, 836, 2842, 654]. Quantity [2908]. Québec [1658, 1665, 212, 1406, 1455, 1421, 1438, 1505, 1486, 1557]. Queen [792]. Queensland [798, 891]. Question [2003]. Questions [1045, 2297, 1448, 1370, 1994].

R [1147, 1542, 2729, 1270, 3284, 2265, 2153, 2185, 1969, 3007, 2654, 2264, 2599, 2293, 2129, 2198, 3212]. R-Based [2264]. R. [1768]. R/V [2729]. Rachel [1587]. Radiocarbon [2597]. Rafinesque [2857, 1347]. Rage [2701]. Ragfish [2527]. Rainbow [3217, 3223, 2878, 2413, 226]. Raise [2036]. Raises [1647]. Raising [3, 1788, 1647]. Raja [2232]. Ranaviruses [1868]. Random [1649]. Range [822, 779, 78, 5]. Ranks [1177]. Rapid [2585, 1157, 613]. Rapids [2680, 3239, 3235, 3103, 3123]. Rare [2628, 447, 1216, 2203, 2655]. Rassam [1142]. Rate [300]. Rates [330, 2701, 1675, 2151, 2104, 2362]. Rather [155]. Rating [261, 2205]. Rationale [3202]. Ratios [870, 586]. Raton [168]. Raw [2750]. Ray [1539]. Rays [2883]. Razorback [2902]. Re [1957, 194, 2450, 744]. Re-Creational [2450]. Re-envisioning [1957]. re-evaluation [744]. Re-Registration [194]. Reach [2531, 2260, 706]. Reaches [1859]. Reaching [1728, 2656]. Reactions [2584]. reader [649]. Readiness [3226]. real [750]. Realistic [1333]. Reality [849, 2491]. Really [2559, 3271, 2021, 913]. Reappraisal [313]. Reared [1721]. Reasons [2140, 2308, 3123, 2504, 2670]. Reassessment [818]. Reauthorization [2427, 2409]. Recalibration [2312]. Recap [1617, 3235]. Recapture [922]. Receive [3236]. Recently [1031]. Recipient [1744, 1660]. Reclaiming [2350]. Reclamation [147]. Recognition [1802, 1296, 894, 188, 2182]. Recolonization [1939]. Recommendation [3027]. Recommendations [1206, 223, 63, 816, 1949, 649, 744]. Reconcile [1312]. Reconnecting [1923]. Record [1694, 3053, 2164, 2991, 1100]. record-breaking [1100]. Recording [224, 25]. Records [2165]. Recovering [738, 2864, 3002]. Recovery [833, 2992, 3263, 821, 1677, 2570, 2835, 1970, 746, 541, 728, 3174, 3018, 577, 2869]. Recreating [3148]. Recreation [7]. Recreational [1172, 1171, 199, 1553, 806, 2748, 627, 2779, 551, 1265, 3074, 2401, 1695, 2866, 1562, 1982, 987, 1089, 2085, 411, 2819, 432, 2809, 621, 2737, 598, 1956, 1820, 1598, 798, 891, 2450, 2040, 668, 2178]. Recruiting [1146]. Recruitment [103, 3139, 984, 1837]. Recycling [3087]. Red [1125, 624, 1041, 1761, 1675, 1794, 1630, 2352, 435]. Redfish [1630]. Redmond [3102]. Reduce [121, 649]. Reduces [2698]. Reducing [169, 170, 2629, 2646, 1252, 167, 166]. Reduction [294]. Reductionism [1461]. Reece [1714]. Reef [2537, 486, 558, 842, 1761, 236]. Reefs [461, 460, 619, 835]. Reeling [2313]. Reestablishment [1565]. Reevaluate [3201]. Reexamination [822]. Reference [51]. References [815]. Reflections [2435, 2598, 239, 2893]. Reflector [2029]. Reform [737, 2736, 1384]. Refuge [1761]. Refuge-Seeking [1761]. Regaining [2867]. Regarding [989, 1640, 631]. Regime [2972, 2299]. Regimens [7]. Regimes [513]. Region [179, 3029, 3188, 2832, 1295, 1555]. Regional [180, 2243, 2970, 460, 3065, 1097, 219, 886, 3164, 398, 264, 401]. Regionally [1144]. Regions [2251, 3019]. Registration [3116, 2858, 194]. Regulate [168]. Regulated [1677]. Regulated-River [1677]. Regulation [2921, 3064, 2082, 155, 2387, 2312, 767]. Regulations [3150, 2532, 1397, 411, 3151, 414, 1646, 2584, 2948, 767, 601, 598, 1893, 2797, 2106]. Regulatory [819, 246, 2702, 507, 1361, 2747]. Rehabilitating [123]. Rehabilitation [1638, 1469]. Rehabilitations [1737]. Reintroduction [1485]. Rejoinder [1433]. Relatability [2493]. Related [767, 7, 857]. Relates [2780]. Relating [1697]. Relationships [1066, 1229, 42, 536, 2493]. Relative [2878, 410, 274, 264]. Relatives [557]. Relativity [1051]. Release [989, 2878, 2671, 1761, 1675, 3087, 1267, 2168, 2519, 562, 2465, 1096]. Released [867]. Releases [3059, 2573]. Relevance [1171, 2921, 3261]. Relevancy [2006]. Relevant [1264, 2672, 1858]. Reliance [1125]. Relief [549]. Remain [2521]. Remaining [2251]. Remains [2672]. Remarkable [1925]. Remedial [421]. Remediation [1469]. Remedies [1759]. Remembering [2112]. Remission [1199]. Remote [2250, 1767, 922, 216, 786]. Removal [1939, 1586, 2621, 2600, 746, 776, 2638]. Removals [2108]. Remove [1820]. Removing [3008]. Renewal [3085]. Renewed [1709]. Renewing [2411]. Reno [2446, 2456, 2613, 2523, 2504]. Reopen [2880]. Repeated [359]. Repeated-Measures [359]. Replace [2374]. Replacement [2880, 704]. Replenishing [1319]. Reply [876, 1501]. Report [1631, 1206, 1641, 1439, 1904, 2100, 1060, 1588, 1813, 1591]. Reporting [3202, 63]. Reports [1407, 25]. Repository [2377, 3043]. Represent [101]. Representative [706]. Representing [802]. Reproducible [2750]. Reproduction [2639, 5]. Reproductive [1057]. Republic [3151]. Require [2557, 2558, 2555, 2556, 2554]. Required [2348]. Requirements [2007, 63]. Reregistration [634]. Rescue [323]. Research [220, 1533, 367, 1115, 2537, 2213, 1658, 819, 486, 2386, 567, 714, 2167, 783, 2671, 2748, 659, 1765, 988, 1298, 491, 576, 1801, 860, 1011, 2009, 411, 1349, 222, 2574, 2923, 1810, 2971, 261, 829, 1599, 2688, 1333, 409, 1074, 1696, 473, 2611, 2962, 3005, 495, 2960, 124, 191, 49, 3163, 168, 1083, 92, 1843, 429, 2768, 1295, 2385, 1111, 1540, 1600]. Research-based [495]. Research/Management [411]. Researchers [717, 1389, 278, 1810, 2822]. Researching [1346]. Reservation [1351, 1903]. Reserve [544]. Reserves [624, 684]. Reservoir [623, 3200, 1736, 1737, 632]. Reservoirs [920, 3238, 632]. Resident [691]. Residents [478]. Resilience [2573, 2596, 1920]. Resiliency [1051]. Resilient [2299]. Resist [3017, 2794]. Resolution [1018]. Resolving [486]. Resource [1206, 543, 2243, 1110, 3226, 449, 155, 198, 2808, 1461, 2400, 2702, 2640, 14, 219, 1231, 1995, 2807, 66, 2595, 704, 2228]. Resources [326, 492, 1978, 1499, 303, 2134, 177, 366, 40, 1615, 18, 2311, 842, 320, 1635, 2455, 369, 172, 282, 141, 1601, 1557, 2118, 1949, 2099, 531, 2263, 564, 7, 1132, 1325, 8, 2292, 1109]. Responding [3251, 2794]. Responds [1792]. Response [1631, 1791, 2843, 296, 1993, 1500, 848, 1636, 1632, 1051, 2948, 2248, 2989]. Responses [1586, 875, 3042, 776, 467]. Responsibility [1341, 884]. Responsible [2961, 1369, 2700, 1966, 2178]. Restoration [801, 923, 1910, 1883, 3251, 1909, 413, 812, 1198, 734, 2295, 646, 508, 1957, 2621, 464, 38, 342, 2600, 829, 800, 944, 759, 1722, 2452, 463, 2932, 478, 1920, 481, 2844, 520, 1779, 372, 608]. Restore [2339, 331]. Restored [1655, 2795].

Restoring [2983, 522, 1894, 2439, 356]. Restrictions [797, 668]. Result [752]. Results [381, 1939, 2363, 1098, 377]. Resurrecting [2971]. Retention [590, 2958, 1144]. Retreat [3154]. Retrieval [1309]. Retrospective [674, 1432, 1956, 1241]. Return [2639, 2084, 2467, 2905, 1820, 520]. Returning [2795]. Reuse [1308]. Reveal [2855, 1780, 2958, 1868]. Reveals [2170, 1922]. Reversed [2971]. Review [1522, 3284, 3064, 1540, 688, 771, 681, 460, 1719, 836, 1826, 500, 200, 1015, 1469, 1470, 816, 141, 1539, 1718, 3267, 1098, 978, 929, 873, 678, 2185, 2362, 168, 3018, 66, 1897, 2662, 2856, 2586, 1717, 883, 1117, 1702]. Reviewers [1710]. Reviewing [2571]. Reviews [2296, 988, 2846, 384]. Revised [1600]. Revision [2479]. Revisions [2867]. Revisited [1583, 864]. Revisiting [1655, 1883, 2161, 1596, 1314]. Revival [1389]. Reviving [2865]. Revolution [314]. Reward [2277]. Rhetoric [2491]. Ribbon [2049, 1836]. Rice [1147]. Richard [1668]. Richness [1030]. Rick [1345]. Right [1595, 1852, 3174]. Rights [1461, 529]. Rigorous [1970]. Rim [1214]. RIP [1513]. Riparian [375, 464, 2399]. Ripples [2930]. Rising [2677, 334, 198]. Risk [638, 2170, 1532, 3152, 646, 826, 755, 546, 581, 556, 273, 1014, 290, 2277]. Risks [1172, 169, 170, 876, 943, 881, 596, 3175, 538, 849, 167, 166]. Risky [2856]. River [2230, 2418, 2375, 1783, 2167, 3264, 689, 189, 1027, 205, 2834, 511, 1586, 1485, 1677, 730, 233, 508, 2058, 1212, 2621, 3065, 2340, 829, 1097, 3238, 3101, 2439, 1599, 1343, 2229, 1098, 42, 2452, 2164, 1938, 3163, 1069, 691, 1227, 1859, 608, 2639, 1939, 1018, 1353, 323, 482, 851, 329, 212, 606, 147, 646, 187, 227, 333, 1212, 2437, 525, 2039, 3173, 433, 2640, 1226, 2056, 435, 746, 541, 2066, 2736, 1085, 2299, 1793, 2764, 2263, 1731, 1415, 577, 2766, 481, 1555, 2844, 864, 520, 3153, 372, 2302]. Riverine [1198, 1808]. Rivers [2177, 2227, 1043, 2531, 1186, 1212, 1923, 3280, 1894, 2118, 2880, 2734, 2121, 574]. Riverscape [2908, 1214]. Road [1484, 1049, 2244, 141]. Road-Stream [1484]. Roadmap [1615]. Roads [718, 331]. Roanoke [3048]. Roar [1045]. Robb [2786]. Robert [1374, 2185, 3186, 1139, 1420, 1502, 2798, 2138, 2221, 2739]. Robots [910]. Rock [1242, 1256, 1269, 1351, 1006, 1141, 1311, 1324, 1336, 1449]. Rockfish [559, 1717, 1706]. Rockfishes [3055, 1718]. Rocky [842, 261, 1281]. Role [794, 1493, 185, 3161, 1184, 1196, 2494, 253, 1644, 154, 2272, 442, 211, 935, 764, 893, 2021, 1005, 1062, 1857, 1132, 616]. Roles [1689, 647, 49]. Romance [1701]. Rome [1966]. Ron [2479]. Ronald [1140, 2755]. Ronde [2844]. Rondegat [1555]. Roots [2630, 2642, 2749, 2713, 2726]. Rose [1243, 1744]. Rosenlund [3268]. Rotenone [1258, 565, 732, 194, 864]. Rough [1435, 2608, 2933]. Roulson [2315]. Round [1985]. Round-the-Coast [1985]. route [689]. Routine [780]. Rule [2451]. Ruling [2349]. RUNNER [1388, 1389]. Running [1894, 1372]. Runs [342, 154]. Ryan [3267, 2793]. Ryder [1668, 1793].

S [1374, 1540, 3209, 2480]. Sacraments [1177]. Sacred [1583]. Sadly [1852, 1747]. Safe [1986, 3127]. Safety [2145, 3015]. Sage [1867]. Said [3138]. Sail [88]. Saila [2330]. Saint [1127, 1173, 1164, 212]. Saint-Laurent [212]. Salamanderfish [238]. salamandroides [238]. Salaries [400]. Salary [1334]. Salient [2137]. Salish [3284, 1403]. Salmo [3009]. Salmon [822, 1939, 1631, 2605, 3258, 2674, 444, 712, 2921, 2152, 586, 3251, 2083, 709, 567, 540, 689, 851, 1076, 2568, 574, 329, 1266, 1721, 2992, 3188, 1485, 2373, 2137, 821, 550, 2578, 513, 2570, 1838, 646, 595, 200, 1632, 233, 293, 779, 1630, 342, 78, 386, 433, 2377, 901, 2192, 2198, 570, 635, 1599, 273, 541, 651, 2298, 873, 1633, 2737, 3009, 473, 2676, 229, 1775, 2251, 827, 869, 445, 356, 549, 2573, 2766, 262, 77, 123, 1214, 2795, 2127, 553, 2991, 329]. Salmonid [522, 682, 2910, 1124, 780, 476, 261, 49, 520, 824, 932]. Salmonids [287, 3264, 212, 423, 2234, 581, 849, 533, 3138, 111, 2572]. Salmonines [211]. Saltwater [235, 700]. Salvelinus [2405, 2815]. Sam [3189]. Same [2621, 2195]. Sample [2107, 641]. Sampling [1445, 626, 2135, 286, 2531, 1884, 671, 1892, 674, 510, 1693, 1837, 3201, 2219, 1332, 936]. Samuel [1347]. San [1577, 811, 822, 811]. Sanctuary [1729]. Sand [2168, 507]. Sandbar [1692]. Sander [2514, 2989]. Sandodden [1045]. Sara [1764]. Sardine [2752]. Satellite [2250]. Satellites [910]. Satisfaction [2780]. Saturday [1959]. Sauger [1115, 2857, 2989]. Saugeye [3201]. Saul [2330]. Save [2627]. Saving [595, 1630]. Savvy [955]. Sawfish [1694, 1907, 2164, 94]. Sawfishes [2165]. Say [2945, 1864]. Says [2506]. Scale [2339, 177, 881, 2855, 1070, 2819, 2600, 2898, 829, 2962, 2452, 2930, 1598, 1564, 1026, 278, 830, 795]. Scales [923, 100, 3065, 1297, 749]. Scaling [1578]. Scan [862, 919]. Scaphirhynchus [2650, 107, 1346]. Scary [1412]. Scenes [1560]. Scenic [2418]. Schedule [1574, 2169, 2355, 1781, 1275, 2511]. Scheduled [1180]. Scheerer [1796]. Scholars [3266, 2763, 2937, 3098, 2022]. School [2845, 2723, 1160, 216]. Schools [129]. Science [1641, 1569, 1609, 2870, 2131, 1422, 1523, 1936, 2280, 1627, 1709, 2637, 2199, 3099, 1390, 2324, 2338, 2687, 1076, 2633, 2823, 3016, 2018, 1615, 439, 821, 3003, 1777, 393, 2532, 2990, 2598, 1024, 3224, 1391, 454, 1654, 1177, 612, 2521, 3100, 3139, 2039, 2519, 1050, 2176, 3076, 1601, 2242, 2629, 2646, 2198, 2041, 935, 2117, 1231, 2842, 2894, 1680, 1240, 3072, 1465, 1853, 759, 2195, 1766, 2919, 2896, 2493, 2676, 249, 2325, 2934, 1757, 1905, 2551, 3252, 1716, 2530, 1297, 1191, 1227, 2034, 2020, 2206, 2409, 2697, 2899, 2147, 2302, 1568]. Science-Based [2532, 2521, 2409]. Science-Policy [1627]. Sciences [3031, 2723, 2328, 2853, 2689, 1070, 1429, 378, 1904, 2768]. Scientific [2891, 1533, 2890, 1358, 1215, 1386, 1528, 101, 2846, 857, 2907, 2427, 2822, 1337, 3237, 1098, 1697, 2193, 226, 1213, 2247, 520, 510]. Scientifically [2374, 1970]. Scientist [1434, 393, 3062, 2792, 3267, 2436, 640]. Scientists [879, 878, 2329, 366, 808, 611, 2893, 2242, 2762, 3079, 3109, 1968, 132, 2807, 2454]. Scope [519]. Scorpionfishes [1718]. Scotia [123]. Scott [2479]. Screen [703]. Script [2599]. Scuppered [1384]. Sea [2705, 3284, 2170, 1798, 2715, 1800, 3073, 1884, 3126, 3082, 1348, 2173, 2766, 2874, 327, 3149, 532, 1404, 1403, 1314, 2170]. Seafood [3066, 2855, 3278, 1922, 1239, 2421]. Seahorses [2105, 2106]. Sean [1914, 2868]. Search [1945, 2105, 2181]. Searches [3222]. Searching [830]. Season [1885, 3085, 1288, 797]. Seaspiracy [2990]. Seattle [1112, 216, 979, 1100, 964, 1046, 1102]. Seaweeds [1681]. Second [1086, 1256, 1421, 1619, 1845, 1690, 1691, 2051, 2255, 2588, 3098, 1901, 1933, 2468, 2975, 2469, 2976, 3283, 1941, 2824, 2644, 3129, 3282, 2153, 2314, 3130, 2825, 2315, 3131, 2645, 2586, 1902, 1934, 2154, 2977, 3281, 1139, 1140, 1320, 1480, 1417]. Section [1283, 1159, 1200, 1454, 1664, 2126, 1188, 2019, 1497, 1991, 1286, 2101, 2324, 2038, 1349, 2013]. Sections [179]. Sector [1334, 1245, 764, 47]. Sectoral [1467]. Security [1467]. Sedative [989]. Sedatives [1096]. Sedell [1270]. See [1935, 1848, 1173]. Seed [2735]. Seeing [3276, 785, 2394]. Seek [1554]. Seeking [1761, 1429, 1737, 2195]. Seeks [1662, 1131, 1457]. Segregation [200]. Selected [723, 488]. Selecting [411]. Selection [140]. Selective [1532]. Seminar [3016]. Senescence [1736]. Sensationalism [2990]. Sense [1932, 3003, 2350]. Sensing [2250]. Sensitive [1986, 475]. Sensors [2923]. Separation [1018]. September [1112, 1351, 2446, 2456, 2673, 3102, 2935, 3199, 1079, 1260, 1248, 1274, 1458, 1423, 1650, 1846, 2393, 2564, 2731, 2888, 3049, 3194]. Sequencing [1349]. Series [3016, 2518, 1070, 1954, 1914, 1928, 1962, 1975, 1999, 2029, 2045, 2061, 2075, 2096, 2231]. Seriously [1950]. Servant [2668]. Service [288, 1298, 133, 1187, 141, 2322, 2372, 2022, 191, 429, 192, 334, 2400, 175, 253, 866, 1174, 2780, 3029, 3018, 2397, 2396, 2395]. Services [519, 2466, 2475, 1702, 2484]. Serving [1075]. Session [1904]. Setting [857, 2106]. Seventh [1337]. Seventy [2715, 1956]. Seventy-Year [2715, 1956]. Sex [1203, 1348, 1856]. Shad [1158, 1157, 589, 749]. Shade [3026]. Shale [1295]. Shall [1529, 2947, 3030]. Shape [2582]. Shaped [1657]. Shapes [2476]. Shaping [2194, 2735]. Sharber [1571]. Share [1569, 1001]. Shared [2274, 2192]. Sharing [2463, 1308, 2438, 2530]. Shark [1842, 2483, 1598, 94, 3225, 750]. Sharks [1436, 1692, 2361, 2168, 557, 1803, 2883]. Sharpening [2161]. she/her [3288]. Sheds [2011]. Shelf [3005]. Shell [1742]. Shellfish [213, 1681, 765, 1325]. Shellfisheries [1823]. Shelton [2963]. Shepard [2935]. Sheraton [1959]. Shift [2933]. Shifting [2897]. Shifts [2130, 1692]. Shiner [2381]. Shining [3235]. Ship [2581]. Ships [1799]. Shoal [2735]. Shooter [3248]. Shooting [846]. Short [1375]. Shortage [879, 878]. Shortened [2866]. Shortnose [930]. Should [1445, 954, 2308, 2359, 1853, 1708, 1693, 2325, 2831, 2504, 2102, 1372]. Show [2089, 1775]. Shows [3173]. Shrimp [1678, 295, 294, 1647]. Shrinking [1344]. Side [2145, 862, 919, 2021, 2195]. Sideways [1925]. Sight [2688]. Sign [1250]. Significance [1835, 641]. Signs [2744]. Silos [2907, 2198]. Silt [1774]. Silver [3045, 3232, 3134, 1894]. Similarities [946]. Simile [2778]. Simon [2139]. Simple [1871, 1097, 2920, 1837]. Simpson [2479]. Simulated [433]. Simulation [1484]. Simulator [1683]. Singing [3149]. Single [101]. Site [1215]. Situ [122]. Situational [1462]. Situations [2633]. SIUC [1298]. Size [185, 2107, 135, 678, 1956, 2833, 641, 668]. Size-based [2833].

Size-Structure [1956]. Sizes [2878, 1761]. Skates [2883]. Sketch [1786]. Skill [2208]. Skills [2654, 2010, 391, 2751]. Skipjack [98]. Slippery [1665]. Slope [1665]. Small [2152, 624, 2855, 278, 3065, 2104, 830, 2821, 71, 2962, 2534, 786, 744]. Small- [278]. Small-Scale [2855, 2962, 830]. Smallmouth [1793, 2764, 3174, 2940, 753]. Smartphone [2085]. Smartphones [1416, 1478, 1780]. Smelt [3036, 2301, 1793]. Smith [1782]. Snack [3094]. Snag [157]. Snagging [329]. Snake [574, 3153, 433, 541, 1085]. Snakehead [2278, 2480]. Snapper [1125, 558, 1041, 1761, 1675, 1794]. Snapshots [1985]. Snot [1006]. Snout [3179]. Snowy [2597]. Social [1515, 1569, 1553, 2686, 1390, 1066, 1229, 1211, 1196, 2327, 612, 3100, 1335, 1644, 1393, 3054, 2326, 1465, 1479, 3109, 662, 2325, 2454, 2734, 2040, 2034]. Social-Ecological [2040]. Societal [1759]. Societally [2374]. Societies [2891, 2980, 1171, 2890, 3161, 1762, 1005, 1062]. Society [794, 907, 1190, 1268, 1813, 1825, 2037, 2051, 2070, 2169, 2239, 2109, 2255, 2446, 2511, 2613, 2456, 2673, 2754, 2770, 2952, 3091, 1155, 3129, 1790, 2733, 1626, 2504, 1362, 961, 976, 1128, 1129, 1284, 1380, 1454, 1453, 1664, 1603, 1619, 1629, 1656, 1663, 1874, 1832, 1845, 1860, 1870, 1876, 1802, 2090, 2072, 2086, 2088, 2200, 2282, 2283, 2442, 2443, 2613, 2615, 2616, 2479, 2783, 3239, 3242, 2721, 1390, 1313, 2363, 1826, 895, 2313, 2006, 2574, 866, 2209, 2547, 546, 2210, 2195, 2504, 1300, 2724, 218, 3153, 899, 970]. Society [997, 1048, 1103, 1134, 1179, 1221, 1274, 1315, 1363, 1423, 1487, 1544, 1607, 1650, 1670, 1723, 1755, 1784, 1846, 1895, 1943, 1997, 2073, 2125, 2187, 2254, 2318, 2368]. Sociocultural [79]. Socioeconomic [848, 2401, 1562]. Sociological [277]. Sociopolitical [2199]. Sockeye [689, 851, 2578, 433, 1599, 2766]. Soft [1855]. Softcover [2662]. Software [318, 1535, 2767, 2182, 2042]. Solar [2880]. Solution [2175, 2889, 2180]. Solutions [2220, 1517, 3126, 1775]. Solving [2641, 409]. Some [486, 101, 536, 3198, 90, 441, 216, 517]. Something [2969]. Sometimes [2277]. Sonar [862, 919, 3110]. Sonic [2170]. Sordid [1791]. SOS [3127]. Sound [1422, 3284, 502, 1024, 2482]. Sounds [679, 3252]. Soundscapes [2482]. Source [2293, 1535, 3075, 2932]. Sourced [2400]. Sources [3072, 340, 616]. South [822, 1041, 2483, 476, 3074, 1228, 294, 2610, 2534, 1555]. Southeast [1349, 295]. Southeastern [444, 157, 1332, 109, 930, 511, 1187, 728]. Southern [822, 2062, 1389, 1801, 779, 1678, 1389, 320, 740, 268, 1694, 632, 579]. Southwest [862, 3152, 268]. Southwestern [1586, 747, 233]. Space [2559, 2350, 2920, 3127]. span [651]. Spanish [2712, 2494]. Sparking [1899]. spathula [80, 3053]. Spatial [923, 707, 2907, 1554, 514, 695, 498, 2832, 1069]. Spatially [300, 1322, 2040]. Spatially-explicit [300]. Spatiotemporal [1297]. Spawn [1085]. Spawned [1404]. Spawners [869]. Spawning [1721, 613, 1868]. Speak [1444, 2947]. Speaker [1576, 1996, 1773]. Speakers [2200, 2378, 2913, 3032, 3176, 2661]. Spearfishing [2698]. Special [2543, 2190, 39, 2427, 1904, 1979, 51, 2396, 242, 1143]. Specialization [240].

Specialized [2907]. Species [2750, 1584, 638, 659, 1986, 1030, 1058, 758, 2585, 2483, 756, 3263, 821, 1838, 1158, 826, 1892, 2058, 741, 1585, 3078, 2036, 3240, 886, 1697, 1989, 1428, 475, 917, 1890, 1760, 1793, 926, 2203, 643, 2426, 3029, 3041, 2276, 1844, 738, 755]. Specifically [2247]. Specimen [2164]. Spencer [1605]. Spending [722]. Spider [14]. Spill [956, 1638, 1794, 1089, 1042, 1702]. Spiny [68, 65, 66, 69]. Split [359]. Split-Plot [359]. Spokane [3020, 2986]. Sponsored [448]. Sponsors [1575, 3176]. Spontaneously [1404]. Sport [323, 747, 291, 160, 780, 206, 767, 654, 452, 2932, 276, 1131, 1300]. Sport-fisheries [452]. Sportfishing [2974]. Spot [2232]. Spotlight [1163, 1405]. Spotted [2224, 1741, 2874, 859]. Spread [470, 743, 1252]. Spring [1548, 2062, 541]. Springs [2344]. St [1180, 1268]. Stabilizing [331]. Stable [586]. Staff [3272, 3288, 2387, 3257, 2260, 1812, 1005, 1062]. Staffing [722]. Stages [2522]. Stakeholder [2069, 2640, 935, 2387, 2517, 2493]. Stakeholder-Centered [935]. Stakeholder-Driven [2387]. Stakeholders [2735]. Stakes [1239]. Stand [1713]. Standard [626, 1871, 2135, 2637, 2522, 612, 1908, 2610]. Standardization [2599]. Standardize [1124]. Standardized [3064, 671, 2819, 827, 3201]. Standards [63]. Star [2562]. Starting [1465]. State [2388, 1643, 1243, 1577, 326, 223, 682, 2386, 2508, 3284, 2463, 2082, 3150, 571, 722, 1313, 1957, 1475, 139, 972, 816, 410, 2184, 71, 593, 531, 2807, 2102, 1227, 448, 744, 2339, 1002, 46, 937, 182, 737, 1642, 194, 492]. State-Level [2386]. State/Federal [194]. Statement [1139, 1140, 1690, 1691, 2891, 1901, 989, 2976, 3283, 2824, 2644, 3282, 2153, 2314, 3130, 2825, 2315, 3131, 2645, 1902, 2977, 3153, 2890, 2154, 1096]. Statements [2468, 2469]. States [3271, 279, 330, 326, 492, 257, 470, 682, 80, 2508, 1645, 747, 109, 930, 771, 571, 734, 511, 657, 2057, 836, 3112, 133, 2349, 3224, 2011, 1360, 464, 209, 3065, 1908, 1334, 1644, 351, 1187, 2197, 507, 414, 535, 764, 570, 635, 298, 2765, 514, 797, 929, 183, 188, 290, 1634, 2807, 809, 124, 1199, 2806, 111, 427, 818, 2102, 353, 579, 345]. Statewide [3140]. Station [2508, 1648, 175, 592]. Statistical [644, 1370]. Status [196, 444, 2386, 339, 213, 930, 771, 571, 2549, 1518, 2867, 86, 854, 1360, 107, 3151, 816, 274, 1343, 188, 230, 276, 445, 427, 818, 353, 579, 327, 448, 345]. Stauffer [2185]. Stay [1004, 1164]. Staying [2476]. Steady [1475, 3271]. Steady-State [1475]. Steelhead [1939, 444, 2377, 1085, 2053, 2773]. Steller [2766]. Stem [888, 2958]. Stennis [175]. Step [1595, 364]. Stepping [2961, 1372]. Steps [1435, 2961, 1375, 2920, 1966]. Sterile [780]. Steve [2479]. Steven [1662, 1661, 1875, 1660, 1087, 1257, 1282, 1451, 1452, 2055, 2252, 2428, 2587, 2753, 2912, 3056, 3211, 3185, 1659, 2709]. Stevens [2427, 1646, 2409, 154]. Stewardship [955, 1995, 2397]. Stewart [1104]. Stickleback [664]. Sticklebacks [665]. Sticks [1873]. Still [2101, 1264, 1820]. Stipends [1199]. Stizostedion [2857]. Stock [686, 879, 878, 50, 502, 880, 1856, 2905, 380, 90, 69, 343]. Stocked [1674]. Stocking [257, 836, 596, 933, 888, 71, 538, 902, 3213, 111, 744]. Stocks [2412, 916, 423, 1990, 581, 556, 273, 589]. Stoking [1273]. Stone [1852]. Stories [2438, 2869]. Storm [1144]. Story [1623, 1271, 1538, 2116, 1796, 2764, 2203]. Strain [2451]. Strains [20]. Strait [2249, 2921]. Straits [3284]. Strangers [2439]. Strategic [1070, 1788, 2479]. Strategies [169, 2078, 170, 821, 1675, 114, 525, 433, 2895, 1252, 167, 166, 373, 2832, 1779]. Strategy [1499, 287, 551, 500, 2234, 2397]. Stratify [1837]. Straying [2152, 646, 2573]. Stream [375, 1910, 1909, 954, 413, 1198, 511, 657, 720, 1484, 336, 200, 232, 464, 925, 149, 378, 211, 1216, 497, 82, 2266, 279, 1057, 42, 3042, 762, 463, 3213, 869, 2042, 706]. Stream-Dwelling [954, 200]. Streambank [812]. Streamflow [3028]. Streaming [2147]. Streamlining [1435]. Streams [157, 296, 2466, 500, 120, 304, 862, 919, 2821, 2439, 97, 800, 3198, 1332, 466, 1779, 3140, 126, 786]. StreamThermal [2042]. Street [2995]. Strengthening [602]. Stress [1803]. Strike [914, 3184]. Striped [623, 2948, 728, 398]. Stripeless [3289]. Striper [2591, 3289]. Strong [1081]. Stronger [2199]. Structure [686, 202, 707, 209, 1657, 531, 1956]. Structured [987, 2600, 1450, 2040]. Structures [1597, 1499, 3043, 632, 768]. Student [1299, 2132, 1700, 956, 1720, 1259, 2148, 1313, 1777, 1145, 2039, 1216, 391, 1821, 1858, 2079, 1314, 1738, 1839, 2013]. Students [2654, 2723, 1215, 956, 2582, 2016, 2326, 2376, 2477]. Studies [3250, 1215, 3202, 3074, 1070, 922, 278, 1672, 163, 695, 845, 761]. Study [2083, 1885, 623, 851, 812, 79, 1677, 646, 144, 508, 815, 1308, 1599, 2822, 1780, 2780, 1633, 1820, 3174, 891, 2844]. Stukel [3189]. Stunted [3201]. Sturgeon [1345, 2447, 2369, 2742, 2650, 914, 771, 756, 1677, 1923, 107, 525, 2439, 1343, 1344, 1346, 2495, 2533, 3111]. Sturgeons [930]. Style [2599]. Styles [1563, 390]. subjectivity [649]. Sublethal [1612]. Submerged [724, 761]. Submission [2119]. Submissions [1299]. Subsection [2013]. Subsidies [932]. Subsistence [637, 2401, 3175, 830]. Substances [2657]. Subunit [1298, 1074, 1313, 1777, 1839]. Suburban [2595]. Succeed [2004]. Success [802, 257, 1796, 1967, 2203, 1952, 824, 1111]. Successes [1447, 2869]. Successful [2464, 1810, 857, 2517, 2162]. Sucker [2902, 2375, 37]. Suckermouth [2698]. Suction [500]. Sufficiency [1332]. Suggestions [2274, 829]. Suggests [1775]. Suitability [2767]. Summaries [2281, 2441, 2614, 2782, 3090, 2951, 2988, 3241]. Summary [1823, 414, 1859]. Summer [2545, 2937, 541, 1803]. Summit [3098]. Sun [2166]. Sunday [1959]. Sunfishes [2994]. Super [1384]. Superheroes [3073]. SuperIDR [1309]. Supplementation [482, 287, 651]. Supplies [143]. Supply [1239]. Support [2088, 1494, 3077, 2992, 2069, 2359, 1715, 2108, 1322, 1711, 1775, 1837, 2495, 3140]. Supporting [2599, 1767, 3278, 2149, 1333]. Supports [2312, 926, 2409]. Suppression [3263]. Sure [963]. Surface [2305]. Surgical [717]. Surgically [3202]. Surpluses [652]. Surrounding [1325]. Survey [381, 631, 1553, 50, 25, 2531, 139, 423, 1334, 2924, 3078, 2004, 535, 183, 809, 111, 632, 182, 2363, 866, 195]. Surveying [833]. Surveys [939, 3099, 1002, 2518, 479, 1375, 862, 1332, 1370, 448, 706, 3140]. Survival [1296, 609, 3264, 574, 1675, 2452, 2572, 2766, 867]. Susceptibility [401]. Suspension [2061]. Suspensions [792]. Susquehanna [1353, 2764]. Sustain [1467]. Sustainability [794, 707, 1637, 2752, 880, 2163, 2129, 584, 3162, 547]. Sustainable [1499, 1076, 2494, 2835, 3127, 1098, 1432, 1711, 1775, 1083, 1213, 1239, 1241, 457, 2228]. Sustained [2437]. Sustaining [3148, 2466, 1695, 544]. SWAPs [1977]. Swedish [120]. Swim [1400]. Swimming [2019]. Swimways [1253]. Sympatric [665]. Symposia [1180, 2281, 2441, 2614, 2782, 3090, 2988, 1132, 3241, 2951]. Symposium [1795, 2362, 2278, 3115, 1417, 1824, 2629, 2646, 2117, 1590, 1141, 2480, 2261]. Synergies [1572]. Synergistic [897]. Synopsis [1141, 2261]. Synthesis [730, 1308, 2266, 2797]. Synthesizing [988, 3139, 824]. System [396, 383, 227, 479, 774, 901, 1144, 2219, 27, 216, 492, 492, 1323, 1187, 2398]. Systematic [3064, 762]. Systematics [387]. Systemic [1552]. Systems [3064, 1430, 491, 439, 550, 454, 285, 1680, 2962]. Szabo [2297].

T [3284, 3114, 1589]. Tablet [1495]. Tablets [1416]. Tabs [3073]. Tackle [1045, 1131, 1300]. Tackled [2079]. Tackling [3162]. Tadpoles [301]. Tag [2351, 562]. Tagged [2472]. Tagging [561, 947, 3075, 3082, 678, 593, 1348, 2571]. Tags [311, 922, 2362]. Tahoe [2456, 2504]. Tail [2950]. Taillight [2381]. Tailwaters [111]. Taimen [3138]. Take [2891, 2890, 590, 544]. Takes [2491]. Taking [2905, 1788]. Tale [3223]. Tales [1757, 2434]. Talk [1733]. Talking [2505]. Talks [1773, 1399]. Taming [187]. Tampa [2037, 2051, 2155, 2280]. Tangled [2199]. Target [2878]. Targeted [1721, 2168]. Targeting [2220]. Tasty [3094]. Tautog [795]. Tavarez [3267]. Taxa [3029, 3018]. Taxes [1835]. Taxonomic [2711]. Taxonomy [2230, 2229]. Taylor [2279]. Teach [1294]. Teacher [2020]. Teaching [618, 390, 1635, 2009, 1287, 2846]. Team [3129, 3174]. Teams [40]. Technical [1011, 240, 2364, 2228]. Technique [1692, 919, 97, 1868, 706]. Techniques [1445, 1540, 717, 3200, 121]. Techno [502]. Techno-arrogance [502]. Technological [137]. Technologies [1523, 805, 454, 239]. Technology [3226, 2531, 2518, 2275, 318, 2019, 3126, 3053, 249, 2219, 216, 838]. Technology-Based [3126]. Teens [2934]. Teeth [3168, 1864]. Telecoupling [2207]. Telemetry [717, 851, 1613, 2293, 2151, 901, 1540]. Tell [2969, 1731]. Temperature [396, 3187, 856, 2042]. Temporal [514, 1370]. Ten [2297, 2961, 808, 1966]. Tennessee [2388, 623]. Tent [1866]. Tenure [3196]. Term [2339, 2992, 2026, 1226, 1696, 3201, 2121]. Terminology [2522]. Tern [3264]. Terra [2294]. Terrestrial [540, 2455]. test [641]. Testimonials [1811]. Testing [1910, 1909, 2866, 749, 1975]. Tethers [2714]. Texas [2185, 398, 1577]. Texoma [398]. Thank [1575, 1449]. Thanked [3209]. Thanos [2519]. Their [787, 511, 1184, 1196, 2657, 815, 1308, 1187, 410, 947, 557, 266, 1938, 2586, 631, 1894, 2004, 2439, 640]. Theirs [1771]. Them [511, 2313, 1161]. Theme [1572, 1602, 1504, 1515, 2557, 2550, 2549, 2558, 2555, 2556, 2554, 2546]. Themed [1628]. Themes [2796]. Theodore [1147]. There [878, 314, 2175, 2021, 2889, 2733, 2752]. Thermal [2042]. these [963]. Thing [2233]. Things [1609, 1400, 1362, 2726, 1173]. Think [1658, 1307, 554, 2205, 2408]. Thinking [1427, 2907, 2792, 584]. Third [1101, 1269, 1438, 1629, 1860, 2072, 366]. Thomas [2139, 2279]. Thompson [2132]. thon [2934]. Thornyheads [1718]. Those [1652]. Thought [1800, 1954, 1852, 1674, 1747, 2421, 2544]. Thoughts [1649, 2544]. Threat [2970]. Threaten [424, 573, 3238]. Threatened [3018, 1555, 3111, 218, 242]. Threatening [3277]. Threats [2508, 1970, 2806, 1564, 3225]. Three [946, 20, 2351, 2036, 294, 1760]. Threespine [664]. Thresholds [1057]. Thrive [1789]. Thriving [3003, 2149]. throughout [1199]. Throw [2247]. Thwarted [16]. Tidal [177]. Tide [2677, 888]. Tied [1388, 1389]. tiered [355]. Tigers [2168]. Tilapia [2838, 1228]. Time [2492, 2246, 2565, 3085, 2820, 2160, 2081, 2248, 3221, 2724, 592]. Timely [1368]. Times [2853, 1752, 1792, 1791]. Timescape [3053]. Timing [689, 1838]. Tips [2594]. TLAs [3271]. Today [941]. Todd [2310]. Together [1547]. Toledo [1763]. Tolerances [396]. Tom [2479]. Tomelleri [3284, 1969]. Tomorrow [1367]. Too [2259, 1827, 2233, 1492, 2675]. Tool [1871, 2207, 394, 2510, 1767, 1872, 2846, 1570, 901, 1309, 473, 3109, 684, 3201, 3252, 2571, 1214]. Toolbox [2532]. Tools [3226, 1416, 2532, 1287, 1413, 2069, 1699, 2108, 1322, 2962, 1693, 2795]. Top [2140, 1745, 2308, 3123, 2504, 2670]. Topics [1858]. Topminnow [1713]. Tougher [1674]. Tourism [627]. Tournament [235, 504, 670, 173]. Tournament-associated [504]. Tournament-caught [670]. Tournaments [1437, 809]. Tours [1220, 1046]. Toxicants [317]. Toy [1514]. Track [3079]. Tracking [2217, 923, 2256, 3202, 1110, 2724]. Trade [792, 3240, 832, 973]. Tradeoffs [1175]. Tradition [3115]. Traditional [1851, 1375, 2865]. Tragedy [1533, 172]. Trail [1791]. Trailered [917]. Training [1296, 717]. Trandahl [2524]. Trans [3115]. Trans-Atlantic [3115]. Transactions [899, 970, 997, 1048, 1103, 1134, 1179, 1221, 1274, 1315, 1363, 1423, 1487, 1544, 1607, 1650, 1670, 1723, 1755, 1784, 1846, 1895, 1943, 1997, 2073, 2125, 2187, 2254, 2318, 2368]. Transboundary [2041]. Transfer [318, 436, 255, 299, 442]. Transformation [2794]. Transformative [2852]. Transgenic [246, 247, 245]. Transition [3183, 3164]. Translate [1215]. Translating [1809]. Translocating [1227]. Transmission [698, 1868]. Transmitters [2923]. Transparency [2750]. Transparent [2199]. Transplant [3112]. Transponder [2351]. Transport [917]. Transportation [27]. Trap [1729, 2234, 2572]. Trapping [1599]. Traps [115]. Travel [3116, 994, 2893, 2436, 1004]. Traveled [1049]. Travelogue [1948]. Trawl [50, 2712]. Trawler [1384]. Trawling [88]. Trawls [294]. Treasure [2056]. Treat [2714, 1589]. Treatment [125, 386, 1856, 864]. Treaty [154, 1639]. Trees [3026, 2394]. trematode [743]. Trend [1434]. Trends [1422, 523, 491, 1537, 836, 2481, 476, 342, 2765, 1751, 797, 1807, 1370].

Triangles [1510]. Tribal [2339, 1485, 1957]. Tribes [154]. Tributaries [2340]. Trip [2710]. Triploid [196]. Trips [1046, 1696]. Trophic [1369, 1681, 1807]. Trophies [3053]. Trophy [631, 2165]. Tropical [1332]. Tropics [150]. Trotline [1808]. Trough [1583]. Trout [833, 2414, 2639, 1792, 2262, 2405, 2966, 2983, 3158, 3245, 3217, 3273, 712, 3223, 2878, 2339, 2843, 296, 1389, 943, 701, 859, 734, 3263, 320, 740, 208, 200, 114, 1281, 424, 160, 2501, 752, 3126, 2971, 890, 2629, 2646, 2117, 1590, 2690, 152, 2413, 1971, 944, 3009, 2815, 3213, 1712, 533, 445, 215, 2874, 1295, 2277, 655, 1779, 643]. Trouts [226]. Truckee [2639]. True [543]. Trushenski [1691, 1941]. Trust [2365]. Trusted [2337]. Try [1693, 1786, 1816, 2608]. Tsunami [2103]. Tuition [1199]. Tuna [1404, 3164]. Tunas [2647, 98]. Turbines [609, 1518]. Turcik [3257]. Turn [2581]. Turned [1852, 2595]. Turner [1764]. Turning [1696]. Tweeting [1816]. Twenty [298]. Twenty-first [298]. Twice [2621]. Twin [1190, 1113, 1086, 1072, 1101, 1164, 1268]. Twist [2937]. Twitch [1402]. Twitter [1496]. Two [2878, 1916, 2923, 2234, 1387, 977]. Two-Way [2234]. Type [2908]. Types [1761].

U.S [141, 2675, 2397]. U.S. [1631, 2729, 2879, 288, 306, 2961, 1041, 744, 2007, 756, 491, 1684, 175, 866, 2924, 2184, 1174, 920, 1062, 728, 2022, 191, 3029, 3018, 2396, 632, 3044, 2832, 1844, 738, 2451]. UC [3184]. Ulrike [1539]. Umatilla [2849, 646]. Uncertainty [777, 3175, 849, 583]. Uncharted [2385]. Unconscious [2970]. Undergraduate [879, 1333, 129, 240, 2768]. Undergraduates [1429]. Underrepresented [1904, 3019]. Underserved [2656]. Understand [3188, 1254, 2482, 1760, 1300]. Understanding [2548, 958, 1499, 1215, 851, 2463, 1517, 361, 3028, 2518, 1281, 2017, 893, 2818, 1085, 1868, 2040]. Underutilization [14]. Underutilized [988]. Underwater [1701, 2655]. Unexpected [2149, 2969]. Unified [2585]. Unique [1795, 2249, 39]. Unit [1020, 1186]. United [326, 80, 747, 109, 930, 734, 511, 464, 3065, 1187, 2197, 414, 2807, 579, 330, 492, 257, 470, 682, 2508, 1645, 771, 571, 657, 2057, 836, 3112, 133, 2349, 3224, 2011, 1360, 209, 1908, 1334, 1644, 351, 507, 535, 764, 570, 635, 298, 2765, 514, 797, 929, 183, 188, 290, 1634, 809, 124, 1199, 2806, 111, 427, 818, 2102, 353, 345]. Units [1595, 1029, 962, 980, 2898, 1047, 2260, 908, 995]. Universities [491, 3277, 2012, 183]. University [1244, 1533, 1313, 618, 1777, 209, 2662, 216, 2768, 1844, 1801]. University-Based [1533]. Unjust [2853]. Unknowns [946, 264]. Unlikely [2410]. Unmanned [2962, 2424]. Unpolluted [147]. Unprecedented [2853, 1601]. Unpredictable [2853]. Unsettling [2853]. Untested [735]. Untold [1271, 1538]. Unusual [160]. Unveiling [3225]. Update [1173, 1164, 1520, 1505, 1486, 900, 1056, 1067, 1954, 340]. Updates [1159, 1200, 1188, 1660]. Updating [2638]. Upper [177, 1776, 811, 492, 1027, 730, 2437, 2340]. Upstream [574, 2906]. Urban [1783, 1469, 1470, 1464, 2039, 1797, 377, 2595, 2709, 667]. Urbanization [1470]. Urbanized [1171]. Urgent [2891, 2065, 2890, 3224, 3138]. US$17.99 [3267]. US$27.95 [2662]. US$36.00 [3284]. US$49.50 [2185]. USA [2710, 1186, 1599]. Usage [1203, 1213]. USDA [220, 192, 2780]. Use [602, 311, 586, 989, 1783, 17, 717, 777, 1636, 1430, 634, 1258, 846, 2311, 79, 3222, 1600, 780, 351, 2700, 359, 565, 732, 219, 135, 2118, 2962, 2398, 1144, 2325, 2247, 363, 1640, 2365, 632, 466, 804, 543]. Used [1794, 925, 149, 816]. Useful [1693]. Usefulness [188]. User [1540, 2767]. User-Friendly [2767]. Users [288]. Uses [143, 406]. USFCS [2658]. USGS [1517]. Using [870, 1910, 1871, 1266, 724, 100, 821, 2518, 2744, 122, 1692, 2151, 1287, 862, 919, 89, 1145, 2567, 2600, 1254, 2907, 1694, 391, 3054, 1344, 3072, 902, 2399, 1322, 1614, 294, 2362, 2182, 2219, 2276, 2832, 2053, 761, 749, 768, 1775]. Utah [205, 175, 864]. Utility [2085, 2006, 2121, 1859]. Utilization [223]. Utilizing [711, 1111].

V [2729, 2662]. Val [2684]. Valentine [1701]. Valid [2857, 2989]. Validating [1519]. Validation [2923, 2597]. validity [749]. Valley [2056, 2573, 553]. Valuable [3074]. Value [1511, 1138, 852, 1794, 1790, 1302, 2098, 105, 1633, 490, 2365, 2533]. Values [943, 511, 1562, 447, 2192, 662]. Vancouver [887]. Varden [859]. Variance [359]. Variation [695]. Varied [3074]. Various [923]. Varying [3127]. Vecsei [2212]. Vector [894]. Vegetation [724, 761]. Vehicle [2962]. Venting [867]. Ventures [2039]. Vernacular [109]. Versus [2177, 2227, 147, 1461, 101]. Vertical [863]. Vessel [1819, 914]. Vessel-Strike [914]. Vessels [527]. via [707, 2120, 436, 2326, 917]. Vice [1139, 1140, 1320, 1690, 1691, 1901, 1933, 2468, 2975, 2469, 2976, 3283, 1941, 2824, 1480, 2644, 3129, 3282, 2153, 2314, 3130, 2825, 2315, 3131, 2645, 1902, 1934, 2154, 2977, 3281, 1126]. Video [1954, 1775, 2219]. Videographer [3044]. Videos [2328]. Vietnam [2039]. View [1446, 1432, 1843, 2436]. Viewed [44]. Viewers [1482]. Views [2744, 2069, 593, 478, 640]. Vigilance [2348, 2697]. Vignette [2402]. Vintage [2628]. Violations [25]. Viral [2722]. Virgin [205]. Virginia [1777, 1205, 2889, 3075, 589, 795, 126, 2662]. Virtual [2781, 3030]. Virus [1631, 681, 1632]. Viruses [698]. Visibility [1020]. Vision [2411, 1026, 903]. Visions [598]. Visitor [712]. Visits [1548, 1215]. Visual [2518, 845, 706]. Visualization [2750, 2351]. Vital [2362]. vitreus [2514]. Voice [2199, 2896, 2182, 2292]. Voices [3063, 2959, 2987, 3016, 2944, 3086, 3097, 3124, 3137, 3172, 3197, 3249]. Volume [1, 2, 4, 6, 9, 10, 11, 12, 13, 15, 19, 21, 22, 24, 26, 30, 32, 34, 36, 41, 48, 52, 53, 54, 56, 59, 62, 64, 70, 72, 74, 76, 81, 83, 84, 87, 91, 93, 95, 96, 99, 102, 106, 108, 110, 112, 113, 116, 119, 127, 130, 134, 138, 142, 145, 148, 151, 153, 156, 158, 161, 164, 171, 174, 178, 181, 184, 186, 190, 193, 197, 201, 204, 207, 210, 214, 217, 221, 225, 228, 231, 234, 237, 241]. Volume [244, 248, 251, 254, 258, 263, 271, 272, 275, 280, 283, 289, 292, 297, 302, 309, 312, 316, 319, 322, 325, 328, 332, 335, 338, 341, 344, 346, 348, 350, 352, 354, 357, 360, 362, 365, 368, 371, 374, 376, 379, 382, 385, 388, 392, 395, 397, 399, 402, 405, 407, 408, 412, 415, 418, 420, 422, 425, 428, 431, 434, 437, 440, 443, 446, 450, 453, 456, 458, 459, 462, 465, 468, 469, 471, 472, 474, 477, 480, 483, 485, 487, 489, 493]. Volume [496, 499, 501, 503, 506, 509, 512, 515, 518, 521, 524, 526, 530, 534, 537, 539, 542, 545, 548, 552, 555, 560, 563, 566, 569, 572, 575, 578, 580, 582, 585, 588, 591, 594, 597, 600, 603, 604, 607, 610, 614, 617, 620, 622, 625, 628, 630, 633, 636, 639, 642, 645, 648, 650, 653, 656, 658, 661, 663, 666, 669, 672, 675, 677, 680, 683, 685, 687, 690, 693, 696, 699, 702, 705, 708, 710, 713, 716, 719, 721, 726, 729, 733, 736]. Volume [739, 742, 745, 748, 751, 754, 757, 760, 763, 766, 770, 773, 775, 778, 781, 784, 788, 790, 793, 796, 799, 803, 807, 810, 814, 817, 820, 823, 825, 828, 831, 834, 837, 840, 844, 847, 850, 853, 855, 858, 861, 865, 868, 871, 874, 877, 882, 885, 889, 892, 896, 942, 945, 1120, 966, 982, 948, 967, 983, 998, 1010, 1023, 1038, 1054, 1065, 1080, 1093, 1107, 1059, 1003, 1021, 1032, 1090, 1116, 997, 1048, 1103, 1121, 1137, 1154, 1168, 1183, 1195, 1210]. Volume [1224, 1236, 1249, 1263, 1278, 1292, 1260, 1150, 1165, 1192, 1209, 1232, 1246, 1289, 1134, 1179, 1221, 1274, 1293, 1306, 1318, 1329, 1340, 1357, 1366, 1381, 1396, 1410, 1426, 1443, 1456, 1338, 1394, 1458, 1303, 1326, 1355, 1378, 1408, 1441, 1315, 1363, 1423, 1460, 1474, 1491, 1509, 1527, 1549, 1561, 1582, 1594, 1610, 1625, 1653, 1669, 1472, 1506, 1524, 1579, 1592, 1622, 1487, 1544, 1607, 1650, 1670, 1877, 1673, 1688, 1707, 1726, 1743, 1758, 1772, 1787, 1817, 1834, 1849, 1865, 1815, 1686, 1704, 1739]. Volume [1769, 1831, 1861, 1878, 1723, 1755, 1784, 1846, 2092, 1881, 1898, 1915, 1929, 1947, 1963, 1976, 2000, 2030, 2046, 2063, 2076, 1926, 1972, 2043, 1912, 1958, 2027, 2059, 2094, 1895, 1943, 1997, 2073, 2287, 2097, 2113, 2128, 2144, 2159, 2174, 2189, 2204, 2225, 2241, 2257, 2271, 2111, 2172, 2238, 2142, 2157, 2202, 2223, 2268, 2286, 2125, 2187, 2254, 2290, 2307, 2321, 2335, 2346, 2358, 2371, 2383, 2393, 2407, 2420, 2433, 2304, 2354, 2404, 2332, 2343, 2380, 2390, 2318, 2368, 2449, 2461, 2474, 2489, 2500]. Volume [2503, 2529, 2542, 2564, 2580, 2593, 2607, 2626, 2636, 2652, 2667, 2682, 2695, 2707, 2720, 2731, 2746, 2760, 2775, 2791, 2803, 2816, 2830, 2840, 2851, 2863, 2876, 2888, 2904, 2918, 2929, 2942, 2957, 2967, 2984, 2999, 3013, 3024, 3037, 3049, 3060, 3070, 3083, 3095, 3107, 3121, 3135, 3146, 3159, 3169, 3180, 3194, 3207, 3218, 3233, 3246, 3259, 3274]. Volunteer [2425, 1798, 3126, 3082]. Volunteering [1799, 1566]. Volunteers [3073, 711]. Vote [1096, 2708]. Voyage [1800]. vs [2990, 869, 795, 2491]. Vu [2743]. Vulnerability [417, 1990].

W [3284, 1540, 2185, 2755, 2739]. W. [2132]. Wadeable [706]. Wades [2409]. Wait [1935]. Wake [1051]. Waldman [1924]. Walking [1925]. Walleye [2514, 3145, 1115, 257, 2256, 185, 2857, 281, 419, 3139, 438, 2989, 1522]. Walls [1931]. Walter [1648, 1542]. Want [2328, 1719, 1528, 1652, 1046, 1837]. War [2715, 2242]. Warming [268, 211, 264]. Warmwater [511, 301, 149, 29, 71, 744]. Warning [1717]. Warranted [1842]. Warsaw [2597]. Was [1954, 1967, 2969, 1100, 175]. Washington [3284, 216, 1112, 331, 737, 273, 2491]. Washoe [2262]. Water [602, 1597, 1499, 1645, 1783, 1993, 205, 1312, 1537, 122, 2509, 1469, 1470, 2660, 143, 1497, 141, 2336, 299, 442, 219, 2301, 1273, 1775, 1760, 168, 856, 2697, 349, 801, 1144, 1827]. Watermilfoil [394]. Waters [2601, 147, 3224, 2168, 3078, 809, 655, 2385, 2451, 1127]. Watershed [3250, 1909, 413, 2339, 1002, 182, 813, 1891, 2600, 762, 466, 616, 730]. Watershed-Scale [2339]. Watersheds [522, 1910, 1565, 212, 386, 104, 932]. Watt [2821]. Wave [805]. Waves [2080]. Way [1100, 1463, 2068, 1800, 1055, 2234, 1747, 3262]. Ways [2088, 1068, 3220, 2733, 1775]. Weak [886]. Web [2199, 1805, 776, 869, 1871, 2510]. Web-Based [1871]. Webs [1806, 2149, 692, 932]. Website [1595]. Webster [2366]. Weeds [33]. Week [1446, 2280]. Wegener [2216]. Weighing [2365]. Weight [274]. Welcome [1730, 2155, 2341, 2523, 2673, 3103]. Welcomes [2456]. Weldon [1768]. Well [1294, 1681, 337, 2031, 2205, 260, 783]. Well-Established [1681]. Well-Managed [260]. Wenatchee [1599]. We’re [1463]. Werner [2798]. Wes [2973]. West [848, 1777, 513, 611, 1583, 203, 488, 126]. West-Central [488]. Western [3210, 2070, 3228, 3214, 3120, 375, 890, 3198, 1779, 1631, 701, 734, 464, 2197, 1971, 3153]. Westslope [3223, 2971]. Wetlands [1789, 2466, 1043, 804, 605]. Weyl [2799]. Whale [1741]. Whales [2170, 2467]. WHAM [725, 762]. Wheaton [2662]. Whelan [1902, 2154]. Where [2845, 1127, 1164, 2273, 1445, 1250, 1004, 212, 1387, 1161, 2022, 1712, 1848, 2008, 136, 2127, 2559]. Whether [1266]. while [2228]. Whirling [1747]. Whisperer [3012]. White [146, 33, 859, 872, 525, 1343, 1939]. Whitefish [1776, 3175, 1387, 2069]. Whitesel [1219]. Who [2945, 717, 366, 2947, 1906, 2064, 2505, 1161, 917, 3026]. Whole [776]. Whole-lake [776]. WI [1243]. Widening [3052]. Widespread [2165, 1747]. Wild [2639, 1243, 2418, 2921, 1388, 2373, 320, 187, 3126, 2629, 2646, 2117, 1590, 152, 849, 1633, 1820, 2053, 1214, 1243]. Wilderness [430, 655]. Wildfire [475]. Wildlife [381, 288, 1547, 192, 2686, 25, 3062, 2349, 63, 139, 1429, 2019, 866, 141, 2184, 829, 1174, 929, 2022, 2150, 389, 191, 1199, 3221, 3029, 3018, 7, 2504, 1300, 2451, 2864, 3002, 605, 2613, 3044]. Wiley [2934]. Will [2299, 1820, 2436, 2899, 665, 1148, 2215]. Willamette [2405]. William [2979, 2963, 3114, 2725]. Win [2242]. Wind [2385]. Windows [1683]. Winkle [2366]. Winner [1885, 1216, 1821, 1387]. Winners [1087, 1283, 1257, 1451, 1661, 1875, 2055, 2428, 2587, 2753, 2912, 3045, 3056, 3211, 1259]. Winning [2934]. Wins [2709, 2697]. Winter [954, 1885]. Wintu [592]. Wisconsin [2779, 774, 1956, 2520, 2960, 2300, 466]. Wisdom [2519]. Withdrawal [765]. within [2545, 925, 104, 410, 902, 3005, 1552, 1227]. Without [2669, 2482, 1149]. Woes [1759, 890, 2348]. Wolff [2995]. Women [1163, 787, 2823, 3063, 2959, 2987, 2944, 3100, 2535, 3086, 2958, 3162, 1321, 3097, 3124, 3137, 3172, 3197, 3249]. Wonders [3173, 1804]. Won’t [2633, 1046]. Wood [1577]. Woods [2552]. Woody [862, 776]. Word [1330]. Work [1503, 1798, 2920, 1536, 883]. Working [1806, 1245, 640]. Works [1918]. Workshop [1573, 377, 2934]. Workshops [2124, 2141, 2156, 2171, 2186, 2201, 2222, 2236, 2267, 879, 318]. World [1100, 1126, 2891, 2627, 2890, 1367, 1945, 366, 1925, 1684, 2661, 1055, 1642, 2672, 2690, 2647, 3053, 3174, 2251, 2436, 2530, 2620, 2734, 2147, 1516, 1483, 3253, 2023, 2661, 2858, 3280, 3276, 2881, 3138]. World-Class [2661, 3174]. Worldwide [1585]. Worries [1900]. Worth [2438, 2112, 2277]. Would [155, 2732]. Woundfin [205]. Wrap [1268, 2781, 2023, 3039, 1449]. Wrap-up [2781, 2023]. Wrapup [3089]. Wrinkles [2160]. Writing [1299, 1259, 1386, 2846, 1216, 391, 1821]. Wrong [1852]. Wyoming [250, 725, 762, 1893, 655, 864].

XP [1513].

Yakima [287]. Yakima/Klickitat [287]. Y’All [2504]. Yasmine [1142]. Year [2696, 1617, 3085, 2568, 2715, 1089, 2384, 1956, 98, 1897]. Years [2492, 2628, 2723, 1553, 1665, 1323, 1630, 2103, 2960, 2477, 818, 913, 2260]. Yeas [486]. Yellow [714]. Yellowstone [2414, 2983, 3273, 3263, 424, 752]. Yes [1932, 2889, 1820]. Yield [432, 35]. Yields [1550, 1432, 1241]. Yolo [605]. York [1792, 1791, 1002, 182, 329, 125]. Yosemite [2413]. Young [2101, 1389, 1715, 2363]. Youth [1737, 505]. Yukon [3188, 445].

Zeb [2620]. Zebra [337, 401]. Zebrafish [2619]. Zero [765]. Zone [1677, 3127, 2675]. Zoning [863]. Zoo [1763]. Zoogeography [2738]. Zooplankton [163, 695].

References

Anonymous:1976:FIAa

[1] Anonymous. Full issue article volume 1, issue 1. Fisheries, 1(1):1–52, 1976. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

      http://www.tandfonline.com/doi/abs/10.1577/1548-8446-1-1.

Anonymous:1976:FIAb

[2] Anonymous. Full issue article volume 1, issue 2. Fisheries, 1(2):1–44, 1976. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

      http://www.tandfonline.com/doi/abs/10.1577/1548-8446-1-2.

Huner:1976:RCF

[3] Jay V. Huner. Raising crawfish for food and fish bait: a new polyculture crop with fish. Fisheries, 1(2):7–9, 1976. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

      doi/abs/10.1577/1548-8446%281976%29001%3C0007%3ARCFFAF%3E2.
      0.CO%3B2.

Anonymous:1976:FIAc

[4] Anonymous. Full issue article volume 1, issue 3. Fisheries, 1(3):1–48, 1976. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-1-3.

Stanley:1976:RGC

[5] Jon G. Stanley. Reproduction of the grass carp (Ctenopharyngodon idella) outside its native range. Fisheries, 1(3):7–10, 1976. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281976%29001%
3C0007%3AROTGCC%3E2.0.CO%3B2.

Anonymous:1976:FIAd

[6] Anonymous. Full issue article volume 1, issue 4. Fisheries, 1(4):1–48, 1976. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-1-4.

Tennant:1976:IFR

[7] Donald Leroy Tennant. Instream flow regimens for fish, wildlife, recreation and related environmental resources. Fisheries, 1(4):6– 10, 1976. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281976%29001%3C0006%3AIFRFFW%3E2.0.CO%3B2.

Walker:1976:PBC

[8] Charles R. Walker. Polychlorinated biphenyl compounds (PCB’s) and fishery resources. Fisheries, 1(4):19–25, 1976. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281976%29001%3C0019%3APBCPAF%
3E2.0.CO%3B2.

Anonymous:1976:FIAe

[9] Anonymous. Full issue article volume 1, issue 5. Fisheries, 1(5):1–52, 1976. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-1-5.

Anonymous:1976:FIAf

[10] Anonymous. Full issue article volume 1, issue 6. Fisheries, 1(6):1–54, 1976. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-1-6.

Anonymous:1977:FIAa

[11] Anonymous. Full issue article volume 2, issue 1. Fisheries, 2(1):1–50, 1977. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-2-1.

Anonymous:1977:FIAb

[12] Anonymous. Full issue article volume 2, issue 2. Fisheries, 2(2):1–50, 1977. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-2-2.

Anonymous:1977:FIAc

[13] Anonymous. Full issue article volume 2, issue 3. Fisheries, 2(3):1–62, 1977. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-2-3.

Miller:1977:RUS

[14] Robert J. Miller. Resource underutilization in a spider crab industry. Fisheries, 2(3):9–30, 1977. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281977%29002%3C0009%3ARUIASC%3E2.0.CO%3B2.

Anonymous:1977:FIAd

[15] Anonymous. Full issue article volume 2, issue 4. Fisheries, 2(4):1–42, 1977. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-2-4.

Armour:1977:HNE

[16] Carl L. Armour. How the National Environmental Policy Act (NEPA) can be thwarted — possible corrective measures. Fisheries, 2(4):14– 17, 1977. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281977%29002%3C0014%3AHTNEPA%3E2.0.CO%3B2.

Clark:1977:CUF

[17] Richard D. Clark, Jr., Donald L. Garling, Jr., and Robert T. Lackey. Computer use in freshwater fisheries management. Fisheries, 2(4):21– 23, 1977. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281977%29002%3C0021%3ACUIFFM%3E2.0.CO%3B2.

Gilbert:1977:CEE

[18] Douglas L. Gilbert. Changing education and employment in natural resources management. Fisheries, 2(4):24–25, 1977. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281977%29002%
3C0024%3ACEAEIN%3E2.0.CO%3B2.

Anonymous:1977:FIAe

[19] Anonymous. Full issue article volume 2, issue 5. Fisheries, 2(5):1–42, 1977. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-2-5.

Inman:1977:GCC

[20] Charles R. Inman, Royce C. Dewey, and Philip P. Durocher. Growth comparisons and catchability of three largemouth bass strains. Fisheries, 2(5):20–25, 1977. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281977%29002%3C0020%3AGCACOT%3E2.0.CO%3B2.

Anonymous:1977:FIAf

[21] Anonymous. Full issue article volume 2, issue 6. Fisheries, 2(6):1–54, 1977. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-2-6.

Anonymous:1978:FIAa

[22] Anonymous. Full issue article volume 3, issue 1. Fisheries, 3(1):1–50, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-3-1.

Bennett:1978:CFF

[23] David H. Bennett, Ed L. Hampton, and Robert T. Lackey. Current and future fisheries management goals: Implications for future management. Fisheries, 3(1):10–14, 1978. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281978%29003%3C0010%3ACAFFMG%3E2.0.CO%3B2.

Anonymous:1978:FIAb

[24] Anonymous. Full issue article volume 3, issue 2. Fisheries, 3(2):1–48, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-3-2.

Cowles:1978:SMR

[25] Cleveland J. Cowles, Kirk H. Beattie, and Robert H. Giles, Jr. A survey of methods of recording reports of fish and wildlife law violations. Fisheries, 3(2):8–11, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281978%29003%3C0008%3AASOMOR%3E2.0.CO%3B2.

Anonymous:1978:FIAc

[26] Anonymous. Full issue article volume 3, issue 3. Fisheries, 3(3):1–46, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-3-3.

Taft:1978:FDT

[27] Edward P. Taft III and Yusuf G. Mussalli. Fish diversion and transportation system for power plant application. Fisheries, 3(3):2–5, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281978%29003%3C0002%3AFDATSF%3E2.0.CO%3B2.

Weithman:1978:MEF

[28] A. Stephen Weithman and Richard O. Anderson. A method of evaluating fishing quality. Fisheries, 3(3):6–10, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281978%29003%3C0006%3AAMOEFQ%3E2.
0.CO%3B2.

McSwain:1978:MCA

[29] Lawrence E. McSwain, Paul M. Loska, and Jerome F. Germann. A method of cost analysis for warmwater fish hatcheries. Fisheries, 3(3): 12–15, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281978%29003%3C0012%3AAMOCAF%3E2.0.CO%3B2.

Anonymous:1978:FIAd

[30] Anonymous. Full issue article volume 3, issue 4. Fisheries, 3(4):1–54, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-3-4.

Matsuda:1978:GAD

[31] Yoshiaki Matsuda. The growth of aquaculture in developing countries: Potentials, patterns, and pitfalls. Fisheries, 3(4):2–6, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281978%29003%
3C0002%3ATGOAID%3E2.0.CO%3B2.

Anonymous:1978:FIAe

[32] Anonymous. Full issue article volume 3, issue 5. Fisheries, 3(5):1–46, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-3-5.

Custer:1978:WAB

[33] Phillip E. Custer, Francis D. Halverson, Jim Malone, and Cesar Von Chong. The white amur as a biological control agent of aquatic weeds in the Panama Canal. Fisheries, 3(5):2–9, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281978%29003%3C0002%3ATWAAAB%
3E2.0.CO%3B2.

Anonymous:1978:FIAf

[34] Anonymous. Full issue article volume 3, issue 6. Fisheries, 3(6):1–58, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-3-6.

Sissenwine:1978:MAF

[35] Michael P. Sissenwine. Is MSY an adequate foundation for optimum yield? Fisheries, 3(6):22–42, 1978. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281978%29003%3C0022%3AIMAAFF%3E2.
0.CO%3B2.

Anonymous:1979:FIAa

[36] Anonymous. Full issue article volume 4, issue 1. Fisheries, 4(1):1–46, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-4-1.

Holey:1979:NGS

[37] Mark Holey, Bruce Hollender, Mark Imhof, Roman Jesien, Richard Konopacky, Michael Toneys, and Daniel Coble. “Never Give a Sucker an Even Break”. Fisheries, 4(1):2–6, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281979%29004%3C0002%3ANGASAE%3E2.
0.CO%3B2.

Keup:1979:FLR

[38] Lowell E. Keup. Fisheries in lake restoration. Fisheries, 4(1):7– 20, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281979%29004%3C0007%3AFILR%3E2.0.CO%3B2.

Klessig:1979:LDU

[39] Lowell Klessig. Lake districts — a unique organization with a special purpose. Fisheries, 4(1):10–14, 1979. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281979%29004%3C0010%3ALDAUOW%3E2.0.CO%3B2.

Everest:1979:HDI

[40] Fred H. Everest. How to demonstrate the importance of fishery resources to interdisciplinary planning teams. Fisheries, 4(1):15–20, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281979%29004%3C0015%3AHTDTIO%3E2.0.CO%3B2.

Anonymous:1979:FIAb

[41] Anonymous. Full issue article volume 4, issue 2. Fisheries, 4(2):1–69, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-4-2.

Platts:1979:RAS

[42] William S. Platts. Relationships among stream order, fish populations, and aquatic geomorphology in an Idaho river drainage. Fisheries, 4(2): 5–9, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281979%29004%3C0005%3ARASOFP%3E2.0.CO%3B2.

Lackey:1979:FEI

[43] Robert T. Lackey. Fisheries education in the 1980’s: The issues. Fisheries, 4(2):16–17, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281979%29004%3C0016%3AFEITST%3E2.0.CO%3B2.

Chapman:1979:FEV

[44] Douglas G. Chapman. Fisheries education as viewed from inside. Fisheries, 4(2):18–21, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281979%29004%3C0018%3AFEAVFI%3E2.0.CO%3B2.

Hester:1979:FEF

[45] F. Eugene Hester. Fisheries education from the Federal perspective. Fisheries, 4(2):22–24, 1979. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281979%29004%3C0022%3AFEFTFP%3E2.0.CO%3B2.

Donaldson:1979:FES

[46] John R. Donaldson. Fisheries education from the State perspective. Fisheries, 4(2):24–26, 1979. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281979%29004%3C0024%3AFEFTSP%3E2.0.CO%3B2.

Olmsted:1979:FEP

[47] Larry L. Olmsted. Fisheries education from the private sector perspective. Fisheries, 4(2):26–28, 1979. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281979%29004%3C0026%3AFEFTPS%3E2.0.CO%3B2.

Anonymous:1979:FIAc

[48] Anonymous. Full issue article volume 4, issue 3. Fisheries, 4(3):1–62, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-4-3.

Schuman:1979:RRE

[49] Sandor P. Schuman, Tommy L. Brown, and Michael W. Duttweiler. The roles of research and extension education in the developing Lake Ontario salmonid fishery. Fisheries, 4(3):6–8, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281979%29004%3C0006%3ATRORAE%
3E2.0.CO%3B2.

Clark:1979:ABT

[50] Stephen H. Clark. Application of bottom-trawl survey data to fish stock assessment. Fisheries, 4(3):9–15, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281979%29004%3C0009%3AAOBSDT%3E2.
0.CO%3B2.

Shafland:1979:NNF

[51] Paul L. Shafland. Non-native fish introductions with special reference to Florida. Fisheries, 4(3):18–24, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281979%29004%3C0018%3ANFIWSR%3E2.
0.CO%3B2.

Anonymous:1979:FIAd

[52] Anonymous. Full issue article volume 4, issue 4. Fisheries, 4(4):1–54, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-4-4.

Anonymous:1979:FIAe

[53] Anonymous. Full issue article volume 4, issue 5. Fisheries, 4(5):1–46, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-4-5.

Anonymous:1979:FIAf

[54] Anonymous. Full issue article volume 4, issue 6. Fisheries, 4(6):1–70, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-4-6.

Thomas:1979:PAB

[55] John M. Thomas and D. H. McKenzie. A procedure for assessing biological effects of power plants on fish. Fisheries, 4(6):23– 27, 1979. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281979%29004%3C0023%3AAPFABE%3E2.0.CO%3B2.

Anonymous:1980:FIAa

[56] Anonymous. Full issue article volume 5, issue 1. Fisheries, 5(1):1–46, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-5-1.

Platts:1980:PFH

[57] William S. Platts. A plea for fishery habitat classification. Fisheries, 5 (1):2–6, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281980%29005%3C0002%3AAPFFHC%3E2.0.CO%3B2.

Kauffman:1980:EMI

[58] John Kauffman. Effect of a mercury-induced consumption ban on angling pressure. Fisheries, 5(1):10–12, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281980%29005%3C0010%3AEOAMCB%3E2.
0.CO%3B2.

Anonymous:1980:FIAb

[59] Anonymous. Full issue article volume 5, issue 2. Fisheries, 5(2):1–46, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-5-2.

Lippson:1980:MMP

[60] A. J. Lippson, M. S. Haire, A. F. Holland, F. Jacobs, J. Jensen, R. L. Moran-Johnson, T. T. Polgar, and W. R. Richkus. Martin marietta produces classic environmental atlas. Fisheries, 5(2):2– 4, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281980%29005%3C0002%3AMMPCEA%3E2.0.CO%3B2.

Wydoski:1980:FA

[61] Richard S. Wydoski. The Fisheries Academy. Fisheries, 5(2):13– 18, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281980%29005%3C0013%3ATFA%3E2.0.CO%3B2.

Anonymous:1980:FIAc

[62] Anonymous. Full issue article volume 5, issue 3. Fisheries, 5(3):1–62, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-5-3.

Horak:1980:FWM

[63] Gerald C. Horak and James E. Olson. Fish and wildlife mitigation: The need for standards and criteria to determine effectiveness and recommendations for reporting requirements. Fisheries, 5(3):2– 6, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281980%29005%3C0002%3AFAWM%3E2.0.CO%3B2.

Anonymous:1980:FIAd

[64] Anonymous. Full issue article volume 5, issue 4. Fisheries, 5(4):1–82, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-5-4.

Labisky:1980:FSL

[65] Ronald F. Labisky, Douglas R. Gregory, Jr., and Joseph A. Conti. Florida’s spiny lobster fishery: an historical perspective. Fisheries, 5 (4):28–37, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281980%29005%3C0028%3AFSLF%3E2.0.CO%3B2.

Simmons:1980:RFS

[66] David C. Simmons. Review of the Florida spiny lobster resource. Fisheries, 5(4):37–43, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281980%29005%3C0037%3AROTFSL%3E2.0.CO%3B2.

Zuboy:1980:LFM

[67] J. R. Zuboy, A. C. Jones, and T. J. Costello. Lobster fishery management under the Fishery Conservation and Management Act. Fisheries, 5 (4):50–52, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281980%29005%3C0050%3ALFMUTF%3E2.0.CO%3B2.

Davis:1980:JSL

[68] Gary E. Davis. Juvenile spiny lobster management or how to make the most of what you get. Fisheries, 5(4):57–59, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281980%29005%
3C0057%3AJSLM%3E2.0.CO%3B2.

Tamm:1980:SLC

[69] George R. Tamm. Spiny lobster culture: an alternative to natural stock assessment. Fisheries, 5(4):59–62, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281980%29005%
3C0059%3ASLC%3E2.0.CO%3B2.

Anonymous:1980:FIAe

[70] Anonymous. Full issue article volume 5, issue 5. Fisheries, 5(5):1–56, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-5-5.

Modde:1980:SSP

[71] Timothy Modde. State stocking policies for small warmwater impoundments. Fisheries, 5(5):13–17, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281980%29005%3C0013%3ASSPFSW%3E2.
0.CO%3B2.

Anonymous:1980:FIAf

[72] Anonymous. Full issue article volume 5, issue 6. Fisheries, 5(6):1–78, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-5-6.

Acheson:1980:ATL

[73] James M. Acheson. Attitudes towards limited entry among finfishermen in Northern New England. Fisheries, 5(6):20–25, 1980. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281980%29005%
3C0020%3AATLEAF%3E2.0.CO%3B2.

Anonymous:1981:FIAa

[74] Anonymous. Full issue article volume 6, issue 1. Fisheries, 6(1):1–46, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-6-1.

Weithman:1981:GPA

[75] A. Stephen Weithman and Ronald J. Ebert. Goal programming to assist in decision making. Fisheries, 6(1):5–8, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281981%29006%3C0005%3AGPTAID%
3E2.0.CO%3B2.

Anonymous:1981:FIAb

[76] Anonymous. Full issue article volume 6, issue 2. Fisheries, 6(2):1–46, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-6-2.

Wen-Hwa:1981:PSG

[77] Kwain Wen-Hwa and Andrew H. Lawrie. Pink salmon in the Great Lakes. Fisheries, 6(2):2–6, 1981. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281981%29006%3C0002%3APSITGL%3E2.0.CO%3B2.

Lee:1981:REP

[78] Emery Lee. Range extension of pink salmon (Oncorhynchus gorbuscha) into the Lower Great Lakes. Fisheries, 6(2):7–10, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281981%29006%
3C0007%3AREOPSO%3E2.0.CO%3B2.

Groth:1981:EUS

[79] Philip G. Groth. Effective use of sociocultural data in fisheries management: a case study. Fisheries, 6(2):11–16, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281981%29006%
3C0011%3AEUOSDI%3E2.0.CO%3B2.

Carlson:1981:PPS

[80] Douglas M. Carlson and Patrick S. Bonislawsky. The paddlefish (Polyodon spathula) fisheries of the Midwestern United States. Fisheries, 6 (2):17–27, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281981%29006%3C0017%3ATPPSFO%3E2.0.CO%3B2.

Anonymous:1981:FIAc

[81] Anonymous. Full issue article volume 6, issue 3. Fisheries, 6(3):1–62, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-6-3.

Nelson:1981:CDE

[82] Kent L. Nelson and Edward Weaver. Criteria for the design and evaluation of stream excavation projects in North Carolina. Fisheries, 6(3): 7–10, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281981%29006%3C0007%3ACFTDAE%3E2.0.CO%3B2.

Anonymous:1981:FIAd

[83] Anonymous. Full issue article volume 6, issue 4. Fisheries, 6(4):1–48, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-6-4.

Anonymous:1981:FIAe

[84] Anonymous. Full issue article volume 6, issue 5. Fisheries, 6(5):1–64, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-6-5.

Benson:1981:FIE

[85] Norman G. Benson. The freshwater-inflow-to-estuaries issue. Fisheries, 6(5):8–10, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281981%29006%3C0008%3ATFI%3E2.0.CO%3B2.

Jagoe:1981:CAF

[86] Charles H. Jagoe, David P. Small, Robert L. Blake, Susan Woodward, David C. Smith, and John R. Moring. Commercial aquaculture of fishes in Maine: Status and future prospects. Fisheries, 6(5): 16–25, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281981%29006%3C0016%3ACAOFIM%3E2.0.CO%3B2.

Anonymous:1981:FIAf

[87] Anonymous. Full issue article volume 6, issue 6. Fisheries, 6(6):1–62, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-6-6.

Seki:1981:HST

[88] Humitake Seki, Atsunobu Hamada, Tetsuo Iwami, and Robin J. Le-Brasseur. Hobikiami sail trawling in Japan. Fisheries, 6(6):2–15, 1981. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281981%29006%3C0002%3AHSTIJ%3E2.0.CO%3B2.

Karr:1981:ABI

[89] James R. Karr. Assessment of biotic integrity using fish communities. Fisheries, 6(6):21–27, 1981. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281981%29006%3C0021%3AAOBIUF%3E2.0.CO%3B2.

Sissenwine:1981:OSM

[90] M. P. Sissenwine. An overview of some methods of fish stock assessment. Fisheries, 6(6):31–35, 1981. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281981%29006%3C0031%3AAOOSMO%3E2.0.CO%3B2.

Anonymous:1982:FIAa

[91] Anonymous. Full issue article volume 7, issue 1. Fisheries, 7(1):1–54, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-7-1.

Sousa:1982:NOF

[92] Robert J. Sousa. A national overview of fishery research funded by the Dingell–Johnson Program. Fisheries, 7(1):3–6, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281982%29007%
3C0003%3AANOOFR%3E2.0.CO%3B2.

Anonymous:1982:FIAb

[93] Anonymous. Full issue article volume 7, issue 2. Fisheries, 7(2):1–46, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-7-2.

Thorson:1982:ICE

[94] Thomas B. Thorson. The impact of commercial exploitation on sawfish and shark populations in Lake Nicaragua. Fisheries, 7(2):2– 10, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281982%29007%3C0002%3ATIOCEO%3E2.0.CO%3B2.

Anonymous:1982:FIAc

[95] Anonymous. Full issue article volume 7, issue 3. Fisheries, 7(3):1–68, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-7-3.

Anonymous:1982:FIAd

[96] Anonymous. Full issue article volume 7, issue 4. Fisheries, 7(4):1–54, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-7-4.

Oswood:1982:AFH

[97] Mark W. Oswood and Willard E. Barber. Assessment of fish habitat in streams: Goals, constraints, and a new technique. Fisheries, 7(4): 8–11, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281982%29007%3C0008%3AAOFHIS%3E2.0.CO%3B2.

Sakagawa:1982:ISY

[98] Gary T. Sakagawa and Philip E. K. Symons. The International Skipjack Year Program of the International Commission for the Conservation of Atlantic Tunas. Fisheries, 7(4):12–17, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281982%29007%3C0012%3ATISYPO%
3E2.0.CO%3B2.

Anonymous:1982:FIAe

[99] Anonymous. Full issue article volume 7, issue 5. Fisheries, 7(5):1–46, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-7-5.

Frie:1982:MFS

[100] Richard V. Frie. Measurement of fish scales and back calculation of body lengths using a digitizing pad and microcomputer. Fisheries, 7 (5):5–8, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281982%29007%3C0005%3AMOFSAB%3E2.0.CO%3B2.

Jennings:1982:NES

[101] Mark R. Jennings. The nomenclatural enigma of single versus double -i endings for scientific patronyms emended to represent a man’s name: a comment and some considerations. Fisheries, 7(5):9–10, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281982%29007%3C0009%3ATNEOSV%3E2.0.CO%3B2.

Anonymous:1982:FIAf

[102] Anonymous. Full issue article volume 7, issue 6. Fisheries, 7(6):1–54, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-7-6.

Davies:1982:PDR

[103] William D. Davies, William L. Shelton, and Stephen P. Malvestuto. Prey-dependent recruitment of largemouth bass: a conceptual model. Fisheries, 7(6):12–15, 1982. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281982%29007%3C0012%3APROLB%3E2.0.CO%3B2.

Marsh:1982:OAL

[104] Paul C. Marsh and James E. Luey. Oases for aquatic life within agricultural watersheds. Fisheries, 7(6):16–24, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281982%29007%3C0016%3AOFALWA%
3E2.0.CO%3B2.

Martin:1982:EVL

[105] William E. Martin, Frank H. Bollman, and Russell L. Gum. Economic value of Lake Mead fishery. Fisheries, 7(6):20–24, 1982. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281982%29007%
3C0020%3AEVOLMF%3E2.0.CO%3B2.

Anonymous:1983:FIAa

[106] Anonymous. Full issue article volume 8, issue 1. Fisheries, 8(1):1–62, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-8-1.

Kallemeyn:1983:SPS

[107] Larry Kallemeyn. Status of the pallid sturgeon: Scaphirhynchus albus. Fisheries, 8(1):3–9, 1983. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281983%29008%3C0003%3ASOTPS%3E2.0.CO%3B2.

Anonymous:1983:FIAb

[108] Anonymous. Full issue article volume 8, issue 2. Fisheries, 8(2):1–46, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-8-2.

Cloutman:1983:VNF

[109] Donald G. Cloutman and Larry L. Olmsted. Vernacular names of freshwater fishes of the Southeastern United States. Fisheries, 8(2): 7–11, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281983%29008%3C0007%3AVNOFFO%3E2.0.CO%3B2.

Anonymous:1983:FIAc

[110] Anonymous. Full issue article volume 8, issue 3. Fisheries, 8(3):1–56, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-8-3.

Swink:1983:NST

[111] William D. Swink. Nonmigratory salmonids and tailwaters — a survey of stocking practices in the United States. Fisheries, 8(3):5– 9, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281983%29008%3C0005%3ANSATSO%3E2.0.CO%3B2.

Anonymous:1983:FIAd

[112] Anonymous. Full issue article volume 8, issue 4. Fisheries, 8(4):1–46, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-8-4.

Anonymous:1983:FIAe

[113] Anonymous. Full issue article volume 8, issue 5. Fisheries, 8(5):1–46, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-8-5.

Hicks:1983:AAA

[114] Charles E. Hicks, Lawrence C. Belusz, Daniel J. Witter, and Pamela S. Haverland. Application of angler attitudes and motives to management strategies at Missouri’s trout parks. Fisheries, 8(5):2–7, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281983%29008%3C0002%3AAOAAAM%3E2.0.CO%3B2.

Miller:1983:CCF

[115] Robert J. Miller. Considerations for conducting field experiments with baited traps. Fisheries, 8(5):14–17, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281983%29008%3C0014%3ACFCFEW%3E2.
0.CO%3B2.

Anonymous:1983:FIAf

[116] Anonymous. Full issue article volume 8, issue 6. Fisheries, 8(6):1–46, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-8-6.

Hartmann:1983:CIP

[117] Jürgen Hartmann. Charr as indicators of pollution. Fisheries, 8(6): 10–12, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281983%29008%3C0010%3ACAIOP%3E2.0.CO%3B2.

Bogardus:1983:CEB

[118] Raymond B. Bogardus. The consulting environmental biologist. Fisheries, 8(6):16–18, 1983. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281983%29008%3C0016%3ATCEB%3E2.0.CO%3B2.

Anonymous:1984:FIAa

[119] Anonymous. Full issue article volume 9, issue 1. Fisheries, 9(1):1–68, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-9-1.

Hasselrot:1984:LAS

[120] Bengt Hasselrot and Hans Hultberg. Liming of acidified Swedish lakes and streams and its consequences for aquatic ecosystems. Fisheries, 9 (1):4–9, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281984%29009%3C0004%3ALOASLA%3E2.0.CO%3B2.

Rosseland:1984:ARE

[121] Bjørn Olav Rosseland and Odd K. Skogheim. Attempts to reduce effects of acidification on fishes in Norway by different mitigation techniques. Fisheries, 9(1):10–16, 1984. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281984%29009%3C0010%3AATREOA%3E2.0.CO%3B2.

Gunn:1984:ESM

[122] J. M. Gunn and W. Keller. In Situ manipulation of water chemistry using crushed limestone and observed effects on fish. Fisheries, 9(1): 19–24, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281984%29009%3C0019%3AISMOWC%3E2.0.CO%3B2.

White:1984:EFR

[123] W. J. White, W. D. Watt, and C. D. Scott. An experiment on the feasibility of rehabilitating acidified Atlantic salmon habitat in Nova Scotia by the addition of lime. Fisheries, 9(1):25–30, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281984%29009%3C0025%3AAEOTFO%3E2.0.CO%3B2.

Schreiber:1984:FPM

[124] R. Kent Schreiber and Paul J. Rago. The federal plan for mitigation of acid precipitation effects in the United States: Opportunities for basic and applied research. Fisheries, 9(1):31–36, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281984%29009%
3C0031%3ATFPFMO%3E2.0.CO%3B2.

Kretser:1984:TNY

[125] Walter A. Kretser and James R. Colquhoun. Treatment of New York’s Adirondack lakes by liming. Fisheries, 9(1):36–41, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281984%29009%
3C0036%3ATONYAL%3E2.0.CO%3B2.

Zurbuch:1984:NAS

[126] Peter E. Zurbuch. Neutralization of acidified streams in West Virginia. Fisheries, 9(1):42–47, 1984. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281984%29009%3C0042%3ANOASIW%3E2.0.CO%3B2.

Anonymous:1984:FIAb

[127] Anonymous. Full issue article volume 9, issue 2. Fisheries, 9(2):1–46, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-9-2.

Dombeck:1984:FMF

[128] Michael Dombeck, James Hammill, and William Bullen. Fisheries management and fish dependent birds. Fisheries, 9(2):2–4, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281984%29009%3C0002%3AFMAFDB%3E2.0.CO%3B2.

Nielsen:1984:UCI

[129] Larry A. Nielsen. Undergraduate curricula of inland fisheries schools. Fisheries, 9(2):5–7, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/abs/10.
1577/1548-8446%281984%29009%3C0005%3AUCOIFS%3E2.0.CO%3B2.

Anonymous:1984:FIAc

[130] Anonymous. Full issue article volume 9, issue 3. Fisheries, 9(3):1–54, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-9-3.

Penkal:1984:COO

[131] Russ F. Penkal and Glenn R. Phillips. Construction and operation of oil and gas pipelines. Fisheries, 9(3):6–8, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281984%29009%3C0006%3ACAOOOA%
3E2.0.CO%3B2.

Royce:1984:PEF

[132] William F. Royce. A professional education for fishery scientists. Fisheries, 9(3):12–17, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281984%29009%3C0012%3AAPEFFS%3E2.0.CO%3B2.

Helfrich:1984:CES

[133] Louis A. Helfrich and Peter T. Bromley. Cooperative extension service education in fisheries in the United States. Fisheries, 9(3):18– 21, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281984%29009%3C0018%3ACESEIF%3E2.0.CO%3B2.

Anonymous:1984:FIAd

[134] Anonymous. Full issue article volume 9, issue 4. Fisheries, 9(4):1–46, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-9-4.

Novinger:1984:OUS

[135] Gary D. Novinger. Observations on the use of size limits for black basses in large impoundments. Fisheries, 9(4):2–6, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281984%29009%
3C0002%3AOOTUOS%3E2.0.CO%3B2.

Voiland:1984:WHC

[136] Michael P. Voiland and Michael W. Duttweiler. Where’s the humanity? A challenge and opportunity for the fisheries community. Fisheries, 9 (4):10–12, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281984%29009%3C0010%3AWTHACA%3E2.0.CO%3B2.

Parker:1984:TIA

[137] Nick C. Parker. Technological innovations in aquaculture. Fisheries, 9 (4):13–16, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281984%29009%3C0013%3ATIIA%3E2.0.CO%3B2.

Anonymous:1984:FIAe

[138] Anonymous. Full issue article volume 9, issue 5. Fisheries, 9(5):1–46, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-9-5.

Hunter:1984:MPS

[139] Richard G. Hunter. Managerial professionalism in state fish and wildlife agencies: a survey of duties, attitudes, and needs. Fisheries, 9(5): 2–7, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281984%29009%3C0002%3AMPISFA%3E2.0.CO%3B2.

Richkus:1984:CSC

[140] William A. Richkus. Considerations in the selection of commercial analytical laboratories. Fisheries, 9(5):12–15, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281984%29009%
3C0012%3ACITSOC%3E2.0.CO%3B2.

Manny:1984:BRP

[141] Bruce A. Manny. Book review: Potential Impacts of Water Diversions on Fishery Resources in the Great Lakes: Contribution No. 610 of the Great Lakes Fishery Laboratory, U.S. Fish and Wildlife Service, 1451 Green Road, Ann Arbor, Michigan 48105. Fisheries, 9(5): 19–23, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281984%29009%3C0019%3APIOWDO%3E2.0.CO%3B2.

Anonymous:1984:FIAf

[142] Anonymous. Full issue article volume 9, issue 6. Fisheries, 9(6):1–58, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-9-6.

Layher:1984:CMU

[143] William G. Layher. Compatibility of multiple uses: Potable water supplies and fisheries. Fisheries, 9(6):2–11, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281984%29009%
3C0002%3ACOMU%3E2.0.CO%3B2.

Herke:1984:CEN

[144] William H. Herke and Barton D. Rogers. Comprehensive estuarine nursery study completed. Fisheries, 9(6):12–16, 1984. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281984%29009%
3C0012%3ACENSC%3E2.0.CO%3B2.

Anonymous:1985:FIAa

[145] Anonymous. Full issue article volume 10, issue 1. Fisheries, 10(1):1–48, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-10-1.

Boileau:1985:EWP

[146] M. G. Boileau. The expansion of white perch, Morone americana, in the Lower Great Lakes. Fisheries, 10(1):6–10, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281985%29010%
3C0006%3ATEOWPM%3E2.0.CO%3B2.

Gunning:1985:RPR

[147] Gerald E. Gunning and Royal D. Suttkus. Reclamation of the Pearl River: a perspective of unpolluted versus polluted waters. Fisheries, 10 (1):14–16, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281985%29010%3C0014%3AROTPR%3E2.0.CO%3B2.

Anonymous:1985:FIAb

[148] Anonymous. Full issue article volume 10, issue 2. Fisheries, 10(2):1–46, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-10-2.

Larimore:1985:AHU

[149] R. Weldon Larimore and Dwight D. Garrels. Assessing habitats used by warmwater stream fishes. Fisheries, 10(2):10–16, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281985%29010%
3C0010%3AAHUBWS%3E2.0.CO%3B2.

Oglesby:1985:MLF

[150] Ray T. Oglesby. Management of lacustrine fisheries in the tropics. Fisheries, 10(2):16–19, 1985. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281985%29010%3C0016%3AMOLFIT%3E2.0.CO%3B2.

Anonymous:1985:FIAc

[151] Anonymous. Full issue article volume 10, issue 3. Fisheries, 10(3):1–55, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-10-3.

Netsch:1985:WTA

[152] Norval Netsch and Robert E. Putz. Wild trout in Alaska-now and future. Fisheries, 10(3):8–12, 1985. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281985%29010%3C0008%3AWTIAAF%3E2.0.CO%3B2.

Anonymous:1985:FIAd

[153] Anonymous. Full issue article volume 10, issue 4. Fisheries, 10(4):1–46, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-10-4.

Marsh:1985:RST

[154] John H. Marsh and James H. Johnson. The role of Stevens Treaty Tribes in the management of anadromous fish runs in the Columbia Basin. Fisheries, 10(4):2–5, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281985%29010%3C0002%3ATROSTT%3E2.0.CO%3B2.

Gale:1985:FWY

[155] Richard P. Gale. Fisherman, would you rather be a logger?: Federal resource regulation in the Pacific Northwest. Fisheries, 10(4): 6–8, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281985%29010%3C0006%3AFWYRBA%3E2.0.CO%3B2.

Anonymous:1985:FIAe

[156] Anonymous. Full issue article volume 10, issue 5. Fisheries, 10(5):1–46, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-10-5.

Benke:1985:ISH

[157] A. C. Benke, R. L. Henry III, D. M. Gillespie, and R. J. Hunter. Importance of snag habitat for animal production in southeastern streams. Fisheries, 10(5):8–13, 1985. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281985%29010%3C0008%3AIOSHFA%3E2.0.CO%3B2.

Anonymous:1985:FIAf

[158] Anonymous. Full issue article volume 10, issue 6. Fisheries, 10(6):1–64, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-10-6.

Marking:1985:BAN

[159] Leif L. Marking and Fred P. Meyer. Are better anesthetics needed in fisheries? Fisheries, 10(6):2–5, 1985. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281985%29010%3C0002%3AABANIF%3E2.0.CO%3B2.

Kircheis:1985:SFM

[160] Frederick W. Kircheis. Sport fishery management of a lake containing lake trout and an unusual dwarf Arctic char. Fisheries, 10(6): 6–8, 1985. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281985%29010%3C0006%3ASFMOAL%3E2.0.CO%3B2.

Anonymous:1986:FIAa

[161] Anonymous. Full issue article volume 11, issue 1. Fisheries, 11(1):1–72, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-11-1.

Meffe:1986:CGM

[162] Gary K. Meffe. Conservation genetics and the management of endangered fishes. Fisheries, 11(1):14–23, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281986%29011%3C0014%3ACGATMO%3E2.
0.CO%3B2.

Mills:1986:CPZ

[163] Edward L. Mills and John L. Confer. Computer processing of zooplankton — application in fisheries studies. Fisheries, 11(1):24–27, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281986%29011%3C0024%3ACPOZIF%3E2.0.CO%3B2.

Anonymous:1986:FIAb

[164] Anonymous. Full issue article volume 11, issue 2. Fisheries, 11(2):1–62, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-11-2.

Welcomme:1986:IMC

[165] R. L. Welcomme. International measures for the control of introductions of aquatic organisms. Fisheries, 11(2):4–9, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281986%29011%
3C0004%3AIMFTCO%3E2.0.CO%3B2.

Sindermann:1986:SRR

[166] Carl J. Sindermann. Strategies for reducing risks from introductions of aquatic organisms: a marine perspective. Fisheries, 11(2):10– 15, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281986%29011%3C0010%3ASFRRFI%3E2.0.CO%3B2.

Shelton:1986:SRR

[167] William L. Shelton. Strategies for reducing risks from introductions of aquatic organisms: an aquaculture perspective. Fisheries, 11(2): 16–19, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281986%29011%3C0016%3ASFRRFI%3E2.0.CO%3B2.

Shafland:1986:RFE

[168] Paul L. Shafland. A review of Florida’s efforts to regulate, assess and manage exotic fishes: Contribution number 41, Non-Native Fish Research Laboratory, Florida Game and Fresh Water Fish Commission, Boca Raton, Florida 33431. Fisheries, 11(2):20–25, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281986%29011%3C0020%3AAROFET%3E2.0.CO%3B2.

Clugston:1986:SRR

[169] James P. Clugston. Strategies for reducing risks from introductions of aquatic organisms: The Federal perspective. Fisheries, 11(2):26– 29, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281986%29011%3C0026%3ASFRRFI%3E2.0.CO%3B2.

Courtenay:1986:SRR

[170] Walter R. Courtenay, Jr. and Jeffrey N. Taylor. Strategies for reducing risks from introductions of aquatic organisms: a philosophical perspective. Fisheries, 11(2):30–33, 1986. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281986%29011%3C0030%3ASFRRFI%3E2.0.CO%3B2.

Anonymous:1986:FIAc

[171] Anonymous. Full issue article volume 11, issue 3. Fisheries, 11(3):1–72, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-11-3.

Johnson:1986:COA

[172] Barry L. Johnson and Roy A. Stein. Competition for open-access resources: a class exercise that demonstrates the tragedy of the commons. Fisheries, 11(3):2–6, 1986. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281986%29011%3C0002%3ACFORAC%3E2.0.CO%3B2.

Willis:1986:KBB

[173] David W. Willis and Robert F. Hartmann. The Kansas Black Bass Tournament Monitoring Program. Fisheries, 11(3):7–10, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281986%29011%
3C0007%3ATKBBTM%3E2.0.CO%3B2.

Anonymous:1986:FIAd

[174] Anonymous. Full issue article volume 11, issue 4. Fisheries, 11(4):1–67, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-11-4.

Kennedy:1986:ECD

[175] James J. Kennedy. Early career development of Forest Service fisheries managers: This project was jointly funded by the U.S. Forest Service and the Utah Agricultural Experiment Station, MacIntire Stennis Project 712 (journal paper 3152). Fisheries, 11(4):8– 13, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281986%29011%3C0008%3AECDOFS%3E2.0.CO%3B2.

Smith:1986:LCF

[176] Courtland L. Smith. The life cycle of fisheries. Fisheries, 11(4):20– 25, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281986%29011%3C0020%3ATLCOF%3E2.0.CO%3B2.

Dadswell:1986:PIL

[177] Michael J. Dadswell, Roger A. Rulifson, and Graham R. Daborn. Potential impact of large-scale tidal power developments in the upper Bay of Fundy on fisheries resources of the Northwest Atlantic. Fisheries, 11 (4):26–35, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281986%29011%3C0026%3APIOLTP%3E2.0.CO%3B2.

Anonymous:1986:FIAe

[178] Anonymous. Full issue article volume 11, issue 5. Fisheries, 11(5):1–45, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-11-5.

HcHugh:1986:HCF

[179] J. L. HcHugh and David Conover. History and condition of food finfisheries in the Middle Atlantic region compared with other sections of the Coast. Fisheries, 11(5):8–13, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281986%29011%3C0008%3AHACOFF%3E2.
0.CO%3B2.

Anderson:1986:REI

[180] Randall S. Anderson, Cynthia J. Schwinden, and Jay A. Leitch. Regional economic impact of the Devils Lake fishery. Fisheries, 11(5): 14–17, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281986%29011%3C0014%3AREIOTD%3E2.0.CO%3B2.

Anonymous:1986:FIAf

[181] Anonymous. Full issue article volume 11, issue 6. Fisheries, 11(6):1–62, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-11-6.

George:1986:IAF

[182] Carl J. George, Robert A. Daniels, and Timothy J. Sinnott. The importance of archives: Fish illustrations of the 1926–39 Watershed Survey in New York State. Fisheries, 11(6):2–11, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281986%29011%3C0002%3ATIOAFI%
3E2.0.CO%3B2.

Pfund:1986:SFP

[183] Rose T. Pfund. A survey of fishery programs in colleges and universities in the United States. Fisheries, 11(6):16–21, 1986. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281986%29011%
3C0016%3AASOFPI%3E2.0.CO%3B2.

Anonymous:1987:FIAa

[184] Anonymous. Full issue article volume 12, issue 1. Fisheries, 12(1):1–48, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-12-1.

Brousseau:1987:RSL

[185] Christopher S. Brousseau and Edward R. Armstrong. The role of size limits in walleye management. Fisheries, 12(1):2–5, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281987%29012%
3C0002%3ATROSLI%3E2.0.CO%3B2.

Anonymous:1987:FIAb

[186] Anonymous. Full issue article volume 12, issue 2. Fisheries, 12(2):1–58, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-12-2.

Hesse:1987:TWM

[187] Larry W. Hesse. Taming the wild Missouri River: What has it cost? Fisheries, 12(2):2–9, 1987. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281987%29012%3C0002%3ATTWMRW%3E2.0.CO%3B2.

Quinn:1987:SUA

[188] Stephen P. Quinn. The status and usefulness of angler recognition programs in the United States. Fisheries, 12(2):10–16, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281987%29012%
3C0010%3ATSAUOA%3E2.0.CO%3B2.

Cross:1987:OIM

[189] David Cross. An opportunity for integrated management of the Flathead River–Lake Ecosystem, Montana. Fisheries, 12(2):17–22, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281987%29012%3C0017%3AAOFIMO%3E2.0.CO%3B2.

Anonymous:1987:FIAc

[190] Anonymous. Full issue article volume 12, issue 3. Fisheries, 12(3):1–64, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-12-3.

Schreiber:1987:APM

[191] R. Kent Schreiber and Douglas L. Britt. Acid precipitation mitigation: Implementation of the U.S. Fish and Wildlife Service Research Program. Fisheries, 12(3):2–6, 1987. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281987%29012%3C0002%3AAPMIOT%3E2.0.CO%3B2.

Brouha:1987:APP

[192] Paul Brouha. Agency policy and practices: Wildlife and fisheries management in the USDA Forest Service. Fisheries, 12(3):8–10, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281987%29012%3C0008%3AAPAPWA%3E2.0.CO%3B2.

Anonymous:1987:FIAd

[193] Anonymous. Full issue article volume 12, issue 4. Fisheries, 12(4):1–56, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-12-4.

Sousa:1987:RRR

[194] Robert J. Sousa, Fred P. Meyer, and Rosalie A. Schnick. Re-registration of rotenone: a State/Federal cooperative effort. Fisheries, 12(4):9– 13, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281987%29012%3C0009%3ARORAFC%3E2.0.CO%3B2.

Lichtkoppler:1987:OCC

[195] Frank R. Lichtkoppler, Leroy J. Hushak, David O. Kelch, and Fred L. Snyder. The 1985 Ohio Charter Captains Survey. Fisheries, 12(4): 14–18, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281987%29012%3C0014%3ATOCCS%3E2.0.CO%3B2.

Allen:1987:TGC

[196] Standish K. Allen, Jr. and Robert J. Wattendorf. Triploid grass carp: Status and management implications. Fisheries, 12(4):20–24, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281987%29012%3C0020%3ATGCSAM%3E2.0.CO%3B2.

Anonymous:1987:FIAe

[197] Anonymous. Full issue article volume 12, issue 5. Fisheries, 12(5):1–54, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-12-5.

Gale:1987:RMR

[198] Richard P. Gale. Resource miracles and rising expectations: a challenge to fishery managers. Fisheries, 12(5):8–13, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281987%29012%
3C0008%3ARMAREA%3E2.0.CO%3B2.

Autore:1987:ERP

[199] David E. Autore, William W. Shaw, and William J. Matter. The educational and recreational potential of fish hatcheries. Fisheries, 12(5): 14–16, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281987%29012%3C0014%3ATEARPO%3E2.0.CO%3B2.

Hearn:1987:ICH

[200] William E. Hearn. Interspecific competition and habitat segregation among stream-dwelling trout and salmon: a review. Fisheries, 12(5): 24–31, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281987%29012%3C0024%3AICAHSA%3E2.0.CO%3B2.

Anonymous:1987:FIAf

[201] Anonymous. Full issue article volume 12, issue 6. Fisheries, 12(6):1–71, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-12-6.

Barber:1987:FMS

[202] Willard E. Barber. The fisheries management structure and process under the MFCMA: a North Pacific perspective. Fisheries, 12(6):10– 17, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281987%29012%3C0010%3ATFMSAP%3E2.0.CO%3B2.

Jennings:1987:FMC

[203] Mark R. Jennings. Frederic Morton Chamberlain (1867–1921), pioneer fishery biologist of the American West. Fisheries, 12(6):22– 29, 1987. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281987%29012%3C0022%3AFMCPFB%3E2.0.CO%3B2.

Anonymous:1988:FIAa

[204] Anonymous. Full issue article volume 13, issue 1. Fisheries, 13(1):1–56, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-13-1.

Deacon:1988:EWW

[205] James E. Deacon. The endangered woundfin and water management in the Virgin River, Utah, Arizona, Nevada. Fisheries, 13(1):18– 24, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281988%29013%3C0018%3ATEWAWM%3E2.0.CO%3B2.

Matlock:1988:IFC

[206] Gary C. Matlock, Gary E. Saul, and C. E. Bryan. Importance of fish consumption to sport fishermen. Fisheries, 13(1):25–26, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281988%29013%3C0025%3AIOFCTS%3E2.0.CO%3B2.

Anonymous:1988:FIAb

[207] Anonymous. Full issue article volume 13, issue 2. Fisheries, 13(2):1–62, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-13-2.

Hartzler:1988:CTF

[208] James R. Hartzler. Catchable trout fisheries: The need for assessment. Fisheries, 13(2):2–8, 1988. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281988%29013%3C0002%3ACTFTNF%3E2.0.CO%3B2.

Kelso:1988:UFP

[209] William E. Kelso and Brian R. Murphy. University fisheries programs in the United States: Structure and needs. Fisheries, 13(2): 9–15, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281988%29013%3C0009%3AUFPITU%3E2.0.CO%3B2.

Anonymous:1988:FIAc

[210] Anonymous. Full issue article volume 13, issue 3. Fisheries, 13(3):1–80, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-13-3.

Meisner:1988:RGI

[211] J. D. Meisner, J. S. Rosenfeld, and H. A. Regier. The role of groundwater in the impact of climate warming on stream salmonines. Fisheries, 13(3):2–8, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281988%29013%3C0002%3ATROGIT%3E2.0.CO%3B2.

Dumont:1988:ISW

[212] Pierre Dumont, Jacques F. Bergeron, Pierre Dulude, Yves Mailhot, Aubin Rouleau, Grégoire Ouellet, and Jean-Pierre Lebel. Introduced salmonids: Where are they going in Quebec watersheds of the Saint-Laurent River? Fisheries, 13(3):9–17, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281988%29013%3C0009%3AISWATG%
3E2.0.CO%3B2.

Caturano:1988:SMS

[213] Suzanne Caturano, Lori S. Glanz, David C. Smith, Leonidas Tsomides, and John R. Moring. Shellfish mariculture: The status of mussel power in Maine. Fisheries, 13(3):18–21, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281988%29013%3C0018%3ASMTSOM%3E2.
0.CO%3B2.

Anonymous:1988:FIAd

[214] Anonymous. Full issue article volume 13, issue 4. Fisheries, 13(4):1–54, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-13-4.

Smith:1988:MFT

[215] C. Lavett Smith. Minnows first, then trout. Fisheries, 13(4):4– 8, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281988%29013%3C0004%3AMFTT%3E2.0.CO%3B2.

Thorne:1988:SIR

[216] Richard E. Thorne. Some impacts of remote system technology on future fisheries management: Contribution no. 740, School of Fisheries, University of Washington, Seattle, Washington 98195. Fisheries, 13(4): 14–17, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281988%29013%3C0014%3ASIORST%3E2.0.CO%3B2.

Anonymous:1988:FIAe

[217] Anonymous. Full issue article volume 13, issue 5. Fisheries, 13(5):1–52, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-13-5.

Williams:1988:AFS

[218] Jack E. Williams, Donald W. Sada, and Cynthia Deacon Williams. American Fisheries Society guidelines for introductions of threatened and endangered fishes. Fisheries, 13(5):5–11, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281988%29013%3C0005%3AAFSGFI%
3E2.0.CO%3B2.

Miller:1988:RAI

[219] David L. Miller, Robert M. Hughes, James R. Karr, Paul M. Leonard, Peter B. Moyle, Lynn H. Schrader, Bruce A. Thompson, Robert A. Daniels, Kurt D. Fausch, Gary A. Fitzhugh, James R. Gammon, David B. Halliwell, Paul L. Angermeier, and Donald J. Orth. Regional applications of an index of biotic integrity for use in water resource management. Fisheries, 13(5):12–20, 1988. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281988%29013%3C0012%3ARAOAIO%3E2.0.CO%3B2.

Ables:1988:UFF

[220] Ernest D. Ables and Kerry P. Reese. USDA funding for fisheries research. Fisheries, 13(5):22–24, 1988. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281988%29013%3C0022%3AUFFFR%3E2.0.CO%3B2.

Anonymous:1988:FIAf

[221] Anonymous. Full issue article volume 13, issue 6. Fisheries, 13(6):1–54, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-13-6.

Kapuscinski:1988:FGI

[222] Anne R. Kapuscinski and David P. Philipp. Fisheries genetics: Issues and priorities for research and policy development. Fisheries, 13(6): 4–10, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281988%29013%3C0004%3AFGIAPF%3E2.0.CO%3B2.

Ashley:1988:POU

[223] Keith W. Ashley and Bobby Buff. Pond owner utilization of management recommendations made by state fishery personnel. Fisheries, 13 (6):12–14, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281988%29013%3C0012%3APOUOMR%3E2.0.CO%3B2.

Bayley:1988:MFR

[224] Peter B. Bayley and Robert F. Illyes. Measuring fish and recording data in the field and laboratory: a microcomputer application. Fisheries, 13 (6):15–18, 1988. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281988%29013%3C0015%3AMFARDI%3E2.0.CO%3B2.

Anonymous:1989:FIAa

[225] Anonymous. Full issue article volume 14, issue 1. Fisheries, 14(1):1–62, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-14-1.

Smith:1989:CSN

[226] Gerald R. Smith and Ralph F. Stearley. The classification and scientific names of rainbow and cutthroat trouts. Fisheries, 14(1):4– 10, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281989%29014%3C0004%3ATCASNO%3E2.0.CO%3B2.

Hesse:1989:MRM

[227] Larry W. Hesse, Gerald R. Chaffin, and Jerry Brabander. Missouri River mitigation: a system approach. Fisheries, 14(1):11–15, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281989%29014%3C0011%3AMRMASA%3E2.0.CO%3B2.

Anonymous:1989:FIAb

[228] Anonymous. Full issue article volume 14, issue 2. Fisheries, 14(2):1–62, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-14-2.

Scarnecchia:1989:HDA

[229] Dennis L. Scarnecchia. The history and development of Atlantic salmon management in Iceland. Fisheries, 14(2):14–21, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281989%29014%
3C0014%3ATHADOA%3E2.0.CO%3B2.

Reiser:1989:SIF

[230] Dudley W. Reiser, Thomas A. Wesche, and Christopher Estes. Status of instream flow legislation and practices in North America. Fisheries, 14 (2):22–29, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281989%29014%3C0022%3ASOIFLA%3E2.0.CO%3B2.

Anonymous:1989:FIAc

[231] Anonymous. Full issue article volume 14, issue 3. Fisheries, 14(3):1–76, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-14-3.

Jackson:1989:GHS

[232] Donald C. Jackson and John R. Jackson. A glimmer of hope for stream fisheries in Mississippi. Fisheries, 14(3):4–9, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281989%29014%
3C0004%3AAGOHFS%3E2.0.CO%3B2.

Hesthagen:1989:EFK

[233] Trygve Hesthagen. Episodic fish kills in an acidified salmon river in Southwestern Norway. Fisheries, 14(3):10–17, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281989%29014%
3C0010%3AEFKIAA%3E2.0.CO%3B2.

Anonymous:1989:FIAd

[234] Anonymous. Full issue article volume 14, issue 4. Fisheries, 14(4):1–44, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-14-4.

Falk:1989:PPM

[235] James M. Falk, Alan R. Graefe, and Robert B. Ditton. Patterns of participation and motivation among saltwater tournament anglers. Fisheries, 14(4):10–17, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281989%29014%3C0010%3APOPAMA%3E2.0.CO%3B2.

McGurrin:1989:AAA

[236] Joseph McGurrin. An assessment of Atlantic artificial reef development. Fisheries, 14(4):19–25, 1989. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281989%29014%3C0019%3AAAOAAR%3E2.0.CO%3B2.

Anonymous:1989:FIAe

[237] Anonymous. Full issue article volume 14, issue 5. Fisheries, 14(5):1–48, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-14-5.

Berra:1989:BEB

[238] Tim M. Berra and Gerald R. Allen. Burrowing, emergence, behavior, and functional morphology of the Australian salamanderfish, Lepidogalaxias salamandroides. Fisheries, 14(5):2–10, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281989%29014%3C0002%3ABEBAFM%
3E2.0.CO%3B2.

Lamb:1989:CTD

[239] Berton L. Lamb. Comprehensive technologies and decision-making: Reflections on the instream flow incremental methodology. Fisheries, 14 (5):12–16, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281989%29014%3C0012%3ACTADRO%3E2.0.CO%3B2.

Oglesby:1989:UFE

[240] Ray T. Oglesby and Charles C. Krueger. Undergraduate fisheries education: Technical specialization or broad foundation? Fisheries, 14(5): 17–21, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281989%29014%3C0017%3AUFETSO%3E2.0.CO%3B2.

Anonymous:1989:FIAf

[241] Anonymous. Full issue article volume 14, issue 6. Fisheries, 14(6):1–80, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-14-6.

Williams:1989:FNA

[242] Jack E. Williams, James E. Johnson, Dean A. Hendrickson, Salvador Contreras-Balderas, James D. Williams, Miguel Navarro-Mendoza, Don E. McAllister, and James E. Deacon. Fishes of North America endangered, threatened, or of special concern: 1989. Fisheries, 14(6): 2–20, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281989%29014%3C0002%3AFONAET%3E2.0.CO%3B2.

Miller:1989:ENA

[243] Robert R. Miller, James D. Williams, and Jack E. Williams. Extinctions of North American fishes during the past century. Fisheries, 14(6): 22–38, 1989. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281989%29014%3C0022%3AEONAFD%3E2.0.CO%3B2.

Anonymous:1990:FIAa

[244] Anonymous. Full issue article volume 15, issue 1. Fisheries, 15(1):1–64, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-15-1.

Kapuscinski:1990:TFP

[245] Anne R. Kapuscinski and Eric M. Hallerman. Transgenic fish and public policy: Anticipating environmental impacts of transgenic fish. Fisheries, 15(1):2–11, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0002%3ATFAPPA%3E2.0.CO%3B2.

Hallerman:1990:TFPa

[246] Eric M. Hallerman and Anne R. Kapuscinski. Transgenic fish and public policy: Regulatory concerns. Fisheries, 15(1):12–20, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0012%3ATFAPPR%3E2.0.CO%3B2.

Hallerman:1990:TFPb

[247] Eric M. Hallerman and Anne R. Kapuscinski. Transgenic fish and public policy: Patenting of transgenic fish. Fisheries, 15(1):21–24, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0021%3ATFAPPP%3E2.0.CO%3B2.

Anonymous:1990:FIAb

[248] Anonymous. Full issue article volume 15, issue 2. Fisheries, 15(2):1–54, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-15-2.

Serchuk:1990:EFM

[249] Fredric M. Serchuk and Ronald J. Smolowitz. Ensuring fisheries management dysfunction: The neglect of science and technology. Fisheries, 15(2):4–7, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0004%3AEFMDTN%3E2.0.CO%3B2.

Hubert:1990:CIF

[250] Wayne A. Hubert, Catherine Raley, and Stanley H. Anderson. Compliance with instream flow agreements in Colorado, Montana, and Wyoming. Fisheries, 15(2):8–10, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281990%29015%3C0008%3ACWIFAI%3E2.
0.CO%3B2.

Anonymous:1990:FIAc

[251] Anonymous. Full issue article volume 15, issue 3. Fisheries, 15(3):1–72, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-15-3.

Railsback:1990:IEF

[252] Steven F. Railsback, Charles C. Coutant, and Michael J. Sale. Improving the effectiveness of fisheries agencies in developing hydropower mitigation. Fisheries, 15(3):3–8, 1990. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281990%29015%3C0003%3AITEOFA%3E2.0.CO%3B2.

Kennedy:1990:RME

[253] James J. Kennedy and Brett B. Roper. The role of mentoring in early career development of Forest Service fisheries biologists. Fisheries, 15 (3):9–13, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0009%3ATROMIE%3E2.0.CO%3B2.

Anonymous:1990:FIAd

[254] Anonymous. Full issue article volume 15, issue 4. Fisheries, 15(4):1–64, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-15-4.

Manci:1990:PCA

[255] William E. Manci. Private consultants in aquaculture and information transfer. Fisheries, 15(4):6–10, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281990%29015%3C0006%3APCIAAI%3E2.
0.CO%3B2.

Gorbman:1990:HJ

[256] Aubrey Gorbman, Hideshi Kobayashi, Yoshiharu Honma, and Michiya Matsuyama. The hagfishery of Japan. Fisheries, 15(4):12–18, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0012%3ATHOJ%3E2.0.CO%3B2.

Bennett:1990:PSW

[257] David H. Bennett and Thomas J. McArthur. Predicting success of walleye stocking programs in the United States and Canada. Fisheries, 15 (4):19–23, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0019%3APSOWSP%3E2.0.CO%3B2.

Anonymous:1990:FIAe

[258] Anonymous. Full issue article volume 15, issue 5. Fisheries, 15(5):1–48, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-15-5.

Wilbur:1990:GLP

[259] Robert L. Wilbur. Gray literature: a professional dilemma. Fisheries, 15(5):2–6, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0002%3AGLAPD%3E2.0.CO%3B2.

Miller:1990:PWM

[260] Robert J. Miller. Properties of a well-managed nearshore fishery. Fisheries, 15(5):7–12, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281990%29015%3C0007%3APOAWNF%3E2.0.CO%3B2.

Marcus:1990:RSH

[261] Michael D. Marcus, Lynn E. Noel, and Michael K. Young. Rating salmonid habitat research needs in the Central Rocky Mountains. Fisheries, 15(5):14–18, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/abs/10.
1577/1548-8446%281990%29015%3C0014%3ARSHRNI%3E2.0.CO%3B2.

Waples:1990:GAM

[262] Robin S. Waples, Gary A. Winans, Fred M. Utter, and Conrad Mahnken. Genetic approaches to the management of Pacific salmon. Fisheries, 15 (5):19–25, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0019%3AGATTMO%3E2.0.CO%3B2.

Anonymous:1990:FIAf

[263] Anonymous. Full issue article volume 15, issue 6. Fisheries, 15(6):1–72, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-15-6.

Smith:1990:GRC

[264] Joel B. Smith. From global to regional climate change: Relative knowns and unknowns about global warming. Fisheries, 15(6):2– 6, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0002%3AFGTRCC%3E2.0.CO%3B2.

Francis:1990:CCM

[265] Robert C. Francis. Climate change and marine fisheries. Fisheries, 15 (6):7–9, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0007%3ACCAMF%3E2.0.CO%3B2.

Regier:1990:AEC

[266] Henry A. Regier and J. Donald Meisner. Anticipated effects of climate change on freshwater fishes and their habitat. Fisheries, 15(6):

10–15, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0010%3AAEOCCO%3E2.0.CO%3B2.

Kennedy:1990:AEC

[267] Victor S. Kennedy. Anticipated effects of climate change on estuarine and coastal fisheries. Fisheries, 15(6):16–24, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281990%29015%3C0016%3AAEOCCO%
3E2.0.CO%3B2.

Matthews:1990:PEG

[268] William J. Matthews and Earl G. Zimmerman. Potential effects of global warming on native fishes of the Southern Great Plains and the Southwest. Fisheries, 15(6):26–32, 1990. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281990%29015%3C0026%3APEOGWO%3E2.0.CO%3B2.

Gucinski:1990:GCC

[269] Hermann Gucinski, Robert T. Lackey, and Brian C. Spence. Global climate change: Policy implications for fisheries. Fisheries, 15(6):33– 38, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0033%3AGCCPIF%3E2.0.CO%3B2.

Glantz:1990:DHF

[270] Michael H. Glantz. Does history have a future? Forecasting climate change effects on fisheries by analogy. Fisheries, 15(6):39– 44, 1990. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281990%29015%3C0039%3ADHHAFF%3E2.0.CO%3B2.

Anonymous:1991:FIAa

[271] Anonymous. Full issue article volume 16, issue 1. Fisheries, 16(1):1–56, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-16-1.

Anonymous:1991:FIAb

[272] Anonymous. Full issue article volume 16, issue 2. Fisheries, 16(2):1–64, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-16-2.

Nehlsen:1991:PSC

[273] Willa Nehlsen, Jack E. Williams, and James A. Lichatowich. Pacific salmon at the crossroads: Stocks at risk from California, Oregon, Idaho, and Washington. Fisheries, 16(2):4–21, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281991%29016%3C0004%3APSATCS%
3E2.0.CO%3B2.

Murphy:1991:RWI

[274] Brian R. Murphy, David W. Willis, and Timothy A. Springer. The relative weight index in fisheries management: Status and needs. Fisheries, 16(2):30–38, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281991%29016%3C0030%3ATRWIIF%3E2.0.CO%3B2.

Anonymous:1991:FIAc

[275] Anonymous. Full issue article volume 16, issue 3. Fisheries, 16(3):1–79, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-16-3.

Schramm:1991:SCS

[276] Harold L. Schramm, Jr., Michael L. Armstrong, Nicholas A. Funicelli, David M. Green, Dennis P. Lee, Ralph E. Manns, Jr., Bruce D. Taubert, and Stephen J. Waters. The status of competitive sport fishing in North America. Fisheries, 16(3):4–12, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281991%29016%3C0004%3ATSOCSF%3E2.
0.CO%3B2.

Schramm:1991:SEB

[277] Harold L. Schramm, Jr., Michael L. Armstrong, Anthony J. Fedler, Nicholas A. Funicelli, David M. Green, Jeffery L. Hahn, Dennis P. Lee, Ralph E. Manns, Jr., Stephen P. Quinn, and Stephen J. Waters. Sociological, economic, and biological aspects of competitive fishing. Fisheries, 16 (3):13–21, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281991%29016%3C0013%3ASEABAO%3E2.0.CO%3B2.

Hinch:1991:SLS

[278] Scott G. Hinch. Small- and large-scale studies in fisheries ecology: The need for cooperation among researchers. Fisheries, 16(3):22– 27, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281991%29016%3C0022%3ASALSIF%3E2.0.CO%3B2.

Osborne:1991:SHA

[279] L. L. Osborne, B. Dickson, M. Ebbers, R. Ford, J. Lyons, D. Kline, E. Rankin, D. Ross, R. Sauer, P. Seelbach, C. Speas, T. Stefanavage, J. Waite, and S. Walker. Stream habitat assessment programs in states of the AFS North Central Division. Fisheries, 16(3):28– 35, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281991%29016%3C0028%3ASHAPIS%3E2.0.CO%3B2.

Anonymous:1991:FIAd

[280] Anonymous. Full issue article volume 16, issue 4. Fisheries, 16(4):1–64, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-16-4.

Buttner:1991:AAP

[281] Joseph K. Buttner, David B. MacNeill, David M. Green, and Richard T. Colesante. Angler associations as partners in walleye management. Fisheries, 16(4):12–17, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/abs/10.
1577/1548-8446%281991%29016%3C0012%3AAAAPIW%3E2.0.CO%3B2.

Keen:1991:OPM

[282] Elmer A. Keen. Ownership and productivity of marine fishery resources. Fisheries, 16(4):18–22, 1991. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281991%29016%3C0018%3AOAPOMF%3E2.0.CO%3B2.

Anonymous:1991:FIAe

[283] Anonymous. Full issue article volume 16, issue 5. Fisheries, 16(5):1–72, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-16-5.

Murawski:1991:CWM

[284] Steven A. Murawski. Can we manage our multispecies fisheries? Fisheries, 16(5):5–13, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281991%29016%3C0005%3ACWMOMF%3E2.0.CO%3B2.

Momot:1991:PEF

[285] Walter T. Momot. Potential for exploitation of freshwater crayfish in coolwater systems: Management guidelines and issues. Fisheries, 16(5):

14–21, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281991%29016%3C0014%3APFEOFC%3E2.0.CO%3B2.

Chittenden:1991:OPS

[286] Mark E. Chittenden, Jr. Operational procedures and sampling in the Chesapeake Bay pound-net fishery. Fisheries, 16(5):22–27, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281991%29016%3C0022%3AOPASIT%3E2.0.CO%3B2.

Clune:1991:YKF

[287] Thomas Clune and Dennis Dauble. The Yakima/Klickitat Fisheries Project: a strategy for supplementation of anadromous salmonids. Fisheries, 16(5):28–34, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/abs/10.
1577/1548-8446%281991%29016%3C0028%3ATKFPAS%3E2.0.CO%3B2.

Armour:1991:EIF

[288] Carl L. Armour and Jonathan G. Taylor. Evaluation of the instream flow incremental methodology by U.S. Fish and Wildlife Service Field Users. Fisheries, 16(5):36–43, 1991. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281991%29016%3C0036%3AEOTIFI%3E2.0.CO%3B2.

Anonymous:1991:FIAf

[289] Anonymous. Full issue article volume 16, issue 6. Fisheries, 16(6):1–56, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-16-6.

Reinert:1991:RAR

[290] Robert E. Reinert, Barbara A. Knuth, Michael A. Kamrin, and Quentin J. Stober. Risk assessment, risk management, and fish consumption advisories in the United States. Fisheries, 16(6):5–12, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281991%29016%3C0005%3ARARMAF%3E2.0.CO%3B2.

Green:1991:IFC

[291] Trellis G. Green. Importance of fish consumption to sport fishermen: an economic analysis. Fisheries, 16(6):13–19, 1991. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281991%29016%
3C0013%3AIOFCTS%3E2.0.CO%3B2.

Anonymous:1992:FIAa

[292] Anonymous. Full issue article volume 17, issue 1. Fisheries, 17(1):1–64, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-17-1.

Hilborn:1992:HFS

[293] Ray Hilborn. Hatcheries and the future of salmon in the Northwest. Fisheries, 17(1):5–8, 1992. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281992%29017%3C0005%3AHATFOS%3E2.0.CO%3B2.

Rulifson:1992:FCR

[294] Roger A. Rulifson, James D. Murray, and James J. Bahen. Finfish catch reduction in South Atlantic shrimp trawls using three designs of by-catch reduction devices. Fisheries, 17(1):9–20, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281992%29017%3C0009%3AFCRISA%
3E2.0.CO%3B2.

Murray:1992:MCC

[295] James D. Murray, James J. Bahen, and Roger A. Rulifson. Management considerations for by-catch in the North Carolina and Southeast shrimp fishery. Fisheries, 17(1):21–26, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281992%29017%3C0021%3AMCFBIT%3E2.
0.CO%3B2.

Carline:1992:CFC

[296] Robert F. Carline, William E. Sharpe, and Charles J. Gagen. Changes in fish communities and trout management in response to acidification of streams in Pennsylvania. Fisheries, 17(1):33–38, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281992%29017%
3C0033%3ACIFCAT%3E2.0.CO%3B2.

Anonymous:1992:FIAb

[297] Anonymous. Full issue article volume 17, issue 2. Fisheries, 17(2):1–64, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-17-2.

Murdock:1992:DCU

[298] Steve H. Murdock, Kenneth Backman, Robert B. Ditton, Md Nazrul Hoque, and David Ellis. Demographic change in the United States in the 1990s and the twenty-first century: Implications for fisheries management. Fisheries, 17(2):6–13, 1992. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281992%29017%3C0006%3ADCITUS%3E2.0.CO%3B2.

Meador:1992:IBW

[299] Michael R. Meador. Inter-basin water transfer: Ecological concerns. Fisheries, 17(2):17–22, 1992. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281992%29017%3C0017%3AIWTEC%3E2.0.CO%3B2.

Brandt:1992:SEM

[300] Stephen B. Brandt, Doran M. Mason, and E. Vincent Patrick. Spatially-explicit models of fish growth rate. Fisheries, 17(2):23– 35, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281992%29017%3C0023%3ASMOFGR%3E2.0.CO%3B2.

Kane:1992:ETW

[301] Andrew S. Kane, Renate Reimschuessel, and Michael M. Lipsky. Effect of tadpoles on warmwater fish pond production. Fisheries, 17(2): 36–39, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281992%29017%3C0036%3AEOTOWF%3E2.0.CO%3B2.

Anonymous:1992:FIAc

[302] Anonymous. Full issue article volume 17, issue 3. Fisheries, 17(3):1–60, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-17-3.

Cairns:1992:BMN

[303] Michael A. Cairns and Robert T. Lackey. Biodiversity and management of natural resources: The issues. Fisheries, 17(3):6–10, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281992%29017%3C0006%3ABAMONR%3E2.0.CO%3B2.

Hughes:1992:BDB

[304] Robert M. Hughes and Reed F. Noss. Biological diversity and biological integrity: Current concerns for lakes and streams. Fisheries, 17(3): 11–19, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281992%29017%3C0011%3ABAMONR%3E2.0.CO%3B2.

Upton:1992:BCM

[305] Harold F. Upton. Biodiversity and conservation of the marine environment. Fisheries, 17(3):20–25, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281992%29017%3C0020%3ABACOTM%3E2.
0.CO%3B2.

Blockstein:1992:APU

[306] David E. Blockstein. An aquatic perspective on U.S. biodiversity policy. Fisheries, 17(3):26–30, 1992. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281992%29017%3C0026%3AAAPOUB%3E2.0.CO%3B2.

Titus:1992:BNN

[307] Timothy R. Titus. Biodiversity: The need for a national policy. Fisheries, 17(3):31–34, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281992%29017%3C0031%3ABTNFAN%3E2.0.CO%3B2.

Schweitzer:1992:CBD

[308] Jeff Schweitzer. Conserving biodiversity in developing countries. Fisheries, 17(3):35–38, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/abs/10.
1577/1548-8446%281992%29017%3C0035%3ACBIDC%3E2.0.CO%3B2.

Anonymous:1992:FIAd

[309] Anonymous. Full issue article volume 17, issue 4. Fisheries, 17(4):1–62, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-17-4.

Hilborn:1992:CFA

[310] Ray Hilborn. Can fisheries agencies learn from experience? Fisheries, 17 (4):6–14, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281992%29017%3C0006%3ACFALFE%3E2.0.CO%3B2.

Bergman:1992:PDU

[311] Peter K. Bergman, Frank Haw, H. Lee Blankenship, and Raymond M. Buckley. Perspectives on design, use, and misuse of fish tags. Fisheries, 17 (4):20–25, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281992%29017%3C0020%3APODUAM%3E2.0.CO%3B2.

Anonymous:1992:FIAe

[312] Anonymous. Full issue article volume 17, issue 5. Fisheries, 17(5):1–64, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-17-5.

Scarnecchia:1992:RGB

[313] Dennis L. Scarnecchia. A reappraisal of gars and bowfins in fishery management. Fisheries, 17(5):6–12, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281992%29017%3C0006%3AAROGAB%3E2.
0.CO%3B2.

Gale:1992:TFM

[314] Richard P. Gale. Is there a fisheries management revolution in your future? Fisheries, 17(5):14–19, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281992%29017%3C0014%3AITAFMR%3E2.
0.CO%3B2.

Weigmann:1992:CAP

[315] Diana L. Weigmann and Louis A. Helfrich. Content analysis of the pond management literature. Fisheries, 17(5):27–34, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281992%29017%
3C0027%3ACAOTPM%3E2.0.CO%3B2.

Anonymous:1992:FIAf

[316] Anonymous. Full issue article volume 17, issue 6. Fisheries, 17(6):1–62, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-17-6.

Marking:1992:ETC

[317] Leif L. Marking. Evaluation of toxicants for the control of carp and other nuisance fishes. Fisheries, 17(6):6–13, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281992%29017%3C0006%3AEOTFTC%
3E2.0.CO%3B2.

Johnson:1992:TCS

[318] Barry L. Johnson. Transfer of computer software technology through workshops: The case of fish bioenergetics modeling. Fisheries, 17(6):

14–18, 1992. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281992%29017%3C0014%3ATOCSTT%3E2.0.CO%3B2.

Anonymous:1993:FIAa

[319] Anonymous. Full issue article volume 18, issue 1. Fisheries, 18(1):1–48, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-1.

Habera:1993:WTR

[320] James W. Habera and Richard J. Strange. Wild trout resources and management in the Southern Appalachian Mountains. Fisheries, 18(1): 6–13, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0006%3AWTRAMI%3E2.0.CO%3B2.

Lichtkoppler:1993:OLE

[321] Frank R. Lichtkoppler and Leroy J. Hushak. Ohio’s 1990 Lake Erie charter fishing industry. Fisheries, 18(1):14–21, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281993%29018%
3C0014%3AOLECFI%3E2.0.CO%3B2.

Anonymous:1993:FIAb

[322] Anonymous. Full issue article volume 18, issue 2. Fisheries, 18(2):1–40, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-2.

Bernard:1993:OPD

[323] David R. Bernard, William D. Arvey, and Rolland A. Holmes. Operational planning: The Dall River and rescue of its sport fishery. Fisheries, 18(2):6–12, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0006%3AOPTDRA%3E2.0.CO%3B2.

Loomis:1993:DJF

[324] David K. Loomis and Robert B. Ditton. Distributive justice in fisheries management. Fisheries, 18(2):14–18, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281993%29018%
3C0014%3ADJIFM%3E2.0.CO%3B2.

Anonymous:1993:FIAc

[325] Anonymous. Full issue article volume 18, issue 3. Fisheries, 18(3):1–44, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-3.

Anthony:1993:SGR

[326] Vaughn C. Anthony. The state of groundfish resources off the Northeastern United States. Fisheries, 18(3):12–17, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281993%29018%
3C0012%3ATSOGRO%3E2.0.CO%3B2.

Wespestad:1993:SBS

[327] Vidar G. Wespestad. The status of Bering sea Pollock and the effect of the “Donut hole” fishery. Fisheries, 18(3):18–24, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281993%29018%
3C0018%3ATSOBSP%3E2.0.CO%3B2.

Anonymous:1993:FIAd

[328] Anonymous. Full issue article volume 18, issue 4. Fisheries, 18(4):1–46, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-4.

Dawson:1993:SSC

[329] Chad P. Dawson, Nancy A. Connelly, and Tommy L. Brown. Salmon snagging controversy: New York’s Salmon River. Fisheries, 18(4):6– 10, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0006%3ASSCNYS%3E2.0.CO%3B2.

Adams:1993:EDA

[330] Clark E. Adams, John K. Thomas, and W. Roy Knowles, Jr. Explaining differences in angling rates in the United States. Fisheries, 18(4): 11–17, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0011%3AEDIARI%3E2.0.CO%3B2.

Harr:1993:SFR

[331] R. Dennis Harr and Roger A. Nichols. Stabilizing forest roads to help restore fish habitats: a Northwest Washington example. Fisheries, 18 (4):18–22, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0018%3ASFRTHR%3E2.0.CO%3B2.

Anonymous:1993:FIAe

[332] Anonymous. Full issue article volume 18, issue 5. Fisheries, 18(5):1–46, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-5.

Hesse:1993:MRH

[333] Larry W. Hesse and Wes Sheets. The Missouri River hydrosystem. Fisheries, 18(5):5–14, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281993%29018%3C0005%3ATMRH%3E2.0.CO%3B2.

Forsgren:1993:RGF

[334] Harv Forsgren and Andrew J. Loftus. Rising to a greater future: Forest Service fisheries program accountability. Fisheries, 18(5):15– 21, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0015%3ARTAGFF%3E2.0.CO%3B2.

Anonymous:1993:FIAf

[335] Anonymous. Full issue article volume 18, issue 6. Fisheries, 18(6):1–47, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-6.

Hawkins:1993:HAC

[336] Charles P. Hawkins, Jeffrey L. Kershner, Peter A. Bisson, Mason D. Bryant, Lynn M. Decker, Stanley V. Gregory, Dale A. McCullough, C. K. Overton, Gordon H. Reeves, Robert J. Steedman, and Michael K. Young. A hierarchical approach to classifying stream habitat features. Fisheries, 18(6):3–12, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0003%3AAHATCS%3E2.0.CO%3B2.

French:1993:HWC

[337] John R. P. French III. How well can fishes prey on zebra mussels in Eastern North America? Fisheries, 18(6):13–19, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281993%29018%
3C0013%3AHWCFPO%3E2.0.CO%3B2.

Anonymous:1993:FIAg

[338] Anonymous. Full issue article volume 18, issue 7. Fisheries, 18(7):1–48, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-7.

Cada:1993:SFP

[339] Glenn F. Cada and Michael J. Sale. Status of fish passage facilities at nonfederal hydropower projects. Fisheries, 18(7):4–12, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0004%3ASOFPFA%3E2.0.CO%3B2.

Rumsey:1993:FMA

[340] Gary L. Rumsey. Fish meal and alternate sources of protein in fish feeds update 1993. Fisheries, 18(7):14–19, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281993%29018%3C0014%3AFMAASO%3E2.
0.CO%3B2.

Anonymous:1993:FIAh

[341] Anonymous. Full issue article volume 18, issue 8. Fisheries, 18(8):1–52, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-8.

Lawson:1993:COP

[342] Peter W. Lawson. Cycles in ocean productivity, trends in habitat quality, and the restoration of salmon runs in Oregon. Fisheries, 18(8): 6–10, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0006%3ACIOPTI%3E2.0.CO%3B2.

Utter:1993:GMM

[343] Fred Utter and Nils Ryman. Genetic markers and mixed stock fisheries. Fisheries, 18(8):11–21, 1993. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281993%29018%3C0011%3AGMAMSF%3E2.0.CO%3B2.

Anonymous:1993:FIAi

[344] Anonymous. Full issue article volume 18, issue 9. Fisheries, 18(9):1–76, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-9.

Williams:1993:CSF

[345] James D. Williams, Melvin L. Warren, Jr., Kevin S. Cummings, John L. Harris, and Richard J. Neves. Conservation status of freshwater mussels of the United States and Canada. Fisheries, 18(9):6– 22, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0006%3ACSOFMO%3E2.0.CO%3B2.

Anonymous:1993:FIAj

[346] Anonymous. Full issue article volume 18, issue 10. Fisheries, 18(10):1–60, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-10.

Sissenwine:1993:MFC

[347] Michael P. Sissenwine and Andrew A. Rosenberg. Marine fisheries at a critical juncture. Fisheries, 18(10):6–14, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281993%29018%3C0006%3AMFAACJ%
3E2.0.CO%3B2.

Anonymous:1993:FIAk

[348] Anonymous. Full issue article volume 18, issue 11. Fisheries, 18(11):1–56, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-11.

Wood:1993:IEW

[349] Christopher A. Wood. Implementation and evaluation of the water budget. Fisheries, 18(11):6–17, 1993. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281993%29018%3C0006%3AIAEOTW%3E2.0.CO%3B2.

Anonymous:1993:FIAl

[350] Anonymous. Full issue article volume 18, issue 12. Fisheries, 18(12):1–64, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-18-12.

Litvak:1993:EFB

[351] Matthew K. Litvak and Nicholas E. Mandrak. Ecology of freshwater baitfish use in Canada and the United States. Fisheries, 18(12): 6–13, 1993. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281993%29018%3C0006%3AEOFBUI%3E2.0.CO%3B2.

Anonymous:1994:FIAa

[352] Anonymous. Full issue article volume 19, issue 1. Fisheries, 19(1):1–64, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-1.

Warren:1994:SFF

[353] Melvin L. Warren, Jr. and Brooks M. Burr. Status of freshwater fishes of the United States: Overview of an imperiled fauna. Fisheries, 19(1): 6–18, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281994%29019%3C0006%3ASOFFOT%3E2.0.CO%3B2.

Anonymous:1994:FIAb

[354] Anonymous. Full issue article volume 19, issue 2. Fisheries, 19(2):1–56, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-2.

Moyle:1994:PAB

[355] Peter B. Moyle and Ronald M. Yoshiyama. Protection of aquatic biodiversity in California: a five-tiered approach. Fisheries, 19(2):6– 18, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281994%29019%3C0006%3APOABIC%3E2.0.CO%3B2.

Smith:1994:CCB

[356] Courtland L. Smith. Connecting cultural and biological diversity in restoring northwest salmon. Fisheries, 19(2):20–26, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281994%29019%
3C0020%3ACCABDI%3E2.0.CO%3B2.

Anonymous:1994:FIAc

[357] Anonymous. Full issue article volume 19, issue 3. Fisheries, 19(3):1–64, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-3.

Hallerman:1994:ANE

[358] Eric M. Hallerman. Assessment of needs for education opportunities in fisheries genetics. Fisheries, 19(3):6–12, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281994%29019%3C0006%3AAONFEO%
3E2.0.CO%3B2.

Maceina:1994:USP

[359] Michael J. Maceina, Phillip W. Bettoli, and Dennis R. DeVries. Use of a split-plot analysis of variance design for repeated-measures fishery data. Fisheries, 19(3):14–20, 1994. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281994%29019%3C0014%3AUOASAO%3E2.0.CO%3B2.

Anonymous:1994:FIAd

[360] Anonymous. Full issue article volume 19, issue 4. Fisheries, 19(4):1–64, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-4.

Fedler:1994:UAM

[361] Anthony J. Fedler and Robert B. Ditton. Understanding angler motivations in fisheries management. Fisheries, 19(4):6–13, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281994%29019%
3C0006%3AUAMIFM%3E2.0.CO%3B2.

Anonymous:1994:FIAe

[362] Anonymous. Full issue article volume 19, issue 5. Fisheries, 19(5):1–64, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-5.

Stickney:1994:UHF

[363] Robert R. Stickney. Use of hatchery fish in enhancement programs. Fisheries, 19(5):6–13, 1994. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281994%29019%3C0006%3AUOHFIE%3E2.0.CO%3B2.

Cunningham:1994:NFC

[364] Patricia A. Cunningham, Susan L. Smith, John P. Tippett, and Alison Greene. A national fish consumption advisory data base: a step toward consistency. Fisheries, 19(5):14–23, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281994%29019%3C0014%3AANFCAD%3E2.
0.CO%3B2.

Anonymous:1994:FIAf

[365] Anonymous. Full issue article volume 19, issue 6. Fisheries, 19(6):1–88, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-6.

Dudley:1994:TWF

[366] Richard G. Dudley. Third world fisheries resources: Who cares? How can North American fisheries scientists help colleagues around the world? Fisheries, 19(6):6–11, 1994. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281994%29019%3C0006%3ATWFRWC%3E2.0.CO%3B2.

Austen:1994:IGC

[367] Douglas J. Austen, Peter B. Bayley, and Bruce W. Menzel. Importance of the guild concept to fisheries research and management. Fisheries, 19 (6):12–20, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281994%29019%3C0012%3AIOTGCT%3E2.0.CO%3B2.

Anonymous:1994:FIAg

[368] Anonymous. Full issue article volume 19, issue 7. Fisheries, 19(7):1–68, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-7.

Huntsman:1994:EMF

[369] Gene R. Huntsman. Endangered marine finfish: Neglected resources or beasts of fiction? Fisheries, 19(7):8–15, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281994%29019%3C0008%3AEMFNRO%
3E2.0.CO%3B2.

Wirgin:1994:WDC

[370] Isaac I. Wirgin and John R. Waldman. What DNA can do for you. Fisheries, 19(7):16–27, 1994. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281994%29019%3C0016%3AWDCDFY%3E2.0.CO%3B2.

Anonymous:1994:FIAh

[371] Anonymous. Full issue article volume 19, issue 8. Fisheries, 19(8):1–88, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-8.

Wunderlich:1994:RER

[372] Robert C. Wunderlich, Brian D. Winter, and John H. Meyer. Restoration of the Elwha River ecosystem. Fisheries, 19(8):11–19, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281994%29019%
3C0011%3AROTERE%3E2.0.CO%3B2.

Turgeon:1994:MSB

[373] Donna D. Turgeon. Managerial strategies for breaking through a “Glass ceiling”. Fisheries, 19(8):21–28, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281994%29019%3C0021%3AMSFBTA%3E2.
0.CO%3B2.

Anonymous:1994:FIAi

[374] Anonymous. Full issue article volume 19, issue 9. Fisheries, 19(9):1–55, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-9.

Armour:1994:ELG

[375] Carl Armour, Don Duff, and Wayne Elmore. The effects of livestock grazing on western riparian and stream ecosystem. Fisheries, 19(9): 9–12, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281994%29019%3C0009%3ATEOLGO%3E2.0.CO%3B2.

Anonymous:1994:FIAj

[376] Anonymous. Full issue article volume 19, issue 10. Fisheries, 19(10):1–72, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-10.

Schramm:1994:PUF

[377] Harold L. Schramm, Jr. and Gary B. Edwards. The perspectives on urban fisheries management: Results of a workshop. Fisheries, 19(10): 9–15, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281994%29019%3C0009%3ATPOUFM%3E2.0.CO%3B2.

McCain:1994:ISM

[378] Katherine W. McCain. Islands in the stream: Mapping the fisheries and aquatic sciences literatures. Fisheries, 19(10):20–27, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281994%29019%
3C0020%3AIITS%3E2.0.CO%3B2.

Anonymous:1994:FIAk

[379] Anonymous. Full issue article volume 19, issue 11. Fisheries, 19(11):1–64, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-11.

Schnute:1994:SAC

[380] Jon T. Schnute and Laura J. Richards. Stock assessment for the 21st century. Fisheries, 19(11):10–16, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281994%29019%3C0010%3ASAFTSC%3E2.
0.CO%3B2.

Adelman:1994:ENF

[381] Ira R. Adelman, David J. Schmidly, and Yosef Cohen. Educational needs of fisheries and wildlife professionals: Results of a survey. Fisheries, 19 (11):17–25, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281994%29019%3C0017%3AENOFAW%3E2.0.CO%3B2.

Anonymous:1994:FIAl

[382] Anonymous. Full issue article volume 19, issue 12. Fisheries, 19(12):1–64, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-19-12.

Hambright:1994:CAP

[383] K. David Hambright and Samuel Fridman. A computer-assisted plankton analysis system for the Macintosh. Fisheries, 19(12):6– 8, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281994%29019%3C0006%3AACPASF%3E2.0.CO%3B2.

Robertson:1994:PCE

[384] John M. Robertson, Dennis A. Conway, and Charles Pecor. Planning and conducting external program reviews. Fisheries, 19(12):12– 18, 1994. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281994%29019%3C0012%3APACEPR%3E2.0.CO%3B2.

Anonymous:1995:FIAa

[385] Anonymous. Full issue article volume 20, issue 1. Fisheries, 20(1):1–64, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-1.

Lichatowich:1995:ADT

[386] James Lichatowich, Lars Mobrand, Larry Lestelle, and Tom Vogel. An approach to the diagnosis and treatment of depleted Pacific salmon populations in Pacific Northwest watersheds. Fisheries, 20(1):10– 18, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281995%29020%3C0010%3AAATTDA%3E2.0.CO%3B2.

Collette:1995:IBF

[387] Bruce B. Collette and Michael Vecchione. Interactions between fisheries and systematics. Fisheries, 20(1):20–25, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281995%29020%
3C0020%3AIBFAS%3E2.0.CO%3B2.

Anonymous:1995:FIAb

[388] Anonymous. Full issue article volume 20, issue 2. Fisheries, 20(2):1–88, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-2.

Scalet:1995:AFW

[389] Charles G. Scalet and Ira R. Adelman. Accreditation of fisheries and wildlife programs. Fisheries, 20(2):8–13, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281995%29020%3C0008%3AAOFAWP%
3E2.0.CO%3B2.

Guy:1995:ILT

[390] Christopher S. Guy and Susan Denson-Guy. An introduction to learning and teaching styles: Making the match. Fisheries, 20(2):18– 20, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281995%29020%3C0018%3AAITLAT%3E2.0.CO%3B2.

Murphy:1995:UMI

[391] Brian R. Murphy. Using microthemes to improve students’ writing skills. Fisheries, 20(2):22–26, 1995. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281995%29020%3C0022%3AUMTISW%3E2.0.CO%3B2.

Anonymous:1995:FIAc

[392] Anonymous. Full issue article volume 20, issue 3. Fisheries, 20(3):1–68, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-3.

Hard:1995:SEF

[393] Jeffrey J. Hard. Science, education, and the fisheries scientist. Fisheries, 20(3):10–16, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281995%29020%3C0010%3ASEATFS%3E2.0.CO%3B2.

Engel:1995:EWF

[394] Sandy Engel. Eurasian watermilfoil as a fishery management tool. Fisheries, 20(3):20–27, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/abs/10.
1577/1548-8446%281995%29020%3C0020%3AEWAAFM%3E2.0.CO%3B2.

Anonymous:1995:FIAd

[395] Anonymous. Full issue article volume 20, issue 4. Fisheries, 20(4):1–64, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-4.

Eaton:1995:FIB

[396] J. G. Eaton, J. H. McCormick, B. E. Goodno, D. G. O’Brien, H. G. Stefany, M. Hondzo, and R. M. Scheller. A field information-based system for estimating fish temperature tolerances. Fisheries, 20(4): 10–18, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281995%29020%3C0010%3AAFISFE%3E2.0.CO%3B2.

Anonymous:1995:FIAe

[397] Anonymous. Full issue article volume 20, issue 5. Fisheries, 20(5):1–56, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-5.

Schorr:1995:REI

[398] Mark S. Schorr, Jaysingh Sah, Dean F. Schreiner, Michael R. Meador, and Loren G. Hill. Regional economic impact of the Lake Texoma (Oklahoma–Texas) striped bass fishery. Fisheries, 20(5):14–18, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281995%29020%3C0014%3AREIOTL%3E2.0.CO%3B2.

Anonymous:1995:FIAf

[399] Anonymous. Full issue article volume 20, issue 6. Fisheries, 20(6):1–80, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-6.

Brouha:1995:CAF

[400] Paul Brouha, John Esson, Walter Harris, and Susan Monseur. A comparison of agencies’ fisheries biologist salaries: 1987, 1990, 1994. Fisheries, 20(6):14–18, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281995%29020%3C0014%3AACOAFB%3E2.0.CO%3B2.

Whittier:1995:RSN

[401] Thomas R. Whittier, Alan T. Herlihy, and Suzanne M. Pierson. Regional susceptibility of northeast lakes to zebra mussel invasion. Fisheries, 20 (6):20–27, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281995%29020%3C0020%3ARSONLT%3E2.0.CO%3B2.

Anonymous:1995:FIAg

[402] Anonymous. Full issue article volume 20, issue 7. Fisheries, 20(7):1–72, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-7.

Murphy:1995:LLA

[403] William F. Murphy, Gerald H. Cross, and Louis A. Helfrich. Lifelong learning for agency fisheries professionals: What are the continuing education needs? Fisheries, 20(7):10–16, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281995%29020%3C0010%3ALLFAFP%
3E2.0.CO%3B2.

Radomski:1995:HML

[404] Paul J. Radomski and Timothy J. Goeman. The homogenizing of Minnesota lake fish assemblages. Fisheries, 20(7):20–23, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281995%29020%
3C0020%3ATHOMLF%3E2.0.CO%3B2.

Anonymous:1995:FIAh

[405] Anonymous. Full issue article volume 20, issue 8. Fisheries, 20(8):1–64, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-8.

Nielsen:1995:PUF

[406] Larry A. Nielsen. The practical uses of fisheries history. Fisheries, 20 (8):16–18, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281995%29020%3C0016%3ATPUOFH%3E2.0.CO%3B2.

Anonymous:1995:FIAi

[407] Anonymous. Full issue article volume 20, issue 9. Fisheries, 20(9):1–64, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-9.

Anonymous:1995:FIAj

[408] Anonymous. Full issue article volume 20, issue 10. Fisheries, 20(10):1–64, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-10.

Parrish:1995:PSR

[409] Donna L. Parrish, Martha E. Mather, and Roy A. Stein. Problemsolving research for management: a perspective. Fisheries, 20(10):6– 12, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281995%29020%3C0006%3APRFMAP%3E2.0.CO%3B2.

Mather:1995:MIT

[410] Martha E. Mather, Donna L. Parrish, Roy A. Stein, and Robert M. Muth. Management issues and their relative priority within state fisheries agencies. Fisheries, 20(10):14–21, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281995%29020%3C0014%3AMIATRP%3E2.
0.CO%3B2.

Johnson:1995:SHR

[411] Brett M. Johnson and Patrick J. Martinez. Selecting harvest regulations for recreational fisheries: Opportunities for research/management cooperation. Fisheries, 20(10):22–29, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281995%29020%3C0022%3ASHRFRF%3E2.
0.CO%3B2.

Anonymous:1995:FIAk

[412] Anonymous. Full issue article volume 20, issue 11. Fisheries, 20(11):1–64, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-11.

Bryant:1995:PMW

[413] Mason D. Bryant. Pulsed monitoring for watershed and stream restoration. Fisheries, 20(11):6–13, 1995. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281995%29020%3C0006%3APMFWAS%3E2.0.CO%3B2.

Meronek:1995:SBR

[414] Thomas G. Meronek, Fred A. Copes, and Daniel W. Coble. A summary of bait regulations in the North Central United States. Fisheries, 20(11): 16–23, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281995%29020%3C0016%3AASOBRI%3E2.0.CO%3B2.

Anonymous:1995:FIAl

[415] Anonymous. Full issue article volume 20, issue 12. Fisheries, 20(12):1–72, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-20-12.

Bevelhimer:1995:RAC

[416] Mark S. Bevelhimer. Recent advances in contaminant assessment offer proactive alternatives for managing contaminated fisheries. Fisheries, 20 (12):6–10, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281995%29020%3C0006%3ARAICAO%3E2.0.CO%3B2.

Currens:1995:FAG

[417] Kenneth P. Currens and Craig A. Busack. A framework for assessing genetic vulnerability. Fisheries, 20(12):24–31, 1995. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281995%29020%
3C0024%3AAFFAGV%3E2.0.CO%3B2.

Anonymous:1996:FIAa

[418] Anonymous. Full issue article volume 21, issue 1. Fisheries, 21(1):1–64, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-1.

Fenton:1996:RFD

[419] Robert Fenton, Jack A. Mathias, and G. Eric E. Moodie. Recent and future demand for walleye in North America. Fisheries, 21(1): 6–12, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281996%29021%3C0006%3ARAFDFW%3E2.0.CO%3B2.

Anonymous:1996:FIAb

[420] Anonymous. Full issue article volume 21, issue 2. Fisheries, 21(2):1–56, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-2.

Hartig:1996:TIR

[421] J. H. Hartig, D. P. Dodge, D. Jester, J. Atkinson, R. Thoma, and K. Cullis. Toward integrating remedial action planning and fishery management planning in Great Lakes areas of concern. Fisheries, 21(2): 6–13, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281996%29021%3C0006%3ATIRAPA%3E2.0.CO%3B2.

Anonymous:1996:FIAc

[422] Anonymous. Full issue article volume 21, issue 3. Fisheries, 21(3):1–64, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-3.

Huntington:1996:SHN

[423] Charles Huntington, Willa Nehlsen, and Jon Bowers. A survey of healthy native stocks of anadromous salmonids in the Pacific Northwest and California. Fisheries, 21(3):6–14, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281996%29021%3C0006%3AASOHNS%3E2.
0.CO%3B2.

Kaeding:1996:LTD

[424] Lynn R. Kaeding, Glenn D. Boltz, and Daniel G. Carty. Lake trout discovered in Yellowstone Lake threaten native cutthroat trout. Fisheries, 21(3):16–20, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281996%29021%3C0016%3ALTDIYL%3E2.0.CO%3B2.

Anonymous:1996:FIAd

[425] Anonymous. Full issue article volume 21, issue 4. Fisheries, 21(4):1–64, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-4.

Ballweber:1996:OEF

[426] Jeffery A. Ballweber and Donald C. Jackson. Opportunities to emphasize fisheries concerns in Federal Agency decision-making: an introduction. Fisheries, 21(4):14–19, 1996. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281996%29021%3C0014%3AOTEFCI%3E2.0.CO%3B2.

Taylor:1996:CSC

[427] Christopher A. Taylor, Melvin L. Warren, Jr., J. F. Fitzpatrick, Jr., Horton H. Hobbs III, Raymond F. Jezerinac, William L. Pflieger, and Henry W. Robison. Conservation status of crayfishes of the United States and Canada. Fisheries, 21(4):25–38, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281996%29021%3C0025%3ACSOCOT%3E2.
0.CO%3B2.

Anonymous:1996:FIAe

[428] Anonymous. Full issue article volume 21, issue 5. Fisheries, 21(5):1–56, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-5.

Thayer:1996:HRP

[429] Gordon W. Thayer, James P. Thomas, and K. V. Koski. The habitat research plan of the national marine fisheries service. Fisheries, 21(5): 6–10, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281996%29021%3C0006%3ATHRPOT%3E2.0.CO%3B2.

Fraley:1996:CCW

[430] John Fraley. Cooperation and controversy in wilderness fisheries management. Fisheries, 21(5):16–21, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281996%29021%3C0016%3ACACIWF%3E2.
0.CO%3B2.

Anonymous:1996:FIAf

[431] Anonymous. Full issue article volume 21, issue 6. Fisheries, 21(6):1–88, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-6.

Malvestuto:1996:OYR

[432] Stephen P. Malvestuto and Michael D. Hudgins. Optimum yield for recreational fisheries management. Fisheries, 21(6):6–17, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281996%29021%
3C0006%3AOYFRFM%3E2.0.CO%3B2.

Luecke:1996:SGP

[433] Chris Luecke, Wayne A. Wurtsbaugh, Phaedra Budy, Howard P. Gross, and Geoff Steinhart. Simulated growth and production of endangered Snake River sockeye salmon: Assessing management strategies for the nursery lakes. Fisheries, 21(6):18–25, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281996%29021%3C0018%3ASGAPOE%
3E2.0.CO%3B2.

Anonymous:1996:FIAg

[434] Anonymous. Full issue article volume 21, issue 7. Fisheries, 21(7):1–64, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-7.

Moring:1996:FDL

[435] John R. Moring. Fish discoveries by the Lewis and Clark and Red River expeditions. Fisheries, 21(7):6–12, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281996%29021%3C0006%3AFDBTLA%3E2.
0.CO%3B2.

Ludwig:1996:ITA

[436] Herbert R. Ludwig, Jr. and Jay A. Leitch. Interbasin transfer of aquatic biota via anglers’ bait buckets. Fisheries, 21(7):14–18, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281996%29021%3C0014%3AITOABV%3E2.0.CO%3B2.

Anonymous:1996:FIAh

[437] Anonymous. Full issue article volume 21, issue 8. Fisheries, 21(8):1–64, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-8.

McMahon:1996:WNP

[438] Thomas E. McMahon and David H. Bennett. Walleye and northern pike. Fisheries, 21(8):6–13, 1996. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281996%29021%3C0006%3AWANP%3E2.0.CO%3B2.

Geoghegan:1996:MQC

[439] Paul Geoghegan. The management of quality control and quality assurance systems in fisheries science. Fisheries, 21(8):14–18, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281996%29021%3C0014%3ATMOQCA%3E2.0.CO%3B2.

Anonymous:1996:FIAi

[440] Anonymous. Full issue article volume 21, issue 9. Fisheries, 21(9):1–64, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-9.

Stiassny:1996:OFB

[441] Melanie L. J. Stiassny. An overview of freshwater biodiversity: With some lessons from African fishes. Fisheries, 21(9):7–13, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281996%29021%
3C0007%3AAOOFB%3E2.0.CO%3B2.

Meador:1996:WTP

[442] Michael R. Meador. Water transfer projects and the role of fisheries biologists. Fisheries, 21(9):18–23, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281996%29021%3C0018%3AWTPATR%3E2.
0.CO%3B2.

Anonymous:1996:FIAj

[443] Anonymous. Full issue article volume 21, issue 10. Fisheries, 21(10):1–72, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-10.

Baker:1996:SPS

[444] Timothy T. Baker, Alex C. Wertheimer, Robert D. Burkett, Ronald Dunlap, Douglas M. Eggers, Ellen I. Fritts, Anthony J. Gharrett, Rolland A. Holmes, and Richard L. Wilmot. Status of Pacific salmon and steelhead escapements in southeastern Alaska. Fisheries, 21(10): 6–18, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281996%29021%3C0006%3ASOPSAS%3E2.0.CO%3B2.

Slaney:1996:SAS

[445] T. L. Slaney, K. D. Hyatt, T. G. Northcote, and R. J. Fielden. Status of anadromous salmon and trout in British Columbia and Yukon. Fisheries, 21(10):20–35, 1996. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281996%29021%3C0020%3ASOASAT%3E2.0.CO%3B2.

Anonymous:1996:FIAk

[446] Anonymous. Full issue article volume 21, issue 11. Fisheries, 21(11):1–64, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-11.

Loomis:1996:EVI

[447] John B. Loomis and Douglas S. White. Economic values of increasingly rare and endangered fish. Fisheries, 21(11):6–10, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281996%29021%
3C0006%3AEVOIRA%3E2.0.CO%3B2.

Wilde:1996:SHD

[448] Gene R. Wilde, Robert B. Ditton, Shepherd R. Grimes, and Robin K. Riechers. Status of human dimensions surveys sponsored by state and provincial fisheries management agencies in North America. Fisheries, 21 (11):12–17, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281996%29021%3C0012%3ASOHDSS%3E2.0.CO%3B2.

Daigle:1996:PJF

[449] Cheryl Perusse Daigle, David K. Loomis, and Robert B. Ditton. Procedural justice in fishery resource allocations. Fisheries, 21(11):18– 23, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281996%29021%3C0018%3APJIFRA%3E2.0.CO%3B2.

Anonymous:1996:FIAl

[450] Anonymous. Full issue article volume 21, issue 12. Fisheries, 21(12):1–64, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-21-12.

Schramm:1996:EMI

[451] Harold L. Schramm, Jr. and Wayne A. Hubert. Ecosystem management: Implications for fisheries management. Fisheries, 21(12):6– 11, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281996%29021%3C0006%3AEMIFFM%3E2.0.CO%3B2.

Radomski:1996:DMM

[452] Paul J. Radomski and Timothy J. Goeman. Decision making and modeling in freshwater sport-fisheries management. Fisheries, 21(12):14– 21, 1996. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281996%29021%3C0014%3ADMAMIF%3E2.0.CO%3B2.

Anonymous:1997:FIAa

[453] Anonymous. Full issue article volume 22, issue 1. Fisheries, 22(1):1–64, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-1.

Isaak:1997:INT

[454] Daniel J. Isaak and Wayne A. Hubert. Integrating new technologies into fisheries science: The application of geographic information systems. Fisheries, 22(1):6–10, 1997. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281997%29022%3C0006%3AINTIFS%3E2.0.CO%3B2.

Lichtkoppler:1997:OLE

[455] Frank R. Lichtkoppler. Ohio’s Lake Erie charter fishing industry: 1985– 1994. Fisheries, 22(1):14–21, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281997%29022%3C0014%3AOLECFI%3E2.
0.CO%3B2.

Anonymous:1997:FIAb

[456] Anonymous. Full issue article volume 22, issue 2. Fisheries, 22(2):1–64, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-2.

Williams:1997:SFE

[457] Cindy Deacon Williams. Sustainable fisheries: Economics, ecology, and ethics. Fisheries, 22(2):6–11, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281997%29022%3C0006%3ASFEEAE%3E2.
0.CO%3B2.

Anonymous:1997:FIAc

[458] Anonymous. Full issue article volume 22, issue 3. Fisheries, 22(3):1–64, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-3.

Anonymous:1997:FIAd

[459] Anonymous. Full issue article volume 22, issue 4. Fisheries, 22(4):1–64, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-4.

Grossman:1997:DAR

[460] Gary D. Grossman, Geoff P. Jones, and William J. Seaman, Jr. Do artificial reefs increase regional fish production? A review of existing data. Fisheries, 22(4):17–23, 1997. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281997%29022%3C0017%3ADARIRF%3E2.0.CO%3B2.

Carr:1997:ARI

[461] Mark H. Carr and Mark A. Hixon. Artificial reefs: The importance of comparisons with natural reefs. Fisheries, 22(4):28–33, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281997%29022%3C0028%3AARTIOC%3E2.0.CO%3B2.

Anonymous:1997:FIAe

[462] Anonymous. Full issue article volume 22, issue 5. Fisheries, 22(5):1–64, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-5.

Roper:1997:SRF

[463] Brett B. Roper, Jeffrey J. Dose, and Jack E. Williams. Stream restoration: Is fisheries biology enough? Fisheries, 22(5):6–11, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281997%29022%3C0006%3ASRIFBE%3E2.0.CO%3B2.

Kauffman:1997:EPR

[464] J. Boone Kauffman, Robert L. Beschta, Nick Otting, and Danna Lytjen. An ecological perspective of riparian and stream restoration in the Western United States. Fisheries, 22(5):12–24, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281997%29022%
3C0012%3AAEPORA%3E2.0.CO%3B2.

Anonymous:1997:FIAf

[465] Anonymous. Full issue article volume 22, issue 6. Fisheries, 22(6):1–88, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-6.

Wang:1997:IWL

[466] Lizhu Wang, John Lyons, Paul Kanehl, and Ronald Gatti. Influences of watershed land use on habitat quality and biotic integrity in Wisconsin streams. Fisheries, 22(6):6–12, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281997%29022%3C0006%3AIOWLUO%3E2.
0.CO%3B2.

Wilde:1997:LBF

[467] Gene R. Wilde. Largemouth bass fishery responses to length limits. Fisheries, 22(6):14–23, 1997. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281997%29022%3C0014%3ALBFRTL%3E2.0.CO%3B2.

Anonymous:1997:FIAg

[468] Anonymous. Full issue article volume 22, issue 7. Fisheries, 22(7):1–56, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-7.

Anonymous:1997:FIAh

[469] Anonymous. Full issue article volume 22, issue 8. Fisheries, 22(8):1–64, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-8.

Bergersen:1997:DSM

[470] Eric P. Bergersen and Dennis E. Anderson. The distribution and spread of Myxobolus cerebralis in the United States. Fisheries, 22(8): 6–7, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281997%29022%3C0006%3ATDASOM%3E2.0.CO%3B2.

Anonymous:1997:FIAi

[471] Anonymous. Full issue article volume 22, issue 9. Fisheries, 22(9):1–56, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-9.

Anonymous:1997:FIAj

[472] Anonymous. Full issue article volume 22, issue 10. Fisheries, 22(10):1–48, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-10.

Rand:1997:NRE

[473] Peter S. Rand, James P. Scandol, and Emily E. Walter. NerkaSim: a research and educational tool to of Pacific salmon in a dynamic environment. Fisheries, 22(10):6–13, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281997%29022%3C0006%3ANARAET%3E2.
0.CO%3B2.

Anonymous:1997:FIAk

[474] Anonymous. Full issue article volume 22, issue 11. Fisheries, 22(11):1–56, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-11.

Rieman:1997:WNF

[475] Bruce Rieman and Jim Clayton. Wildfire and native fish: Issues of forest health and conservation of sensitive species. Fisheries, 22(11): 6–15, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281997%29022%3C0006%3AWANFIO%3E2.0.CO%3B2.

Larkin:1997:ITM

[476] Gillian A. Larkin and Pat A. Slaney. Implications of trends in marinederived nutrient influx to south coastal British Columbia salmonid production. Fisheries, 22(11):16–24, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281997%29022%3C0016%3AIOTIMN%3E2.
0.CO%3B2.

Anonymous:1997:FIAl

[477] Anonymous. Full issue article volume 22, issue 12. Fisheries, 22(12):1–64, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-22-12.

Smith:1997:CVC

[478] Courtland L. Smith, Jennifer D. Gilden, Joseph S. Cone, and Brent S. Steel. Contrasting views of coastal residents and coastal Coho restoration planners. Fisheries, 22(12):8–15, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281997%29022%3C0008%3ACVOCRA%3E2.
0.CO%3B2.

Jeffrey:1997:GPS

[479] Jay D. Jeffrey and David R. Edds. A global positioning system for aquatic surveys. Fisheries, 22(12):16–20, 1997. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281997%29022%3C0016%3AAGPSFA%3E2.
0.CO%3B2.

Anonymous:1998:FIAa

[480] Anonymous. Full issue article volume 23, issue 1. Fisheries, 23(1):1–48, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-1.

Weiner:1998:KRR

[481] Arthur H. Weiner. Kenai River restoration and management. Fisheries, 23(1):6–10, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281998%29023%3C0006%3AKRRAM%3E2.0.CO%3B2.

Bugert:1998:MSC

[482] Robert M. Bugert. Mechanics of supplementation in the Columbia River. Fisheries, 23(1):11–20, 1998. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281998%29023%3C0014%3AMOSITC%3E2.0.CO%3B2.

Anonymous:1998:FIAb

[483] Anonymous. Full issue article volume 23, issue 2. Fisheries, 23(2):1–50, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-2.

Olson:1998:MMI

[484] Mark H. Olson, Stephen R. Carpenter, Paul Cunningham, Sarig Gafny, Brian R. Herwig, Nathan P. Nibbelink, Tom Pellett, Christine Storlie, Anett S. Trebitz, and Karen A. Wilson. Managing macrophytes to improve fish growth: a multi-lake experiment. Fisheries, 23(2):6– 12, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281998%29023%3C0006%3AMMTIFG%3E2.0.CO%3B2.

Anonymous:1998:FIAc

[485] Anonymous. Full issue article volume 23, issue 3. Fisheries, 23(3):1–48, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-3.

Bortone:1998:RAP

[486] Stephen A. Bortone. Resolving the attraction–production dilemma in artificial reef research: Some yeas and nays. Fisheries, 23(3):6– 10, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281998%29023%3C0006%3ARTADIA%3E2.0.CO%3B2.

Anonymous:1998:FIAd

[487] Anonymous. Full issue article volume 23, issue 4. Fisheries, 23(4):1–56, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-4.

Lyons:1998:DFF

[488] John Lyons, Georgina González-Hernández, Eduardo Soto-Galera, and Manuel Guzmán-Arroyo. Decline of freshwater fishes and fisheries in selected drainages of West-Central Mexico. Fisheries, 23(4):10– 18, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281998%29023%3C0010%3ADOFFAF%3E2.0.CO%3B2.

Anonymous:1998:FIAe

[489] Anonymous. Full issue article volume 23, issue 5. Fisheries, 23(5):1–52, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-5.

Rivers:1998:VA

[490] P. J. Rivers and W. R. Ardren. The value of archives. Fisheries, 23 (5):6–9, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281998%29023%3C0006%3ATVOA%3E2.0.CO%3B2.

Fisher:1998:RTG

[491] William L. Fisher and Conrad S. Toepfer. Recent trends in geographic information systems education and fisheries research applications at U.S. universities. Fisheries, 23(5):10–13, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281998%29023%3C0010%3ARTIGIS%3E2.
0.CO%3B2.

Beard:1998:MSA

[492] T. Douglas Beard, Jr., Douglas Austen, Stephen J. Brady, Mike E. Costello, Henry G. Drewes, Constance H. Young-Dubovsky, Curtis H. Flather, Thomas W. Gengerke, Chris Larson, Andrew J. Loftus, and Michael J. Mac. The Multi-state Aquatic Resources Information System: an Internet system to access fisheries information in the upper midwestern United States. Fisheries, 23(5):14– 18, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281998%29023%3C0014%3ATMARIS%3E2.0.CO%3B2.

Anonymous:1998:FIAf

[493] Anonymous. Full issue article volume 23, issue 6. Fisheries, 23(6):1–88, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-6.

Crowder:1998:FBI

[494] Larry B. Crowder and Steven A. Murawski. Fisheries bycatch: Implications for management. Fisheries, 23(6):8–17, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281998%29023%
3C0008%3AFBIFM%3E2.0.CO%3B2.

Rupert:1998:FPR

[495] Jeffery D. Rupert and Shari L. Dann. Fishing in the parks: a research-based outreach program. Fisheries, 23(6):19–27, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281998%29023%
3C0019%3AFITPAR%3E2.0.CO%3B2.

Anonymous:1998:FIAg

[496] Anonymous. Full issue article volume 23, issue 7. Fisheries, 23(7):1–60, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-7.

Moyle:1998:FHD

[497] Peter B. Moyle, Michael P. Marchetti, Jean Baldrige, and Thomas L. Taylor. Fish health and diversity: Justifying flows for a California stream. Fisheries, 23(7):6–15, 1998. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281998%29023%3C0006%3AFHADJF%3E2.0.CO%3B2.

Rubec:1998:SMB

[498] Peter J. Rubec, Michael S. Coyne, Robert H. McMichael, Jr., and Mark E. Monaco. Spatial methods being developed in Florida to determine essential fish habitat. Fisheries, 23(7):21–25, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281998%29023%
3C0021%3ASMBDIF%3E2.0.CO%3B2.

Anonymous:1998:FIAh

[499] Anonymous. Full issue article volume 23, issue 8. Fisheries, 23(8):1–64, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-8.

Harvey:1998:ESD

[500] Bret C. Harvey and Thomas E. Lisle. Effects of suction dredging on streams: a review and an evaluation strategy. Fisheries, 23(8): 8–17, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281998%29023%3C0008%3AEOSDOS%3E2.0.CO%3B2.

Anonymous:1998:FIAi

[501] Anonymous. Full issue article volume 23, issue 9. Fisheries, 23(9):1–60, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-9.

Grimes:1998:MSE

[502] Churchill B. Grimes. Marine stock enhancement: Sound management or techno-arrogance? Fisheries, 23(9):18–23, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281998%29023%
3C0018%3AMSE%3E2.0.CO%3B2.

Anonymous:1998:FIAj

[503] Anonymous. Full issue article volume 23, issue 10. Fisheries, 23(10):1–64, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-10.

Wilde:1998:TAM

[504] Gene R. Wilde. Tournament-associated mortality in black bass. Fisheries, 23(10):12–22, 1998. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281998%29023%3C0012%3ATMIBB%3E2.0.CO%3B2.

Zint:1998:NAF

[505] Michaela Zint and Alan Crook. A needs assessment of fisheries education materials for youth. Fisheries, 23(10):24–34, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281998%29023%
3C0024%3AANAOFE%3E2.0.CO%3B2.

Anonymous:1998:FIAk

[506] Anonymous. Full issue article volume 23, issue 11. Fisheries, 23(11):1–56, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-11.

Meador:1998:ISG

[507] Michael R. Meador and April O. Layher. Instream sand and gravel mining: Environmental issues and regulatory process in the United States. Fisheries, 23(11):6–13, 1998. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281998%29023%3C0006%3AISAGM%3E2.0.CO%3B2.

Hill:1998:ERC

[508] Mark T. Hill and William S. Platts. Ecosystem restoration: a case study in the Owens River Gorge, California. Fisheries, 23(11):18– 27, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281998%29023%3C0018%3AER%3E2.0.CO%3B2.

Anonymous:1998:FIAl

[509] Anonymous. Full issue article volume 23, issue 12. Fisheries, 23(12):1–56, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-23-12.

Nielsen:1998:ESS

[510] Jennifer L. Nielsen. Scientific Sampling Effects: Electrofishing California’s endangered fish populations. Fisheries, 23(12):6–12, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281998%29023%
3C0006%3ASSEECE%3E2.0.CO%3B2.

Fisher:1998:WSR

[511] William L. Fisher, Albert F. Surmont, and Craig D. Martin. Warmwater stream and river fisheries in the Southeastern United States: Are we managing them in proportion to their values? Fisheries, 23(12): 16–24, 1998. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281998%29023%3C0016%3AWSARFI%3E2.0.CO%3B2.

Anonymous:1999:FIAa

[512] Anonymous. Full issue article volume 24, issue 1. Fisheries, 24(1):1–64, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-1.

Hare:1999:IPR

[513] Steven R. Hare, Nathan J. Mantua, and Robert C. Francis. Inverse production regimes: Alaska and West Coast Pacific salmon. Fisheries, 24 (1):6–14, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0006%3AIPR%3E2.0.CO%3B2.

Nico:1999:STP

[514] Leo G. Nico and Pam L. Fuller. Spatial and temporal patterns of nonindigenous fish introductions in the United States. Fisheries, 24(1): 16–27, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0016%3ASATPON%3E2.0.CO%3B2.

Anonymous:1999:FIAb

[515] Anonymous. Full issue article volume 24, issue 2. Fisheries, 24(2):1–48, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-2.

Barse:1999:ENP

[516] Ann M. Barse and David H. Secor. An exotic nematode parasite of the American eel. Fisheries, 24(2):6–10, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281999%29024%3C0006%3AAENPOT%3E2.
0.CO%3B2.

Waples:1999:DSM

[517] Robin S. Waples. Dispelling some myths about hatcheries. Fisheries, 24 (2):12–21, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0012%3ADSMAH%3E2.0.CO%3B2.

Anonymous:1999:FIAc

[518] Anonymous. Full issue article volume 24, issue 3. Fisheries, 24(3):1–48, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-3.

Brown:1999:SDP

[519] Michael L. Brown and David W. Willis. Scope and diversity of privatized services in the North Central Division. Fisheries, 24(3):6– 9, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0006%3ASADOPS%3E2.0.CO%3B2.

Williams:1999:ERR

[520] Richard N. Williams, Peter A. Bisson, Daniel L. Bottom, Lyle D. Calvin, Charles C. Coutant, Michael W. Erho, Jr., Christopher A. Frissell, James A. Lichatowich, William J. Liss, Willis E. McConnaha, Phillip R. Mundy, Jack A. Stanford, and Richard R. Whitney. Return to the River: Scientific issues in the restoration of salmonid fishes in the Columbia River. Fisheries, 24(3):10–19, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%281999%29024%3C0010%3ARTTRSI%3E2.
0.CO%3B2.

Anonymous:1999:FIAd

[521] Anonymous. Full issue article volume 24, issue 4. Fisheries, 24(4):1–52, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-4.

Beechie:1999:ARS

[522] Tim Beechie and Susan Bolton. An approach to restoring salmonid habitat-forming processes in Pacific Northwest watersheds. Fisheries, 24 (4):6–15, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0006%3AAATRSH%3E2.0.CO%3B2.

Bain:1999:TMA

[523] Mark B. Bain, Thomas C. Hughes, and Kristin K. Arend. Trends in methods for assessing freshwater habitats. Fisheries, 24(4):16– 21, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0016%3ATIMFAF%3E2.0.CO%3B2.

Anonymous:1999:FIAe

[524] Anonymous. Full issue article volume 24, issue 5. Fisheries, 24(5):1–56, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-5.

LaPatra:1999:ADM

[525] Scott E. LaPatra, Susan C. Ireland, Joseph M. Groff, Kathy M. Clemens, and John T. Siple. Adaptive disease management strategies for the endangered population of Kootenai River white sturgeon. Fisheries, 24 (5):6–13, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0006%3AADMSFT%3E2.0.CO%3B2.

Anonymous:1999:FIAf

[526] Anonymous. Full issue article volume 24, issue 6. Fisheries, 24(6):1–72, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-6.

Hall-Arber:1999:HOD

[527] Madeleine Hall-Arber and Judith Pederson. Habitat observed from the decks of fishing vessels. Fisheries, 24(6):6–13, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281999%29024%
3C0006%3AHOFTDO%3E2.0.CO%3B2.

Langton:1999:MFH

[528] Richard W. Langton and Peter J. Auster. Marine fishery and habitat interactions: To what extent are fisheries and habitat interdependent? Fisheries, 24(6):14–21, 1999. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281999%29024%3C0014%3AMFAHI%3E2.0.CO%3B2.

Radomski:1999:COP

[529] Paul J. Radomski. Commercial overfishing and property rights. Fisheries, 24(6):22–29, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/abs/10.
1577/1548-8446%281999%29024%3C0022%3ACOAPR%3E2.0.CO%3B2.

Anonymous:1999:FIAg

[530] Anonymous. Full issue article volume 24, issue 7. Fisheries, 24(7):1–48, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-7.

Ross:1999:SMF

[531] Michael R. Ross and David K. Loomis. State management of freshwater fisheries resources: Its organizational structure, funding, and programmatic emphases. Fisheries, 24(7):8–14, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%281999%29024%3C0008%3ASMOFFR%
3E2.0.CO%3B2.

Moore:1999:DSF

[532] Jon A. Moore. Deep-sea finfish fisheries: Lessons from history. Fisheries, 24(7):16–21, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281999%29024%3C0016%3ADFFLFH%3E2.0.CO%3B2.

Schmetterling:1999:MAI

[533] David A. Schmetterling and Mack H. Long. Montana anglers’ inability to identify bull trout and other salmonids. Fisheries, 24(7):24– 27, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0024%3AMAITIB%3E2.0.CO%3B2.

Anonymous:1999:FIAh

[534] Anonymous. Full issue article volume 24, issue 8. Fisheries, 24(8):1–48, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-8.

Michaletz:1999:SCM

[535] Paul H. Michaletz and Joe G. Dillard. A survey of catfish management in the United States and Canada. Fisheries, 24(8):6–11, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0006%3AASOCMI%3E2.0.CO%3B2.

Rinne:1999:FGR

[536] John N. Rinne. Fish and grazing relationships: The facts and some pleas. Fisheries, 24(8):12–21, 1999. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281999%29024%3C0012%3AFAGR%3E2.0.CO%3B2.

Anonymous:1999:FIAi

[537] Anonymous. Full issue article volume 24, issue 9. Fisheries, 24(9):1–48, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-9.

Pearsons:1999:PAA

[538] Todd N. Pearsons and Charles W. Hopley. A practical approach for assessing ecological risks associated with fish stocking programs. Fisheries, 24(9):16–23, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%281999%29024%3C0016%3AAPAFAE%3E2.0.CO%3B2.

Anonymous:1999:FIAj

[539] Anonymous. Full issue article volume 24, issue 10. Fisheries, 24(10):1–48, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-10.

Cederholm:1999:PSC

[540] C. Jeff Cederholm, Matt D. Kunze, Takeshi Murota, and Atuhiro Sibatani. Pacific salmon carcasses: Essential contributions of nutrients and energy for aquatic and terrestrial ecosystems. Fisheries, 24(10):

6–15, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0006%3APSC%3E2.0.CO%3B2.

Nemeth:1999:SRS

[541] Douglas J. Nemeth and Russell B. Kiefer. Snake River spring and summer Chinook salmon — the choice for recovery. Fisheries, 24(10): 16–23, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0016%3ASRSASC%3E2.0.CO%3B2.

Anonymous:1999:FIAk

[542] Anonymous. Full issue article volume 24, issue 11. Fisheries, 24(11):1–52, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-11.

Berkson:1999:HCB

[543] Jim Berkson and Carl N. Shuster, Jr. The horseshoe crab: The battle for a true multiple-use resource. Fisheries, 24(11):6–10, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%281999%29024%
3C0006%3ATHCTBF%3E2.0.CO%3B2.

Murray:1999:NTR

[544] Steven N. Murray, Richard F. Ambrose, James A. Bohnsack, Louis W. Botsford, Mark H. Carr, Gary E. Davis, Paul K. Dayton, Dan Gotshall, Don R. Gunderson, Mark A. Hixon, Jane Lubchenco, Marc Mangel, Alec MacCall, Deborah A. McArdle, John C. Ogden, Joan Roughgarden, Richard M. Starr, Mia J. Tegner, and Mary M. Yoklavich. No-take reserve networks: Sustaining fishery populations and marine ecosystems. Fisheries, 24(11):11–25, 1999. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%281999%29024%3C0011%3ANRN%3E2.0.CO%3B2.

Anonymous:1999:FIAl

[545] Anonymous. Full issue article volume 24, issue 12. Fisheries, 24(12):1–48, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-24-12.

Musick:1999:CDE

[546] J. A. Musick. Criteria to define extinction risk in marine fishes: The American Fisheries Society initiative. Fisheries, 24(12):6–14, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0006%3ACTDERI%3E2.0.CO%3B2.

Sherman:1999:LME

[547] Kenneth Sherman and Alfred M. Duda. Large marine ecosystems: an emerging paradigm for fishery sustainability. Fisheries, 24(12):15– 26, 1999. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%281999%29024%3C0015%3ALME%3E2.0.CO%3B2.

Anonymous:2000:FIAa

[548] Anonymous. Full issue article volume 25, issue 1. Fisheries, 25(1):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-1.

Smith:2000:HHN

[549] Courtland L. Smith and Jennifer Gilden. Human and habitat needs in disaster relief for Pacific Northwest salmon fisheries. Fisheries, 25(1): 6–14, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0006%3AHAHNID%3E2.0.CO%3B2.

Gresh:2000:EHC

[550] Ted Gresh, Jim Lichatowich, and Peter Schoonmaker. An estimation of historic and current levels of salmon production in the Northeast Pacific Ecosystem: Evidence of a nutrient deficit in the freshwater systems of the Pacific Northwest. Fisheries, 25(1):15– 21, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0015%3AAEOHAC%3E2.0.CO%3B2.

Fedler:2000:DNO

[551] Anthony J. Fedler and Robert B. Ditton. Developing a national outreach strategy for recreational fishing and boating. Fisheries, 25(1): 22–28, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0022%3ADANOSF%3E2.0.CO%3B2.

Anonymous:2000:FIAb

[552] Anonymous. Full issue article volume 25, issue 2. Fisheries, 25(2):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-2.

Yoshiyama:2000:CSC

[553] Ronald M. Yoshiyama, Peter B. Moyle, Eric R. Gerstung, and Frank W. Fisher. Chinook salmon in the California Central Valley: an assessment. Fisheries, 25(2):6–20, 2000. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282000%29025%3C0006%3ACSITCC%3E2.0.CO%3B2.

Connelly:2000:DAF

[554] Nancy A. Connelly, Tommy L. Brown, and Barbara A. Knuth. Do anglers and fishery professionals think alike? Fisheries, 25(2):21– 25, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0021%3ADAAFPT%3E2.0.CO%3B2.

Anonymous:2000:FIAc

[555] Anonymous. Full issue article volume 25, issue 3. Fisheries, 25(3):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-3.

Musick:2000:PMF

[556] J. A. Musick, S. A. Berkeley, G. M. Cailliet, M. Camhi, G. Huntsman, M. Nammack, and M. L. Warren, Jr. Protection of marine fish stocks at risk of extinction. Fisheries, 25(3):6–8, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%282000%29025%3C0006%3APOMFSA%
3E2.0.CO%3B2.

Musick:2000:MST

[557] J. A. Musick, G. Burgess, G. Cailliet, M. Camhi, and S. Fordham. Management of sharks and their relatives (Elasmobranchii). Fisheries, 25 (3):9–13, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0009%3AMOSATR%3E2.0.CO%3B2.

Coleman:2000:LLR

[558] F. C. Coleman, C. C. Koenig, G. R. Huntsman, J. A. Musick, A. M. Eklund, J. C. McGovern, G. R. Sedberry, R. W. Chapman, and C. B. Grimes. Long-lived reef fishes: The grouper–snapper complex. Fisheries, 25(3):14–21, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%282000%29025%3C0014%3ALRF%3E2.0.CO%3B2.

Parker:2000:MPR

[559] S. J. Parker, S. A. Berkeley, J. T. Golden, D. R. Gunderson, J. Heifetz, M. A. Hixon, R. Larson, B. M. Leaman, M. S. Love, J. A. Musick, V. M. O’Connell, S. Ralston, H. J. Weeks, and M. M. Yoklavich. Management of Pacific rockfish. Fisheries, 25(3):22–30, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0022%3AMOPR%3E2.0.CO%3B2.

Anonymous:2000:FIAd

[560] Anonymous. Full issue article volume 25, issue 4. Fisheries, 25(4):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-4.

Loftus:2000:OAB

[561] Andrew J. Loftus, Jeff Waldon, Virginia Fay, Kay Davy, and Jon Lucy. Overview of angler-based tagging programs and management issues. Fisheries, 25(4):8–13, 2000. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282000%29025%3C0008%3AOOATPA%3E2.0.CO%3B2.

Lucy:2000:BAA

[562] Jon Lucy and Kay Davy. Benefits of angler-assisted tag and release programs. Fisheries, 25(4):18–23, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%282000%29025%3C0018%3ABOATAR%3E2.
0.CO%3B2.

Anonymous:2000:FIAe

[563] Anonymous. Full issue article volume 25, issue 5. Fisheries, 25(5):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-5.

Stockner:2000:COC

[564] J. G. Stockner, E. Rydin, and P. Hyenstrand. Cultural oligotrophication: Causes and consequences for fisheries resources. Fisheries, 25(5): 7–14, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0007%3ACO%3E2.0.CO%3B2.

McClay:2000:RUN

[565] William McClay. Rotenone use in North America (1988–1997). Fisheries, 25(5):15–21, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%282000%29025%3C0015%3ARUINA%3E2.0.CO%3B2.

Anonymous:2000:FIAf

[566] Anonymous. Full issue article volume 25, issue 6. Fisheries, 25(6):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-6.

Brodeur:2000:CRP

[567] Richard D. Brodeur, William T. Peterson, George W. Boehlert, Ed Casillas, Michael H. Schiewe, Maxwell B. Eldridge, Steven T. Lindley, John H. Helle, and William R. Heard. A coordinated research plan for estuarine and ocean research on Pacific salmon. Fisheries, 25(6):7– 16, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0007%3AACRPFE%3E2.0.CO%3B2.

Beamesderfer:2000:MFP

[568] Raymond C. P. Beamesderfer. Managing fish predators and competitors: Deciding when intervention is effective and appropriate. Fisheries, 25 (6):18–23, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0018%3AMFPACD%3E2.0.CO%3B2.

Anonymous:2000:FIAg

[569] Anonymous. Full issue article volume 25, issue 7. Fisheries, 25(7):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-7.

Moring:2000:CFP

[570] John R. Moring. The creation of the first public salmon hatchery in the United States. Fisheries, 25(7):6–12, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%282000%29025%3C0006%3ATCOTFP%
3E2.0.CO%3B2.

Epifanio:2000:SCF

[571] John Epifanio. The status of coldwater fishery management in the United States: an overview of state programs. Fisheries, 25(7):13– 27, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0013%3ATSOCFM%3E2.0.CO%3B2.

Anonymous:2000:FIAh

[572] Anonymous. Full issue article volume 25, issue 8. Fisheries, 25(8):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-8.

Lodge:2000:NCT

[573] David M. Lodge, Christopher A. Taylor, David M. Holdich, and Jostein Skurdal. Nonindigenous crayfishes threaten North American freshwater biodiversity: Lessons from Europe. Fisheries, 25(8):7– 20, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0007%3ANCTNAF%3E2.0.CO%3B2.

Dauble:2000:UPM

[574] Dennis D. Dauble and Robert P. Mueller. Upstream passage monitoring: Difficulties in estimating survival for adult Chinook salmon in the Columbia and Snake Rivers. Fisheries, 25(8):24–34, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282000%29025%
3C0024%3AUPMDIE%3E2.0.CO%3B2.

Anonymous:2000:FIAi

[575] Anonymous. Full issue article volume 25, issue 9. Fisheries, 25(9):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-9.

Haro:2000:PDA

[576] Alex Haro, William Richkus, Kevin Whalen, Alex Hoar, W-Dieter Busch, Sandra Lary, Tim Brush, and Douglas Dixon. Population decline of the American eel: Implications for research and management. Fisheries, 25 (9):7–16, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0007%3APDOTAE%3E2.0.CO%3B2.

Tyus:2000:NFC

[577] Harold M. Tyus and James F. Saunders III. Nonnative fish control and endangered fish recovery: Lessons from the Colorado River. Fisheries, 25 (9):17–24, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0017%3ANFCAEF%3E2.0.CO%3B2.

Anonymous:2000:FIAj

[578] Anonymous. Full issue article volume 25, issue 10. Fisheries, 25(10):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-10.

Warren:2000:DDC

[579] Melvin L. Warren, Jr., Brooks M. Burr, Stephen J. Walsh, Henry L. Bart, Jr., Robert C. Cashner, David A. Etnier, Byron J. Freeman, Bernard R. Kuhajda, Richard L. Mayden, Henry W. Robison, Stephen T. Ross, and Wayne C. Starnes. Diversity, distribution, and conservation status of the native freshwater fishes of the Southern United States. Fisheries, 25 (10):7–31, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0007%3ADDACSO%3E2.0.CO%3B2.

Anonymous:2000:FIAk

[580] Anonymous. Full issue article volume 25, issue 11. Fisheries, 25(11):1–56, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-11.

Musick:2000:MED

[581] J. A. Musick, M. M. Harbin, S. A. Berkeley, G. H. Burgess, A. M. Eklund, L. Findley, R. G. Gilmore, J. T. Golden, D. S. Ha, G. R. Huntsman, J. C. McGovern, G. R. Sedberry, S. J. Parker, S. G. Poss, E. Sala, T. W. Schmidt, H. Weeks, and S. G. Wright. Marine, estuarine, and diadromous fish stocks at risk of extinction in North America (exclusive of Pacific salmonids). Fisheries, 25(11):6– 30, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0006%3AMEADFS%3E2.0.CO%3B2.

Anonymous:2000:FIAl

[582] Anonymous. Full issue article volume 25, issue 12. Fisheries, 25(12):1–48, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-25-12.

Weeks:2000:UPM

[583] Hal Weeks and Steve Berkeley. Uncertainty and precautionary management of marine fisheries: Can the old methods fit the new mandates? Fisheries, 25(12):6–15, 2000. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282000%29025%3C0006%3AUAPMOM%3E2.0.CO%3B2.

Pajak:2000:SEM

[584] Paul Pajak. Sustainability, ecosystem management, and indicators: Thinking globally and acting locally in the 21st century. Fisheries, 25 (12):16–30, 2000. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282000%29025%3C0016%3ASEMAIT%3E2.0.CO%3B2.

Anonymous:2001:FIAa

[585] Anonymous. Full issue article volume 26, issue 1. Fisheries, 26(1):1–48, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-1.

Bilby:2001:PEU

[586] Robert E. Bilby, Brian R. Fransen, Jason K. Walter, C. Jeff Cederholm, and Warren J. Scarlett. Preliminary evaluation of the use of nitrogen stable isotope ratios to establish escapement levels for Pacific salmon. Fisheries, 26(1):6–14, 2001. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282001%29026%3C0006%3APEOTUO%3E2.0.CO%3B2.

Lichtkoppler:2001:OLE

[587] Frank R. Lichtkoppler and Leroy Hushak. Ohio’s 1998 Lake Erie charter fishing industry. Fisheries, 26(1):15–23, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%282001%29026%3C0015%3AOLECFI%
3E2.0.CO%3B2.

Anonymous:2001:FIAb

[588] Anonymous. Full issue article volume 26, issue 2. Fisheries, 26(2):1–48, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-2.

Olney:2001:MFU

[589] John E. Olney and John M. Hoenig. Managing a fishery under moratorium: Assessment opportunities for Virginia’s stocks of American shad. Fisheries, 26(2):6–12, 2001. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282001%29026%3C0006%3AMAFUMA%3E2.0.CO%3B2.

Ross:2001:PTF

[590] Michael R. Ross and David K. Loomis. Put-and-take fisheries: Investigating catch and retention assumptions. Fisheries, 26(2):13–18, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282001%29026%3C0013%3APFICAR%3E2.0.CO%3B2.

Anonymous:2001:FIAc

[591] Anonymous. Full issue article volume 26, issue 3. Fisheries, 26(3):1–48, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-3.

Yoshiyama:2001:LTP

[592] Ronald M. Yoshiyama and Frank W. Fisher. Long time past: Baird Station and the McCloud Wintu. Fisheries, 26(3):6–22, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282001%29026%
3C0006%3ALTPBSA%3E2.0.CO%3B2.

Pope:2001:ATM

[593] Kevin L. Pope. Anglers tagging and marking fish: Provincial and state fishery agency views. Fisheries, 26(3):23–27, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282001%29026%
3C0023%3AATAMFP%3E2.0.CO%3B2.

Anonymous:2001:FIAd

[594] Anonymous. Full issue article volume 26, issue 4. Fisheries, 26(4):1–48, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-4.

Healey:2001:SEF

[595] Michael Healey, Paul Kline, and Tsai Chu-Fa. Saving the endangered Formosa landlocked salmon. Fisheries, 26(4):6–14, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282001%29026%
3C0006%3ASTEFLS%3E2.0.CO%3B2.

Ham:2001:PAC

[596] Kenneth D. Ham and Todd N. Pearsons. A practical approach for containing ecological risks associated with fish stocking programs. Fisheries, 26(4):15–23, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%282001%29026%3C0015%3AAPAFCE%3E2.0.CO%3B2.

Anonymous:2001:FIAe

[597] Anonymous. Full issue article volume 26, issue 5. Fisheries, 26(5):1–48, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-5.

Radomski:2001:VRF

[598] Paul J. Radomski, Gerold C. Grant, Peter C. Jacobson, and Mark F. Cook. Visions for recreational fishing regulations. Fisheries, 26(5): 7–18, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282001%29026%3C0007%3AVFRFR%3E2.0.CO%3B2.

Cook:2001:CLM

[599] Mark F. Cook, Timothy J. Goeman, Paul J. Radomski, Jerry A. Younk, and Peter C. Jacobson. Creel limits in Minnesota: a proposal for change. Fisheries, 26(5):19–26, 2001. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282001%29026%3C0019%3ACLIM%3E2.0.CO%3B2.

Anonymous:2001:FIAf

[600] Anonymous. Full issue article volume 26, issue 6. Fisheries, 26(6):1–72, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-6.

Paukert:2001:ONP

[601] Craig P. Paukert, Joel A. Klammer, Rodney B. Pierce, and Timothy D. Simonson. An overview of northern pike regulations in North America. Fisheries, 26(6):6–13, 2001. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282001%29026%3C0006%3AAOONPR%3E2.0.CO%3B2.

Bauer:2001:SUA

[602] Stephen B. Bauer and Stephen C. Ralph. Strengthening the use of aquatic habitat indicators in Clean Water Act programs. Fisheries, 26 (6):14–25, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282001%29026%3C0014%3ASTUOAH%3E2.0.CO%3B2.

Anonymous:2001:FIAg

[603] Anonymous. Full issue article volume 26, issue 7. Fisheries, 26(7):1–48, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-7.

Anonymous:2001:FIAh

[604] Anonymous. Full issue article volume 26, issue 8. Fisheries, 26(8):1–56, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-8.

Sommer:2001:CYB

[605] Ted Sommer, Bill Harrell, Matt Nobriga, Randall Brown, Peter Moyle, Wim Kimmerer, and Larry Schemel. California’s Yolo Bypass: Evidence that flood control can be compatible with fisheries, wetlands, wildlife, and agriculture. Fisheries, 26(8):6–16, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282001%29026%3C0006%3ACYB%3E2.0.CO%3B2.

Grimes:2001:FPM

[606] Churchill B. Grimes. Fishery production and the Mississippi River discharge. Fisheries, 26(8):17–26, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%282001%29026%3C0017%3AFPATMR%3E2.
0.CO%3B2.

Anonymous:2001:FIAi

[607] Anonymous. Full issue article volume 26, issue 9. Fisheries, 26(9):1–48, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-9.

Parasiewicz:2001:MCA

[608] Piotr Parasiewicz. MesoHABSIM: a concept for application of instream flow models in river restoration planning. Fisheries, 26(9):6– 13, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282001%29026%3C0006%3AM%3E2.0.CO%3B2.

Cada:2001:DAH

[609] Glenn F. Cada. The development of advanced hydroelectric turbines to improve fish passage survival. Fisheries, 26(9):14–23, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282001%29026%
3C0014%3ATDOAHT%3E2.0.CO%3B2.

Anonymous:2001:FIAj

[610] Anonymous. Full issue article volume 26, issue 10. Fisheries, 26(10):1–48, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-10.

Harms:2001:CCP

[611] John Harms and Gil Sylvia. A comparison of conservation perspectives between scientists, managers, and industry in the West Coast groundfish fishery. Fisheries, 26(10):6–15, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%282001%29026%3C0006%3AACOCPB%3E2.
0.CO%3B2.

Jacob:2001:LDF

[612] Steve Jacob, Frank L. Farmer, Michael Jepson, and Charles Adams. Landing a definition of fishing dependent communities: Potential social science contributions to meeting National Standard 8. Fisheries, 26(10): 16–22, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282001%29026%3C0016%3ALADOFD%3E2.0.CO%3B2.

Sala:2001:RDN

[613] Enric Sala, Enric Ballesteros, and Richard M. Starr. Rapid decline of Nassau grouper spawning aggregations in Belize: Fishery management and conservation needs. Fisheries, 26(10):23–30, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282001%29026%3C0023%3ARDONGS%3E2.0.CO%3B2.

Anonymous:2001:FIAk

[614] Anonymous. Full issue article volume 26, issue 11. Fisheries, 26(11):1–48, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-11.

Scott:2001:NIH

[615] Mark C. Scott and Gene S. Helfman. Native invasions, homogenization, and the mismeasure of integrity of fish assemblages. Fisheries, 26(11): 6–15, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282001%29026%3C0006%3ANIHATM%3E2.0.CO%3B2.

Wright:2001:RIS

[616] Angela S. Wright and Bruce Shindler. The role of information sources in watershed management. Fisheries, 26(11):16–23, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282001%29026%
3C0016%3ATROISI%3E2.0.CO%3B2.

Anonymous:2001:FIAl

[617] Anonymous. Full issue article volume 26, issue 12. Fisheries, 26(12):1–48, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-26-12.

Gemeinhardt:2001:FIT

[618] Todd R. Gemeinhardt, Douglas B. Noltie, Justin M. Nolan, and Michael C. Robbins. Fishing for ideas: Teaching and learning in a university ichthyology course. Fisheries, 26(12):6–14, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282001%29026%
3C0006%3AFFI%3E2.0.CO%3B2.

Kim:2001:ARK

[619] Chang Gil Kim. Artificial reefs in Korea. Fisheries, 26(12):15– 18, 2001. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282001%29026%3C0015%3AARIK%3E2.0.CO%3B2.

Anonymous:2002:FIAa

[620] Anonymous. Full issue article volume 27, issue 1. Fisheries, 27(1):1–48, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-1.

Post:2002:CRF

[621] John R. Post, Michael Sullivan, Sean Cox, Nigel P. Lester, Carl J. Walters, Eric A. Parkinson, Andrew J. Paul, Leyland Jackson, and Brian J. Shuter. Canada’s recreational fisheries: The invisible collapse? Fisheries, 27(1):6–17, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282002%29027%3C0006%3ACRF%3E2.0.CO%3B2.

Anonymous:2002:FIAb

[622] Anonymous. Full issue article volume 27, issue 2. Fisheries, 27(2):1–52, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-2.

Churchill:2002:ACF

[623] Timothy N. Churchill, Phillip W. Bettoli, Douglas C. Peterson, William C. Reeves, and Bob Hodge. Angler conflicts in fisheries management: a case study of the striped bass controversy at Norris Reservoir, Tennessee. Fisheries, 27(2):10–19, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282002%29027%
3C0010%3AACIFM%3E2.0.CO%3B2.

Collins:2002:SMR

[624] Mark R. Collins, Theodore J. I. Smith, Wallace E. Jenkins, and Michael R. Denson. Small marine reserves may increase escapement of red drum. Fisheries, 27(2):20–24, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%282002%29027%3C0020%3ASMRMIE%3E2.
0.CO%3B2.

Anonymous:2002:FIAc

[625] Anonymous. Full issue article volume 27, issue 3. Fisheries, 27(3):1–52, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-3.

Bonar:2002:SSI

[626] Scott A. Bonar and Wayne A. Hubert. Standard sampling of inland fish: Benefits, challenges, and a call for action. Fisheries, 27(3):10– 16, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282002%29027%3C0010%3ASSOIF%3E2.0.CO%3B2.

Ditton:2002:RFT

[627] Robert B. Ditton, Stephen M. Holland, and David K. Anderson. Recreational fishing as tourism. Fisheries, 27(3):17–24, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282002%29027%
3C0017%3ARFAT%3E2.0.CO%3B2.

Anonymous:2002:FIAd

[628] Anonymous. Full issue article volume 27, issue 4. Fisheries, 27(4):1–48, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-4.

Link:2002:ECF

[629] Jason S. Link. Ecological considerations in fisheries management: When does it matter? Fisheries, 27(4):10–17, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282002%29027%
3C0010%3AECIFM%3E2.0.CO%3B2.

Anonymous:2002:FIAe

[630] Anonymous. Full issue article volume 27, issue 5. Fisheries, 27(5):1–48, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-5.

Arterburn:2002:SAA

[631] John E. Arterburn, Daniel J. Kirby, and Charles R. Berry, Jr. A survey of angler attitudes and biologist opinions regarding trophy catfish and their management. Fisheries, 27(5):10–21, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%282002%29027%3C0010%3AASOAAA%
3E2.0.CO%3B2.

Tugend:2002:UAH

[632] Kimberly I. Tugend, Micheal S. Allen, and Mark Webb. Use of artificial habitat structures in U.S. lakes and reservoirs: a survey from the Southern Division AFS Reservoir Committee. Fisheries, 27(5):22– 27, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282002%29027%3C0022%3AUOAHSI%3E2.0.CO%3B2.

Anonymous:2002:FIAf

[633] Anonymous. Full issue article volume 27, issue 6. Fisheries, 27(6):1–72, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-6.

Finlayson:2002:AAU

[634] Brian J. Finlayson, Rosalie A. Schnick, Richard L. Cailteux, Leo De-Mong, William D. Horton, William McClay, and Charles W. Thompson. Assessment of antimycin a use in fisheries and its potential for reregistration. Fisheries, 27(6):10–18, 2002. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282002%29027%3C0010%3AAOAAUI%3E2.0.CO%3B2.

Moring:2002:MEA

[635] John R. Moring. Marking experiments with Atlantic salmon in the United States in the 1870s and 1880s. Fisheries, 27(6):21–25, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282002%29027%3C0021%3AMEWASI%3E2.0.CO%3B2.

Anonymous:2002:FIAg

[636] Anonymous. Full issue article volume 27, issue 7. Fisheries, 27(7):1–48, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-7.

Buklis:2002:SFM

[637] Lawrence S. Buklis. Subsistence fisheries management on Federal public lands in Alaska. Fisheries, 27(7):10–18, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%282002%29027%3C0010%3ASFMOFP%
3E2.0.CO%3B2.

Close:2002:ECI

[638] David A. Close, Martin S. Fitzpatrick, and Hiram W. Li. The ecological and cultural importance of a species at risk of extinction, Pacific lamprey. Fisheries, 27(7):19–25, 2002. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282002%29027%3C0019%3ATEACIO%3E2.0.CO%3B2.

Anonymous:2002:FIAh

[639] Anonymous. Full issue article volume 27, issue 8. Fisheries, 27(8):1–56, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-8.

Wilson:2002:IDA

[640] Douglas Clyde Wilson, Bonnie J. McCay, Veronica Rowan, and Barbara Grandin. Institutional differences among marine fisheries scientists’ views of their working conditions, discipline, and fisheries management. Fisheries, 27(8):14–24, 2002. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282002%29027%3C0014%3AIDAMFS%3E2.0.CO%3B2.

Chittenden:2002:GST

[641] Mark E. Chittenden, Jr. Given a significance test, how large a sample size is large enough? Fisheries, 27(8):25–29, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%282002%29027%3C0025%3AGASTHL%
3E2.0.CO%3B2.

Anonymous:2002:FIAi

[642] Anonymous. Full issue article volume 27, issue 9. Fisheries, 27(9):1–48, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-9.

Schmetterling:2002:NSC

[643] David A. Schmetterling and Tina Bernd-Cohen. Native species conservation through education: The Adopt-A-Trout program in Montana. Fisheries, 27(9):10–15, 2002. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282002%29027%3C0010%3ANSCTE%3E2.0.CO%3B2.

Paukert:2002:AMS

[644] Craig P. Paukert and Timothy A. Wittig. Applications of multivariate statistical methods in fisheries. Fisheries, 27(9):16–22, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282002%29027%3C0016%3AAOMSMI%3E2.0.CO%3B2.

Anonymous:2002:FIAj

[645] Anonymous. Full issue article volume 27, issue 10. Fisheries, 27(10):1–48, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-10.

Hayes:2002:SRU

[646] Michael C. Hayes and Richard W. Carmichael. Salmon restoration in the Umatilla River: a study of straying and risk containment. Fisheries, 27 (10):10–19, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282002%29027%3C0010%3ASRITUR%3E2.0.CO%3B2.

Conway:2002:CCR

[647] Flaxen D. L. Conway, Jennifer Gilden, and Anisa Zvonkovic. Changing communication and roles: Innovations in Oregon’s fishing families, communities, and management. Fisheries, 27(10):20–29, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446%282002%29027%
3C0020%3ACCAR%3E2.0.CO%3B2.

Anonymous:2002:FIAk

[648] Anonymous. Full issue article volume 27, issue 11. Fisheries, 27(11):1–48, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-11.

Buckmeier:2002:ARA

[649] David L. Buckmeier. Assessment of reader accuracy and recommendations to reduce subjectivity in age estimation. Fisheries, 27(11): 10–14, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282002%29027%3C0010%3AAORAAR%3E2.0.CO%3B2.

Anonymous:2002:FIAl

[650] Anonymous. Full issue article volume 27, issue 12. Fisheries, 27(12):1–48, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-27-12.

Pearsons:2002:CEI

[651] Todd N. Pearsons. Chronology of ecological interactions associated with the life-span of salmon supplementation programs. Fisheries, 27(12): 10–15, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282002%29027%3C0010%3ACOEIAW%3E2.0.CO%3B2.

Bisson:2002:HSP

[652] Peter A. Bisson, Charles C. Coutant, Daniel Goodman, Robert Gramling, Dennis Lettenmaier, James Lichatowich, William Liss, Eric Loudenslager, Lyman McDonald, David Philipp, and Brian Riddell. Hatchery surpluses in the Pacific Northwest. Fisheries, 27(12):16– 27, 2002. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282002%29027%3C0016%3AHSITPN%3E2.0.CO%3B2.

Anonymous:2003:FIAa

[653] Anonymous. Full issue article volume 28, issue 1. Fisheries, 28(1):1–48, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-1.

Peterson:2003:QDA

[654] James T. Peterson and James W. Evans. Quantitative decision analysis for sport fisheries management. Fisheries, 28(1):10–21, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282003%2928%5B10%3AQDAFSF%5D2.0.CO%3B2.

Wiley:2003:PTW

[655] Robert W. Wiley. Planting trout in Wyoming high-elevation wilderness waters. Fisheries, 28(1):22–27, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282003%2928%5B22%
3APTIWHW%5D2.0.CO%3B2.

Anonymous:2003:FIAb

[656] Anonymous. Full issue article volume 28, issue 2. Fisheries, 28(2):1–48, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-2.

Fisher:2003:SFM

[657] William L. Fisher and Jim P. Burroughs. Stream fisheries management in the United States. Fisheries, 28(2):10–18, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282003%2928%
5B10%3ASFMITU%5D2.0.CO%3B2.

Anonymous:2003:FIAc

[658] Anonymous. Full issue article volume 28, issue 3. Fisheries, 28(3):1–48, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-3.

Cooperman:2003:ESA

[659] Michael S. Cooperman and Douglas F. Markle. The Endangered Species Act and the National Research Council’s interim judgment in Klamath Basin. Fisheries, 28(3):10–19, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282003%2928%5B10%
3ATESAAT%5D2.0.CO%3B2.

Kerr:2003:CFF

[660] Steven J. Kerr and Kendall K. Kamke. Competitive fishing in freshwaters of North America. Fisheries, 28(3):26–31, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282003%2928%5B26%
3ACFIFON%5D2.0.CO%3B2.

Anonymous:2003:FIAd

[661] Anonymous. Full issue article volume 28, issue 4. Fisheries, 28(4):1–48, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-4.

Sharp:2003:ISV

[662] Shayla B. Sharp and Denise Lach. Integrating social values into fisheries management. Fisheries, 28(4):10–15, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282003%2928%5B10%
3AISVIFM%5D2.0.CO%3B2.

Anonymous:2003:FIAe

[663] Anonymous. Full issue article volume 28, issue 5. Fisheries, 28(5):1–60, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-5.

Foster:2003:CCT

[664] Susan A. Foster, John A. Baker, and Michael A. Bell. The case for conserving threespine stickleback populations. Fisheries, 28(5):10– 18, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282003%2928%5B10%3ATCFCTS%5D2.0.CO%3B2.

Wood:2003:WCP

[665] Paul M. Wood. Will Canadian policies protect British Columbia’s endangered pairs of sympatric sticklebacks? Fisheries, 28(5):19– 26, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282003%2928%5B19%3AWCPPBC%5D2.0.CO%3B2.

Anonymous:2003:FIAf

[666] Anonymous. Full issue article volume 28, issue 6. Fisheries, 28(6):1–72, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-6.

Arlinghaus:2003:MPU

[667] Robert Arlinghaus and Thomas Mehner. Management preferences of urban anglers. Fisheries, 28(6):10–17, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%282003%2928%5B10%3AMPOUA%5D2.
0.CO%3B2.

Wilde:2003:LSR

[668] Gene R. Wilde, Kevin L. Pope, and Bart W. Durham. Luresize restrictions in recreational fisheries. Fisheries, 28(6):18–26, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282003%2928%5B18%3ALRIRF%5D2.0.CO%3B2.

Anonymous:2003:FIAg

[669] Anonymous. Full issue article volume 28, issue 7. Fisheries, 28(7):1–48, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-7.

Wilde:2003:DTC

[670] Gene R. Wilde. Dispersal of tournament-caught black bass. Fisheries, 28 (7):10–17, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282003%2928%5B10%3ADOTBB%5D2.0.CO%3B2.

Hayes:2003:DSS

[671] D. Hayes, E. Baker, R. Bednarz, D. Borgeson, Jr., J. Braunscheidel, A. Harrington, R. Hay, J. Waybrant, J. Breck, A. Nuhfer, M. Bremigan, R. Lockwood, J. Schneider, P. Seelbach, and T. Zorn. Developing a standardized sampling program. Fisheries, 28(7):18–25, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282003%2928%5B18%3ADASSP%5D2.0.CO%3B2.

Anonymous:2003:FIAh

[672] Anonymous. Full issue article volume 28, issue 8. Fisheries, 28(8):1–56, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-8.

Epifanio:2003:ICG

[673] John Epifanio, Gordon Haas, Karen Pratt, Bruce Rieman, Paul Spruell, Craig Stockwell, Fred Utter, and William Young. Integrating conservation genetic considerations into conservation planning. Fisheries, 28(8): 10–24, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282003%2928%5B10%3AICGCIC%5D2.0.CO%3B2.

Irvine:2003:RS

[674] James R. Irvine, Din (Ding Geng) Chen, and Jon T. Schnute. Retrospective sampling. Fisheries, 28(8):25–30, 2003. CODEN ????

ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282003%2928%5B25%
3ARS%5D2.0.CO%3B2.

Anonymous:2003:FIAi

[675] Anonymous. Full issue article volume 28, issue 9. Fisheries, 28(9):1–48, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-9.

Latour:2003:TEB

[676] Robert J. Latour, Mark J. Brush, and Christopher F. Bonzek. Toward ecosystem-based fisheries management. Fisheries, 28(9):10–22, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282003%2928%5B10%3ATEFM%5D2.0.CO%3B2.

Anonymous:2003:FIAj

[677] Anonymous. Full issue article volume 28, issue 10. Fisheries, 28(10):1–48, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-10.

Pine:2003:RTM

[678] William E. Pine, Kenneth H. Pollock, Joseph E. Hightower, Thomas J. Kwak, and James A. Rice. A review of tagging methods for estimating fish population size and components of mortality. Fisheries, 28(10): 10–23, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282003%2928%5B10%3AAROTMF%5D2.0.CO%3B2.

Popper:2003:EAS

[679] Arthur N. Popper. Effects of anthropogenic sounds on fishes. Fisheries, 28(10):24–31, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%282003%2928%5B24%3AEOASOF%5D2.0.CO%3B2.

Anonymous:2003:FIAk

[680] Anonymous. Full issue article volume 28, issue 11. Fisheries, 28(11):1–48, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-11.

Grizzle:2003:RLB

[681] John M. Grizzle and Cindy J. Brunner. Review of largemouth bass virus. Fisheries, 28(11):10–14, 2003. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282003%2928%5B10%3AROLBV%5D2.0.CO%3B2.

Born:2003:OSF

[682] Stephen M. Born and G. Simeon Stairs. An overview of salmonid fisheries planning by state agencies in the United States. Fisheries, 28(11): 15–25, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282003%2928%5B15%3AAOOSFP%5D2.0.CO%3B2.

Anonymous:2003:FIAl

[683] Anonymous. Full issue article volume 28, issue 12. Fisheries, 28(12):1–48, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-28-12.

Shipp:2003:PMR

[684] Robert L. Shipp. A perspective on marine reserves as a fishery management tool. Fisheries, 28(12):10–21, 2003. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282003%2928%5B10%
3AAPOMRA%5D2.0.CO%3B2.

Anonymous:2004:FIAa

[685] Anonymous. Full issue article volume 29, issue 1. Fisheries, 29(1):1–48, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-1.

Ames:2004:ACS

[686] Edward P. Ames. Atlantic cod stock structure in the Gulf of Maine. Fisheries, 29(1):10–28, 2004. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282004%2929%5B10%3AACSSIT%5D2.0.CO%3B2.

Anonymous:2004:FIAb

[687] Anonymous. Full issue article volume 29, issue 2. Fisheries, 29(2):1–48, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-2.

Coleman:2004:RGM

[688] Felicia C. Coleman, Pamela B. Baker, and Christopher C. Koenig. A review of Gulf of Mexico marine protected areas. Fisheries, 29(2):10– 21, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282004%2929%5B10%3AAROGOM%5D2.0.CO%3B2.

Cooke:2004:AMT

[689] Steven J. Cooke, Scott G. Hinch, Anthony P. Farrell, Michael F. Lapointe, Simon R. M. Jones, J. Steveson Macdonald, David A. Patterson, Michael C. Healey, and Glen Van Der Kraak. Abnormal migration timing and high en route mortality of sockeye salmon in the Fraser River, British Columbia. Fisheries, 29(2):22– 33, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282004%2929%5B22%3AAMTAHE%5D2.0.CO%3B2.

Anonymous:2004:FIAc

[690] Anonymous. Full issue article volume 29, issue 3. Fisheries, 29(3):1–48, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-3.

Ward:2004:RFC

[691] Neil E. Ward and David L. Ward. Resident fish in the Columbia River Basin. Fisheries, 29(3):10–18, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%282004%2929%5B10%3ARFITCR%5D2.0.CO%
3B2.

Sala:2004:FCF

[692] Enric Sala, Octavio Aburto-Oropeza, Miriam Reza, Gustavo Paredes, and Luis G. López-Lemus. Fishing down coastal food webs in the Gulf of California. Fisheries, 29(3):19–25, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282004%2929%5B19%
3AFDCFWI%5D2.0.CO%3B2.

Anonymous:2004:FIAd

[693] Anonymous. Full issue article volume 29, issue 4. Fisheries, 29(4):1–48, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-4.

Jesus:2004:CFP

[694] Marcos Javier De Jesús and Christopher C. Kohler. The commercial fishery of the Peruvian Amazon. Fisheries, 29(4):10–16, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282004%2929%5B10%3ATCFOTP%5D2.0.CO%3B2.

Olson:2004:ESV

[695] Nathan W. Olson, Stephen K. Wilson, and David W. Willis. Effect of spatial variation on zooplankton community assessment in fishery studies. Fisheries, 29(4):17–22, 2004. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282004%2929%5B17%3AEOSVOZ%5D2.0.CO%3B2.

Anonymous:2004:FIAe

[696] Anonymous. Full issue article volume 29, issue 5. Fisheries, 29(5):1–64, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-5.

Cimo:2004:EEC

[697] Laura Faitel Cimo and Tracy Dobson. Evaluation of the effectiveness of court-ordered fisheries management policies. Fisheries, 29(5): 10–18, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282004%2929%5B10%3AEOTEOC%5D2.0.CO%3B2.

Goodwin:2004:TEF

[698] Andrew E. Goodwin, James E. Peterson, Theodore R. Meyers, and David J. Money. Transmission of exotic fish viruses. Fisheries, 29(5): 19–23, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282004%2929%5B19%3ATOEFV%5D2.0.CO%3B2.

Anonymous:2004:FIAf

[699] Anonymous. Full issue article volume 29, issue 6. Fisheries, 29(6):1–72, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-6.

Salz:2004:SAA

[700] Ronald J. Salz and David K. Loomis. Saltwater anglers’ attitudes towards marine protected areas. Fisheries, 29(6):10–17, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282004%2929%
5B10%3ASAATMP%5D2.0.CO%3B2.

Dunham:2004:ACN

[701] Jason B. Dunham, David S. Pilliod, and Michael K. Young. Assessing the consequences of nonnative trout in headwater ecosystems in Western North America. Fisheries, 29(6):18–26, 2004. CODEN ????

ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282004%2929%5B18%
3AATCONT%5D2.0.CO%3B2.

Anonymous:2004:FIAg

[702] Anonymous. Full issue article volume 29, issue 7. Fisheries, 29(7):1–48, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-7.

McMichael:2004:CFS

[703] Geoffrey A. McMichael, Jessica A. Vucelick, Cary S. Abernethy, and Duane A. Neitzel. Comparing fish screen performance to physical design criteria. Fisheries, 29(7):10–16, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282004%2929%5B10%
3ACFSPTP%5D2.0.CO%3B2.

Strange:2004:HBR

[704] Elizabeth M. Strange, P. David Allen, Douglas Beltman, Joshua Lipton, and David Mills. The habitat-based replacement cost method for assessing monetary damages for fish resource injuries. Fisheries, 29(7): 17–24, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282004%2929%5B17%3ATHRCMF%5D2.0.CO%3B2.

Anonymous:2004:FIAh

[705] Anonymous. Full issue article volume 29, issue 8. Fisheries, 29(8):1–60, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-8.

Williams:2004:BVE

[706] Lance R. Williams, Melvin L. Warren, Jr., Susan B. Adams, Joseph L. Arvai, and Christopher M. Taylor. Basin Visual Estimation Technique (BVET) and representative reach approaches to wadeable stream surveys. Fisheries, 29(8):12–22, 2004. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282004%2929%5B12%3ABVETBA%5D2.0.CO%3B2.

Berkeley:2004:FSP

[707] Steven A. Berkeley, Mark A. Hixon, Ralph J. Larson, and Milton S. Love. Fisheries sustainability via protection of age structure and spatial distribution of fish populations. Fisheries, 29(8):23– 32, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282004%2929%5B23%3AFSVPOA%5D2.0.CO%3B2.

Anonymous:2004:FIAi

[708] Anonymous. Full issue article volume 29, issue 9. Fisheries, 29(9):1–48, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-9.

Brannon:2004:CAS

[709] Ernest L. Brannon, Donald F. Amend, Matthew A. Cronin, James E. Lannan, Scott LaPatra, William J. McNeil, Richard E. Noble, Charlie E. Smith, André J. Talbot, Gary A. Wedemeyer, and Harry Westers. The controversy about salmon hatcheries. Fisheries, 29(9):12– 31, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282004%2929%5B12%3ATCASH%5D2.0.CO%3B2.

Anonymous:2004:FIAj

[710] Anonymous. Full issue article volume 29, issue 10. Fisheries, 29(10):1–48, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-10.

Leslie:2004:UVF

[711] Laura L. Leslie, Cristina E. Velez, and Scott A. Bonar. Utilizing volunteers on fisheries projects. Fisheries, 29(10):10–14, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282004%2929%
5B10%3AUVOFP%5D2.0.CO%3B2.

Barnes:2004:VPP

[712] Michael E. Barnes and Gary E. Whelan. Visitor policies and practices at public trout and salmon hatcheries in North America. Fisheries, 29(10): 16–21, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282004%2929%5B16%3AVPAPAP%5D2.0.CO%3B2.

Anonymous:2004:FIAk

[713] Anonymous. Full issue article volume 29, issue 11. Fisheries, 29(11):1–48, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-11.

Clapp:2004:YPR

[714] David F. Clapp and John M. Dettmers. Yellow perch research and management in Lake Michigan. Fisheries, 29(11):11–19, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282004%2929%
5B11%3AYPRAMI%5D2.0.CO%3B2.

Fraidenburg:2004:DC

[715] Michael Fraidenburg and Linda Strever. Diagnosing conflict. Fisheries, 29(11):20–29, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%282004%2929%5B20%3ADC%5D2.0.CO%3B2.

Anonymous:2004:FIAl

[716] Anonymous. Full issue article volume 29, issue 12. Fisheries, 29(12):1–48, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-29-12.

Cooke:2004:TEO

[717] Steven J. Cooke and Glenn N. Wagner. Training, experience, and opinions of researchers who use surgical techniques to implant telemetry devices into fish. Fisheries, 29(12):10–18, 2004. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282004%2929%5B10%
3ATEAOOR%5D2.0.CO%3B2.

Angermeier:2004:CFA

[718] Paul L. Angermeier, Andrew P. Wheeler, and Amanda E. Rosenberger. A conceptual framework for assessing impacts of roads on aquatic biota. Fisheries, 29(12):19–29, 2004. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282004%2929%5B19%3AACFFAI%5D2.0.CO%3B2.

Anonymous:2005:FIAa

[719] Anonymous. Full issue article volume 30, issue 1. Fisheries, 30(1):1–48, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-1.

Gibson:2005:LFH

[720] R. John Gibson, Richard L. Haedrich, and C. Michael Wernerheim. Loss of fish habitat as a consequence of inappropriately constructed stream crossings. Fisheries, 30(1):10–17, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282005%2930%5B10%
3ALOFHAA%5D2.0.CO%3B2.

Anonymous:2005:FIAb

[721] Anonymous. Full issue article volume 30, issue 2. Fisheries, 30(2):1–48, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-2.

Gabelhouse:2005:SSF

[722] Donald W. Gabelhouse, Jr. Staffing, spending, and funding of state inland fisheries programs. Fisheries, 30(2):10–17, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282005%2930%5B10%
3ASSAFOS%5D2.0.CO%3B2.

Harper:2005:CAE

[723] David J. Harper and Jason T. Quigley. A comparison of the areal extent of fish habitat gains and losses associated with selected compensation projects in Canada. Fisheries, 30(2):18–25, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B18%3AACOTAE%5D2.0.CO%3B2.

Fitzgerald:2005:QSA

[724] D. G. Fitzgerald, B. Zhu, S. B. Hoskins, D. E. Haddad, K. N. Green, L. G. Rudstam, and E. L. Mills. Quantifying submerged aquatic vegetation using aerial photograph interpretation. Fisheries, 30(2):61– 73, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B61%3AQSAVUA%5D2.0.CO%3B2.

Quist:2005:WHA

[725] Michael C. Quist, Wayne A. Hubert, Mark Fowden, Steven W. Wolff, and Michael R. Bower. The Wyoming Habitat Assessment Methodology (WHAM). Fisheries, 30(2):75–81, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282005%2930%5B75%
3ATWHAMW%5D2.0.CO%3B2.

Anonymous:2005:FIAc

[726] Anonymous. Full issue article volume 30, issue 3. Fisheries, 30(3):1–48, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-3.

Kuehn:2005:GLC

[727] Diane Kuehn, Frank Lichtkoppler, and Charles Pistis. The Great Lakes charter fishing industry. Fisheries, 30(3):10–17, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282005%2930%5B10%
3ATGLCFI%5D2.0.CO%3B2.

Reinert:2005:DPR

[728] Thomas R. Reinert, Cecil A. Jennings, Ted A. Will, and Julie E. Wallin. Decline and potential recovery of striped bass in a Southeastern U.S. estuary. Fisheries, 30(3):18–25, 2005. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282005%2930%5B18%3ADAPROS%5D2.0.CO%3B2.

Anonymous:2005:FIAd

[729] Anonymous. Full issue article volume 30, issue 4. Fisheries, 30(4):1–48, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-4.

Hamilton:2005:DAF

[730] John B. Hamilton, Gary L. Curtis, Scott M. Snedaker, and David K. White. Distribution of anadromous fishes in the Upper Klamath River Watershed prior to hydropower dams — a synthesis of the historical evidence. Fisheries, 30(4):10–20, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282005%2930%5B10%
3ADOAFIT%5D2.0.CO%3B2.

Habron:2005:ICL

[731] Geoffrey Habron. Infusing constructivist learning in fisheries education. Fisheries, 30(4):21–26, 2005. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282005%2930%5B21%3AICLIFE%5D2.0.CO%3B2.

McClay:2005:RUN

[732] William McClay. Rotenone use in North America (1988–2002). Fisheries, 30(4):29–31, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%282005%2930%5B29%3ARUINA%5D2.0.CO%3B2.

Anonymous:2005:FIAe

[733] Anonymous. Full issue article volume 30, issue 5. Fisheries, 30(5):1–64, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-5.

Finlayson:2005:NIT

[734] Brian Finlayson, William Somer, Dan Duffield, David Propst, Chad Mellison, Tom Pettengill, Hilda Sexauer, Tom Nesler, Scott Gurtin, John Elliot, Fred Partridge, and Don Skaar. Native inland trout restoration on National Forests in the Western United States. Fisheries, 30 (5):10–19, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B10%3ANITRON%5D2.0.CO%3B2.

Moyle:2005:UA

[735] Peter B. Moyle and Joshua A. Israel. Untested assumptions. Fisheries, 30 (5):20–28, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B20%3AUA%5D2.0.CO%3B2.

Anonymous:2005:FIAf

[736] Anonymous. Full issue article volume 30, issue 6. Fisheries, 30(6):1–72, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-6.

Mobrand:2005:HRW

[737] Lars E. Mobrand, John Barr, Lee Blankenship, Donald E. Campton, Trevor T. P. Evelyn, Tom A. Flagg, Conrad V. W. Mahnken, Lisa W. Seeb, Paul R. Seidel, and William W. Smoker. Hatchery reform in Washington State. Fisheries, 30(6):11–23, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%282005%2930%5B11%3AHRIWS%5D2.
0.CO%3B2.

Williams:2005:PRE

[738] Jack E. Williams, Catherine A. Macdonald, Cindy Deacon Williams Fishery consultant in Medford, Hal Weeks, Georgina Lampman, and Don W. Sada. Prospects for recovering endemic fishes pursuant to the U.S. Endangered Species Act. Fisheries, 30(6):24–29, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282005%2930%5B24%
3APFREFP%5D2.0.CO%3B2.

Anonymous:2005:FIAg

[739] Anonymous. Full issue article volume 30, issue 7. Fisheries, 30(7):1–48, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-7.

Habera:2005:MSA

[740] Jim Habera and Steve Moore. Managing Southern Appalachian brook trout. Fisheries, 30(7):10–20, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446%282005%2930%5B10%3AMSABT%5D2.
0.CO%3B2.

Kerr:2005:IAS

[741] Steven J. Kerr, Christopher S. Brousseau, and Mark Muschett. Invasive aquatic species in Ontario. Fisheries, 30(7):21–30, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282005%2930%
5B21%3AIASIO%5D2.0.CO%3B2.

Anonymous:2005:FIAh

[742] Anonymous. Full issue article volume 30, issue 8. Fisheries, 30(8):1–56, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-8.

Mitchell:2005:SEF

[743] Andrew J. Mitchell, Robin M. Overstreet, Andrew E. Goodwin, and Thomas M. Brandt. Spread of an exotic fish-gill trematode. Fisheries, 30 (8):11–16, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B11%3ASOAEFT%5D2.0.CO%3B2.

Dauwalter:2005:REU

[744] Daniel C. Dauwalter and John R. Jackson. A re-evaluation of U.S. state fish-stocking recommendations for small, private, warmwater impoundments. Fisheries, 30(8):18–28, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282005%2930%5B18%
3AAROUSF%5D2.0.CO%3B2.

Anonymous:2005:FIAi

[745] Anonymous. Full issue article volume 30, issue 9. Fisheries, 30(9):1–48, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-9.

Mueller:2005:PFR

[746] Gordon A. Mueller. Predatory fish removal and native fish recovery in the Colorado River mainstem. Fisheries, 30(9):10–19, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B10%3APFRANF%5D2.0.CO%3B2.

Clarkson:2005:CBN

[747] Robert W. Clarkson, Paul C. Marsh, Sally E. Stefferud, and Jerome A. Stefferud. Conflicts between native fish and nonnative sport fish management in the Southwestern United States. Fisheries, 30(9):20– 27, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B20%3ACBNFAN%5D2.0.CO%3B2.

Anonymous:2005:FIAj

[748] Anonymous. Full issue article volume 30, issue 10. Fisheries, 30(10):1–48, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-10.

McBride:2005:TVC

[749] Richard S. McBride, Michael L. Hendricks, and John E. Olney. Testing the validity of Cating’s (1953) method for age determination of American shad using scales. Fisheries, 30(10):10–18, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B10%3ATTVOCM%5D2.0.CO%3B2.

Burgess:2005:CSP

[750] George H. Burgess, Lawrence R. Beerkircher, Gregor M. Cailliet, John K. Carlson, Enris Cortés, Kenneth J. Goldman, R. Dean Grubbs, John A. Musick, Michael K. Musyl, and Colin A. Simpfendorfer. Is the collapse of shark populations in the Northwest Atlantic Ocean and Gulf of Mexico real? Fisheries, 30(10):19–26, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B19%3AITCOSP%5D2.0.CO%3B2.

Anonymous:2005:FIAk

[751] Anonymous. Full issue article volume 30, issue 11. Fisheries, 30(11):1–48, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-11.

Koel:2005:NLT

[752] Todd M. Koel, Patricia E. Bigelow, Philip D. Doepke, Brian D. Ertel, and Daniel L. Mahony. Nonnative lake trout result in Yellowstone cutthroat trout decline and impacts to bears and anglers. Fisheries, 30(11): 10–19, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B10%3ANLTRIY%5D2.0.CO%3B2.

Warner:2005:SBI

[753] Kendall Warner. Smallmouth bass introductions in Maine. Fisheries, 30 (11):20–26, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B20%3ASBIIM%5D2.0.CO%3B2.

Anonymous:2005:FIAl

[754] Anonymous. Full issue article volume 30, issue 12. Fisheries, 30(12):1–48, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-30-12.

Irvine:2005:CSR

[755] James R. Irvine, Mart R. Gross, Chris C. Wood, L. Blair Holtby, Neil D. Schubert, and Peter G. Amiro. Canada’s Species at Risk Act. Fisheries, 30(12):11–19, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%282005%2930%5B11%3ACSARA%5D2.0.CO%3B2.

Fernandez:2005:MAS

[756] Stephen M. Fernandez. A market approach to sturgeon conservation under the U.S. Endangered Species Act. Fisheries, 30(12):20– 27, 2005. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282005%2930%5B20%3AAMATSC%5D2.0.CO%3B2.

Anonymous:2006:FIAa

[757] Anonymous. Full issue article volume 31, issue 1. Fisheries, 31(1):1–52, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-31-1.

Drake:2006:FPD

[758] John M. Drake and David M. Lodge. Forecasting potential distributions of nonindigenous species with a genetic algorithm. Fisheries, 31 (1):9–16, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282006%2931%5B9%3AFPDONS%5D2.0.CO%3B2.

Reeve:2006:BSA

[759] Todd Reeve, Jim Lichatowich, William Towey, and Angus Duncan. Building science and accountability into community-based restoration. Fisheries, 31(1):17–24, 2006. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282006%2931%5B17%3ABSAAIC%5D2.0.CO%3B2.

Anonymous:2006:FIAb

[760] Anonymous. Full issue article volume 31, issue 2. Fisheries, 31(2):53–104, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-31-2.

Fitzgerald:2006:FFH

[761] D. G. Fitzgerald, B. Zhu, S. B. Hoskins, D. E. Haddad, K. N. Green, L. G. Rudstam, and E. L. Mills. Feature: Fish habitat — quantifying submerged aquatic vegetation using aerial photograph interpretation: Application in studies assessing fish habitat in freshwater ecosystems. Fisheries, 31(2):61–73, 2006. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282006%2931%5B61%3AFHQSAV%5D2.0.CO%3B2.

Quist:2006:FFH

[762] Michael C. Quist, Wayne A. Hubert, Mark Fowden, Steven W. Wolff, and Michael R. Bower. Feature — fish habitat — the Wyoming Habitat Assessment Methodology (WHAM): a systematic approach to evaluating watershed conditions and stream habitat. Fisheries, 31(2):75– 81, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282006%2931%5B75%3AFFHTWH%5D2.0.CO%3B2.

Anonymous:2006:FIAc

[763] Anonymous. Full issue article volume 31, issue 3. Fisheries, 31(3):105– 156, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-31-3.

Mitchell:2006:PSR

[764] Andrew J. Mitchell and Anita M. Kelly. The public sector role in the establishment of grass carp in the United States. Fisheries, 31(3):113– 121, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282006%2931%5B113%3ATPSRIT%5D2.0.CO%3B2.

Schnick:2006:ZWA

[765] Rosalie “Roz” Schnick. Zero withdrawal anesthetic for all finfish and shellfish. Fisheries, 31(3):122–126, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%282006%2931%5B122%3AZWAFAF%5D2.0.
CO%3B2.

Anonymous:2006:FIAd

[766] Anonymous. Full issue article volume 31, issue 4. Fisheries, 31(4):157– 208, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-31-4.

Page:2006:CSF

[767] Kevin S. Page and Paul Radomski. Compliance with sport fishery regulations in Minnesota as related to regulation awareness. Fisheries, 31(4): 166–178, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282006%2931%5B166%3ACWSFRI%5D2.0.CO%3B2.

Wagner:2006:AMD

[768] Tyler Wagner, Daniel B. Hayes, and Mary T. Bremigan. Accounting for multilevel data structures in fisheries data using mixed models. Fisheries, 31(4):180–187, 2006. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282006%2931%5B180%3AAFMDSI%5D2.0.CO%3B2.

Hyun:2006:MC

[769] Karen Hyun. Matters of consequence. Fisheries, 31(4):188, 2006. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282006%2931%
5B188%3AMOC%5D2.0.CO%3B2.

Anonymous:2006:FIAe

[770] Anonymous. Full issue article volume 31, issue 5. Fisheries, 31(5):209– 260, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-31-5.

Dadswell:2006:RSA

[771] Michael J. Dadswell. A review of the status of Atlantic sturgeon in Canada, with comparisons to populations in the United States and Europe. Fisheries, 31(5):218–229, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%282006%2931%5B218%3AAROTSO%5D2.0.
CO%3B2.

Latta:2006:EHF

[772] W. C. Latta. The early history of fisheries management in Michigan. Fisheries, 31(5):230–234, 2006. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282006%2931%5B230%3ATEHOFM%5D2.0.CO%3B2.

Anonymous:2006:FIAf

[773] Anonymous. Full issue article volume 31, issue 6. Fisheries, 31(6):261– 312, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-31-6.

Lyons:2006:PBC

[774] John Lyons, Paul Hanson, and Elizabeth White. A photo-based computer system for identifying Wisconsin fishes. Fisheries, 31(6):269– 275, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282006%2931%5B269%3AAPCSFI%5D2.0.CO%3B2.

Anonymous:2006:FIAg

[775] Anonymous. Full issue article volume 31, issue 7. Fisheries, 31(7):313– 364, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-31-7.

Sass:2006:FCF

[776] Greg G. Sass, James F. Kitchell, Stephen R. Carpenter, Thomas R. Hrabik, Anna E. Marburg, and Monica G. Turner. Fish community and food web responses to a whole-lake removal of coarse woody habitat. Fisheries, 31(7):321–330, 2006. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282006%2931%5B321%3AFCAFWR%5D2.0.CO%3B2.

Dame:2006:UUN

[777] James K. Dame and Robert R. Christian. Uncertainty and the use of network analysis for ecosystem-based fishery management. Fisheries, 31(7):

331–341, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282006%2931%5B331%3AUATUON%5D2.0.CO%3B2.

Anonymous:2006:FIAh

[778] Anonymous. Full issue article volume 31, issue 8. Fisheries, 31(8):365– 424, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-31-8.

Kaczynski:2006:ASR

[779] V. W. Kaczynski and Fabian Alvarado. Assessment of the southern range limit of North American Coho salmon. Fisheries, 31(8):374– 391, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282006%2931%5B374%3AAOTSRL%5D2.0.CO%3B2.

Kozfkay:2006:RUS

[780] Joseph R. Kozfkay, Jeffrey C. Dillon, and Daniel J. Schill. Routine use of sterile fish in salmonid sport fisheries. Fisheries, 31(8):392– 401, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282006%2931%5B392%3ARUOSFI%5D2.0.CO%3B2.

Anonymous:2006:FIAi

[781] Anonymous. Full issue article volume 31, issue 9. Fisheries, 31(9):425– 476, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-31-9.

Rountree:2006:LF

[782] Rodney A. Rountree, R. Grant Gilmore, Clifford A. Goudey, Anthony D. Hawkins, Joseph J. Luczkovich, and David A. Mann. Listening to fish. Fisheries, 31(9):433–446, 2006. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282006%2931%5B433%3ALTF%5D2.0.CO%3B2.

Conway:2006:EHW

[783] Flaxen D. L. Conway and Carrie Pomeroy. Evaluating the human-as well as the biological-objectives of cooperative fisheries research. Fisheries, 31(9):447–454, 2006. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282006%2931%5B447%3AETHWAT%5D2.0.CO%3B2.

Anonymous:2006:FIAj

[784] Anonymous. Full issue article volume 31, issue 10. Fisheries, 31(10): 477–528, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-31-10.

Smith:2006:SE

[785] David R. Smith and Stewart F. Michels. Seeing the elephant. Fisheries, 31(10):485–491, 2006. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282006%2931%5B485%3ASTE%5D2.0.CO%3B2.

Zydlewski:2006:RMF

[786] Gayle Barbin Zydlewski, Gregg Horton, Todd Dubreuil, Benjamin Letcher, Sean Casey, and Joseph Zydlewski. Remote monitoring of fish in small streams. Fisheries, 31(10):492–502, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282006%2931%
5B492%3ARMOFIS%5D2.0.CO%3B2.

Connelly:2006:AMW

[787] Nancy A. Connelly, Tommy L. Brown, and Jill M. Hardiman. AFS men and women differ most in their lifestyle choices. Fisheries, 31(10): 503–506, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282006%2931%5B503%3AAMAWDM%5D2.0.CO%3B2.

Anonymous:2006:FIAk

[788] Anonymous. Full issue article volume 31, issue 11. Fisheries, 31(11): 529–580, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-31-11.

Saunders:2006:MDF

[789] Rory Saunders, Michael A. Hachey, and Clem W. Fay. Maine’s diadromous fish community. Fisheries, 31(11):537–547, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282006%2931%
5B537%3AMDFC%5D2.0.CO%3B2.

Anonymous:2006:FIAl

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-31-12.

Brown:2006:CIE

[791] Stephen K. Brown, Manoj Shivlani, David Die, David B. Sampson, and Tina A. Ting. The center for independent experts. Fisheries, 31(12): 590–600, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282006%2931%5B590%3ATCFIE%5D2.0.CO%3B2.

Acosta:2006:ITS

[792] Charles A. Acosta. Impending trade suspensions of Caribbean queen conch under CITES. Fisheries, 31(12):601–606, 2006. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282006%2931%
5B601%3AITSOCQ%5D2.0.CO%3B2.

Anonymous:2007:FIAa

[793] Anonymous. Full issue article volume 32, issue 1. Fisheries, 32(1):1–52, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-32-1.

Angermeier:2007:RFB

[794] Paul L. Angermeier. The role of fish biologists in helping society build ecological sustainability. Fisheries, 32(1):9–20, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282007%2932%
5B9%3ATROFBI%5D2.0.CO%3B2.

Tuckey:2007:ELV

[795] Troy Tuckey, Noëlle Yochum, John Hoenig, Jon Lucy, and Joseph Cimino. Evaluating localized vs. large-scale management: The example of Tautog in Virginia. Fisheries, 32(1):21–28, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282007%2932%
5B21%3AELVLME%5D2.0.CO%3B2.

Anonymous:2007:FIAb

[796] Anonymous. Full issue article volume 32, issue 2. Fisheries, 32(2):53–104, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-32-2.

Paukert:2007:HTC

[797] Craig Paukert, Michael McInerny, and Randall Schultz. Historical trends in creel limits, length-based limits, and season restrictions for black basses in the United States and Canada. Fisheries, 32(2):62– 72, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B62%3AHTICLB%5D2.0.CO%3B2.

Sutton:2007:CRF

[798] Stephen G. Sutton. Constraints on recreational fishing participation in Queensland, Australia. Fisheries, 32(2):73–83, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446%282007%2932%5B73%
3ACORFPI%5D2.0.CO%3B2.

Anonymous:2007:FIAc

[799] Anonymous. Full issue article volume 32, issue 3. Fisheries, 32(3):105– 156, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-32-3.

Pearsons:2007:DCA

[800] Todd N. Pearsons, Dennis D. Roley, and Christopher L. Johnson. Development of a carcass analog for nutrient restoration in streams. Fisheries, 32(3):114–124, 2007. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282007%2932%5B114%3ADOACAF%5D2.0.CO%3B2.

Anders:2007:CWP

[801] Paul J. Anders and Ken I. Ashley. The clear-water paradox of aquatic ecosystem restoration. Fisheries, 32(3):125–128, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282007%2932%
5B125%3ATWPOAE%5D2.0.CO%3B2.

Bailey:2007:PAS

[802] Paul E. Bailey. Proportional angling success: an alternative approach to representing angling success. Fisheries, 32(3):129–135, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B129%3APASAAT%5D2.0.CO%3B2.

Anonymous:2007:FIAd

[803] Anonymous. Full issue article volume 32, issue 4. Fisheries, 32(4):157– 208, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-32-4.

Wei:2007:UII

[804] Anhua Wei and Patricia Chow-Fraser. Use of IKONOS imagery to map coastal wetlands of Georgian Bay. Fisheries, 32(4):167–173, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B167%3AUOIITM%5D2.0.CO%3B2.

Cada:2007:PIH

[805] Glenn Cada, James Ahlgrimm, Michael Bahleda, Tom Bigford, Stefanie Damiani Stavrakas, Douglas Hall, Russell Moursund, and Michael Sale. Potential impacts of hydrokinetic and wave energy conversion technologies on aquatic environments. Fisheries, 32(4):174–181, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B174%3APIOHAW%5D2.0.CO%3B2.

Conway:2007:CIO

[806] Flaxen Conway and Laura Opsommer. Communicating and interacting with Oregon’s coastal marine recreational fishing community. Fisheries, 32(4):182–188, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%282007%2932%5B182%3ACAIWSC%5D2.0.CO%3B2.

Anonymous:2007:FIAe

[807] Anonymous. Full issue article volume 32, issue 5. Fisheries, 32(5):209– 260, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-32-5.

Francis:2007:TCE

[808] Robert C. Francis, Mark A. Hixon, M. Elizabeth Clarke, Steven A. Murawski, and Stephen Ralston. Ten commandments for ecosystembased fisheries scientists. Fisheries, 32(5):217–233, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282007%2932%
5B217%3ATCFBFS%5D2.0.CO%3B2.

Schramm:2007:IBP

[809] Harold L. Schramm, Jr. and Kevin M. Hunt. Issues, benefits, and problems associated with fishing tournaments in inland waters of the United States: a survey of fishery agency administrators. Fisheries, 32(5):234– 243, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B234%3AIBAPAW%5D2.0.CO%3B2.

Anonymous:2007:FIAf

[810] Anonymous. Full issue article volume 32, issue 6. Fisheries, 32(6):261– 312, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-32-6.

Sommer:2007:CPF

[811] Ted Sommer, Chuck Armor, Randall Baxter, Richard Breuer, Larry Brown, Mike Chotkowski, Steve Culberson, Fredrick Feyrer, Marty Gingras, Bruce Herbold, Wim Kimmerer, Anke Mueller-Solger, Matt Nobriga, and Kelly Souza. The collapse of pelagic fishes in the upper San Francisco Estuary: El colapso de los peces pelagicos en la cabecera del Estuario San Francisco. Fisheries, 32(6):270–277, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B270%3ATCOPFI%5D2.0.CO%3B2.

Cooperman:2007:SRE

[812] M. S. Cooperman, S. G. Hinch, S. Bennett, M. A. Branton, R. V. Galbraith, J. T. Quigley, and B. A. Heise. Streambank restoration effectiveness: Lessons learned from a comparative study. Fisheries, 32(6): 278–291, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B278%3ASRELFA%5D2.0.CO%3B2.

Harvey:2007:EMF

[813] Bret C. Harvey and Steven F. Railsback. Estimating multi-factor cumulative watershed effects on fish populations with an individual-based model. Fisheries, 32(6):292–298, 2007. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446%282007%2932%5B292%3AEFCWEO%5D2.0.CO%3B2.

Anonymous:2007:FIAg

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-32-7.

Jackson:2007:ERA

[815] James R. Jackson. Earliest references to age determination of fishes and their early application to the study of fisheries. Fisheries, 32(7): 321–328, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B321%3AERTADO%5D2.0.CO%3B2.

Maceina:2007:CSR

[816] Michael J. Maceina, Jeff Boxrucker, David L. Buckmeier, R. Scott Gangl, David O. Lucchesi, Daniel A. Isermann, James R. Jackson, and Patrick J. Martinez. Current status and review of freshwater fish aging procedures used by state and provincial fisheries agencies with recommendations for future directions. Fisheries, 32(7):329– 340, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B329%3ACSAROF%5D2.0.CO%3B2.

Anonymous:2007:FIAh

[817] Anonymous. Full issue article volume 32, issue 8. Fisheries, 32(8):365– 416, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-32-8.

Taylor:2007:RCS

[818] Christopher A. Taylor, Guenter A. Schuster, John E. Cooper, Robert J. Distefano, Arnold G. Eversole, Premek Hamr, Horton H. Hobbs III, Henry W. Robison, Christopher E. Skelton, and Roger F. Thoma. A reassessment of the conservation status of crayfishes of the United States and Canada after 10+ years of increased awareness. Fisheries, 32(8): 372–389, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B372%3AAROTCS%5D2.0.CO%3B2.

Bettoli:2007:PPO

[819] Phillip W. Bettoli, George D. Scholten, and Willliam C. Reeves. Protecting paddlefish from overfishing: a case history of the research and regulatory process. Fisheries, 32(8):390–397, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1577/1548-8446%282007%2932%
5B390%3APPFOCH%5D2.0.CO%3B2.

Anonymous:2007:FIAi

[820] Anonymous. Full issue article volume 32, issue 9. Fisheries, 32(9):417– 468, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-32-9.

Good:2007:RPE

[821] Thomas P. Good, Timothy J. Beechie, Paul McElhany, Michelle M. Mc-Clure, and Mary H. Ruckelshaus. Recovery planning for Endangered Species Act-listed Pacific salmon: Using science to inform goals and strategies. Fisheries, 32(9):426–440, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446%282007%2932%5B426%3ARPFESL%5D2.0.
CO%3B2.

Adams:2007:CSN

[822] Peter B. Adams, Louis W. Botsford, Kenneth W. Gobalet, Robert A. Leidy, Dennis R. McEwan, Peter B. Moyle, Jerry J. Smith, John G. Williams, and Ronald M. Yoshiyama. Coho salmon are native south of San Francisco Bay: a reexamination of North American Coho salmon’s southern range limit. Fisheries, 32(9):441–451, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B441%3ACSANSO%5D2.0.CO%3B2.

Anonymous:2007:FIAj

[823] Anonymous. Full issue article volume 32, issue 10. Fisheries, 32(10): 469–520, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-32-10.

Williams:2007:CSI

[824] Jack E. Williams, Amy L. Haak, Nathaniel G. Gillespie, and Warren T. Colyer. The conservation success index: Synthesizing and communicating salmonid condition and management needs. Fisheries, 32(10): 477–493, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B477%3ATCSIAC%5D2.0.CO%3B2.

Anonymous:2007:FIAk

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-32-11.

Hill:2007:NAS

[826] Jeffrey E. Hill and Paul Zajicek. National aquatic species risk analysis: a call for improved implementation. Fisheries, 32(11):530– 538, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B530%3ANASRAC%5D2.0.CO%3B2.

Seeb:2007:DSD

[827] L. W. Seeb, A. Antonovich, M. A. Banks, T. D. Beacham, M. R. Bellinger, S. M. Blankenship, M. R. Campbell, N. A. Decovich, J. C. Garza, C. M. Guthrie III, T. A. Lundrigan, P. Moran, S. R. Narum, J. J. Stephenson, K. J. Supernault, D. J. Teel, W. D. Templin, J. K. Wenburg, S. F. Young, and C. T. Smith. Development of a standardized DNA database for Chinook salmon. Fisheries, 32(11):540– 552, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B540%3ADOASDD%5D2.0.CO%3B2.

Anonymous:2007:FIAl

[828] Anonymous. Full issue article volume 32, issue 12. Fisheries, 32(12): 573–624, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-32-12.

McDonald:2007:RME

[829] Lyman L. McDonald, Robert Bilby, Peter A. Bisson, Charles C. Coutant, John M. Epifanio, Daniel Goodman, Susan Hanna, Nancy Hundy, Erik Merrill, Brian Riddell, William Liss, Eric J. Loudenslager, David P. Philipp, William Smoker, Richard R. Whitney, and Richard N. Williams. Research, monitoring, and evaluation of fish and wildlife restoration projects in the Columbia River Basin: Lessons learned and suggestions for large-scale monitoring programs. Fisheries, 32(12):582–590, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446%282007%2932%5B582%3ARMAEOF%5D2.0.CO%3B2.

Macinko:2007:SSF

[830] Seth Macinko and Sarah Schumann. Searching for subsistence: In the field in pursuit of an elusive concept in small-scale fisheries. Fisheries, 32(12):592–600, 2007. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446%282007%2932%5B592%3ASFSTFI%5D2.0.CO%3B2.

Anonymous:2008:FIAa

[831] Anonymous. Full issue article volume 33, issue 1. Fisheries, 33(1):1–52, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-33-1.

Rogers:2008:HIL

[832] Mark W. Rogers and Michael S. Allen. Hurricane impacts to Lake Okeechobee: Altered hydrology creates difficult management trade offs. Fisheries, 33(1):11–17, 2008. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-33.1.11.

Al-Chokhachy:2008:SPO

[833] Robert Al-Chokhachy, Wade Fredenberg, and Shelley Spalding. Surveying professional opinion to inform bull trout recovery and management decisions. Fisheries, 33(1):18–28, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446-33.1.18.

Anonymous:2008:FIAb

[834] Anonymous. Full issue article volume 33, issue 2. Fisheries, 33(2):53–104, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-33-2.

Kim:2008:PPM

[835] Chang Gil Kim and Ho Sang Kim. Post-placement management of artificial reefs in Korea. Fisheries, 33(2):61–68, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-33.2.61.

Halverson:2008:STQ

[836] M. Anders Halverson. Stocking trends: a quantitative review of governmental fish stocking in the United States, 1931 to 2004. Fisheries, 33 (2):69–75, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33.2.69.

Anonymous:2008:FIAc

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33-3.

Thresher:2008:ATC

[838] Ronald E. Thresher. Autocidal technology for the control of invasive fish. Fisheries, 33(3):114–121, 2008. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-33.3.114.

Whittier:2008:IF

[839] Thomas R. Whittier and J. Kevin Aitkin. Introduced fish. Fisheries, 33 (3):122–128, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446-33.3.122.

Anonymous:2008:FIAd

[840] Anonymous. Full issue article volume 33, issue 4. Fisheries, 33(4):157– 208, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33-4.

Morgan:2008:OFM

[841] Mark Morgan. Outlaw fishing in Missouri. Fisheries, 33(4):165– 171, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33.4.165.

Gunderson:2008:CMN

[842] Donald R. Gunderson, Ana M. Parma, Ray Hilborn, Jason M. Cope, David L. Fluharty, Marc L. Miller, Russell D. Vetter, Selina S. Heppell, and H. Gary Greene. The challenge of managing nearshore rocky reef resources. Fisheries, 33(4):172–179, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446-33.4.172.

Briggs:2008:NAO

[843] John C. Briggs. The North Atlantic Ocean: Need for proactive management. Fisheries, 33(4):180–185, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446-33.4.180. See reply [876].

Anonymous:2008:FIAe

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33-5.

Railsback:2008:DFA

[845] Steven F. Railsback and John Kadvany. Demonstration flow assessment: Judgment and visual observation in instream flow studies. Fisheries, 33 (5):217–227, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446-33.5.217.

Goddard:2008:MCA

[846] Chris I. Goddard, Nancy J. Leonard, Doug L. Stang, P. Jack Wingate, Barnett A. Rattner, J. Christian Franson, and Steven R. Sheffield. Management concerns about known and potential impacts of lead use in shooting and in fishing activities. Fisheries, 33(5):228–236, 2008. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1577/1548-8446-33.5.228.

Anonymous:2008:FIAf

[847] Anonymous. Full issue article volume 33, issue 6. Fisheries, 33(6):261– 312, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33-6.

Conway:2008:SLL

[848] Flaxen Conway and Wesley Shaw. Socioeconomic lessons learned from the response to the federally-declared West Coast groundfish disaster. Fisheries, 33(6):269–277, 2008. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-33.6.269.

Pearsons:2008:MRU

[849] Todd N. Pearsons. Misconception, reality, and uncertainty about ecological interactions and risks between hatchery and wild salmonids. Fisheries, 33(6):278–290, 2008. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-33.6.278.

Anonymous:2008:FIAg

[850] Anonymous. Full issue article volume 33, issue 7. Fisheries, 33(7):313– 364, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33-7.

Cooke:2008:DMU

[851] Steven J. Cooke, S. G. Hinch, A. P. Farrell, D. A. Patterson, K. Miller-Saunders, D. W. Welch, M. R. Donaldson, K. C. Hanson, G. T. Crossin, M. T. Mathes, A. G. Lotto, K. A. Hruska, I. C. Olsson, G. N. Wagner, R. Thomson, R. Hourston, K. K. English, S. Larsson, J. M. Shrimpton, and G. Van der Kraak. Developing a mechanistic understanding of fish migrations by linking telemetry with physiology, behavior, genomics and experimental biology: an interdisciplinary case study on adult Fraser River sockeye salmon. Fisheries, 33(7):321–339, 2008. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1577/1548-8446-33.7.321.

Hansen:2008:VIF

[852] Gretchen J. A. Hansen and Michael L. Jones. The value of information in fishery management. Fisheries, 33(7):340–348, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-33.7.340.

Anonymous:2008:FIAh

[853] Anonymous. Full issue article volume 33, issue 8. Fisheries, 33(8):365– 424, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33-8.

Jelks:2008:CSI

[854] Howard L. Jelks, Stephen J. Walsh, Noel M. Burkhead, Salvador Contreras-Balderas, Edmundo Diaz-Pardo, Dean A. Hendrickson, John Lyons, Nicholas E. Mandrak, Frank McCormick, Joseph S. Nelson, Steven P. Platania, Brady A. Porter, Claude B. Renaud, Juan Jacobo Schmitter-Soto, Eric B. Taylor, and Melvin L. Warren, Jr. Conservation status of imperiled North American freshwater and diadromous fishes. Fisheries, 33(8):372–407, 2008. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-33.8.372.

Anonymous:2008:FIAi

[855] Anonymous. Full issue article volume 33, issue 9. Fisheries, 33(9):425– 476, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33-9.

Todd:2008:DNW

[856] Andrew S. Todd, Mark A. Coleman, Aimee M. Konowal, Melynda K. May, Sarah Johnson, Nicole K. M. Vieira, and James F. Saunders. Development of new water temperature criteria to protect Colorado’s fisheries. Fisheries, 33(9):433–443, 2008. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-33.9.433.

Mather:2008:MCL

[857] Martha E. Mather, Donna L. Parrish, and John M. Dettmers. Mapping the changing landscape of fish-related journals: Setting a course for successful communication of scientific information. Fisheries, 33(9): 444–453, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33.9.444.

Anonymous:2008:FIAj

[858] Anonymous. Full issue article volume 33, issue 11. Fisheries, 33(11): 529–580, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33-11.

Dunham:2008:EEC

[859] Jason Dunham, Colden Baxter, Kurt Fausch, Wade Fredenberg, Satoshi Kitano, Itsuro Koizumi, Kentaro Morita, Tomoyuki Nakamura, Bruce Rieman, Ksenia Savvaitova, Jack Stanford, Eric Taylor, and Shoichiro Yamamoto. Evolution, ecology, and conservation of Dolly Varden, white spotted char, and bull trout. Fisheries, 33(11):537–550, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-33.11.537.

Hartley:2008:CRP

[860] Troy W. Hartley and Robert A. Robertson. Cooperative research program goals in New England: Perceptions of active commercial fishermen. Fisheries, 33(11):551–559, 2008. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-33.11.551.

Anonymous:2008:FIAk

[861] Anonymous. Full issue article volume 33, issue 12. Fisheries, 33(12): 581–632, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33-12.

Kaeser:2008:ADL

[862] Adam J. Kaeser and Thomas L. Litts. An assessment of deadhead logs and large woody debris using side scan sonar and field surveys in streams of southwest Georgia. Fisheries, 33(12):589–597, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-33.12.589.

Grober-Dunsmore:2008:VZM

[863] Rikki Grober-Dunsmore, Lisa Wooninck, John Field, Cameron Ainsworth, Jim Beets, Steve Berkeley, Jim Bohnsack, Rafe Boulon, Richard Brodeur, John Brodziak, Larry Crowder, Danny Gleason, Mark Hixon, Les Kaufman, Bill Lindberg, Marc Miller, Lance Morgan, and Charles Wahle. Vertical zoning in marine protected areas: Ecological considerations for balancing pelagic fishing with conservation of benthic communities. Fisheries, 33(12):598–610, 2008. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-33.12.598.

Wiley:2008:RTG

[864] Robert W. Wiley. The 1962 rotenone treatment of the Green River, Wyoming and Utah, revisited: Lessons learned. Fisheries, 33(12):611– 617, 2008. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-33.12.611.

Anonymous:2009:FIAa

[865] Anonymous. Full issue article volume 34, issue 1. Fisheries, 34(1):1–52, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-34-1.

Lauber:2009:FIM

[866] T. Bruce Lauber, Eric J. Taylor, and Barbara A. Knuth. Factors influencing membership of U.S. Fish and Wildlife Service and U.S. Geological Survey biologists in the American Fisheries Society. Fisheries, 34 (1):9–19, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34.1.9.

Wilde:2009:DVP

[867] Gene R. Wilde. Does venting promote survival of released fish? Fisheries, 34(1):20–28, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446-34.1.20.

Anonymous:2009:FIAb

[868] Anonymous. Full issue article volume 34, issue 2. Fisheries, 34(2):53–104, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-34-2.

Shaff:2009:DIN

[869] Courtney D. Shaff and Jana E. Compton. Differential incorporation of natural spawners vs. artificially planted salmon carcasses in a stream food web: Evidence from δ ^{15}N of juvenile Coho salmon. Fisheries, 34 (2):62–72, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34.2.62.

Barrett:2009:EDH

[870] Paul J. Barrett. Estimating devils hole pupfish lifestage ratios using the Delphi method. Fisheries, 34(2):73–79, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446-34.2.73.

Anonymous:2009:FIAc

[871] Anonymous. Full issue article volume 34, issue 3. Fisheries, 34(3):105– 156, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34-3.

Kerr:2009:PMF

[872] Lisa A. Kerr, David H. Secor, and Philip M. Piccoli. Partial migration of fishes as exemplified by the estuarine-dependent white perch. Fisheries, 34(3):114–123, 2009. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-34.3.114.

Piccolo:2009:LAS

[873] John J. Piccolo, Milo D. Adkison, and Frank Rue. Linking Alaskan salmon fisheries management with ecosystem-based escapement goals: a review and prospectus. Fisheries, 34(3):124–134, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446-34.3.124.

Anonymous:2009:FIAd

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34-4.

Pine:2009:CRF

[875] William E. Pine III, Steven J. D. Martell, Carl J. Walters, and James F. Kitchell. Counterintuitive responses of fish populations to management actions. Fisheries, 34(4):165–180, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446-34.4.165.

Courtenay:2009:RIM

[876] Walter R. Courtenay, Jr., Bruce B. Collette, Timothy E. Essington, Ray Hilborn, James W. Orr, Daniel Pauly, John E. Randall, and William F. Smith-Vaniz. Risks of introductions of marine fishes: Reply to Briggs. Fisheries, 34(4):181–186, 2009. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-34.4.181. See [843].

Anonymous:2009:FIAe

[877] Anonymous. Full issue article volume 34, issue 5. Fisheries, 34(5):209– 260, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34-5.

Berkson:2009:TSF

[878] J. Berkson, K. M. Hunt, J. C. Whitehead, D. J. Murie, T. J. Kwak, and J. Boreman. Is there a shortage of fisheries stock assessment scientists? Fisheries, 34(5):217–219, 2009. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-34.5.217.

Berkson:2009:ASS

[879] J. Berkson, S. F. Hudy, S. L. McMullin, B. R. Murphy, M. H. Prager, M. J. Kelly, and N. B. Thompson. Addressing the shortage of stock assessment scientists through undergraduate workshops. Fisheries, 34(5): 220–227, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34.5.220.

Jacob:2009:CCC

[880] Steve Jacob and Michael Jepson. Creating a community context for the Fishery Stock Sustainability Index. Fisheries, 34(5):228–231, 2009. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1577/1548-8446-34.5.228.

Fayram:2009:ECB

[881] Andrew H. Fayram, Dennis A. Schenborn, Joseph M. Hennessy, Nancy A. Nate, and Patrick J. Schmalz. Exploring the conflict between broad scale and local inland fisheries management: The risks to agency credibility. Fisheries, 34(5):232–236, 2009. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-34.5.232.

Anonymous:2009:FIAf

[882] Anonymous. Full issue article volume 34, issue 6. Fisheries, 34(6):261– 312, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34-6.

Vries:2009:EPR

[883] Dennis R. De Vries, Elizabeth A. Marschall, and Roy A. Stein. Exploring the peer review process: What is it, does it work, and can it be improved? Fisheries, 34(6):270–279, 2009. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-34.6.270.

Bromley:2009:ARD

[884] Daniel W. Bromley. Abdicating responsibility: The deceits of fisheries policy. Fisheries, 34(6):280–290, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446-34.6.280.

Anonymous:2009:FIAg

[885] Anonymous. Full issue article volume 34, issue 8. Fisheries, 34(8):365– 416, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34-8.

Peters:2009:ISP

[886] Jody A. Peters and David M. Lodge. Invasive species policy at the regional level: a multiple weak links problem. Fisheries, 34(8):373– 380, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34.8.373.

Bothwell:2009:BFH

[887] Max L. Bothwell, Donovan R. Lynch, Harlan Wright, and John Deniseger. On the boots of fishermen: The history of didymo blooms on Vancouver Island, British Columbia. Fisheries, 34(8):382–388, 2009. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1577/1548-8446-34.8.382.

Johnson:2009:WDA

[888] Brett M. Johnson, Robert Arlinghaus, and Patrick J. Martinez. Are we doing all we can to stem the tide of illegal fish stocking? Fisheries, 34(8): 389–394, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34.8.389.

Anonymous:2009:FIAh

[889] Anonymous. Full issue article volume 34, issue 9. Fisheries, 34(9):417– 468, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34-9.

Martinez:2009:WLT

[890] Patrick J. Martinez, Patricia E. Bigelow, Mark A. Deleray, Wade A. Fredenberg, Barry S. Hansen, Ned J. Horner, Stafford K. Lehr, Roger W. Schneidervin, Scott A. Tolentino, and Art E. Viola. Western lake trout woes. Fisheries, 34(9):424–442, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446-34.9.424.

Sutton:2009:WDP

[891] Stephen G. Sutton, Kara Dew, and Jim Higgs. Why do people drop out of recreational fishing? A study of lapsed fishers from Queensland, Australia. Fisheries, 34(9):443–452, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446-34.9.443.

Anonymous:2009:FIAi

[892] Anonymous. Full issue article volume 34, issue 12. Fisheries, 34(12): 573–624, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34-12.

Monroe:2009:FPE

[893] Jeremy B. Monroe, Colden V. Baxter, Julian D. Olden, and Paul L. Angermeier. Freshwaters in the public eye: Understanding the role of images and media in aquatic conservation. Fisheries, 34(12):581– 585, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-34.12.581.

DiStefano:2009:BIP

[894] Robert J. DiStefano, Mary E. Litvan, and Paul T. Horner. The bait industry as a potential vector for alien crayfish introductions: Problem recognition by fisheries agencies and a Missouri evaluation. Fisheries, 34(12):586–597, 2009. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-34.12.586.

Hewitt:2009:MCA

[895] David A. Hewitt, Jason S. Link, David H. Wahl, Steven J. Cooke, and Martha E. Mather. Maintaining the competitiveness of the American Fisheries Society journals: an assessment based on influence and costeffectiveness. Fisheries, 34(12):598–606, 2009. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446-34.12.598.

Anonymous:2010:FIAa

[896] Anonymous. Full issue article volume 35, issue 1. Fisheries, 35(1):1–52, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-35.1.

Jackson:2010:MSM

[897] Donald C. Jackson. Mentoring: a synergistic multi-directional endeavor. Fisheries, 35(1):4–47, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.1.4.

Anonymous:2010:NF

[898] Anonymous. News: Fisheries. Fisheries, 35(1):5, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/1548-8446-35.1.5.

Anonymous:2010:JHT

[899] Anonymous. Journal highlights: Transactions of the American Fisheries Society. Fisheries, 35(1):6, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.1.6.

Hawkes:2010:ULP

[900] Elden Hawkes, Jr. Update: Legislation and policy. Fisheries, 35(1):8, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1577/1548-8446-35.1.8.

McMichael:2010:JSA

[901] Geoffrey A. McMichael, M. Brad Eppard, Thomas J. Carlson, Jessica A. Carter, Blaine D. Ebberts, Richard S. Brown, Mark Weiland, Gene R. Ploskey, Ryan A. Harnish, and Z. Daniel Deng. The juvenile salmon acoustic telemetry system: a new tool. Fisheries, 35(1):9–22, 2010. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1577/1548-8446-35.1.9.

Pearsons:2010:OHW

[902] Todd N. Pearsons. Operating hatcheries within an ecosystem context using the adaptive stocking concept. Fisheries, 35(1):23–31, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.1.23.

Rassam:2010:LFV

[903] Gus Rassam. Looking forward with a 2020 vision. … Fisheries, 35 (1):32, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.1.32.

Anonymous:2010:GFG

[904] Anonymous. Guidelines: Fisheries 2010 guide for authors. Fisheries, 35 (1):33–39, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.1.33.

Anonymous:2010:CAC

[905] Anonymous. Calendar: 2010 AFS Chapter and Division Meetings. Fisheries, 35(1):40, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446-35.1.40.

Anonymous:2010:CFE

[906] Anonymous. Calendar: Fisheries events. Fisheries, 35(1):41–42, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.1.41.

Anonymous:2010:AFS

[907] Anonymous. American Fisheries Society 140th Annual Meeting: a diverse array of gatherings. Fisheries, 35(1):43, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.1.43.

Anonymous:2010:NAU

[908] Anonymous. News: AFS Units. Fisheries, 35(1):44–45, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.1.44.

Anonymous:2010:EFP

[909] Anonymous. Expert in fish passage and protection. Fisheries, 35(1): 46, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.1.46.

Gray:2010:RSF

[910] Steven Gray. Are robots and satellites the future of fisheries management? Fisheries, 35(1):48, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.1.48.

Anonymous:2010:AJC

[911] Anonymous. Announcements: Job center. Fisheries, 35(1):49–50, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.1.49.

Anonymous:2010:FIPa

[912] Anonymous. Full issue PDF. Fisheries, 35(2):53–104, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.2.

Vinson:2010:PIA

[913] Mark R. Vinson, Eric C. Dinger, and Deanna K. Vinson. Piscicides and invertebrates: After 70 years, does anyone really know? Fisheries, 35 (2):61–71, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.2.61.

Brown:2010:ASV

[914] J. Jed Brown and Gregory W. Murphy. Atlantic sturgeon vessel-strike mortalities in the Delaware Estuary. Fisheries, 35(2):72–83, 2010. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1577/1548-8446-35.2.72.

Anonymous:2010:FIPb

[915] Anonymous. Full issue PDF. Fisheries, 35(3):105–156, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.3.

Hilborn:2010:DOS

[916] Ray Hilborn and Kevin Stokes. Defining overfished stocks: Have we lost the plot? Fisheries, 35(3):113–120, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446-35.3.113.

Rothlisberger:2010:AIS

[917] John D. Rothlisberger, W. Lindsay Chadderton, Joanna McNulty, and David M. Lodge. Aquatic invasive species transport via trailered boats: What is being moved, who is moving it, and what can be done. Fisheries, 35(3):121–132, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.3.121.

Anonymous:2010:FIPc

[918] Anonymous. Full issue PDF. Fisheries, 35(4):157–208, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.4.

Kaeser:2010:NTM

[919] Adam J. Kaeser and Thomas L. Litts. A novel technique for mapping habitat in navigable streams using low-cost side scan sonar. Fisheries, 35(4):163–174, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
1548-8446-35.4.163.

Miranda:2010:FHD

[920] L. E. Miranda, M. Spickard, T. Dunn, K. M. Webb, J. N. Aycock, and K. Hunt. Fish habitat degradation in U.S. reservoirs. Fisheries, 35(4): 175–184, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.4.175.

Anonymous:2010:FIPd

[921] Anonymous. Full issue PDF. Fisheries, 35(5):209–260, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.5.

Hewitt:2010:IIF

[922] David A. Hewitt, Eric C. Janney, Brian S. Hayes, and Rip S. Shively. Improving inferences from fisheries capture-recapture studies through remote detection of PIT tags. Fisheries, 35(5):217–231, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.5.217.

Barnas:2010:CTH

[923] Katie Barnas and Stephen L. Katz. The challenges of tracking habitat restoration at various spatial scales. Fisheries, 35(5):232–241, 2010. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1577/1548-8446-35.5.232.

Anonymous:2010:FIPe

[924] Anonymous. Full issue PDF. Fisheries, 35(6):261–312, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.6.

Kershner:2010:EMO

[925] Jeffrey L. Kershner and Brett B. Roper. An evaluation of management objectives used to assess stream habitat conditions on Federal lands within the interior Columbia Basin. Fisheries, 35(6):269–278, 2010. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1577/1548-8446-35.6.269.

Scheerer:2010:SCM

[926] P. D. Scheerer and T. A. O’Neill. A species crediting methodology that supports conservation banking for an endangered floodplain minnow. Fisheries, 35(6):280–291, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.6.280.

Anonymous:2010:FIPf

[927] Anonymous. Full issue PDF. Fisheries, 35(7):313–364, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.7.1.

Woody:2010:MLC

[928] Carol Ann Woody, Robert M. Hughes, Eric J. Wagner, Thomas P. Quinn, Leanne H. Roulson, Lori M. Martin, and Kitty Griswold. The Mining Law of 1872: Change is overdue. Fisheries, 35(7):321–331, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.7.321.

Pender:2010:ERA

[929] Danielle R. Pender and Fred A. Harris. Environmental review approaches by fish and wildlife agencies in the United States and Canada. Fisheries, 35(7):332–339, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.7.332.

Collins:2010:PAA

[930] Mark R. Collins. Problems associated with assessing the status of populations of Atlantic and shortnose sturgeons in the Southeastern United States. Fisheries, 35(7):340–341, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1577/1548-8446-35.7.340.

Anonymous:2010:FIPg

[931] Anonymous. Full issue PDF. Fisheries, 35(8):365–416, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.8.1.

Wipfli:2010:LEF

[932] Mark S. Wipfli and Colden V. Baxter. Linking ecosystems, food webs, and fish production: Subsidies in salmonid watersheds. Fisheries, 35(8): 373–387, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.8.373.

Jennings:2010:IGC

[933] Martin J. Jennings, Brian L. Sloss, Gene R. Hatzenbeler, Jeffrey M. Kampa, Timothy D. Simonson, Steven P. Avelallemant, Gary A. Lindenberger, and Bruce D. Underwood. Implementation of genetic conservation practices in a muskellunge propagation and stocking program. Fisheries, 35(8):388–395, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.8.388.

Anonymous:2010:FIPh

[934] Anonymous. Full issue PDF. Fisheries, 35(9):417–468, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.9.1.

Miller:2010:FIR

[935] Thomas J. Miller, Jeff A. Blair, Thomas F. Ihde, Robert M. Jones, David H. Secor, and Michael J. Wilberg. FishSmart: an innovative role for science in stakeholder-centered approaches to fisheries management. Fisheries, 35(9):424–433, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.9.422.

Williams:2010:SEF

[936] John G. Williams. Sampling for environmental flow assessments. Fisheries, 35(9):434–443, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.9.434.

Gabelhouse:2010:NPF

[937] Donald William Gabelhouse, Jr. Needs and proficiencies of fisheries hires by State agencies. Fisheries, 35(9):445–448, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.9.445.

Anonymous:2010:F

[938] Anonymous. Fisheries. Fisheries, 35(10):1–52, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1577/1548-8446-35.10.1.

Al-Chokhachy:2010:DAH

[939] Robert Al-Chokhachy and Brett B. Roper. Different approaches to habitat surveys can impact fisheries management and conservation decisions. Fisheries, 35(10):476–488, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.10.476.

Radomski:2010:PID

[940] Paul Radomski, Lyn A. Bergquist, Michael Duval, and Andrew Williquett. Potential impacts of docks on littoral habitats in Minnesota lakes. Fisheries, 35(10):489–495, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.10.489.

Seitz:2010:DLT

[941] Andrew C. Seitz and Trent M. Sutton. Distance learning in today’s classroom. Fisheries, 35(10):501–505, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446-35.10.501.

Anonymous:2010:FIAb

[942] Anonymous. Full issue article volume 35, issue 11. Fisheries, 35(11): 521–572, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.11.1.

Director:2010:CPT

[943] Amy L. Haak Resource Information Director, Jack E. Williams Senior Scientist, Helen M. Neville Research Scientists, Daniel C. Dauwalter Research Scientists, and Warren T. Colyer Restoration Ecologist. Conserving peripheral trout populations: The values and risks of life on the edge. Fisheries, 35(11):530–549, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.11.530.

Pister:2010:CGT

[944] Edwin P. Pister. California golden trout: Perspectives on restoration and management. Fisheries, 35(11):550–553, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1577/1548-8446-35.11.550.

Anonymous:2010:FIAc

[945] Anonymous. Full issue article volume 35, issue 12. Fisheries, 35(12): 573–624, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.12.1.

Clemens:2010:SDU

[946] Benjamin J. Clemens, Thomas R. Binder, Margaret F. Docker, Mary L. Moser, and Stacia A. Sower. Similarities, differences, and unknowns in biology and management of three parasitic lampreys of North America. Fisheries, 35(12):580–594, 2010. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/1548-8446-35.12.580.

Miller:2010:TLF

[947] Robert J. Miller. Tagging by lobster fishermen to estimate abundance of the discarded portions of their catch. Fisheries, 35(12):595– 600, 2010. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1577/1548-
8446-35.12.595.

Anonymous:2011:FIPa

[948] Anonymous. Full issue PDF volume 36, issue 1. Fisheries, 36 (1):1–52, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
03632415.2011.580717.

Hubert:2011:PHN

[949] Wayne Hubert. President’s hook: New frontiers in fisheries management and ecology: Leadership in collaboration and coalition. Fisheries, 36(1):4–40, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389047.

Rassam:2011:DLP

[950] Gus Rassam. Director’s line: Plus ça change …. Fisheries, 36 (1):5, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389048.

Anonymous:2011:JAAa

[951] Anonymous. Journal of aquatic animal health. Fisheries, 36(1): 6, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389050.

Anonymous:2011:NAJa

[952] Anonymous. North American Journal of Aquaculture. Fisheries, 36 (1):6, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389049.

Hawkes:2011:LPa

[953] Elden Hawkes, Jr. Legislation and policy. Fisheries, 36(1):7, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1577/03632415.2011.10389051.

Brown:2011:PWI

[954] Richard S. Brown, Wayne A. Hubert, and Steven F. Daly. A primer on winter, ice, and fish: What fisheries biologists should know about winter ice processes and stream-dwelling fish. Fisheries, 36(1):8–26, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1577/03632415.2011.10389052.

Limburg:2011:HPI

[955] Karin E. Limburg, Robert M. Hughes, Donald C. Jackson, and Brian Czech. Human population increase, economic growth, and fish conservation: Collision course or savvy stewardship? Fisheries, 36(1): 27–35, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389053.

Cathcart:2011:SAG

[956] C. Nate Cathcart and Elliot M. Broder. Students’ angle: The Gulf oil spill: What it means to the Gulf and the future of fisheries biology students. Fisheries, 36(1):36–37, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1577/03632415.2011.10389054.

Amezcua:2011:FA

[957] Felipe Amezcua. Felipe Amezcua. Fisheries, 36(1):38–39, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1577/03632415.2011.10389055.

Blazer:2011:FCU

[958] Vicki S. Blazer. Fisheries currents: Understanding fish mortalities in the Potomac Drainage. Fisheries, 36(1):40, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1577/03632415.2011.10389056.

Anonymous:2011:FEa

[959] Anonymous. Fisheries events. Fisheries, 36(1):41, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1577/03632415.2011.10389057.

Anonymous:2011:FGA

[960] Anonymous. Fisheries 2011 guide for authors. Fisheries, 36(1): 42–43, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389058.

Anonymous:2011:AFSa

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389059.

Anonymous:2011:UAC

[962] Anonymous. Units: The 2010 Alaska Chapter Annual Meeting. Fisheries, 36(1):47, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389060.

Anonymous:2011:ACD

[963] Anonymous. 2011 AFS Chapter and Division Meetings: Members: Be sure to mark your calendars now, so you don’t miss these important local events! Fisheries, 36(1):48, 2011. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/abs/
10.1577/03632415.2011.10389061.

Cronin:2011:ASE

[964] Amanda Cronin. AFS 2011 Seattle: an extraordinary meeting in an extraordinary place. Fisheries, 36(1):49, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1577/03632415.2011.10389062.

Anonymous:2011:JCa

[965] Anonymous. Job center. Fisheries, 36(1):50, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1577/03632415.2011.10389063.

Anonymous:2011:EBPa

[966] Anonymous. Editorial Board page for Fisheries, volume 36, number 1. Fisheries, 36(1):??, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1577/03632415.2011.10389046.

Anonymous:2011:FIPb

[967] Anonymous. Full issue PDF volume 36, issue 2. Fisheries, 36(2): 55–102, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2011.557301.

Hubert:2011:PH

[968] Wayne Hubert. President’s hook. Fisheries, 36(2):56–94, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1577/03632415.2011.10389065.

Anonymous:2011:BMC

[969] Anonymous. Biology, management, and conservation of lampreys in North America. Fisheries, 36(2):57, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1577/03632415.2011.10389066.

Anonymous:2011:TAFa

[970] Anonymous. Transactions of the American Fisheries Society. Fisheries, 36(2):58, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389067.

Hawkes:2011:LPb

[971] Elden Hawkes, Jr. Legislation and policy. Fisheries, 36(2):59, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1577/03632415.2011.10389068.

Larson:2011:SCP

[972] Eric R. Larson and Julian D. Olden. The state of crayfish in the Pacific Northwest. Fisheries, 36(2):60–73, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1577/03632415.2011.10389069.

Strecker:2011:ATI

[973] Angela L. Strecker, Philip M. Campbell, and Julian D. Olden. The aquarium trade as an invasion pathway in the Pacific Northwest. Fisheries, 36(2):74–85, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389070.

Rassam:2011:DL

[974] Gus Rassam. Director’s line. Fisheries, 36(2):86, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1577/03632415.2011.10389071.

Anonymous:2011:FEb

[975] Anonymous. Fisheries events. Fisheries, 36(2):87, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1577/03632415.2011.10389072.

Anonymous:2011:AFSb

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389073.

Schiff:2011:TIC

[977] Danielle Joanne Schiff and Larry Barrett. Two Idaho Chapter AFS members die in helicopter crash. Fisheries, 36(2):92–93, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1577/03632415.2011.10389074.

Parasiewicz:2011:BR

[978] Piotr Parasiewicz. Book review. Fisheries, 36(2):95, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1577/03632415.2011.10389075.

Anonymous:2011:ASE

[979] Anonymous. AFS 2011 Seattle: an extraordinary meeting in an extraordinary place. Fisheries, 36(2):96–97, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1577/03632415.2011.10389076.

Anonymous:2011:UAM

[980] Anonymous. Units: AFS members in Hall of Fame. Fisheries, 36 (2):98–101, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1577/
03632415.2011.10389077.

Anonymous:2011:JCb

[981] Anonymous. Job center. Fisheries, 36(2):102, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1577/03632415.2011.10389078.

Anonymous:2011:EBPb

[982] Anonymous. Editorial Board page for Fisheries, volume 36, number 2. Fisheries, 36(2):??, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1577/03632415.2011.10389064.

Anonymous:2011:FIPc

[983] Anonymous. Full issue PDF volume 36, issue 3. Fisheries, 36(3): 107–154, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1577/
03632415.2011.571535.

President:2011:PHN

[984] Wayne A. Hubert Afs President. President’s hook: New frontiers in fisheries management and ecology: Leadership in recruitment of fisheries professionals. Fisheries, 36(3):109–147, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2011.10389080.

Fox:2011:F

[985] Sarah Fox. Fisheries. Fisheries, 36(3):110–111, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.10389081.

Hawkes:2011:LPc

[986] Elden W. Hawkes, Jr. and Policy Coordinator. Legislation and policy. Fisheries, 36(3):112, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2011.10389082.

Irwin:2011:ASD

[987] Brian J. Irwin, Michael J. Wilberg, Michael L. Jones, and James R. Bence. Applying structured decision making to recreational fisheries management. Fisheries, 36(3):113–122, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2011.10389083.

Donaldson:2011:CMR

[988] M. R. Donaldson, D. D. Aday, and S. J. Cooke. A call for mini-reviews: an effective but underutilized method of synthesizing knowledge to inform and direct fisheries management, policy, and research. Fisheries, 36(3):123–129, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.10389084.

Bowker:2011:GDL

[989] Jim Bowker and Jesse Trushenski. Guest Director’s line: AFS policy statement regarding the need for an immediate-release anesthetic/sedative for use in the fisheries disciplines. Fisheries, 36(3): 132–135, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.10389085.

Anonymous:2011:FEc

[990] Anonymous. Fisheries events. Fisheries, 36(3):136, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.10389086.

Anonymous:2011:MHM

[991] Anonymous. Margaret H. Murphy. Fisheries, 36(3):137–138, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.10389087.

Anonymous:2011:DP

[992] Anonymous. Donna L. Parrish. Fisheries, 36(3):139–140, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.10389088.

Loury:2011:FMC

[993] Erin Loury. Fishing with a mission: Collaborating to monitor California’s marine protected areas. Fisheries, 36(3):141–142, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.10389089.

Cronin:2011:THA

[994] Amanda Cronin. Travel and hotel accomodations easy in, easy out. Fisheries, 36(3):145–146, 2011. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2011.10389090.

Anonymous:2011:AU

[995] Anonymous. AFS Units. Fisheries, 36(3):148–149, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.10389091.

Anonymous:2011:MJa

[996] Anonymous. March 2011: Jobs. Fisheries, 36(3):150–151, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.10389092.

Anonymous:2011:TAFb

[997] Anonymous. Transactions of the American Fisheries Society, volume 140, issue 1. Fisheries, 36(3):152, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.10389093.

Anonymous:2011:FIPd

[998] Anonymous. Full issue PDF volume 36, issue 4. Fisheries, 36(4): 157–208, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2011.581164.

Hubert:2011:NFFa

[999] Wayne A. Hubert. New frontiers in fisheries management and ecology: Addressing contemporary environmental issues. Fisheries, 36 (4):161, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.564514.

Hawkes:2011:LPd

[1000] Elden W. Hawkes, Jr. Legislation and Policy. Fisheries, 36(4): 162–163, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.564510.

Fina:2011:ECS

[1001] Mark Fina. Evolution of catch share management: Lessons from catch share management in the North Pacific. Fisheries, 36(4): 164–177, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.564509.

Daniels:2011:LNY

[1002] Robert A. Daniels. Legacy of New York State’s watershed surveys, 1926– 1939. Fisheries, 36(4):179–189, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.564504.

Anonymous:2011:NAJb

[1003] Anonymous. North American Journal of Aquaculture, volume 73, number 1. Fisheries, 36(4):191, 2011. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2011.585556.

Cronin:2011:MTG

[1004] Amanda Cronin. A mini-travel guide on what to do, where to eat, where to stay and why. Fisheries, 36(4):192–194, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.564515.

Rassam:2011:RUFa

[1005] Gus Rassam. The role in US Federal Fisheries Staff and professional societies. Fisheries, 36(4):195, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.564500.

James:2011:RSB

[1006] Daniel James. Rock snot and boulder boogers. Fisheries, 36(4): 196, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.564516.

Winkler:2011:HJF

[1007] Kathryn Winkler. Hutton Junior Fisheries Biology Program. Fisheries, 36(4):197–199, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.564502.

Anonymous:2011:FEd

[1008] Anonymous. Fisheries events. Fisheries, 36(4):202, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.585562.

Anonymous:2011:AJa

[1009] Anonymous. April 2011: Jobs. Fisheries, 36(4):203–204, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.585563.

Anonymous:2011:FIPe

[1010] Anonymous. Full issue PDF volume 36, issue 5. Fisheries, 36(5): 209–260, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2011.583857.

Hubert:2011:NFFb

[1011] Wayne A. Hubert. New frontiers in fisheries management and ecology: Communicating research and technical information at our meetings. Fisheries, 36(5):213, 2011. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2011.574576.

Hawkes:2011:LPe

[1012] Elden W. Hawkes, Jr. Legislation and policy. Fisheries, 36(5): 214, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.574577.

Cucherousset:2011:EIN

[1013] Julien Cucherousset and Julian D. Olden. Ecological impacts of nonnative freshwater fishes. Fisheries, 36(5):215–230, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.574578.

Powles:2011:ARE

[1014] Howard Powles. Assessing risk of extinction of marine fishes in Canada — the COSEWIC experience. Fisheries, 36(5):231–246, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.574582.

Hilborn:2011:BRM

[1015] Ray Hilborn. Book review: Mismanagement of Marine Fisheries by Alan Longhurst. Fisheries, 36(5):247, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.564494.

Anonymous:2011:FEe

[1016] Anonymous. Fisheries events. Fisheries, 36(5):248, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.583850.

Anonymous:2011:ACA

[1017] Anonymous. AFS constitutional amendment. Fisheries, 36(5): 249, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.583851.

Anonymous:2011:RES

[1018] Anonymous. Resolution on ecological separation of the Great Lakes and Mississippi River drainage basins. Fisheries, 36(5):250, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.583852.

Anonymous:2011:CEP

[1019] Anonymous. Continuing education program. Fisheries, 36(5):251, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.574590.

Goldberg:2011:ACP

[1020] Gail S. Goldberg. AFS certification provides more credibility and more visibility and help for unit leaders. Fisheries, 36(5):252–253, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2011.574588.

Anonymous:2011:NAJc

[1021] Anonymous. North American Journal of Fisheries Management, volume 31, number 1. Fisheries, 36(5):256, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.583855.

Anonymous:2011:MJb

[1022] Anonymous. May 2011: Jobs. Fisheries, 36(5):257–258, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.583856.

Anonymous:2011:FIPf

[1023] Anonymous. Full issue PDF volume 36, issue 6. Fisheries, 36(6): 261–312, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2011.591657.

Hubert:2011:NFFc

[1024] Wayne A. Hubert. New frontiers in fisheries management and ecology: Advancing sound science. Fisheries, 36(6):265, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.582391.

Hawkes:2011:LPf

[1025] Elden W. Hawkes, Jr. Legislation and policy. Fisheries, 36(6): 266, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.591624.

Williams:2011:NFC

[1026] Jack E. Williams, Richard N. Williams, Russell F. Thurow, Leah Elwell, David P. Philipp, Fred A. Harris, Jeffrey L. Kershner, Patrick J. Martinez, Dirk Miller, Gordon H. Reeves, Christopher A. Frissell, and James R. Sedell. Native fish conservation areas: a vision for large-scale conservation of native fish communities. Fisheries, 36(6): 267–277, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.582398.

Dauwalter:2011:IIN

[1027] Daniel C. Dauwalter, John S. Sanderson, Jack E. Williams, and James R. Sedell. Identification and implementation of native fish conservation areas in the Upper Colorado River Basin. Fisheries, 36(6): 278–288, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.582411.

Anonymous:2011:FP

[1028] Anonymous. Fish and people. Fisheries, 36(6):289, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.591634.

Anonymous:2011:U

[1029] Anonymous. Units. Fisheries, 36(6):290–293, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.591635.

Dembkowski:2011:FSR

[1030] Dan J. Dembkowski. Fish species richness in oxbow lakes. Fisheries, 36(6):295, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.574589.

Anonymous:2011:RCP

[1031] Anonymous. Recently certified professionals. Fisheries, 36(6):296, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.582423.

Anonymous:2011:NAJd

[1032] Anonymous. North American Journal of Aquaculture, volume 73, number 2. Fisheries, 36(6):297, 2011. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2011.591642.

Anonymous:2011:NAMa

[1033] Anonymous. New AFS members. Fisheries, 36(6):298, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.582424.

Anonymous:2011:A

[1034] Anonymous. From the archives. Fisheries, 36(6):299, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.591648.

Millenbah:2011:EEM

[1035] Kelly F. Millenbah, Bjørn H. K. Wolter, and William W. Taylor. Education in the era of the millennials and implications for future fisheries professionals and conservation. Fisheries, 36(6):300–304, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.582428.

Anonymous:2011:FEf

[1036] Anonymous. Fisheries events. Fisheries, 36(6):307, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.582429.

Anonymous:2011:JJa

[1037] Anonymous. June 2011: Jobs. Fisheries, 36(6):308–309, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.582430.

Anonymous:2011:FIPg

[1038] Anonymous. Full issue PDF volume 36, issue 7. Fisheries, 36(7): 313–364, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2011.599690.

Hubert:2011:NFFd

[1039] Wayne A. Hubert and William Fisher. New frontiers in fisheries management and ecology: Affiliates of AFS. Fisheries, 36(7):317, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.589752.

Hawkes:2011:LPg

[1040] Elden W. Hawkes, Jr. Legislation and policy. Fisheries, 36(7): 318, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.599691.

Cowan:2011:RSG

[1041] James H. Cowan, Jr. Red snapper in the Gulf of Mexico and U.S. South Atlantic: Data, doubt, and debate. Fisheries, 36(7):319–331, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2011.589318.

McCrea-Strub:2011:PIE

[1042] A. McCrea-Strub, K. Kleisner, U. R. Sumaila, W. Swartz, R. Watson, D. Zeller, and D. Pauly. Potential impact of the Deepwater Horizon oil spill on commercial fisheries in the Gulf of Mexico. Fisheries, 36 (7):332–336, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.589334.

Douglas:2011:CBI

[1043] Marlis R. Douglas, Yuriy V. Slyn’ko, Steven Kohl, Charles R. Lane, Elena E. Slyn’ko, and Michael E. Douglas. Crossroad blues: an intersection of rivers, wetlands, and public policy. Fisheries, 36(7): 337–339, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.574585.

Anonymous:2011:NMa

[1044] Anonymous. New members. Fisheries, 36(7):340, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.599693.

Anonymous:2011:TBQ

[1045] Anonymous. The tackle box — 15 questions with Roar Sandodden. Fisheries, 36(7):341–342, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2011.589342.

Ostberg:2011:STF

[1046] Carl Ostberg and Jeff Duda. Seattle tours and field trips you won’t want to miss. Fisheries, 36(7):343–345, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.589344.

Vidergar:2011:U

[1047] Dmitri Vidergar and Randall Schultz. Units. Fisheries, 36(7):346– 349, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.589358.

Anonymous:2011:TAFc

[1048] Anonymous. Transactions of the American Fisheries Society, volume 140, number 2. Fisheries, 36(7):350, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2011.599694.

James:2011:ARL

[1049] Daniel A. James. The academic road less traveled. Fisheries, 36(7): 351–352, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.589750.

Lynch:2011:FNP

[1050] Abigail J. Lynch. Forging new perspectives of fisheries science and management. Fisheries, 36(7):353–354, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.599696.

Jackson:2011:RRR

[1051] Donald C. Jackson. Relativity, resiliency, and response in the wake of environmental challenges. Fisheries, 36(7):355–356, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.589751.

Anonymous:2011:FEg

[1052] Anonymous. Fisheries events. Fisheries, 36(7):358, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.589753.

Anonymous:2011:JJb

[1053] Anonymous. July 2011: Jobs. Fisheries, 36(7):360–361, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.589754.

Anonymous:2011:FIPh

[1054] Anonymous. Full issue PDF volume 36, issue 8. Fisheries, 36(8): 365–420, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2011.607336.

Hubert:2011:NFFe

[1055] Wayne A. Hubert. New frontiers in fisheries management and ecology: Leading the way in a changing world. Fisheries, 36(8):369, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.597665.

Hawkes:2011:ULPa

[1056] Elden W. Hawkes, Jr. Update: Legislation and policy. Fisheries, 36(8):370, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.607316.

Perkin:2011:SFT

[1057] Joshuah S. Perkin and Keith B. Gido. Stream fragmentation thresholds for a reproductive guild of Great Plains fishes. Fisheries, 36(8): 371–383, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.597666.

Donaldson:2011:CGG

[1058] Michael R. Donaldson, Connie M. O’Connor, Lisa A. Thompson, Andrew J. Gingerich, Sascha E. Danylchuk, René R. Duplain, and Steven J. Cooke. Contrasting global game fish and non-game fish species. Fisheries, 36(8):385–397, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.597672.

Anonymous:2011:JHN

[1059] Anonymous. Journal highlights: North American Journal of Fisheries Management, volume 31, number 2. Fisheries, 36(8):398, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.607324.

Anonymous:2011:AAR

[1060] Anonymous. AFS 2010 Annual Report. Fisheries, 36(8):399– 409, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.597675.

Anonymous:2011:NMb

[1061] Anonymous. New members. Fisheries, 36(8):410, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.597676.

Rassam:2011:RUFb

[1062] Gus Rassam. The role of U.S. Federal Fisheries Staff in professional societies — part II. Fisheries, 36(8):411, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2011.597673.

Anonymous:2011:CFE

[1063] Anonymous. Calendar: Fisheries events. Fisheries, 36(8):414, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2011.607332.

Anonymous:2011:AJb

[1064] Anonymous. August 2011: Jobs. Fisheries, 36(8):415–418, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2011.607334.

Anonymous:2011:FIPi

[1065] Anonymous. Full issue PDF volume 36, issue 9. Fisheries, 36(9): 421–472, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2011.616810.

Fisher:2011:FNB

[1066] Bill Fisher. Fisheries networks: Building ecological, social, and professional relationships. Fisheries, 36(9):425, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2011.610700.

Hawkes:2011:ULPb

[1067] Elden W. Hawkes, Jr. Update: Legislation and policy. Fisheries, 36 (9):426–427, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.609387.

Fraidenburg:2011:EWI

[1068] Michael E. Fraidenburg. Eleven ways to improve executive management of conflict. Fisheries, 36(9):428–435, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2011.607086.

Wang:2011:HSF

[1069] Lizhu Wang, Dana Infante, Peter Esselman, Arthur Cooper, Dayong Wu, William Taylor, Doug Beard, Gary Whelan, and Andrea Ostroff. A hierarchical spatial framework and database for the National River Fish Habitat Condition Assessment. Fisheries, 36(9):436–449, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.607075.

Hasler:2011:NAA

[1070] Caleb T. Hasler, Gavin C. Christie, Jack Imhof, Michael Power, and Steven J. Cooke. A network approach to addressing strategic fisheries, aquaculture, and aquatic sciences issues at a national scale: an introduction to a series of case studies from Canada. Fisheries, 36 (9):450–453, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.607739.

Anonymous:2011:NAMb

[1071] Anonymous. New AFS members. Fisheries, 36(9):454, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.609404.

Anonymous:2011:TC

[1072] Anonymous. Twin Cities 2012. Fisheries, 36(9):455–456, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.616804.

Anonymous:2011:MDH

[1073] Anonymous. In memoriam: David “Homer” Buck. Fisheries, 36 (9):457, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.607084.

Pate:2011:CRB

[1074] William M. Pate, William L. Stacy, Eric I. Gardunio, and Jesse M. Lepak. Collaborative research between current and future fisheries professionals: Facilitating AFS subunit participation. Fisheries, 36(9): 458–460, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.607083.

Foster:2011:FDC

[1075] Michael J. Foster, Chanda Bennett, Eleanor J. Sterling, and Nora Bynum. Fostering the development of conservation leadership at minority-serving institutions. Fisheries, 36(9):461–463, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.607741.

Cox:2011:PSE

[1076] Sean Cox. Pacific salmon environmental and life history models: Advancing science for sustainable salmon in the future. Fisheries, 36 (9):464, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.607742.

Anonymous:2011:JAAb

[1077] Anonymous. Journal of Aquatic Animal Health. Fisheries, 36(9): 465, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.609415.

Anonymous:2011:FEh

[1078] Anonymous. Fisheries events. Fisheries, 36(9):466, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2011.616806.

Anonymous:2011:SJ

[1079] Anonymous. September 2011: Jobs. Fisheries, 36(9):469–470, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2011.609418.

Anonymous:2011:FIPj

[1080] Anonymous. Full issue PDF volume 36, issue 10. Fisheries, 36(10): 473–524, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2011.626742.

Fisher:2011:CNA

[1081] Bill Fisher. Collaborative networks and AFS: How strong are our connections? Fisheries, 36(10):477, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.619873.

Hawkes:2011:LPh

[1082] Elden W. Hawkes, Jr. Legislation and policy. Fisheries, 36(10): 478–479, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.615656.

Smokorowski:2011:NHN

[1083] Karen E. Smokorowski, Normand Bergeron, Daniel Boisclair, Keith Clarke, Steven Cooke, Rick Cunjak, Jeff Dawson, Brett Eaton, Faye Hicks, Paul Higgins, Chris Katopodis, Michel Lapointe, Pierre Legendre, Michael Power, Robert Randall, Joseph Rasmussen, George Rose, Andre Saint-Hilaire, Brent Sellars, Gary Swanson, Nicholas Winfield, Roger Wysocki, and David Zhu. NSERC’s HydroNet: a national research network to promote sustainable hydropower and healthy aquatic ecosystems. Fisheries, 36(10):480–488, 2011. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2011.616459.

Lieb:2011:CMC

[1084] David A. Lieb, Raymond W. Bouchard, Robert F. Carline, Ted R. Nuttall, John R. Wallace, and Carrie L. Burkholder. Conservation and management of crayfishes: Lessons from Pennsylvania. Fisheries, 36 (10):489–507, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.607080.

Penney:2011:LSA

[1085] Zachary L. Penney. Live to spawn another day: Understanding the fuel efficiency of Snake River steelhead. Fisheries, 36(10):508, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.617220.

Anonymous:2011:SCP

[1086] Anonymous. Second call for papers: Twin Cities. Fisheries, 36(10): 509–510, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.626731.

Anonymous:2011:SBM

[1087] Anonymous. The Steven Berkeley Marine Conservation Fellowship winners. Fisheries, 36(10):511–512, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.617615.

Anonymous:2011:MJN

[1088] Anonymous. In memoriam: Joseph Nelson. Fisheries, 36(10):513– 514, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.618056.

Jackson:2011:ORF

[1089] Donald C. Jackson. Observations from recreational fishing in the Northern Gulf of Mexico one year after the Deepwater Horizon oil spill. Fisheries, 36(10):515–517, 2011. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2011.617223.

Anonymous:2011:NAJe

[1090] Anonymous. North American Journal of Aquaculture, volume 73, number 3. Fisheries, 36(10):518, 2011. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2011.626735.

Anonymous:2011:FEi

[1091] Anonymous. Fisheries events. Fisheries, 36(10):519, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.626736.

Anonymous:2011:OJ

[1092] Anonymous. October 2011: Jobs. Fisheries, 36(10):521, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.626737.

Anonymous:2011:FIPk

[1093] Anonymous. Full issue PDF volume 36, issue 11. Fisheries, 36(11): 525–576, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2011.634297.

Fisher:2011:HLA

[1094] Bill Fisher. How large is AFS? Fisheries, 36(11):529, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.634300.

Hawkes:2011:NNA

[1095] Elden W. Hawkes, Jr. NOAA/NMFS approves the Mid-Atlantic ACL/ AM Omnibus Amendment. Fisheries, 36(11):530–531, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.626725.

Anonymous:2011:CYV

[1096] Anonymous. Cast your vote on immediate-release sedatives policy statement. Fisheries, 36(11):532–533, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.634301.

McGarvey:2011:SMP

[1097] Daniel J. McGarvey and John M. Johnston. A simple method to predict regional fish abundance: an example in the McKenzie River Basin, Oregon. Fisheries, 36(11):534–546, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.626659.

Paquet:2011:HCS

[1098] P. J. Paquet, T. Flagg, A. Appleby, J. Barr, L. Blankenship, D. Campton, M. Delarm, T. Evelyn, D. Fast, J. Gislason, P. Kline, D. Maynard, L. Mobrand, G. Nandor, P. Seidel, and S. Smith. Hatcheries, conservation, and sustainable fisheries-achieving multiple goals: Results of the Hatchery Scientific Review Group’s Columbia River Basin review. Fisheries, 36(11):547–561, 2011. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2011.626661.

Pandolfo:2011:LEF

[1099] Tamara Pandolfo. Living on the edge: Freshwater mussels on the brink of extinction. Fisheries, 36(11):562, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2011.626662.

Anonymous:2011:NFS

[1100] Anonymous. New frontiers in Seattle: New frontiers in fisheries management and ecology: Leading the way in a changing world — was an AFS record-breaking event! Fisheries, 36(11):563–564, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.634303.

Anonymous:2011:TCP

[1101] Anonymous. Third call for papers: Twin Cities 2012. Fisheries, 36 (11):565–566, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.634305.

Rassam:2011:SAH

[1102] Gus Rassam. Seattle 2011: AFS highlights. Fisheries, 36(11): 567, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.634306.

Anonymous:2011:TAFd

[1103] Anonymous. Transactions of the American Fisheries Society, volume 140, number 4. Fisheries, 36(11):569, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2011.634307.

Scarnecchia:2011:MPA

[1104] Dennis Scarnecchia. In memoriam: Phillip Allen Stewart. Fisheries, 36(11):570, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.634308.

Anonymous:2011:FEj

[1105] Anonymous. Fisheries events. Fisheries, 36(11):572, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.634309.

Anonymous:2011:NJ

[1106] Anonymous. November 2011: Jobs. Fisheries, 36(11):573, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.634310.

Anonymous:2011:FIPl

[1107] Anonymous. Full issue PDF volume 36, issue 12. Fisheries, 36(12): 577–628, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2011.640237.

Fisher:2011:CFE

[1108] Bill Fisher. The changing fisheries education and employment connection. Fisheries, 36(12):581, 2011. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2011.640225.

Hawkes:2011:HCN

[1109] Elden W. Hawkes, Jr. House Committee on Natural Resources holds hearing on national ocean policy. Fisheries, 36(12):582, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.633468.

Cooke:2011:OTN

[1110] Steven J. Cooke, Sara J. Iverson, Michael J. W. Stokesbury, Scott G. Hinch, Aaron T. Fisk, David L. VanderZwaag, Richard Apostle, and Fred Whoriskey. Ocean Tracking Network Canada: a network approach to addressing critical issues in fisheries and resource management with implications for ocean governance. Fisheries, 36(12):583–592, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2011.633464.

Yochum:2011:UFK

[1111] Noëlle Yochum, Richard M. Starr, and Dean E. Wendt. Utilizing fishermen knowledge and expertise: Keys to success for collaborative fisheries research. Fisheries, 36(12):593–605, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2011.633467.

Anonymous:2011:AAM

[1112] Anonymous. AFS 141^{st} Annual Meeting, September 4–8, 2011, Seattle, Washington. Fisheries, 36(12):606–613, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2011.640227.

Anonymous:2011:FCP

[1113] Anonymous. Fourth call for papers: Twin Cities. Fisheries, 36(12): 614–615, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.640228.

Anonymous:2011:NAMc

[1114] Anonymous. New AFS members. Fisheries, 36(12):616–617, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2011.640229.

Baccante:2011:FRN

[1115] D. A. Baccante, B. A. Barton, M. A. Bozek, and J. C. Bruner. Future research needs of walleye and sauger. Fisheries, 36(12):618–619, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2011.640231.

Anonymous:2011:NAJf

[1116] Anonymous. North American Journal of Fisheries Management, volume 31, number 3. Fisheries, 36(12):620, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2011.640232.

Waldman:2011:BR

[1117] John Waldman. Book review. Fisheries, 36(12):621, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.644096.

Anonymous:2011:FEk

[1118] Anonymous. Fisheries events. Fisheries, 36(12):623, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2011.640233.

Anonymous:2011:DJ

[1119] Anonymous. December 2011: Jobs. Fisheries, 36(12):624–625, 2011. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2011.640235.

Anonymous:2011:AIF

[1120] Anonymous. Annual index: Fisheries volume 36 2011. Fisheries, 36 (12):626, 2011. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2011.640236.

Anonymous:2012:FIPa

[1121] Anonymous. Full issue PDF volume 37, issue 1. Fisheries, 37 (1):1–52, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.655554.

Fisher:2012:WJA

[1122] Bill Fisher. Why join AFS? Fisheries, 37(1):3, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.639677.

Rassam:2012:H

[1123] A. Yasmine Rassam. Headliners. Fisheries, 37(1):4–5, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.639678.

Hamm:2012:DED

[1124] David E. Hamm. Development and evaluation of a data dictionary to standardize salmonid habitat assessments in the Pacific Northwest. Fisheries, 37(1):6–18, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.639679.

Alsharif:2012:DEA

[1125] Kamal A. Alsharif and Nathan Miller. Data envelopment analysis to evaluate the reliance and engagement of Florida communities on Gulf of Mexico commercial red snapper fisheries. Fisheries, 37(1):19–26, 2012. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2012.639680.

Anonymous:2012:QWF

[1126] Anonymous. Q&A on 6th World Fisheries Congress: an interview with Michel Kaiser, Vice President of the World Fisheries Congress. Fisheries, 37(1):27–28, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.639681.

Anonymous:2012:AMSa

[1127] Anonymous. AFS 2012 Minneapolis–Saint Paul: Where waters meet, people greet, and networks are born. Fisheries, 37(1):29, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.639682.

Anonymous:2012:AFSa

[1128] Anonymous. 2012 American Fisheries Society awards. Fisheries, 37 (1):30–33, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.639685.

Anonymous:2012:AFSb

[1129] Anonymous. American Fisheries Society application for committee appointment. Fisheries, 37(1):34, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.639688.

Anonymous:2012:FGA

[1130] Anonymous. Fisheries 2012 guide for authors. Fisheries, 37(1): 35–37, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.639689.

Trushenski:2012:ASC

[1131] Jesse Trushenski, Thomas E. Bigford, and Paul J. Radomski. AFS seeks comment on lead in sport fishing tackle policy. Fisheries, 37 (1):38, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.654421.

Trushenski:2012:SHR

[1132] Jesse Trushenski. Symposia on hatcheries and the role of fish culture in the future of aquatic resources mark milestones at AFS 2011 Annual Meeting. Fisheries, 37(1):39, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.639690.

Baccante:2012:HFF

[1133] Dominic Baccante. Hydraulic fracturing: a fisheries biologist’s perspective. Fisheries, 37(1):40–41, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.639691.

Anonymous:2012:TAFa

[1134] Anonymous. Transactions of the American Fisheries Society, volume 140, number 5. Fisheries, 37(1):42, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.639692.

Anonymous:2012:FEa

[1135] Anonymous. Fisheries events. Fisheries, 37(1):45, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.639693.

Anonymous:2012:JJa

[1136] Anonymous. January 2012 jobs. Fisheries, 37(1):46–47, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.639695.

Anonymous:2012:FIPb

[1137] Anonymous. Full issue PDF volume 37, issue 2. Fisheries, 37(2): 47–98, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.660090.

Fisher:2012:VM

[1138] Bill Fisher. The value of meetings. Fisheries, 37(2):51, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.660068.

Anonymous:2012:RCS

[1139] Anonymous. Robert L. Curry: Second Vice President candidate statement. Fisheries, 37(2):52, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.660071.

Anonymous:2012:RJE

[1140] Anonymous. Ronald J. Essig: Second Vice President candidate statement. Fisheries, 37(2):53, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.660072.

ONeal:2012:FHR

[1141] Sarah O’Neal and Robert M. Hughes. Fisheries and hard rock mining: AFS Symposium synopsis. Fisheries, 37(2):54–55, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.661318.

Anonymous:2012:HYR

[1142] Anonymous. Headliners: A. Yasmine Rassam. Fisheries, 37(2): 56–59, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.660074.

Wuellner:2012:PCA

[1143] Melissa R. Wuellner and Donald C. Jackson. Population characteristics of AFS membership: Special focus on the millennial generation of fisheries professionals. Fisheries, 37(2):60–65, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2012.650991.

Schaeffer:2012:USW

[1144] Jeffrey S. Schaeffer, James K. Bland, and John Janssen. Use of a storm water retention system for conservation of regionally endangered fishes. Fisheries, 37(2):66–75, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.650992.

Lauer:2012:ILP

[1145] Thomas E. Lauer. Improving the learning process for both instructor and student using assessments. Fisheries, 37(2):76–79, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.650993.

Seitz:2012:RNG

[1146] Andrew C. Seitz, Katherine M. Straub, and Trent M. Sutton. Recruiting the next generation of fisheries professionals. Fisheries, 37 (2):80–83, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.650996.

Anonymous:2012:MDT

[1147] Anonymous. In memoriam: Dr. Theodore R. Rice. Fisheries, 37 (2):84, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.650997.

Anonymous:2012:BGW

[1148] Anonymous. Being green will be easy for AFS 2012. Fisheries, 37 (2):85, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.660078.

Knudsen:2012:AFD

[1149] Eric Knudsen, Larry Dominguez, and Cleve Steward. AFS facilitates development of ecologists without borders. Fisheries, 37(2): 86–87, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.650998.

Anonymous:2012:NAJa

[1150] Anonymous. North American Journal of Aquaculture, volume 73, number 4 and Journal of Aquatic Animal Health volume 23, number 4. Fisheries, 37(2):88, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.660082.

Anonymous:2012:NAMa

[1151] Anonymous. New AFS members. Fisheries, 37(2):89, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.660083.

Anonymous:2012:FEb

[1152] Anonymous. Fisheries events. Fisheries, 37(2):93, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.660085.

Anonymous:2012:FJ

[1153] Anonymous. February 2012 jobs. Fisheries, 37(2):94–95, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.660089.

Anonymous:2012:FIPc

[1154] Anonymous. Full issue PDF volume 37, issue 3. Fisheries, 37(3): 97–144, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.671045.

Fisher:2012:OCP

[1155] Bill Fisher. Our changing professional society. Fisheries, 37(3): 99, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.660029.

Anonymous:2012:Ha

[1156] Anonymous. Headliners. Fisheries, 37(3):100–102, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.669610.

Hasselman:2012:RED

[1157] Daniel J. Hasselman, Richard A. Hinrichsen, Barbara A. Shields, and Curtis C. Ebbesmeyer. The rapid establishment, dispersal, and increased abundance of invasive American shad in the Pacific Northwest. Fisheries, 37(3):103–114, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.659938.

Hasselman:2012:ASP

[1158] Daniel J. Hasselman, Richard A. Hinrichsen, Barbara A. Shields, and Curtis C. Ebbesmeyer. American shad of the Pacific Coast: a harmful invasive species or benign introduction? Fisheries, 37(3): 115–122, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.659941.

Anonymous:2012:SUa

[1159] Anonymous. Section updates. Fisheries, 37(3):123–125, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.669611.

OConnor:2012:HFG

[1160] Constance M. O’Connor. How to find a good graduate advisor and make the most of graduate school. Fisheries, 37(3):126–128, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.659942.

Rassam:2012:AAM

[1161] Gus Rassam. AFS Annual Meetings: Why and where we hold them; why they are expensive; and who does what. Fisheries, 37 (3):129–130, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.659947.

Avyle:2012:MJJ

[1162] Mike Van Den Avyle, Mike Allen, Bill Pine, and Rich Noble. In memoriam: John (Jack) F. Dequine, 1917–2011. Fisheries, 37(3): 131, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.659944.

Anonymous:2012:SWF

[1163] Anonymous. Spotlight on women in fisheries. Fisheries, 37(3): 132–135, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.669617.

Anonymous:2012:TCM

[1164] Anonymous. Twin Cities meeting update: AFS 2012 Minneapolis– Saint Paul: Where to stay and how to get around. Fisheries, 37(3): 136–137, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.669620.

Anonymous:2012:NAJb

[1165] Anonymous. North American Journal of Fisheries Management, volume 31, number 6. Fisheries, 37(3):139, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.669621.

Anonymous:2012:FEc

[1166] Anonymous. Fisheries events. Fisheries, 37(3):142, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.669623.

Anonymous:2012:MJa

[1167] Anonymous. March 2012 jobs. Fisheries, 37(3):143, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.669638.

Anonymous:2012:FIPd

[1168] Anonymous. Full issue PDF volume 37, issue 4. Fisheries, 37(4): 145–192, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.676473.

Fisher:2012:OIC

[1169] Bill Fisher. Our international connections. Fisheries, 37(4):147, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.676474.

Anonymous:2012:Hb

[1170] Anonymous. Headliners. Fisheries, 37(4):148–152, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.676475.

Arlinghaus:2012:PAA

[1171] Robert Arlinghaus, Alexander Schwab, Carsten Riepe, and Tara Teel. A primer on anti-angling philosophy and its relevance for recreational fisheries in urbanized societies. Fisheries, 37(4):153–164, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.666472.

Arlinghaus:2012:BRA

[1172] Robert Arlinghaus, T. Douglas Beard, Jr., Steven J. Cooke, and Ian G. Cowx. Benefits and risks of adopting the global code of practice for recreational fisheries. Fisheries, 37(4):165–172, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.666473.

Anonymous:2012:AMSb

[1173] Anonymous. AFS 2012 Minneapolis–Saint Paul update: things to see and do. Fisheries, 37(4):173–176, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.676476.

McGarvey:2012:UFW

[1174] Daniel McGarvey. U.S. Fish and Wildlife Service is expecting multiples. Fisheries, 37(4):177, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2012.666476.

Belgrano:2012:EBF

[1175] Andrea Belgrano and Charles W. Fowler. Ecosystem-based fisheries management: Confronting tradeoffs. Fisheries, 37(4):178, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.666478.

Higley:2012:FFN

[1176] Corrine Higley. The 2010–2011 Fenske Fellowship: New perspectives in fisheries management. Fisheries, 37(4):179–180, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.666486.

Jackson:2012:LSS

[1177] Donald C. Jackson. The liturgy and sacraments of science for naturalists, poets, and artists in our ranks. Fisheries, 37(4):181–182, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.666488.

Anonymous:2012:NAMb

[1178] Anonymous. New AFS members. Fisheries, 37(4):183, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.676477.

Anonymous:2012:TAFb

[1179] Anonymous. Transactions of the American Fisheries Society, volume 141, number 1, January 2012. Fisheries, 37(4):185, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.676478.

Anonymous:2012:AAMa

[1180] Anonymous. AFS 2012 Annual Meeting Minneapolis–St. Paul, August 19–23, 2012 scheduled symposia. Fisheries, 37(4):186, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.676479.

Anonymous:2012:FEd

[1181] Anonymous. Fisheries events. Fisheries, 37(4):189, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.676480.

Anonymous:2012:AJa

[1182] Anonymous. April 2012 jobs. Fisheries, 37(4):190, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.676481.

Anonymous:2012:FIPe

[1183] Anonymous. Full issue PDF volume 37, issue 5. Fisheries, 37(5): 193–240, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.686833.

Fisher:2012:ENT

[1184] Bill Fisher. Ecological networks: Their role in fisheries. Fisheries, 37(5):195, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.676830.

Anonymous:2012:Hc

[1185] Anonymous. Headliners. Fisheries, 37(5):196–200, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.676831.

Hughes:2012:PCP

[1186] Robert M. Hughes and Alan T. Herlihy. Patterns in catch per unit effort of native prey fish and alien piscivorous fish in 7 Pacific Northwest USA rivers. Fisheries, 37(5):201–211, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.676833.

Long:2012:AFF

[1187] James M. Long, Nathan P. Nibbelink, Kevin T. McAbee, and Julie W. Stahli. Assessment of freshwater fish assemblages and their habitats in the National Park Service System of the Southeastern United States. Fisheries, 37(5):212–225, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.676835.

Gause:2012:SU

[1188] Brian Gause. Section updates. Fisheries, 37(5):226–227, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.676832.

Anonymous:2012:MMB

[1189] Anonymous. In memoriam: Mark B. Bain (1955–2012). Fisheries, 37(5):229, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.676844.

Anonymous:2012:CEP

[1190] Anonymous. Continuing education program: American Fisheries Society 142^{nd} Annual Meeting, Twin Cities 2012, August 18^{th}–August 19^{th}. Fisheries, 37(5):230–232, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.686840.

Varanasi:2012:FSA

[1191] Usha Varanasi. Frontloading the science in anticipation of environmental disasters. Fisheries, 37(5):233–235, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.676837.

Anonymous:2012:NAJc

[1192] Anonymous. North American Journal of Aquaculture Volume 74, Number 1, January 2012/Journal of Aquatic Animal Health Volume 24, Number 1, March 2012. Fisheries, 37(5):236, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2012.686841.

Anonymous:2012:FEe

[1193] Anonymous. Fisheries events. Fisheries, 37(5):237, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.686842.

Anonymous:2012:MJb

[1194] Anonymous. May 2012 jobs. Fisheries, 37(5):238–239, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.686843.

Anonymous:2012:FIPf

[1195] Anonymous. Full issue PDF volume 37, issue 6. Fisheries, 37(6): 241–288, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.698160.

Fisher:2012:SNT

[1196] Bill Fisher. Social networks: Their role in fisheries. Fisheries, 37 (6):243, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.687246.

Anonymous:2012:Hd

[1197] Anonymous. Headliners. Fisheries, 37(6):244–245, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.687250.

DeVries:2012:ERL

[1198] Paul DeVries, Kevin L. Fetherston, Angelo Vitale, and Sue Madsen. Emulating riverine landscape controls of beaver in stream restoration. Fisheries, 37(6):246–255, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.687263.

Schultz:2012:CSH

[1199] Luke D. Schultz and Justin A. VanDeHey. A comparison of stipends, health insurance, and tuition remission policies at fisheries and wildlife graduate programs throughout the United States. Fisheries, 37(6): 257–263, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.687259.

Anonymous:2012:SUb

[1200] Anonymous. Section updates. Fisheries, 37(6):264–267, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.687253.

Anonymous:2012:NAMc

[1201] Anonymous. New AFS Members. Fisheries, 37(6):268, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.698152.

Bohnsack:2012:AMD

[1202] Brian L. Bohnsack, Ronald J. Essig, Robert L. Curry, and Douglas D. Nygren. A 75th anniversary must-do. Fisheries, 37(6):269–270, 2012. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2012.687275.

Ogle:2012:USG

[1203] Derek H. Ogle and Kevin F. Schanning. Usage of “Sex” and “Gender”. Fisheries, 37(6):271–272, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.687265.

Slaughter:2012:GCH

[1204] Joey Slaughter. Georgia Chapter holds 20th anniversary meeting in Macon. Fisheries, 37(6):273–274, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.687267.

Kittrell:2012:VCH

[1205] William B. Kittrell, Jr. Virginia Chapter holds its 22nd Annual Meeting in Blacksburg. Fisheries, 37(6):275–276, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2012.687268.

Anonymous:2012:NRE

[1206] Anonymous. Natural resource education and employment conference report and recommendations. Fisheries, 37(6):277–284, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.698159.

Anonymous:2012:FEf

[1207] Anonymous. Fisheries events. Fisheries, 37(6):285, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.687279.

Anonymous:2012:JJb

[1208] Anonymous. June 2012 jobs. Fisheries, 37(6):286, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.687281.

Anonymous:2012:NAJd

[1209] Anonymous. North American Journal of Fisheries Management volume 32, number 2, April 2012. Fisheries, 37(6):287, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.687278.

Anonymous:2012:FIPg

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.706500.

Fisher:2012:IES

[1211] Bill Fisher. Integrating ecological and social networks in fisheries. Fisheries, 37(7):291, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.695995.

Jackson:2012:HMP

[1212] Don Jackson. Headliners: Mississippi’s Pascagoula River designated as a model river in America’s Great Outdoor Rivers Program. Fisheries, 37(7):294–295, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.695993.

Snelgrove:2012:CHO

[1213] Paul V. R. Snelgrove, Philippe Archambault, S. Kim Juniper, Peter Lawton, Anna Metaxas, Pierre Pepin, Jake C. Rice, and Verena Tunnicliffe. Canadian healthy oceans network (CHONe): an academic-government partnership to develop scientific guidelines for conservation and sustainable usage of marine biodiversity. Fisheries, 37(7):296–304, 2012. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2012.696002.

Whited:2012:RAT

[1214] Diane C. Whited, John S. Kimball, John A. Lucotch, Niels K. Maumenee, Huan Wu, Samantha D. Chilcote, and Jack A. Stanford. A riverscape analysis tool developed to assist wild salmon conservation across the North Pacific Rim. Fisheries, 37(7):305–314, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.696009.

Burt:2012:IFS

[1215] J. M. Burt, M. R. Donaldson, K. A. Hruska, S. G. Hinch, and J. S. Richardson. Interactive field site visits can help students translate scientific studies into contextual understanding. Fisheries, 37(7): 315–319, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.696016.

Midway:2012:ASW

[1216] Steve Midway. AFS 2011 Student Writing Contest — honorable mention winner: Filth, flows, and family: Pressures mount on a rare stream catfish. Fisheries, 37(7):320–321, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.696019.

DuPont:2012:ICH

[1217] Joe DuPont. Idaho Chapter holds annual meeting in Coeur d’Alene. Fisheries, 37(7):322–324, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.696020.

Martin:2012:LHC

[1218] Jim Martin, William W. Taylor, and Kelsey M. Schlee. Leadership in the hurricane of change. Fisheries, 37(7):325–328, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.696022.

Skud:2012:LEW

[1219] Bernard Einar Skud. Leslie Edward Whitesel. Fisheries, 37(7): 329, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.696025.

Anonymous:2012:AAMb

[1220] Anonymous. AFS 2012 Annual Meeting tours and eats. Fisheries, 37(7):330, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.696026.

Anonymous:2012:TAFc

[1221] Anonymous. Transactions of the American Fisheries Society, volume 141, number 3, May 2012. Fisheries, 37(7):331, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.696027.

Anonymous:2012:FEg

[1222] Anonymous. Fisheries events. Fisheries, 37(7):333, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.696028.

Anonymous:2012:JJc

[1223] Anonymous. July 2012 jobs. Fisheries, 37(7):335–336, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.696030.

Anonymous:2012:FIPh

[1224] Anonymous. Full issue PDF volume 37, issue 8. Fisheries, 37(8): 337–384, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.716293.

Fisher:2012:FAN

[1225] Bill Fisher. The functioning AFS network. Fisheries, 37(8):339, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.704820.

McClelland:2012:LTI

[1226] Michael A. McClelland, Greg G. Sass, Thad R. Cook, Kevin S. Irons, Nerissa N. Michaels, T. Matthew O’Hara, and Cammy S. Smith. The long-term Illinois River fish population monitoring program. Fisheries, 37(8):340–350, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.704815.

Ward:2012:TAP

[1227] David L. Ward, Benjamin J. Clemens, David Clugston, Aaron D. Jackson, Mary L. Moser, Chris Peery, and David P. Statler. Translocating adult Pacific lamprey within the Columbia River Basin: State of the science. Fisheries, 37(8):351–361, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.704818.

Roberts:2012:TSI

[1228] Mervin F. Roberts. Tilapia in South India. Fisheries, 37(8): 362, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.704828.

Fisher:2012:FNB

[1229] William A. Fisher and Gus Rassam. Fisheries networks: Building ecological, social, and professional relationships. Fisheries, 37(8): 363–374, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.716296.

Yanke:2012:OCA

[1230] Jeff Yanke and Allison Evans. Oregon Chapter Annual Meeting. Fisheries, 37(8):375, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.704833.

Moffitt:2012:DNR

[1231] Christine M. Moffitt. Diversity in natural resource science professions. Fisheries, 37(8):376–377, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.716321.

Anonymous:2012:NAJe

[1232] Anonymous. North American Journal of Aquaculture volume 74, number 2, April 2012. Fisheries, 37(8):378, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.716323.

Anonymous:2012:NAMd

[1233] Anonymous. New AFS Members. Fisheries, 37(8):379, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.716325.

Anonymous:2012:FEh

[1234] Anonymous. Fisheries events. Fisheries, 37(8):380, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.716327.

Anonymous:2012:AJb

[1235] Anonymous. August 2012 jobs. Fisheries, 37(8):381, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.716329.

Anonymous:2012:FIPi

[1236] Anonymous. Full issue PDF volume 37, issue 9. Fisheries, 37(9): 385–432, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.724643.

Boreman:2012:PCA

[1237] John Boreman. Preparing for the challenges ahead. Fisheries, 37 (9):387, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.714313.

Anonymous:2012:He

[1238] Anonymous. Headliners. Fisheries, 37(9):388–389, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.724635.

Trushenski:2012:AAA

[1239] Jesse Trushenski, Lorenzo Juarez, Gary L. Jensen, Mike Freeze, Michael Schwarz, Jeff Silverstein, Joel Bader, Jill Rolland, and Michael Rubino. AFS and aquaculture — addressing the high stakes of a sustainable seafood supply. Fisheries, 37(9):390–396, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2012.714315.

ODor:2012:CFE

[1240] Ron O’Dor, Andre M. Boustany, Cedar M. Chittenden, Mark J. Costello, Hassan Moustahfid, John Payne, Dirk Steinke, Michael J. W. Stokesbury, and Edward Vanden Berghe. A census of fishes and everything they eat: How the census of marine life advanced fisheries science. Fisheries, 37 (9):398–409, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.714323.

Wang:2012:RES

[1241] You-Gan Wang and Na Wang. A retrospective evaluation of sustainable yields for Australia’s northern prawn fishery. Fisheries, 37(9): 410–416, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.714326.

Anonymous:2012:LR

[1242] Anonymous. Little Rock 2013. Fisheries, 37(9):418–419, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.724637.

Anonymous:2012:WRS

[1243] Anonymous. Wild Rose State Fish Hatchery, Wild Rose, WI. Fisheries, 37(9):420–421, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.714329.

Anonymous:2012:DFA

[1244] Anonymous. Department of Fisheries and Allied Aquacultures Auburn University, Auburn, Alabama. Fisheries, 37(9):422, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.724639.

Long:2012:CBA

[1245] James M. Long and Joe E. Slaughter Iv. Certification benefits AFS members working in the private sector. Fisheries, 37(9):424–425, 2012. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2012.714330.

Anonymous:2012:NAJf

[1246] Anonymous. North American Journal of Fisheries Management, volume 32, number 4, August 2012. Fisheries, 37(9):426, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.724640.

Anonymous:2012:FEi

[1247] Anonymous. Fisheries events. Fisheries, 37(9):427, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2012.724641.

Anonymous:2012:SJ

[1248] Anonymous. September 2012 jobs. Fisheries, 37(9):428–429, 2012. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2012.724642.

Anonymous:2012:FIPj

[1249] Anonymous. Full issue PDF volume 37, issue 10. Fisheries, 37(10): 433–480, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.733244.

Boreman:2012:WDS

[1250] John Boreman. Where do I sign up? Fisheries, 37(10):435, 2012. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2012.723968.

Bowker:2012:H

[1251] J. Bowker and J. Trushenski. Headliners. Fisheries, 37(10):436– 439, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.723966.

Root:2012:DCI

[1252] Samantha Root and Catherine M. O’Reilly. Didymo control: Increasing the effectiveness of decontamination strategies and reducing spread. Fisheries, 37(10):440–448, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.722873.

Pracheil:2012:SPP

[1253] Brenda M. Pracheil, Mark A. Pegg, Larkin A. Powell, and Gerald E. Mestl. Swimways: Protecting paddlefish through movement-centered management. Fisheries, 37(10):449–457, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2012.722877.

Martin:2012:UIU

[1254] Dustin R. Martin, Brenda M. Pracheil, Jason A. DeBoer, Gene R. Wilde, and Kevin L. Pope. Using the Internet to understand angler behavior in the information age. Fisheries, 37(10):458–463, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.722875.

Roberts:2012:EE

[1255] Mervin F. Roberts. East is East and …. Fisheries, 37(10):464– 465, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.722879.

Anonymous:2012:SCP

[1256] Anonymous. Second call for papers: Little Rock 2013. Fisheries, 37 (10):466–468, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.733238.

Anonymous:2012:SBMa

[1257] Anonymous. The Steven Berkeley Marine Conservation Fellowship winners. Fisheries, 37(10):469–470, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.723965.

Finlayson:2012:RUF

[1258] Brian Finlayson, Rosalie Schnick, Don Skaar, Jon Anderson, Leo Demong, Dan Duffield, William Horton, Jarle Steinkjer, and Chris Van-Maaren. Rotenone use in fish management and Parkinson’s disease: Another look. Fisheries, 37(10):471–474, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2012.723963.

Cooney:2012:SWC

[1259] Patrick Cooney and Brandon Peoples. 2012 student writing contest cowinners. Fisheries, 37(10):475–476, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.723969.

Anonymous:2012:JAA

[1260] Anonymous. Journal of Aquatic Animal Health volume 24, number 2, June 2012, and Journal of Aquatic Animal Health volume 24, number 3, September 2012. Fisheries, 37(10):477, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2012.733240.

Anonymous:2012:FEj

[1261] Anonymous. Fisheries events. Fisheries, 37(10):478, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.733242.

Anonymous:2012:OJ

[1262] Anonymous. October 2012 jobs. Fisheries, 37(10):479, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.733243.

Anonymous:2012:FIPk

[1263] Anonymous. Full issue PDF volume 37, issue 11. Fisheries, 37(11): 481–528, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.743843.

Boreman:2012:ASR

[1264] John Boreman. Is AFS still relevant? Fisheries, 37(11):483, 2012. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2012.731882.

Freire:2012:OIR

[1265] Kátia M. F. Freire, Michel L. Machado, and Daniel Crepaldi. Overview of inland recreational fisheries in Brazil. Fisheries, 37(11):484– 494, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.731867.

Dettmers:2012:MAU

[1266] John M. Dettmers, Christopher I. Goddard, and Kelley D. Smith. Management of alewife using Pacific salmon in the Great Lakes: Whether to manage for economics or the ecosystem? Fisheries, 37(11): 495–501, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.731875. See response [1500].

Kerns:2012:IAP

[1267] Janice A. Kerns, Micheal S. Allen, and Julianne E. Harris. Importance of assessing population-level impact of catch-and-release mortality. Fisheries, 37(11):502–503, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.731878.

Anonymous:2012:TCW

[1268] Anonymous. Twin Cities wrap up: 142nd Annual Meeting of the American Fisheries Society, Minneapolis–St. Paul, MN, August 19-=23, 2012. Fisheries, 37(11):504–508, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.743837.

Anonymous:2012:TCP

[1269] Anonymous. Third call for papers: Little Rock 2013. Fisheries, 37 (11):509–511, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.731918.

Bisson:2012:JRS

[1270] Pete Bisson, Gordie Reeves, and Stan Gregory. James R. Sedell (July 5, 1944–August 18, 2012). Fisheries, 37(11):512–513, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.731889.

Bowker:2012:USB

[1271] Jim Bowker. The untold story of Bob Piper — AFS Honorary Member. Fisheries, 37(11):514, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.731896.

Schaeffer:2012:N

[1272] Jeff Schaeffer. Needlefish. Fisheries, 37(11):515, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.731911.

Piccolo:2012:SGF

[1273] John J. Piccolo. Stoking the “Green Fire”: Bringing the land ethic to the water. Fisheries, 37(11):516–518, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2012.731916.

Anonymous:2012:TAFd

[1274] Anonymous. Transactions of the American Fisheries Society, volume 141, number 5, September 2012. Fisheries, 37(11):519, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.743839.

Roberts:2012:NDD

[1275] Mervin F. Roberts. The night I disrupted a Danish airliner schedule. Fisheries, 37(11):520, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.731892.

Anonymous:2012:FEk

[1276] Anonymous. Fisheries events. Fisheries, 37(11):522, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.743840.

Anonymous:2012:NJ

[1277] Anonymous. November 2012 jobs. Fisheries, 37(11):523–524, 2012. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2012.743841.

Anonymous:2012:FIPl

[1278] Anonymous. Full issue PDF volume 37, issue 12. Fisheries, 37(12): 529–576, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2012.755036.

Boreman:2012:DLT

[1279] John Boreman. Developing a legacy through mentoring. Fisheries, 37(12):531, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.742803.

Lucente:2012:OLE

[1280] Joseph E. Lucente, Tory Gabriel, Gregory Davis, Colleen Wellington, and Frank Lichtkoppler. Ohio’s 2010 Lake Erie charter fishing industry. Fisheries, 37(12):532–541, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.742805.

Isaak:2012:PPF

[1281] Daniel J. Isaak, Clint C. Muhlfeld, Andrew S. Todd, Robert Alchokhachy, James Roberts, Jeffrey L. Kershner, Kurt D. Fausch, and Steven W. Hostetler. The past as prelude to the future for understanding 21st-century climate effects on Rocky Mountain trout. Fisheries, 37 (12):542–556, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.742808.

Anonymous:2012:SBMb

[1282] Anonymous. The Steven Berkeley Marine Conservation Fellowship. Fisheries, 37(12):557, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.742810.

Anonymous:2012:AFSc

[1283] Anonymous. 2012 American fisheries section award winners. Fisheries, 37(12):558, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.742811.

Anonymous:2012:CAN

[1284] Anonymous. Call for award nominations: American Fisheries Society awards 2013. Fisheries, 37(12):559–562, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2012.742819.

Anonymous:2012:FCP

[1285] Anonymous. Fourth call for papers. Fisheries, 37(12):563–565, 2012. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2012.742824.

Bellgraph:2012:NIF

[1286] Brian Bellgraph. News from the International Fisheries Section. Fisheries, 37(12):567, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.742827.

Juanes:2012:UFD

[1287] Francis Juanes. Using film documentaries as fisheries teaching tools. Fisheries, 37(12):568, 2012. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2012.742828.

Jackson:2012:GHS

[1288] Donald C. Jackson. Gifts for the holiday season. Fisheries, 37(12): 569–570, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.742830.

Anonymous:2012:NAJg

[1289] Anonymous. North American Journal of Aquaculture, volume 74, number 4, October 2012. Fisheries, 37(12):571, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2012.742832.

Anonymous:2012:FEl

[1290] Anonymous. Fisheries events. Fisheries, 37(12):572, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.742833.

Anonymous:2012:DJ

[1291] Anonymous. December 2012 jobs. Fisheries, 37(12):573, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2012.742834.

Anonymous:2012:FV

[1292] Anonymous. Fisheries volume 37 — 2012. Fisheries, 37(12): 575, 2012. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2012.742829.

Anonymous:2013:FIPa

[1293] Anonymous. Full issue PDF volume 38, issue 1. Fisheries, 38 (1):1–48, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.762297.

Boreman:2013:TYC

[1294] John Boreman. Teach your children well. Fisheries, 38(1):3, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.748554.

Weltman-Fahs:2013:HFB

[1295] Maya Weltman-Fahs and Jason M. Taylor. Hydraulic fracturing and brook trout habitat in the Marcellus Shale Region: Potential impacts and research needs. Fisheries, 38(1):4–15, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.750112.

Brown:2013:AFW

[1296] Grant E. Brown, Maud C. O. Ferrari, and Douglas P. Chivers. Adaptive forgetting: Why predator recognition training might not enhance poststocking survival. Fisheries, 38(1):16–25, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.750133.

Utz:2013:NEO

[1297] Ryan M. Utz, Michael R. Fitzgerald, Keli J. Goodman, Stephanie M. Parker, Heather Powell, and Charlotte L. Roehm. The National Ecological Observatory Network: an observatory poised to expand spatiotemporal scales of inquiry in aquatic and fisheries science. Fisheries, 38 (1):26–35, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.748551.

Fenn:2013:SSB

[1298] Carlin Fenn, Jeffrey Hillis, and Jesse Trushenski. SIUC subunit blends research and service in pursuit of professional development. Fisheries, 38(1):36, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.751850.

Anonymous:2013:SWC

[1299] Anonymous. Student writing contest now accepting submissions. Fisheries, 38(1):37, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.751859.

Trushenski:2013:AFS

[1300] Jesse Trushenski and Paul Radomski. American Fisheries Society adopts new policy, encourages efforts to understand and limit effects of lead in sport fishing tackle on fish and wildlife. Fisheries, 38(1):38, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.750141.

Anonymous:2013:FGA

[1301] Anonymous. Fisheries 2013 guide for authors. Fisheries, 38(1): 39–41, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.750146.

Lynch:2013:FFF

[1302] Abigail J. Lynch and William W. Taylor. The four Fs of fish: Communicating the public value of fish and fisheries. Fisheries, 38(1): 43–44, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.762305.

Anonymous:2013:NAJa

[1303] Anonymous. North American Journal of Fisheries Management, volume 32, number 6, December 2012. Fisheries, 38(1):45, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.750149.

Anonymous:2013:FEa

[1304] Anonymous. Fisheries events. Fisheries, 38(1):47, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.750151.

Anonymous:2013:JJ

[1305] Anonymous. January 2013 jobs. Fisheries, 38(1):48, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.750154.

Anonymous:2013:FIPb

[1306] Anonymous. Full issue PDF volume 38, issue 2. Fisheries, 38(2): 49–96, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.771544.

Boreman:2013:TGA

[1307] John Boreman. Think globally, act globally. Fisheries, 38(2): 51–52, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.757971.

Kolb:2013:HMD

[1308] Tracy L. Kolb, E. Agnes Blukacz-Richards, Andrew M. Muir, Randall M. Claramunt, Marten A. Koops, William W. Taylor, Trent M. Sutton, Michael T. Arts, and Ed Bissel. How to manage data to enhance their potential for synthesis, preservation, sharing, and reuse — a Great Lakes case study. Fisheries, 38(2):52–64, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.757975.

Murthy:2013:STF

[1309] Uma Murthy, Edward A. Fox, Yinlin Chen, Eric M. Hallerman, Donald J. Orth, Ricardo da S. Torres, Lin Tzy Li, Nádia P. Kozievitch, Felipe S. P. Andrade, Tiago R. C. Falcao, and Evandro Ramos. SuperIDR: a tool for fish identification and information retrieval. Fisheries, 38(2):65–75, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.757982.

Christensen:2013:ENF

[1310] Villy Christensen. Ecological networks in fisheries: Predicting the future? Fisheries, 38(2):76–81, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.757987.

Anonymous:2013:ALRa

[1311] Anonymous. AFS 2013 Little Rock: Preparing for the challenges ahead. Fisheries, 38(2):83, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2013.771536.

Dominguez:2013:BBC

[1312] Larry Dominguez. Beyond biology: Can we reconcile the land-water-humankind interface? Fisheries, 38(2):84–86, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.760279.

Galloway:2013:MSU

[1313] Ben Galloway. The Montana State University Student Subunit of the American Fisheries Society (MSUAFS). Fisheries, 38(2):87, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.771537.

Wuellner:2013:RPQ

[1314] Melissa R. Wuellner and Shannon Fisher. Revisiting the process for qualifying for the AFS/Sea Grant Best Student Presentation and Poster Awards. Fisheries, 38(2):88–90, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.757989.

Anonymous:2013:TAFa

[1315] Anonymous. Transactions of the American Fisheries Society, volume 142, number 1, January 2013. Fisheries, 38(2):93, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.771540.

Anonymous:2013:FEb

[1316] Anonymous. Fisheries events. Fisheries, 38(2):95, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.771541.

Anonymous:2013:FJ

[1317] Anonymous. February 2013 jobs. Fisheries, 38(2):96, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.771543.

Anonymous:2013:FIPc

[1318] Anonymous. Full issue PDF volume 38, issue 3. Fisheries, 38(3): 97–144, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.782734.

Boreman:2013:RFP

[1319] John Boreman. Replenishing the fisheries profession. Fisheries, 38 (3):99, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.767243.

Anonymous:2013:SVP

[1320] Anonymous. Second Vice President. Fisheries, 38(3):100–101, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.767242.

White:2013:DJF

[1321] Gwen White, Julie Claussen, Christine Moffitt, Brenda Norcross, and Donna Parrish. Dr. J. Frances Allen: Pioneer of women in fisheries.

Fisheries, 38(3):103–111, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.769433.

Peterson:2013:LCC

[1322] Douglas P. Peterson, Seth J. Wenger, Bruce E. Rieman, and Daniel J. Isaak. Linking climate change and fish conservation efforts using spatially explicit decision support tools. Fisheries, 38(3):112–127, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.769157.

Fuller:2013:NAS

[1323] Pam Fuller, Matt Neilson, and Dane H. Huge. The NAS Alert System: a look at the first eight years. Fisheries, 38(3):128–138, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.767241.

Anonymous:2013:ALRb

[1324] Anonymous. AFS 2013 Little Rock: Preparing for the challenges ahead. Fisheries, 38(3):139, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2013.782724.

Trushenski:2013:HMA

[1325] Jesse Trushenski and Jim Bowker. Hatcheries and Management of Aquatic Resources (HaMAR) Group focuses attention on issues surrounding hatcheries and hatchery-origin fish and shellfish. Fisheries, 38(3):140–141, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.776854.

Anonymous:2013:NAJb

[1326] Anonymous. North American Journal of Aquaculture, volume 75, number 1, January 2013. Fisheries, 38(3):142, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.782726.

Anonymous:2013:NAMa

[1327] Anonymous. New AFS Members. Fisheries, 38(3):143, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.769434.

Anonymous:2013:FEc

[1328] Anonymous. Fisheries events. Fisheries, 38(3):144, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.782728.

Anonymous:2013:FIPd

[1329] Anonymous. Full issue PDF volume 38, issue 4. Fisheries, 38(4): 145–192, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.792176.

Boreman:2013:GW

[1330] John Boreman. Getting the word out. Fisheries, 38(4):147, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.775573.

Anonymous:2013:LPO

[1331] Anonymous. Letter to President Obama. Fisheries, 38(4):148, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.777667.

Terra:2013:SSF

[1332] Bianca de Freitas Terra, Robert M. Hughes, and Francisco Gerson Araújo. Sampling sufficiency for fish assemblage surveys of tropical Atlantic forest streams, southeastern Brazil. Fisheries, 38(4): 150–158, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.775572.

Myrick:2013:SUE

[1333] Christopher A. Myrick, Zachary E. Underwood, and Kristoph D. Kinzli. Supporting undergraduate education with realistic laboratory exercises and research experience. Fisheries, 38(4):160–168, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.767240.

Kurzawski:2013:SSP

[1334] Kenneth F. Kurzawski, William L. Fisher, Dirk Miller, and James M. Long. The 2012 salary survey of public and private sector fisheries employers in the United States and Canada. Fisheries, 38(4): 169–178, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.777668.

Kopaska:2013:ASM

[1335] Jeff Kopaska and Sarah Gilbert Fox. AFS and social media. Fisheries, 38(4):179–184, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.779174.

Anonymous:2013:ALRc

[1336] Anonymous. AFS 2013 Little Rock: Preparing for the challenges ahead. Fisheries, 38(4):187, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2013.777666.

Page:2013:NSE

[1337] Lawrence M. Page, Héctor Espinosa-Pérez, Lloyd T. Findley, Carter R. Gilbert, Robert N. Lea, Nicholas E. Mandrak, and Richard L. Mayden. New seventh edition of Common and Scientific Names of Fishes: Changes include capitalization of common names. Fisheries, 38(4): 188–189, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.767244.

Anonymous:2013:JAAa

[1338] Anonymous. Journal of Aquatic Animal Health, volume 25, number 1, March 2013. Fisheries, 38(4):190, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.792173.

Anonymous:2013:FEd

[1339] Anonymous. Fisheries events. Fisheries, 38(4):191, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.792174.

Anonymous:2013:FIPe

[1340] Anonymous. Full issue PDF volume 38, issue 5. Fisheries, 38(5): 193–240, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.800351.

Boreman:2013:MOM

[1341] John Boreman. Meeting our meeting responsibility. Fisheries, 38 (5):195, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.785394.

Bigford:2013:HCU

[1342] Thomas E. Bigford. Habitat connects us all. Fisheries, 38(5): 196, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.792608.

Nelson:2013:SWS

[1343] Troy C. Nelson, William J. Gazey, Karl K. English, and Marvin L. Rosenau. Status of white sturgeon in the Lower Fraser River, British Columbia. Fisheries, 38(5):197–209, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.777664.

Oliver:2013:SHU

[1344] Matthew J. Oliver, Matthew W. Breece, Dewayne A. Fox, Danielle E. Haulsee, Josh T. Kohut, John Manderson, and Tom Savoy. Shrinking the haystack: Using an AUV in an integrated ocean observatory to map Atlantic sturgeon in the coastal ocean. Fisheries, 38(5):210–216, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.782861.

Anonymous:2013:RHS

[1345] Anonymous. Rick Hansen: Sturgeon hero. Fisheries, 38(5):217– 220, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.787856.

Small:2013:RPC

[1346] Brian C. Small and Elliott C. Kittel. Researching the physiology and culture of Scaphirhynchus sturgeon. Fisheries, 38(5):221–223, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.787611.

Coble:2013:CSR

[1347] Daniel W. Coble. Constantine Samuel Rafinesque 1783–1840. Fisheries, 38(5):224–226, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.787612.

Provost:2013:TBS

[1348] Mikaela Provost. Tagging black sea bass for sex change data. Fisheries, 38(5):227, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.779179.

Jones:2013:AGS

[1349] Marissa Jones. AFS Genetics Section members attend meeting on “Next Generation” DNA sequencing for fisheries research in Southeast Asia. Fisheries, 38(5):228–229, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.800352.

Anonymous:2013:GDD

[1350] Anonymous. Gary D. “Dox” Doxtater, 1936–2012 and John D. (Jack) McIntyre, 1938–2012. Fisheries, 38(5):230–232, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.779186.

Anonymous:2013:BRF

[1351] Anonymous. Booth reservation form: AFS–Little Rock, AR September 8–12, 2013. Fisheries, 38(5):233, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/abs/10.1080/03632415.2013.800353.

Anonymous:2013:CEP

[1352] Anonymous. Continuing education program. Fisheries, 38(5): 234, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.787865.

Arway:2013:SRF

[1353] John A. Arway and Geoffrey Smith. The Susquehanna River — a fishery in decline. Fisheries, 38(5):235–236, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.785395.

Anonymous:2013:NAMb

[1354] Anonymous. New AFS Members. Fisheries, 38(5):237, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.800354.

Anonymous:2013:NAJc

[1355] Anonymous. North American Journal of Fisheries Management, volume 33, number 2, April 2013. Fisheries, 38(5):239, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.800360.

Anonymous:2013:FEe

[1356] Anonymous. Fisheries events. Fisheries, 38(5):240, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.800361.

Anonymous:2013:FIPf

[1357] Anonymous. Full issue PDF volume 38, issue 6. Fisheries, 38(6): 241–292, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.808501.

Boreman:2013:SME

[1358] John Boreman. Scientific meetings are essential. Fisheries, 38(6): 245, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.797863.

Bigford:2013:WEF

[1359] Thomas E. Bigford. What exactly is fish habitat and why must we care? Fisheries, 38(6):246, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.797862.

Johnson:2013:CSF

[1360] Paul D. Johnson, Arthur E. Bogan, Kenneth M. Brown, Noel M. Burkhead, James R. Cordeiro, Jeffrey T. Garner, Paul D. Hartfield, Dwayne A. W. Lepitzki, Gerry L. Mackie, Eva Pip, Thomas A. Tarpley, Jeremy S. Tiemann, Nathan V. Whelan, and Ellen E. Strong. Conservation status of freshwater gastropods of Canada and the United States. Fisheries, 38(6):247–282, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.785396.

Meyer:2013:EAC

[1361] Joseph S. Meyer and Greg G. Pyle. Effects of anthropogenic chemicals on chemosensation and behavior in fish: Organismal, ecological, and regulatory implications. Fisheries, 38(6):283–284, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.797299.

Trushenski:2013:NOS

[1362] Jesse T. Trushenski. New and ongoing society initiatives to craft a lasting partnership between AFS and all things aquaculture. Fisheries, 38(6):285, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.797300.

Anonymous:2013:TAFb

[1363] Anonymous. Transactions of the American Fisheries Society, volume 142, number 3, May 2013. Fisheries, 38(6):286, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.808508.

Anonymous:2013:FEf

[1364] Anonymous. Fisheries events. Fisheries, 38(6):287, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.808509.

Anonymous:2013:NAMc

[1365] Anonymous. New AFS Members. Fisheries, 38(6):288, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.808510.

Anonymous:2013:FIPg

[1366] Anonymous. Full issue PDF volume 38, issue 7. Fisheries, 38(7): 293–340, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.820962.

Boreman:2013:WT

[1367] John Boreman. The world of tomorrow. Fisheries, 38(7):295, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.808083.

Bigford:2013:ITA

[1368] Thomas E. Bigford. The imperative for timely action. Fisheries, 38 (7):296, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.805691.

Chopin:2013:CIM

[1369] Thierry Chopin, Bruce MacDonald, Shawn Robinson, Stephen Cross, Christopher Pearce, Duncan Knowler, Anthony Noce, Gregor Reid, Andrew Cooper, David Speare, Les Burridge, Curran Crawford, Manav Sawhney, Keng Pee Ang, Clare Backman, and Marilyn Hutchinson. The Canadian Integrated Multi-Trophic Aquaculture Network (CIMTAN) — a network for a new era of ecosystem responsible aquaculture. Fisheries, 38(7):297–308, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.791285.

Wagner:2013:DTT

[1370] Tyler Wagner, Brian J. Irwin, James R. Bence, and Daniel B. Hayes. Detecting temporal trends in freshwater fisheries surveys: Statistical power and the important linkages between management questions and monitoring objectives. Fisheries, 38(7):309–319, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.799466.

Bunch:2013:MEF

[1371] Aaron J. Bunch, Carl J. Walters, and Lewis G. Coggins, Jr. Measurement error in fish lengths: Evaluation and management implications. Fisheries, 38(7):320–326, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.803472.

Trushenski:2013:SWY

[1372] Jesse T. Trushenski and Margaret H. Murphy. Stepping up: Why you should consider running for an AFS Office. Fisheries, 38(7): 327, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.820963.

Anonymous:2013:MFO

[1373] Anonymous. Meet five of our new members! Fisheries, 38(7): 328–330, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.806308.

Anonymous:2013:DRC

[1374] Anonymous. Dr. Robert S. Campbell, 1913–2012. Fisheries, 38 (7):331, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.808086.

Jones:2013:EDG

[1375] Marissa Jones. Environmental DNA: Genetics steps forward when traditional ecological surveys fall short. Fisheries, 38(7):332–333, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.810984.

Conroy:2013:FNP

[1376] Joseph D. Conroy and David I. Wellman. Forming new partnerships… and eating hotdogs! Fisheries, 38(7):334, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.787931.

Anonymous:2013:NAMd

[1377] Anonymous. New AFS Members/newly certified professionals. Fisheries, 38(7):335, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.806305.

Anonymous:2013:NAJd

[1378] Anonymous. North American Journal of Aquaculture, volume 75, number 2, April 2013. Fisheries, 38(7):336, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.820966.

Anonymous:2013:FEg

[1379] Anonymous. Fisheries events. Fisheries, 38(7):339, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.806306.

Anonymous:2013:AFSa

[1380] Anonymous. American Fisheries Society application for committee appointment. Fisheries, 38(7):340, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.820969.

Anonymous:2013:FIPh

[1381] Anonymous. Full issue PDF volume 38, issue 8. Fisheries, 38(8): 341–388, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.830008.

Boreman:2013:BF

[1382] John Boreman. On behalf of the fish. Fisheries, 38(8):343, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.813484.

Bigford:2013:EP

[1383] Thomas E. Bigford. The ecosystem perspective. Fisheries, 38(8): 344, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.813485.

Tracey:2013:STS

[1384] Sean Tracey, Colin Buxton, Caleb Gardner, Bridget Green, Klaas Hartmann, Marcus Haward, Julia Jabour, Jeremy Lyle, and Jan McDonald. Super trawler scuppered in Australian fisheries management reform. Fisheries, 38(8):345–350, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.813486.

Anonymous:2013:C

[1385] Anonymous. Correction. Fisheries, 38(8):351, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.830012.

Hayer:2013:PPC

[1386] Cari-Ann Hayer, Mark Kaemingk, Jason J. Breeggemann, Daniel Dembkowski, David Deslauriers, and Tobias Rapp. Pressures to publish: Catalysts for the loss of scientific writing integrity? Fisheries, 38 (8):352–355, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.813845.

Lynch:2013:EWO

[1387] Abigail J. Lynch. WINNER: One fish, two fish, where fish for whitefish? Fisheries, 38(8):356, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2013.816680.

Carmona-Catot:2013:ERT

[1388] Gerard Carmona-Catot. RUNNER UP (Tied) fish go wild in California. Fisheries, 38(8):357, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2013.816681.

Cooney:2013:ERT

[1389] Patrick Cooney. RUNNER UP (Tied): a southern revival: Researchers and young anglers contribute to the revival of Southern Appalachian trout fishing. Fisheries, 38(8):358, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.816682.

Claussen:2013:SCD

[1390] Julie E. Claussen, Patrick B. Cooney, Julie M. Defilippi, Sarah Gilbert Fox, Sarah Michele Glaser, Elden Hawkes, Clifford Hutt, Marissa H. Jones, Iris M. Kemp, Aaron Lerner, Stephen R. Midway, Shivonne Nesbit, Jeremiah Osborne-Gowey, Ryan Roberts, and Cleve Steward. Science communication in a digital age: Social media and the American Fisheries Society. Fisheries, 38(8):359–362, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.816289.

Ihde:2013:IJL

[1391] Tom Ihde and Howard Townsend. Interview with Jason Link: Champion for ecosystem science and management. Fisheries, 38(8):363– 369, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.813487.

Anonymous:2013:PCA

[1392] Anonymous. Preparing for the challenges ahead. Fisheries, 38(8): 370–381, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.817272.

Midway:2013:MCD

[1393] Steve Midway and Patrick Cooney. Membership and communication: The dual benefits of social media for AFS. Fisheries, 38(8): 382–383, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.813488.

Anonymous:2013:JAAb

[1394] Anonymous. Journal of Aquatic Animal Health, volume 25, number 2, June 2013. Fisheries, 38(8):384, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.830018.

Anonymous:2013:FEh

[1395] Anonymous. Fisheries events. Fisheries, 38(8):387, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.830020.

Anonymous:2013:FIPi

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.839439.

Hughes:2013:CBM

[1397] Bob Hughes. A call for better mining regulations. Fisheries, 38 (9):391, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.826203.

Love:2013:LE

[1398] Milton Love and Jim Reynolds. Letters to the Editor. Fisheries, 38(9):392, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.828041.

Slacum:2013:ATC

[1399] Ward Slacum and Lee Benaka. AFS talks climate change and fisheries on Capitol Hill. Fisheries, 38(9):393–394, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.828043.

Bowzer:2013:GST

[1400] John C. Bowzer and Jesse T. Trushenski. Getting in the swim of things. Fisheries, 38(9):396–398, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.828045.

Bigford:2013:CHD

[1401] Thomas E. Bigford. Connecting the habitat dots: Perceptions and expectations. Fisheries, 38(9):399, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.828040.

Jackson:2013:T

[1402] Donald C. Jackson. The twitch. Fisheries, 38(9):400–401, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.828044.

Hershberger:2013:IDF

[1403] Paul Hershberger, Linda Rhodes, Gael Kurath, and James Winton. Infectious diseases of fishes in the Salish Sea. Fisheries, 38(9):402– 409, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.826202.

Grubisic:2013:MGI

[1404] Leon Grubisić, Tanja Segvić-Bubić, Ivana Lepen Pleić, Krstina Mislov-Jelavić, Vjeko Ticina, Ivan Katavić, and Ivona Mladineo. Morphological and genetic identification of spontaneously spawned larvae of captive bluefin tuna in the Adriatic Sea. Fisheries, 38(9):410–417, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.826201.

Anonymous:2013:MS

[1405] Anonymous. Member spotlight. Fisheries, 38(9):418–419, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.828282.

Anonymous:2013:FCPa

[1406] Anonymous. First call for papers: Québec City 2014. Fisheries, 38 (9):420–422, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.828046.

Anonymous:2013:NFF

[1407] Anonymous. New FAO fisheries reports. Fisheries, 38(9):423– 424, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.829721.

Anonymous:2013:NAJe

[1408] Anonymous. North American Journal of Fisheries Management, volume 33, number 4, August 2013. Fisheries, 38(9):425, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.839520.

Anonymous:2013:FEi

[1409] Anonymous. Fisheries events. Fisheries, 38(9):427, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.828283.

Anonymous:2013:FIPj

[1410] Anonymous. Full issue PDF volume 38, issue 10. Fisheries, 38(10): 429–476, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.851552.

Hughes:2013:WMD

[1411] Bob Hughes. We must do better — we have to do better. Fisheries, 38(10):431, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.833506.

Bigford:2013:SN

[1412] Thomas E. Bigford. Scary numbers. Fisheries, 38(10):432, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.836502.

Kopaska:2013:HT

[1413] Jeff Kopaska. Hiring tools. Fisheries, 38(10):433, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2013.836503.

OKeefe:2013:FPA

[1414] Catherine E. O’Keefe and Gregory R. DeCelles. Forming a partnership to avoid bycatch. Fisheries, 38(10):434–444, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.838122.

Tsehaye:2013:PFI

[1415] Iyob Tsehaye, Matthew Catalano, Greg Sass, David Glover, and Brian Roth. Prospects for fishery-induced collapse of invasive Asian carp in the Illinois River. Fisheries, 38(10):445–454, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.836501.

Gutowsky:2013:SDT

[1416] Lee F. G. Gutowsky, Jenilee Gobin, Nicholas J. Burnett, Jacqueline M. Chapman, Lauren J. Stoot, and Shireen Bliss. Smartphones and digital tablets: Emerging tools for fisheries professionals. Fisheries, 38 (10):455–461, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.838133.

Long:2013:CEM

[1417] James M. Long. Conservation, ecology, and management of catfish: The Second International Symposium. Fisheries, 38(10):462, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.833508.

Link:2013:EAF

[1418] Jason S. Link. Ecosystem approaches to fisheries: a global perspective. Fisheries, 38(10):463, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.851556.

Emlen:2013:EBM

[1419] John M. Emlen. Ecosystem-based management for marine fisheries: an evolving perspective. Fisheries, 38(10):464, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.851558.

Anonymous:2013:RBH

[1420] Anonymous. Robert “Bob” L. Hunt. Fisheries, 38(10):465, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.828281.

Anonymous:2013:SCP

[1421] Anonymous. Second call for papers: Québec City 2014. Fisheries, 38 (10):466–468, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.836504.

Austen:2013:SSF

[1422] Doug Austen. Sound science and future trends. Fisheries, 38(10): 469, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.838138.

Anonymous:2013:TAFc

[1423] Anonymous. Transactions of the American Fisheries Society, volume 142, number 5, September 2013. Fisheries, 38(10):470, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.851562.

Anonymous:2013:NAMe

[1424] Anonymous. New AFS Members. Fisheries, 38(10):471, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.838136.

Anonymous:2013:FEj

[1425] Anonymous. Fisheries events. Fisheries, 38(10):473, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.851563.

Anonymous:2013:FIPk

[1426] Anonymous. Full issue PDF volume 38, issue 11. Fisheries, 38(11): 477–524, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.859467.

Hughes:2013:TAG

[1427] Bob Hughes. Thinking and acting globally and locally. Fisheries, 38 (11):479, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.848348.

Riecke:2013:FFP

[1428] Dennis K. Riecke, Kristen H. Ferry, Jill M. Hardiman, Robert M. Hughes, Cynthia S. Kolar, Philip Moy, Donna L. Parrish, Gregory D. Pitchford, and Kirk Schroeder. Federal funding for programs to prevent, control, and manage aquatic invasive species. Fisheries, 38(11):480, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.797298.

Kaemingk:2013:IUS

[1429] Mark A. Kaemingk, Daniel J. Dembkowski, Hilary A. Meyer, and Larry M. Gigliotti. Insights for undergraduates seeking an advanced degree in wildlife and fisheries sciences. Fisheries, 38(11): 483–490, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.826200.

Eder:2013:UGI

[1430] Brandon L. Eder and Ben C. Neely. Use of geographic information systems by fisheries management agencies. Fisheries, 38(11): 491–496, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.848344.

Hutchings:2013:GCF

[1431] Jeffrey A. Hutchings and John R. Post. Gutting Canada’s Fisheries Act: No fishery, no fish habitat protection. Fisheries, 38(11): 497–501, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.848345.

Pascoe:2013:RES

[1432] Sean Pascoe, Cathy M. Dichmont, Simon Vieira, Tom Kompas, Rik C. Buckworth, and David Carter. A retrospective evaluation of sustainable yields for Australia’s northern prawn fishery: an alternative view. Fisheries, 38(11):502–508, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.848342.

Wang:2013:RCP

[1433] You-Gan Wang and Na Wang. Rejoinder to comment paper. Fisheries, 38(11):509, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.848346.

Chenoweth:2013:TLH

[1434] Ellen Chenoweth. Trend lines to headlines: a scientist’s foray into journalism. Fisheries, 38(11):510–511, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.848401.

Austen:2013:RGD

[1435] Doug Austen and Bob Hughes. A rough guide to decision making at AFS — first steps in streamlining the process. Fisheries, 38 (11):512, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.850000.

Bigford:2013:SEW

[1436] Thomas E. Bigford. Sharks eat what? Fisheries, 38(11):513, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.848349.

Kopaska:2013:FT

[1437] Jeff Kopaska. Fishing tournaments. Fisheries, 38(11):514, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.848350.

Anonymous:2013:TCP

[1438] Anonymous. Third call for papers: Québec City 2014. Fisheries, 38 (11):515–517, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.850003.

Inkley:2013:NRA

[1439] Doug Inkley. New report addresses climate change and freshwater fish. Fisheries, 38(11):518, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.848402.

Anonymous:2013:NAMf

[1440] Anonymous. New AFS Members. Fisheries, 38(11):520, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.859472.

Anonymous:2013:NAJf

[1441] Anonymous. North American Journal of Aquaculture, volume 75, number 4, October 2013. Fisheries, 38(11):521, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.859473.

Anonymous:2013:FEk

[1442] Anonymous. Fisheries events. Fisheries, 38(11):522, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.859474.

Anonymous:2013:FIPl

[1443] Anonymous. Full issue PDF volume 38, issue 12. Fisheries, 38(12): 525–584, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2013.873249.

Hughes:2013:WAW

[1444] Bob Hughes. We are the AFS, we speak for the fish. Fisheries, 38 (12):527, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.859608.

Bodine:2013:CST

[1445] Kristopher A. Bodine, Daniel E. Shoup, Jason Olive, Zachary L. Ford, Rebecca Krogman, and Tyler J. Stubbs. Catfish sampling techniques: Where we are now and where we should go. Fisheries, 38 (12):529–546, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.848343.

Austen:2013:VBW

[1446] Doug Austen. The view from Bethesda — week #6. Fisheries, 38 (12):547, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.859607.

Bigford:2013:CSC

[1447] Thomas E. Bigford. Communicating successes and convincing others. Fisheries, 38(12):548, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.859603.

Bowker:2013:FDQ

[1448] Jim Bowker and Jesse Trushenski. Fish drug questions answered by the FDA. Fisheries, 38(12):549–552, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.863091.

Schaeffer:2013:AMW

[1449] Jeff Schaeffer. 143rd Annual Meeting wrap up: Thank you Little Rock. Fisheries, 38(12):553–567, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.859975.

Peterson:2013:FFS

[1450] Lisa Peterson. The 2012–2013 Fenske Fellowship: Structured decision making in dam management. Fisheries, 38(12):568–569, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.859612.

Anonymous:2013:SBMa

[1451] Anonymous. The Steven Berkeley Marine Conservation Fellowship winners. Fisheries, 38(12):570, 2013. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.859614.

Anonymous:2013:SBMb

[1452] Anonymous. The Steven Berkeley Marine Conservation Fellowship. Fisheries, 38(12):571, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.873250.

Anonymous:2013:CAN

[1453] Anonymous. Call for award nominations: 2014 American Fisheries Society awards. Fisheries, 38(12):572–574, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.859610.

Anonymous:2013:AFSb

[1454] Anonymous. American Fisheries Society section awards for 2013. Fisheries, 38(12):575, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.859611.

Anonymous:2013:FCPb

[1455] Anonymous. Fourth call for papers: Québec City 2014. Fisheries, 38 (12):576–578, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.873251.

Anonymous:2013:FV

[1456] Anonymous. Fisheries volume 38 2013. Fisheries, 38(12):579, 2013. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.859615.

Trushenski:2013:ASC

[1457] Jesse Trushenski. AFS seeks comment on mining and oil and gas extraction policy. Fisheries, 38(12):580, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.873252.

Anonymous:2013:JAAc

[1458] Anonymous. Journal of Aquatic Animal Health, volume 25, number 3, September 2013. Fisheries, 38(12):581, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2013.873253.

Anonymous:2013:FEl

[1459] Anonymous. Fisheries events. Fisheries, 38(12):582, 2013. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.859609.

Anonymous:2014:FIPa

[1460] Anonymous. Full issue PDF volume 39, issue 1. Fisheries, 39 (1):1–48, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.881154.

Hughes:2014:RVN

[1461] Bob Hughes. Reductionism versus natural rights or why is effective natural resource management so difficult? Fisheries, 39(1):3, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.865607.

Austen:2014:SA

[1462] Doug Austen. Situational awareness. Fisheries, 39(1):4, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.866492.

Bigford:2014:WAL

[1463] Thomas E. Bigford. We’re all leading the way. Fisheries, 39 (1):5, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.865608.

Kopaska:2014:UF

[1464] Jeff Kopaska. Urban fishing. Fisheries, 39(1):6, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.865609.

Osborne-Gowey:2014:SSM

[1465] Jeremiah Osborne-Gowey. Starting the social media science communicationonline and in fisheries. Fisheries, 39(1):7, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.866510.

Anonymous:2014:ANa

[1466] Anonymous. AFS news. Fisheries, 39(1):8–9, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.865611.

Anonymous:2014:CSC

[1467] Anonymous. Cross-sectoral conference to sustain livelihoods, food security, and aquatic ecosystems. Fisheries, 39(1):10–11, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.865610.

Marsie-Hazen:2014:DF

[1468] Rahel Marsie-Hazen. Diversity in fisheries. Fisheries, 39(1):12– 17, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.865612.

Hughes:2014:RUWa

[1469] Robert M. Hughes, Susie Dunham, Kathleen G. Maas-Hebner, J. Alan Yeakley, Carl Schreck, Michael Harte, Nancy Molina, Clinton C. Shock, Victor W. Kaczynski, and Jeff Schaeffer. A review of urban water body challenges and approaches: (1) rehabilitation and remediation. Fisheries, 39(1):18–29, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.836500.

Hughes:2014:RUWb

[1470] Robert M. Hughes, Susie Dunham, Kathleen G. Maas-Hebner, J. Alan Yeakley, Michael Harte, Nancy Molina, Clinton C. Shock, and Victor W. Kaczynski. A review of urban water body challenges and approaches: (2) mitigating effects of future urbanization. Fisheries, 39 (1):30–40, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.866507.

Ortiz:2014:CMF

[1471] Pedro B. Ortiz. 21st century metropolises and fisheries. Fisheries, 39(1):41, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.833507.

Anonymous:2014:NAJa

[1472] Anonymous. North American Journal of Fisheries Management, volume 33, number 6, December 2013. Fisheries, 39(1):42, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.881157.

Anonymous:2014:FEa

[1473] Anonymous. Fisheries events. Fisheries, 39(1):44, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.881158.

Anonymous:2014:FIPb

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.890488.

Hughes:2014:PSS

[1475] Robert Hughes. A plea for a steady-state economy. Fisheries, 39 (2):51, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.874528.

Austen:2014:AMM

[1476] Doug Austen and Bob Hughes. Annual meetings and money. Fisheries, 39(2):52, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.875004.

Bigford:2014:GE

[1477] Thomas E. Bigford. Great expectations. Fisheries, 39(2):53, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.875005.

Kopaska:2014:SF

[1478] Jeff Kopaska. Smartphones in fisheries. Fisheries, 39(2):54, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2013.875006.

Osborne-Gowey:2014:WSM

[1479] Jeremiah Osborne-Gowey. What is social media. Fisheries, 39 (2):55, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.876883.

Harrell:2014:SVP

[1480] Reginal M. Harrell and Steve L. McMullin. Second Vice President. Fisheries, 39(2):56–57, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.877274.

Anonymous:2014:ANb

[1481] Anonymous. AFS news. Fisheries, 39(2):58–59, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2013.875007.

Trushenski:2014:FMG

[1482] Jesse Trushenski. Fish Meat gives viewers plenty to digest. Fisheries, 39(2):60, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.876887.

Anonymous:2014:WFM

[1483] Anonymous. World Fish Migration Day. Fisheries, 39(2):61, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.890494.

Gillespie:2014:FER

[1484] Nathaniel Gillespie, Amy Unthank, Lauren Campbell, Paul Anderson, Robert Gubernick, Mark Weinhold, Daniel Cenderelli, Brian Austin, Daniel McKinley, Susan Wells, Janice Rowan, Curt Orvis, Mark Hudy, Alison Bowden, Amy Singler, Eileen Fretz, Jessica Levine, and Richard Kirn. Flood effects on road-stream crossing infrastructure: Economic and ecological benefits of stream simulation designs. Fisheries, 39 (2):62–76, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2013.874527.

Galbreath:2014:ETR

[1485] Peter F. Galbreath, Michael A. Bisbee, Jr., Douglas W. Dompier, Cory M. Kamphaus, and Todd H. Newsome. Extirpation and tribal reintroduction of Coho salmon to the interior Columbia River Basin. Fisheries, 39(2):77–87, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2013.874526.

Anonymous:2014:QMU

[1486] Anonymous. Québec Meeting update: The Québec Convention Center. Fisheries, 39(2):88–89, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.890495.

Anonymous:2014:TAFa

[1487] Anonymous. Transactions of the American Fisheries Society, volume 143, number 1, January 2014. Fisheries, 39(2):90, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.890496.

Anonymous:2014:FEb

[1488] Anonymous. Fisheries events. Fisheries, 39(2):92, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.890497.

Anonymous:2014:NAMa

[1489] Anonymous. New AFS Members. Fisheries, 39(2):93, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.890498.

Anonymous:2014:C

[1490] Anonymous. Correction. Fisheries, 39(2):94, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.890499.

Anonymous:2014:FIPc

[1491] Anonymous. Full issue PDF volume 39, issue 3. Fisheries, 39(3): 97–144, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.902705.

Hughes:2014:HMP

[1492] Bob Hughes. How many people are enough (too many)? Fisheries, 39(3):99, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.883970.

Austen:2014:FRC

[1493] Doug Austen. The foundational role of chapters. Fisheries, 39 (3):101, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.885261.

Bigford:2014:EMS

[1494] Thomas E. Bigford. Ecosystem modeling to support fishery management. Fisheries, 39(3):102, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.883971.

Kopaska:2014:TCF

[1495] Jeff Kopaska. Tablet computers in fisheries. Fisheries, 39(3): 103, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.883972.

Osborne-Gowey:2014:WDY

[1496] Jeremiah Osborne-Gowey. What to do with your New Twitter account (or Facebook, or …). Fisheries, 39(3):104, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.883973.

Lomnicky:2014:WQS

[1497] Gregg A. Lomnicky, Robert H. Gray, and John W. Meldrim. Water quality section introduction and history. Fisheries, 39(3):105–106, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.885302.

Anonymous:2014:ANc

[1498] Anonymous. AFS news. Fisheries, 39(3):107, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.885307.

Brown:2014:UBF

[1499] Richard S. Brown, Alison H. Colotelo, Brett D. Pflugrath, Craig A. Boys, Lee J. Baumgartner, Z. Daniel Deng, Luiz G. M. Silva, Colin J. Brauner, Martin Mallen-Cooper, Oudom Phonekhampeng, Garry Thorncraft, and Douangkham Singhanouvong. Understanding barotrauma in fish passing hydro structures: a global strategy for sustainable development of water resources. Fisheries, 39(3):108–122, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.883570.

Claramunt:2014:RDA

[1500] Randall M. Claramunt and David F. Clapp. Response to Dettmers et al. (2012): Great Lakes Fisheries Managers Are Pursuing Appropriate Goals. Fisheries, 39(3):123–125, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.883968. See [1266, 1501].

Dettmers:2014:CWD

[1501] John M. Dettmers, Christopher I. Goddard, and Kelley D. Smith. Considerations when determining appropriate management goals: a reply to Claramunt and Clapp. Fisheries, 39(3):126–127, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.883969.

See [1500].

Anonymous:2014:MRJ

[1502] Anonymous. In memoriam: [Robert J. Behnke, Ph.D. 1929–2013]. Fisheries, 39(3):128–132, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.883571.

Anonymous:2014:KLW

[1503] Anonymous. Killed in the line of work. Fisheries, 39(3):133– 134, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.902706.

Anonymous:2014:GIFa

[1504] Anonymous. Global Inland Fisheries Conference: Theme 1 — biological assessment. Fisheries, 39(3):135–136, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.885368.

Anonymous:2014:MUE

[1505] Anonymous. Meeting update: Exploring Québec City. Fisheries, 39 (3):137–138, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.902707.

Anonymous:2014:NAJb

[1506] Anonymous. North American Journal of Aquaculture, volume 76, number 1, January 2014. Fisheries, 39(3):139, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.902708.

Anonymous:2014:FEc

[1507] Anonymous. Fisheries events. Fisheries, 39(3):142, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.902709.

Anonymous:2014:NAMb

[1508] Anonymous. New AFS Members. Fisheries, 39(3):143, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.902710.

Anonymous:2014:FIPd

[1509] Anonymous. Full issue PDF volume 39, issue 4. Fisheries, 39(4): 145–192, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.911032.

Hughes:2014:ITF

[1510] Bob Hughes. Iron triangles and fisheries. Fisheries, 39(4):147, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.891990.

Austen:2014:VPC

[1511] Doug Austen. The value proposition challenge of AFS. Fisheries, 39(4):149, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.897232.

Bigford:2014:ABC

[1512] Thomas E. Bigford. Aquatic blue carbon. Fisheries, 39(4):150, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.897231.

Litts:2014:RX

[1513] Thom Litts and Jeff Kopaska. RIP XP. Fisheries, 39(4):151, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.893871.

Osborne-Gowey:2014:NTM

[1514] Jeremiah Osborne-Gowey. New toy mania and manageable bytes. Fisheries, 39(4):152, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.893869.

Anonymous:2014:GIFb

[1515] Anonymous. Global Inland Fisheries Conference: Theme 2 — economic and social assessment. Fisheries, 39(4):153, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.893872.

Larsen:2014:MTM

[1516] Andrea M. Larsen and Cova R. Arias. More than mucus: The hidden world of the fish microbiota. Fisheries, 39(4):154, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.893870.

Deweber:2014:IUL

[1517] Jefferson Tyrell Deweber, Yin-Phan Tsang, Damon M. Krueger, Joanna B. Whittier, Tyler Wagner, Dana M. Infante, and Gary Whelan. Importance of understanding landscape biases in USGS gage locations: Implications and solutions for managers. Fisheries, 39(4): 155–163, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.891503.

Hogan:2014:SEE

[1518] Timothy W. Hogan, Glenn F. Cada, and Stephen V. Amaral. The status of environmentally enhanced hydropower turbines. Fisheries, 39 (4):164–172, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.897195.

Patrick:2014:VIL

[1519] Wesley S. Patrick, Jason M. Cope, and James T. Thorson. Validating and improving life history data in FishBase. Fisheries, 39(4): 173–176, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.891504.

Anonymous:2014:MU

[1520] Anonymous. Meeting update. Fisheries, 39(4):177–180, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.897163.

Anonymous:2014:ANd

[1521] Anonymous. AFS news. Fisheries, 39(4):181–183, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.911037.

Bettoli:2014:BRB

[1522] Phillip W. Bettoli. Book review: Biology, Management, and Culture of Walleye, edited by Bruce A. Barton. Fisheries, 39(4):184, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.911038.

Bai:2014:KFF

[1523] Sungchul C. Bai. Korean Federation of Fisheries Science and Technologies Annual Meeting. Fisheries, 39(4):185–186, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.911040.

Anonymous:2014:NAJc

[1524] Anonymous. North American Journal of Fisheries Management volume 34, number 1, February 2014. Fisheries, 39(4):188, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.911042.

Anonymous:2014:NAMc

[1525] Anonymous. New AFS Members. Fisheries, 39(4):189, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.911043.

Anonymous:2014:FEd

[1526] Anonymous. Fisheries events. Fisheries, 39(4):191, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.911044.

Anonymous:2014:FIPe

[1527] Anonymous. Full issue PDF volume 39, issue 5. Fisheries, 39(5): 193–240, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.921475.

Hughes:2014:FEW

[1528] Bob Hughes. Fisheries ethics, or what do you want to do with your scientific knowledge in addition to earning a living? Fisheries, 39 (5):195, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.903842.

Bigford:2014:SWA

[1529] Thomas E. Bigford. Shall we advocate? Fisheries, 39(5):196, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.905473.

Anonymous:2014:J

[1530] Anonymous. In the journals. Fisheries, 39(5):197–198, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.907154.

Moffitt:2014:FEH

[1531] Sarah Moffitt. Fisheries ecosystems and hypoxia: a future informed by the past. Fisheries, 39(5):199–200, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2013.875008.

Drake:2014:ERL

[1532] D. Andrew R. Drake and Nicholas E. Mandrak. Ecological risk of live bait fisheries: a new angle on selective fishing. Fisheries, 39(5): 201–211, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.903835.

Arlinghaus:2014:CRE

[1533] Robert Arlinghaus. Are current research evaluation metrics causing a tragedy of the scientific commons and the extinction of universitybased fisheries programs? Fisheries, 39(5):212–215, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.903837.

Anonymous:2014:GIFc

[1534] Anonymous. Global Inland Fisheries Conference: Call for abstracts. Fisheries, 39(5):216, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.903843.

Litts:2014:WOS

[1535] Thom Litts. Why open source GIS software is not for me. Fisheries, 39(5):217, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.903174.

Osborne-Gowey:2014:DCP

[1536] Jeremiah Osborne-Gowey. Developing communication policies that work. Fisheries, 39(5):218, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.907152.

Geist:2014:TDW

[1537] Juergen Geist. Trends and directions in water quality and habitat management in the context of the European Water Framework Directive. Fisheries, 39(5):219–220, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.903838.

Buckel:2014:BDF

[1538] Jeffrey A. Buckel. Billion-dollar fish: The untold story of Alaska pollock. Fisheries, 39(5):221, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.903841.

Naumann:2014:BRO

[1539] Cayla Naumann, James P. W. Robinson, Cameron Freshwater, Eric Hertz, David Stormer, Amy K. Teffer, and Francis Juanes. Book review: Overfishing: What Everyone Needs to Know, by Ray Hilborn with Ulrike Hilborn. Fisheries, 39(5):222, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.903839.

Chittenden:2014:BRT

[1540] Cedar Chittenden. Book review: Telemetry Techniques: a User Guide for Fisheries Research, edited by Noah S. Adams, John W. Beeman, and John H. Eiler. Fisheries, 39(5):223, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.903840.

Anonymous:2014:BKH

[1541] Anonymous. Better know a hatchery. Fisheries, 39(5):224–226, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.903177.

Anonymous:2014:MDJ

[1542] Anonymous. In memoriam [Dr. J. (“Jady”) Frances Allen, Lochie Jo Allen, Walter R. Courtenay, Michael Joseph Mason]. Fisheries, 39 (5):227–229, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.907155.

Anonymous:2014:CEP

[1543] Anonymous. Continuing education program. Fisheries, 39(5):230– 231, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.907153.

Anonymous:2014:TAFb

[1544] Anonymous. Transactions of the American Fisheries Society, volume 143, number 2, March 2014. Fisheries, 39(5):232, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.921478.

Anonymous:2014:NAMd

[1545] Anonymous. New AFS Members. Fisheries, 39(5):233, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.921480.

Anonymous:2014:FEe

[1546] Anonymous. Fisheries events. Fisheries, 39(5):234, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.921479.

Austen:2014:BTF

[1547] Doug Austen. Bringing together fish and wildlife in 2017. Fisheries, 39(5):235, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.907157.

Anonymous:2014:HVJ

[1548] Anonymous. Hughes visits JSFS Spring Meeting. Fisheries, 39 (5):237, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.921481.

Anonymous:2014:FIPf

[1549] Anonymous. Full issue PDF volume 39, issue 6. Fisheries, 39(6): 241–288, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.931119.

Hughes:2014:MGY

[1550] Bob Hughes. Monitoring: Garbage in yields garbage out. Fisheries, 39(6):243, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.915813.

Bigford:2014:MJA

[1551] Thomas E. Bigford. Maybe it’s not just about the fish. Fisheries, 39(6):245, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.914926.

Siitari:2014:IFF

[1552] Kiira Siitari, Jim Martin, and William W. Taylor. Information flow in fisheries management: Systemic distortion within agency hierarchies. Fisheries, 39(6):246–250, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.915814.

Brownscombe:2014:CRF

[1553] Jacob W. Brownscombe, Shannon D. Bower, William Bowden, Liane Nowell, Jonathan D. Midwood, Neville Johnson, and Steven J. Cooke. Canadian recreational fisheries: 35 years of social, biological, and economic dynamics from a national survey. Fisheries, 39(6):251– 260, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.915811.

Matthias:2014:HSI

[1554] Bryan G. Matthias, Micheal S. Allen, Robert N. M. Ahrens, T. Douglas Beard, Jr., and Janice A. Kerns. Hide and seek: Interplay of fish and anglers influences spatial fisheries management. Fisheries, 39 (6):261–269, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.903836.

Weyl:2014:TEF

[1555] Olaf L. F. Weyl, Brian Finlayson, N. Dean Impson, Darragh J. Woodford, and Jarle Steinkjer. Threatened endemic fishes in South Africa’s Cape Floristic Region: a new beginning for the Rondegat River. Fisheries, 39(6):270–279, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.914924.

Anonymous:2014:CMF

[1556] Anonymous. Carlos M. Fetterolf, Jr. 1926–2014. Fisheries, 39(6): 280, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.915812.

Murphy:2014:DIF

[1557] Margaret H. Murphy. Dam impacts on fishery resources — join us in Québec. Fisheries, 39(6):281, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.916551.

Anonymous:2014:DM

[1558] Anonymous. Draft 1, May 2014. Fisheries, 39(6):282–284, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.931121.

Anonymous:2014:FEf

[1559] Anonymous. Fisheries events. Fisheries, 39(6):285, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.931122.

Austen:2014:BSM

[1560] Doug Austen. Behind the scenes at Mazatlan. Fisheries, 39(6): 286, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.922346.

Anonymous:2014:FIPg

[1561] Anonymous. Full issue PDF volume 39, issue 7. Fisheries, 39(7): 289–336, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.939930.

Hughes:2014:SVR

[1562] Bob Hughes. The socioeconomic values of recreational fishing. Fisheries, 39(7):291, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.926894.

Bigford:2014:LSW

[1563] Thomas E. Bigford. Leadership styles we can appreciate. Fisheries, 39(7):293, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.926893.

Utz:2014:WOL

[1564] Ryan M. Utz. Are we overlooking landscape-scale threats to common freshwater fishes? Fisheries, 39(7):294–297, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.923768.

Busch:2014:CAR

[1565] Wolf-Dieter N. Busch and David P. Braun. A case for accelerated reestablishment of American eel in the Lake Ontario and Champlain watersheds. Fisheries, 39(7):298–304, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.923769.

Schuett:2014:AMV

[1566] Michael A. Schuett, Gerard T. Kyle, Jeremy Leitz, Ken Kurzawski, and Kyunghee Lee. Anglers’ motivations for volunteering with fishing or conservation organizations. Fisheries, 39(7):305–311, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.924407.

Anonymous:2014:QEL

[1567] Anonymous. Q&A: The 2014 emerging leaders. Fisheries, 39(7): 312–315, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.925450.

Sass:2014:QBE

[1568] Greg G. Sass and Micheal S. Allen. Q&A: Book editors; Foundations of Fisheries Science. Fisheries, 39(7):316–317, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.925448.

Anonymous:2014:SSM

[1569] Anonymous. Science and social media: How, what, and when to share. Fisheries, 39(7):318, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.923770.

Kopaska:2014:DFA

[1570] Jeff Kopaska. Drones — a fisheries assessment tool? Fisheries, 39 (7):319, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.923771.

Anonymous:2014:MNG

[1571] Anonymous. In memoriam: [Norman G. Sharber, 1925–2013]. Fisheries, 39(7):320, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.923772.

Anonymous:2014:GCIa

[1572] Anonymous. Global conference on inland fisheries: Theme 3 — drivers and synergies. Fisheries, 39(7):321, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.925449.

Anonymous:2014:NWL

[1573] Anonymous. National Workshop on Large Landscape Conservation. Fisheries, 39(7):322, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.928109.

Anonymous:2014:SG

[1574] Anonymous. Schedule at a glance. Fisheries, 39(7):323–326, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.928110.

Anonymous:2014:TYO

[1575] Anonymous. Thank you to our generous sponsors! Fisheries, 39 (7):327, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.939924.

Anonymous:2014:PSA

[1576] Anonymous. Plenary speakers’ abstracts for the 144th Annual Meeting. Fisheries, 39(7):328, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.939925.

Anonymous:2014:WSF

[1577] Anonymous. A. E. Wood State Fish Hatchery, San Marcos, Texas. Fisheries, 39(7):329–330, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.903176.

Austen:2014:SC

[1578] Doug Austen. Scaling up conservation. Fisheries, 39(7):331, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.926892.

Anonymous:2014:NAJd

[1579] Anonymous. North American Journal of Aquaculture volume 76, number 2, April 2014; Journal of Aquatic Animal Health volume 26, number 2, June 2014. Fisheries, 39(7):332, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.939927.

Anonymous:2014:FEg

[1580] Anonymous. Fisheries events. Fisheries, 39(7):333, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.939928.

Anonymous:2014:NAMe

[1581] Anonymous. New AFS Members. Fisheries, 39(7):334, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.939929.

Anonymous:2014:FIPh

[1582] Anonymous. Full issue PDF volume 39, issue 8. Fisheries, 39(8): 337–388, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.946330.

Hughes:2014:LGW

[1583] Bob Hughes. Livestock grazing in the West: Sacred cows at the public trough revisited. Fisheries, 39(8):339, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.932775.

Bigford:2014:FSI

[1584] Thomas E. Bigford. Forage species and issues. Fisheries, 39(8): 340, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.933425.

Kuehne:2014:ECM

[1585] Lauren M. Kuehne and Julian D. Olden. Ecology and conservation of mudminnow species worldwide. Fisheries, 39(8):341–351, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.933318.

Franssen:2014:FCR

[1586] Nathan R. Franssen, Jason E. Davis, Dale W. Ryden, and Keith B. Gido. Fish community responses to mechanical removal of nonnative fishes in a large southwestern river. Fisheries, 39(8):352–363, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.924409.

Edelstein:2014:RAM

[1587] Bailey Edelstein. Rachel Anne Morrison 1987–2014. Fisheries, 39 (8):364, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.933427.

Anonymous:2014:AR

[1588] Anonymous. 2013 Annual Report. Fisheries, 39(8):365–377, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.933426.

Trushenski:2014:QHA

[1589] Jesse T. Trushenski and James D. Bowker. Q&A: Halamid(R) aqua (chloramine-t) approved by FDA to treat fish diseases-what it means for fisheries. Fisheries, 39(8):378–379, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.924929.

Moore:2014:IWT

[1590] Stephen E. Moore, Daniel J. Schill, and Robert F. Carline. The International Wild Trout Symposium: The past and the future. Fisheries, 39(8):380–382, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.934641.

Austen:2014:OPA

[1591] Doug Austen. The oyster and the pearl: The Annual Report and your input. Fisheries, 39(8):383, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.933428.

Anonymous:2014:NAJe

[1592] Anonymous. North American Journal of Fisheries Management, volume 34, number 3, June 2014. Fisheries, 39(8):384, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.946333.

Anonymous:2014:FEh

[1593] Anonymous. Fisheries events. Fisheries, 39(8):385, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.946334.

Anonymous:2014:FIPi

[1594] Anonymous. Full issue PDF volume 39, issue 9. Fisheries, 39(9): 389–436, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.961858.

Allison:2014:USR

[1595] Jordan Allison. Units, step right up: Get your free Website. Fisheries, 39(9):391, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.940079.

Parrish:2014:RAL

[1596] Donna Parrish. Revisiting the AFS lexicon. Fisheries, 39(9): 393, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.944974.

Bigford:2014:EIW

[1597] Thomas E. Bigford. Effects of industrial water intake structures. Fisheries, 39(9):394, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.943743.

Shiffman:2014:ASP

[1598] David Samuel Shiffman and Neil Hammerschlag. An assessment of the scale, practices, and conservation implications of Florida’s charter boatbased recreational shark fishery. Fisheries, 39(9):395–407, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.941439.

Murauskas:2014:TEF

[1599] Joshua G. Murauskas, Jeffery K. Fryer, Bryan Nordlund, and Joseph L. Miller. Trapping effects and fisheries research: a case study of sockeye salmon in the Wenatchee River, USA. Fisheries, 39(9): 408–414, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.943366.

Jenkins:2014:GUF

[1600] Jill A. Jenkins, Henry L. Bart, Jr., James D. Bowker, Paul R. Bowser, J. Randy MacMillan, John G. Nickum, Joseph W. Rachlin, James D. Rose, Peter W. Sorensen, Barbara E. Warkentine, and Greg W. Whitledge. Guidelines for Use of Fishes in Research — Revised and Expanded, 2014. Fisheries, 39(9):415–416, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.924408.

McManamay:2014:OAD

[1601] Ryan A. McManamay and Ryan M. Utz. Open-access databases as unprecedented resources and drivers of cultural change in fisheries science. Fisheries, 39(9):417–425, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.946128.

Anonymous:2014:GCIb

[1602] Anonymous. Global conference on inland fisheries: Theme 4 — policy and governance. Fisheries, 39(9):426, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.938156.

Anonymous:2014:AMAa

[1603] Anonymous. 145th Annual Meeting of the American Fisheries Society: First call for papers. Fisheries, 39(9):427–428, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.943368.

Anonymous:2014:QBE

[1604] Anonymous. Q&A: Book Editors. Fisheries, 39(9):429, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.946131.

Anonymous:2014:DS

[1605] Anonymous. Denise Spencer. Fisheries, 39(9):430, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.946129.

Austen:2014:FIB

[1606] Doug Austen. Fisheries inside the Beltway. Fisheries, 39(9): 432, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.946130.

Anonymous:2014:TAFc

[1607] Anonymous. Transactions of the American Fisheries Society, volume 143, number 4, July 2014. Fisheries, 39(9):433, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.961862.

Anonymous:2014:FEi

[1608] Anonymous. Fisheries events. Fisheries, 39(9):435, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.961863.

Anonymous:2014:TWD

[1609] Anonymous. The things we do for science. Fisheries, 39(9): 436, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.941058.

Anonymous:2014:FIPj

[1610] Anonymous. Full issue PDF volume 39, issue 10. Fisheries, 39(10): 437–480, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.978726.

Parrish:2014:ACA

[1611] Donna Parrish. AFS communications for the ages. Fisheries, 39 (10):439, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.955570.

Bigford:2014:WSD

[1612] Thomas E. Bigford. When is sublethal deadly? Fisheries, 39 (10):440, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.955091.

Goulette:2014:OAT

[1613] Graham S. Goulette, James P. Hawkes, John F. Kocik, James P. Manning, Paul A. Music, John P. Wallinga, and Gayle Barbin Zydlewski. Opportunistic acoustic telemetry platforms: Benefits of collaboration in the Gulf of Maine. Fisheries, 39(10):441–450, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.943740.

Pracheil:2014:UHP

[1614] Brenda M. Pracheil, J. Derek Hogan, John Lyons, and Peter B. McIntyre. Using hard-part microchemistry to advance conservation and management of North American freshwater fishes. Fisheries, 39(10): 451–465, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.937858.

Fink:2014:RSE

[1615] Wendy Fink, Daniel Edge, and Eric Hallerman. A roadmap for science, education, and outreach for natural resources. Fisheries, 39(10): 466–467, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.960332.

Gonsalves:2014:GGF

[1616] Lonnie Gonsalves, Marybeth Brey, and Cecilia Lewis. The Gullah/ Geechee Fishing Association. Fisheries, 39(10):468–470, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.954038.

Chandler:2014:RYA

[1617] Jim Chandler. A recap of this year’s annual meeting in Idaho Falls, Idaho. Fisheries, 39(10):471, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.941057.

Anonymous:2014:GCIc

[1618] Anonymous. Global Conference on Inland Fisheries: Bringing a new international focus. Fisheries, 39(10):472, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.960755.

Anonymous:2014:AMAb

[1619] Anonymous. 145th Annual Meeting of the American Fisheries Society: Second call for papers http://fisheries.org/meetings. Fisheries, 39(10):473, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.973319.

Anonymous:2014:MBF

[1620] Anonymous. In memoriam: Billy Frank, Jr. Fisheries, 39(10): 474, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.955094.

Austen:2014:FNN

[1621] Doug Austen. Fish have no nationalities, but harvest does. Fisheries, 39(10):475–476, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.960776.

Anonymous:2014:NAJf

[1622] Anonymous. North American Journal of Fisheries Management, volume 34, number 5, October 2014. Fisheries, 39(10):477, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.973320.

Anonymous:2014:DFS

[1623] Anonymous. A deckhand “Fester” story. Fisheries, 39(10):478, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.955571.

Anonymous:2014:FEj

[1624] Anonymous. Fisheries events. Fisheries, 39(10):480, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.973322.

Anonymous:2014:FIPk

[1625] Anonymous. Full issue PDF volume 39, issue 11. Fisheries, 39(11): 481–572, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.988555.

Parrish:2014:SC

[1626] Donna Parrish. Society by committee. Fisheries, 39(11):483, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.968247.

Bigford:2014:ASP

[1627] Thomas E. Bigford. Aquaculture and the science-policy continuum. Fisheries, 39(11):485, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.964394.

Trushenski:2014:IAT

[1628] Jesse Trushenski and Jim Bowker. Introduction to the aquaculturethemed issue. Fisheries, 39(11):486, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.966093.

Anonymous:2014:AMAc

[1629] Anonymous. 145th Annual Meeting of the American Fisheries Society: Third call for papers 2015.fisheries.org. Fisheries, 39 (11):487, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.985973.

Kline:2014:PRB

[1630] Paul A. Kline and Thomas A. Flagg. Putting the red back in Redfish Lake, 20 years of progress toward saving the Pacific Northwest’s most endangered salmon population. Fisheries, 39(11):488–500, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.966087.

Amos:2014:URR

[1631] Kevin H. Amos, Lori Gustafson, Janet Warg, Janet Whaley, Maureen Purcell, Jill Rolland, James Winton, Kevin Snekvik, Theodore Meyers, Bruce Stewart, John Kerwin, Marilyn Blair, Joel Bader, and Joy Evered. U.S. response to a report of infectious salmon anemia virus in Western North America. Fisheries, 39(11):501–506, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.967348.

Hershberger:2014:RPC

[1632] Paul Hershberger. In response — peer commentary on infectious salmon anemia virus paper. Fisheries, 39(11):507, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.988472.

Pitcher:2014:CAN

[1633] Trevor E. Pitcher. Canadian aquaculture news: Grant to study the value of incorporating wild salmon genes into an organic aquaculture industry Partner’s practice. Fisheries, 39(11):507, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.984560.

Rust:2014:EPM

[1634] Michael B. Rust, Kevin H. Amos, April L. Bagwill, Walton W. Dickhoff, Lorenzo M. Juarez, Carol S. Price, James A. Morris, Jr., and Michael C. Rubino. Environmental performance of marine net-pen aquaculture in the United States. Fisheries, 39(11):508–524, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.966818.

Hart:2014:MEA

[1635] Emily R. Hart, James B. Webb, Craig Hollingsworth, and Andy J. Danylchuk. Managing expectations for aquaponics in the classroom: Enhancing academic learning and teaching an appreciation for aquatic resources. Fisheries, 39(11):525–530, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.966353.

Dembkowski:2014:RUA

[1636] Daniel J. Dembkowski, Landon L. Pierce, and Craig P. Paukert. In response — the use of aquaponics in the classroom. Fisheries, 39 (11):530, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.964397.

Byelashov:2014:FFP

[1637] Oleksandr A. Byelashov and Mark E. Griffin. Fish in, fish out: Perception of sustainability and contribution to public health. Fisheries, 39 (11):531–535, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.967765.

Gothreaux:2014:ALF

[1638] Craig Gothreaux and Patrick Banks. Aquaculture and Louisiana fisheries: Innovative oil spill rehabilitation efforts. Fisheries, 39(11): 536–540, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.966088.

Smith:2014:HHM

[1639] Carlos Smith. Hatcheries and harvest: Meeting treaty obligations through artificial propagation. Fisheries, 39(11):541–542, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.966090.

Trushenski:2014:ACA

[1640] Jesse Trushenski, Lee Blankenship, Jim Bowker, Tom Flagg, Jay Hesse, Ken Leber, Kai Lorenzen, Don MacKinlay, Des Maynard, Christine Moffitt, Vince Mudrak, Kim Scribner, Scott Stuewe, John Sweka, Gary Whelan, and Connie Young-Dubovsky. AFS completes assessment, issues new guidance regarding hatchery operation and the use of hatchery-origin fish. Fisheries, 39(11):543–547, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.964395.

Anonymous:2014:HIN

[1641] Anonymous. HSRG issues new report, “On the Science of Hatcheries … ”. Fisheries, 39(11):548–551, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.964396.

Moffitt:2014:BGF

[1642] Christine M. Moffitt and Lubia Cajas-Cano. Blue growth: The 2014 FAO State of world fisheries and aquaculture. Fisheries, 39(11): 552–553, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.966265.

Amezcua:2014:CSA

[1643] Felipe Amezcua and Martin F. Soto. Current state of aquaculture in México. Fisheries, 39(11):554–555, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.984561.

Lang:2014:RHE

[1644] Tom Lang. The role of hatcheries in ensuring social and economic benefits of fisheries in the United States. Fisheries, 39(11):556–557, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.966091.

Chen:2014:AWQ

[1645] Yushun Chen, Jesse T. Trushenski, and Margaret H. Murphy. Aquaculture and water quality management in the United States. Fisheries, 39 (11):558–561, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.964398.

Montanez:2014:OAR

[1646] José L. Montañez. Offshore aquaculture regulations under the Magnuson–Stevens Fishery Conservation and Management Act. Fisheries, 39(11):562–563, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.978169.

Trushenski:2014:RSR

[1647] Jesse Trushenski. Raising Shrimp raises concerns, offers answers for an increasingly invertivorous public. Fisheries, 39(11):564, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.966092.

Anonymous:2014:LPW

[1648] Anonymous. Little Port Walter Marine Station. Fisheries, 39(11): 565–566, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.903182.

Fayram:2014:ALR

[1649] Andrew Fayram. Aquaculture 101 or … random thoughts on aquaculture. Fisheries, 39(11):567, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.966819.

Anonymous:2014:TAFd

[1650] Anonymous. Transactions of the American Fisheries Society, volume 143, number 5, September 2014. Fisheries, 39(11):570, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.985975.

Anonymous:2014:FEk

[1651] Anonymous. Fisheries events. Fisheries, 39(11):571, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.985972.

Love:2014:DYW

[1652] Milton Love. Do you want hash browns with those? Fisheries, 39 (11):572, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.967349.

Anonymous:2014:FIPl

[1653] Anonymous. Full issue PDF volume 39, issue 12. Fisheries, 39(12): 573–624, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2014.997596.

Isaak:2014:BFS

[1654] Dan Isaak. Blogging fish science. Fisheries, 39(12):575, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.979916.

Bigford:2014:RHC

[1655] Thomas E. Bigford. Revisiting hopes for created and restored habitats. Fisheries, 39(12):576, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.976859.

Anonymous:2014:AMAd

[1656] Anonymous. 145th Annual Meeting of the American Fisheries Society: Final call for papers 2015.fisheries.org. Fisheries, 39 (12):577, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.992724.

Miralles:2014:PSB

[1657] Laura Miralles, Francis Juanes, Antonio F. Pardiñas, and Eva Garcia-Vazquez. Paleoclimate shaped bluefish structure in the Northern Hemisphere. Fisheries, 39(12):578–586, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.976701.

Beard:2014:AMQ

[1658] Beth Beard. 144th Annual Meeting in Québec City — from fisheries research to management: Think and act locally and globally. Fisheries, 39(12):587–603, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.976860.

Sogard:2014:SBM

[1659] Susan Sogard. Steven Berkeley’s marine conservation legacy lives on. Fisheries, 39(12):604–605, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.976858.

Williams:2014:SBM

[1660] Howard Williams. 2008–2013 Steven Berkeley Marine Conservation Fellowship recipient updates. Fisheries, 39(12):605–607, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.992727.

Anonymous:2014:SBM

[1661] Anonymous. The 2014 Steven Berkeley Marine Conservation Fellowship winners. Fisheries, 39(12):608–609, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.976703.

Anonymous:2014:CSB

[1662] Anonymous. Call for 2015 Steven Berkeley applications, and AFS seeks an Editor for JAAH. Fisheries, 39(12):609, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.976702.

Anonymous:2014:CAN

[1663] Anonymous. Call for award nominations: 2015 American Fisheries Society awards. Fisheries, 39(12):610–611, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.979343.

Anonymous:2014:AFS

[1664] Anonymous. American Fisheries Society section awards for 2014. Fisheries, 39(12):612, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.984152.

Dekker:2014:QDC

[1665] Willem Dekker and John M. Casselman. The 2003 Québec Declaration of Concern About Eel Declines — 11 years later: Are eels climbing back up the slippery slope? Fisheries, 39(12):613–614, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.979342.

Anonymous:2014:KJ

[1666] Anonymous. Keith Jefferts 1931–2014. Fisheries, 39(12):615–616, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.976861.

Anonymous:2014:GKM

[1667] Anonymous. Grace Klein-MacPhee 1939–2014. Fisheries, 39(12): 617, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.979683.

Anonymous:2014:APP

[1668] Anonymous. AFS Past President Richard (Dick) Ryder 1931–2014. Fisheries, 39(12):618–619, 2014. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.979684.

Anonymous:2014:FV

[1669] Anonymous. Fisheries volume 39 2014. Fisheries, 39(12):620, 2014. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.973558.

Anonymous:2014:TAFe

[1670] Anonymous. Transactions of the American Fisheries Society, volume 143, number 6, November 2014. Fisheries, 39(12):621, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.992725.

Anonymous:2014:FEl

[1671] Anonymous. Fisheries events. Fisheries, 39(12):622, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.992726.

Love:2014:MFH

[1672] Milton Love. A meditation on food habit studies. Fisheries, 39 (12):624, 2014. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.976862.

Anonymous:2015:FIPa

[1673] Anonymous. Full issue PDF volume 40, issue 1. Fisheries, 40 (1):1–44, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1011567.

Schaeffer:2015:SPM

[1674] Jeff Schaeffer. Stocked piscivores may be tougher than we thought. Fisheries, 40(1):3, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.992088.

Harrison:2015:ISR

[1675] Sarah Harrison. Increasing survival rates of discarded red snapper: Best release strategies. Fisheries, 40(1):3–4, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.992087.

Bigford:2015:CCC

[1676] Thomas E. Bigford. Culverts as culprit and conservation choice. Fisheries, 40(1):5, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.986572.

Guy:2015:BRR

[1677] Christopher S. Guy, Hilary B. Treanor, Kevin M. Kappenman, Eric A. Scholl, Jason E. Ilgen, and Molly A. H. Webb. Broadening the regulatedriver management paradigm: a case study of the forgotten dead zone hindering pallid sturgeon recovery. Fisheries, 40(1):6–14, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.987236.

Arreguin-Sanchez:2015:CCE

[1678] Francisco Arreguín-Sánchez, Pablo del Monte-Luna, and Manuel J. Zetina-Rejón. Climate change effects on aquatic ecosystems and the challenge for fishery management: Pink shrimp of the Southern Gulf of Mexico. Fisheries, 40(1):15–19, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.988075.

Zollett:2015:GPD

[1679] Erika A. Zollett, Robert J. Trumble, Jill H. Swasey, and Shawn B. Stebbins. Guiding principles for development of effective commercial fishery monitoring programs. Fisheries, 40(1):20–25, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.989124.

Neely:2015:IFG

[1680] Ben C. Neely and Brandon L. Eder. An introduction to free geographic information systems in fisheries science. Fisheries, 40(1): 26–27, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.985377.

Chopin:2015:MAC

[1681] Thierry Chopin. Marine aquaculture in Canada: Well-established monocultures of finfish and shellfish and an emerging integrated multi-trophic aquaculture (IMTA) approach including seaweeds, other invertebrates, and microbial communities. Fisheries, 40(1):28–31, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.986571.

Sopinka:2015:FMO

[1682] Natalie Sopinka. Fishes on the move as oceans heat up. Fisheries, 40(1):32–35, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.989356.

Loftus:2015:FAM

[1683] Andrew Loftus. Fishery Analysis and Modeling Simulator (FAMS) now available for Windows 7 and 8. Fisheries, 40(1):36, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2014.987235.

Gruss:2015:QPF

[1684] Arnaud Grüss. Q&A: The present and the future of world and U.S. fisheries — interview with Daniel Pauly. Fisheries, 40(1):37–41, 2015. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.980903.

Anonymous:2015:Ca

[1685] Anonymous. Correction. Fisheries, 40(1):42, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1002287.

Anonymous:2015:NAJa

[1686] Anonymous. North American Journal of Fisheries Management, volume 34, number 6, December 2014. Fisheries, 40(1):42, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1002288.

Anonymous:2015:FEa

[1687] Anonymous. Fisheries events. Fisheries, 40(1):43, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.992717.

Anonymous:2015:FIPb

[1688] Anonymous. Full issue PDF volume 40, issue 2. Fisheries, 40(2): 45–92, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1019772.

Bigford:2015:RE

[1689] Thomas E. Bigford. Roles for everyone. Fisheries, 40(2):47, 2015. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2014.1000487.

Anonymous:2015:CSAa

[1690] Anonymous. Candidate statement: AFS Second Vice President: Dale P. Burkett. Fisheries, 40(2):48, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1006500.

Anonymous:2015:CSAb

[1691] Anonymous. Candidate statement: AFS Second Vice President: Jesse Trushenski. Fisheries, 40(2):49, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1006501.

Harrison:2015:DSD

[1692] Sarah Harrison. Diet shifts detected in sandbar sharks using a nonlethal technique. Fisheries, 40(2):52, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1003276.

Schaeffer:2015:CNU

[1693] Jeff Schaeffer. Cast nets are useful sampling tools, and you should try one. Fisheries, 40(2):53, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.1001025.

Mendoza-Carranza:2015:HRS

[1694] Manuel Mendoza-Carranza and Alejandro Espinoza-Tenorio. A historical record of sawfish in the Southern Gulf of Mexico: Evidence of diversity loss using old photos. Fisheries, 40(2):54–55, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.996807.

Hansen:2015:LMM

[1695] Gretchen J. A. Hansen, Jereme W. Gaeta, Jonathan F. Hansen, and Stephen R. Carpenter. Learning to manage and managing to learn: Sustaining freshwater recreational fisheries in a changing environment. Fisheries, 40(2):56–64, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.996804.

Quinn:2015:TCF

[1696] Thomas P. Quinn. Turning class field trips into long-term research: a great idea with a few pitfalls. Fisheries, 40(2):65–68, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.996805.

Ramirez-Rodriguez:2015:RSN

[1697] Mauricio Ramírez-Rodríguez. Relating scientific names to common names for important fisheries species of the Mexican Pacific. Fisheries, 40(2):69–71, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.996806.

Wang:2015:CPL

[1698] Christina Wang and Howard Schaller. Conserving Pacific lamprey through collaborative efforts. Fisheries, 40(2):72–79, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.996871.

Mawdsley:2015:EDC

[1699] Jonathan Mawdsley and Arpita Choudhury. Environmental DNA clearinghouse: New tools for managers. Fisheries, 40(2):80, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2014.999047.

Carlson:2015:PHI

[1700] Andrew K. Carlson, Jesse R. Fischer, Landon L. Pierce, Dan J. Dembkowski, Michael E. Colvin, Janice A. Kerns, and Jeffrey D. Fore. Purpose, history, and importance of the student angle. Fisheries, 40 (2):81–83, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.964399.

Sopinka:2015:URV

[1701] Natalie Sopinka. Underwater romance: a Valentine’s Day ode to fishes. Fisheries, 40(2):84–85, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.1001026.

Wiens:2015:RES

[1702] John A. Wiens. Review of an ecosystem services approach to assessing the impacts of the Deepwater Horizon oil spill in the Gulf of Mexico. Fisheries, 40(2):86, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.972107.

Anonymous:2015:Cb

[1703] Anonymous. Correction. Fisheries, 40(2):89, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1016812.

Anonymous:2015:NAJb

[1704] Anonymous. North American Journal of Aquaculture, volume 77, number 1, January 2015. Fisheries, 40(2):89, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1012374.

Anonymous:2015:Cc

[1705] Anonymous. Calendar. Fisheries, 40(2):90, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1012375.

Love:2015:DYK

[1706] Milton Love. Did you know? Rockfish. Fisheries, 40(2):92, 2015. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/03632415.2014.987234.

Anonymous:2015:FIPc

[1707] Anonymous. Full issue PDF volume 40, issue 3. Fisheries, 40(3): 93–140, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1029803.

Parrish:2015:AMH

[1708] Donna Parrish. AFS membership: How much should it cost? Fisheries, 40(3):95, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1010720.

Bigford:2015:ORE

[1709] Thomas E. Bigford. Our renewed effort to make the most of science. Fisheries, 40(3):96, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2015.1009046.

Schaeffer:2015:CAR

[1710] Jeff Schaeffer. Calling all reviewers — the AFS Editorial Board. Fisheries, 40(3):97, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1006769.

Schaeffer:2015:AJS

[1711] Jeff Schaeffer. AFS journals support sustainable aquaculture. Fisheries, 40(3):100, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1013387.

Schaeffer:2015:PBT

[1712] Jeff Schaeffer. Predicting brook trout presence at the local level, where it counts. Fisheries, 40(3):101, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1013388.

Bettoli:2015:LSB

[1713] Phillip W. Bettoli. The last stand of the Barrens topminnow. Fisheries, 40(3):102–103, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1006324.

Kopaska:2015:BGP

[1714] Jeff Kopaska. Brushes with greatness: Preserving original Maynard Reece fish art. Fisheries, 40(3):104–107, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1009047.

Meyer:2015:ASY

[1715] Hilary Meyer, Quinton Phelps, Jesse Fischer, Mark Fincel, Marty Hamel, Rebecca Krogman, Tom Lang, Travis Neebling, Landon Pierce, and Tyler Stubbs. Aiming to support young professionals. Fisheries, 40 (3):108, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1008695.

Tringali:2015:QBB

[1716] Michael D. Tringali, James M. Long, Timothy W. Birdsong, and Micheal S. Allen. Q&A: Black bass diversity: Multidisciplinary science for conservation. Fisheries, 40(3):109, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.955093.

Utter:2015:RRW

[1717] Fred Utter. Review of The Rockfish’s Warning. Fisheries, 40(3): 110, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2014.947202.

Orr:2015:RGR

[1718] James W. Orr. Review of A Guide to the Rockfishes, Thornyheads, and Scorpionfishes of the Northeast Pacific. Fisheries, 40(3):111, 2015. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2015.1007208.

Gunderson:2015:RCM

[1719] Donald R. Gunderson. Review of Certainly More Than You Want to Know About The Fishes of The Pacific Coast: a Postmodern Experience. Fisheries, 40(3):112, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2014.943746.

Cooke:2015:IIN

[1720] Steven J. Cooke and Jack Imhof. Inaugural inductees and New AFS Director of Student and Professional Development. Fisheries, 40(3):113, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1013389.

Dittman:2015:IHR

[1721] Andrew H. Dittman, Todd N. Pearsons, Darran May, Ryan B. Couture, and David L. G. Noakes. Imprinting of hatchery-reared salmon to targeted spawning locations: a new embryonic imprinting paradigm for hatchery programs. Fisheries, 40(3):114–123, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1007206.

Rieman:2015:CAH

[1722] Bruce E. Rieman, Courtland L. Smith, Robert J. Naiman, Gregory T. Ruggerone, Chris C. Wood, Nancy Huntly, Erik N. Merrill, J. Richard Alldredge, Peter A. Bisson, James Congleton, Kurt D. Fausch, Colin Levings, William Pearcy, Dennis Scarnecchia, and Peter Smouse. A comprehensive approach for habitat restoration in the Columbia Basin. Fisheries, 40(3):124–135, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1007205.

Anonymous:2015:TAFa

[1723] Anonymous. Transactions of the American Fisheries Society, volume 144, number 1, January 2015. Fisheries, 40(3):137, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1025018.

Anonymous:2015:FEb

[1724] Anonymous. Fisheries events. Fisheries, 40(3):138, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1025019.

Sopinka:2015:LFF

[1725] Natalie Sopinka. Living fossil fishes. Fisheries, 40(3):140, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1007500.

Anonymous:2015:FIPd

[1726] Anonymous. Full issue PDF volume 40, issue 4. Fisheries, 40(4): 141–188, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1039888.

Parrish:2015:HMP

[1727] Donna Parrish. How to be a meeting participant and not just an attendee. Fisheries, 40(4):143, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2015.1021922.

Bigford:2015:ROA

[1728] Thomas E. Bigford. Reaching outside AFS to heighten our impact. Fisheries, 40(4):144, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1020374.

Harrison:2015:LTD

[1729] Sarah Harrison. Lobster trap debris in a marine sanctuary — more than you could imagine. Fisheries, 40(4):145, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1027158.

Anonymous:2015:WP

[1730] Anonymous. Welcome to Portland! Fisheries, 40(4):146, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1027159.

Stallsmith:2015:WCT

[1731] Bruce Stallsmith. What can I tell you about the Flint River in Alabama? Fisheries, 40(4):148–149, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1021921.

Schaeffer:2015:QID

[1732] Jeff Schaeffer. Q&A: Interview with David Fielder. Fisheries, 40(4): 150–151, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1027160.

Harrison:2015:QFT

[1733] Sarah Harrison. Q&A: Fish talk with Kristen Dahl. Fisheries, 40 (4):152–154, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1027161.

Patrick:2015:MCI

[1734] Wesley S. Patrick and Jason S. Link. Myths that continue to impede progress in ecosystem-based fisheries management. Fisheries, 40 (4):155–160, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1024308.

Shi:2015:DFP

[1735] Xiaotao Shi, Boyd Kynard, Defu Liu, Ye Qiao, and Qiuwen Chen. Development of fish passage in China. Fisheries, 40(4):161–169, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1017634.

Miranda:2015:FAI

[1736] L. E. Miranda and Rebecca M. Krogman. Functional age as an indicator of reservoir senescence. Fisheries, 40(4):170–176, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1007207.

Pegg:2015:RRS

[1737] Mark A. Pegg, Kevin L. Pope, Larkin A. Powell, Kelly C. Turek, Jonathan J. Spurgeon, Nathaniel T. Stewart, Nick P. Hogberg, and Mark T. Porath. Reservoir rehabilitations: Seeking the Fountain of youth. Fisheries, 40(4):177–181, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1017635.

Lang:2015:SLF

[1738] Tom Lang. Student Loan Forgiveness Program aims to keep great minds in government. Fisheries, 40(4):182, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1021923.

Anonymous:2015:NAJc

[1739] Anonymous. North American Journal of Fisheries Management, volume 35, number 1, February 2015. Fisheries, 40(4):183, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1037650.

Anonymous:2015:FEc

[1740] Anonymous. Fisheries events. Fisheries, 40(4):185, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1037651.

Sopinka:2015:HHB

[1741] Natalie Sopinka. Hitch-hiking beaver spotted napping atop humpback whale. Fisheries, 40(4):187, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1018382.

Morgan:2015:FEH

[1742] Annie B. Morgan. Fisheries enhancement on the half-shell. Fisheries, 40(4):188, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1018383.

Anonymous:2015:FIPe

[1743] Anonymous. Full issue PDF volume 40, issue 5. Fisheries, 40(5): 189–240, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1050759.

Parrish:2015:HAA

[1744] Donna L. Parrish. Highlighting the 2014 AFS Award of Excellence recipient — Kenneth A. Rose. Fisheries, 40(5):191, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1031886.

Bigford:2015:TPI

[1745] Thomas E. Bigford. Top 10 policy issues for AFS. Fisheries, 40 (5):192, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1033055.

Schaeffer:2015:AJO

[1746] Jeff Schaeffer. AFS journals offer insights into conservation of apex predators. Fisheries, 40(5):194, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1031887.

Schaeffer:2015:WDF

[1747] Jeff Schaeffer. Whirling disease is fortunately not a permanent condition, a novel and efficient way to assess net pen impacts, and sadly, lionfish are more widespread than we thought. Fisheries, 40(5):195, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1037661.

Carlson:2015:HNF

[1748] Andrew K. Carlson, Karen M. Dunmall, Ross E. Boucek, Nicholas W. Cole, Janice A. Kerns, Rebecca M. Krogman, M. Clint Lloyd, Vivian M. Nguyen, Tracy R. Wendt, Shannon L. White, and Kyle L. Wilson. How to navigate fisheries education and employment. Fisheries, 40 (5):196–197, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1025952.

Bigford:2015:AHC

[1749] Thomas E. Bigford. AFS hosts Capitol Hill briefings on marine fisheries management. Fisheries, 40(5):198–199, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1038435.

Lichtkoppler:2015:GLC

[1750] Frank Lichtkoppler, Daniel M. O’Keefe, Joseph Lucente, and Tory Gabriel. The Great Lakes charter fishing industry: 2002 to 2011. Fisheries, 40(5):200–212, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2015.1029574.

OKeefe:2015:FIC

[1751] Daniel M. O’Keefe, Donna L. Wesander, and Chi-Ok Oh. Factors influencing charter fishing effort trends in Lake Huron. Fisheries, 40 (5):214–221, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1025951.

DellApa:2015:NCC

[1752] Andrea Dell’Apa, Evan Knight, Anthony S. Overton, Craig E. Landry, Christopher F. Dumas, John C. Whitehead, and James H. Herstine. The North Carolina charter boat fishery changing with the times: a comparative analysis of the catch composition (1978 and 2007–2008). Fisheries, 40(5):222–233, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1026332.

Anonymous:2015:CEP

[1753] Anonymous. Continuing education program. Fisheries, 40(5):234– 235, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1038436.

Anonymous:2015:Cd

[1754] Anonymous. Correction. Fisheries, 40(5):237, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1046781.

Anonymous:2015:TAFb

[1755] Anonymous. Transactions of the American Fisheries Society, volume 144, number 2, March 2015. Fisheries, 40(5):237, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1046778.

Anonymous:2015:FEd

[1756] Anonymous. Fisheries events. Fisheries, 40(5):238, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1046779.

Sopinka:2015:FTS

[1757] Natalie Sopinka. Fishy tales from science communicators in the Pacific Northwest. Fisheries, 40(5):240, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1030500.

Anonymous:2015:FIPf

[1758] Anonymous. Full issue PDF volume 40, issue 6. Fisheries, 40(6): 241–292, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1055972.

Bigford:2015:NBR

[1759] Thomas E. Bigford. Nature-based remedies to societal woes. Fisheries, 40(6):243, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1038783.

Schaeffer:2015:CCL

[1760] Jeff Schaeffer. Coping with crisis: Lessons from the Maine lobster fishery, “All Water Has pH!” Except now you can understand what that means and ozone disinfection has minimal impacts on larvae of three important species in aquaculture. Fisheries, 40(6):245–246, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1041200.

Harrison:2015:ART

[1761] Sarah Harrison. Artificial reef types affect red snapper sizes, and catchand-release practices inhibit refuge-seeking behavior in fish. Fisheries, 40(6):246–247, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1041201.

Hoyer:2015:ICC

[1762] Mark V. Hoyer, Terry McNabb, Micheal Allen, and Michael D. Netherland. Improving communication/cooperation among aquatic professional societies. Fisheries, 40(6):248–251, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1041335.

Schaeffer:2015:QIJ

[1763] Jeff Schaeffer. Q&A: Interview with Jay Hemdal, Toledo Zoo Fish Curator. Fisheries, 40(6):252–255, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1037667.

Harrison:2015:QIS

[1764] Sarah Harrison. Q&A: Interview with Sara Turner. Fisheries, 40 (6):256, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1041202.

Crane:2015:MNP

[1765] Derek P. Crane, Loren M. Miller, James S. Diana, John M. Casselman, John M. Farrell, Kevin L. Kapuscinski, and Joe K. Nohner. Muskellunge and northern pike ecology and management: Important issues and research needs. Fisheries, 40(6):258–267, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1038382.

Rice:2015:SFI

[1766] Jake Rice, Michael J. Bradford, Keith D. Clarke, Marten A. Koops, Robert G. Randall, and Roger Wysocki. The science framework for implementing the fisheries protection provisions of Canada’s Fisheries Act. Fisheries, 40(6):268–275, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1038381.

Greenberg:2015:TSE

[1767] Saul Greenberg and Theresa Godin. A tool supporting the extraction of angling effort data from remote camera images. Fisheries, 40 (6):276–287, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1038380.

Anonymous:2015:RWL

[1768] Anonymous. R. Weldon “Larry” Larimore. Fisheries, 40(6):288, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1049920.

Anonymous:2015:NAJd

[1769] Anonymous. North American Journal of Aquaculture, volume 77, number 2, April 2015. Fisheries, 40(6):289, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1053730.

Anonymous:2015:FEe

[1770] Anonymous. Fisheries events. Fisheries, 40(6):290, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1053732.

Sopinka:2015:YBT

[1771] Natalie Sopinka. From your body to theirs: How exposure to pharmaceuticals is altering gene expression in fishes. Fisheries, 40(6):292, 2015. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2015.1038784.

Anonymous:2015:FIPg

[1772] Anonymous. Full issue PDF volume 40, issue 7. Fisheries, 40(7): 293–340, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1066628.

Bonar:2015:PCF

[1773] Scott A. Bonar. Point–counterpoint: Be flexible in the number of talks per speaker at meetings! Fisheries, 40(7):295, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1051416.

Bigford:2015:SHL

[1774] Thomas E. Bigford. Silt happens … literally. Fisheries, 40(7): 296, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1049695.

Schaeffer:2015:BCM

[1775] Jeff Schaeffer. Bluegill can be managed more effectively with bag limits! Solutions to California’s drought issues: Modeling suggests ways to conserve Chinook salmon in the face of water diversion. Probiotics show promise for aquaculture. More evidence that plant-based feeds can support sustainable aquaculture. Using a video lander to assess difficult marine habitats. Fisheries, 40(7):297–299, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1051415.

Hammond:2015:FFL

[1776] Marissa L. Hammond. The 2013–2014 Fenske Fellowship: Learning the ins and outs of lake whitefish management in the upper Great Lakes. Fisheries, 40(7):300–301, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1049694.

Hanks:2015:WVU

[1777] Daniel Hanks, Ross Andrew, and Alison Anderson. West Virginia University AFS Student Subunit leads a citizen science program. Fisheries, 40(7):302, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1049696.

Cooper:2015:ABM

[1778] John Cooper. The 63rd Annual Business Meeting of the Northeastern Division. Fisheries, 40(7):303, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1051417.

Williams:2015:CCA

[1779] Jack E. Williams, Helen M. Neville, Amy L. Haak, Warren T. Colyer, Seth J. Wenger, and Stan Bradshaw. Climate change adaptation and restoration of western trout streams: Opportunities and strategies. Fisheries, 40(7):304–317, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2015.1049692.

Papenfuss:2015:SRA

[1780] Jason T. Papenfuss, Nicholas Phelps, David Fulton, and Paul A. Venturelli. Smartphones reveal angler behavior: a case study of a popular mobile fishing application in Alberta, Canada. Fisheries, 40(7): 318–327, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1049693.

Harrison:2015:PSG

[1781] Sarah Harrison. Portland 2015 schedule at a glance. Fisheries, 40 (7):328–331, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1053749.

Waldman:2015:CLS

[1782] John Waldman. C. Lavett Smith, Jr. Fisheries, 40(7):332, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1053748.

Chen:2015:ULU

[1783] Yushun Chen, Kathryn Herzog, Sagar Shrestha, Daniel Grigas, John Farrelly, Christopher Laskodi, and Matthew Skoog. Urban land use, water quality, and biological conditions in the Lower Mississippi River Basin bayous. Fisheries, 40(7):334–335, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2014.943367.

Anonymous:2015:TAFc

[1784] Anonymous. Transactions of the American Fisheries Society, volume 144, number 3, May 2015. Fisheries, 40(7):336, 2015. CODEN ????

ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1059718.

Anonymous:2015:FEf

[1785] Anonymous. Fisheries events. Fisheries, 40(7):337, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1059717.

Sopinka:2015:TSY

[1786] Natalie Sopinka. Try this! Sketch your conference notes. Fisheries, 40 (7):338–339, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1050095.

Anonymous:2015:FIPh

[1787] Anonymous. Full issue PDF volume 40, issue 8. Fisheries, 40(8): 341–428, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1078269.

White:2015:RBT

[1788] Gwen White. Raising the bar: Taking a more strategic and integrated approach to AFS communications. Fisheries, 40(8):344, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1065138.

Bigford:2015:WTC

[1789] Thomas E. Bigford. Wetlands thrive on connectivity. Fisheries, 40 (8):345, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1060225.

Lucey:2015:VAB

[1790] Sean Lucey. Value added by being a member of a professional society. Fisheries, 40(8):346, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2015.1060224.

Austen:2015:STA

[1791] Douglas Austen. The sordid trail of an AFS response to New York Times op-ed piece. Fisheries, 40(8):348, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1067566.

Anonymous:2015:ARO

[1792] Anonymous. AFS responds to an op-ed in the New York Times on trout fishing in the Northeast US. Fisheries, 40(8):349–350, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1073537.

Schaeffer:2015:FCR

[1793] Jeff Schaeffer. Fisheries classics: Ryder’s morphoedaphic index. Identifying endocrine disrupters as an emerging issue: Smallmouth bass in the Potomac River. Avoiding domestication of an endangered species: Conservation culture of California’s delta smelt. Fisheries, 40(8): 351, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1067564.

Harrison:2015:OSI

[1794] Sarah Harrison. Oil spill impacts to red snapper in the Gulf and economic impact and value data can be used to assess policy changes. Fisheries, 40(8):352, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1067565.

Bowker:2015:WMA

[1795] Jim Bowker and Nancy Leonard. What not to miss at the Annual Meeting (including the most unique symposium). Fisheries, 40 (8):353, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1065148.

Hughes:2015:QSS

[1796] Bob Hughes. Q&A: The success story of the Oregon chub: an interview with Paul Scheerer. Fisheries, 40(8):354–355, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1065149.

Moore:2015:AUA

[1797] Christopher Moore. Adventures of an urban angler. Fisheries, 40(8): 356–357, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1065656.

Dawicki:2015:NFS

[1798] Shelley Dawicki. NOAA fisheries at-sea volunteer opportunities: Hands-on, hard work, and a great experience. Fisheries, 40(8):

359–360, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1062373.

Kunkel:2015:MAV

[1799] Joseph Kunkel. My adventure volunteering on NOAA ships. Fisheries, 40(8):360–361, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1068642.

Glor:2015:VCW

[1800] Carol Glor. A voyage that changed the way I thought about life at sea. Fisheries, 40(8):361–362, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1068643.

Harrison:2015:USM

[1801] Sarah Harrison. The University of Southern Mississippi — Gulf Coast Research Laboratory: Interview with Jim Franks. Fisheries, 40(8): 363–365, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1069059.

Anonymous:2015:RAF

[1802] Anonymous. Recognition of the American Fisheries Society professional certification program. Fisheries, 40(8):366, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1065659.

Robinson:2015:LLS

[1803] Claire E. Robinson. From landlocked to longlining: a summer with sharks puts stress into perspective. Fisheries, 40(8):367–368, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1060965.

Trana:2015:EWA

[1804] Marci Trana. Exploring the wonders of the Arctic from a lab bench. Fisheries, 40(8):369, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2015.1060415.

Christensen:2015:FTF

[1805] Villy Christensen. Fishing down through the food web. Fisheries, 40 (8):370–372, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1065155.

Branch:2015:FIF

[1806] Trevor A. Branch. Fishing impacts on food webs: Multiple working hypotheses. Fisheries, 40(8):373–375, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1059825.

Rountos:2015:WCW

[1807] Konstantine J. Rountos, Michael G. Frisk, and Ellen K. Pikitch. Are we catching what they eat? Moving beyond trends in the mean trophic level of catch. Fisheries, 40(8):376–385, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1061509.

Dickinson:2015:CHD

[1808] Benjamin D. Dickinson, Donald J. Orth, and Steve L. McMullin. Characterizing the human dimensions of a hidden fishery: Riverine trotline fishers. Fisheries, 40(8):386–394, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1062372.

DeWeber:2015:TCC

[1809] Jefferson T. DeWeber and Tyler Wagner. Translating climate change effects into everyday language: an example of more driving and less angling. Fisheries, 40(8):395–398, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1065252.

Maitland:2015:FPS

[1810] Bryan M. Maitland, Steven J. Cooke, and Mark Poesch. Finding the path to a successful graduate and research career: Advice for early career researchers. Fisheries, 40(8):399–403, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1065253.

Anonymous:2015:MT

[1811] Anonymous. Membership testimonials. Fisheries, 40(8):404–405, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1067560.

Anonymous:2015:MAS

[1812] Anonymous. Meet the AFS Staff. Fisheries, 40(8):406–409, 2015. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2015.1068071.

Anonymous:2015:AFSa

[1813] Anonymous. American Fisheries Society 2014 Annual Report. Fisheries, 40(8):411–425, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1067561.

Anonymous:2015:FEg

[1814] Anonymous. Fisheries events. Fisheries, 40(8):426, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1071104.

Anonymous:2015:JAA

[1815] Anonymous. Journal of Aquatic Animal Health, volume 27, number 2, June 2015. Fisheries, 40(8):427, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1074847.

Sopinka:2015:TTM

[1816] Natalie Sopinka. Try this! “Tweeting the Meeting”. Fisheries, 40 (8):428, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1065157.

Anonymous:2015:FIPi

[1817] Anonymous. Full issue PDF volume 40, issue 9. Fisheries, 40(9): 429–476, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1090249.

Essig:2015:FCM

[1818] Ron Essig. Fisheries conservation and management: Making connections and building partnerships. Fisheries, 40(9):431, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1076677.

Bigford:2015:VNA

[1819] Thomas E. Bigford. Vessel noise alters fish behavior. Fisheries, 40 (9):432, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1074571.

Schaeffer:2015:NGI

[1820] Jeff Schaeffer. NAJFM goes international: a recreational fishing study in Australia; if you remove a dam will anadromous fish still return? YES!; fish farms alter wild fish diets; all fish anesthetics are not alike!; aquaculture continues to evolve. Fisheries, 40(9):433–434, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1080518.

Ng:2015:SWC

[1821] Elizabeth Ng. 2014 student writing contest winner: Learning from the past. Fisheries, 40(9):436, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2014.955092.

deSouza:2015:AAG

[1822] Lesley S. de Souza. Arapaima adventures in Guyana. Fisheries, 40 (9):437–438, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1074074.

Bayer:2015:SNS

[1823] Skylar R. Bayer, Bonnie J. Becker, and P. Sean McDonald. Summary of the 2015 National Shellfisheries Association Meeting. Fisheries, 40 (9):439–440, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1083384.

McBride:2015:IOS

[1824] Richard McBride, Joseph O’Malley, Trika Gerard, and Beverly Barnett. 5th International Otolith Symposium. Fisheries, 40(9):440, 2015. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2015.1083382.

Anonymous:2015:IAF

[1825] Anonymous. Introducing the American Fisheries Society Fellows Program. Fisheries, 40(9):441, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1076682.

Handley:2015:EGD

[1826] Grace Handley, Cynthia M. Frantz, Patrick M. Kocovsky, Dennis R. DeVries, Steven J. Cooke, and Julie Claussen. An examination of gender differences in the American Fisheries Society peer-review process. Fisheries, 40(9):442–451, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1059824.

Cook:2015:FWH

[1827] Katrina V. Cook, Robert J. Lennox, Scott G. Hinch, and Steven J. Cooke. FISH out of WATER: How much air is too much? Fisheries, 40(9):452–461, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1074570.

Secor:2015:AEW

[1828] David H. Secor. American eel: When does diversity matter? Fisheries, 40(9):462–463, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1073152.

Douglas:2015:IEI

[1829] Marlis R. Douglas, Yury V. Slynko, Yury Yu. Dgebuadze, Sergej Olenin, Borys Aleksandrov, Alexander Boltachev, Elena E. Slynko, Dima Khristenko, Dan Minchin, Dmitry F. Pavlov, Andrey N. Reshetnikov, Dmitry A. Vekhov, Christopher J. Ware, and Michael E. Douglas. Invasion ecology: an international perspective centered in the Holarctic. Fisheries, 40(9):464–470, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1075344.

Anonymous:2015:FEh

[1830] Anonymous. Fisheries events. Fisheries, 40(9):472, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1077083.

Anonymous:2015:NAJe

[1831] Anonymous. North American Journal of Fisheries Management, Volume 35, Number 4, August 2015. Fisheries, 40(9):473, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1083740.

Anonymous:2015:AMAa

[1832] Anonymous. 146th Annual Meeting of the American Fisheries Society: First call for papers. Fisheries, 40(9):474–475, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1083739.

Sopinka:2015:FFD

[1833] Natalie Sopinka. Fish and fisheries in the dark. Fisheries, 40 (9):476, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1074800.

Anonymous:2015:FIPj

[1834] Anonymous. Full issue PDF volume 40, issue 10. Fisheries, 40(10): 477–524, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1105018.

Essig:2015:SGT

[1835] Ron Essig. The significance of gasoline taxes to fisheries. Fisheries, 40(10):479, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1080690.

Bigford:2015:BRE

[1836] Thomas E. Bigford. A blue-ribbon effort to identify creative financing. Fisheries, 40(10):480, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1079520.

Schaeffer:2015:IYW

[1837] Jeff Schaeffer. If you want better recruitment estimates, stratify your sampling by habitat; your favorite microbrew may support fish production; assays are important but not simple to create. Fisheries, 40 (10):482, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1084776.

Harrison:2015:MEI

[1838] Sarah Harrison. Mechanical eradication of invasive fish species can be effective; timing is everything during salmon migration in estuaries. Fisheries, 40(10):483, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1100489.

Lederman:2015:HES

[1839] Nathan J. Lederman and Andrew K. Carlson. How to establish a Student Subunit. Fisheries, 40(10):484–485, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1089097.

Fenn:2015:MNF

[1840] Carlin Fenn. Mora National Fish Hatchery. Fisheries, 40(10):486– 489, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1100490.

Harrison:2015:QIB

[1841] Sarah Harrison. Q&A with Ivy Baremore. Fisheries, 40(10):490– 491, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1080691.

Braccini:2015:GQA

[1842] Matias Braccini. Is a global quantitative assessment of shark populations warranted? Fisheries, 40(10):492–501, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1080689.

Struthers:2015:ACB

[1843] Daniel P. Struthers, Andy J. Danylchuk, Alexander D. M. Wilson, and Steven J. Cooke. Action cameras: Bringing aquatic and fisheries research into view. Fisheries, 40(10):502–512, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1082472.

Westhoff:2015:PAI

[1844] Jacob T. Westhoff and Travis A. Kobermann. Prevalence of aquatic introduced species prevention protocols at U.S. college and university fisheries programs. Fisheries, 40(10):513–519, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1083431.

Anonymous:2015:AMAb

[1845] Anonymous. 146th Annual Meeting of the American Fisheries Society: Second call for papers. Fisheries, 40(10):520, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1100486.

Anonymous:2015:TAFd

[1846] Anonymous. Transactions of the American Fisheries Society, volume 144, number 5, September 2015. Fisheries, 40(10):521, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1100487.

Anonymous:2015:FEi

[1847] Anonymous. Fisheries events. Fisheries, 40(10):522, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1100488.

Sopinka:2015:LFS

[1848] Natalie Sopinka. Listening to fishes to see where they go. Fisheries, 40(10):524, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1082473.

Anonymous:2015:FIPk

[1849] Anonymous. Full issue PDF volume 40, issue 11. Fisheries, 40(11): 525–572, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1115707.

Essig:2015:PM

[1850] Ron Essig. The power of a motion. Fisheries, 40(11):527, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1093474.

Bigford:2015:TKB

[1851] Thomas E. Bigford. Traditional knowledge and biodiversity. Fisheries, 40(11):528, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1095184.

Schaeffer:2015:FCM

[1852] Jeff Schaeffer. Fisheries classics: The model that turned out to be, sadly, but absolutely right; and, everything you thought about stone crab fisheries is wrong. Fisheries, 40(11):529, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1100484.

Osborne-Gowey:2015:DCS

[1853] Jeremiah Osborne-Gowey. Doing cool science? Why you should be blogging about it. Fisheries, 40(11):530, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2014.935345.

Anonymous:2015:CBN

[1854] Anonymous. Craig Brook National Fish Hatchery. Fisheries, 40(11): 532–533, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1100481.

Skaggs:2015:DNA

[1855] Jordan Skaggs and Micheal S. Allen. Data needs to assess effects of soft plastic lure ingestion on fish populations. Fisheries, 40(11): 534–535, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1100477.

Provost:2015:IFH

[1856] Mikaela M. Provost and Olaf P. Jensen. The impacts of fishing on hermaphoditic and treatment of sex change in stock assessments. Fisheries, 40(11):536–545, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1093471.

Thompson:2015:RIP

[1857] Paul D. Thompson, P. Aaron Webber, and Cassie D. Mellon. The role of introduced populations in the management and conservation of least chub. Fisheries, 40(11):546–556, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1093472.

Rombenso:2015:RTK

[1858] Artur Rombenso, Karma Kissinger, Michael Ciaramella, and Jesse Trushenski. Relevant topics to keep in mind as a graduate student. Fisheries, 40(11):557–561, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1093473.

Wieferich:2015:EUN

[1859] Daniel J. Wieferich, Wesley M. Daniel, and Dana M. Infante. Enhancing the utility of the NHDPlus river coverage: Characterizing ecological river reaches for improved management and summary of information. Fisheries, 40(11):562–564, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1090977.

Anonymous:2015:AMAc

[1860] Anonymous. 146th Annual Meeting of the American Fisheries Society: Third call for papers. Fisheries, 40(11):565, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1108757.

Anonymous:2015:NAJf

[1861] Anonymous. North American Journal of Aquaculture, volume 77, number 4, October 2015. Fisheries, 40(11):566, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1108753.

Anonymous:2015:Ce

[1862] Anonymous. Correction. Fisheries, 40(11):568, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1105557.

Anonymous:2015:FEj

[1863] Anonymous. Fisheries events. Fisheries, 40(11):568, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1108754.

Sopinka:2015:TFS

[1864] Natalie Sopinka. The teeth of fishes: Say ahhh! Fisheries, 40(11): 571–572, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1100045.

Anonymous:2015:FIPl

[1865] Anonymous. Full issue PDF, volume 40, issue 12. Fisheries, 40(12): 573–628, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2015.1130505.

Essig:2015:BAT

[1866] Ron Essig. The big AFS tent. Fisheries, 40(12):575, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1109422.

Bigford:2015:LGS

[1867] Thomas E. Bigford. Lessons from the greater sage grouse and Oregon chub. Fisheries, 40(12):576, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1104304.

Schaeffer:2015:NTS

[1868] Jeff Schaeffer. A novel technique for spawning channel catfish based on new understanding of both physiology and genetics; phenology: One of the least known consequences of climate change; experiments reveal cross-class transmission of ranaviruses. Fisheries, 40(12):577– 578, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1113090.

Anonymous:2015:Cf

[1869] Anonymous. Correction. Fisheries, 40(12):578, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1125705.

Anonymous:2015:AMAd

[1870] Anonymous. 146th Annual Meeting of the American Fisheries Society: Final call for papers. Fisheries, 40(12):579, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1117274.

Bonar:2015:SWB

[1871] Scott A. Bonar, Norman Mercado-Silva, Matt Rahr, Yuta T. Torrey, and Averill Cate, Jr. A simple Web-based tool to compare freshwater fish data collected using AFS standard methods. Fisheries, 40 (12):580–589, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1106944.

Hartman:2015:BIA

[1872] Kyle J. Hartman, F. Joseph Margraf, Andrew W. Hafs, and M. Keith Cox. Bioelectrical impedance analysis: a new tool for assessing fish condition. Fisheries, 40(12):590–600, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1106943.

Beard:2015:AMP

[1873] Beth Beard. 2015 145th Annual Meeting in Portland: Fish sticks, doughnuts, and haruspication. Fisheries, 40(12):601–611, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2015.1112164.

Anonymous:2015:AFSb

[1874] Anonymous. American Fisheries Society awards for 2015. Fisheries, 40 (12):612–618, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1109420.

Anonymous:2015:SBM

[1875] Anonymous. The 2015 Steven Berkeley Marine Conservation Fellowship winners. Fisheries, 40(12):619–620, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2015.1109418.

Anonymous:2015:CAN

[1876] Anonymous. Call for award nominations: 2016 American Fisheries Society awards. Fisheries, 40(12):621–622, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1109419.

Anonymous:2015:AIF

[1877] Anonymous. Annual index Fisheries volume 40 2015. Fisheries, 40 (12):623, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1109423.

Anonymous:2015:NAJg

[1878] Anonymous. North American Journal of Fisheries Management, volume 35, number 5, October 2015. Fisheries, 40(12):624, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1117273.

Anonymous:2015:FEk

[1879] Anonymous. Fisheries events. Fisheries, 40(12):625, 2015. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1117272.

Sopinka:2015:OGK

[1880] Natalie Sopinka. Otoliths: The gift that keeps on giving new data. Fisheries, 40(12):627–628, 2015. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1112167.

Anonymous:2016:FIPa

[1881] Anonymous. Full issue PDF, volume 41, issue 1. Fisheries, 41 (1):1–56, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1134950.

Jackson:2016:FU

[1882] Gary Jackson and Ron Essig. Fisheries down under. Fisheries, 41 (1):3, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1119585.

Bigford:2016:RPR

[1883] Thomas E. Bigford. Revisiting the protection/restoration debate. Fisheries, 41(1):4, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1119593.

Harrison:2016:PGM

[1884] Sarah Harrison. Postsmolt growth more important than initial growth at sea; cluster sampling made easy. Fisheries, 41(1):5, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1119606.

Candidate:2016:WWF

[1885] Michelle Lavery Msc Candidate. Winner: Winter: The forgotten study season. Fisheries, 41(1):6–7, 2016. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2016.1128380.

Dunmall:2016:HMM

[1886] Karen M. Dunmall. Honorable mention: My Arctic alter ego. Fisheries, 41(1):8, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1130488.

Schaeffer:2016:CF

[1887] Jeff Schaeffer. Changes at Fisheries. Fisheries, 41(1):9, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1130495.

Lynch:2016:NDD

[1888] Abigail J. Lynch. No damsel in distress. Fisheries, 41(1):10– 11, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1111875.

Hughes:2016:APP

[1889] Robert M. Hughes, Felipe Amezcua, David M. Chambers, Wesley M. Daniel, James S. Franks, William Franzin, Donald MacDonald, Eric Merriam, George Neall, Paulo dos Santos Pompeu, Lou Reynolds, and Carol Ann Woody. AFS position paper and policy on mining and fossil fuel extraction. Fisheries, 41(1):12–15, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1121742.

Rountos:2016:DFS

[1890] Konstantine J. Rountos. Defining forage species to prevent a management dilemma. Fisheries, 41(1):16–17, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1110791.

Hilborn:2016:CCF

[1891] Ray Hilborn. Correlation and causation in fisheries and watershed management. Fisheries, 41(1):18–25, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1119600.

Hoffman:2016:SDE

[1892] Joel C. Hoffman, Joshua Schloesser, Anett S. Trebitz, Greg S. Peterson, Michelle Gutsch, Henry Quinlan, and John R. Kelly. Sampling design for early detection of aquatic invasive species in Great Lakes ports. Fisheries, 41(1):26–37, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1114926.

Rahel:2016:CPF

[1893] Frank J. Rahel. Changing philosophies of fisheries management as illustrated by the history of fishing regulations in Wyoming. Fisheries, 41(1):38–48, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1116444.

Kocik:2016:RSR

[1894] John F. Kocik. Running silver: Restoring Atlantic rivers and their great fish migrations. Fisheries, 41(1):49, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1119597.

Anonymous:2016:TAFa

[1895] Anonymous. Transactions of the American Fisheries Society, volume 144, number 6, November 2015. Fisheries, 41(1):52, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1130494.

Anonymous:2016:FEa

[1896] Anonymous. Fisheries events. Fisheries, 41(1):53, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1119607.

Sopinka:2016:YR

[1897] Natalie Sopinka. Year in review: 2015. Fisheries, 41(1):55–56, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1119602.

Anonymous:2016:FIPb

[1898] Anonymous. Full issue PDF, volume 41, issue 2. Fisheries, 41(2): 57–108, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1145527.

Essig:2016:SIF

[1899] Ron Essig. Sparking interest in fisheries. Fisheries, 41(2):59, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1128761.

Bigford:2016:ELF

[1900] Thomas E. Bigford. Emerging and legacy fish worries. Fisheries, 41(2):60, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1127228.

Bonar:2016:CSA

[1901] Scott Bonar. Candidate statement: AFS Second Vice President: Scott Bonar. Fisheries, 41(2):61–62, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1128767.

Whelan:2016:CSA

[1902] Gary Whelan. Candidate statement AFS Second Vice President: Gary Whelan. Fisheries, 41(2):63–64, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1128768.

Anonymous:2016:AKC

[1903] Anonymous. AFS 2016 Kansas City Booth Reservation form. Fisheries, 41(2):65, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1133196.

Moffitt:2016:ADC

[1904] Christine M. Moffitt. Actions and dialog to change perceptions and increase engagement of underrepresented minorities in fisheries and aquatic sciences: Report to membership from a special session in Portland. Fisheries, 41(2):66–67, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1133192.

Sopinka:2016:HDC

[1905] Natalie Sopinka. How does creativity fit into science communication? Fisheries, 41(2):68–69, 2016. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2016.1130530.

Margraf:2016:WKI

[1906] Joe Margraf. Who knew? Inconnu. Fisheries, 41(2):70, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1130535.

Morgan:2016:WFA

[1907] David L. Morgan, Barbara E. Wueringer, Mark G. Allen, Brendan C. Ebner, Jeff M. Whitty, Adrian C. Gleiss, and Stephen J. Beatty. What is the fate of amputee sawfish? Fisheries, 41(2):71–73, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2015.1125887.

Krogman:2016:SMA

[1908] Rebecca M. Krogman. A standard method for abbreviating the names of fishes of the United States, Canada, and Mexico. Fisheries, 41 (2):74–82, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1125886.

Bouwes:2016:AAM

[1909] Nicolaas Bouwes, Stephen Bennett, and Joe Wheaton. Adapting adaptive management for testing the effectiveness of stream restoration: an intensively monitored watershed example. Fisheries, 41(2): 84–91, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1127806.

Bennett:2016:PCT

[1910] Stephen Bennett, George Pess, Nicolaas Bouwes, Phil Roni, Robert E. Bilby, Sean Gallagher, Jim Ruzycki, Thomas Buehrens, Kirk Krueger, William Ehinger, Joseph Anderson, Chris Jordan, Brett Bowersox, and Correigh Greene. Progress and challenges of testing the effectiveness of stream restoration in the Pacific Northwest using intensively monitored watersheds. Fisheries, 41(2):92–103, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1127805.

Anonymous:2016:FME

[1911] Anonymous. Fred M. Eiserman, October 24, 1923–November 14, 2015. Fisheries, 41(2):104, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1142751.

Anonymous:2016:NAJa

[1912] Anonymous. North American Journal of Fisheries Management, volume 35, number 6, December 2015. Fisheries, 41(2):105, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1135724.

Anonymous:2016:FEb

[1913] Anonymous. Fisheries events. Fisheries, 41(2):106, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1135723.

Sopinka:2016:BPPa

[1914] Natalie Sopinka. Back page photo series: Blocked: an interview with Sean Landsman. Fisheries, 41(2):108, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1130551.

Anonymous:2016:FIPc

[1915] Anonymous. Full issue PDF, volume 41, issue 3. Fisheries, 41(3): 109–156, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1156386.

Essig:2016:OFT

[1916] Ron Essig. One fish, two fish. Fisheries, 41(3):111, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1142242.

Bigford:2016:CCF

[1917] Thomas E. Bigford. Climate change and fish. Fisheries, 41(3): 112, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1142254.

Austen:2016:AAM

[1918] Doug Austen. AFS Annual Meetings: What works and what doesn’t. Fisheries, 41(3):113, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1142253.

Minke-Martin:2016:NN

[1919] Vanessa Minke-Martin. Narratives of nature. Fisheries, 41(3):114– 115, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1135691.

Waldman:2016:RAC

[1920] John Waldman, Karen A. Wilson, Martha Mather, and Noah P. Snyder. A resilience approach can improve anadromous fish restoration. Fisheries, 41(3):116–126, 2016. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2015.1134501.

Anonymous:2016:AAM

[1921] Anonymous. 146th AFS Annual Meeting in Kansas City, MO. Fisheries, 41(3):127, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1152634.

Munoz-Colmenero:2016:DAF

[1922] Marta Muñoz-Colmenero, Oscar Blanco, Vanessa Arias, Jose Luis Martinez, and Eva Garcia-Vazquez. DNA authentication of fish products reveals mislabeling associated with seafood processing. Fisheries, 41 (3):128–138, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2015.1132706.

Jager:2016:RFS

[1923] Henriette I. Jager, Michael J. Parsley, Joseph J. Cech, Jr., Robert L. McLaughlin, Patrick S. Forsythe, Robert F. Elliott, and Brenda M. Pracheil. Reconnecting fragmented sturgeon populations in North American rivers. Fisheries, 41(3):140–148, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2015.1132705.

Waldman:2016:QJW

[1924] John Waldman. Q&A with John Waldman. Fisheries, 41(3): 149, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1119598.

Duguid:2016:WSR

[1925] W. D. P. Duguid. Walking sideways: The remarkable world of crabs. Fisheries, 41(3):150, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2015.1024309.

Anonymous:2016:JAAa

[1926] Anonymous. Journal of Aquatic Animal Health volume 27, number 4; North American Journal of Aquaculture volume 78, number 1. Fisheries, 41(3):151, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1152635.

Anonymous:2016:FEc

[1927] Anonymous. Fisheries events. Fisheries, 41(3):153, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1152636.

Sopinka:2016:BPPb

[1928] Natalie Sopinka and Allison Barner. Back page photo series: Intertidal after dark. Fisheries, 41(3):155–156, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1135692.

Anonymous:2016:FIPd

[1929] Anonymous. Full issue PDF, volume 41, issue 4. Fisheries, 41(4): 157–212, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1172453.

Essig:2016:APC

[1930] Ron Essig. AFS Professional Certification 101. Fisheries, 41(4): 159, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1153890.

Bigford:2016:WAH

[1931] Thomas E. Bigford. Walls, aquatic habitat, and fish. Fisheries, 41 (4):160, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1150264.

Austen:2016:DSY

[1932] Doug Austen. The dollars and sense (yes) of the Portland AFS Meeting. Fisheries, 41(4):161, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1164526.

Bonar:2016:QCA

[1933] Scott Bonar. Q&A with the candidates: AFS Second Vice President. Fisheries, 41(4):162–164, 2016. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/10962247.2016.1163148.

Whelan:2016:QCA

[1934] Gary Whelan. Q&A with the candidates: AFS Second Vice President. Fisheries, 41(4):165–167, 2016. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2016.1163155.

Anonymous:2016:WCW

[1935] Anonymous. We can’t wait to see you in Kansas City! Fisheries, 41(4):168, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1153909.

Baumwell:2016:CSB

[1936] Leah Baumwell. Citizen science at its best. Fisheries, 41(4):169– 170, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1159063.

Good:2016:MFF

[1937] Molly J. Good. The MSU Fenske Fellowship: Fresh perspectives on fish, management, and law. Fisheries, 41(4):171–172, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1159065.

Scribner:2016:AGD

[1938] Kim T. Scribner, Winsor H. Lowe, Erin Landguth, Gordon Luikart, Dana M. Infante, Gary E. Whelan, and Clint C. Muhlfeld. Applications of genetic data to improve management and conservation of river fishes and their habitats. Fisheries, 41(4):174–188, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1150838.

Allen:2016:SSW

[1939] M. Brady Allen, Rod O. Engle, Joseph S. Zendt, Frank C. Shrier, Jeremy T. Wilson, and Patrick J. Connolly. Salmon and steelhead in the White Salmon River after the removal of Condit Dam — planning efforts and recolonization results. Fisheries, 41(4):190–203, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1150839.

Haw:2016:PKB

[1940] Frank Haw and Geraldine Vander Haegen. Peter K. Bergman 1933– 2016. Fisheries, 41(4):204, 2016. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2016.1159062.

Fox-Tolson:2016:QJT

[1941] Sarah Fox-Tolson. Q&A with Jesse Trushenski, AFS Second Vice President. Fisheries, 41(4):205, 2016. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2016.1168669.

Anonymous:2016:Ca

[1942] Anonymous. Correction. Fisheries, 41(4):206, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1164528.

Anonymous:2016:TAFb

[1943] Anonymous. Transactions of the American Fisheries Society, volume 145, number 1, January 2016. Fisheries, 41(4):206, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1153911.

Anonymous:2016:FEd

[1944] Anonymous. Fisheries events. Fisheries, 41(4):208, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1161375.

Darimont:2016:WMD

[1945] Chris T. Darimont. A world for my daughter: an ecologist’s search for optimism. Fisheries, 41(4):210, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1168660.

Sopinka:2016:P

[1946] Natalie Sopinka. Paired up. Fisheries, 41(4):212, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1153897.

Anonymous:2016:FIPe

[1947] Anonymous. Full issue PDF, volume 41, issue 5. Fisheries, 41(5): 213–264, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1181512.

Essig:2016:AOT

[1948] Ron Essig. AFS Officer travelogue. Fisheries, 41(5):215, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1162557.

Pool:2016:FON

[1949] Taylor Pool. Future of our nation’s aquatic resources: Recommendations for the next administration. Fisheries, 41(5):216, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1160895.

Nguyen:2016:SGK

[1950] Vivian Nguyen. Seriously, get to know your audience. Fisheries, 41 (5):217–218, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1162568.

Lunde:2016:HH

[1951] Mike Lunde. Headwater heaven. Fisheries, 41(5):219–221, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1167503.

Taylor:2016:MFS

[1952] William W. Taylor and Sarah Harrison. Mentoring for future success. Fisheries, 41(5):223, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1162566.

Harris:2016:PAF

[1953] Fred Harris and Christian Waters. Perspectives on the administration of fisheries programs. Fisheries, 41(5):223–226, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1162567.

Meyer:2016:FHE

[1954] Hilary Meyer. FMS Hall of Excellence video interview series update: I bet you thought it was all about the fish. Fisheries, 41(5):227, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1167498.

Schaeffer:2016:QBK

[1955] Jeff Schaeffer. Q&A with Bruce Koike, this month’s cover artist. Fisheries, 41(5):228–229, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1172889.

Rypel:2016:SYR

[1956] Andrew L. Rypel, John Lyons, Joanna D. Tober Griffin, and Timothy D. Simonson. Seventy-year retrospective on size-structure changes in the recreational fisheries of Wisconsin. Fisheries, 41(5):230–243, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1160894.

Holtgren:2016:RES

[1957] J. Marty Holtgren and Nancy A. Auer. Re-envisioning state and tribal collaboration in fishery assessment and restoration. Fisheries, 41(5): 244–257, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1162159.

Anonymous:2016:NAJb

[1958] Anonymous. North American Journal of Fisheries Management volume 36, number 1, February 2016. Fisheries, 41(5):258, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1164562.

Anonymous:2016:CEP

[1959] Anonymous. Continuing education program: Sheraton Crown Center, Kansas City Saturday August 20–Sunday August 21, 2016. Fisheries, 41(5):259–260, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1162553.

Anonymous:2016:Cb

[1960] Anonymous. Correction. Fisheries, 41(5):262, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1178526.

Anonymous:2016:FEe

[1961] Anonymous. Fisheries events. Fisheries, 41(5):262, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1164563.

Sopinka:2016:BPPc

[1962] Natalie Sopinka. Back page photo series: On the inside. Fisheries, 41 (5):263–264, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1162562.

Anonymous:2016:FIPf

[1963] Anonymous. Full issue PDF, volume 41, issue 6. Fisheries, 41(6): 265–324, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1195212.

Essig:2016:CFI

[1964] Ron Essig. Communicating fisheries information. Fisheries, 41(6): 267, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1179505.

Bigford:2016:IF

[1965] Thomas E. Bigford. Instream flow. Fisheries, 41(6):268, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1181442.

Taylor:2016:CAR

[1966] William W. Taylor and Devin M. Bartley. Call to action — the “Rome Declaration”: Ten steps to responsible inland fisheries. Fisheries, 41 (6):269, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1183993.

Neville:2016:AIC

[1967] Helen Neville. AFS Idaho Chapter Annual Meeting was a great success. Fisheries, 41(6):270, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1175845.

Riley:2016:PCW

[1968] Betsy Riley, Jim Martin, and William W. Taylor. Politics of change: What fisheries scientists and professionals can learn from political campaigns. Fisheries, 41(6):271–273, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1175846.

Schaeffer:2016:QJR

[1969] Jeff Schaeffer. Q&A with Joseph R. Tomelleri. Fisheries, 41(6): 274–275, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1175848.

Maas-Hebner:2016:SDF

[1970] Kathleen G. Maas-Hebner, Carl Schreck, Robert M. Hughes, J. Alan Yeakley, and Nancy Molina. Scientifically defensible fish conservation and recovery plans: Addressing diffuse threats and developing rigorous adaptive management plans. Fisheries, 41(6):276–285, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1175346.

Penaluna:2016:CNP

[1971] Brooke E. Penaluna, Alicia Abadía-Cardoso, Jason B. Dunham, Francisco J. García-Dé León, Robert E. Gresswell, Arturo Ruiz Luna, Eric B. Taylor, Bradley B. Shepard, Robert Al-Chokhachy, Clint C. Muhlfeld, Kevin R. Bestgen, Kevin Rogers, Marco A. Escalante, Ernest R. Keeley, Gabriel M. Temple, Jack E. Williams, Kathleen R. Matthews, Ron Pierce, Richard L. Mayden, Ryan P. Kovach, John Carlos Garza, and Kurt D. Fausch. Conservation of native Pacific trout diversity in Western North America. Fisheries, 41(6):286–300, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1175888.

Anonymous:2016:JAAb

[1972] Anonymous. Journal of Aquatic Animal Health, volume 28, number 1, March 2016. Fisheries, 41(6):301, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1175851.

Jacobson:2016:FHC

[1973] Peter C. Jacobson, Timothy K. Cross, Donna L. Dustin, and Michael Duval. A fish habitat conservation framework for Minnesota lakes. Fisheries, 41(6):302–317, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1172482.

Anonymous:2016:FEf

[1974] Anonymous. Fisheries events. Fisheries, 41(6):320, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1175850.

Sopinka:2016:BPPd

[1975] Natalie Sopinka. Back page photo series: Onboard testing. Fisheries, 41(6):322–324, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1175849.

Anonymous:2016:FIPg

[1976] Anonymous. Full issue PDF, volume 41, issue 7. Fisheries, 41(7): 325–428, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1205860.

Essig:2016:CCS

[1977] Ron Essig. Climate change: SWAPs and AFS. Fisheries, 41(7): 327, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1187008.

Bigford:2016:DFA

[1978] Thomas E. Bigford. Drought, flow, and aquatic resources. Fisheries, 41(7):328, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1182515.

Paukert:2016:ECC

[1979] Craig P. Paukert, Abigail J. Lynch, and James E. Whitney. Effects of climate change on North American inland fishes: Introduction to the special issue. Fisheries, 41(7):329–330, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1187011.

Whitney:2016:PBC

[1980] James E. Whitney, Robert Al-Chokhachy, David B. Bunnell, Colleen A. Caldwell, Steven J. Cooke, Erika J. Eliason, Mark Rogers, Abigail J. Lynch, and Craig P. Paukert. Physiological basis of climate change impacts on North American inland fishes. Fisheries, 41(7): 332–345, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1186656.

Lynch:2016:CCE

[1981] Abigail J. Lynch, Bonnie J. E. Myers, Cindy Chu, Lisa A. Eby, Jeffrey A. Falke, Ryan P. Kovach, Trevor J. Krabbenhoft, Thomas J. Kwak, John Lyons, Craig P. Paukert, and James E. Whitney. Climate change effects on North American inland fish populations and assemblages. Fisheries, 41(7):346–361, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1186016.

Hunt:2016:IAP

[1982] Len M. Hunt, Eli P. Fenichel, David C. Fulton, Robert Mendelsohn, Jordan W. Smith, Tyler D. Tunney, Abigail J. Lynch, Craig P. Paukert, and James E. Whitney. Identifying alternate pathways for climate change to impact inland recreational fishers. Fisheries, 41(7):362–372, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1187015.

Paukert:2016:AIF

[1983] Craig P. Paukert, Bob A. Glazer, Gretchen J. A. Hansen, Brian J. Irwin, Peter C. Jacobson, Jeffrey L. Kershner, Brian J. Shuter, James E. Whitney, and Abigail J. Lynch. Adapting inland fisheries management to a changing climate. Fisheries, 41(7):374–384, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1185009.

Poesch:2016:CCI

[1984] Mark S. Poesch, Louise Chavarie, Cindy Chu, Shubha N. Pandit, and William Tonn. Climate change impacts on freshwater fishes: a Canadian perspective. Fisheries, 41(7):385–391, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1180285.

Limburg:2016:RCS

[1985] Karin Limburg, Randy Brown, Rachel Johnson, Bill Pine, Roger Rulifson, David Secor, Kelly Timchak, Ben Walther, and Karen Wilson. Round-the-coast: Snapshots of estuarine climate change effects. Fisheries, 41(7):392–394, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1182506.

Dege:2016:PSH

[1986] Michael Dege, Eric Jones, Mark Clifford, and Carl Kittel. Providing safe haven for sensitive aquatic species in a changing climate. Fisheries, 41 (7):395–396, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1180286.

Chiaramonte:2016:CCC

[1987] Luciano Chiaramonte, Doug Munson, and Jesse Trushenski. Climate change and considerations for fish health and fish health professionals. Fisheries, 41(7):396–399, 2016. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2016.1182508.

Winfield:2016:IPE

[1988] Ian J. Winfield, Claudio Baigún, Pavel A. Balykin, Barbara Becker, Yushun Chen, Ana F. Filipe, Yuri V. Gerasimov, Alexandre L. Godinho, Robert M. Hughes, John D. Koehn, Dmitry N. Kutsyn, Verónica Mendoza-Portillo, Thierry Oberdorff, Alexei M. Orlov, Andrey P. Pedchenko, Florian Pletterbauer, Ivo G. Prado, Roland Rösch, and Shane J.

Vatland. International perspectives on the effects of climate change on inland fisheries. Fisheries, 41(7):399–405, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1182513.

Rehage:2016:WCW

[1989] J. S. Rehage and J. R. Blanchard. What can we expect from climate change for species invasions? Fisheries, 41(7):405–407, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1180287.

Morrison:2016:MAV

[1990] Wendy E. Morrison, Mark W. Nelson, Roger B. Griffis, and Jonathan A. Hare. Methodology for assessing the vulnerability of marine and anadromous fish stocks in a changing climate. Fisheries, 41(7): 407–409, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1182507.

Nelson:2016:EPP

[1991] Jay A. Nelson and Adalberto L. Val. From the equator to the poles, a physiology section perspective on climate change. Fisheries, 41(7): 409–411, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1182511.

Carlson:2016:CCF

[1992] Andrew K. Carlson and Nathan J. Lederman. Climate change and fisheries education. Fisheries, 41(7):411–412, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1182510.

Chen:2016:AWQ

[1993] Yushun Chen, Andrew S. Todd, Margaret H. Murphy, and Gregg Lomnicky. Anticipated water quality changes in response to climate change and potential consequences for inland fishes. Fisheries, 41(7): 413–416, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1182509.

Whittier:2016:LBD

[1994] Joanna Whittier, Nick Sievert, Andrew Loftus, Julie M. Defilippi, Rebecca M. Krogman, Jeffrey Ojala, Thom Litts, Jeff Kopaska, and Nicole Eiden. Leveraging BIG data from BIG databases to answer BIG questions. Fisheries, 41(7):417–419, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1191911.

Porath:2016:ESI

[1995] Mark T. Porath. Effective stewardship incorporates expertise and innovative approaches to aquatic resource management. Fisheries, 41 (7):419–420, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1191913.

Anonymous:2016:PSA

[1996] Anonymous. Plenary speakers’ abstracts for the 146th Annual Meeting. Fisheries, 41(7):421, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1187020.

Anonymous:2016:TAFc

[1997] Anonymous. Transactions of the American Fisheries Society, volume 145, number 3, May 2016. Fisheries, 41(7):422, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1187021.

Anonymous:2016:FEg

[1998] Anonymous. Fisheries events. Fisheries, 41(7):424, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1187022.

Sopinka:2016:BPPe

[1999] Natalie Sopinka. Back page photo series: Dichotomy. Fisheries, 41 (7):427–428, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1187023.

Anonymous:2016:FIPh

[2000] Anonymous. Full issue PDF, volume 41, issue 8. Fisheries, 41(8): 429–500, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1219166.

Essig:2016:PFF

[2001] Ron Essig. Preparing future fisheries professionals. Fisheries, 41 (8):431, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1199208.

Bigford:2016:CIF

[2002] Thomas E. Bigford. The case for interdisciplinary fisheries education. Fisheries, 41(8):432, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1203163.

Good:2016:EEQ

[2003] Molly J. Good and William W. Taylor. To evolve, or not to evolve? — That is the question. Fisheries, 41(8):433–435, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1203186.

McMullin:2016:WPN

[2004] Steve L. McMullin, Vic DiCenzo, Ron Essig, Craig Bonds, Robin L. De-Bruyne, Mark A. Kaemingk, Martha E. Mather, Christopher Myrick, Quinton E. Phelps, Trent M. Sutton, and James R. Triplett. Are we preparing the next generation of fisheries professionals to succeed in their careers?: a survey of AFS members. Fisheries, 41(8):436–449, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1199218.

Grossman:2016:IAF

[2005] Gary D. Grossman, Donald J. Orth, and Jason R. Neuswanger. Innovative approaches to fisheries education and outreach. Fisheries, 41 (8):450–457, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1204836.

Kaemingk:2016:ERU

[2006] Mark A. Kaemingk, Ron Essig, Steve L. McMullin, Craig Bonds, Robin L. DeBruyne, Christopher Myrick, Quinton E. Phelps, Trent M. Sutton, and James R. Triplett. Examining the relevancy and utility of the American Fisheries Society professional certification program to prepare future fisheries professionals. Fisheries, 41(8):458–461, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1199231.

Essig:2016:UFF

[2007] Ron Essig. U.S. Federal fish biologist educational requirements. Fisheries, 41(8):462, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1199834.

Sutton:2016:WDW

[2008] Trent M. Sutton, Katie N. Bertrand, James R. Jackson, Jeffrey C. Jolley, Quinton E. Phelps, James B. Reynolds, and Melissa R. Wuellner. Where do we go from here? ICE connects employers and educators to bring fisheries to the next level. Fisheries, 41(8):463, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1199831.

Hughes:2016:PIR

[2009] R. M. Hughes, C. B. M. Alves, F. Amezcua, J. Claussen, S. J. Cooke, I. G. Cowx, K. D. Fausch, M. T. Ferreira, W. G. Franzin, D. Hall, M. O. Ipinmoroti, R. Ligeiro, N. Mercado-Silva, M. Miller, P. S. Pompeu, D. C. Schmidt, L. G. M. Silva, and B. F. Terra. Preparing for international research and teaching. Fisheries, 41(8):464–465, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1199222.

Bonar:2016:BCS

[2010] Scott A. Bonar. Biological and communication skills needed for introduced fish biologists. Fisheries, 41(8):466–467, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1199223.

Jensen:2016:NAS

[2011] Gary Jensen, Michael Schwarz, Sandra Shumway, Jesse Trushenski, L. Curry Woods III, Thomas Broyles, and Maxwell Mayeaux. National assessment sheds light on educational needs for aquaculture in the United States. Fisheries, 41(8):467–469, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1199830.

Jackson:2016:UOF

[2012] James R. Jackson, David W. Willis, and Douglas L. Stang. Are universities offering fewer fisheries courses? Fisheries, 41(8):470–471, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1199226.

Lederman:2016:PNG

[2013] Nathan J. Lederman and Andrew K. Carlson. Preparing the next generation of fisheries professionals: Insights from the Student Subsection of the Education Section. Fisheries, 41(8):471–472, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1199227.

Terre:2016:PNG

[2014] David R. Terre. Preparing the next generation of fisheries professionals. Fisheries, 41(8):473, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1199230.

Colvin:2016:PFF

[2015] Michael E. Colvin and James T. Peterson. Preparing future fisheries professionals to make good decisions. Fisheries, 41(8):473–474, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1199233.

Hansen:2016:IKT

[2016] Gretchen J. A. Hansen. If I knew then what I know now: Advice for fisheries students from recent fisheries graduates. Fisheries, 41(8): 475–476, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1199229.

Jolley:2016:NGF

[2017] Jeffrey C. Jolley. The next generation of fisheries professionals —- understanding generations. Fisheries, 41(8):476, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1199835.

Dunmall:2016:NBD

[2018] Karen M. Dunmall and Steven J. Cooke. Narrow or broad: Diverse academic pathways to a career in fisheries science and management. Fisheries, 41(8):477–478, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2016.1199228.

Kopaska:2016:FIT

[2019] Jeff Kopaska, Julie M. Defilippi, Keith Hurley, Rebecca M. Krogman, and Andrew Loftus. The fisheries information and technology section and the organization of fish and wildlife information managers guide to swimming in the ocean of information. Fisheries, 41(8):478–480, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1199833.

Winans:2016:NTL

[2020] Gary A. Winans, Jon Baker, and Jennifer Hammond. NOAA’s teacher in the laboratory program — Northwest Fisheries Science Center. Fisheries, 41(8):481–483, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1199224.

Murphy:2016:CRF

[2021] Margaret Murphy, Bob Hughes, Carlos Alves, Joao Oliveira, Don Mac-Donald, Danielle Reich, J. Fred Heitman, and Doug Bradley. Consultants’ role in fisheries (is there really a dark side?). Fisheries, 41 (8):484–487, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1204831.

Rowell:2016:UFW

[2022] Denise Rowell. U.S. Fish and Wildlife Service’s Alabama Field Office Hutton Scholars: Where are they now? Fisheries, 41(8): 488–489, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1211438.

Beard:2016:WFC

[2023] Beth Beard. World Fisheries Congress wrap-up. Fisheries, 41(8): 490, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1215177.

Phelps:2016:WMK

[2024] Quinton E. Phelps and Sara Tripp. What not to miss at the 2016 KC Annual Meeting. Fisheries, 41(8):492, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1203176.

Anonymous:2016:OFF

[2025] Anonymous. Otto F. Fajen 1927–2016. Fisheries, 41(8):493, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1199836.

Houde:2016:OLL

[2026] Edward D. Houde. Oneida Lake: Long-term dynamics of a managed ecosystem and its fishery. Fisheries, 41(8):494, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1199841.

Anonymous:2016:NAJc

[2027] Anonymous. North American Journal of Aquaculture volume 78, number 2, April 2016. Fisheries, 41(8):495, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1199838.

Anonymous:2016:FEh

[2028] Anonymous. Fisheries events. Fisheries, 41(8):497, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1199839.

Sopinka:2016:BPPf

[2029] Natalie Sopinka. Back page photo series: Reflector. Fisheries, 41 (8):499–500, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1199837.

Anonymous:2016:FIPi

[2030] Anonymous. Full issue PDF, volume 41, issue 9. Fisheries, 41(9): 501–560, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1227614.

Margraf:2016:WHW

[2031] Joe Margraf. Well, here we go! Fisheries, 41(9):503, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1210948.

Pool:2016:FFE

[2032] Taylor Pool. For forage fish, In dubio pro natura (when in doubt, favor nature). Fisheries, 41(9):504, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1210967.

White:2016:LE

[2033] Ray J. White. Letter to the Editor. Fisheries, 41(9):505, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1210950.

Wilson:2016:GNA

[2034] Martha Wilson. The good news about science communication in the social media age. Fisheries, 41(9):506, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1220224.

Good:2016:EO

[2035] Molly J. Good, William W. Taylor, John Beck, Robert Lambe, and John Robertson. On the evolution of organizations. Fisheries, 41(9): 507–510, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1220222.

Million:2016:DFB

[2036] Kara Million. Darter festival and BioBlitz raise public awareness of three endangered fish species in Alabama. Fisheries, 41(9):511–512, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1210953.

Anonymous:2016:AFSa

[2037] Anonymous. American Fisheries Society 147th Annual Meeting Tampa, Florida August 20–24, 2017. Fisheries, 41(9):513–514, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1210949.

Contreras:2016:BHA

[2038] G. P. Contreras and R. M. Hughes. A brief history of the AFS International Fisheries Section 1987–2016: Past, current, and future bridgebuilding. Fisheries, 41(9):515–518, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1208091.

Linh:2016:DUA

[2039] Vo Le Gia Linh. Describing urban artisanal fisheries in Vietnam’s Lower Mekong River: a student’s ventures into fisheries science. Fisheries, 41 (9):519–523, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1210966.

Ward:2016:UMS

[2040] Hillary G. M. Ward, Micheal S. Allen, Edward V. Camp, Nick Cole, Len M. Hunt, Bryan Matthias, John R. Post, Kyle Wilson, and Robert Arlinghaus. Understanding and managing social-ecological feedbacks in spatially structured recreational fisheries: The overlooked behavioral dimension. Fisheries, 41(9):524–535, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1207632.

Midway:2016:TFS

[2041] Stephen R. Midway, Tyler Wagner, Joseph D. Zydlewski, Brian J. Irwin, and Craig P. Paukert. Transboundary fisheries science: Meeting the challenges of inland fisheries management in the 21st century. Fisheries, 41(9):536–546, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1208090.

Tsang:2016:PSS

[2042] Yin-Phan Tsang, Dana M. Infante, Jana Stewart, Lizhu Wang, Ralph W. Tingly III, Darren Thornbrugh, Arthur R. Cooper, and Wesley M. Daniel. StreamThermal: a software package for calculating thermal metrics from stream temperature data. Fisheries, 41(9):548–554, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1210517.

Anonymous:2016:JAAc

[2043] Anonymous. Journal of Aquatic Animal Health, volume 28, number 2, June 2016. Fisheries, 41(9):555, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1210961.

Anonymous:2016:FEi

[2044] Anonymous. Fisheries events. Fisheries, 41(9):558, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2016.1210960.

Sopinka:2016:BPPg

[2045] Natalie Sopinka. Back page photo series: Pickled. Fisheries, 41(9): 559–560, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1210958.

Anonymous:2016:FIPj

[2046] Anonymous. Full issue PDF, volume 41, issue 10. Fisheries, 41(10): 561–616, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1239463.

Margraf:2016:MH

[2047] Joe Margraf. It’s much harder. Fisheries, 41(10):563, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1222133.

Bigford:2016:IA

[2048] Thomas E. Bigford. Inspirational actions. Fisheries, 41(10):564, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1222136.

Austen:2016:PBB

[2049] Douglas Austen. Put a bow on this blue-ribbon idea. Fisheries, 41 (10):565, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1229970.

Good:2016:DFL

[2050] Molly J. Good, William W. Taylor, John Robertson, William Demmer, and Jordan Burroughs. Developing future leaders. Fisheries, 41 (10):566–568, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1229971.

Anonymous:2016:AFSb

[2051] Anonymous. American Fisheries Society 147th Annual Meeting Tampa, Florida August 20–24, 2017: Second Call for Papers. Fisheries, 41 (10):569–570, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1221705.

Lennox:2016:BP

[2052] Robert J. Lennox. Bonefish paradise. Fisheries, 41(10):571–572, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1225477.

Watson:2016:WDN

[2053] Nicole M. Watson. Wild: Determining natal origins of juvenile steelhead using otolith chemical analysis. Fisheries, 41(10):572–573, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1229969.

Carlson:2016:FHH

[2054] Andrew K. Carlson. In the footsteps of a heroine: Honoring Janice Lee Fenske. Fisheries, 41(10):574–575, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1221704.

Anonymous:2016:SBM

[2055] Anonymous. Steven Berkeley Marine Conservation Fellowship: 2016 winners and call for 2017 applications. Fisheries, 41(10):576–577, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1225478.

Miranda:2016:FPL

[2056] L. E. Miranda. Fishes in paleochannels of the Lower Mississippi River alluvial valley: a national treasure. Fisheries, 41(10):578–588, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1219949.

Frimpong:2016:PID

[2057] Emmanuel A. Frimpong, Jian Huang, and Yu Liang. IchthyMaps: a database of historical distributions of freshwater fishes of the United States. Fisheries, 41(10):590–599, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1219948.

Hupfeld:2016:MRB

[2058] Ryan N. Hupfeld, Quinton E. Phelps, Sara J. Tripp, and David P. Herzog. Mississippi River Basin paddlefish population dynamics: Implications for the management of a highly migratory species. Fisheries, 41 (10):600–610, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1221701.

Anonymous:2016:NAJd

[2059] Anonymous. North American Journal of Fisheries Management, volume 36, number 3, June 2016. Fisheries, 41(10):613, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1222138.

Anonymous:2016:FEj

[2060] Anonymous. Fisheries events. Fisheries, 41(10):614, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1222139.

Sopinka:2016:BPPh

[2061] Natalie Sopinka. Back page photo series: Suspension. Fisheries, 41 (10):615–616, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1221703.

Anonymous:2016:ASD

[2062] Anonymous. 2017 AFS Southern Division Spring Meeting. Fisheries, 41(10):616, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1221706.

Anonymous:2016:FIPk

[2063] Anonymous. Full issue PDF, volume 41, issue 11. Fisheries, 41(11): 617–660, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1255481.

Margraf:2016:WWW

[2064] Joe Margraf. Who and what are we? Fisheries, 41(11):619, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1240504.

Bigford:2016:UNI

[2065] Thomas E. Bigford. The urgent need for intern opportunities. Fisheries, 41(11):620, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1240500.

Pearsons:2016:EFP

[2066] Todd N. Pearsons, Thomas J. Dresser, Peter J. Graf, Michael S. Nicholls, Shannon Lowry, Michael C. Clement, and Curtis L. Dotson. Emergency fish passage at a fractured mainstem Columbia River dam. Fisheries, 41 (11):622–629, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1232953.

Brodeur:2016:EEC

[2067] Richard D. Brodeur, Jason S. Link, B. E. Smith, M. D. Ford, D. R. Kobayashi, and T. T. Jones. Ecological and economic consequences of ignoring jellyfish: a plea for increased monitoring of ecosystems. Fisheries, 41(11):630–637, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1232964.

Evans:2016:GFM

[2068] Mary Anne Evans. Graphical function mapping as a new way to explore cause-and-effect chains. Fisheries, 41(11):638–643, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1221404.

Lynch:2016:SVM

[2069] Abigail J. Lynch, William W. Taylor, and Aaron M. McCright. Stakeholder views of management and decision support tools to integrate climate change into Great Lakes lake whitefish management. Fisheries, 41 (11):644–652, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1232960.

Anonymous:2016:AFSc

[2070] Anonymous. American Fisheries Society Western Division Meeting: Call for papers. Fisheries, 41(11):652, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1246312.

Anonymous:2016:PCN

[2071] Anonymous. Paul C. Neth, 1930–2016. Fisheries, 41(11):653, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1240519.

Anonymous:2016:AMAa

[2072] Anonymous. 147th Annual Meeting of the American Fisheries Society: Third call for papers. Fisheries, 41(11):654–655, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1232950.

Anonymous:2016:TAFd

[2073] Anonymous. Transactions of the American Fisheries Society, volume 145, number 4, July 2016. Fisheries, 41(11):656, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1232966.

Anonymous:2016:FEk

[2074] Anonymous. Fisheries events. Fisheries, 41(11):657, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1232967.

Sopinka:2016:BPPi

[2075] Natalie Sopinka. Back page photo series: Boundaries. Fisheries, 41 (11):659–660, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1232940.

Anonymous:2016:FIPl

[2076] Anonymous. Full issue PDF, volume 41, issue 12. Fisheries, 41(12): 709–800, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
03632415.2016.1266879.

Margraf:2016:HDG

[2077] Joe Margraf. How did I get here? Fisheries, 41(12):711, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1262148.

Coakley:2016:IHE

[2078] Jessica M. Coakley and Christopher M. Moore. Incorporating habitat into ecosystem-based fisheries management strategies. Fisheries, 41(12):712, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1245980.

Shamaskin:2016:EPH

[2079] Andrew Shamaskin. The end parenthesis: How a fisheries student tackled programming. Fisheries, 41(12):714–715, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1252624.

Wiedenhoft:2016:AWL

[2080] Heather Wiedenhoft. Amber waves of … algae? Fisheries, 41(12): 716–719, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1252623.

Railsback:2016:WIT

[2081] Steven F. Railsback. Why it is time to put PHABSIM out to pasture. Fisheries, 41(12):720–725, 2016. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2016.1245991.

DiStefano:2016:EOS

[2082] Robert J. DiStefano, Ronald A. Reitz, and Emily M. Imhoff. Examining one state’s regulation development process to manage alien crayfish introductions. Fisheries, 41(12):726–737, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1246871.

Bowerman:2016:PSP

[2083] Tracy Bowerman, Matthew L. Keefer, and Christopher C. Caudill. Pacific salmon prespawn mortality: Patterns, methods, and study design considerations. Fisheries, 41(12):738–749, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1245993.

OShaughnessy:2016:CLE

[2084] Eric O’Shaughnessy, Matthew Landi, Stephanie R. Januchowski-Hartley, and Matthew Diebel. Conservation leverage: Ecological design culverts also return fiscal benefits. Fisheries, 41(12):750–757, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1246875.

Jiorle:2016:AUS

[2085] Ryan P. Jiorle, Robert N. M. Ahrens, and Micheal S. Allen. Assessing the utility of a smartphone app for recreational fishery catch data. Fisheries, 41(12):758–766, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1249709.

Anonymous:2016:AMAb

[2086] Anonymous. 147th Annual Meeting of the American Fisheries Society: Final call for papers. Fisheries, 41(12):767–768, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1245986.

Anonymous:2016:HJF

[2087] Anonymous. Hutton Junior Fisheries Biology Program. Fisheries, 41 (12):769, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1262172.

Anonymous:2016:FEW

[2088] Anonymous. Four easy ways to support the American Fisheries Society. Fisheries, 41(12):770, 2016. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2016.1262173.

Beard:2016:AMK

[2089] Beth Beard. 2016 146th Annual Meeting in Kansas City: Show me another great meeting in Missouri. Fisheries, 41(12):771–784, 2016. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2016.1245984.

Anonymous:2016:AFSd

[2090] Anonymous. American Fisheries Society awards for 2016. Fisheries, 41 (12):785–792, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1245974.

Anonymous:2016:CAA

[2091] Anonymous. Call for 2017 AFS award nominations. Fisheries, 41 (12):793–794, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1245973.

Anonymous:2016:AIF

[2092] Anonymous. Annual index Fisheries volume 41 2016. Fisheries, 41 (12):795, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1245976.

Sheaves:2016:MFH

[2093] Marcus Sheaves. Mangroves as fish habitat. Fisheries, 41(12): 796, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1245561.

Anonymous:2016:NAJe

[2094] Anonymous. North American Journal of Aquaculture, volume 78, number 4, October 2016. Fisheries, 41(12):797, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2016.1245977.

Anonymous:2016:FEl

[2095] Anonymous. Fisheries events. Fisheries, 41(12):798, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2016.1245978.

Sopinka:2016:BPPj

[2096] Natalie Sopinka. Back page photo series: In flux. Fisheries, 41(12): 799–800, 2016. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1245982.

Anonymous:2017:FIPa

[2097] Anonymous. Full issue PDF, volume 42, issue 1. Fisheries, 42(1): 1–72, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1286112.

Margraf:2017:VIC

[2098] Joe Margraf. The value of international connections. Fisheries, 42 (1):3, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1262677.

Pool:2017:FNF

[2099] Taylor Pool. Future of the nation’s fisheries and aquatic resources: Challenges we face in 2017 and beyond. Fisheries, 42(1):4, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1259942.

Roberts:2017:NNF

[2100] Ryan Roberts. New national fish habitat assessment report provides key information for fish conservation planning. Fisheries, 42 (1):5–6, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1262674.

Bigford:2017:AFH

[2101] Thomas E. Bigford. The AFS Fish Habitat Section — still young and still growing. Fisheries, 42(1):8, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1259941.

Troia:2017:WSY

[2102] Matthew J. Troia. What should your state fish be? Improving public awareness of endemic fishes in the United States. Fisheries, 42 (1):9–15, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1259940.

Miller:2017:NAG

[2103] Michael J. Miller and Jun Aoyama. A note about the Great East Japan Earthquake and Tsunami in 2011 from a fisheries perspective five years later. Fisheries, 42(1):16–18, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1262670.

Lawson:2017:LBR

[2104] Julia M. Lawson, Sarah J. Foster, and Amanda C. J. Vincent. Low bycatch rates add up to big numbers for a genus of small fishes. Fisheries, 42(1):19–33, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1259944.

Lawson:2017:GSS

[2105] Julia M. Lawson. The global search for seahorses in bycatch. Fisheries, 42(1):34–39, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1273660.

Vincent:2017:SPE

[2106] Amanda C. J. Vincent and Sarah J. Foster. Setting precedent in export regulations for marine fishes with seahorses. Fisheries, 42(1): 40–43, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1273658.

Croce:2017:EDN

[2107] Patrick Della Croce, Geoffrey C. Poole, Robert A. Payn, and Robert E. Gresswell. Early detection of nonnative alleles in fish populations: When sample size actually matters. Fisheries, 42(1):44–56, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1259947.

Moody:2017:PPB

[2108] Allison T. Moody, Thomas M. Neeson, Steve Wangen, Jeff Dischler, Matthew W. Diebel, Austin Milt, Matthew Herbert, Mary Khoury, Eugene Yacobson, Patrick J. Doran, Michael C. Ferris, Jesse R. O’Hanley, and Peter B. McIntyre. Pet project or best project? Online decision support tools for prioritizing barrier removals in the Great Lakes and beyond. Fisheries, 42(1):57–65, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2016.1263195.

Anonymous:2017:AMA

[2109] Anonymous. 147th Annual Meeting of the American Fisheries Society August 20–24, 2017. Fisheries, 42(1):67, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2017.1259937.

Anonymous:2017:FE

[2110] Anonymous. Fisheries events. Fisheries, 42(1):68, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2017.1267989.

Anonymous:2017:JAAa

[2111] Anonymous. Journal of Aquatic Animal Health, volume 28, number 4, December 2016. Fisheries, 42(1):69, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2017.1267987.

Waldman:2017:EFL

[2112] John Waldman. An eel fishing legacy worth remembering. Fisheries, 42(1):71–72, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1259939.

Anonymous:2017:FIPb

[2113] Anonymous. Full issue PDF, volume 42, issue 2. Fisheries, 42(2):73– 132, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1295730.

Margraf:2017:DEA

[2114] Joe Margraf. Diversity and the ecology of AFS. Fisheries, 42 (2):75, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1281057.

Bigford:2017:POA

[2115] Thomas E. Bigford. Pursuing opportunities as an AFS Liaison. Fisheries, 42(2):76, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1276327.

Gillespie:2017:AOP

[2116] Nathaniel Gillespie and Joseph Encinas. An aquatic organism passage story map. Fisheries, 42(2):77–79, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1276325.

Mitro:2017:SPW

[2117] Matthew Mitro. Science, politics, and wild trout management at the 12th International Wild Trout Symposium. Fisheries, 42(2):81–82, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1276324.

Pennock:2017:BGP

[2118] Casey A. Pennock. Beautification of Great Plains rivers: a perspective on the use and appreciation of aquatic resources. Fisheries, 42 (2):83–87, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1280284.

Anonymous:2017:SDA

[2119] Anonymous. Submission deadline for all paper and poster abstracts is March 17th! Fisheries, 42(2):88, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1285170.

Evans:2017:CCE

[2120] Nathan T. Evans, Patrick D. Shirey, Jamin G. Wieringa, Andrew R. Mahon, and Gary A. Lamberti. Comparative cost and effort of fish distribution detection via environmental DNA analysis and electrofishing. Fisheries, 42(2):90–99, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1276329.

Ward:2017:LTF

[2121] David L. Ward, Andrew F. Casper, Timothy D. Counihan, Jennifer M. Bayer, Ian R. Waite, John J. Kosovich, Colin G. Chapman, Elise R. Irwin, Jennifer S. Sauer, Brian S. Ickes, and Alexa J. McKerrow. Longterm fish monitoring in large rivers: Utility of “Benchmarking” across basins. Fisheries, 42(2):100–114, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1276330.

Lynch:2017:GCM

[2122] Abigail J. Lynch, Steven J. Cooke, T. Douglas Beard, Jr., Yu-Chun Kao, Kai Lorenzen, Andrew M. Song, Micheal S. Allen, Zeenatul Basher, David B. Bunnell, Edward V. Camp, Ian G. Cowx, Jonathan A. Freedman, Vivian M. Nguyen, Joel K. Nohner, Mark W. Rogers, Zachary A. Siders, William W. Taylor, and So-Jung Youn. Grand challenges in the management and conservation of North American inland fishes and fisheries. Fisheries, 42(2):115–124, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1259945.

Wilson:2017:GAC

[2123] Chris Wilson. Genetic aspects of climate change influences on inland fishes and fisheries. Fisheries, 42(2):125–126, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2017.1259946.

Anonymous:2017:CWa

[2124] Anonymous. Conferences and workshops. Fisheries, 42(2):128, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1284522.

Anonymous:2017:TAFa

[2125] Anonymous. Transactions of the American Fisheries Society, volume 146, issue 1, January 2017. Fisheries, 42(2):129, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1284521.

Anonymous:2017:AIFa

[2126] Anonymous. 2017 AFS International Fisheries Section Fellow Award: Call for applicants. Fisheries, 42(2):131, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2017.1292088.

Wipfli:2017:WSB

[2127] Mark S. Wipfli. Where salmon, bears, and people gather. Fisheries, 42(2):132, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1281665.

Anonymous:2017:FIPc

[2128] Anonymous. Full issue PDF, volume 42, issue 3. Fisheries, 42(3):133– 180, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1301737.

Margraf:2017:SRU

[2129] Joe Margraf. Sustainability R us. Fisheries, 42(3):135, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1284496.

Bigford:2017:ASN

[2130] Thomas E. Bigford. AFS shifts to new approach in policy. Fisheries, 42(3):136, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1284497.

Atkins:2017:SAM

[2131] Rebecca L. Atkins. Science and art: a mutualism. Fisheries, 42(3): 137–139, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1284499.

Anonymous:2017:WFT

[2132] Anonymous. W. F. Thompson Award for best student paper published in 2015. Fisheries, 42(3):140, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1293961.

Anonymous:2017:GMA

[2133] Anonymous. 2016 Golden Member awardees. Fisheries, 42(3): 141, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1299990.

Cooke:2017:CEB

[2134] Steven J. Cooke, Sean Wesch, Lisa A. Donaldson, Alexander D. M. Wilson, and Neal R. Haddaway. A call for evidence-based conservation and management of fisheries and aquatic resources. Fisheries, 42 (3):143–149, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1276343.

Bonar:2017:SMS

[2135] Scott A. Bonar, Norman Mercado-Silva, Wayne A. Hubert, T. Douglas Beard, Jr., Göran Dave, Jan Kubecka, Brian D. S. Graeb, Nigel P. Lester, Mark Porath, and Ian J. Winfield. Standard methods for sampling freshwater fishes: Opportunities for international collaboration. Fisheries, 42 (3):150–156, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1276352.

Deslauriers:2017:HIE

[2136] David Deslauriers. How to incorporate experiential learning in the classroom to address complex problems in ecology. Fisheries, 42(3): 157–162, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1276354.

Gibson:2017:SNC

[2137] Robert John Gibson. Salient needs for conservation of Atlantic salmon. Fisheries, 42(3):163–174, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2016.1276331.

Haedrich:2017:RJG

[2138] Richard L. Haedrich and Patricia Gallaugher. Robert John Gibson. Fisheries, 42(3):175, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2016.1246888.

Kane:2017:TPS

[2139] Douglas D. Kane. Thomas Paul Simon. Fisheries, 42(3):176, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1293964.

Anonymous:2017:TRA

[2140] Anonymous. Top 6 reasons to attend. Fisheries, 42(3):177, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1284507.

Anonymous:2017:CWb

[2141] Anonymous. Conferences and workshops. Fisheries, 42(3):178, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1284509.

Anonymous:2017:NAJa

[2142] Anonymous. North American Journal of Aquaculture, volume 79, number 1, January 2017. Fisheries, 42(3):179, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2017.1284511.

Anonymous:2017:MM

[2143] Anonymous. Mountain migrators. Fisheries, 42(3):180, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1284510.

Anonymous:2017:FIPd

[2144] Anonymous. Full issue PDF, volume 42, issue 4. Fisheries, 42(4):181– 244, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1310486.

Bonar:2017:DSS

[2145] Scott A. Bonar. The dark side of safety. Fisheries, 42(4):183, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1297113.

Bigford:2017:AEF

[2146] Thomas E. Bigford. Adding ecosystems to fisheries management. Fisheries, 42(4):184, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1294901.

Wood:2017:EWP

[2147] Nicole Wood. The engaging world of periscope: Live-streaming science to the masses. Fisheries, 42(4):185–186, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2017.1294900.

Dippold:2017:MYM

[2148] David A. Dippold, Grant D. Adams, Troy M. Farmer, and Benjamin A. Staton. Maximize your meeting: a student’s guide to AFS meetings. Fisheries, 42(4):187–189, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1288472.

McFarland:2017:TUP

[2149] Jason McFarland. Thriving in unexpected places: Food webs supporting Arctic grayling in Alaska’s Far North. Fisheries, 42(4): 190–196, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1301726.

Sass:2017:IFH

[2150] Greg G. Sass, Andrew L. Rypel, and Joshua D. Stafford. Inland fisheries habitat management: Lessons learned from wildlife ecology and a proposal for change. Fisheries, 42(4):197–209, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1276344.

Hightower:2017:EFM

[2151] Joseph E. Hightower and Julianne E. Harris. Estimating fish mortality rates using telemetry and multistate models. Fisheries, 42(4): 210–219, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1276347.

Bett:2017:CCS

[2152] Nolan N. Bett, Scott G. Hinch, Nicholas J. Burnett, Michael R. Donaldson, and Sean M. Naman. Causes and consequences of straying into small populations of Pacific salmon. Fisheries, 42(4):220–230, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1276356.

Murphy:2017:ASV

[2153] Brian R. Murphy. AFS Second Vice President candidate statement: Brian R. Murphy. Fisheries, 42(4):231–232, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1297645.

Whelan:2017:ASV

[2154] Gary Whelan. AFS Second Vice President Candidate Statement: Gary Whelan. Fisheries, 42(4):233–234, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1304787.

Anonymous:2017:WTA

[2155] Anonymous. Welcome to Tampa: Annual Meeting preview. Fisheries, 42(4):235–240, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1297138.

Anonymous:2017:CWc

[2156] Anonymous. Conferences and workshops. Fisheries, 42(4):242, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1297093.

Anonymous:2017:NAJb

[2157] Anonymous. North American Journal of Fisheries Management, volume 37, number 1, February 2017. Fisheries, 42(4):243, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1297092.

Martin:2017:AD

[2158] Rene P. Martin. Art of the deep. Fisheries, 42(4):244, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1296737.

Anonymous:2017:FIPe

[2159] Anonymous. Full issue PDF, volume 42, issue 5. Fisheries, 42(5):245– 296, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1326221.

Jackson:2017:WT

[2160] Donald C. Jackson. Wrinkles in time. Fisheries, 42(5):247, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1311108.

Bigford:2017:RSO

[2161] Thomas E. Bigford. Revisiting (or sharpening) our policy mandate. Fisheries, 42(5):248, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1305200.

Walton:2017:BBE

[2162] Sarah Walton. Budding biologists + experienced ecologists = successful conference! Fisheries, 42(5):249, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1305204.

Keefe:2017:AS

[2163] Alicia Keefe, Emily Bosanquet, and Katherine Cheney. The art of sustainability. Fisheries, 42(5):250–253, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2017.1305206.

Schmid:2017:NRL

[2164] Kurt Schmid and Tommaso Giarrizzo. New record of a largetooth sawfish specimen from the Amazon River Estuary in Northern Brazil. Fisheries, 42(5):254–255, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1276357.

Bonfil:2017:FWA

[2165] Ramón Bonfil, Oscar Uriel Mendoza-Vargas, Melina Ricaño-Soriano, Paola Yanira Palacios-Barreto, and Nataly Bolaño-Martínez. Former widespread abundance and recent downfall of sawfishes in Mexico as evidenced by historical photographic and trophy records. Fisheries, 42 (5):256–259, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1276339.

Heim:2017:FMU

[2166] Kurt Heim. Fish movement under the Midnight Sun: Behavioral adaptions to a harsh aquatic environment. Fisheries, 42(5):260– 267, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1311105.

Clemens:2017:CCR

[2167] Benjamin J. Clemens, Richard J. Beamish, Kelly C. Coates, Margaret F. Docker, Jason B. Dunham, Ann E. Gray, Jon E. Hess, Jeffrey C. Jolley, Ralph T. Lampman, Brian J. McIlraith, Mary L. Moser, Joshua G. Murauskas, David L. G. Noakes, Howard A. Schaller, Carl B. Schreck, Steven J. Starcevich, Bianca Streif, Stan J. van de Wetering, Joy Wade, Laurie A. Weitkamp, and Lance A. Wyss. Conservation challenges and research needs for Pacific lamprey in the Columbia River Basin. Fisheries, 42(5):268–280, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2017.1305857.

Kilfoil:2017:TCR

[2168] James P. Kilfoil, Bradley M. Wetherbee, John K. Carlson, and Dewayne A. Fox. Targeted catch-and-release of prohibited sharks: Sand tigers in coastal Delaware waters. Fisheries, 42(5):281–287, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1306974.

Anonymous:2017:AFSa

[2169] Anonymous. American Fisheries Society Annual Meeting 2017 continuing education schedule. Fisheries, 42(5):288–290, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1305209.

Cullis-Suzuki:2017:SSR

[2170] Sarika Cullis-Suzuki. Sonic Sea reveals whales are at risk in an ocean of noise … but what about other sea life? Fisheries, 42(5):291, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1305856.

Anonymous:2017:CWd

[2171] Anonymous. Conferences and workshops. Fisheries, 42(5):294, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1311100.

Anonymous:2017:JAAb

[2172] Anonymous. Journal of Aquatic Animal Health volume 29, number 1, March 2017. Fisheries, 42(5):295, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1311099.

Schnell:2017:FFF

[2173] Nalani Schnell and G. David Johnson. The first functional fish head joint discovered in deep sea dragonfishes. Fisheries, 42(5):296, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1305195.

Anonymous:2017:FIPf

[2174] Anonymous. Full issue PDF, volume 42, issue 6. Fisheries, 42(6):297– 344, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1335525.

Margraf:2017:TS

[2175] Joe Margraf. Is there a solution? Fisheries, 42(6):299, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1330601.

Marshak:2017:HSE

[2176] Anthony R. Marshak and Stephen K. Brown. Habitat science is an essential element of ecosystem-based fisheries management. Fisheries, 42(6):300, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1323523.

Bigford:2017:RFRa

[2177] Thomas E. Bigford. Rivers and fish, not rivers versus fish. Fisheries, 42(6):301, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1317140.

Yang:2017:RRF

[2178] Zi jiang Yang, Yushun Chen, Dongchun Wang, Long teng Liu, Cong Liu, Robert M. Hughes, and Yadan Liu. Responsible recreational fisheries: a Chinese perspective. Fisheries, 42(6):303–307, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1324703.

AFCS:2017:CFI

[2179] AFS Fish Culture Section. Conservation Fisheries, Inc. Fisheries, 42 (6):308–311, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1319210.

Wiedenhoft:2017:BAE

[2180] Heather Wiedenhoft. Biofiltration: an attractive extractive solution. Fisheries, 42(6):312–314, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1305854.

Leppi:2017:SAB

[2181] Jason Leppi, Michael Lunde, Mark Wipfli, and Dan Rinella. In search of Arctic bonefish. Fisheries, 42(6):315–319, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1323525.

Sitar:2017:UVR

[2182] Shawn P. Sitar and Daniel J. Traynor. Using voice recognition software to collect fisheries biodata: a proof of concept. Fisheries, 42 (6):320–326, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1314149.

Clancy:2017:CAH

[2183] Niall G. Clancy. Can amphibians help conserve native fishes? Fisheries, 42(6):327–331, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1314150.

Mawdsley:2017:ABU

[2184] Jonathan Mawdsley, Davia Palmeri, and Mark Humpert. Aquatic biodiversity in the U.S. state wildlife action plans. Fisheries, 42(6): 332–334, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1316148.

Rachlin:2017:BRF

[2185] Joseph W. Rachlin. Book review: The Fishes of Pennsylvania, by Jay R. Stauffer, Jr., Robert W. Criswell, and Douglas P. Fischer. Cichlid Press, El Paso, Texas. 2016. 556 pages. US$49.50 . Fisheries, 42(6):336, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1319208.

Anonymous:2017:CWe

[2186] Anonymous. Conferences and workshops. Fisheries, 42(6):340, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1326229.

Anonymous:2017:TAFb

[2187] Anonymous. Transactions of the American Fisheries Society, volume 146, number 2, March 2017. Fisheries, 42(6):341, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2017.1323529.

Gloor:2017:HC

[2188] Christian Gloor. A hairy couple. Fisheries, 42(6):344, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1324120.

Anonymous:2017:FIPg

[2189] Anonymous. Full issue PDF, volume 42, issue 7. Fisheries, 42(7):345– 388, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1348869.

Hughes:2017:ISI

[2190] Robert M. Hughes. Introduction to the special issue on ethics and advocacy. Fisheries, 42(7):347–349, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1330049.

Margraf:2017:DWA

[2191] Joe Margraf. Do we advocate? Fisheries, 42(7):350, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1338451.

Meyer:2017:KFB

[2192] Benjamin Meyer. The King of Fish “Book Drop”: Generating conversation and shared values about salmon in Alaska. Fisheries, 42 (7):351–352, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1330048.

Ruch:2017:ELS

[2193] Jeff Ruch. Emerging law of scientific integrity — a bumpy birth. Fisheries, 42(7):353–356, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2017.1330051.

Pool:2017:SAA

[2194] Taylor Pool, Thomas E. Bigford, and Owen Mulvey-McFerron. Shaping AFS advocacy based on history and need. Fisheries, 42(7): 357–360, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1330059.

Reiser:2017:SAA

[2195] Dudley W. Reiser. Science and advocacy in the American Fisheries Society: Seeking the same side of the fence. Fisheries, 42(7): 361–365, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1330057.

Fraidenburg:2017:AA

[2196] Michael E. Fraidenburg. Always advocating. Fisheries, 42(7):366– 368, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1330050.

Martin:2017:ECA

[2197] Lori M. Martin, Leanne H. Roulson, and Hilda Sexauer. Engagement and challenges in advocacy: Experiences from the Western United States. Fisheries, 42(7):369–372, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1330054.

Michael:2017:MSE

[2198] J. Hal Michael, Jr. Managing salmon for ecosystem needs in the Pacific Northwest: Limiting science input in ecosystem management — silos R us. Fisheries, 42(7):373–376, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1330055.

Burroughs:2017:TSW

[2199] Jim Burroughs and Shannon K. Brewer. The tangled sociopolitical web: How do we give science a stronger voice and make our decisions more transparent? Fisheries, 42(7):377–379, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2017.1276349.

Anonymous:2017:AFSb

[2200] Anonymous. American Fisheries Society: Annual Meeting 2017 plenary speakers. Fisheries, 42(7):380–381, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1338452.

Anonymous:2017:CWf

[2201] Anonymous. Conferences and workshops. Fisheries, 42(7):383, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1338453.

Anonymous:2017:NAJc

[2202] Anonymous. North American Journal of Aquaculture, volume 79, number 2, April 2017. Fisheries, 42(7):384, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1332300.

Scheerer:2017:OCR

[2203] Paul Scheerer. Oregon chub — a rare success story for the Endangered Species Act. Fisheries, 42(7):388, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1330047.

Anonymous:2017:FIPh

[2204] Anonymous. Full issue PDF, volume 42, issue 8. Fisheries, 42(8):389– 448, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1364518.

Margraf:2017:RMA

[2205] Joe Margraf. Rating myself as AFS President, or how well I think I did. Fisheries, 42(8):391, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1342467.

Winters:2017:MCS

[2206] Drue Winters. Making the case for science. Fisheries, 42(8): 393, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1355185.

Carlson:2017:TFI

[2207] Andrew K. Carlson, William W. Taylor, Jianguo Liu, and Ivan Orlic. The telecoupling framework: an integrative tool for enhancing fisheries management. Fisheries, 42(8):395–397, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2017.1342491.

Bonar:2017:SWA

[2208] Scott A. Bonar and Jesse Trushenski. The skill we all need. Fisheries, 42(8):398, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1355186.

McMullin:2017:NPA

[2209] Steve L. McMullin. Nine proposed action areas to enhance diversity and inclusion in the American Fisheries Society — an introduction. Fisheries, 42(8):399, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1345548.

Penaluna:2017:NPA

[2210] Brooke E. Penaluna, Ivan Arismendi, Christine M. Moffitt, and Zachary L. Penney. Nine proposed action areas to enhance diversity and inclusion in the American Fisheries Society. Fisheries, 42 (8):399–402, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1345549.

Anonymous:2017:DP

[2211] Anonymous. Diane Peebles. Fisheries, 42(8):403–405, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1342470.

Harrison:2017:QPP

[2212] Sarah Harrison and Paul Vecsei. Q&A with photographer Paul Vecsei. Fisheries, 42(8):406–408, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1342471.

Baumwell:2017:BRQ

[2213] Leah Baumwell. Behind the research: Q&A with artist and marine conservationist Guy Harvey. Fisheries, 42(8):409–412, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1363642.

Nagid:2017:EN

[2214] Eric Nagid. Eric Nagid. Fisheries, 42(8):413, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1342472.

Patterson:2017:WP

[2215] Will Patterson. Will Patterson. Fisheries, 42(8):414, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1346995.

Wegener:2017:MW

[2216] Matt Wegener. Matt Wegener. Fisheries, 42(8):415, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1346997.

Adams:2017:TFM

[2217] Aaron Adams. Tracking fish movements to inform conservation. Fisheries, 42(8):416–420, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2017.1342474.

Chagaris:2017:EBA

[2218] David Chagaris, Samantha Binion-Rock, Alex Bogdanoff, Kristen Dahl, Jennifer Granneman, Holden Harris, John Mohan, Merrill B. Rudd, Mary Kate Swenarton, Rob Ahrens, William F. Patterson III, James A. Morris, Jr., and Micheal Allen. An ecosystem-based approach to evaluating impacts and management of invasive lionfish. Fisheries, 42(8): 421–431, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1340273.

Stokesbury:2017:TIS

[2219] Kevin D. E. Stokesbury, Steven X. Cadrin, Nick Calabrese, Emily Keiley, Travis M. Lowery, Brian J. Rothschild, and Gregory R. DeCelles.

Towards an improved system for sampling New England groundfish using video technology. Fisheries, 42(8):432–439, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1342630.

Dahl:2017:TIH

[2220] Kristen Dahl and Will Patterson. Targeting an invader: The hunt for lionfish and management solutions. Fisheries, 42(8):440–444, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1348049.

Waldman:2017:RHB

[2221] John Waldman. Robert H. Boyle. Fisheries, 42(8):445, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1348065.

Anonymous:2017:CWg

[2222] Anonymous. Conferences and workshops. Fisheries, 42(8):446, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1353357.

Anonymous:2017:NAJd

[2223] Anonymous. North American Journal of Fisheries Management, volume 37, number 3, June 2017. Fisheries, 42(8):447, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1353358.

David:2017:SG

[2224] Solomon R. David. Spotted gar. Fisheries, 42(8):448, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1346999.

Anonymous:2017:FIPi

[2225] Anonymous. Full issue PDF, volume 42, issue 9. Fisheries, 42(9):449– 492, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1377479.

Mcmullin:2017:GMY

[2226] Steve L. Mcmullin. Getting more out of your AFS dues. Fisheries, 42(9):451, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1356118.

Bigford:2017:RFRb

[2227] Thomas E. Bigford. Rivers and fish, not rivers versus fish. Fisheries, 42(9):452, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1342473.

Youn:2017:IBP

[2228] So-Jung Youn, William W. Taylor, Jim Martin, and John Robertson. Interactions between politics and natural resource organizations: How do you address increasing politicization while maintaining technical expertise for sustainable resource management? Fisheries, 42 (9):453–455, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1356117.

Page:2017:TGL

[2229] Lawrence M. Page, Carole C. Baldwin, Héctor Espinosa-Pérez, Lloyd T. Findley, Carter R. Gilbert, Karsten E. Hartel, Robert N. Lea, Nicholas E. Mandrak, Juan J. Schmitter-Soto, and H. J. Walker, Jr. Taxonomy of Gila in the Lower Colorado River Basin of Arizona and New Mexico. Fisheries, 42(9):456–460, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1356108. See comment [2230].

Anonymous:2017:CTG

[2230] Anonymous. Comment to “Taxonomy of Gila in the Lower Colorado River Basin of Arizona and New Mexico”. Fisheries, 42(9): 460, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1356115. See [2229].

Walton-Rabideau:2017:PFS

[2231] Sarah Walton-Rabideau. Primitive fish series — bowfin Amia calva. Fisheries, 42(9):461, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2017.1358563.

Cox:2017:RAB

[2232] Kieran Cox and Jenna Bright. Raja Ampat: a biodiversity hot spot and the future of marine conservation. Fisheries, 42(9):462–467, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1356119.

Dorr:2017:DCC

[2233] Brian S. Dorr and David G. Fielder. Double-crested cormorants: Too much of a good thing? Fisheries, 42(9):468–477, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2017.1356121.

Lusardi:2017:TWT

[2234] Robert A. Lusardi and Peter B. Moyle. Two-way trap and haul as a conservation strategy for anadromous salmonids. Fisheries, 42(9): 478–487, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1356124.

Ross:2017:CFF

[2235] Stephen T. Ross. Conservation of freshwater fishes. Fisheries, 42 (9):488, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1369797.

Anonymous:2017:CWh

[2236] Anonymous. Conferences and workshops. Fisheries, 42(9):489, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1370261.

Anonymous:2017:C

[2237] Anonymous. Correction. Fisheries, 42(9):490, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1369800.

Anonymous:2017:JAAc

[2238] Anonymous. Journal of Aquatic Animal Health, volume 29, number 2, June 2017. Fisheries, 42(9):490, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1356161.

Anonymous:2017:AFSc

[2239] Anonymous. American Fisheries Society 148th Annual Meeting, Atlantic City, New Jersey August 19–23, 2018. Fisheries, 42(9):491, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1375740.

Boczek:2017:ID

[2240] Dustin Boczek. Iowa darter. Fisheries, 42(9):492, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03632415.2017.1358557.

Anonymous:2017:FIPj

[2241] Anonymous. Full issue PDF, volume 42, issue 10. Fisheries, 42(10): 493–560, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1394126.

Mcmullin:2017:CFS

[2242] Steve Mcmullin. Can fisheries scientists win the war on science? Fisheries, 42(10):495, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1358561.

Bigford:2017:RAM

[2243] Thomas E. Bigford. Regional approaches to marine resource management. Fisheries, 42(10):497, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1358559.

Link:2017:CAN

[2244] Jason Link. A conversation about NMFS’ ecosystem-based fisheries management policy and road map. Fisheries, 42(10):498–503, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1391596.

Crouse:2017:CFF

[2245] Curtis Crouse. The Conservation Fund’s Freshwater Institute. Fisheries, 42(10):504–507, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1368295.

Beecher:2017:CWI

[2246] Hal A. Beecher. Comment 1: Why it is time to put PHABSIM out to pasture. Fisheries, 42(10):508–510, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1380985.

Stalnaker:2017:CDT

[2247] Clair B. Stalnaker, Ian Chisholm, and Andrew Paul. Comment 2: Don’t throw out the baby (PHABSIM) with the bathwater: Bringing scientific credibility to use of hydraulic habitat models, specifically PHABSIM. Fisheries, 42(10):510–516, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1380986.

Railsback:2017:WIT

[2248] Steven F. Railsback. Why it is time to put PHABSIM out to pasture: Response to comments 1 and 2. Fisheries, 42(10):517–518, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1368296.

Cox:2017:UEL

[2249] Kieran Cox. The unique ecology of Lembeh Strait, Indonesia. Fisheries, 42(10):519–525, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1358558.

Dauwalter:2017:SAR

[2250] Daniel C. Dauwalter, Kurt A. Fesenmyer, Robin Bjork, Douglas R. Leasure, and Seth J. Wenger. Satellite and airborne remote sensing applications for freshwater fisheries. Fisheries, 42(10):526–537, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1357911.

Schoen:2017:FPS

[2251] Erik R. Schoen, Mark S. Wipfli, E. Jamie Trammell, Daniel J. Rinella, Angelica L. Floyd, Jess Grunblatt, Molly D. McCarthy, Benjamin E. Meyer, John M. Morton, James E. Powell, Anupma Prakash, Matthew N. Reimer, Svetlana L. Stuefer, Horacio Toniolo, Brett M. Wells, and Frank D. W. Witmer. Future of Pacific salmon in the face of environmental change: Lessons from one of the world’s remaining productive salmon regions. Fisheries, 42(10):538–553, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1374251.

Anonymous:2017:SBM

[2252] Anonymous. Steven Berkeley Marine Conservation Fellowship. Fisheries, 42(10):554–555, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1368298.

Anonymous:2017:CEa

[2253] Anonymous. Conferences and events. Fisheries, 42(10):556, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1370664.

Anonymous:2017:TAFc

[2254] Anonymous. Transactions of the American Fisheries Society, volume 146, number 4, July 2017. Fisheries, 42(10):557, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1370666.

Anonymous:2017:SCP

[2255] Anonymous. Second call for papers: American Fisheries Society 148th Annual Meeting, Atlantic City, New Jersey August 19–23, 2018. Fisheries, 42(10):559, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1391598.

Brooks:2017:TW

[2256] Jill Brooks and Steven Cooke. Tracking walleye. Fisheries, 42 (10):560, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1373544.

Anonymous:2017:FIPk

[2257] Anonymous. Full issue PDF, volume 42, issue 11. Fisheries, 42(11): 561–604, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1401806.

Mcmullin:2017:FAJ

[2258] Steve Mcmullin. The future of AFS journals. Fisheries, 42(11): 563, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1381529.

Bigford:2017:FNI

[2259] Thomas E. Bigford. Fish need infrastructure, too. Fisheries, 42 (11):564, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1379298.

White:2017:EOR

[2260] Gwen White. Expanding our reach: How AFS staff and units have advanced communications in just a few years. Fisheries, 42(11): 565–566, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1379827.

Schmidt:2017:SFI

[2261] Brian Schmidt, Dana Castle, Ed Roseman, Toomas Saat, and Hannu Lehtonen. Synopsis of the Fourth International Percid Fish Symposium. Fisheries, 42(11):567–570, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1377562.

Anonymous:2017:WPT

[2262] Anonymous. Washoe Park Trout Hatchery. Fisheries, 42(11):571– 573, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1379314.

Schramm:2017:FRM

[2263] Harold L. Schramm, Jr. The fishery resources of the Mississippi River: a model for conservation and management. Fisheries, 42(11): 574–585, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1377554.

Deslauriers:2017:PFB

[2264] David Deslauriers, Steven R. Chipps, James E. Breck, James A. Rice, and Charles P. Madenjian. Fish Bioenergetics 4.0: an R-based modeling application. Fisheries, 42(11):586–596, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2017.1377558.

Moffitt:2017:GJR

[2265] Christine M. Moffitt, Gary Wedemeyer, John Colt, Ron Garton, Chris West, and Steve Bouck. Gerald (Jerry) R. Bouck: August 22, 1934 to May 5, 2017. Fisheries, 42(11):597, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1377564.

Olden:2017:SFC

[2266] Julian D. Olden. Stream fish community dynamics — a critical synthesis. Fisheries, 42(11):598, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03632415.2017.1381530.

Anonymous:2017:CWi

[2267] Anonymous. Conferences and workshops. Fisheries, 42(11):599, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic).  URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1386913.

Anonymous:2017:NAJe

[2268] Anonymous. North American Journal of Aquaculture, volume 79, issue 3, July 2017. Fisheries, 42(11):600, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.com/

doi/full/10.1080/03632415.2017.1383845.

Anonymous:2017:CP

[2269] Anonymous. Call for papers. Fisheries, 42(11):601, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1383759.

Harrison:2017:PP

[2270] Sarah Harrison. The painted predator. Fisheries, 42(11):604, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1386914.

Anonymous:2017:FIPl

[2271] Anonymous. Full issue PDF, volume 42, issue 12. Fisheries, 42(12): 605–672, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1080/03632415.2017.1415590.

Mcmullin:2017:ARH

[2272] Steve Mcmullin. The AFS role in high-priority policy issues. Fisheries, 42(12):607, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1394774.

Bigford:2017:WDW

[2273] Thomas E. Bigford. Where do we go from here? Fisheries, 42 (12):608, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1391583.

Jawad:2017:SES

[2274] Laith A. Jawad. Shared experiences and suggestions. Fisheries, 42 (12):610–613, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1391579.

Gutierrez:2017:HCA

[2275] Nicolas L. Gutierrez. Harnessing citizenry awareness and technology to improve fisheries information: The power of data. Fisheries, 42 (12):615–620, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1383904.

Treanor:2017:UCD

[2276] Hilary B. Treanor, Andrew M. Ray, Megan Layhee, Barnaby J. Watten, Jackson A. Gross, Robert E. Gresswell, and Molly A. H. Webb. Using carbon dioxide in fisheries and aquatic invasive species management. Fisheries, 42(12):621–628, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1383903.

White:2017:BTS

[2277] Shannon White. In brook trout, sometimes the risk is worth the reward. Fisheries, 42(12):629, 2017. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03632415.2017.1396050.

Anonymous:2017:FIS

[2278] Anonymous. The First International Snakehead Symposium. Fisheries, 42(12):630, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1391571.

Taylor:2017:PTT

[2279] Thomas Taylor. Photography by Thomas Taylor. Fisheries, 42(12): 631–634, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1391581.

Beard:2017:AMT

[2280] Beth Beard. 147th Annual Meeting in Tampa: Eclipse highlights week of celebrating science. Fisheries, 42(12):635–636, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1394777.

Anonymous:2017:SS

[2281] Anonymous. Symposia summaries. Fisheries, 42(12):637–652, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1391585.

Anonymous:2017:AFSd

[2282] Anonymous. 2017 American Fisheries Society awards. Fisheries, 42 (12):653–662, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1391567.

Anonymous:2017:CAN

[2283] Anonymous. Call for award nominations: 2018 American Fisheries Society awards. Fisheries, 42(12):663–664, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/03632415.2017.1391577.

Moyle:2017:LAF

[2284] Peter B. Moyle. A life among fishes: The art of Gyotaku. Fisheries, 42(12):665, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1391575.

Anonymous:2017:CEb

[2285] Anonymous. Conferences and events. Fisheries, 42(12):666, 2017. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/03632415.2017.1398984.

Anonymous:2017:NAJf

[2286] Anonymous. North American Journal of Fisheries Management, volume 37, number 5, October 2017. Fisheries, 42(12):667, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/03632415.2017.1396526.

Anonymous:2017:AIFb

[2287] Anonymous. Annual index Fisheries volume 42 2017. Fisheries, 42 (12):668, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1396527.

Candidate:2017:B

[2288] Natalie Klinard M. Sc. Candidate. Bowfin. Fisheries, 42(12): 672, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
03632415.2017.1396569.

Anonymous:2018:IIa

[2289] Anonymous. Issue information. Fisheries, 43(1):1–2, January 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10053.

Anonymous:2018:FVNa

[2290] Anonymous. Fisheries volume 43 number 1 January 2018. Fisheries, 43 (1):1–68, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10001.

McMullin:2018:WAD

[2291] Steve L. McMullin. What is AFS doing about diversity? Fisheries, 43 (1):3, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10021.

Winters:2018:VAR

[2292] Drue Banta Winters. Be the voice for aquatic resources. Fisheries, 43 (1):4, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10026.

Harding:2018:PPN

[2293] Joel M. S. Harding, Douglas C. Braun, Nicholas J. Burnett, and Annika Putt. PITR: a new open source R package for PIT telemetry data. Fisheries, 43(1):5, January 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10027.

Varisco:2018:DCT

[2294] Martín Varisco, Julia Colombo, and Julio Vinuesa. Deadliest catch, Terra Australis edition. Fisheries, 43(1):6–10, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10016.

Fryxell:2018:COR

[2295] David C. Fryxell, Katie M. Kobayashi, Katherine N. McElroy, and Hayley Nuetzel. Creating outreach and restoration programs to promote the development of fisheries professionals: The primacy of cross-institutional collaboration. Fisheries, 43(1):11–14, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10020.

Anonymous:2018:JR

[2296] Anonymous. Journal reviews. Fisheries, 43(1):15–16, January 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10024.

Anonymous:2018:TQI

[2297] Anonymous. Ten questions for Isaac Szabo. Fisheries, 43(1):17–21, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10023.

Peiman:2018:OPS

[2298] Kathryn Peiman. An overview of Pacific salmon. Fisheries, 43(1):22– 25, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10025.

Powell:2018:WAF

[2299] James E. Powell, Mark S. Wipfli, Keith R. Criddle, and Erik R. Schoen. Will Alaska’s fisheries regime prove resilient? Kenai River fishery management as a model for adaptive governance. Fisheries, 43(1):26–30, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10022.

Stevens:2018:DME

[2300] Andrew L. Stevens, Ian G. Baird, and Peter B. McIntyre. Differences in mercury exposure among Wisconsin anglers arising from fish consumption preferences and advisory awareness. Fisheries, 43(1):31–41, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10013.

Moyle:2018:DSW

[2301] Peter B. Moyle, James A. Hobbs, and John R. Durand. Delta smelt and water politics in California. Fisheries, 43(1):42–50, January 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10014.

Young:2018:EDN

[2302] Nathan Young, Marianne Corriveau, Vivian M. Nguyen, Steven J. Cooke, and Scott G. Hinch. Embracing disruptive new science? Biotelemetry meets co-management in Canada’s Fraser River. Fisheries, 43(1):51– 60, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10015.

Anonymous:2018:CEa

[2303] Anonymous. Conferences and events. Fisheries, 43(1):66, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10035.

Anonymous:2018:JAAa

[2304] Anonymous. Journal of Aquatic Animal Health, volume 29, issue 4, December 2017. Fisheries, 43(1):67, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10036.

Wipfli:2018:BBS

[2305] Mark S. Wipfli. Beauty below the surface. Fisheries, 43(1):68, January 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10028.

Anonymous:2018:IIb

[2306] Anonymous. Issue information. Fisheries, 43(2):69–70, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10051.

Anonymous:2018:FVNb

[2307] Anonymous. Fisheries volume 43 number 2 February 2018. Fisheries, 43(2):69–116, February 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10002.

Brust:2018:TOM

[2308] Jeff Brust. Top 10 (OK, maybe 13) reasons you should go to the AFS Annual Meeting in Atlantic City. Fisheries, 43(2):71–72, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10033.

Winters:2018:POA

[2309] Drue Banta Winters. Policy opportunities abound in Atlantic City. Fisheries, 43(2):73, February 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10037.

Anonymous:2018:PIT

[2310] Anonymous. Photography interview with Todd Pearsons. Fisheries, 43 (2):75–78, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10040.

Grossman:2018:NUO

[2311] Gary D. Grossman and Kirill Y. Chernoff. The need and use of open educational resources in fisheries, environmental education, and conservation. Fisheries, 43(2):79–82, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10029.

Moran:2018:GMA

[2312] Paul Moran, Joe Dazey, Larrie LaVoy, and Sewall Young. Genetic mixture analysis supports recalibration of the Fishery Regulation Assessment Model. Fisheries, 43(2):83–97, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10017.

Honsey:2018:RTI

[2313] Andrew E. Honsey, Grace L. Loppnow, Timothy J. Martin, Leslie A. Schroeder, Megan M. Tomamichel, Nathan T. Huempfner, and Paul A. Venturelli. Reeling them in: Initial insight into the factors affecting presentation attendance at American Fisheries Society meetings. Fisheries, 43(2):98–105, February 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10018.

Murphy:2018:ASV

[2314] Margaret H. Murphy. AFS Second Vice President candidate statement: Margaret H. Murphy. Fisheries, 43(2):106–107, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10043.

Roulson:2018:ASV

[2315] Leanne H. Roulson. AFS Second Vice President candidate statement: Leanne H. Roulson. Fisheries, 43(2):108–109, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10042.

Anonymous:2018:GLL

[2316] Anonymous. Gary L. Larson, limnologist. Fisheries, 43(2):110–111, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10039.

Anonymous:2018:CEb

[2317] Anonymous. Conferences and events. Fisheries, 43(2):114, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10049.

Anonymous:2018:TAFa

[2318] Anonymous. Transactions of the American Fisheries Society, volume 147, issue 1, January 2018. Fisheries, 43(2):115, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10050.

Wipfli:2018:CI

[2319] Mark S. Wipfli. Colorful inside and out. Fisheries, 43(2):116, February 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10052.

Anonymous:2018:IIc

[2320] Anonymous. Issue information. Fisheries, 43(3):117–118, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10058.

Anonymous:2018:FVNc

[2321] Anonymous. Fisheries volume 43 number 3 March 2018. Fisheries, 43(3): 117–168, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10003.

McMullin:2018:AYS

[2322] Steve L. McMullin. AFS, at your service. Fisheries, 43(3):119, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10044.

Winters:2018:GAM

[2323] Drue Banta Winters. Grassroots advocacy: Make policy personal to bring home your message. Fisheries, 43(3):120, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10045.

Claussen:2018:WIA

[2324] Julie Claussen, Solomon David, Katherine O’Reilly, Edward Henry, and Natalie Sopinka. #Welcome: Introducing the AFS Science Communication Section. Fisheries, 43(3):121–122, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10046.

Shiffman:2018:SMF

[2325] David S. Shiffman. Social media for fisheries science and management professionals: How to use it and why you should. Fisheries, 43(3):123– 129, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10031.

Orth:2018:SMM

[2326] Donald J. Orth. Social media may empower fisheries students via learning networks. Fisheries, 43(3):130–138, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10034.

Henry:2018:IPH

[2327] Ed Henry. The International Pacific Halibut Commission and social media. Fisheries, 43(3):139–142, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10032.

Danylchuk:2018:YWM

[2328] Andy J. Danylchuk, Campbell Morgan, and Nichole Ring. So you want to make a film: an introduction to creating videos for broader impacts in fisheries and aquatic sciences. Fisheries, 43(3):144–151, March 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10059.

Doll:2018:IBM

[2329] Jason C. Doll and Stephen J. Jacquemin. Introduction to Bayesian modeling and inference for fisheries scientists. Fisheries, 43(3):152–161, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10038.

Anonymous:2018:SBS

[2330] Anonymous. Saul B. Saila (1923–2017). Fisheries, 43(3):162–163, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10048.

Anonymous:2018:CEc

[2331] Anonymous. Conferences and events. Fisheries, 43(3):166, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10060.

Anonymous:2018:NAJa

[2332] Anonymous. North American Journal of Aquaculture, volume 80, number 1, January 2018. Fisheries, 43(3):167, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10064.

Orth:2018:CL

[2333] Donald J. Orth. Claytor Lake. Fisheries, 43(3):168, March 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10057.

Anonymous:2018:IId

[2334] Anonymous. Issue information. Fisheries, 43(4):169–170, April 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10068.

Anonymous:2018:FVNd

[2335] Anonymous. Fisheries volume 43 number 4 April 2018. Fisheries, 43(4): 169–220, April 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10004.

McMullin:2018:CCW

[2336] Steve L. McMullin. Climate change, water, and fisheries. Fisheries, 43 (4):171, April 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10061.

Murchie:2018:AFB

[2337] Karen J. Murchie, Charles R. Knapp, and Peter B. McIntyre. Advancing freshwater biodiversity conservation by collaborating with public aquaria: Making the most of an engaged audience and trusted arena. Fisheries, 43(4):172–178, April 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10056.

Cockrell:2018:EAS

[2338] Marcy Cockrell, Kate Dubickas, Megan Hepner, Alex Ilich, and Matthew McCarthy. Embracing advocacy in science. Fisheries, 43(4):179–182, April 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10055.

Buktenica:2018:LTW

[2339] M. W. Buktenica, D. K. Hering, N. Scott, C. Lambert, J. McKee, E. Maletis, J. Pellissier, W. Tinniswood, and N. Banish. A longterm watershed-scale partnership to restore bull trout across Federal, State, private, and historic tribal land near Crater Lake National Park, Oregon. Fisheries, 43(4):183–193, April 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10047.

Laub:2018:ECP

[2340] Brian G. Laub, Gary P. Thiede, William W. Macfarlane, and Phaedra Budy. Evaluating the conservation potential of tributaries for native fishes in the Upper Colorado River Basin. Fisheries, 43(4):194–206, April 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10054.

Anonymous:2018:WAC

[2341] Anonymous. Welcome to Atlantic City: Annual Meeting preview. Fisheries, 43(4):208–214, April 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10088.

Anonymous:2018:CEd

[2342] Anonymous. Conferences and events. Fisheries, 43(4):217, April 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10069.

Anonymous:2018:NAJb

[2343] Anonymous. North American Journal of Fisheries Management, volume 38, number 1, February 2018. Fisheries, 43(4):218, April 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10072.

Wheaton:2018:GBS

[2344] Derek Wheaton. Gilchrist Blue Springs. Fisheries, 43(4):220, April 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10070.

Anonymous:2018:IIe

[2345] Anonymous. Issue information. Fisheries, 43(5):221–222, May 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10080.

Anonymous:2018:FVNe

[2346] Anonymous. Fisheries volume 43 number 5 May 2018. Fisheries, 43(5): 221–260, May 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10005.

Bonar:2018:HUC

[2347] Scott Bonar and Steve L. McMullin. Help us celebrate our 150th anniversary. Fisheries, 43(5):223, May 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10073.

Winters:2018:BWC

[2348] Drue Banta Winters. Budget woes continue, vigilance required on key issues. Fisheries, 43(5):224, May 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10074.

Hill:2018:CRC

[2349] Jeffrey E. Hill, Quenton M. Tuckett, and Craig A. Watson. Court ruling creates opportunity to improve management of nonnative fish and wildlife in the United States. Fisheries, 43(5):225–230, May 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10071.

Liebich:2018:RSD

[2350] Katrina B. Liebich, John F. Kocik, and William W. Taylor. Reclaiming a space for diadromous fish in the public psyche and sense of place. Fisheries, 43(5):231–240, May 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10063.

Letcher:2018:TVA

[2351] Benjamin H. Letcher, Jeffrey D. Walker, Matthew J. O’Donnell, Andrew R. Whiteley, Keith H. Nislow, and Jason A. Coombs. Three visualization approaches for communicating and exploring passive integrated transponder tag data. Fisheries, 43(5):241–248, May 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10067.

Miller:2018:CIF

[2352] Chris Miller. Commercial ice fishing for red king crab in Nome, Alaska. Fisheries, 43(5):249–254, May 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10065.

Anonymous:2018:CEe

[2353] Anonymous. Conferences and events. Fisheries, 43(5):255, May 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10076.

Anonymous:2018:JAAb

[2354] Anonymous. Journal of Aquatic Animal Health, volume 30, number 1, March 2018. Fisheries, 43(5):256, May 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10075.

Anonymous:2018:CES

[2355] Anonymous. Continuing education schedule. Fisheries, 43(5):257–259, May 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10107.

Harrison:2018:DL

[2356] Sarah Harrison. Digital lines. Fisheries, 43(5):260, May 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10081.

Anonymous:2018:IIf

[2357] Anonymous. Issue information. Fisheries, 43(6):261–262, June 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10101.

Anonymous:2018:FVNf

[2358] Anonymous. Fisheries volume 43 number 6 June 2018. Fisheries, 43(6): 261–298, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10006.

McMullin:2018:WAS

[2359] Steve L. McMullin. Why agencies should support participation in AFS. Fisheries, 43(6):263–264, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10086.

Winters:2018:DMP

[2360] Drue Banta Winters. Disaster mitigation provides opportunities to benefit fish and communities. Fisheries, 43(6):265, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10096.

Juanes:2018:SHA

[2361] Francis Juanes. On sharks and humanity, an art exhibit at the Hong Kong Maritime Museum. Fisheries, 43(6):266–267, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10078.

Runde:2018:SRU

[2362] Brendan J. Runde, Julianne E. Harris, and Jeffrey A. Buckel. Symposium review: Using electronic tags to estimate vital rates in fishes. Fisheries, 43(6):268–270, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10077.

Hamel:2018:YPS

[2363] Martin J. Hamel, Mark T. Porath, and Landon Pierce. Young Professional Survey results: Member and nonmember perspectives on decisions to join the American Fisheries Society. Fisheries, 43(6):271–277, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10066.

Reiser:2018:PPC

[2364] Dudley W. Reiser and Phillip J. Hilgert. A practitioner’s perspective on the continuing technical merits of PHABSIM. Fisheries, 43(6):278–283, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10082.

Trushenski:2018:WKH

[2365] Jesse T. Trushenski, Gary E. Whelan, and James D. Bowker. Why keep hatcheries? Weighing the economic cost and value of fish production for public use and public trust purposes. Fisheries, 43(6):284–293, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10084.

Coutant:2018:WVW

[2366] Charles C. Coutant and Henriette I. Jager. Webster Van Winkle, Jr., fish population modeler. Fisheries, 43(6):294–295, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10085.

Anonymous:2018:CEf

[2367] Anonymous. Conferences and events. Fisheries, 43(6):296, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10099.

Anonymous:2018:TAFb

[2368] Anonymous. Transactions of the American Fisheries Society, volume 147, number 2, March 2018. Fisheries, 43(6):297, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10100.

Anonymous:2018:PSH

[2369] Anonymous. Pallid sturgeon hatching. Fisheries, 43(6):298, June 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10092.

Anonymous:2018:IIg

[2370] Anonymous. Issue information. Fisheries, 43(7):299–300, July 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10123.

Anonymous:2018:FVNg

[2371] Anonymous. Fisheries volume 43 number 7 July 2018. Fisheries, 43(7): 299–338, July 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10007.

McMullin:2018:SA

[2372] Steve L. McMullin. In service to AFS. Fisheries, 43(7):301–302, July 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10125.

Gayeski:2018:FWS

[2373] Nick J. Gayeski, Jack A. Stanford, David R. Montgomery, Jim Lichatowich, Randall M. Peterman, and Richard N. Williams. The failure of wild salmon management: Need for a place-based conceptual foundation. Fisheries, 43(7):303–309, July 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10062.

Kocovsky:2018:ACS

[2374] Patrick M. Kočovský, Duane C. Chapman, and Song Qian. “Asian carp” is societally and scientifically problematic. Let’s replace it. Fisheries, 43 (7):311–316, July 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10087.

Cathcart:2018:FSI

[2375] C. Nate Cathcart. Flannelmouth sucker: The ironhorse of the Colorado River Basin. Fisheries, 43(7):318–321, July 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10104.

Skelley:2018:SWE

[2376] Emma Skelley. Students: What to expect at the 148th AFS Annual Meeting. Fisheries, 43(7):323–324, July 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10130.

McCane:2018:FNO

[2377] Jesse McCane, Chris Adam, Bill Fleming, Mark Bricker, and Matthew R. Campbell. FishGen.net: an online genetic repository for salmon and steelhead genetic baselines. Fisheries, 43(7):326–330, July 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10105.

Anonymous:2018:PS

[2378] Anonymous. Plenary speakers. Fisheries, 43(7):331–334, July 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10128.

Anonymous:2018:CEg

[2379] Anonymous. Conferences and events. Fisheries, 43(7):336, July 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10121.

Anonymous:2018:NAJc

[2380] Anonymous. North American Journal of Aquaculture, volume 80, number 2, April 2018. Fisheries, 43(7):337, July 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10122.

Rosenberger:2018:TSN

[2381] Amanda Rosenberger. Taillight shiner Notropis maculatus. Fisheries, 43(7):338, July 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10103.

Anonymous:2018:IIh

[2382] Anonymous. Issue information. Fisheries, 43(8):339–340, August 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10133.

Anonymous:2018:FVNh

[2383] Anonymous. Fisheries volume 43 number 8 August 2018. Fisheries, 43 (8):339–378, August 2018. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10008.

McMullin:2018:WWA

[2384] Steve L. McMullin. What have we accomplished this year? Fisheries, 43 (8):341–342, August 2018. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10135.

Winters:2018:OWE

[2385] Drue Banta Winters. Offshore wind energy impacts on fisheries: Investigating uncharted waters in research and monitoring. Fisheries, 43(8): 343–344, August 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10126.

Bouska:2018:SLF

[2386] Kristen L. Bouska, Amanda Rosenberger, Stephen E. McMurray, Garth A. Lindner, and Kayla N. Key. State-level freshwater mussel programs: Current status and a research framework to aid in mussel management and conservation. Fisheries, 43(8):345–360, August 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10106.

Morales:2018:ESD

[2387] Nia A. Morales and Allen Martin. Evaluating a stakeholder-driven bass regulation change in Florida: Agency staff and stakeholder perspectives. Fisheries, 43(8):361–367, August 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10129.

Amacker:2018:TMB

[2388] Todd Amacker. Tennessee — the most biodiverse inland state. Fisheries, 43(8):369–373, August 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10124.

Anonymous:2018:C

[2389] Anonymous. Calendar. Fisheries, 43(8):375, August 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10132.

Anonymous:2018:NAJd

[2390] Anonymous. North American Journal of Fisheries Management, volume 38, number 3, June 2018. Fisheries, 43(8):376–377, August 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10134.

Anonymous:2018:LMB

[2391] Anonymous. Lake Michigan bloaters. Fisheries, 43(8):378, August 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10089.

Anonymous:2018:IIi

[2392] Anonymous. Issue information. Fisheries, 43(9):379–380, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10162.

Anonymous:2018:FVNi

[2393] Anonymous. Fisheries volume 43 number 9 September 2018. Fisheries, 43(9):379–450, September 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10009.

Trushenski:2018:SFT

[2394] Jesse T. Trushenski. Seeing the forest and the trees. Fisheries, 43(9): 381, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10151.

Winters:2018:FSI

[2395] Drue Banta Winters. Forest Service impacts extend across the nation. Fisheries, 43(9):383–384, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10153.

Shively:2018:UFS

[2396] Dan Shively, John Rothlisberger, and Nat Gillespie. U.S. Forest Service forward to this special edition of Fisheries. Fisheries, 43(9):385, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10152.

Shively:2018:HPN

[2397] Dan Shively, John D. Rothlisberger, Nathaniel Gillespie, Mike Dombeck, Virgil Moore, Craig Roghair, Devin Demario, Frank McCormick, and Leslie Weldon. Historical perspectives and a New U.S. Forest Service strategy for fish and aquatic stewardship. Fisheries, 43(9):386–395, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10098.

Roper:2018:CAB

[2398] Brett B. Roper, James M. Capurso, Yvette Paroz, and Michael K. Young. Conservation of aquatic biodiversity in the context of multiple-use management on National Forest System lands. Fisheries, 43(9):396–405, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10168.

Penaluna:2018:UND

[2399] Brooke E. Penaluna, Gordon H. Reeves, Zanethia C. Barnett, Peter A. Bisson, John M. Buffington, C. Andrew Dolloff, Rebecca L. Flitcroft, Charles H. Luce, Keith H. Nislow, John D. Rothlisberger, and Melvin L. Warren, Jr. Using natural disturbance and portfolio concepts to guide aquatic–riparian ecosystem management. Fisheries, 43(9):406– 422, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10097.

Isaak:2018:CSD

[2400] Daniel J. Isaak, Michael K. Young, Callie McConnell, Brett B. Roper, Eric K. Archer, Brian Staab, Christine Hirsch, David E. Nagel, Michael K. Schwartz, and Gwynne L. Chandler. Crowd-sourced databases as essential elements for Forest Service partnerships and aquatic resource conservation. Fisheries, 43(9):423–430, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10083.

Gillespie:2018:SBR

[2401] Nathaniel Gillespie, Jonas Epstein, Susan Alexander, J. M. Bowker, Ron Medel, Mike Leonard, and Andrew Thoms. Socioeconomic benefits of recreational, commercial, and subsistence fishing associated with National Forests. Fisheries, 43(9):432–439, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10127.

Claggett:2018:V

[2402] Sally Claggett, Andy Dolloff, Rebecca Flitcroft, Bart Gamett, and Sheila Jacobson. Vignette. Fisheries, 43(9):441–445, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10120.

Anonymous:2018:CEh

[2403] Anonymous. Conferences and events. Fisheries, 43(9):448, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10160.

Anonymous:2018:JAAc

[2404] Anonymous. Journal of Aquatic Animal Health, volume 30, number 2, June 2018. Fisheries, 43(9):449, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10161.

Anonymous:2018:BTS

[2405] Anonymous. Bull trout Salvelinus confluentus, Willamette National Forest. Fisheries, 43(9):450, September 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10159.

Anonymous:2018:IIj

[2406] Anonymous. Issue information. Fisheries, 43(10):451–452, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10174.

Anonymous:2018:FVNj

[2407] Anonymous. Fisheries volume 43 number 10 October 2018. Fisheries, 43(10):451–502, October 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10010.

Trushenski:2018:TGC

[2408] Jesse T. Trushenski. Think globally, communicate locally. Fisheries, 43 (10):453, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10155.

Winters:2018:AWM

[2409] Drue Banta Winters. AFS wades into Magnuson–Stevens Act reauthorization; supports science-based fisheries management. Fisheries, 43(10): 454–456, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10172.

Wiedenhoft:2018:UH

[2410] Heather Wiedenhoft. An unlikely hero. Fisheries, 43(10):457–459, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10158.

Aday:2018:RVA

[2411] D. Derek Aday, Daniel J. Daugherty, and Debra J. Murie. Renewing the vision for AFS journals. Fisheries, 43(10):460–462, October 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10163.

Harford:2018:HCE

[2412] William J. Harford, Arnaud Grüss, Michael J. Schirripa, Skyler R. Sagarese, Meaghan Bryan, and Mandy Karnauskas. Handle with care: Establishing catch limits for fish stocks experiencing episodic natural mortality events. Fisheries, 43(10):463–471, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10131.

Pearse:2018:AAR

[2413] Devon E. Pearse and Matthew A. Campbell. Ancestry and adaptation of rainbow trout in Yosemite National Park. Fisheries, 43(10):472–484, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10136.

Al-Chokhachy:2018:PFP

[2414] Robert Al-Chokhachy, Bradley B. Shepard, Jason C. Burckhardt, Dan Garren, Scott Opitz, Todd M. Koel, Lee Nelson, and Robert E. Gresswell. A portfolio framework for prioritizing conservation efforts for Yellowstone cutthroat trout populations. Fisheries, 43(10):485–496, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10137.

Black:2018:FLA

[2415] Morgan J. Black. Fishing lessons — artisanal fisheries and the future of our oceans. Fisheries, 43(10):497, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10154.

Anonymous:2018:CEi

[2416] Anonymous. Conferences and events. Fisheries, 43(10):499, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10175.

Anonymous:2018:JHa

[2417] Anonymous. Journal highlights. Fisheries, 43(10):500–501, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10166.

Anonymous:2018:DWS

[2418] Anonymous. Deschutes wild and scenic river, Oregon. Fisheries, 43(10): 502, October 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10090.

Anonymous:2018:IIk

[2419] Anonymous. Issue information. Fisheries, 43(11):503–504, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10188.

Anonymous:2018:FVNk

[2420] Anonymous. Fisheries volume 43 number 11 November 2018. Fisheries, 43(11):503–550, November 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10011.

Trushenski:2018:ST

[2421] Jesse T. Trushenski. Seafood for thought. Fisheries, 43(11):506, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10173.

Winters:2018:WMC

[2422] Drue Banta Winters. We must be the change. Fisheries, 43(11):507, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10183.

Juanes:2018:ALF

[2423] Francis Juanes, Pierre Pepin, and John Dower. The 42nd Annual Larval Fish Conference. Fisheries, 43(11):508–509, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10177.

Tyler:2018:PAU

[2424] S. Tyler, O. P. Jensen, Z. Hogan, S. Chandra, L. M. Galland, J. Simmons, and The 2017 Taimen Research Team. Perspectives on the application of unmanned aircraft for freshwater fisheries census. Fisheries, 43(11):510– 516, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10167.

Crandall:2018:HBH

[2425] Chelsey A. Crandall, Martha Monroe, Jynessa Dutka-Gianelli, Brett Fitzgerald, and Kai Lorenzen. How to bait the hook: Identifying what motivates anglers to participate in a volunteer angler data program. Fisheries, 43(11):517–526, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10156.

Shirey:2018:ISP

[2426] Patrick D. Shirey, Leanne H. Roulson, and Thomas E. Bigford. Imperiled species policy is a critical issue for AFS. Fisheries, 43(11):527– 532, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10157.

Miller:2018:SCI

[2427] Thomas J. Miller, Cynthia M. Jones, Chad Hanson, Selina Heppell, Olaf P. Jensen, Patricia Livingston, Kai Lorenzen, Katherine Mills, Will Patterson, Patrick Sullivan, and Richard Wong. Scientific considerations informing Magnuson–Stevens Fishery Conservation and Management Act reauthorization: AFS Special Committee. Fisheries, 43(11): 533–541, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10179.

Anonymous:2018:SBM

[2428] Anonymous. Steven Berkeley Marine Conservation Fellowship: 2018 winners. Fisheries, 43(11):542–545, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10165.

Anonymous:2018:CEj

[2429] Anonymous. Conferences and events. Fisheries, 43(11):546, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10189.

Anonymous:2018:JHb

[2430] Anonymous. Journal highlights. Fisheries, 43(11):547–549, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10185.

Boczek:2018:F

[2431] Dustin Boczek. Faceoff. Fisheries, 43(11):550, November 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10182.

Anonymous:2018:IIl

[2432] Anonymous. Issue information. Fisheries, 43(12):551–552, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10210.

Anonymous:2018:FVNl

[2433] Anonymous. Fisheries volume 43 number 12 December 2018. Fisheries, 43(12):551–662, December 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10012.

Trushenski:2018:FT

[2434] Jesse T. Trushenski. Fish tales. Fisheries, 43(12):553–554, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10181.

Austen:2018:RA

[2435] Doug Austen. Reflections on AFS 1870. Fisheries, 43(12):555, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10202.

Swales:2018:FWT

[2436] Stephen Swales. Have fish will travel — a scientist’s view of the world. Fisheries, 43(12):556–562, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10170.

Klein:2018:CFU

[2437] Zachary B. Klein, Michael C. Quist, Leandro E. Miranda, Michelle M. Marron, Michael J. Steuck, and Kirk A. Hansen. Commercial fisheries of the Upper Mississippi River: a century of sustained harvest. Fisheries, 43(12):563–574, December 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10176.

Moore:2018:SWS

[2438] M. J. Moore and D. J. Orth. Stories worth sharing. Fisheries, 43(12):575– 576, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10169.

Moore:2018:STO

[2439] Michael J. Moore. Strangers in their own streams: Restoring lake sturgeon populations to the Lower Missouri River Basin. Fisheries, 43(12): 577–578, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10199.

Turcik:2018:AMA

[2440] Peter Turcik and Beth Beard. 148th Annual Meeting in Atlantic City. Fisheries, 43(12):579–582, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10197.

Anonymous:2018:SS

[2441] Anonymous. Symposia summaries. Fisheries, 43(12):583–601, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10195.

Anonymous:2018:AFSa

[2442] Anonymous. 2018 American Fisheries Society awards. Fisheries, 43 (12):603–614, December 2018. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10192.

Anonymous:2018:CAN

[2443] Anonymous. Call for award nominations: 2019 American Fisheries Society awards. Fisheries, 43(12):615–616, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10201.

Anonymous:2018:CEk

[2444] Anonymous. Conferences and events. Fisheries, 43(12):617, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10211.

Anonymous:2018:JHc

[2445] Anonymous. Journal highlights. Fisheries, 43(12):618–619, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10196.

Anonymous:2018:AFSb

[2446] Anonymous. American Fisheries Society 149^{th} Annual Meeting Reno, Nevada September 29–October 3. Fisheries, 43(12):620, December 2018. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10204.

Anonymous:2018:GS

[2447] Anonymous. Gulf sturgeon. Fisheries, 43(12):622, December 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10193.

Anonymous:2019:IIa

[2448] Anonymous. Issue information. Fisheries, 44(1):1–2, January 2019. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10222.

Anonymous:2019:FVNa

[2449] Anonymous. Fisheries volume 44 number 1 January 2019. Fisheries, 44 (1):1–48, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10234.

Trushenski:2019:RRC

[2450] Jesse T. Trushenski. Recreational and re-creational fishing. Fisheries, 44 (1):4–5, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10208.

Winters:2019:PCW

[2451] Drue Banta Winters. Proposed changes to waters of the U.S. rule could strain fisheries and wildlife. Fisheries, 44(1):6, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10207.

Roni:2019:DRR

[2452] Philip Roni. Does river restoration increase fish abundance and survival or concentrate fish? The effects of project scale, location, and fish life history. Fisheries, 44(1):7–19, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10180.

Taylor:2019:CBB

[2453] Andrew T. Taylor, James M. Long, Michael D. Tringali, and Brandon L. Barthel. Conservation of black bass diversity: an emerging management paradigm. Fisheries, 44(1):20–36, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10187.

Taylor:2019:BGB

[2454] Andrew T. Taylor and Steven M. Sammons. Bridging the gap between scientists and anglers: The Black Bass Conservation Committee’s social media outreach efforts. Fisheries, 44(1):37–41, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10186.

Hughes:2019:BBA

[2455] Robert M. Hughes. Bristol Bay Alaska: Natural resources of the aquatic and terrestrial ecosystems. Fisheries, 44(1):42, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10191.

Anonymous:2019:AMA

[2456] Anonymous. 149th Annual Meeting of the American Fisheries Society September 29–October 3, 2019: Reno–Tahoe Welcomes You! Fisheries, 44(1):43, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10206.

Anonymous:2019:CEa

[2457] Anonymous. Conferences and events. Fisheries, 44(1):44, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10223.

Anonymous:2019:JHa

[2458] Anonymous. Journal highlights. Fisheries, 44(1):45–47, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10212.

Anonymous:2019:GH

[2459] Anonymous. Great hogfish. Fisheries, 44(1):48, January 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10200.

Anonymous:2019:IIb

[2460] Anonymous. Issue information. Fisheries, 44(2):49–50, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10139.

Anonymous:2019:FVNb

[2461] Anonymous. Fisheries volume 44 number 2 February 2019. Fisheries, 44(2):49–108, February 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10237.

Trushenski:2019:DF

[2462] Jesse T. Trushenski. Decent fisheries. Fisheries, 44(2):51–52, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10213.

Cooke:2019:SFF

[2463] Steven J. Cooke, Abigail J. Lynch, Shane Vatland, Danielle A. Reich, and Julie E. Claussen. On the state of fish, fisheries, and fisheries management practices around the globe: Sharing national perspectives to build understanding. Fisheries, 44(2):53–55, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10209.

High:2019:MSA

[2464] Brett High. A motivational and successful AFS Idaho Chapter 2018 Meeting. Fisheries, 44(2):56–58, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10217.

Saul:2019:CRE

[2465] Steven Saul and Cristalís Capielo Rosario. Catch and release: The effectiveness of an educational class for anglers caught with fishing citations. Fisheries, 44(2):59–72, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10178.

Colvin:2019:HSW

[2466] Susan A. R. Colvin, S. Mažeika P. Sullivan, Patrick D. Shirey, Randall W. Colvin, Kirk O. Winemiller, Robert M. Hughes, Kurt D. Fausch, Dana M. Infante, Julian D. Olden, Kevin R. Bestgen, Robert J. Danehy, and Lisa Eby. Headwater streams and wetlands are critical for sustaining fish, fisheries, and ecosystem services. Fisheries, 44(2):73–91, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10229.

Reidy:2019:BCR

[2467] Rhonda Reidy. British Columbia: Return of the great whales, and what it may mean for fisheries. Fisheries, 44(2):92–97, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10150.

Bowker:2019:ASV

[2468] Jim Bowker. AFS Second Vice President candidate statements: Jim Bowker. Fisheries, 44(2):99–100, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10225.

Croxton:2019:ASV

[2469] April Croxton. AFS Second Vice President candidate statements: April Croxton. Fisheries, 44(2):101–102, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10221.

Anonymous:2019:CEb

[2470] Anonymous. Conferences and events. Fisheries, 44(2):103, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10233.

Anonymous:2019:JHb

[2471] Anonymous. Journal highlights. Fisheries, 44(2):106–107, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10230.

Anonymous:2019:PTH

[2472] Anonymous. PIT tagged Higgins eye pearlymussels. Fisheries, 44(2): 108, February 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10091.

Anonymous:2019:IIc

[2473] Anonymous. Issue information. Fisheries, 44(3):109–110, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10138.

Anonymous:2019:FVNc

[2474] Anonymous. Fisheries volume 44 number 3 March 2019. Fisheries, 44(3): 109–156, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10249.

Trushenski:2019:ESW

[2475] Jesse T. Trushenski. Ecosystem services — what’s in it for me? Fisheries, 44(3):111, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10226.

Winters:2019:LGS

[2476] Drue Banta Winters. The long game: Staying the course shapes policy outcomes. Fisheries, 44(3):113, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10231.

Stang:2019:LLY

[2477] Stephen A. Stang, Corbin D. Hilling, and Donald J. Orth. Lessons learned from 35 years of students organizing the mudbass classic. Fisheries, 44(3):115–117, March 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10203.

Anonymous:2019:AMC

[2478] Anonymous. AFS meetings code of conduct. Fisheries, 44(3):118, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10241.

Anonymous:2019:CCA

[2479] Anonymous. Charting the course for the American Fisheries Society: The Strategic Plan Revision Committee includes: Jim Bowker, Tom Bigford, Scott Bonar, Julie Defilippi-Simpson, Ron Essig, John Jackson, Steve Midway, and Margaret Murphy. Fisheries, 44(3):119–122, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10244.

Odenkirk:2019:FIS

[2480] John S. Odenkirk. The First International Snakehead Symposium: News from the front(s). Fisheries, 44(3):123–128, March 2019. CODEN ????

ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10218.

Kahn:2019:TAF

[2481] Desmond M. Kahn. Trends in abundance and fishing mortality of American eels. Fisheries, 44(3):129–136, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10184.

Rountree:2019:HCW

[2482] Rodney A. Rountree, Marta Bolgan, and Francis Juanes. How can we understand freshwater soundscapes without fish sound descriptions? Fisheries, 44(3):137–143, March 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10190.

Feitosa:2019:DSE

[2483] Leonardo Manir Feitosa, Ana Paula Barbosa Martins, Rosangela Paula Teixeira Lessa, Ricardo Barbieri, and Jorge Luiz Silva Nunes. Daggernose shark: an elusive species from Northern South America. Fisheries, 44(3):144–147, March 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10205.

Winemiller:2019:FPM

[2484] Reviewed by Kirk O. Winemiller. Floodplains — processes and management for ecosystem services. Fisheries, 44(3):148, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10198.

Anonymous:2019:CEc

[2485] Anonymous. Conferences and events. Fisheries, 44(3):151, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10248.

Anonymous:2019:JHc

[2486] Anonymous. Journal highlights. Fisheries, 44(3):152–155, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10243.

Anonymous:2019:CCC

[2487] Anonymous. Common carp Cyprinus carpio. Fisheries, 44(3):156, March 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10220.

Anonymous:2019:IId

[2488] Anonymous. Issue information. Fisheries, 44(4):157–158, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10255.

Anonymous:2019:FVNd

[2489] Anonymous. Fisheries volume 44 number 4 April 2019. Fisheries, 44(4): 157–204, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10257.

Miller:2019:CCM

[2490] Darren Miller. Communicating that conservation matters. Fisheries, 44 (4):159–160, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10242.

Winters:2019:WTC

[2491] Drue Banta Winters. Washington takes on climate change? Rhetoric vs. reality. Fisheries, 44(4):161, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10246.

Austen:2019:TNY

[2492] Doug Austen. The time is now: 150 years in the making. Fisheries, 44 (4):162, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10245.

Runnebaum:2019:CRR

[2493] Jocelyn M. Runnebaum, Elisabeth A. Maxwell, Joshua S. Stoll, Karen E. Pianka, and Noah G. Oppenheim. Communication, relationships, and relatability influence stakeholder perceptions of credible science. Fisheries, 44(4):164–171, April 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10214.

Herrera-Racionero:2019:SMF

[2494] Paloma Herrera-Racionero, Emmánuel Lizcano, Lluis Miret-Pastor, and Yesmina Mascarell. The Spanish Mediterranean fishing guilds (Cofradías): An example of collaborative management with a key role in sustainable fisheries. Fisheries, 44(4):172–182, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10224.

Tripp:2019:SPM

[2495] Sara J. Tripp, Quinton E. Phelps, Ryan N. Hupfeld, David P. Herzog, David E. Ostendorf, Travis L. Moore, Ron C. Brooks, and James E. Garvey. Sturgeon and paddlefish migration: Evidence to support the need for interjurisdictional management. Fisheries, 44(4):183–193, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10215.

Anonymous:2019:PIJ

[2496] Anonymous. Photography interview with Jon Anderson. Fisheries, 44(4): 194–199, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10239.

Anonymous:2019:CEd

[2497] Anonymous. Conferences and events. Fisheries, 44(4):200, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10256.

Anonymous:2019:JHd

[2498] Anonymous. Journal highlights. Fisheries, 44(4):201–203, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10250.

Tyson:2019:OMA

[2499] Julie A. Tyson. Olympic mudminnow all dressed up. Fisheries, 44(4): 204, April 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10093.

Anonymous:2019:FVNe

[2500] Anonymous. Fisheries volume 44 number 5 May 2019. Fisheries, 44(5): 205–260, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10270.

Knepp:2019:CT

[2501] Timothy Knepp. Cutthroat trout. Fisheries, 44(5):348, July 2019. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10288.

Anonymous:2019:IIe

[2502] Anonymous. Issue information. Fisheries, 44(6):261–262, June 2019. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10143.

Anonymous:2019:FVNf

[2503] Anonymous. Fisheries volume 44 number 6 June 2019. Fisheries, 44(6): 261–304, June 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10282.

Titus:2019:TRY

[2504] Rob Titus and Jackie Watson. Top 10 reasons Y’All should come to Reno–Tahoe for the Joint Annual Meeting of the American Fisheries Society and The Wildlife Society. Fisheries, 44(6):263–265, June 2019. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10276.

Quinn:2019:WTG

[2505] Danielle Quinn. Who’s talking? Gender discrepancy in conversation at AFS 148. Fisheries, 44(6):266–268, June 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10253.

Waldman:2019:WPS

[2506] John Waldman. What a postcard says about anadromous fish management. Fisheries, 44(6):269, June 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10265.

Farady:2019:FCC

[2507] Susan E. Farady and Thomas E. Bigford. Fisheries and climate change: Legal and management implications. Fisheries, 44(6):270–275, June 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10263.

Carlson:2019:TFF

[2508] Andrew K. Carlson, William W. Taylor, Michael T. Kinnison, S. Mažeika P. Sullivan, Michael J. Weber, Richard T. Melstrom, Paul A. Venturelli, Melissa R. Wuellner, Raymond M. Newman, Kyle J. Hartman, Gayle B. Zydlewski, Dennis R. DeVries, Suzanne M. Gray, Dana M. Infante, Mark A. Pegg, and Reggie M. Harrell. Threats to freshwater fisheries in the United States: Perspectives and investments of state fisheries administrators and agricultural experiment station directors. Fisheries, 44(6):276–287, June 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10238.

Guthrie:2019:LWQ

[2509] Amanda G. Guthrie, William W. Taylor, Andrew M. Muir, Henry A. Regier, and Marc Gaden. Linking water quality and fishery management facilitated the development of ecosystem-based management in the Great Lakes Basin. Fisheries, 44(6):288–292, June 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10240.

Gaeta:2019:WTE

[2510] Jereme W. Gaeta. A Web tool to estimate prey fish lengths consumed by common freshwater piscivorous fishes. Fisheries, 44(6):293– 296, June 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10235.

Anonymous:2019:AFSa

[2511] Anonymous. American Fisheries Society: Annual Meeting 2019 Continuing Education Schedule. Fisheries, 44(6):297–299, June 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10284.

Anonymous:2019:CEe

[2512] Anonymous. Conferences and events. Fisheries, 44(6):300, June 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10283.

Anonymous:2019:JHe

[2513] Anonymous. Journal highlights. Fisheries, 44(6):301–303, June 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10280.

Anonymous:2019:WSV

[2514] Anonymous. Walleye Sander vitreus. Fisheries, 44(6):304, June 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10279.

Anonymous:2019:IIf

[2515] Anonymous. Issue information. Fisheries, 44(7):205–206, May 2019. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10268.

Trushenski:2019:PP

[2516] Jesse T. Trushenski. Pricing the priceless. Fisheries, 44(7):207, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10258.

Runde:2019:SEP

[2517] Brendan J. Runde. Stakeholder engagement is the path to successful management. Fisheries, 44(7):209–211, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10251.

Grasty:2019:UQB

[2518] Sarah E. Grasty and Matthew D. Campbell. Understanding and quantifying bias in visual fisheries surveys using advanced technology: Fisheries technology series. Fisheries, 44(7):212–215, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10236.

Love:2019:WTF

[2519] Joseph W. Love. The wisdom of Thanos in fishery science: Why manage catch-and-release fisheries? Fisheries, 44(7):216–224, May 2019. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10219.

Rypel:2019:FCW

[2520] Andrew L. Rypel, Timothy D. Simonson, Daniel L. Oele, Joanna D. T. Griffin, Timothy P. Parks, David Seibel, Craig M. Roberts, Scott Toshner, Lori S. Tate, and John Lyons. Flexible classification of Wisconsin lakes for improved fisheries conservation and management. Fisheries, 44 (7):225–238, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10228.

Karnauskas:2019:ESB

[2521] Mandy Karnauskas, Rebecca J. Allee, J. Kevin Craig, Michael Jepson, Christopher R. Kelble, Morgan Kilgour, Richard D. Methot, and Seann D. Regan. Effective science-based fishery management is good for Gulf of Mexico’s “Bottom Line” — but evolving challenges remain. Fisheries, 44(7):239–242, May 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10216.

Clemens:2019:CST

[2522] Benjamin J. Clemens. A call for standard terminology for lamprey life stages. Fisheries, 44(7):243–245, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10227.

Anonymous:2019:WRA

[2523] Anonymous. Welcome to Reno: Annual Meeting preview. Fisheries, 44 (7):246–252, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10277.

Martinez:2019:MAJ

[2524] Carlos R. Martinez. In memoriam: [Arden J. Trandahl]. Fisheries, 44(7): 253–254, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10252.

Anonymous:2019:CEf

[2525] Anonymous. Conferences and events. Fisheries, 44(7):255, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10269.

Anonymous:2019:JHf

[2526] Anonymous. Journal highlights. Fisheries, 44(7):256–257, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10266.

Charter:2019:RE

[2527] Sharon Charter. Ragfish egg. Fisheries, 44(7):260, May 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10094.

Anonymous:2019:IIg

[2528] Anonymous. Issue information. Fisheries, 44(7):305–306, July 2019. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10144.

Anonymous:2019:FVNg

[2529] Anonymous. Fisheries volume 44 number 7 July 2019. Fisheries, 44(7): 305–348, July 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10317.

Trushenski:2019:SSP

[2530] Jesse T. Trushenski. Sharing science in a post-fact world. Fisheries, 44 (7):307–308, July 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10281.

Grasty:2019:PSB

[2531] Sarah Grasty, J. Christopher Taylor, and John Janssen. Pushing the sampling boundaries: Advanced technology is allowing us to survey deeper, longer, and harder to reach areas of our oceans, lakes, and rivers. Fisheries, 44(7):309–311, July 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10313.

Hardy:2019:AIC

[2532] Ryan Hardy. AFS Idaho Chapter 2019 Annual Meeting: Science-based regulations and angler ethics. Are tools missing from the toolbox? Fisheries, 44(7):312–313, July 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10264.

Webb:2019:MVP

[2533] Tess Webb and Andrew Meyer. The mystery and value of place names: Can history help us get to the bottom of sturgeon pool? Fisheries, 44(7): 314–320, July 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10259.

Sundmark:2019:EAG

[2534] Aaron P. Sundmark and Larry M. Gigliotti. Economic activity generated by angling at small South Dakota lakes. Fisheries, 44(7):321–330, July 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10261.

Lavoie:2019:EWK

[2535] Anna Lavoie, Jean Lee, Kim Sparks, Gayla Hoseth, and Sarah Wise. Engaging with women’s knowledge in Bristol Bay fisheries through oral history and participatory ethnography. Fisheries, 44(7):331–337, July 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10271.

Bourret:2019:EBI

[2536] Samuel L. Bourret. The evolution of a biologist: an interview with Clint Muhlfeld. Fisheries, 44(7):338–340, July 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10254.

Barans:2019:MAR

[2537] Charles A. Barans. Marine artificial reef research and development — integrating fisheries management objectives. Fisheries, 44(7):341, July 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10274.

Anonymous:2019:CEg

[2538] Anonymous. Conferences and events. Fisheries, 44(7):342, July 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10316.

Anonymous:2019:JHg

[2539] Anonymous. Journal highlights. Fisheries, 44(7):343–345, July 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10306.

Anonymous:2019:Ca

[2540] Anonymous. Correction. Fisheries, 44(7):345, July 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10338.

Anonymous:2019:IIh

[2541] Anonymous. Issue information. Fisheries, 44(8):349–350, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10145.

Anonymous:2019:FVNh

[2542] Anonymous. Fisheries volume 44 number 8 August 2019. Fisheries, 44 (8):349–388, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10340.

Brix:2019:ISI

[2543] Kaja Brix. Introduction to special issue on diversity and inclusion. Fisheries, 44(8):351–352, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10326.

Trushenski:2019:TDD

[2544] Jesse T. Trushenski. Thoughts on diversity, diversity in thought. Fisheries, 44(8):353–354, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10321.

Banning:2019:HJF

[2545] Mary Webb Banning. The Hutton Junior Fisheries Biology Program: an innovative summer experience to promote diversity within the fisheries profession. Fisheries, 44(8):355–356, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10319.

Parker:2019:TWD

[2546] Mamie Parker. Theme 1: Why diversity? Fisheries, 44(8):357–358, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10322.

McMullin:2019:TMD

[2547] Steve L. McMullin. Toward a more diverse and inclusive American Fisheries Society. Fisheries, 44(8):359–360, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10311.

Beamish:2019:UMD

[2548] Richard Beamish. Understanding the mechanisms of diversity. Fisheries, 44(8):361–362, August 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10327.

Hillgartner:2019:TSD

[2549] Megan Hillgartner and Sorina Stalla. Theme 2: Status of diversity in the fisheries profession: Demographics. Fisheries, 44(8):363–364, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10323.

Hillgartner:2019:TAD

[2550] Megan Hillgartner and Sorina Stalla. Theme 3: Actions on diversity in the fisheries profession. Fisheries, 44(8):365–366, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10328.

Spangler:2019:ANS

[2551] Elizabeth Spangler. The Alaska Native Science and Engineering Program. Fisheries, 44(8):367–368, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10335.

Liles:2019:PEP

[2552] George Liles. The partnership education program in Woods Hole, Massachusetts. Fisheries, 44(8):369–370, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10334.

Stalla:2019:PEP

[2553] Sorina Stalla, Megan Hillgartner, and Vladimir Alexeev. Partnership education program Alaska. Fisheries, 44(8):371–372, August 2019. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10333.

Moncada:2019:TWD

[2554] Angelica Moncada. Theme 4: What does diversity require of us? Fisheries, 44(8):373–374, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10312.

Lee:2019:TWD

[2555] Olivia Lee. Theme 4: What does diversity require of us? Fisheries, 44 (8):375, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10324.

Mecum:2019:TWD

[2556] Doug Mecum. Theme 4: What does diversity require of us? Fisheries, 44 (8):376, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10320.

Gonzalez:2019:TWD

[2557] Rosalinda Gonzalez. Theme 4: What does diversity require of us? Fisheries, 44(8):378, August 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10314.

Karp:2019:TWD

[2558] Bill Karp. Theme 4: What does diversity require of us? Fisheries, 44(8): 380–381, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10309.

Fiske:2019:WYR

[2559] Shanyn Fiske. “Where are you really from?” and other invasions of narrative space. Fisheries, 44(8):382–383, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10308.

Anonymous:2019:Cb

[2560] Anonymous. Calendar. Fisheries, 44(8):385, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10341.

Anonymous:2019:JHh

[2561] Anonymous. Journal highlights. Fisheries, 44(8):386–387, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10339.

Anonymous:2019:SH

[2562] Anonymous. Star hands. Fisheries, 44(8):388, August 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10331.

Anonymous:2019:IIi

[2563] Anonymous. Issue information. Fisheries, 44(9):389–390, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10146.

Anonymous:2019:FVNi

[2564] Anonymous. Fisheries volume 44 number 9 September 2019. Fisheries, 44(9):389–456, September 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10353.

Bonar:2019:NTA

[2565] Scott A. Bonar. Now is the time for action. Fisheries, 44(9):391, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10336.

Winters:2019:CNY

[2566] Drue Banta Winters. Conservation needs you now. Fisheries, 44(9): 393, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10305.

Lerner:2019:WWU

[2567] Aaron Lerner. Why are we using plastic to mail Fisheries magazine? Fisheries, 44(9):395, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10337.

Curtis:2019:NEY

[2568] Gary Curtis, Michael R. Donaldson, and Peter Turcik. Note from the Editors: The year of the salmon. Fisheries, 44(9):397, September 2019. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10355.

Cathcart:2019:FFP

[2569] C. Nate Cathcart. Flying the frontier: Perspectives from the Great Alaska Freshwater Fish Inventory. Fisheries, 44(9):398–402, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10287.

Hare:2019:ASR

[2570] Jonathan A. Hare, John F. Kocik, and Jason S. Link. Atlantic salmon recovery informing and informed by ecosystem-based fisheries management. Fisheries, 44(9):403–411, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10262.

Steele:2019:PBT

[2571] Craig A. Steele, Maureen Hess, Shawn Narum, and Matthew Campbell. Parentage-based tagging: Reviewing the implementation of a new tool for an old problem. Fisheries, 44(9):412–422, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10260.

Tuohy:2019:SSE

[2572] Adrian M. Tuohy, John R. Skalski, and Nick J. Gayeski. Survival of salmonids from an experimental commercial fish trap. Fisheries, 44(9): 423–432, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10292.

Sturrock:2019:EDH

[2573] Anna M. Sturrock, William H. Satterthwaite, Kristina M. Cervantes-Yoshida, Eric R. Huber, Hugh J. W. Sturrock, Sébastien Nusslé, and Stephanie M. Carlson. Eight decades of hatchery salmon releases in the California Central Valley: Factors influencing straying and resilience. Fisheries, 44(9):433–444, September 2019. CODEN ????

ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10267.

Kocovsky:2019:EGP

[2574] Patrick M. Kočovský, Patricia S. Gaunt, Brandon K. Peoples, and Emmanuel A. Frimpong. Ethical guidelines for publication of fisheries research: Publications Overview Committee, American Fisheries Society. Fisheries, 44(9):445–448, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10329.

Anonymous:2019:AFCa

[2575] Anonymous. AFS 150 first call. Fisheries, 44(9):449, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.13057.

Anonymous:2019:Cc

[2576] Anonymous. Calendar. Fisheries, 44(9):450, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10352.

Anonymous:2019:JHi

[2577] Anonymous. Journal highlights. Fisheries, 44(9):454–455, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10354.

Hagerty:2019:SS

[2578] Ryan Hagerty. Sockeye salmon. Fisheries, 44(9):456, September 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10348.

Anonymous:2019:IIj

[2579] Anonymous. Issue information. Fisheries, 44(10):457–458, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10147.

Anonymous:2019:FVNj

[2580] Anonymous. Fisheries volume 44 number 10 October 2019. Fisheries, 44(10):457–508, October 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10362.

Bonar:2019:TSA

[2581] Scott A. Bonar. Turn this ship around! Fisheries, 44(10):459–460, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10347.

Ecker:2019:DDH

[2582] Jayne K. Ecker. Dakota to Dominica: How multiple experiences can shape the future of fisheries students. Fisheries, 44(10):461–464, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10275.

Morgan:2019:RED

[2583] Cheryl A. Morgan, Brian R. Beckman, Laurie A. Weitkamp, and Kurt L. Fresh. Recent ecosystem disturbance in the Northern California Current. Fisheries, 44(10):465–474, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10273.

Murphy:2019:AAE

[2584] Robert Murphy, Jr., Steven Scyphers, Steven Gray, and Jonathan H. Grabowski. Angler attitudes explain disparate behavioral reactions to fishery regulations. Fisheries, 44(10):475–487, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10286.

Faucheux:2019:RAI

[2585] Nicky M. H. Faucheux, Cody A. Craig, and Timothy H. Bonner. Rapid assessment for identifying species of greatest conservation need: Towards a unified approach. Fisheries, 44(10):488–497, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10289.

Thorgaard:2019:BRF

[2586] Gary H. Thorgaard. Book review: Fishes of the Columbia Basin: a Guide to Their Natural History and Identification. Second Edition. Fisheries, 44(10):498, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10278.

Anonymous:2019:SBM

[2587] Anonymous. Steven Berkeley Marine Conservation Fellowship: 2019 winners. Fisheries, 44(10):499–500, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10332.

Anonymous:2019:ASC

[2588] Anonymous. AFS 150 second call. Fisheries, 44(10):501, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10361.

Anonymous:2019:Cd

[2589] Anonymous. Calendar. Fisheries, 44(10):505, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10360.

Anonymous:2019:JHj

[2590] Anonymous. Journal highlights. Fisheries, 44(10):506–507, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10358.

Pribanic:2019:FFS

[2591] J. B. Pribanic. Fishing the fall striper blitz. Fisheries, 44(10):508, October 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10356.

Anonymous:2019:IIk

[2592] Anonymous. Issue information. Fisheries, 44(11):509–510, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10148.

Anonymous:2019:FVNk

[2593] Anonymous. Fisheries volume 44 number 11 November 2019. Fisheries, 44(11):509–560, November 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10371.

Bonar:2019:MDF

[2594] Scott A. Bonar. Making a difference: a few tips. Fisheries, 44(11):511– 512, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10363.

Stafford:2019:FIF

[2595] Austin J. Stafford. The future of inland fisheries management is urban: a suburban fisherman turned natural resource intern. Fisheries, 44(11): 513–514, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10345.

Valdez:2019:PRF

[2596] Rene Xavier Valdez, Nils Peterson, Alyssa Chen, Michael Steward, Katrina Hannameyer, Hans Seebaluck, Kaj Hulthén, and R. Brian Langerhans. Perceptions of resilience in fishery-dependent Bahamian communities following a category 4 hurricane. Fisheries, 44(11):515–523, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10310.

Sanchez:2019:BRA

[2597] Phillip J. Sanchez, Jeffrey P. Pinsky, and Jay R. Rooker. Bomb radiocarbon age validation of Warsaw grouper and snowy grouper. Fisheries, 44(11):524–533, November 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10291.

Hasler:2019:RLC

[2598] Caleb T. Hasler, Graham D. Raby, Emmanuelle Chrétien, Margot Stockwell, Steven J. Cooke, Erin Rechisky, David W. Welch, Natalie M. Sopinka, and Nicholas E. Mandrak. Reflections on the legends of Canadian fisheries science and management. Fisheries, 44(11):534–538, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10290.

Glassic:2019:CFR

[2599] Hayley C. Glassic, Kurt C. Heim, and Christopher S. Guy. Creating figures in R that meet the AFS Style Guide: Standardization and supporting script. Fisheries, 44(11):539–544, November 2019. CODEN ????

ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10272.

Lin:2019:USD

[2600] Hsien-Yung Lin, Kelly F. Robinson, Michael L. Jones, and Lisa Walter. Using structured decision making to overcome scale mismatch challenges in barrier removal for watershed restoration. Fisheries, 44(11):545–550, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10342.

Allen:2019:FFW

[2601] Micheal S. Allen. Fishes in the fresh waters of Florida: an identification guide and atlas. Fisheries, 44(11):552, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10325.

Anonymous:2019:AFCb

[2602] Anonymous. AFS 150 first call. Fisheries, 44(11):554, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10370.

Anonymous:2019:Cf

[2603] Anonymous. Calendar. Fisheries, 44(11):555, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10369.

Anonymous:2019:JHk

[2604] Anonymous. Journal highlights. Fisheries, 44(11):557–558, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10367.

Anonymous:2019:SN

[2605] Anonymous. Salmon net. Fisheries, 44(11):560, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10359.

Anonymous:2019:IIl

[2606] Anonymous. Issue information. Fisheries, 44(12):561–562, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10149.

Anonymous:2019:FVNl

[2607] Anonymous. Fisheries volume 44 number 12 December 2019. Fisheries, 44(12):561–624, December 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10385.

Bonar:2019:HRC

[2608] Scott A. Bonar. Having a rough conversation? Try this! Fisheries, 44 (12):563, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10381.

Winters:2019:CHC

[2609] Drue Banta Winters. Celebrate the holidays by celebrating fisheries. Fisheries, 44(12):565–566, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10380.

Smith:2019:AAS

[2610] Bradley J. Smith and Brian G. Blackwell. Adoption of AFS standard methods: The South Dakota experience. Fisheries, 44(12):567–575, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10346.

Reid:2019:EFR

[2611] Connor H. Reid, Christopher S. Vandergoot, Jonathan D. Midwood, E. Don Stevens, Jim Bowker, and Steven J. Cooke. On the electroimmobilization of fishes for research and practice: Opportunities, challenges, and research needs. Fisheries, 44(12):576–585, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10307.

Gervasi:2019:PJP

[2612] Carissa Gervasi. Poor Jack: The plight of a forgotten fish. Fisheries, 44 (12):586–587, December 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10344.

Anonymous:2019:AFSb

[2613] Anonymous. American Fisheries Society — the Wildlife Society Joint Conference in Reno, Nevada. Fisheries, 44(12):588–592, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10376.

Anonymous:2019:SS

[2614] Anonymous. Symposia summaries. Fisheries, 44(12):593–605, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10373.

Anonymous:2019:AFSc

[2615] Anonymous. 2019 American Fisheries Society awards. Fisheries, 44 (12):606–615, December 2019. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10372.

Anonymous:2019:CAN

[2616] Anonymous. Call for award nominations: 2020 American Fisheries Society awards. Fisheries, 44(12):616–617, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10375.

Anonymous:2019:Cg

[2617] Anonymous. Calendar. Fisheries, 44(12):621, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10384.

Anonymous:2019:JHl

[2618] Anonymous. Journal highlights. Fisheries, 44(12):622–623, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10378.

Anonymous:2019:Z

[2619] Anonymous. Zebrafish. Fisheries, 44(12):624, December 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10377.

Turcik:2020:WFM

[2620] Peter Turcik. World fish migration day: an interview with Zeb Hogan. Fisheries, 45(1):46–48, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10368.

Karr:2020:SRT

[2621] Reviewed by James R. Karr. Same river twice: The politics of dam removal and river restoration. Fisheries, 45(1):50, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10343.

Anonymous:2020:Ca

[2622] Anonymous. Calendar. Fisheries, 45(1):53, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10399.

Anonymous:2020:JHa

[2623] Anonymous. Journal highlights. Fisheries, 45(1):54–55, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10398.

Anonymous:2020:FHA

[2624] Anonymous. Fisheries history: Aquaculture. Fisheries, 45(1):56, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10394.

Anonymous:2020:IIa

[2625] Anonymous. Issue information. Fisheries, 45(1):1–2, January 2020. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10293.

Anonymous:2020:FVNa

[2626] Anonymous. Fisheries volume 45 number 1 January 2020. Fisheries, 45 (1):1–56, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10400.

Bonar:2020:LBS

[2627] Scott A. Bonar. Learn the basics, save the world. Fisheries, 45(1): 3, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10383.

Austen:2020:RVA

[2628] Doug Austen. A rare vintage — AFS at 150 years. Fisheries, 45(1): 4, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10386.

Meyer:2020:RGBa

[2629] Kevin Meyer. Reducing the gap between science and public opinion at the 13th International Wild Trout Symposium. Fisheries, 45(1):5, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10382.

Murphy:2020:ARFa

[2630] Brian R. Murphy. AFS roots: The founding five. Fisheries, 45(1):6– 19, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10397.

Koenig:2020:AGG

[2631] Christopher C. Koenig, Felicia C. Coleman, and Christopher R. Malinowski. Atlantic goliath grouper of Florida: To fish or not to fish. Fisheries, 45(1):20–32, January 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10349.

Skov:2020:DFA

[2632] Christian Skov, Søren Berg, Ole Ritzau Eigaard, Troels Kjeldberg Jessen, and Peter Vilhelm Skov. Danish fisheries and aquaculture: Past, present, and future. Fisheries, 45(1):33–41, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10330.

Crandall:2020:WWT

[2633] Chelsey A. Crandall, Martha C. Monroe, and Kai Lorenzen. Why won’t they listen to us? Communicating science in contentious situations. Fisheries, 45(1):42–45, January 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10315.

Anonymous:2020:Cb

[2634] Anonymous. Correction. Fisheries, 45(1):45, January 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10407.

Anonymous:2020:IIb

[2635] Anonymous. Issue information. Fisheries, 45(2):57–58, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10294.

Anonymous:2020:FVNb

[2636] Anonymous. Fisheries volume 45 number 2 February 2020. Fisheries, 45(2):57–112, February 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10415.

Bonar:2020:MYS

[2637] Scott A. Bonar. Make your science go further with standard procedures! Fisheries, 45(2):59–60, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10402.

Shirey:2020:UAD

[2638] Patrick Douglas Shirey. Updating AFS dam removal policy to engage with policymakers. Fisheries, 45(2):61–62, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10408.

Al-Chokhachy:2020:RGC

[2639] Robert Al-Chokhachy, Lisa Heki, Tim Loux, and Roger Peka. Return of a giant: Coordinated conservation leads to the first wild reproduction of Lahontan cutthroat trout in the Truckee River in nearly a century. Fisheries, 45(2):63–73, February 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10350.

Mac:2020:SIN

[2640] Michael J. Mac and Sarah Palmer. Stakeholder involvement in natural resource decisions: The Missouri River. Fisheries, 45(2):74–83, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10364.

Jacobson:2020:IEA

[2641] Susan K. Jacobson, Jennifer Seavey, Janis Goodman, Owen C. Nichols, Lindsey C. Williams, Marcela Márquez-García, and Olga Barbosa. Integrating entrepreneurship and art to improve creative problem solving in fisheries education. Fisheries, 45(2):84–89, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10351.

Murphy:2020:ARFb

[2642] Brian R. Murphy. AFS roots: The father of all the fishes. Fisheries, 45(2):90–100, February 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10413.

Danila:2020:LB

[2643] Don Danila. Linda E. Bireley. Fisheries, 45(2):101, February 2020. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10401.

Jennings:2020:ASV

[2644] Cecil A. Jennings. AFS Second Vice President candidate statement. Fisheries, 45(2):102–103, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10403.

Simpson:2020:ASV

[2645] Julie Defilippi Simpson. AFS Second Vice President candidate statement. Fisheries, 45(2):104–105, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10409.

Meyer:2020:RGBb

[2646] Kevin Meyer. Reducing the gap between science and public opinion at the 13th International Wild Trout Symposium. Fisheries, 45(2):106, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10396.

Muhling:2020:TBW

[2647] Reviewed by Barbara Muhling. Tunas and billfishes of the world. Fisheries, 45(2):107, February 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10388.

Anonymous:2020:CEa

[2648] Anonymous. Conferences and events. Fisheries, 45(2):109, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10416.

Anonymous:2020:JHb

[2649] Anonymous. Journal highlights. Fisheries, 45(2):110–111, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10412.

Billings:2020:PSS

[2650] Brett Billings. Pallid sturgeon Scaphirhynchus albus. Fisheries, 45(2): 112, February 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10421.

Anonymous:2020:IIc

[2651] Anonymous. Issue information. Fisheries, 45(3):113–114, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10295.

Anonymous:2020:FVNc

[2652] Anonymous. Fisheries volume 45 number 3 March 2020. Fisheries, 45(3): 113–172, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10424.

Bonar:2020:EIY

[2653] Scott A. Bonar. An ecosystem, if you can keep it. Fisheries, 45(3): 115, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10420.

Aldridge:2020:CSB

[2654] Caleb A. Aldridge. Computing skills beyond R for fisheries students. Fisheries, 45(3):116–120, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10392.

Ulrich:2020:IUF

[2655] Taylor L. Ulrich and Scott A. Bonar. Inexpensive, underwater filming of rare fishes in high definition. Fisheries, 45(3):121–130, March 2020. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10391.

Clark:2020:RUP

[2656] Kalysha Clark, Danielle Pender, M. Nils Peterson, Kathryn Stevenson, Danielle Lawson, and Rachel Szczytko. Reaching underserved populations through a fisheries education program. Fisheries, 45(3):131–137, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10390.

Fraley:2020:PPS

[2657] Kevin M. Fraley, Heather N. Fraley, Donald Arthur, and Eric J. Walther. Per- and polyfluoroalkyl substances (PFAS): Anglers may be exposed to harmful chemicals in their catch. Fisheries, 45(3):138–144, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10389.

Bonar:2020:EYP

[2658] Scott A. Bonar. Exploring your personal fisheries history: The US-FCS Albatross and Clinton K. Curtis. Fisheries, 45(3):145–150, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10417.

Holmes:2020:AEQ

[2659] Keith Holmes, Kieran Cox, Amy R. Cline, Marco B. A. Hatch, Morgan J. Black, Anne K. Salomon, Dana Lepofsky, Nicole F. Smith, and Sarah Dudas. Ancient ecology: The Quadra Island clam gardens. Fisheries, 45 (3):151–156, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10374.

Kusnierz:2020:CCA

[2660] Paul C. Kusnierz, Henriette I. Jager, and Andrew S. Todd. A call for collaboration among water quality and fisheries professionals. Fisheries, 45(3):157–162, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10366.

Ham:2020:WCP

[2661] Jane Ham. World-class plenary speakers announced for World Fisheries Congress 2020. Fisheries, 45(3):163, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10422.

Stallsmith:2020:BRF

[2662] Bruce W. Stallsmith. Book review: Field Guide to Freshwater Fishes of Virginia, Bugas, P. E., Hilling, C. D., Kells, V., Pinder, M. J., Wheaton, D. A., and Orth., D. J. Johns Hopkins University Press, 2019. 195. pages. US$27.95 (Softcover). Fisheries, 45(3):164, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10405.

Anonymous:2020:CEb

[2663] Anonymous. Conferences and events. Fisheries, 45(3):165, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10423.

Anonymous:2020:JHc

[2664] Anonymous. Journal highlights. Fisheries, 45(3):166–171, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10426.

Stukel:2020:BLL

[2665] Sam Stukel. Burbot Lota lota. Fisheries, 45(3):172, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10404.

Anonymous:2020:IId

[2666] Anonymous. Issue information. Fisheries, 45(4):173–174, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10296.

Anonymous:2020:FVNd

[2667] Anonymous. Fisheries volume 45 number 4 April 2020. Fisheries, 45(4): 173–220, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10285.

Bonar:2020:OCS

[2668] Scott A. Bonar. Ode to the Civil Servant. Fisheries, 45(4):175–176, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10433.

Lerner:2020:AJF

[2669] Aaron Lerner. AFS journals and Fisheries magazine to mail without polybags. Fisheries, 45(4):176, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10442.

Wagner:2020:TRA

[2670] Curtis Wagner and Janice Kerns. Top 10 reasons to attend AFS 150 in Columbus. Fisheries, 45(4):177, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10430.

Cooke:2020:ICB

[2671] Steven J. Cooke, Alice E. I. Abrams, Aaron J. Zolderdo, and Cory D. Suski. On improved care of black bass during live-release competitive angling events — recent innovations and associated research needs. Fisheries, 45(4):178–183, April 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10393.

Moffitt:2020:AAR

[2672] Christine Moffitt, Robert M. Hughes, Doug Austen, Tom Bigford, Beth Beard, Dan Cassidy, and Aaron Lerner. AFS adapts and remains relevant in the digital world. Fisheries, 45(4):184–193, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10425.

Anonymous:2020:AFSa

[2673] Anonymous. American Fisheroes Society 150th Annual Meeting Columbus, Ohio, August 30–September 3, 2020: Welcome to Columbus: Annual Meeting preview. Fisheries, 45(4):194–198, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10428.

Armstrong:2020:KKL

[2674] Kyle M. Armstrong and Leigh P. Gaffney. Kamchatka’s Kurile Lake: Linking bears, salmon, and humans for better ecosystem management. Fisheries, 45(4):200–206, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10395.

Mizuta:2020:FMN

[2675] Darien D. Mizuta and Gary H. Wikfors. Farmed mussels in the Northeastern U.S. Exclusive Economic Zone: an opportunity too good to ignore. Fisheries, 45(4):207–213, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10365.

Sawyer:2020:CSS

[2676] Alexandra C. Sawyer, Jonathan W. Moore, Daniel E. Schindler, and Peter A. H. Westley. Connecting salmon science in an era of global change. Fisheries, 45(4):214–215, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10432.

Bartholomew:2020:OOC

[2677] Jerri L. Bartholomew. Ocean outbreak: Confronting the rising tide of marine disease. Fisheries, 45(4):216, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10387.

Anonymous:2020:CEc

[2678] Anonymous. Conferences and events. Fisheries, 45(4):217, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10431.

Anonymous:2020:JHd

[2679] Anonymous. Journal highlights. Fisheries, 45(4):218–219, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10435.

Anonymous:2020:GRA

[2680] Anonymous. Grand Rapids Annual Meeting 1936. Fisheries, 45(4):220, April 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10434.

Anonymous:2020:IIe

[2681] Anonymous. Issue information. Fisheries, 45(5):221–222, May 2020. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10446.

Anonymous:2020:FVNe

[2682] Anonymous. Fisheries volume 45 number 5 May 2020. Fisheries, 45(5): 221–287, May 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10448.

Bonar:2020:CCC

[2683] Scott A. Bonar. Coronavirus, climate change, and a bit of hope. Fisheries, 45(5):223–224, May 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10444.

Hendee:2020:IIV

[2684] Laura Hendee. An interview with illustrator Val Kells. Fisheries, 45(5): 225–228, May 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10414.

Reynolds:2020:DEF

[2685] James B. Reynolds and Jan C. Dean. Development of electrofishing for fisheries management. Fisheries, 45(5):229–237, May 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10440.

Carlson:2020:SEO

[2686] Andrew K. Carlson, William W. Taylor, Melissa R. Cronin, Mitchell J. Eaton, Lauren E. Eckert, Mark A. Kaemingk, Andrea J. Reid, and Ashley Trudeau. A social–ecological odyssey in fisheries and wildlife management. Fisheries, 45(5):238–243, May 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10439.

Copeland:2020:AIC

[2687] Tim Copeland. AFS Idaho Chapter 2020 Annual Meeting — navigating the challenges of science communication. Fisheries, 45(5):244–246, May 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10436.

Murphy:2020:HPS

[2688] Christina A. Murphy, William Gerth, Kelci Pauk, Peter Konstantinidis, and Ivan Arismendi. Hiding in plain sight: Historical fish collections aid contemporary parasite research. Fisheries, 45(5):263–270, May 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10411.

Gallo:2020:FGM

[2689] Benjamin D. Gallo, John M. Farrell, and Brian F. Leydet. Fish gut microbiome: a primer to an emerging discipline in the fisheries sciences. Fisheries, 45(5):271–282, May 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10379.

Moyle:2020:TCW

[2690] Reviewed by Peter B. Moyle. Trout and char of the world. Fisheries, 45(5):283, May 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10406.

Anonymous:2020:CEd

[2691] Anonymous. Conferences and events. Fisheries, 45(5):284, May 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10447.

Anonymous:2020:JHe

[2692] Anonymous. Journal highlights. Fisheries, 45(5):285–286, May 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10450.

Anonymous:2020:FFD

[2693] Anonymous. Fly fishing during COVID-19. Fisheries, 45(5):287, May 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10445.

Anonymous:2020:IIf

[2694] Anonymous. Issue information. Fisheries, 45(6):289–290, June 2020. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10458.

Anonymous:2020:FVNf

[2695] Anonymous. Fisheries volume 45 number 6 June 2020. Fisheries, 45(6): 289–338, June 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10297.

Bonar:2020:AMH

[2696] Scott A. Bonar. AFS making history during a historic year. Fisheries, 45(6):291, June 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10482.

Winters:2020:SWD

[2697] Drue Banta Winters. Science wins the day in Clean Water Act groundwater battle: Continued vigilance needed to protect water quality. Fisheries, 45(6):292, June 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10479.

Blanton:2020:GAS

[2698] Caleb S. Blanton, Joshuah S. Perkin, Nick Menchaca, and Kristy A. Kollaus. A gap in the armor: Spearfishing reduces biomass of invasive suckermouth armored catfish. Fisheries, 45(6):293–302, June 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10410.

Castaneda:2020:BHF

[2699] Rowshyra A. Castañeda, Colleen M. M. Burliuk, John M. Casselman, Steven J. Cooke, Karen M. Dunmall, L. Scott Forbes, Caleb T. Hasler, Kimberly L. Howland, Jeffrey A. Hutchings, Geoff M. Klein, Vivian M. Nguyen, Michael H. H. Price, Andrea J. Reid, James D. Reist, John D. Reynolds, Alexander Van Nynatten, and Nicholas E. Mandrak. A brief history of fisheries in Canada. Fisheries, 45(6):303–318, June 2020. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10449.

Lynch:2020:PIP

[2700] Abigail J. Lynch, Devin M. Bartley, Thomas Douglas Beard, Jr., David B. Bunnell, Steven J. Cooke, Ian G. Cowx, Simon Funge-Smith, Craig P. Paukert, Mark W. Rogers, and William W. Taylor. InFish: a professional network to promote global conservation and responsible use of inland fish. Fisheries, 45(6):319–326, June 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10419.

Feiner:2020:ARA

[2701] Zachary S. Feiner, Alexander W. Latzka, Max H. Wolter, Lawrence D. Eslinger, and Gene R. Hatzenbeler. Assessing the rage against the machines: Do ice anglers’ electronics improve catch and harvest rates? Fisheries, 45(6):327–333, June 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10427.

Jennings:2020:PEA

[2702] Cecil A. Jennings. Paddlefish: Ecological, aquacultural, and regulatory challenges of managing a global resource. Fisheries, 45(6):334, June 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10418.

Anonymous:2020:CEe

[2703] Anonymous. Conferences and events. Fisheries, 45(6):335, June 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10459.

Anonymous:2020:JHf

[2704] Anonymous. Journal highlights. Fisheries, 45(6):336–337, June 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10477.

Anonymous:2020:SLP

[2705] Anonymous. Sea lamprey Petromyzon marinus. Fisheries, 45(6):338, June 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10462.

Anonymous:2020:IIg

[2706] Anonymous. Issue information. Fisheries, 45(7):341–342, July 2020. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10498.

Anonymous:2020:FVNg

[2707] Anonymous. Fisheries volume 45 number 7 July 2020. Fisheries, 45(7): 341–384, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10299.

Bonar:2020:V

[2708] Scott A. Bonar. VOTE! Fisheries, 45(7):343, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10495.

Turcik:2020:SKW

[2709] Peter Turcik. Steven Kambouris wins awards for urban planning and kayak fishing. Fisheries, 45(7):344, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10492.

Gallagher:2020:IFT

[2710] Brian K. Gallagher. Ice fishing trip promotes collaboration and networking across the USA–Canada border. Fisheries, 45(7):345–346, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10453.

Pickett:2020:LTC

[2711] Brandon D. Pickett, Elizabeth M. Wallace, Perry G. Ridge, and John S. K. Kauwe. Lingering taxonomic challenges hinder conservation and management of global bonefishes. Fisheries, 45(7):347–358, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10438.

Garcia-Santiago:2020:EID

[2712] Xela García-Santiago, Amaya Franco-Uría, Luis T. Antelo, M. Teresa Moreira, Gumersindo Feijoo, and Ricardo I. Pérez-Martín. Environmental implications of discarding fish in Northern Spanish coastal bottom otter trawl fisheries. Fisheries, 45(7):359–368, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10429.

Murphy:2020:ARM

[2713] Christina A. Murphy, Lauren Zatkos, Kelsi Antonelli, A. Brent Cárdenas, Christina Linkem, Michael Manning, Juan Andres Olivos, Brooke E. Penaluna, Amanda Pollock, and Ivan Arismendi. AFS roots: Mothers of fishes. Fisheries, 45(7):369–376, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10485.

Drymon:2020:DDT

[2714] J. Marcus Drymon, Amanda E. Jefferson, Crystal Louallen-Hightower, and Sean P. Powers. Descender devices or treat tethers: Does barotrauma mitigation increase opportunities for depredation? Fisheries, 45 (7):377–379, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10476.

Docker:2020:GLS

[2715] Margaret F. Docker. Great Lakes sea lamprey: the seventy-year war on a biological invader. Fisheries, 45(7):380, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10460.

Anonymous:2020:CEf

[2716] Anonymous. Conferences and events. Fisheries, 45(7):381, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10499.

Anonymous:2020:JHg

[2717] Anonymous. Journal highlights. Fisheries, 45(7):382–383, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10493.

Anonymous:2020:CD

[2718] Anonymous. Crappie diem. Fisheries, 45(7):384, July 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10494.

Anonymous:2020:IIh

[2719] Anonymous. Issue information. Fisheries, 45(8):385–386, August 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10503.

Anonymous:2020:FVNh

[2720] Anonymous. Fisheries volume 45 number 8 August 2020. Fisheries, 45 (8):385–447, August 2020. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10300.

Bonar:2020:AAA

[2721] Scott A. Bonar. The amazing 150th anniversary of the American Fisheries Society! Fisheries, 45(8):387, August 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10454.

Austen:2020:VPL

[2722] Doug Austen. A viral push to look to the future. Fisheries, 45(8): 388, August 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10452.

Banning:2020:HJFa

[2723] Mary Webb Banning. The Hutton Junior Fisheries Biology Program: 20 years of inspiring and introducing high school students to the field of fisheries and aquatic sciences. Fisheries, 45(8):389–391, August 2020. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10505.

Whelan:2020:TFT

[2724] Gary E. Whelan, Diana M. Day, John M. Casselman, Laura Gephart, Carolyn J. Hall, James Lichatowich, Michael Matylewich, Leandro Miranda, Leanne Roulson, Patrick D. Shirey, Norman Mercado-Silva, John Waldman, and Drue Winters. Tracking fisheries through time: The American Fisheries Society as a historical lens. Fisheries, 45(8):392– 426, August 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10457.

Shaw:2020:WMA

[2725] Daniel Shaw. William E. Meehan: an AFS legacy builder. Fisheries, 45 (8):427–434, August 2020. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10484.

Zatkos:2020:ARE

[2726] Lauren Zatkos, Christina A. Murphy, Amanda Pollock, Brooke E. Penaluna, J. Andres Olivos, Emilee Mowlds, Christine Moffitt, Michael Manning, Christina Linkem, Lisa Holst, A. Brent Cárdenas, and Ivan Arismendi. AFS roots: Emmeline Moore, all things to all fishes. Fisheries, 45(8):435–443, August 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10501.

Anonymous:2020:CEg

[2727] Anonymous. Conferences and events. Fisheries, 45(8):444, August 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10504.

Anonymous:2020:JHh

[2728] Anonymous. Journal highlights. Fisheries, 45(8):445–446, August 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10508.

Anonymous:2020:UFC

[2729] Anonymous. U.S. Fish Commission R/V Fish Hawk. Fisheries, 45(8): 447, August 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10481.

Anonymous:2020:IIi

[2730] Anonymous. Issue information. Fisheries, 45(9):449–450, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10517.

Anonymous:2020:FVNi

[2731] Anonymous. Fisheries volume 45 number 9 September 2020. Fisheries, 45(9):449–500, September 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10301.

Murphy:2020:WWJ

[2732] Brian R. Murphy. What would John do? Fisheries, 45(9):451–452, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10509.

Olden:2020:TMF

[2733] Julian D. Olden, Jean R. S. Vitule, Julien Cucherousset, and Mark J. Kennard. There’s more to fish than just food: Exploring the diverse ways that fish contribute to human society. Fisheries, 45(9):453–464, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10443.

Twardek:2020:WFM

[2734] William M. Twardek, Herman Wanningen, Pao Fernández Garrido, Kerry Brink, Josh Royte, Arjan Berkhuysen, Bart Geenen, and Steven J. Cooke. World fish migration day connects fish, rivers, and people — from a one-day event to a broader social movement. Fisheries, 45(9):465– 474, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10451.

Trushenski:2020:JFS

[2735] Jesse T. Trushenski, James D. Bowker, Gary E. Whelan, and Jeff A. Heindel. From Johnny fish-seed to hatchery-bashing to shaping the shoal of aquaculture stakeholders. Fisheries, 45(9):475–483, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10510.

Pearsons:2020:EPI

[2736] Todd N. Pearsons, Alf H. Haukenes, Paul A. Hoffarth, and Steven P. Richards. Expanding partnerships and innovations to implement reform of a large Columbia River hatchery program. Fisheries, 45(9):484–491, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10437.

Quindazzi:2020:ERS

[2737] Micah J. Quindazzi, William D. P. Duguid, Katie G. Innes, Jessica Qualley, and Francis Juanes. Engaging recreational salmon anglers in fisheries ecology. Fisheries, 45(9):492–494, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10478.

Mundy:2020:FBZ

[2738] Reviewed by Bruce C. Mundy. Frogfishes. biodiversity, zoogeography, and behavioral ecology. Fisheries, 45(9):495, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10497.

Wiley:2020:MRW

[2739] Don Wiley. In memoriam: Robert W. Wiley, December 12, 1940– July 20, 2020. Fisheries, 45(9):496, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10515.

Anonymous:2020:CEh

[2740] Anonymous. Conferences and events. Fisheries, 45(9):497, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10518.

Anonymous:2020:JHi

[2741] Anonymous. Journal highlights. Fisheries, 45(9):498–499, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10519.

Anonymous:2020:DNL

[2742] Anonymous. Drift netting for lake sturgeon fry. Fisheries, 45(9):500, September 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10496.

Murphy:2020:DVA

[2743] Brian R. Murphy. Déjà vu all over again? Fisheries, 45(10):303–304, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10528.

Gray:2020:AVU

[2744] Kathleen M. Gray, Catherine E. LePrevost, and W. Gregory Cope. Anglers’ views on using signs to communicate fish consumption advisories. Fisheries, 45(10):307–316, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10463.

Anonymous:2020:IIj

[2745] Anonymous. Issue information. Fisheries, 45(10):501–502, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10530.

Anonymous:2020:FVNj

[2746] Anonymous. Fisheries volume 45 number 10 October 2020. Fisheries, 45(10):501–568, October 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10302.

Winters:2020:ACC

[2747] Drue Banta Winters. AFS calls for a clear, predictable, regulatory framework to enable growth of offshore aquaculture. Fisheries, 45(10):505– 506, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10524.

Cooke:2020:MFE

[2748] Steven J. Cooke, Robert J. Lennox, Ben Cantrell, and Andy J. Danylchuk. Micro-fishing as an emerging form of recreational angling: Research gaps and policy considerations. Fisheries, 45(10):517–521, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10487.

Murphy:2020:ARJ

[2749] Brian R. Murphy. AFS roots: James Alexander Henshall, the apostle of the black bass. Fisheries, 45(10):522–534, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10521.

Anderson:2020:RVR

[2750] Sean C. Anderson, Elise A. Keppel, and Andrew M. Edwards. Reproducible visualization of raw fisheries data for 113 species improves transparency, assessment efficiency, and monitoring. Fisheries, 45(10):535– 543, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10441.

Triezenberg:2020:MHE

[2751] Heather A. Triezenberg, Diane Doberneck, Henry Campa III, and William W. Taylor. Mid- and high-engagement programs to develop future fisheries management professionals’ skills. Fisheries, 45(10):544– 553, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10480.

Izquierdo-Pena:2020:TFS

[2752] Vanessa Izquierdo-Peña, Salvador E. Lluch-Cota, Francisco P. Chavez, Daniel B. Lluch-Cota, Enrique Morales-Bojórquez, and Germán Ponce-Díaz. Is there a future in the sustainability certification of sardine and anchovy fisheries? Fisheries, 45(10):554–560, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10461.

Anonymous:2020:SBM

[2753] Anonymous. Steven Berkeley Marine Conservation Fellowship: 2020 winners. Fisheries, 45(10):561–562, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10511.

Anonymous:2020:AFSb

[2754] Anonymous. American Fisheries Society 151st Annual Meeting Baltimore, Maryland, August 8–12, 2021. Fisheries, 45(10):563, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10537.

Wagner:2020:MRW

[2755] Eric J. Wagner and Lisa Duskin Goede. In memoriam: [Ronald W. Goede, 3 April 1934–29 August 2020]. Fisheries, 45(10):564, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10523.

Anonymous:2020:JHj

[2756] Anonymous. Journal highlights. Fisheries, 45(10):565–566, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10533.

Anonymous:2020:CEi

[2757] Anonymous. Conferences and events. Fisheries, 45(10):567, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10531.

Hagerty:2020:ODE

[2758] Ryan Hagerty. Okaloosa darter Etheostoma okaloosae. Fisheries, 45(10): 568, October 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10522.

Anonymous:2020:IIk

[2759] Anonymous. Issue information. Fisheries, 45(11):569–570, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10541.

Anonymous:2020:FVNk

[2760] Anonymous. Fisheries volume 45 number 11 November 2020. Fisheries, 45(11):569–632, November 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10303.

Anonymous:2020:IIl

[2761] Anonymous. Issue information. Fisheries, 45(11):569–570, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Murphy:2020:ESC

[2762] Brian R. Murphy. Even scientists can be biased. Fisheries, 45(11):571– 573, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10543.

Banning:2020:HJFb

[2763] Mary Webb Banning. The Hutton Junior Fisheries Biology Program: a look into the 2020 class of Hutton Scholars. Fisheries, 45(11):574– 575, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10540.

Schall:2020:FAD

[2764] Megan K. Schall, Geoffrey D. Smith, Vicki S. Blazer, Heather L. Walsh, Yan Li, and Tyler Wagner. A fishery after the decline: The Susquehanna River smallmouth bass story. Fisheries, 45(11):576–584, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10491.

Murray:2020:TIC

[2765] Devin N. Murray, David B. Bunnell, Mark W. Rogers, Abigail J. Lynch, T. Douglas Beard, Jr., and Simon Funge-Smith. Trends in inland commercial fisheries in the United States. Fisheries, 45(11):585–596, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10483.

Walters:2020:SSL

[2766] Carl J. Walters, Murdoch K. McAllister, and Villy Christensen. Has Steller sea lion predation impacted survival of Fraser River sockeye salmon? Fisheries, 45(11):597–604, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10488.

Naman:2020:BHS

[2767] Sean M. Naman, Jordan S. Rosenfeld, Jason R. Neuswanger, Eva C. Enders, John W. Hayes, Eric O. Goodwin, Ian.G. Jowett, and Brett C. Eaton. Bioenergetic habitat suitability curves for instream flow modeling: Introducing user-friendly software and its potential applications. Fisheries, 45(11):605–613, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10489.

Thornton:2020:CPU

[2768] Blake P. Thornton, Andy M. Laughlin, Douglas J. Swartz, and Bart W. Durham. A contemporary perspective on undergraduate research mentoring in university fisheries and aquatic sciences programs. Fisheries, 45(11):615–624, November 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10490.

Bettoli:2020:MMJ

[2769] Phil Bettoli, Micheal Allen, Jeff Slipke, and Steve Sammons. In memoriam: Michael J. Maceina. Fisheries, 45(11):626, November 2020. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10525.

Anonymous:2020:AFSd

[2770] Anonymous. American Fisheries Society 151st Annual Meeting Baltimore, Maryland, August 8–12, 2021. Fisheries, 45(11):627, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10550.

Anonymous:2020:CEk

[2771] Anonymous. Conferences and events. Fisheries, 45(11):628, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10542.

Anonymous:2020:JHl

[2772] Anonymous. Journal highlights. Fisheries, 45(11):629–630, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10545.

Boucher:2020:MSO

[2773] Will Boucher. Migrating steelhead Oncorhynchus mykiss. Fisheries, 45 (11):632, November 2020. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10539.

Anonymous:2020:IIm

[2774] Anonymous. Issue information. Fisheries, 45(12):633–634, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10556.

Anonymous:2020:FVNl

[2775] Anonymous. Fisheries volume 45 number 12 December 2020. Fisheries, 45(12):633–696, December 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10304.

Murphy:2020:WWA

[2776] Brian R. Murphy. Why we advocate. Fisheries, 45(12):635–636, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10555.

Austen:2020:PAC

[2777] Douglas Austen. The postponed AFS 150th Celebration — the best is yet to come. Fisheries, 45(12):637, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10554.

Orth:2020:HSM

[2778] Donald J. Orth, Joseph D. Schmitt, and Corbin D. Hilling. Hyperbole, simile, metaphor, and invasivore: Messaging about non-native blue catfish expansion. Fisheries, 45(12):638–646, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10502.

Embke:2020:FFQ

[2779] Holly S. Embke, T. Douglas Beard, Jr., Abigail J. Lynch, and M. Jake Vander Zanden. Fishing for food: Quantifying recreational fisheries harvest in Wisconsin lakes. Fisheries, 45(12):647–655, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10486.

Penaluna:2020:MRJ

[2780] Brooke E. Penaluna, Dan Shively, Brett B. Roper, Lee K. Cerveny, Shelly Witt, and John D. Rothlisberger. Mentoring relates to job satisfaction for fish biologists: a longitudinal study of the USDA Forest Service. Fisheries, 45(12):656–663, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10500.

Anonymous:2020:VAM

[2781] Anonymous. 2020 Virtual Annual meeting wrap-up. Fisheries, 45(12): 664–670, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10559.

Anonymous:2020:SS

[2782] Anonymous. Symposia summaries. Fisheries, 45(12):672–678, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10547.

Anonymous:2020:AFSc

[2783] Anonymous. 2020 American Fisheries Society awards. Fisheries, 45 (12):679–686, December 2020. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10538.

Anonymous:2020:CAN

[2784] Anonymous. Call for award nominations. Fisheries, 45(12):687–688, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10558.

Beamish:2020:MDJ

[2785] Dick Beamish. In memoriam: [Donald James Noakes, November 22, 1954–October 19, 2020]. Fisheries, 45(12):689, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10551.

Shepard:2020:MRF

[2786] Bradley B. Shepard, Steve Leary, Fred Allendorf, Matt Boyer, Ryan Kovach, David Schmetterling, Lee Nelson, Steve Amish, Travis Horton, and Don Skaar. In memoriam: [Robb F. Leary, April 11, 1955– July 27, 2020]. Fisheries, 45(12):691, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10532.

Anonymous:2020:CEj

[2787] Anonymous. Conferences and events. Fisheries, 45(12):692, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10557.

Anonymous:2020:JHk

[2788] Anonymous. Journal highlights. Fisheries, 45(12):693–695, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10561.

Turcik:2020:BCI

[2789] Peter Turcik. Blue catfish Ictalurus furcatus. Fisheries, 45(12):696, December 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10552.

Anonymous:2021:IIa

[2790] Anonymous. Issue information. Fisheries, 46(1):1–2, January 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10572.

Anonymous:2021:FVNa

[2791] Anonymous. Fisheries volume 46 number 1 January 2021. Fisheries, 46 (1):1–52, January 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10464.

Murphy:2021:TLS

[2792] Brian R. Murphy. Thinking like a scientist. Fisheries, 46(1):3–4, January 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10570.

Anonymous:2021:IRH

[2793] Anonymous. Interview with Ryan Hagerty. Fisheries, 46(1):5–7, January 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10546.

Thompson:2021:RET

[2794] Laura M. Thompson, Abigail J. Lynch, Erik A. Beever, Augustin C. Engman, Jeffrey A. Falke, Stephen T. Jackson, Trevor J. Krabbenhoft, David J. Lawrence, Douglas Limpinsel, Robert T. Magill, Tracy A. Melvin, John M. Morton, Robert A. Newman, Jay O. Peterson, Mark T. Porath, Frank J. Rahel, Suresh A. Sethi, and Jennifer L. Wilkening. Responding to ecosystem transformation: Resist, accept, or direct? Fisheries, 46(1):8–21, January 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10506.

Willmes:2021:GTI

[2795] Malte Willmes, Emily E. Jacinto, Levi S. Lewis, Rachel A. Fichman, Zachary Bess, Gabriel Singer, Anna Steel, Peter Moyle, Andrew L. Rypel, Nann Fangue, Justin J. G. Glessner, James A. Hobbs, and Eric D. Chapman. Geochemical tools identify the origins of Chinook salmon returning to a restored creek. Fisheries, 46(1):22–32, January 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10516.

Aldridge:2021:TCI

[2796] Caleb A. Aldridge, Leandro E. Miranda, and Michael E. Colvin. Themes of contemporary inland fisheries goals. Fisheries, 46(1):34–39, January 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10507.

Sinopoli:2021:SMR

[2797] Daniel A. Sinopoli and Donald J. Stewart. A synthesis of management regulations for bowfin, and conservation implications of a developing caviar fishery. Fisheries, 46(1):40–43, January 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10526.

Farrell:2021:MRG

[2798] John M. Farrell and Lee A. Fuiman. In memoriam: [Robert G. Werner 1936–2020]. Fisheries, 46(1):45, January 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10564.

Hargrove:2021:MOL

[2799] John S. Hargrove, Micheal S. Allen, Ryan J. Wasserman, Darragh J. Woodford, Bruce R. Ellender, and Tatenda Dalu. In memoriam: [Olaf Laurence Friedrich Weyl 1972–2020]. Fisheries, 46(1):47, January 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10563.

Anonymous:2021:CEb

[2800] Anonymous. Conferences and events. Fisheries, 46(1):49, January 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10573.

Anonymous:2021:JHb

[2801] Anonymous. Journal highlights. Fisheries, 46(1):50, January 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10569.

Turcik:2021:IBA

[2802] Peter Turcik. Illustration of a bowfin Amia calva. Fisheries, 46(1): 52, January 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10567.

Anonymous:2021:FVNb

[2803] Anonymous. Fisheries volume 46 number 2 February 2021. Fisheries, 46(2):53–100, February 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10465.

Anonymous:2021:IIb

[2804] Anonymous. Issue information. Fisheries, 46(2):53–54, February 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10586.

Austen:2021:LF

[2805] Doug Austen. Lessons from Fauci. Fisheries, 46(2):55, February 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10574.

Seguy:2021:PET

[2806] Lauren Seguy and James M. Long. Perceived ecological threats and economic benefits of non-native black bass in the United States. Fisheries, 46(2):56–65, February 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10520.

Safiq:2021:PEB

[2807] Alexandrea Safiq, Christopher M. Free, Joseph Caracappa, Jessica L. Valenti, and Olaf Jensen. Perceptions of ecosystem-based fisheries management among state natural resource agency scientists in the Northeastern United States. Fisheries, 46(2):66–75, February 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10527.

Grossman:2021:OER

[2808] Gary D. Grossman, Bryan B. Bozeman, Ridge W. Sliger, Troy N. Simon, and George I. Matsumoto. Open educational resource exercises for fisheries classes. Fisheries, 46(2):76–80, February 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10529.

Paradis:2021:CYF

[2809] Yves Paradis, Simon Bernatchez, Dominique Lapointe, and Steven J. Cooke. Can you fish in a pandemic? An overview of recreational fishing management policies in North America during the COVID-19 crisis. Fisheries, 46(2):81–85, February 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10544.

Anonymous:2021:E

[2810] Anonymous. Erratum. Fisheries, 46(2):85, February 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10590.

Osborn:2021:FLL

[2811] Katherine Osborn. Fisherwomen: Lessons learned from launching a new fisheries podcast. Fisheries, 46(2):86–88, February 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10566.

Cooke:2021:KCP

[2812] Steven J. Cooke, Vivian M. Nguyen, Jacqueline M. Chapman, Andrea J. Reid, Sean J. Landsman, Nathan Young, Scott G. Hinch, Stephan Schott, Nicholas E. Mandrak, and Christina A. D. Semeniuk. Knowledge coproduction: a pathway to effective fisheries management, conservation, and governance. Fisheries, 46(2):89–97, February 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10512.

Anonymous:2021:CEa

[2813] Anonymous. Conferences and events. Fisheries, 46(2):98, February 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10587.

Anonymous:2021:JHa

[2814] Anonymous. Journal highlights. Fisheries, 46(2):99, February 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10585.

Sartore:2021:BTS

[2815] Joel Sartore. Bull trout Salvelinus confluentus. Fisheries, 46(2):100, February 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10571.

Anonymous:2021:FVNc

[2816] Anonymous. Fisheries volume 46 number 3 March 2021. Fisheries, 46(3): 101–152, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2021:IIc

[2817] Anonymous. Issue information. Fisheries, 46(3):101–102, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Murphy:2021:CKU

[2818] Brian R. Murphy, Drue Banta Winters, and Julie Claussen. Communication is the key to understanding and change. Fisheries, 46(3):103–104, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Lester:2021:SBS

[2819] Nigel P. Lester, Steve Sandstrom, Derrick T. de Kerckhove, Kim Armstrong, Helen Ball, Jeff Amos, Tal Dunkley, Mike Rawson, Peter Addison, Alan Dextrase, Dan Taillon, Blair Wasylenko, Preston Lennox III, Henrique Corrêa Giacomini, and Cindy Chu. Standardized broadscale management and monitoring of inland lake recreational fisheries: An overview of the Ontario experience. Fisheries, 46(3):107–118, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Harvey:2021:AFA

[2820] Bret C. Harvey and Steven F. Railsback. All fish, all the time: A good general objective for fish passage projects? Fisheries, 46(3):119–124, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Meyer:2021:WMP

[2821] Kevin A. Meyer, Luciano V. Chiaramonte, and James B. Reynolds. The 100-watt method: a protocol for backpack electrofishing in small streams. Fisheries, 46(3):125–130, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Newcomb:2021:BPC

[2822] Tammy J. Newcomb, Paul W. Simonin, Felix A. Martinez, W. Lindsay Chadderton, Jon M. Bossenbroek, Becky Cudmore, Michael H. Hoff, Reuben P. Keller, Berkley D. Ridenhour, John D. Rothlisberger, Edward S. Rutherford, Scott Van Egeren, and David M. Lodge. A best practices case study for scientific collaboration between researchers and managers. Fisheries, 46(3):131–138, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Crandall:2021:HAW

[2823] Chelsey Crandall, Jessica Baumann, Patrick Cooney, Amanda Croteau, April Croxton, Kerry Flaherty-Walia, Janice Kerns, Richard Kraus, and Nia Morales. How to be an ally to women in fisheries science. Fisheries, 46(3):140–144, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Garcia-Bermudez:2021:ASV

[2824] Miguel A. García-Bermudez. AFS Second Vice President candidate statement. Fisheries, 46(3):145–146, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Parsons:2021:ASV

[2825] Brad Parsons. AFS Second Vice President candidate statement. Fisheries, 46(3):147–148, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2021:CEc

[2826] Anonymous. Conferences and events. Fisheries, 46(3):149, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10592.

Anonymous:2021:JHc

[2827] Anonymous. Journal highlights. Fisheries, 46(3):150–151, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10593.

Anonymous:2021:P

[2828] Anonymous. The pike. Fisheries, 46(3):152, March 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10577.

Anonymous:2021:IId

[2829] Anonymous. Issue information. Fisheries, 46(4):153–154, April 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10600.

Anonymous:2021:FVNd

[2830] Anonymous. Fisheries volume 46 number 4 April 2021. Fisheries, 46(4): 153–204, April 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10467.

Stevens:2021:BWY

[2831] Bradley Stevens and Bob Murphy. Baltimore 2021 — why you should attend the AFS Annual Meeting. Fisheries, 46(4):155, April 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10581.

Tzadik:2021:UOD

[2832] Orian E. Tzadik, William S. Arnold, Mallory Brooks, Juan Jose Cruz-Motta, Graciela Garcia-Moliner, Maria del Mar Lopez, Tauna L. Rankin, Alexis M. Sabine, and Sarah P. Stephenson. Using opportunistic datasets to infer spatial management strategies of local fisheries in the U.S. Caribbean Region. Fisheries, 46(4):156–169, April 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10548.

Shephard:2021:ACM

[2833] Samuel Shephard, John Valbo-Jorgensen, Jorge Abadía, Claudio Baigún, Carolina R. C. Doria, Nidia N. Fabré, Victoria J. Isaac, Peng Bun Ngor, Mauro L. Ruffino, and Simon Funge-Smith. From Amazon catfish to Mekong money fish: Size-based assessment of data-limited commercial inland fisheries. Fisheries, 46(4):170–187, April 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10553.

Dwivedi:2021:MAG

[2834] Arvind Kumar Dwivedi. Morphometric assessment of golden mahseer populations in the Ganga River Basin, India. Fisheries, 46(4):188–200, April 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10560.

Kitchel:2021:ORS

[2835] Zoë J. Kitchel, R. M. W. J. Bandara, Jaelyn T. Bos, René D. Clark, Daniel L. Forrest, and Malin L. Pinsky. Ocean recovery: a sustainable future for global fisheries? Fisheries, 46(4):201, April 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10580.

Anonymous:2021:CEd

[2836] Anonymous. Conferences and events. Fisheries, 46(4):202, April 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10601.

Anonymous:2021:JHd

[2837] Anonymous. Journal highlights. Fisheries, 46(4):203, April 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10602.

Anonymous:2021:NTO

[2838] Anonymous. Nile tilapia Oreochromis niloticus. Fisheries, 46(4):204, April 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10595.

Anonymous:2021:IIe

[2839] Anonymous. Issue information. Fisheries, 46(5):205–206, May 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10616.

Anonymous:2021:FVNe

[2840] Anonymous. Fisheries volume 46 number 5 May 2021. Fisheries, 46(5): 205–256, May 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10468.

Murphy:2021:LNN

[2841] Brian R. Murphy. Let nature nurture. Fisheries, 46(5):207–208, May 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10596.

Murphy:2021:QQF

[2842] Robert Murphy, Jr., Curry Cunningham, Bradley P. Harris, and Caroline Brown. Qualitative and quantitative Fisher perceptions to complement natural science data for managing fisheries. Fisheries, 46(5): 209–219, May 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10568.

Campbell:2021:MTR

[2843] Ron Campbell, Phil Hilgert, Keith Binkley, and Hal Beecher. Monitoring trout response to instream flow. Fisheries, 46(5):220–228, May 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10578.

White:2021:PTC

[2844] Seth M. White, Sarah Brandy, Casey Justice, Kayla A. Morinaga, Leslie Naylor, James Ruzycki, Edwin R. Sedell, Jesse Steele, Alexandra Towne, James G. Webster, and Ian Wilson. Progress towards a comprehensive approach for habitat restoration in the Columbia Basin: Case study in the Grande Ronde River. Fisheries, 46(5):229–243, May 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10562.

Aldridge:2021:HDW

[2845] Caleb A. Aldridge. How to decide where to go to grad school. Fisheries, 46(5):245–249, May 2021. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10605.

Juanes:2021:WPS

[2846] Francis Juanes. Writing and publishing scientific book reviews as a teaching tool in fisheries. Fisheries, 46(5):251–252, May 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10579.

Anonymous:2021:C

[2847] Anonymous. Calendar. Fisheries, 46(5):253, May 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10617.

Anonymous:2021:JHe

[2848] Anonymous. Journal highlights. Fisheries, 46(5):254–255, May 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10610.

Anonymous:2021:UNF

[2849] Anonymous. Umatilla National Forest. Fisheries, 46(5):256, May 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10599.

Anonymous:2021:IIf

[2850] Anonymous. Issue information. Fisheries, 46(6):257–258, June 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10639.

Anonymous:2021:FVNf

[2851] Anonymous. Fisheries volume 46 number 6 June 2021. Fisheries, 46(6): 257–308, June 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10469.

Murphy:2021:TAD

[2852] Brian R. Murphy. Transformative action for diversity, equity, and inclusion. Fisheries, 46(6):259–260, June 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10613.

Danylchuk:2021:MLA

[2853] Andy J. Danylchuk and Steven J. Cooke. Mentoring and leadership in aquatic field sciences during unjust, unsettling, unpredictable, and unprecedented times. Fisheries, 46(6):261–265, June 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10589.

Sass:2021:EED

[2854] Greg G. Sass, Zachary S. Feiner, and Stephanie L. Shaw. Empirical evidence for depensation in freshwater fisheries. Fisheries, 46(6):266–276, June 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10584.

Ferrer:2021:MSS

[2855] Erica M. Ferrer, Octavio Aburto-Oropeza, Victoria Jiménez-Esquivel, Juan José Cota-Nieto, Ismael Mascareñas-Osorio, and Catalina López-Sagástegui. Mexican small-scale fisheries reveal new insights into lowcarbon seafood and “Climate-Friendly” fisheries management. Fisheries, 46(6):277–287, June 2021. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10597.

Third:2021:PRU

[2856] Laura C. Third, David R. Browne, and Nicolas W. R. Lapointe. Project review under Canada’s 2012 Fisheries Act: Risky business for fisheries protection. Fisheries, 46(6):288–297, June 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10594.

Bruner:2021:SRP

[2857] John Clay Bruner. Stizostedion Rafinesque, 1820 (Percidae) is the valid generic name for walleye, sauger, and Eurasian pikeperch. Fisheries, 46 (6):298–302, June 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10582.

Ham:2021:WFC

[2858] Jane Ham. World Fisheries Congress 2021 — registration now open. Fisheries, 46(6):303–304, June 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10609.

Anonymous:2021:CEe

[2859] Anonymous. Conferences and events. Fisheries, 46(6):305, June 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10640.

Anonymous:2021:JHf

[2860] Anonymous. Journal highlights. Fisheries, 46(6):306–307, June 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10620.

Anonymous:2021:PP

[2861] Anonymous. Picke-perch. Fisheries, 46(6):308, June 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10598.

Anonymous:2021:IIg

[2862] Anonymous. Issue information. Fisheries, 46(7):309–310, July 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10647.

Anonymous:2021:FVNg

[2863] Anonymous. Fisheries volume 46 number 7 July 2021. Fisheries, 46(7): 309–352, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10470.

Winters:2021:RAW

[2864] Drue Banta Winters. Recovering America’s Wildlife Act and other initiatives provide hope from Capitol Hill. Fisheries, 46(7):311, July 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10638.

Stewart:2021:NOR

[2865] Ian J. Stewart, Jonathan J. Scordino, Joseph R. Petersen, Alex W. Wise, Cole I. Svec, Reginald H. Buttram, Joshua L. Monette, Maria R. Gonzales, Russell Svec, Joe Scordino, Kaeden Butterfield, William Parker, and Laurence A. Buzzell. Out with the new and in with the old: Reviving a traditional Makah halibut hook for modern fisheries management challenges. Fisheries, 46(7):313–320, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10603.

Harris:2021:TBS

[2866] Holden E. Harris, Brian K. Whalen, Andrew G. Gude, and Micheal S. Allen. Testing a bite-shortened hook to minimize fish handling in a recreational fishery. Fisheries, 46(7):321–328, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10608.

Imhof:2021:RLP

[2867] J. G. Imhof, N. W. Lapointe, S. Metikosh, and C. T. Hasler. Regaining lost protections: Status of the revisions to the Canadian Fisheries Act. Fisheries, 46(7):329–334, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10614.

Turcik:2021:IJJ

[2868] Peter Turcik and Mary Webb Banning. An interview with Jordyn Jones and Sean Landsman, Hutton Junior Fisheries Biology Program alumni. Fisheries, 46(7):336–339, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10637.

Zydlewski:2021:CRS

[2869] Joseph Zydlewski. From catastrophe to recovery: Stories of fisheries management successes. Fisheries, 46(7):340, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10604.

Anonymous:2021:AMB

[2870] Anonymous. 151th Annual Meeting Baltimore, November 6–10, 2021: Investing in people, habitat, and science. Fisheries, 46(7):341–345, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10649.

Kirk:2021:AIC

[2871] Robert W. Van Kirk. AFS Idaho Chapter 2021 Annual Meeting — “Aquaculture: Putting the fish in fisheries”. Fisheries, 46(7):347–348, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10612.

Anonymous:2021:CEf

[2872] Anonymous. Conferences and events. Fisheries, 46(7):349, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10648.

Anonymous:2021:JHg

[2873] Anonymous. Journal highlights. Fisheries, 46(7):350–351, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10652.

Weimer:2021:OSS

[2874] Mike Weimer. Osprey with spotted sea trout. Fisheries, 46(7):352, July 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10615.

Anonymous:2021:IIh

[2875] Anonymous. Issue information. Fisheries, 46(8):353–354, August 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10661.

Anonymous:2021:FVNh

[2876] Anonymous. Fisheries volume 46 number 8 August 2021. Fisheries, 46 (8):353–404, August 2021. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10471.

Argent:2021:PEF

[2877] David Argent. The pandemic’s effect on fisheries — an academic’s perspective. Fisheries, 46(8):355–356, August 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10644.

Branigan:2021:RCP

[2878] Philip R. Branigan, Kevin A. Meyer, and John D. Cassinelli. Relative cost and post-release performance of hatchery catchable rainbow trout grown to two target sizes. Fisheries, 46(8):357–371, August 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10621.

Arismendi:2021:UAF

[2879] Ivan Arismendi, Katherine R. McLaughlin, and Brooke E. Penaluna. The U.S. Academic Fisheries Co-Authorship Network under the lens of diversity and inclusion. Fisheries, 46(8):372–382, August 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10618.

Sharma:2021:PSR

[2880] Shailesh Sharma and John Waldman. Potential solar replacement of hydroelectricity to reopen rivers: Maine as a case example. Fisheries, 46 (8):383–390, August 2021. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10619.

McMillan:2021:WFC

[2881] Biela McMillan. The World Fisheries Congress 2021. Fisheries, 46(8): 392–393, August 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10650.

Anonymous:2021:AAMa

[2882] Anonymous. The 151st AFS Annual Meeting November 6–10, 2021, Baltimore Maryland: Events and attractions. Fisheries, 46(8):395–399, August 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10665.

Wallace:2021:SSR

[2883] S. Scott Wallace. Sharks, skates, rays and chimeras of British Columbia. Fisheries, 46(8):400, August 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10611.

Anonymous:2021:CEg

[2884] Anonymous. Conferences and events. Fisheries, 46(8):401, August 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10662.

Anonymous:2021:JHh

[2885] Anonymous. Journal highlights. Fisheries, 46(8):402–403, August 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10655.

Heilmann:2021:EDB

[2886] Tami Heilmann. Embrey Dam breach. Fisheries, 46(8):404, August 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10643.

Anonymous:2021:IIi

[2887] Anonymous. Issue information. Fisheries, 46(9):405–406, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10672.

Anonymous:2021:FVNi

[2888] Anonymous. Fisheries volume 46 number 9 September 2021. Fisheries, 46(9):405–476, September 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10472.

Murphy:2021:YVO

[2889] Brian R. Murphy. Yes Virginia, our climate is changing (and there is but one solution). Fisheries, 46(9):407–410, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10663.

Bonar:2021:DSW

[2890] Scott A. Bonar. Development of the “Statement of world aquatic scientific societies on the need to take urgent action against human-caused climate change, based on scientific evidence”. Fisheries, 46(9):411–412, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10641.

Anonymous:2021:SWA

[2891] Anonymous. Statement of world aquatic scientific societies on the need to take urgent action against human-caused climate change, based on scientific evidence. Fisheries, 46(9):413–422, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10642.

Staudinger:2021:HDC

[2892] Michelle D. Staudinger, Abigail J. Lynch, Sarah K. Gaichas, Michael G. Fox, Daniel Gibson-Reinemer, Joseph A. Langan, Amy K. Teffer, Stephen J. Thackeray, and Ian J. Winfield. How does climate change affect emergent properties of aquatic ecosystems? Fisheries, 46(9):423– 441, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10606.

Limburg:2021:CCC

[2893] Karin E. Limburg and Françoise Daverat. COVID, climate change, and professional travel: Reflections by fisheries scientists during the pandemic. Fisheries, 46(9):442–444, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10607.

Neal:2021:BPC

[2894] J. Wesley Neal, Julie E. Claussen, Marlis R. Douglas, Erin T. Spencer, Erin Tracy, Heidi Blasius, Theresa Mackey M. S., Carolyn J. Hall, Paul C. Kusnierz, Michael E. Douglas, and Scott Bonar. Best practices for communicating climate science for fisheries professionals. Fisheries, 46(9):445–448, September 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10635.

Paukert:2021:CCE

[2895] Craig Paukert, Julian D. Olden, Abigail J. Lynch, David D. Breshears, R. Christopher Chambers, Cindy Chu, Margaret Daly, Kimberly L. Dibble, Jeff Falke, Dan Issak, Peter Jacobson, Olaf P. Jensen, and Daphne Munroe. Climate change effects on North American fish and fisheries to inform adaptation strategies. Fisheries, 46(9):449–464, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10668.

Roulson:2021:WMV

[2896] Leanne H. Roulson. WE must be the voice for science. Fisheries, 46 (9):465–466, September 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10656.

Winters:2021:SNB

[2897] Drue Banta Winters. Shifting the narrative for a better outcome in the fight against climate change. Fisheries, 46(9):467–468, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10657.

Mawdsley:2021:CCS

[2898] Jonathan Mawdsley. Climate, continental-scale conservation, and coop units. Fisheries, 46(9):469–470, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10659.

Winters:2021:ACA

[2899] Drue Banta Winters, Julie Claussen, Carolyn Hall, and Katie O’Reilly. The AFS Climate Ambassador Program will help bridge the communication gap in climate change science. Fisheries, 46(9):471–472, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10658.

Anonymous:2021:CEh

[2900] Anonymous. Conferences and events. Fisheries, 46(9):473, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10673.

Anonymous:2021:JHi

[2901] Anonymous. Journal highlights. Fisheries, 46(9):474–475, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10667.

Anonymous:2021:ERS

[2902] Anonymous. Endangered razorback sucker. Fisheries, 46(9):476, September 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10664.

Anonymous:2021:IIj

[2903] Anonymous. Issue information. Fisheries, 46(10):477–478, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10680.

Anonymous:2021:FVNj

[2904] Anonymous. Fisheries volume 46 number 10 October 2021. Fisheries, 46(10):477–532, October 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10473.

Roulson:2021:TSW

[2905] Leanne H. Roulson. Taking stock as we hope to return to in-person meetings. Fisheries, 46(10):479–480, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10670.

Grasty:2021:UDE

[2906] Sarah Grasty, Michael Messina, and Janine Bryan. Upstream and downstream — exciting advances to modernize fish passage and improve data collection. Fisheries, 46(10):481–484, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10513.

Mather:2021:MSS

[2907] Martha E. Mather, Joseph M. Smith, Kayla M. Boles, Ryland B. Taylor, Cristina G. Kennedy, Sean M. Hitchman, Jane S. Rogosch, and Holly J. Frank. Merging scientific silos: Integrating specialized approaches for thinking about and using spatial data that can provide new directions for persistent fisheries problems. Fisheries, 46(10):485– 494, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10645.

Hitchman:2021:DTQ

[2908] Sean M. Hitchman, Martha E. Mather, and Joseph M. Smith. Does type, quantity, and location of habitat matter for fish diversity in a Great Plains riverscape? Fisheries, 46(10):495–504, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10634.

Tracy:2021:ICF

[2909] Erin E. Tracy, Chad N. Teal, Steven J. Ingram, Christopher J. Jenney, Joshua D. Grant, and Scott A. Bonar. The impact of COVID-19 on freshwater fisheries fieldwork and data collection. Fisheries, 46(10):505– 511, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10636.

Cathcart:2021:OAI

[2910] Charles Cathcart. Olfactory activation: Imprinting as an emerging frontier in the conservation of non-salmonid migratory fishes. Fisheries, 46 (10):513–518, October 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10646.

Reynolds:2021:MLL

[2911] James B. Reynolds. In memoriam: [A. Lawrence “Larry” Kolz February 2, 1936–August 1, 2021]. Fisheries, 46(10):520, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10666.

Anonymous:2021:SBM

[2912] Anonymous. Steven Berkeley Marine Conservation Fellowship: 2021 winners. Fisheries, 46(10):521–522, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10653.

Anonymous:2021:AAMb

[2913] Anonymous. The 151st AFS Annual Meeting, November 6–10, 2021, Baltimore, Maryland: Plenary speakers and exhibits. Fisheries, 46(10): 524–528, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10685.

Anonymous:2021:CEi

[2914] Anonymous. Conferences and events. Fisheries, 46(10):529, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10681.

Anonymous:2021:JHj

[2915] Anonymous. Journal highlights. Fisheries, 46(10):530–531, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10676.

Sexton:2021:ECP

[2916] Tyler Sexton. Endangered Colorado pikeminnow. Fisheries, 46(10):532, October 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10675.

Anonymous:2021:IIk

[2917] Anonymous. Issue information. Fisheries, 46(11):533–534, November 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10689.

Anonymous:2021:FVNk

[2918] Anonymous. Fisheries volume 46 number 11 November 2021. Fisheries, 46(11):533–588, November 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10474.

Roulson:2021:FSP

[2919] Leanne H. Roulson. Fisheries science and policy go hand in hand. Fisheries, 46(11):535–536, November 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10682.

Miles:2021:SSC

[2920] Cassidy Miles. 3 simple steps to create a more inclusive work space for our LGBTQ+ community. Fisheries, 46(11):537–538, November 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10677.

Beamish:2021:NRR

[2921] Richard Beamish and Chrys Neville. The natural regulation and relevance of wild and hatchery Coho salmon production in the Strait of Georgia. Fisheries, 46(11):539–551, November 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10651.

Bolser:2021:OAA

[2922] Derek G. Bolser, Jack. P. Egerton, Arnaud Grüss, and Brad E. Erisman. Optic–acoustic analysis of fish assemblages at petroleum platforms. Fisheries, 46(11):552–563, November 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10654.

Lennox:2021:LFE

[2923] Robert J. Lennox, Cecilie I. Nilsen, Ainslie Nash, Erlend M. Hanssen, Hilde L. Johannesen, Saron Berhe, Bjørn Barlaup, and Knut Wiik Vollset. Laboratory and field experimental validation of two different predation sensors for instrumenting acoustic transmitters in fisheries research. Fisheries, 46(11):565–573, November 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10669.

Lynch:2021:UIC

[2924] Abigail J. Lynch, Nicholas A. Sievert, Holly S. Embke, Ashley M. Robertson, Bonnie J. E. Myers, Micheal S. Allen, Zachary S. Feiner, Frederick Hoogakker, Scott Knoche, Rebecca M. Krogman, Stephen R. Midway, Chelsey L. Nieman, Craig P. Paukert, Kevin L. Pope, Mark W. Rogers, Lyndsie S. Wszola, and T. Douglas Beard, Jr. The U.S. Inland Creel and Angler Survey Catalog (CreelCat): Development, applications, and opportunities. Fisheries, 46(11):574–583, November 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10671.

Anonymous:2021:CEj

[2925] Anonymous. Conferences and events. Fisheries, 46(11):585, November 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10690.

Anonymous:2021:JHk

[2926] Anonymous. Journal highlights. Fisheries, 46(11):586–587, November 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10683.

Shaull:2021:CP

[2927] Lorie Shaull. Up close and personal. Fisheries, 46(11):588, November 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10688.

Anonymous:2021:IIl

[2928] Anonymous. Issue information. Fisheries, 46(12):589–590, December 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10704.

Anonymous:2021:FVNl

[2929] Anonymous. Fisheries volume 46 number 12 December 2021. Fisheries, 46(12):589–640, December 2021. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10475.

Roulson:2021:MRH

[2930] Leanne H. Roulson. Making ripples: How individual action can contribute to large scale effects. Fisheries, 46(12):591–592, December 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10694.

Bunch:2021:WDI

[2931] Aaron Bunch. What does inclusion look like? Fisheries, 46(12):593– 594, December 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10693.

Scarnecchia:2021:SFR

[2932] Dennis L. Scarnecchia, Jason D. Schooley, Alec R. Lackmann, Steven J. Rider, Dennis K. Riecke, Joseph McMullen, J. Eric Ganus, Kirk D. Steffensen, Nicholas W. Kramer, and Zachary R. Shattuck. The sport fish restoration program as a funding source to manage and monitor bowfishing and monitor inland commercial fisheries. Fisheries, 46(12):595– 604, December 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10679.

Rypel:2021:GRF

[2933] Andrew L. Rypel, Parsa Saffarinia, Caryn C. Vaughn, Larry Nesper, Katherine O’Reilly, Christine A. Parisek, Matthew L. Miller, Peter B. Moyle, Nann A. Fangue, Miranda Bell-Tilcock, David Ayers, and Solomon R. David. Goodbye to “Rough Fish”: Paradigm shift in the conservation of native fishes. Fisheries, 46(12):605–616, December 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10660.

Simpson:2021:AIS

[2934] Julie DeFilippi Simpson, Sarah M. Glaser, Karen J. Murchie, and Patrick D. Shirey. Angling for ideas in science communication to hosting a photo-essay workshop for teens: Implementing the winning idea from Wiley Publishing’s first impact-a-thon. Fisheries, 46(12):617–628, December 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10674.

Muhlfeld:2021:MBB

[2935] Clint Muhlfeld, Lee Nelson, Pat Clancey, Chris Guy, Joe DosSantos, Pat Graham, Matt Boyer, Chris Downs, Robert Al-Chokhachy, Brian Marotz, Paul Hamlin, Al Zale, Wade Fredenberg, and Mary Lennon. In memoriam: [Bradley B. Shepard, June 12, 1952–September 23, 2021]. Fisheries, 46(12):629, December 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10691.

Rose:2021:MJJ

[2936] Kenneth Rose, William F. Patterson III, Stephen R. Midway, and David L. Nieland. In memoriam: [James “Jim” H. Cowan, Jr. March 9, 1954–August 11, 2021]. Fisheries, 46(12):630, December 2021. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10692.

Banning:2021:ASH

[2937] Mary Webb Banning. Another summer of Hutton Scholars, with a twist! Fisheries, 46(12):631–632, December 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10698.

Anonymous:2021:CEk

[2938] Anonymous. Conferences and events. Fisheries, 46(12):637, December 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10702.

Anonymous:2021:JHl

[2939] Anonymous. Journal highlights. Fisheries, 46(12):638–639, December 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10699.

Schooley:2021:SB

[2940] Jason D. Schooley. Smallmouth buffalo. Fisheries, 46(12):640, December 2021. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10700.

Anonymous:2022:IIa

[2941] Anonymous. Issue information. Fisheries, 47(1):1–2, January 2022. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10718.

Anonymous:2022:FVNa

[2942] Anonymous. Fisheries volume 47 number 1 January 2022. Fisheries, 47 (1):1–48, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10622.

Roulson:2022:LC

[2943] Leanne H. Roulson. Let’s connect. Fisheries, 47(1):3–4, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10716.

Guo:2022:AVC

[2944] Lian Guo. Amplified voices: a collection of discussions from women of fisheries. Fisheries, 47(1):5, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10715.

Anonymous:2022:WDN

[2945] Anonymous. What does the new logo say about AFS and who we are? Fisheries, 47(1):7, January 2022. CODEN ???? ISSN 0363-2415 (print),

1548-8446 (electronic). URL https://afspubs.onlinelibrary.wiley.
com/doi/10.1002/fsh.10717.

Murawski:2022:FF

[2946] Steven Murawski, Larry McKinney, and William Hogarth. Fisheries forum. Fisheries, 47(1):8, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10712.

Kopaska:2022:WSS

[2947] Jeff Kopaska and Joseph D. Conroy. Who shall speak for us? Fisheries, 47(1):9, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10709.

Murphy:2022:DMB

[2948] Robert Murphy, Jr., Steven Scyphers, Steven Gray, and Jonathan H. Grabowski. Diversity in motivations and behavioral response to regulations in the striped bass commercial fishery. Fisheries, 47(1):10–17, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10678.

Anonymous:2022:AAMa

[2949] Anonymous. AFS Annual Meeting. Fisheries, 47(1):18–25, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10720.

Stokes:2022:IFG

[2950] Gretchen L. Stokes. Inland fisheries go to the “Fishician” for a head to tail exam. Fisheries, 47(1):26–27, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10710.

Anonymous:2022:ASS

[2951] Anonymous. AFS 151 Symposia summaries. Fisheries, 47(1):28–34, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10708.

Anonymous:2022:AFSa

[2952] Anonymous. 2021 American Fisheries Society Awards. Fisheries, 47(1): 35–42, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10711.

Anonymous:2022:Ca

[2953] Anonymous. Calendar. Fisheries, 47(1):44, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://afspubs.

onlinelibrary.wiley.com/doi/10.1002/fsh.10719.

Anonymous:2022:JHa

[2954] Anonymous. Journal highlights. Fisheries, 47(1):45–46, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10714.

Neal:2022:GDP

[2955] J. Wesley Neal. Gilt darter Percina evides print in acrylic. Fisheries, 47 (1):48, January 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10713.

Anonymous:2022:IIb

[2956] Anonymous. Issue information. Fisheries, 47(2):49–50, February 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10732.

Anonymous:2022:FVNb

[2957] Anonymous. Fisheries volume 47 number 2 February 2022. Fisheries, 47 (2):49–96, February 2022. CODEN ???? ISSN 0363-2415 (print), 1548-

8446 (electronic). URL https://afspubs.onlinelibrary.wiley.com/
doi/10.1002/fsh.10623.

Roulson:2022:WCE

[2958] Leanne H. Roulson. What can the effects of COVID reveal about lingering obstacles to retention for women in STEM? Fisheries, 47(2): 51–52, February 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10728.

Crandall:2022:AVCa

[2959] Chelsey Crandall. Amplified voices: a collection of discussions from women of fisheries. Fisheries, 47(2):53–54, February 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10724.

Sass:2022:CYW

[2960] Greg G. Sass, Stephanie L. Shaw, and Kathryn M. Renik. Celebrating 75 years of Wisconsin’s Northern Highland Fishery Research Area: The past, present, and future. Fisheries, 47(2):55–67, February 2022. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10687.

Carlson:2022:SUP

[2961] Andrew K. Carlson, William W. Taylor, Dennis R. DeVries, C. Paola Ferreri, Michael J. Fogarty, Kyle J. Hartman, Dana M. Infante, Michael T. Kinnison, Simon A. Levin, Richard T. Melstrom, Raymond M. Newman, Malin L. Pinsky, Daniel I. Rubenstein, S. Mažeika P. Sullivan, Paul A. Venturelli, Michael J. Weber, Melissa R. Wuellner, and Gayle B. Zydlewski. Stepping up: a U.S. perspective on the ten steps to responsible inland fisheries. Fisheries, 47(2):68–77, February 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10695.

Reis-Filho:2022:PUU

[2962] José Amorim Reis-Filho and Tommaso Giarrizzo. Perspectives on the use of unmanned aerial vehicle systems as tools for small-scale fisheries research and management. Fisheries, 47(2):78–89, February 2022. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10696.

Matthews:2022:MWS

[2963] William J. Matthews, Edie Marsh-Matthews, and Ana Hiott. In memoriam: [William L. Shelton, May 28, 1939–October 14, 2021]. Fisheries, 47 (2):90, February 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446

(electronic). URL https://afspubs.onlinelibrary.wiley.com/doi/
10.1002/fsh.10707.

Anonymous:2022:Cb

[2964] Anonymous. Calendar. Fisheries, 47(2):92, February 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https://

afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10733.

Anonymous:2022:JHb

[2965] Anonymous. Journal highlights. Fisheries, 47(2):93–94, February 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (elec-

tronic). URL https://afspubs.onlinelibrary.wiley.com/doi/10.
1002/fsh.10729.

Anonymous:2022:HTE

[2966] Anonymous. Hatched trout eggs. Fisheries, 47(2):96, February 2022. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:

//afspubs.onlinelibrary.wiley.com/doi/10.1002/fsh.10095.

Anonymous:2022:FVNc

[2967] Anonymous. Fisheries volume 47 number 3 March 2022. Fisheries, 47(3): 97–144, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIc

[2968] Anonymous. Issue information. Fisheries, 47(3):97–98, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Roulson:2022:TMA

[2969] Leanne H. Roulson. Tell me about something that was unexpected. Fisheries, 47(3):99–100, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Favrot:2022:URI

[2970] Scott Favrot. Unconscious regional incivility: a threat to diversity. Fisheries, 47(3):101–103, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Mallet:2022:RII

[2971] Jerry Mallet and Russell F. Thurow. Resurrecting an Idaho icon: How research and management reversed declines of native westslope cutthroat trout. Fisheries, 47(3):104–117, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Pennock:2022:NFN

[2972] Casey A. Pennock, Phaedra Budy, William W. Macfarlane, Matthew J. Breen, Justin Jimenez, and John C. Schmidt. Native fish need a natural flow regime. Fisheries, 47(3):118–123, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Turcik:2022:WNG

[2973] Peter Turcik. Wes Neal: Gyotaku printing to fish painting. Fisheries, 47 (3):124–127, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Leonard:2022:SIP

[2974] Mike Leonard. Sportfishing industry perspective on funding bowfishing management. Fisheries, 47(3):128, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Conroy:2022:FPN

[2975] Joseph D. Conroy, Elsa Goerig, John Jackson, Jeff Kopaska, Tom Lang, Jessica Mistak, Brian Murphy, Todd Pearsons, and Macey Rowan. Focused process: Nominating the next AFS Second Vice President Candidates. Fisheries, 47(3):129, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Dockendorf:2022:ASV

[2976] Kevin J. Dockendorf. AFS Second Vice President Candidate statement. Fisheries, 47(3):130–131, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Whelan:2022:ASV

[2977] Gary Whelan. AFS Second Vice President Candidate statement. Fisheries, 47(3):132–134, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Avila:2022:FED

[2978] Alexandra M. Avila, Jill E. Munger, Claire Rosemond, Samara M. Haver, Olivia Boisen, Janelle Monet Layton, Mee ya H. Monneedy, Katherine S. Lasdin, and Scott Heppell. Fishery ecosystem dynamics. Fisheries, 47 (3):135, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Dillard:2022:MWP

[2979] Joe G. Dillard. In memoriam: [William L. Pflieger (October 26, 1932– December 15, 2021)]. Fisheries, 47(3):136, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JASa

[2980] Anonymous. Joint Aquatic Societies Meeting 2022. Fisheries, 47(3): 137–140, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:Cc

[2981] Anonymous. Calendar. Fisheries, 47(3):141, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JHc

[2982] Anonymous. Journal highlights. Fisheries, 47(3):142–143, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:RYC

[2983] Anonymous. Restoring Yellowstone cutthroat trout. Fisheries, 47(3): 144, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:FVNd

[2984] Anonymous. Fisheries volume 47 number 4 April 2022. Fisheries, 47(4): 145–192, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IId

[2985] Anonymous. Issue information. Fisheries, 47(4):145–146, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Haukenes:2022:AJU

[2986] Alf Haukenes and Dan Brauch. AFS 2022, join us in Spokane. Fisheries, 47(4):147, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Crandall:2022:AVCb

[2987] Chelsey Crandall. Amplified voices: a collection of discussions from women of fisheries. Fisheries, 47(4):148–149, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:SS

[2988] Anonymous. Symposia summaries. Fisheries, 47(4):150, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Scharpf:2022:SOP

[2989] Christopher Scharpf and Ronald Fricke. Sander Oken 1817 (Percidae) is the valid generic name for walleye, sauger, and European pikeperches: a response to Bruner (2021). Fisheries, 47(4):151–153, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Harris:2022:SVS

[2990] Holden E. Harris. Science vs. sensationalism: Lessons for science communication in fisheries from Netflix’s Seaspiracy. Fisheries, 47(4):154–156, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

vonBiela:2022:PMO

[2991] Vanessa R. von Biela, Christopher J. Sergeant, Michael P. Carey, Zachary Liller, Charles Russell, Stephanie Quinn-Davidson, Peter S. Rand, Peter A. H. Westley, and Christian E. Zimmerman. Premature mortality observations among Alaska’s Pacific salmon during record heat and drought in 2019. Fisheries, 47(4):157–168, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Doyle:2022:IEM

[2992] Eric G. Doyle, John E. Arterburn, and Ryan S. Klett. Integrating ecosystem models with long-term monitoring to support salmon recovery. Fisheries, 47(4):169–179, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:AMP

[2993] Anonymous. Annual meeting preview. Fisheries, 47(4):181–184, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Sequeira:2022:OSE

[2994] Ana M. M. Sequeira. The ocean sunfishes: Evolution, biology, and conservation. Fisheries, 47(4):185, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:M

[2995] Anonymous. In memoriam: [Michael Wolff Street (February 4, 1943– December 8, 2021)]. Fisheries, 47(4):187, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:CEa

[2996] Anonymous. Conferences and events. Fisheries, 47(4):189, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JHd

[2997] Anonymous. Journal highlights. Fisheries, 47(4):190–191, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:MM

[2998] Anonymous. Mola mola. Fisheries, 47(4):192, April 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:FVNe

[2999] Anonymous. Fisheries volume 47 number 5 May 2022. Fisheries, 47(5): 193–236, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIe

[3000] Anonymous. Issue information. Fisheries, 47(5):193–194, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Roulson:2022:CE

[3001] Leanne H. Roulson. Cumulative effects. Fisheries, 47(5):195–196, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Winters:2022:CFL

[3002] Drue Banta Winters. Crossing the finish line: Help pass the Recovering America’s Wildlife Act this Congress. Fisheries, 47(5):197, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Guo:2022:KTF

[3003] Lian Guo. A key to thriving in fisheries science is a sense of belonging. Fisheries, 47(5):198–199, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Bade:2022:ACH

[3004] Andrew P. Bade, David A. Dippold, Brian A. Schmidt, Mark R. DuFour, Travis J. Hartman, and Stuart A. Ludsin. Angler choices that help catch lots of big fish. Fisheries, 47(5):200–212, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Rivers:2022:CRN

[3005] James W. Rivers, Jennifer Bailey Guerrero, Richard D. Brodeur, Gregory K. Krutzikowsky, Lindsay J. Adrean, Scott A. Heppell, Kym C. Jacobson, Kristen Milligan, S. Kim Nelson, Daniel D. Roby, William J. Sydeman, Leigh G. Torres, and John A. Barth. Critical research needs for forage fish within inner shelf marine ecosystems. Fisheries, 47(5): 213–221, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:CEb

[3006] Anonymous. Continuing education 2022. Fisheries, 47(5):222–224, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Yoder:2022:MJR

[3007] Chris Yoder, Bob Hughes, Mark Pyron, and C. Lee Bridges. In memoriam: [James R. Gammon: April 24, 1930–February 18, 2022]. Fisheries, 47(5):225–226, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Meyer:2022:AIC

[3008] Kevin A. Meyer. AFS Idaho Chapter 2022 Annual Meeting: Removing barriers: Opening pathways to the fisheries profession. Fisheries, 47(5): 229–230, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Quinn:2022:TSG

[3009] Thomas P. Quinn. Trout and salmon of the genus Salmo. Fisheries, 47 (5):231, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:CEc

[3010] Anonymous. Conferences and events. Fisheries, 47(5):233, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JHe

[3011] Anonymous. Journal highlights. Fisheries, 47(5):234–235, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Grimes:2022:BW

[3012] Greg Grimes. The bluegill whisperer. Fisheries, 47(5):236, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:FVNf

[3013] Anonymous. Fisheries Volume 47 Number 6 June 2022. Fisheries, 47(6): 237–284, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIf

[3014] Anonymous. Issue information. Fisheries, 47(6):237–238, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Roulson:2022:SB

[3015] Leanne H. Roulson. Safety break. Fisheries, 47(6):239–240, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Dale:2022:DVF

[3016] Katherine E. Dale, Laura C. Goetz, Katie M. Kobayashi, Molly R. Lane, and Emily C. Nazario. Diverse voices in fisheries science: Lessons learned from an insightful seminar series. Fisheries, 47(6):241–244, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Rahel:2022:MFF

[3017] Frank J. Rahel. Managing freshwater fish in a changing climate: Resist, accept, or direct. Fisheries, 47(6):245–255, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Shirey:2022:RRP

[3018] Patrick D. Shirey, Susan A. R. Colvin, Leanne H. Roulson, and Thomas E. Bigford. A review of recovery plan criteria for threatened and endangered fish taxa managed by the U.S. Fish and Wildlife Service. Fisheries, 47(6):256–261, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Turcik:2022:LAC

[3019] Peter Turcik. Latin American and Caribbean Fisheries Congress to convene fisheries professionals from underrepresented regions. Fisheries, 47 (6):262–263, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:SA

[3020] Anonymous. Spokane August 21–25, 2022. Fisheries, 47(6):265–279, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:CEd

[3021] Anonymous. Conferences and events. Fisheries, 47(6):280, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JHf

[3022] Anonymous. Journal highlights. Fisheries, 47(6):282–283, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:OD

[3023] Anonymous. Okaloosa darter. Fisheries, 47(6):284, June 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:FVNg

[3024] Anonymous. Fisheries Volume 47 Number 7 July 2022. Fisheries, 47(7): 285–328, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIg

[3025] Anonymous. Issue information. Fisheries, 47(7):285–286, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Roulson:2022:OWP

[3026] Leanne H. Roulson. The ones who plant the trees may not always get to enjoy the shade, and that’s ok. Fisheries, 47(7):287–288, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Hargrove:2022:GBL

[3027] John S. Hargrove. Gender bias in letters of recommendation. Fisheries, 47(7):289, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Freeman:2022:TIU

[3028] Mary C. Freeman, Kevin R. Bestgen, Daren Carlisle, Emmanuel A. Frimpong, Nathan R. Franssen, Keith B. Gido, Elise Irwin, Yoichiro Kanno, Charles Luce, S. Kyle McKay, Meryl C. Mims, Julian D. Olden, N. LeRoy Poff, David L. Propst, Laura Rack, Allison H. Roy, Edward S. Stowe, Annika Walters, and Seth J. Wenger. Toward improved understanding of streamflow effects on freshwater fishes. Fisheries, 47(7):290–298, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Shirey:2022:ESA

[3029] Patrick D. Shirey and Susan A. R. Colvin. Endangered Species Act expenditures for fish taxa managed by the U.S. Fish and Wildlife Service are predicted by lawsuits, captive propagation, and region. Fisheries, 47 (7):299–303, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Rolls:2022:HSW

[3030] Robert J. Rolls, Jane S. Rogosch, and Lauren M. Kuehne. How shall we meet? Embracing the opportunities of virtual conferencing. Fisheries, 47 (7):304–306, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JASb

[3031] Anonymous. The Joint Aquatic Sciences Meeting (JASM). Fisheries, 47 (7):307–315, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:PS

[3032] Anonymous. Plenary speakers. Fisheries, 47(7):317–320, July 2022. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:AAMb

[3033] Anonymous. AFS Annual Meeting exhibitors. Fisheries, 47(7):322–323, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:CEe

[3034] Anonymous. Conferences and events. Fisheries, 47(7):325, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JHg

[3035] Anonymous. Journal highlights. Fisheries, 47(7):326–327, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:DS

[3036] Anonymous. Delta smelt. Fisheries, 47(7):328, July 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:FVNh

[3037] Anonymous. Fisheries Volume 47 Number 8 August 2022. Fisheries, 47(8):329–372, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIh

[3038] Anonymous. Issue information. Fisheries, 47(8):329–330, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Roulson:2022:TW

[3039] Leanne H. Roulson. That’s a wrap. Fisheries, 47(8):331–332, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Chen:2022:TBM

[3040] Emily K. Chen, Jacquelyn R. Galvez, and Stephanie M. Carlson. Towards building a more diverse and inclusive culture in the fisheries classroom. Fisheries, 47(8):333–336, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Tracy:2022:WFS

[3041] Bryn H. Tracy. What’s in a fish species name and when to change it? Fisheries, 47(8):337–345, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Polivka:2022:IYB

[3042] Carlos M. Polivka. “If you build it… ”: Methodological approaches to detect postrestoration responses in stream fishes. Fisheries, 47(8):346– 355, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Pegg:2022:CDR

[3043] Mark A. Pegg, David L. Buckmeier, Martin J. Hamel, and Jeff D. Koch. Creating a digital repository of calcified structures from known-age fishes, a century in the making. Fisheries, 47(8):357–360, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Turcik:2022:BBU

[3044] Peter Turcik. Brett Billings: U.S. Fish and Wildlife videographer. Fisheries, 47(8):362–366, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:CGS

[3045] Anonymous. 2022 classes of Golden and Silver Membership Award winners. Fisheries, 47(8):367, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:Cd

[3046] Anonymous. Calendar. Fisheries, 47(8):369, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JHh

[3047] Anonymous. Journal highlights. Fisheries, 47(8):370–371, August 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:RL

[3048] Anonymous. Roanoke logperch. Fisheries, 47(8):372, August 2022. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:FVNi

[3049] Anonymous. Fisheries volume 47 number 9 September 2022. Fisheries, 47(9):373–416, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIi

[3050] Anonymous. Issue information. Fisheries, 47(9):373–374, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Croxton:2022:NB

[3051] April Croxton. New beginnings. Fisheries, 47(9):375–376, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zipp:2022:WCA

[3052] Kaylyn Zipp. Widening cracks in an already leaky pipeline. Fisheries, 47(9):377–379, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Scarnecchia:2022:TTT

[3053] Dennis L. Scarnecchia and Jason D. Schooley. Trophies, technology, and timescape in fisheries management, as exemplified through Oklahoma’s world record paddlefish Polyodon spathula. Fisheries, 47(9):381– 394, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

OReilly:2022:BLL

[3054] Katherine E. O’Reilly. It’s beginning to look a lot like #25DaysofFishmas: Communicating freshwater biodiversity using social media. Fisheries, 47(9):395–405, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Davies:2022:CRB

[3055] Hailey L. Davies, Shane Gross, Dana R. Haggarty, and Francis Juanes. Conserving rockfishes: Barotrauma and descending devices in the Northeast Pacific. Fisheries, 47(9):406–411, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:SBM

[3056] Anonymous. Steven Berkeley Marine Conservation Fellowship: 2022 winners. Fisheries, 47(9):412, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:Ce

[3057] Anonymous. Calendar. Fisheries, 47(9):413, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JHi

[3058] Anonymous. Journal highlights. Fisheries, 47(9):414–415, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:FBR

[3059] Anonymous. Fishery biologist releases an American paddlefish. Fisheries, 47(9):416, September 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:FVNj

[3060] Anonymous. Fisheries volume 47 number 10 October 2022. Fisheries, 47(10):417–456, October 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIj

[3061] Anonymous. Issue information. Fisheries, 47(10):417–418, October 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Hayes:2022:ECB

[3062] Daniel Hayes. The emotional challenges of being a fisheries (or wildlife) scientist. Fisheries, 47(10):419–420, October 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Crandall:2022:AVC

[3063] Chelsey Crandall. Amplified voices: a collection of discussions from women of fisheries. Fisheries, 47(10):421–422, October 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Chestnut-Faull:2022:REE

[3064] Kristen Chestnut-Faull, Martha Mather, Quinton Phelps, and Dan Shoup. A review of empirical evidence that examines the effectiveness of harvest regulation evaluations in freshwater systems: a systematic, standardized collaborative approach. Fisheries, 47(10):423–434, October 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Kritzer:2022:MSF

[3065] Jacob P. Kritzer, Carolyn J. Hall, Bruce Hoppe, Curtis Ogden, and Jamie M. Cournane. Managing small fish at large scales: The emergence of regional policies for river herring in the Eastern United States. Fisheries, 47(10):435–445, October 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Collie:2022:BEH

[3066] Jeremy Collie, Sarah Schumann, Kate Masury, Hirotsugu Uchida, and Claire Collie. Balancing ecosystems, harvests, and seafood markets. Fisheries, 47(10):446–450, October 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:CEf

[3067] Anonymous. Conferences and events. Fisheries, 47(10):453, October 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JHj

[3068] Anonymous. Journal highlights. Fisheries, 47(10):454–455, October 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:ADC

[3069] Anonymous. Alewife in Deer Creek. Fisheries, 47(10):456, October 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:FVNk

[3070] Anonymous. Fisheries volume 47 number 11 November 2022. Fisheries, 47(11):457–520, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIk

[3071] Anonymous. Issue information. Fisheries, 47(11):457–458, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Oremland:2022:HFM

[3072] Laura Oremland, Abigail Furnish, Julia Byrd, and Richard Cody. How fishery managers can harness the power of the crowd: Using citizen science and nontraditional data sources in fisheries management. Fisheries, 47(11):459–462, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Hall-Scharf:2022:VKT

[3073] Brittany J. Hall-Scharf. Volunteers keeping tabs on superheroes of the sea: Florida’s adventure coast. Fisheries, 47(11):464–467, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Gervasi:2022:RAC

[3074] Carissa L. Gervasi, Jordan A. Massie, Jonathan Rodemann, Shakira Trabelsi, Rolando O. Santos, and Jennifer S. Rehage. Recreational angler contributions to fisheries management are varied and valuable: Case studies from South Florida. Fisheries, 47(11):469–477, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Musick:2022:VGF

[3075] Susanna Musick, Lewis Gillingham, Matt Perkinson, and Thom Teears. Virginia Game Fish Tagging Program: a nontraditional data source for fisheries management. Fisheries, 47(11):478–481, November 2022. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Martin:2022:CSB

[3076] Karen L. M. Martin and Melissa Studer. Citizen science on the beach: Grunion greeters in California. Fisheries, 47(11):483–490, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Byrd:2022:DFP

[3077] Julia Byrd, William C. Collier, and Alyson Iberle. Designing the FISHstory Project to support fisheries management. Fisheries, 47(11):492– 498, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Lynn:2022:CAS

[3078] Kirk Lynn, Emmanis Dorval, Dianna Porzio, and Trung Nguyen. A collaborative aerial survey of coastal pelagic species in nearshore California waters. Fisheries, 47(11):500–508, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Ra:2022:KTH

[3079] Hana Ra, Benjamin L. Richards, Audrey Rollo, Dianna Miller-Greene, and Jeremy Taylor. Keeping track of Hawaii’s bottomfish populations with the help of citizen scientists. Fisheries, 47(11):510–515, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:Cf

[3080] Anonymous. Calendar. Fisheries, 47(11):516, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JHk

[3081] Anonymous. Journal highlights. Fisheries, 47(11):518–519, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Oremland:2022:VTB

[3082] Laura Oremland. Volunteer tagging black sea bass. Fisheries, 47(11): 520, November 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:FVNl

[3083] Anonymous. Fisheries volume 47 number 12 December 2022. Fisheries, 47(12):521–576, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIl

[3084] Anonymous. Issue information. Fisheries, 47(12):521–522, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Croxton:2022:TYA

[3085] April Croxton. It’s that time of the year again: “Membership renewal season”. Fisheries, 47(12):523, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Moncrief-Cox:2022:AVC

[3086] Heather Moncrief-Cox. Amplified voices: a collection of discussions from women of fisheries. Fisheries, 47(12):524–527, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Jones:2022:ERF

[3087] Thomas S. Jones, Matthew I. Jones, Melissa Treml, and Thomas Heinrich. Estimating recycling of fish in catch-and-release fisheries. Fisheries, 47(12):529–536, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Hughes:2022:MJM

[3088] Bob Hughes, Glenn Griffith, and Chris Yoder. In memoriam: James M. Omernik, March 22, 1937–June 22, 2022. Fisheries, 47(12):537–538, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:AMW

[3089] Anonymous. Annual meeting wrapup. Fisheries, 47(12):540–547, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:AMS

[3090] Anonymous. Annual meeting: Symposia summaries 2022. Fisheries, 47 (12):549–560, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:AFSb

[3091] Anonymous. 2022 American Fisheries Society Awards. Fisheries, 47 (12):561–570, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:Cg

[3092] Anonymous. Calendar. Fisheries, 47(12):571, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:JHl

[3093] Anonymous. Journal highlights. Fisheries, 47(12):572–575, December 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:TS

[3094] Anonymous. A tasty snack. Fisheries, 47(12):576, December 2022. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNa

[3095] Anonymous. Fisheries volume 48 number 1 January 2023. Fisheries, 48 (1):1–44, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIa

[3096] Anonymous. Issue information. Fisheries, 48(1):1–2, January 2023. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zipp:2023:AVCa

[3097] Kaylyn Zipp. Amplified voices: a collection of discussions from women of fisheries. Fisheries, 48(1):3–5, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Berniche:2023:HSG

[3098] Ashley Berniche. Hutton Scholars gather for second annual summit. Fisheries, 48(1):6–7, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Campbell:2023:CSS

[3099] Jillian Campbell, Jennifer Yakimishyn, Dana Haggarty, Francis Juanes, and Sarah Dudas. Citizen science surveys provide novel nearshore data. Fisheries, 48(1):8–19, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

King:2023:SII

[3100] Brittany D. King. Social identities, intersectionality, and the experiences of women and women of color in marine, aquatic, and fisheries science professions. Fisheries, 48(1):20–28, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Miranda:2023:AOP

[3101] Leandro E. Miranda. Abundance–occupancy patterns of black bass in an impounded river. Fisheries, 48(1):29–37, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Dillard:2023:MLC

[3102] Joe Dillard. In memoriam: Lee Charles Redmond: September 15, 1939– September 20, 2022. Fisheries, 48(1):38, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:WGR

[3103] Anonymous. Welcome to Grand Rapids. Fisheries, 48(1):39, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Ca

[3104] Anonymous. Calendar. Fisheries, 48(1):40–41, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHa

[3105] Anonymous. Journal highlights. Fisheries, 48(1):42–43, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:LEP

[3106] Anonymous. Lake Eustis pupfish. Fisheries, 48(1):44, January 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNb

[3107] Anonymous. Fisheries volume 48 number 2 February 2023. Fisheries, 48(2):45–88, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIb

[3108] Anonymous. Issue information. Fisheries, 48(2):45–46, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Rice:2023:PFS

[3109] Emma Rice and Edith Gondwe. Perspectives from fisheries social scientists: Mixed methods as a DEIJ tool. Fisheries, 48(2):47–48, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Neely:2023:EEL

[3110] Ben C. Neely, Jeff D. Koch, and Keith B. Gido. Evaluating effects of live-imaging sonar on angler catch of crappies in a Kansas impoundment. Fisheries, 48(2):49–53, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

White:2023:DGS

[3111] Shannon L. White, Dewayne A. Fox, Tamar Beridze, Stephania K. Bolden, Robin L. Johnson, Thomas F. Savoy, Fleur Scheele, Andrea D. Schreier, and David C. Kazyak. Decades of global sturgeon conservation efforts are threatened by an expanding captive culture industry. Fisheries, 48(2):54–61, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Hartman:2023:OIE

[3112] Jordan H. Hartman and Eric R. Larson. Overlooked invaders? Ecological impacts of non-game, native transplant fishes in the United States. Fisheries, 48(2):62–71, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Franzin:2023:MCC

[3113] William G. Franzin, Richard A. Bodaly, Redmond McV. Clarke, G. Eric E. Moodie, Patricia F. Moodie, Erich V. Kliewer, John S. Loch, Chris J. Foote, Allan H. Kristofferson, and G. Burton Ayles. In memoriam: Casimir C. Lindsey, March 22, 1923–April 27, 2022. Fisheries, 48 (2):72, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Murawski:2023:MWH

[3114] Steve Murawski, John Oliver, Mike Sissenwine, Frank Almeida, and Tom Frazer. In memoriam: William T. Hogarth (March 7, 1939–November 5, 2022). Fisheries, 48(2):73–74, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

DeBruyne:2023:IPF

[3115] Robin DeBruyne, Edward Roseman, and Martin Čech. 5th International Percid Fish Symposium: Continuing a great Trans-Atlantic tradition. Fisheries, 48(2):75–78, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:LACa

[3116] Anonymous. Latin American and Caribbean Fisheries Congress: Registration and travel. Fisheries, 48(2):79–80, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:AD

[3117] Anonymous. AFS 2022 donors. Fisheries, 48(2):82–83, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Cb

[3118] Anonymous. Calendar. Fisheries, 48(2):85, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHb

[3119] Anonymous. Journal highlights. Fisheries, 48(2):86–87, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:WM

[3120] Anonymous. Western mosquitofish. Fisheries, 48(2):88, February 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNc

[3121] Anonymous. Fisheries volume 48 number 3 March 2023. Fisheries, 48(3): 89–132, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIc

[3122] Anonymous. Issue information. Fisheries, 48(3):89–90, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Johnson:2023:TRA

[3123] Jennifer Johnson and Patrick Hanchin. Top 10 reasons to attend the annual meeting in Grand Rapids. Fisheries, 48(3):91, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zipp:2023:AVCb

[3124] Kaylyn Zipp. Amplified voices: A collection of discussions from women of fisheries. Fisheries, 48(3):92–94, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Roulson:2023:AA

[3125] Leanne H. Roulson. AFS has an award for that. Fisheries, 48(3):95–96, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Losee:2023:VAT

[3126] James P. Losee, Derek R. Dapp, Gabe Madel, and Greg Shimek. Volunteer angling and technology-based solutions provide the first estimate of sea lice infections for wild coastal cutthroat trout Oncorhynchus clarkii clarkii. Fisheries, 48(3):97–106, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Ofir:2023:ASO

[3127] Eyal Ofir, Tal Silver, Jeroen G. Steenbeek, Noam Shachar, and Gideon Gal. Applying the Safe Operating Space (SOS) approach to sustainable commercial fishing under varying lake levels and littoral zone conditions. Fisheries, 48(3):107–120, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:LACb

[3128] Anonymous. Latin American and Caribbean Fisheries Congress. Fisheries, 48(3):121–122, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Kopaska:2023:BTN

[3129] Jeff Kopaska, Dan Brauch, Joseph D. Conroy, Lawrence Dorsey, Elsa Goerig, Mark Fincel, Paul Kusnierz, Brian Misseldine, Leanne Roulson, and Kaylyn Zipp. Building the team: Nominating the 2023 AFS Second Vice President candidates: 2022–2023 American Fisheries Society Nominating Committee. Fisheries, 48(3):123–124, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Murphy:2023:ASV

[3130] Margaret H. Murphy. AFS Second Vice President candidate statement. Fisheries, 48(3):125–126, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Shirey:2023:ASV

[3131] Patrick D. Shirey. AFS Second Vice President candidate statement. Fisheries, 48(3):127–128, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Cc

[3132] Anonymous. Calendar. Fisheries, 48(3):129, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHc

[3133] Anonymous. Journal highlights. Fisheries, 48(3):130–131, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:SA

[3134] Anonymous. Silver arowana. Fisheries, 48(3):132, March 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNd

[3135] Anonymous. Fisheries volume 48 number 4 April 2023. Fisheries, 48(4): 133–180, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IId

[3136] Anonymous. Issue information. Fisheries, 48(4):133–134, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zipp:2023:AVCc

[3137] Kaylyn Zipp. Amplified voices: a collection of discussions from women of fisheries. Fisheries, 48(4):135–136, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Sloat:2023:WTS

[3138] Matthew R. Sloat. What the Taimen said: an urgent call for conservation of the World’s largest salmonids. Fisheries, 48(4):137–140, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Krabbenhoft:2023:SPO

[3139] Corey A. Krabbenhoft, Stuart A. Ludsin, Elizabeth A. Marschall, Richard R. Budnik, L. Zoe Almeida, Christopher L. Cahill, Holly S. Embke, Zachary S. Feiner, Patrick J. Schmalz, Matt J. Thorstensen, Michael J. Weber, Melissa R. Wuellner, and Gretchen J. A. Hansen. Synthesizing professional opinion and published science to build a conceptual model of walleye recruitment. Fisheries, 48(4):141–156, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zorn:2023:CSS

[3140] Troy Zorn, Todd Wills, Jan-Michael Hessenauer, Joel Lenz, Ed Bissell, Ashley DePottey, Danielle Forsyth Kilijanczyk, and Anila Francis. Combining statewide surveys and classification to support management of streams. Fisheries, 48(4):157–167, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:LAC

[3141] Anonymous. Latin America and Caribbean Fisheries Congress. Fisheries, 48(4):169–171, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:AAM

[3142] Anonymous. AFS annual meeting. Fisheries, 48(4):172–175, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Cd

[3143] Anonymous. Calendar. Fisheries, 48(4):176, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHd

[3144] Anonymous. Journal highlights. Fisheries, 48(4):177–179, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:WE

[3145] Anonymous. Walleye eggs. Fisheries, 48(4):180, April 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNe

[3146] Anonymous. Fisheries volume 48 number 5 May 2023. Fisheries, 48(5): 181–224, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIe

[3147] Anonymous. Issue information. Fisheries, 48(5):181–182, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Cohn:2023:RLS

[3148] Rachel M. Cohn. Recreating labels, sustaining fishing communities. Fisheries, 48(5):183–184, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Woods:2023:SFS

[3149] Mackenzie B. Woods. Singing fish in a sea of noise:: Assessing the impact of anthropogenic noise on plainfin midshipman fish. Fisheries, 48(5): 185–189, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

DiStefano:2023:MPI

[3150] Robert J. DiStefano, Mael G. Glon, and Bailey M. O’Brian. Managing pathways for introduction of alien crayfish: Evaluation of a midwestern state’s regulations on the pet industry. Fisheries, 48(5):190–196, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Lyach:2023:FAC

[3151] Roman Lyach. Fisheries and aquaculture in the Czech Republic: History, current status, management, and regulations. Fisheries, 48(5):197–203, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Gido:2023:MPM

[3152] Keith B. Gido, Megan J. Osborne, David L. Propst, Thomas F. Turner, and Julian D. Olden. Megadroughts pose mega-risk to native fishes of the American Southwest. Fisheries, 48(5):204–214, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Winters:2023:SAF

[3153] Drue Banta Winters. Statement of the American Fisheries Society and the Western Division of AFS about the need to breach the four dams on the Lower Snake River. Fisheries, 48(5):215–217, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:AJR

[3154] Anonymous. AFS journals retreat. Fisheries, 48(5):218, May 2023. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:CEC

[3155] Anonymous. Continuing education course descriptions. Fisheries, 48(5): 219–220, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Ce

[3156] Anonymous. Calendar. Fisheries, 48(5):221, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHe

[3157] Anonymous. Journal highlights. Fisheries, 48(5):222–223, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:GT

[3158] Anonymous. Gila trout. Fisheries, 48(5):224, May 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNf

[3159] Anonymous. Fisheries volume 48 number 6 June 2023. Fisheries, 48(6): 225–272, June 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIf

[3160] Anonymous. Issue information. Fisheries, 48(6):225–226, June 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Croxton:2023:RPS

[3161] April Croxton. The role of professional societies in a post-COVID environment. Fisheries, 48(6):227, June 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Shellock:2023:WLE

[3162] Rebecca Jane Shellock and Katherine Maltby. Women leaders are essential for tackling ocean sustainability challenges. Fisheries, 48(6):228–232, June 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Seegert:2023:ORE

[3163] Greg Seegert, Joe T. Vondruska, Ron King, and Doug Dixon (Retired). The Ohio River Ecological Research Program (ORERP): History and accomplishments. Fisheries, 48(6):233–246, June 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Pipernos:2023:TRF

[3164] Sara Pipernos, Grantly Galland, and Shana Miller. Tuna regional fisheries management organizations need to prioritize transition to management procedures. Fisheries, 48(6):247–254, June 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:AM

[3165] Anonymous. 153rd annual meeting. Fisheries, 48(6):255–268, June 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Cf

[3166] Anonymous. Calendar. Fisheries, 48(6):269, June 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHf

[3167] Anonymous. Journal highlights. Fisheries, 48(6):270–271, June 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FDT

[3168] Anonymous. Freshwater drum teeth. Fisheries, 48(6):272, June 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNg

[3169] Anonymous. Fisheries volume 48 number 7 July 2023. Fisheries, 48(7): 273–316, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIg

[3170] Anonymous. Issue information. Fisheries, 48(7):273–274, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Croxton:2023:IMN

[3171] April Croxton. The importance of mentoring and networking for midcareer professionals. Fisheries, 48(7):275, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zipp:2023:AVCd

[3172] Kaylyn Zipp. Amplified voices: a collection of discussions from women of fisheries. Fisheries, 48(7):276–277, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Lovgren:2023:WMP

[3173] Stefan Lovgren. The wonders of the Mekong Project shows both the wonders and conservation challenges of the Mekong River. Fisheries, 48 (7):279–286, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Schall:2023:ARP

[3174] Megan K. Schall, Geoffrey D. Smith, Vicki S. Blazer, Heather L. Walsh, Timothy Wertz, Dustin Shull, and Tyler Wagner. Assembling the right pieces: Developing an interdisciplinary team to study disease, decline, and recovery of a world-class smallmouth bass fishery. Fisheries, 48(7): 287–294, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Leppi:2023:CCR

[3175] Jason C. Leppi, Daniel J. Rinella, Mark S. Wipfli, Anna K. Liljedahl, Andrew C. Seitz, and Jeffrey A. Falke. Climate change risks to freshwater subsistence fisheries in Arctic Alaska: Insights and uncertainty from broad whitefish Coregonus nasus. Fisheries, 48(7):295–306, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:PSS

[3176] Anonymous. Plenary speakers, sponsors, child care. Fisheries, 48(7): 308–310, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Cg

[3177] Anonymous. Calendar. Fisheries, 48(7):312, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHg

[3178] Anonymous. Journal highlights. Fisheries, 48(7):313–315, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:MAC

[3179] Anonymous. Male Arctic char snout. Fisheries, 48(7):316, July 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNh

[3180] Anonymous. Fisheries volume 48 number 8 August 2023. Fisheries, 48(8):317–356, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIh

[3181] Anonymous. Issue information. Fisheries, 48(8):317–318, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Croxton:2023:PLB

[3182] April Croxton. Passing the leadership baton. Fisheries, 48(8):319, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Cooke:2023:FT

[3183] Steven J. Cooke. Fisheries in transition. Fisheries, 48(8):320, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Platzer:2023:USD

[3184] Elsie Platzer. UC strike demonstrates the importance of DEIJA in fisheries. Fisheries, 48(8):321, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Flores:2023:MSP

[3185] Bobbie Flores. In memoriam: Steven Putman Cramer. Fisheries, 48(8): 322, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Aarrestad:2023:MRB

[3186] Peter Aarrestad, Stephen Gephard, William Hyatt, Richard Jacobson, and Eric Smith. In memoriam: Robert A. “Bob” Jones. Fisheries, 48 (8):323, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Kusnierz:2023:ACT

[3187] Paul C. Kusnierz, Jonathan Leiman, and Henriette I. Jager. Addressing coldwater temperature impairment in a changing climate. Fisheries, 48 (8):324–330, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Feddern:2023:KNB

[3188] Megan L. Feddern, Erik R. Schoen, Rebecca Shaftel, Curry J. Cunningham, Craig Chythlook, Brendan M. Connors, Alyssa D. Murdoch, Vanessa R. von Biela, and Brooke Woods. Kings of the North: Bridging disciplines to understand the effects of changing climate on Chinook salmon in the Arctic–Yukon–Kuskokwim Region. Fisheries, 48(8):331– 343, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Turcik:2023:PSS

[3189] Peter Turcik. Photographer Sam Stukel. Fisheries, 48(8):344–347, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:LH

[3190] Anonymous. LACFC highlights. Fisheries, 48(8):348–351, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Ch

[3191] Anonymous. Calendar. Fisheries, 48(8):353, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHh

[3192] Anonymous. Journal highlights. Fisheries, 48(8):354–355, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Ci

[3193] Anonymous. Coregonines. Fisheries, 48(8):356, August 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNi

[3194] Anonymous. Fisheries volume 48 number 9 September 2023. Fisheries, 48(9):357–400, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIi

[3195] Anonymous. Issue information. Fisheries, 48(9):357–358, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Jennings:2023:MJA

[3196] Cecil A. Jennings. My journey to AFS leadership and goals during my tenure. Fisheries, 48(9):359–360, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zipp:2023:AVCe

[3197] Kaylyn Zipp. Amplified voices: a collection of discussions from women of fisheries. Fisheries, 48(9):361–364, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Roulson:2023:NSS

[3198] Leanne Roulson, Niall Clancy, Adam Price, Cienna Hanson, Zachary Hooley-Underwood, Skylar Rousseau, and Chris Walser. Nonpermanent streams: Some insights from the AFS Western Division. Fisheries, 48 (9):365–366, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Taylor:2023:MGF

[3199] Gerry Taylor and Ross Peterson. In memoriam: Gordon F. Hartman, September 20, 1927–April 21, 2021. Fisheries, 48(9):367, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Grant:2023:IEB

[3200] Joshua D. Grant, Steven J. Ingram, and Scott A. Bonar. Influence of electrofishing boat operation and driving techniques on reservoir fish catches. Fisheries, 48(9):368–376, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Shoup:2023:LTS

[3201] Daniel E. Shoup and Dray D. Carl. Long-term standardized sampling as a tool for evaluating management actions: BACI analysis to reevaluate predatory control of stunted crappie with saugeye. Fisheries, 48(9):377– 386, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Clemens:2023:NRR

[3202] Benjamin J. Clemens, Jordan K. Matley, Natalie V. Klinard, Robert J. Lennox, Lene K. Sortland, and Steven J. Cooke. The need for reporting rationale and detailed methods in studies that surgically implant fish with electronic tracking devices. Fisheries, 48(9):388–394, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FCP

[3203] Anonymous. First call for papers. Fisheries, 48(9):395, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Cj

[3204] Anonymous. Calendar. Fisheries, 48(9):397, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHi

[3205] Anonymous. Journal highlights. Fisheries, 48(9):398–399, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:BCF

[3206] Anonymous. Black crappie fingerlings. Fisheries, 48(9):400, September 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNj

[3207] Anonymous. Fisheries volume 48 number 10 October 2023. Fisheries, 48(10):401–444, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIj

[3208] Anonymous. Issue information. Fisheries, 48(10):401–402, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Jennings:2023:YTY

[3209] Cecil A. Jennings. Have you thanked your mentor(s) lately … ever? Fisheries, 48(10):403–404, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Andrews:2023:AWD

[3210] Lauren Andrews. AFS Western Division and Idaho Chapter 2023 joint annual meeting. Fisheries, 48(10):405–407, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:SBM

[3211] Anonymous. Steven Berkeley Marine Conservation Fellowship: 2023 winners. Fisheries, 48(10):408–409, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Wheeler:2023:NIR

[3212] A. Powell Wheeler, Kyle T. Rachels, and Kevin J. Dockendorf. NCIFD — an internal R package for a fisheries agency. Fisheries, 48(10):411– 417, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Scarnecchia:2023:ETS

[3213] Dennis L. Scarnecchia, Shaylee Martling, and Michael Barnes. The evolution of trout stream management in the Black Hills, 1883–2023, as evidenced through hatchery activities and stocking. Fisheries, 48(10): 419–439, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:CHA

[3214] Anonymous. Co-hosted by the AFS Western Division and the Pacific Islands Chapter. Fisheries, 48(10):440, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Ck

[3215] Anonymous. Calendar. Fisheries, 48(10):441, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:NJD

[3216] Anonymous. New journal design. Fisheries, 48(10):442–443, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:RTP

[3217] Anonymous. The rainbow trout pond. Fisheries, 48(10):444, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNk

[3218] Anonymous. Fisheries volume 48 number 11 November 2023. Fisheries, 48(11):445–488, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIk

[3219] Anonymous. Issue information. Fisheries, 48(11):445–446, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Jennings:2023:EFM

[3220] Cecil A. Jennings. Enhancing fisheries management through other ways of knowing. Fisheries, 48(11):447–448, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Servheen:2023:TCC

[3221] Gregg Servheen. The time has come: a commentary on the North American model of wildlife management. Fisheries, 48(11):449, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Gundermann:2023:DFY

[3222] Daria Gundermann. Don’t be fickle with your fish and climate change literature searches: Use FiCli! Fisheries, 48(11):450–452, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Bourret:2023:IHR

[3223] Samuel L. Bourret, Jill M. Janak, Tim J. Linley, Megan K. Nims, and Geoffrey A. McMichael. Inadvertent hatchery rainbow trout Oncorhynchus mykiss discovered in native westslope cutthroat trout Oncorhynchus clarkii lewisi habitat: a cautionary tale. Fisheries, 48(11): 453–464, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Hughes:2023:WUS

[3224] Robert M. Hughes, Royal C. Gardner, Patrick D. Shirey, S. Mažeika P. Sulliván, Susan A. R. Colvin, and Drue B. Winters. Waters of the United States: an urgent call for action by fisheries and aquatic science professionals. Fisheries, 48(11):465–468, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Wosnick:2023:UTB

[3225] Natascha Wosnick, Patricia Charvet, Rachel Ann Hauser-Davis, Getulio Rincon, Ana Rita Onodera Palmeira Nunes, and Jorge Luiz Silva Nunes. Unveiling the threats beneath: Fish mislabeling in the Brazilian Amazon Coast and its impacts on the critically endangered daggernose shark. Fisheries, 48(11):469–473, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Cupp:2023:ATR

[3226] Aaron Cupp, Andrea Fritts, Marybeth Brey, Christa Woodley, David Smith, Mark Cornish, Amy McGovern, Rob Simmonds, and Neal Jackson. Application of the technology readiness levels framework to natural resource management tools. Fisheries, 48(11):474–479, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Ransom:2023:MBH

[3227] Sonia Ransom. In memoriam: Bruce Hickok Ransom, Jr., January 29, 1949–June 11, 2023. Fisheries, 48(11):481, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:CAW

[3228] Anonymous. Cohosted by the AFS Western Division and the Pacific Islands Chapter. Fisheries, 48(11):482, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Chakrabarty:2023:IBD

[3229] Prosanta Chakrabarty. Ichthyopedia: a Biographical Dictionary of Ichthyologists. Fisheries, 48(11):483, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Cl

[3230] Anonymous. Calendar. Fisheries, 48(11):485, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHj

[3231] Anonymous. Journal highlights. Fisheries, 48(11):486–487, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:ESC

[3232] Anonymous. Electrofishing for silver carp. Fisheries, 48(11):488, November 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:FVNl

[3233] Anonymous. Fisheries volume 48 number 12 December 2023. Fisheries, 48(12):489–544, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:IIl

[3234] Anonymous. Issue information. Fisheries, 48(12):489–490, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zipp:2023:SLE

[3235] Kaylyn Zipp and Grace Colvin. Shining a light on equality: Grand Rapids Annual Meeting recap. Fisheries, 48(12):491–494, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Turcik:2023:NYM

[3236] Peter Turcik. New for 2024: You must opt-in to receive Fisheries magazine in the mail! Fisheries, 48(12):495–496, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Page:2023:EEC

[3237] Lawrence M. Page, Katherine E. Bemis, Thomas E. Dowling, Héctor S. Espinosa-Pérez, Lloyd T. Findley, Carter R. Gilbert, Karsten E. Hartel, Robert N. Lea, Nicholas E. Mandrak, Margaret A. Neighbors, Juan J. Schmitter-Soto, and Harold J. Walker, Jr. The eighth edition of Common and Scientific Names of Fishes. Fisheries, 48(12):497–498, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Miranda:2023:ACI

[3238] Leandro E. Miranda. Aging, climate, and invasions threaten reservoirs in the Mississippi River Basin. Fisheries, 48(12):499–514, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:AFSa

[3239] Anonymous. American Fisheries Society 153rd annual meeting Grand Rapids, Michigan. Fisheries, 48(12):515–518, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Mulligan:2023:HIH

[3240] Hannah Mulligan. Hidden invaders: How the live bait trade can move invasive species. Fisheries, 48(12):519, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Turcik:2023:SS

[3241] Peter Turcik. Symposia summaries 2023. Fisheries, 48(12):520–527, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:AFSb

[3242] Anonymous. 2023 American Fisheries Society awards. Fisheries, 48(12): 529–536, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:Cm

[3243] Anonymous. Calendar. Fisheries, 48(12):538, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:JHk

[3244] Anonymous. Journal highlights. Fisheries, 48(12):540–543, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2023:LT

[3245] Anonymous. Lake trout. Fisheries, 48(12):544, December 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:FVNa

[3246] Anonymous. Fisheries volume 49 number 1 January 2024. Fisheries, 49 (1):1–48, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:IIa

[3247] Anonymous. Issue information. Fisheries, 49(1):1–2, January 2024. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Hayes:2024:ASC

[3248] Daniel Hayes, Mary Tate Bremigan, and Amber Peters. Active shooter on campus. Fisheries, 49(1):3–5, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zipp:2024:AVC

[3249] Kaylyn Zipp. Amplified voices: a collection of discussions from women of fisheries. Fisheries, 49(1):6–7, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Bisson:2024:MEI

[3250] Peter Bisson, Tracy Hillman, Tim Beechie, and George Pess. Managing expectations from intensively monitored watershed studies. Fisheries, 49 (1):8–15, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Bilby:2024:WAS

[3251] Robert E. Bilby, Ken P. Currens, Kurt L. Fresh, Derek B. Booth, Robert R. Fuerstenberg, and Gino L. Lucchetti. Why aren’t salmon responding to habitat restoration in the Pacific Northwest? Fisheries, 49(1):16–27, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Spriel:2024:FSE

[3252] Brittnie Spriel, Hailey L. Davies, Audrey Looby, Hailey Shafer, Sarah Vela, Francis Juanes, and Kieran D. Cox. Fish sounds as an effective tool in marine science communication. Fisheries, 49(1):28–34, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:WFC

[3253] Anonymous. World Fisheries Congress 2024. Fisheries, 49(1):36–39, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:AAM

[3254] Anonymous. AFS annual meeting. Fisheries, 49(1):41, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:Ca

[3255] Anonymous. Calendar. Fisheries, 49(1):43, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:JHa

[3256] Anonymous. Journal highlights. Fisheries, 49(1):44–46, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Turcik:2024:GKY

[3257] Peter Turcik. Get to know your AFS staff: Peter Turcik (he/him). Fisheries, 49(1):47, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:CSF

[3258] Anonymous. Chinook salmon flight. Fisheries, 49(1):48, January 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:FVNb

[3259] Anonymous. Fisheries volume 49 number 2 February 2024. Fisheries, 49(2):49–100, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:IIb

[3260] Anonymous. Issue information. Fisheries, 49(2):49–50, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Jennings:2024:RPF

[3261] Cecil A. Jennings. Relevance of professional fisheries certification in the 21st Century. Fisheries, 49(2):51–52, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zipp:2024:FWF

[3262] Kaylyn Zipp. Finding a way: Fishing has no boundaries casts a broader net, liberating all anglers. Fisheries, 49(2):53–54, February 2024. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Glassic:2024:YCT

[3263] Hayley C. Glassic, David D. Chagaris, Christopher S. Guy, Lusha M. Tronstad, Dominique R. Lujan, Michelle A. Briggs, Lindsey K. Albertson, Travis O. Brenden, Timothy E. Walsworth, and Todd M. Koel. Yellowstone cutthroat trout recovery in Yellowstone Lake: Complex interactions among invasive species suppression, disease, and climate change. Fisheries, 49(2):55–70, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Collis:2024:CTM

[3264] Ken Collis, Daniel D. Roby, Allen F. Evans, Timothy J. Lawes, and Donald E. Lyons. Caspian tern management to increase survival of juvenile salmonids in the Columbia River Basin: Progress and adaptive management considerations. Fisheries, 49(2):71–84, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Neal:2024:PBD

[3265] Reviewed by J. Wesley Neal. Peacock bass: Diversity, ecology and conservation. Fisheries, 49(2):85, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Berniche:2024:EHP

[3266] Ashley Berniche. Experiencing the Hutton Program effect through the perspective of past Hutton scholars. Fisheries, 49(2):87–89, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Orth:2024:BRM

[3267] Donald J. Orth. Book review: My Life in Fish: One Scientist’s Journey Gary D. Grossman. Illustrated by Ryan Tavarez. Published by Impspired Press. 2023. US$17.99 (Paperback). Fisheries, 49(2):90, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Rogers:2024:MBD

[3268] Kevin B. Rogers, Lori M. Martin, Greg Langer, Bill Atkinson, F. Boyd Wright, Harry Crockett, Chris Kennedy, and Doug Krieger. In memoriam: Bruce D. Rosenlund, June 6, 1946–August 31, 2023. Fisheries, 49(2):91–92, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:Cb

[3269] Anonymous. Calendar. Fisheries, 49(2):94, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:JHb

[3270] Anonymous. Journal highlights. Fisheries, 49(2):95–96, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Link:2024:TFA

[3271] Jason Link. TLAs, FLAs, and alternate steady states: Do we really need that many acronyms in our profession? Fisheries, 49(2):97–98, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Austen:2024:GKY

[3272] Doug Austen. Get to know your AFS staff: Doug Austen. Fisheries, 49 (2):99, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:YCT

[3273] Anonymous. Yellowstone cutthroat trout. Fisheries, 49(2):100, February 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:FVNc

[3274] Anonymous. Fisheries volume 49 number 3 March 2024. Fisheries, 49(3): 101–148, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:IIc

[3275] Anonymous. Issue information. Fisheries, 49(3):101–102, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Jennings:2024:WFC

[3276] Cecil A. Jennings. World Fisheries Congress: Seeing fisheries through an international lens. Fisheries, 49(3):103–104, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Harried:2024:BDE

[3277] Brittany Harried. Bans on diversity, equity, and inclusion programming at public universities nationwide are threatening future directions in fisheries. Fisheries, 49(3):105–107, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Jescovitch:2024:GLF

[3278] Lauren N. Jescovitch, Sharon Moen, Fatima Abdl-Haleem, and Titus S. Seilheimer. The Great Lakes future fishers initiative: Supporting the next generation of Great Lakes commercial fishing and seafood processing industries. Fisheries, 49(3):108–116, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Frater:2024:IHB

[3279] Paul N. Frater, Zachary S. Feiner, Gretchen J. A. Hansen, Daniel A. Isermann, Alexander W. Latzka, and Olaf P. Jensen. The incredible HALK: Borrowing data for age assignment. Fisheries, 49(3):117–128, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Januchowski-Hartley:2024:WFM

[3280] Stephanie Januchowski-Hartley, Adela Baratech, Demar Bon, and Beatrice Alducci. World Fish Migration Day: Celebrating a decade of building community around fish and rivers. Fisheries, 49(3):129–134, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Zipp:2024:FFL

[3281] Kaylyn Zipp, Jeff Kopaska, April Croxton, Scott Decker, Lawrence Dorsey, Laurie Earley, Eric Fetherman, Paul Kusnierz, Brian Missildine, Brian Nerbonne, and Jodi Whittier. Future-focused leadership: 2024 AFS Second Vice President nominations. Fisheries, 49(3):135–136, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Martin:2024:ASV

[3282] Lori M. Martin. AFS Second Vice President candidate statement. Fisheries, 49(3):137–138, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Douglas:2024:ASV

[3283] Marlis R. Douglas. AFS Second Vice President candidate statement. Fisheries, 49(3):139–140, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Castillo:2024:BRF

[3284] Pete Castillo, Brittnie Spriel, Bridget Maher, Nathanael Tabert, Talen Rimmer, and Francis Juanes. Book review: A Field Guide to Fishes of the Salish Sea: Puget Sound and the Straits of Georgia and Juan de Fuca, Washington State and British Columbia, T. W. Pietsch and J. W. Orr. Illustrated by J. R. Tomelleri. Published by Chatwin Books. 2023. 372 pages. US$36.00 (Paperback). Fisheries, 49(3):141–142, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:Cc

[3285] Anonymous. Calendar. Fisheries, 49(3):143, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2024:JHc

[3286] Anonymous. Journal highlights. Fisheries, 49(3):144–145, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Link:2024:EA

[3287] Jason Link. Elbows over arse. Fisheries, 49(3):146, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Lehman:2024:GKY

[3288] Debby Lehman. Get to know your AFS staff: Debby (she/her). Fisheries, 49(3):147, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Waldman:2024:SS

[3289] John Waldman. The stripeless striper. Fisheries, 49(3):148, March 2024. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).